A Bibliography of Publications in Computers and Graphics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.80

Title word cross-reference

(n + 1) [SA87]. + [JHPhR11]. 1 [RPM97]. 2 [ASS22, ATAG^{+}21, AA07, BDRV01, BPKG07, BGLA18, BD97, CLH^{+}16, CMB17, CYWM23, Dur89, Dur91, DGR93, EK22, GWX^{+}18, HWSW19, JJL97, KD11, KP95, Kor90, KCS22, KCS24, Laf94, LHS87, LWD^{+}18, LBTM15, MHLB16, Mar09, MS08, MB97a, NG88, Oik98, OdlCA02, PVCM22, PCWD23, RJS01, RE22, RdCVL16, RCG^{+}05, Sak02, SVP82, SK06, WBRV16, WUH^{+}15, WR02, YF09, ZPP^{+}23]. 2(1/2) [Oik98]. 2.5 [ST97]. 21/2 [BG91, NP96]. 3 [AT08, AMHWW16, AK21, ACG15, ASS22, ATAG^{+}21, ASS^{+}19, Ano03f, Ano03-62, Ano05-41, Ano07-39, Ano13o, ATZM19, AA13, ABCO12, ALM19, ABM^{+}06, BM03, BK89a, BF02a, BND^{+}17, BCS^{+}99, BBCG11, BY88, BWdBP13, BTD^{+}22, BAPD23, BCC^{+}22, BSF13, BTS19, BFLP20, BDL^{+}22, BYQZ22, BPKG07, BN03, BGLA18, Bou09, BJS01, BS01b, BK89b, BHL^{+}15, BKL15, BAC14, BD17, CCCS08, CKS98, CTP^{+}21, CPC^{+}18, CGG^{+}20, Car96, CMSF11, CVB16, CB97, CW03, CC08, CSG^{+}17, CSX^{+}19, CYWM23, CCM^{+}18, CSK97, CBNJ^{+}15, CGH97a, CCW01, CB10, CH12, CHSD95, CMDS17, CS18, DN22, DGV^{+}24, DCJH13, DEST95, DCLB19, DMS08, DG01, DSR11, DKY97, DGR93, Elb22, EK22, EKP93, ET07, EBC^{+}15, ERB^{+}14, EME15, FAZ21, FH11a, FH11b, FTB12, FDGM18, FHM98, FS98, FJW11, FWX^{+}18, FGZ^{+}22, FSS^{+}02, FCSB90, Fer01]. 3 [FIC21, FIC23, Fou11, FM22, FFP^{+}21, GPTB02, GVVJ99, GD95, Gin02, GF09, GMd^{+}13, Gom85, GA12, GZZS06, GvK18, GLC20, GBD88, GN89, GLA23, GYL^{+}13, HZ15, HLY^{+}19, HKHP11, HG21, HWSW19, HKBA17, HR88, HHCM17, HHKF10, HR07, HCLC16, HHN^{+}23, HYP^{+}24, HS99, IR06, IO91, JH89, JK15, JXW^{+}22, JTT01, JHPhR11, JA84b, JRS21, JCFN18, dSJdML18, JRJP^{+}22, KS98, Kan85, KSM07, Kas87, Kau88, K^{+}00a, KD11, KRK^{+}06, KCK17, KP95, Kle86, Kor90, KYT^{+}17, KVB^{+}20, KGB^{+}21, KLL^{+}15, KEVD18, KD15, KCS22, KCS24, KYM12, Laf94, LDS^{+}21, LPD^{+}18, LBD17, LNL^{+}23b, LNL^{+}23a, LMC13, LK18, LDLD22, LKL^{+}20, LBLV16, LVM^{+}11, LYW^{+}10, LLLC11, LLLZ16, LYL^{+}17, LZZ^{+}19, LWW^{+}20, LCL^{+}21, LSZQ21, LG94, LZT^{+}24, Lin97, LJWcH07, LXB^{+}15, LZL^{+}15, LM16, LXCW18, LWD^{+}18, LXJL21, LSCJ23, LRY^{+}24, LSS21, LSGFRC^{+}13, LUB^{+}13, LYS^{+}16, LCXL20, LB19]. 3 [LR90, LBTM15, LAE^{+}19, MAFL16, MHLB16, Maj98, MDM^{+}21, MKDM22, MOS^{+}21, MLC^{+}22, Mar79, MG09, MAG^{+}12, Mar10, MWA^{+}13, MMM^{+}20, MHYN23, MBST22, MFP11, MCKS06, MCP^{+}22, MY16, MSRB17, MS09b, NG88, NC12, NTAI20, NP88, NAK13, OK99, OK03, OP13, OCR^{+}19, ÖT21, OC21, PPS20, PSBD19, PRRR13, PK91, PS12, PS13b, PA07, PRW^{+}22, Pic88a, PPD22, PECW22, PPG^{+}18, PTY^{+}16, PCD^{+}15, PCWD23, PPVT03, QYC^{+}22, RV01, RTB^{+}18, RBB^{+}11, RB06, RO13, RMG15, RBF20, RHBS95, SMU22, SSB04, ST20, SDT21, ST22, Sak02, SBWS11, SVNB99, SS22, SGC00, SM99, STP^{+}20, SJT20, SF98, SD15, SPT18, SA17, SLL^{+}17, SLX^{+}16, Sin87, SBS13, SET^{+}88, Sla92, SJ09, SJZ^{+}23, SHBSS17, SGES12, SK06, SS75b, SCFF16, SK13, SHS^{+}23, TBS^{+}23, TZT^{+}22, TT19, THL15, US20, VBTW13, VGP04, VZP22, VAGT08]. 3 [VDOK19, WBA16, WS12, WBB^{+}08, WTM12, WZZZ18, WGS^{+}18a, WGS^{+}18b, WLYH19, WWH^{+}21, WRLZ23, WPB^{+}23, WUH^{+}15, WH96, WBL^{+}97, Wol02b, WWS^{+}13, WYZ20, WYC^{+}23, WLG04, XWW^{+}21, XLL^{+}18, XWWK21, YZ17, YPLL19, YSZ22, YLT^{+}18, YNS94, YHNC22, YHX10, ZTAP21, ZIP^{+}19, ZLLL21, ZPP^{+}23, ZLLG18, ZLZ^{+}20, ZM07, ZMK18, ZLL^{+}20, ZDL22, ZZDZ10, ZSM^{+}19, ZPL^{+}15, ZHG^{+}21, dGGDV11, dlLC99, vBT20a, vBT20b, vBT21]. 3x + 1 [CDR01]. 4 [KD11, KCK17, KGGP19, RAK^{+}15, XCL^{+}19]. 567 [VLD15]. 6 [ZMK18]. ^{2} [LaV07, GSY94]. ^{3} [How88]. ^{∗} [PF16]. ^{∘} [ZZLZ21]. PFH [CH91].^{}˚ [BBMGM22]. ^{T M} [DW89]. A [ALC06, Ano04-56]. α [BLNZ22, LLGA12, SMM20, TY24, VM15]. alpha < 0 [DBG93]. C [CZ98, GB91]. C^{0} [BFRA11]. C^{1} [Wil03, Pie84]. C^{2} [KP18, SZ09, Wil03, YCZ04, Sar92a, Lan88]. d [DW89]. ẋ(t) = −f(y(t)), ẏ(t) = f(x(t)) [Pic87a]. E^{2} [KKŽ04, Ska94, Ska96]. E^{3} [Ska97]. ϵ [CA17, JS09]. exp(−αζ + Z/ζ − Z) − 1 = 0 [YHHS93]. F (z) = e^{zw} + c [CZC02]. G [SP16]. G^{1} [ASC17, ASC18, LJH18, Pet18]. G^{2} [LJH18]. G^{n} [ZQ12]. J [CZ98, CZC02]. J_{1}^{a} [CNS^{+}06]. kd [CCI12]. L [HP01, OdlCA02, QLCV96]. L0 [LGZ^{+}21]. L_{0} [CZL14, CFZL16]. L_{∞} [BW98]. M [SK13, Car99a, CZ98, CZC02, Sob89]. μ [PDS21]. N [PGVACN06, ZGC15, And98, Liv18, SA87, Tar22, Wüt98]. O(1) [Ska96]. O(lg N) [Ska94]. P [YS97, WMRA^{+}15]. R^{4} [BSL^{+}13, CP13]. SL2(R) [NdSV20b]. T [KP19, AMHWW16]. xy [Baw97]. Z [CZ98, DBG93, GBV92]. z ← z^{2} + c [BGV93]. z ← z^{α} + c [GB91, GBV92, DBG93]. z ⟵ zw + c (w = a + ib) [CZ98]. z → z^{−n} + c [Shi93c]. z → z^{2} + μ [Ent89b].

-a [GTTC03]. -ary [ZGC15]. -B-splines [BLNZ22]. -bases [PDS21]. -best [SK13]. -blending [ZQ12]. -Buffer [YS97]. -Constrained [BW98]. -continuous [Wil03, ZQ12]. -Curves [TY24]. -D [CB97, ZPL^{+}15, BD97, Car96, CSK97, CGH97a, DEST95, DKY97, GD95, JJL97, Lin97, NP96, RPM97, ST97, WH96, WBL^{+}97]. -Decomposition [LLGA12]. -dimensional [And98, MB97a, Oik98, OdlCA02, SA87, Wüt98]. -gons [KP19]. -hull [SMM20, VM15]. -machines [JS09]. -manifolds [ACG15, EK22, VBTW13]. -maps [CA17]. -multigrid [WMRA^{+}15]. -plane [GB91, GBV92, DBG93]. -point [KD11]. -remeshing [VLD15]. -scale [FTB12]. -set [Car99a]. -sided [PGVACN06, Tar22]. -space [CCW01]. -splines [SP16]. -surfaces [Lan88]. -systems [HP01, QLCV96]. -trees [CCI12].

/Future [Ano03-41]. /PEX [PH90]. /PEX-environment [PH90].

’04 [Ano04-59]. ’05 [Ano04-58]. ‘07 [Ano07o]. 08 [CIW09].

1000 [Par86]. 10th [Ano01-59, Ano01-65, Ano05-41]. 111 [KRA^{+}23a]. 113C [Jor23d]. 1200 [Par86]. 12TH [Ano07-29, Ano07-28, Ano07a, Ano07b]. 13th [Ano04a, BA23]. ’16 [TSC16]. 16-bit [MDJ^{+}95]. 17th [Ano07-33]. 180VF [BGV93]. 19 [AAK^{+}22b, ASR^{+}22]. 1986 [Hop86]. 1990 [CL92]. 1994 [Ano93d, Ano94f, Ano94g]. 1996 [Ano95w, Ano95a, Ano95b, Ano95c]. 19991/2000 [Ano01-48].

2 [Cho77]. 2.0 [vWdlHFK23]. 2000/2001 [Ano03-38]. 2001 [Ano01j, Ano01k, Ano01l, Ano01q, Ano01r, Ano01p, Ano01t, Ano01m, Ano01n, Ano01-27, Ano01-29, Ano01-33, Ano01-34, Ano01-35, Ano01-38, Ano01-37, Ano01-39, Ano01-42, Ano01c, Ano01-43, Ano01-57, Ano01-55, Ano01-56, Ano01-61, Ano01-63, FT02, LWP02, MLPB02, NA02, Shi02, ZS02].

2002 [Ano01-28, Ano01-32, Ano01-58, Ano02e, Ano02i, Ano02k, Ano02b, Ano02c, Ano02m, Ano02q, Ano02d, Ano02r, Ano02-35, Ano02-29, Ano02-34, Ano02-37, Ano02-42, Ano02-43, Ano02-46, Ano02-51, Ano02-52, Ano02-54]. 2002/2003 [Ano04-44]. 2003 [Ano03e, Ano03q, Ano03y, Ano03-28, Ano03a, Ano03b, Ano03d, Ano03-37, Ano03-32, Ano03-39, Ano03-40]. 2003/2004 [Ano05-33]. 2004 [Ano03x, Ano03z, Ano04f, Ano04-28, Ano04c, Ano04-35, Ano04-62]. 2004/2005 [Ano06-36]. 2005 [Ano05e, Ano05f, Ano05i, Ano05w, Ano05x, Ano05y]. 2006 [Ano06c, Ano06h, Ano06p, Ano06z, Ano06y, Ano06a, Ano06b, Ano06-44, BCMM07, Oli08]. 2007 [Ano06n, Ano06-29, Ano06-30, Ano06-31, Ano06-51, Ano07k, Ano07-27, Ano07c, Ano07-39, Ano07-30, Ano07-38, BP08, QGW08, vdPS08]. 2007E [Ano07x]. 2007E-Learning [Ano07x]. 2008 [Ano08a]. 2009 [AL10, Ano10a]. 2010 [Ano11a, Ano11b, Ano11c, CI11]. 2011 [Ano11k, Ano12s, Ano13c, SK12]. 2012 [Ano13a, Ano13b, SH12]. 2013 [Ano13k]. 2014 [Ano15k, Mou15, NN15]. 2015 [BGT16, JZZ16, MS16, PDK16, Sez16, ZT16]. 2016 [AS16, AF16b, FS17, GW17, GB18, PK17]. 2017 [ALP17, EK18, GPS18, JZ17, Joh18, LPD^{+}18, MLX18, SUF^{+}18, TP17, VB17]. 2018 [AEJZ18, BK19, GHK18, Kiy19, NTB18, RGJdQ18, SPK19, SA19]. 2019 [Ano19i, Ano19j, CJO19, CP19b, IK21, KPBR20, TMH20]. 2020 [Ano20k, Ano20m, Ano20i, Ano20o, Ano20l, Ano20p, Ano20j, Ano20n, BPH20, EZ22, GCvdS^{+}20, KKN^{+}21, KVB^{+}20, LPL^{+}20, Liu21, MCPW21, TBG^{+}20]. 2021 [Ano21i, Ano21m, Ano21q, Ano21k, Ano21n, Ano21p, Ano21r, Ano21l, Ano21o, Ano22j, Ano22k, BV22, CGS^{+}21, KGB^{+}21, OOC22, PB22, RFB^{+}21, RS22, SK22, SLL^{+}21, TKdJO22]. 2022 [Ano22r, Ano22o, Ano22n, Ano22m, Ano22p, Ano22q, Ano22s, Ano22l, AdSMD23, BA23, CWT^{+}23, ECG^{+}22, FT24, GRF^{+}22, IB22, RSK^{+}24, RRB^{+}22, SPO22, TRB^{+}22]. 2023 [Ano23o, Ano23q, Ano23n, Ano23m, Ano23r, Ano23t, Ano23s, Ano23p, GYK^{+}23, SCG23]. 2024 [Ano24b, Ano24h]. 21st [Wol00]. 23rd [Ano05b, SUF^{+}18]. 24th [GMM18]. 25/1 [Ano01-51]. 25/3 [Ano01z]. 25th [KNDT20]. 28 [Ano03i]. 29/6 [Ano06g, Ano06h, Ano06p, Ano06y, Ano06-28, Ano06-44]. 2D [CMLH21, Mar04, WLW05]. 2nd [Ano94a, Ano02-55, Ano07-39, Ano94a, Ano04b, Ano05c].

3 [FR92a, GM86, GV89, HL93, MG86, SG92, SKO83, TMK94]. 3-5 [Ano03a]. 3-D [Ano05d, ÜT99, FR92a, GM86, GV89, HL93, MG86, SG92, SKO83, TMK94]. 30/4 [Ano06e]. 3090 [BGV93]. 3090/180VF [BGV93]. 32 [Cad08]. 33 [CEPS13, SMS09a]. 34th [Ano07e, Ano07d]. 3D [WGS^{+}18c, CMLH21, CCCP04, HR04, KD00, LZP^{+}04, RM05, Shi04, ZD04]. 3DCascade [ALR23]. 3DCascade-GAN [ALR23]. 3DIM [Ano05d]. 3DOR2018 [TT19]. 3DOR2020 [STP^{+}20]. 3DOR2021 [BDL^{+}22]. 3DOR2022 [BTD^{+}22]. 3DPMesh [KCS24]. 3DToothSeg [JXW^{+}22]. 3IA [Ano07f]. 3rd [Ano13n].

4 [HEG98, KMS^{+}97]. 41 [FB15a]. 4D [LWG^{+}23]. 4th [Ano01f, Ano03c, Ano01a, Ano04c].

5-6-7 [AYZ12]. 5th [Ano95v, Ano93a, Ano07g].

60GHz [RGGB02]. 68 [SOdSC18]. 69 [YS21]. 6D [ZHW^{+}21]. 6Dof [ZLL^{+}20]. 6th [Ano01b, Ano01c, Ano03d].

70 [WGS^{+}18c]. 7th [Ano02a].

85 [MDSU88]. ’86 [Ano85b, Ano86g, Hop86]. ’87 [Ano86e, Ano86h]. 8th [DKFC20].

90 [PM22a]. ’91 [Ano91a, Ano91c]. ’93 [Ano91b]. ’94 [Ano94-29, Ano94-30, Ano93c, Ano94c, Ano94h, Ano94l, Ano94t, Ano94s, Ano94-32, Ano94-31]. ’94/CADDM’94 [Ano94t]. ’95 [Ano94p, Ano95k, Ano95l, Ano95m, Ano95u, Ano95v, Ano95-29, Ano95-30, Ano95-27, Ano95j, Ano95z]. ’96 [Ano95q, Ano95r, Ano95s, Ano95-28, Ano95x, Ano95y, Ano95a, Ano95b, Ano95c, Ano95n, Ano95o, Ano95p]. ’98 [SK99, Ska99, SDWE99, SKP99, SKCP99, YJC99, MDS^{+}22]. 9th [Ano94b].

A1 [Ano22r, Ano22o, Ano22n, Ano22m, Ano22p, Ano22q, Ano22s, Ano22l, Ano23o, Ano23q, Ano23n, Ano23m, Ano23r, Ano23t, Ano23s, Ano23p, Ano24h]. A12 [Ano22o, Ano22q, Ano23t, Ano24h]. A14 [Ano22m]. A16 [Ano22r]. A2 [Ano23m]. A2X [SLM^{+}22]. A3GC [PR23]. A3GC-IP [PR23]. A4 [Ano23p]. A6 [Ano22n, Ano22p, Ano22s, Ano22l, Ano23o, Ano23q, Ano23r, Ano23s]. A8 [Ano23n]. abeille [AKF^{+}20]. abilities [COM^{+}94, MGSC^{+}10]. ability [Car93, LWJ^{+}22]. ablation [PSK^{+}11, RGH^{+}19]. abnormal [GWL^{+}20, WXZ^{+}18]. abnormalities [HO12]. above [DCJH13]. absolute [MSMK19, PCKB23]. Absorption [ABG^{+}18]. Abstract [WWL^{+}12, Che06]. abstraction [GDA^{+}13, GNL^{+}15, McW87, Mou13, SD15, dGGDV11]. Abstractions [Eas85]. academia [UKW23]. academic [SGPC20, SET^{+}88]. accelerate [BCG^{+}94, Cas88, GBKG04, KPFT03, Oik98]. Accelerated [KJ08, XPL90, CMM16, FC00a, FBP96, HJW97, LF22, MLM03, SHCW22, TyZfTM12, YK18]. Accelerating [WZW97, LCD15, NIH08]. Acceleration [PDL^{+}21, CF13, Ng95b, WBK98]. accelerators [Mar10]. acceptance [MD99a]. Access [RJS98, BL96, CLH^{+}16, FN99, KPB96]. accessible [MGH13]. account [HJ03]. accumulation [ZNT^{+}18]. accuracy [CS98, FGP^{+}10, KSH17, LZW^{+}21b, Mar02b]. Accurate [CLXJ22, HHCM17, KK16, PvdSLJ99, WTWT18, YZWJ24, FM17, GM02, HHL99, HKYM01, MMS15, TSY11, TVL16, WBP92]. AccuStripes [HGW^{+}24]. Achievement [WF22]. achieving [CGWW16]. Acivs [Ano04d, Ano05e]. Acknowledging [Ano24a]. Acknowledgments [Ano23a]. ACM [Ano01d, Ano01-42, Ano01-57, Ano02-50, Ano04-58, Ano07h, Ano07i, Ano07a, Ano07b, SIG02, SUF^{+}18, Ano94c, Ano01g, Ano01d, Ano02e, Ano02b, Ano02c, Ano02d, Ano03e, Ano03f, Ano04e, Ano04f, Ano04-39, Ano05y, Ano06c, Ano07h, Ano07i, Ano07j, Ano07k, Cou92a, DKFC20]. ACM-CIVR [Ano07i]. ACM/EG [DKFC20]. ACM/SIGGRAPH [Cou92a]. acoustic [ZPIS23]. acquisition [BRdSOS17, CMDS17, FV13, JJPP^{+}22, RZY^{+}20, RLS^{+}12]. acrophobic [JP10]. across [GD11, IKTS22, MPTA^{+}22, ST22]. acting [Per02]. action [FGL23, HWR^{+}23, KSF15, MZCD21, SH94, SWH^{+}17, SKL^{+}13, TC24]. action-based [SWH^{+}17]. actions [GD04]. activation [MZCD21]. Active [BB15, BB03, Fiu89, McC96, YS15, GZSZ20]. activities [GN94, NC07, TMK94]. activity [BSMG24, DBLC02, Mar02b]. actors [NRTT95, PCS00, RCBS10]. Acts [BMdSVR18]. actuator [BMR23]. AD [KNC11]. adapt [dGGV08]. Adaptability [HS03, ZXH^{+}12]. Adaptable [CP10]. Adaptation [DJG^{+}04, NFHS06, WZZ^{+}18]. adapted [KD11]. Adapter [XLW^{+}24]. Adapting [HS00, SKL^{+}13, KA22]. Adaptive [BF02a, BD13, CH94, GSA89, HEK22, LGRP14, LFY^{+}21, LZ14, MZ23, MVS14, PM91, PC23, SMFF04, WCLT21, BSJC02, BTS19, BAS^{+}15, CNS^{+}06, ÇB22, CFZC19, CGMS00, FTB16, FJ17, Fou11, FN99, GA07, JYC^{+}23, KK16, LLLZ16, LLW^{+}19, LOdF02, LWP02, MX14, MYL^{+}23, MWA^{+}13, PMTK01, PHLW15, PAE^{+}21, PR23, Pur87b, SKSZ99, SM99, SK04, SJB^{+}21, SCC11, SWL^{+}16, TF18, WEWL99, WPH^{+}14, WM24, XXHM21, ZPN^{+}21, dF24]. adaptively [ZZCY22]. Adaptivity [Enc95b, OO04, TNFG14]. AdaptMVSNet [JYC^{+}23]. adapts [II22]. ADDI [ESFGDZ97]. adding [ZQ12]. Additional [HL02, Dra98]. additive [HBA13, LCA19, SHD^{+}17]. Addressing [TC24, GBKG04]. Adequate [CGB13]. adjacency [CDW11, PR23]. adjacency-based [CDW11]. adjustable [ZZC20a]. Adjusting [COPR17]. adjustment [FK82, HCV^{+}22, HLO16, LTH^{+}19]. Advanced [Ano94a, BS09, CSS^{+}24, FBT93, FASS16, GS04, HKS00, KDS04, Mit77, Wes94, Ano04d, Ano04h, DWH^{+}15, Gag95, GHM^{+}96, HR97, SR95, WKO12, Ano02f, Hol94]. Advances [All77, Ano07e, Ano07d, Man99, RMSB22, Ano13l, CLMA19, KCU^{+}22, MLM03]. advancing [CKCK09]. advantage [RWD14]. advection [WWL16]. adventure [KCR02]. Adversarial [FNM20, XWF^{+}20, dSMBG23, FCG^{+}21, MRWL23, NZZ^{+}21, OBD^{+}23, ZZF^{+}23, ZLL^{+}23b]. advertisement [Gro92]. aeolian [Pic94]. aerial [ACC^{+}18, SBD15b]. aeronautical [Tes84]. Aesthetic [JWL12, Nap95, McW91b, SCNT03, TY24]. Aesthetics [Ano12s, CIW09, BCMD17, Nik06]. affect [CPS^{+}22, DAG22]. affected [KP95]. Affective [LB12]. Affine [RBB^{+}11, FPC10, Jac95, JSG04, Lan97, Mar01, Mar03, Mar04, Mar09, dCNPdFS14, PdFS06, VT06, VMAL16]. Affine-invariant [RBB^{+}11, VMAL16]. affinity [HZC^{+}22, WWS^{+}13]. Africa [AMC03]. AFRIGRAPH [Ano01e]. After [Jac93, WB24]. afterimage [MSR^{+}13]. against [MD99a, WHH06]. agency [OO04]. Agent [vTCB^{+}21, G^{+}01, OP15, RP20, SBS19, SLM^{+}22]. Agent-based [vTCB^{+}21]. Agents [CF96, dHT01, Ano03-37, Ano03c, BCS08, GS01a, GKT02, MSAR01, RJS98, TB18, ZI00]. AGF [Wis87]. aggregated [HZLC22]. aggregation [Bou06a, WPL^{+}23, WWS^{+}13]. aging [CP10, MG08, MB14]. agnostic [BNS24]. ago [Wil86]. Aid [Per84, Pic84]. Aided [DIE78, Dro78, GS84, GSME04, LS79, Mar79, ME83, Str85b, Tan94, CGH97a, CGH97b, Coo79, Cor84, DKM^{+}20, GHM^{+}96, HKPL98, JK90, JZR88, KAFB18, KAV^{+}88, LKL^{+}20, MMV88, PGB86, PGR83, Shi06, Tan80, TMH20, VBVS88, WF88]. AIM2003 [Ano03g]. Aims [EHBM82]. air [CPC^{+}18, CS18, KK21, OK20, VR16]. airborne [HLZ^{+}17]. Aircraft [BSPR77, DIE78, Dro78, SS75a, Tes84]. aircraft/store [SS75a]. airflow [KSH^{+}19]. airfoils [RAHA88]. airport [AD85]. Alan [Hol94]. Alexander [Adz22]. Alexis [CL96]. algebra [ANE17, Gin02, Pra92, ZD04]. Algebraic [PK85, MSE20, Mok87, Mok88, PMZS97, PDS21, WYXM22, WF11, WM89, WGLS00, ZLS98, ZLS99]. ALGOL [DEW75]. Algorithm [Ger86, MK89, PR93, SEDT^{+}03, VM15, AnD19, AC89, AS91, Ang97, AFW^{+}18, APSS01, AÇÖ96, BP94, BJP97, dLBRM^{+}12, BC13, BK93, Car92, CBG22, Cet23, CS98, CYCL09, CC01b, CGZZ15, CH94, CY94, CN05, Day92, DR15, DWH^{+}15, Eas75, FS86a, FRWW14, FS98, Fle91, Fou11, FDA03, FBP96, Fun99, Goe95, Gom14, GS89, GZL21, Hor82, HH91, IKB00, JV91, Jas88, JPP01, KL07, rKC93, KLP01, Kla91, Kle86, Koh96b, Kru99a, LŽ03, LL92, LJCW04, LWP02, MMALRA01, MFOK94, Muk86, MTM22, NY06, PMM18, PBN97, PR82b, PYD^{+}05, Pie83, Ran87a, Ran91, RMG15, RGGB02, RM91, Sah15, SSV07, SMMS01, SR02, SGC00, SF98, SH94, Sen98, Sen99, SZ95, SLL^{+}17, SDD95, SEC90, SZEG93, SG15, Ska93, Ska94, Ska97, Sla92, Smo03, SPY87, Str86, TSD87, TFY00, TPN95, TT82]. algorithm [TVS^{+}03, TC00, VHS12, WYZ^{+}11, WJ91, WBP92, WTL^{+}11, WGLS00, WHH06, XD08, XL10, YHX10, YS97, ŽC99, ZSW08, ZP92, ZS02, ZCZ^{+}18]. Algorithmic [VPBY02]. Algorithms [And85, DGR93, FS98, GP86, GD87, Haz79a, HL93, HN85, Mok88, SGM97b, VD98, ATB98, AEA13, AG13, Boe91, Bor91, Cas88, CCM^{+}07, CMS98, Fos87, GIZ95, HP03, JALS03, Kau88, KHK18, KNMP14, LPV95, LT95, LWG^{+}23, MFL11, Nav89, PGR83, PP20, PS91, PF80, PPSS96, QN98, SGM97a, SM92, SA86, SKM98, VK07, WEWL99, YR98, vTP20]. aliased [CL06]. aliasing [IV93, MA18, SG92, SMSS13]. Aligned [CSL23, CGH94, LXW^{+}10, LSHL18, MCS^{+}18]. Aligning [LBLD11]. alignment [BND^{+}17, CSL18, KCS22, LWD^{+}18, LSS21, MG09, Med86, PCGS15, SJZ^{+}23, WYZ20, ZSS^{+}18]. allows [Par75b]. along [CGH94, FPR92, MS07, WA02]. Alphabetic [Ano15a]. alpine [LWZ^{+}23a]. alteration [QD03]. altered [YSDG24]. Alternate [PMPR^{+}16, Sar03]. Alternately [XSW23]. alternating [CBM^{+}22]. alternative [SGC^{+}19, Sar92a]. Ambient [TKOD24, Ano04g, AMGA12, CC19, DVND10, LVVC06, WCH^{+}11]. ambiguity [MAH00, ZCT95]. AMDO [Ano06-35]. AMDO-e [Ano06-35]. American [PPV03]. Ames [EHBM82]. among [GKLM07]. amount [Car92]. amounts [SKH^{+}05]. amplitude [Sen99]. amplitude- [Sen99]. Amsterdam [Ano86h]. analogs [CR07]. analogy [Cho06, GRIG12]. Analyse [CF77, CV77, RPM96]. analyser [VH02]. analysing [MRG^{+}19]. Analysis [Ano77, Ano01-58, Ano02g, Ano02h, Ano03-56, Ano04-60, Ano06n, Ano06-51, BRSP15, BSPR77, BS77, BG80, BS82, BKS23, CS85, DBG93, JH11, LM22, Loh95, MAdS^{+}19, MPS85, NYKN83, PPS20, Per84, SKM98, AMFH21, Ade86, AYZ12, ABG^{+}18, AJ94, ATAG^{+}21, ALC06, AK13, ANGH11, Ano01-28, Ano03-55, Ano04n, ASR^{+}22, AM19, ALM19, APSS01, AF11, BNPS10, BHZ^{+}21, BKS21, CP21, CTJ^{+}14, CTLG94, CH96, CEN^{+}23, Cor76, COSEV22, Cot75a, DDM^{+}06, DBS^{+}11, DSB96, EME15, FP75, GZSZ20, GYD75, GD95, Gro91, Gro92, HJW^{+}08, Han84, H^{+}00b, HBG14, HK15, HHZ^{+}22, IMMS82, ID17, KBL22, KGM75, KF88, KGGP19, KAVM23, KD00, KD15, LBLD11, LXB^{+}15, LM16, LWG^{+}23, LRR87, LRHS14, LSR22, MLM^{+}17, MJEG21, MTS^{+}22, MCMT14, MS82, MSD75, MBST22, MPL21, NC07]. analysis [NFLYCO99, OST^{+}16, OL96, Paq05, PiP00, PPP88, PS13b, PPL91, PM84b, PWJ^{+}18, PWV^{+}18, Pol83, PRM^{+}24, RAHA88, RAK^{+}15, RBG^{+}09, RdCVL16, RŞH^{+}22, RGH^{+}19, SRA^{+}19, SMS09a, SMS09b, SJB^{+}21, SC97, SHD^{+}17, Ste75, SBR^{+}22, TMP07, VBS^{+}15, VT07, WHH21, WZL^{+}22, WUH^{+}15, Wel76, WZZ^{+}18, WSX12, XLM12, YHW23, ZLL^{+}21, ZC07, ZZDZ10, vTTK^{+}20, Ano04b, Ano05c, Ano07-37]. analysis/synthesis [MS82]. Analytic [AMGA12]. Analytical [Haz79b, JT02, ABCO12, RH85, RAF21]. Analytics [BA23, CEN^{+}23, FGM^{+}20, KKMT06, NHR^{+}22, VT22, AS22, BN07, CTJ^{+}14, EBST14, FMCM^{+}21, GSF^{+}19, GTdS^{+}18, GPTP10, GWBD17, HMA23, HK15, JSMK14, MRW^{+}21, MRG^{+}19, MA14, RWD14, SVVS^{+}17, WK14, WJD^{+}09, WWO^{+}23, ZZL21]. analyze [Cas96]. Analyzing [CLH^{+}23, EHM84, EBST14, SFS^{+}21, BSC^{+}21, PL97, SK23]. anamorphic [PJP23]. anamorphosis [Lan97]. anatomic [DEST95]. Anatomical [BND^{+}17, FMCM^{+}21, LD11, PPS20, STBG19]. anatomy [ILLC01, PS18, SCSG18, ZSM^{+}19]. anatomy-based [ILLC01]. anchors [LLH17]. ancient [APA^{+}11, CHZ^{+}23, LXPP06, RM22]. Andrew [Cyc93]. Andries [Mil92b]. Android [NSL16]. anesthesia [SHL^{+}24]. Aneulysis [MPL21]. aneurysm [KGK^{+}07, MPL21, WCA^{+}11]. aneurysms [MVPL18]. angiocardiographic [K^{+}00a]. angiography [HTW^{+}19]. angle [XGC18]. angles [MA94, SRZK23]. angular [KLP01, SVW23]. animal [LNL^{+}23b, LNL^{+}23a]. animals [EBC^{+}15]. animate [TPM14]. Animated [LCL^{+}21, MTT84, BCS08, BAC14, CYJ^{+}13, GS01a, JM88, LAL11, LLX^{+}15, MSAR01, PA07, STBG19, WIP08, ZO07]. Animating [Fos87, HCLC16, LZL^{+}19, LDG96, MP89, NLS07, PQ10, PP02]. ANIMATION [Ano01-34, Ano94n, Ano94o, Ano12s, CI11, Cou92b, CS85, DM79, Gom85, GÖT93, Hun77, ILLC01, Kit77, Lin97, MK85, MRR98, Sug83, WAM17, AK06, ACSW75, ANE17, Ano01-33, Ano06-47, AN99, BTC94, dPCOO^{+}05, CBU^{+}15, Com85, CO88, CF96, DMV06, DJG^{+}04, Dur89, Dur91, ET07, GKW^{+}24, HEG98, Hol94, HY03, IC96, JK90, KG04, KR92, LW89, LL11, LD12, LKC98, LZLS18, LG89, MPQG18, MHK99, MKHN01, MAO^{+}12, MY16, Mor75, NLG20, NK01, NUM24, PF97, Par75b, PCS00, PSS04, SOC^{+}19, SOG08, SCSG18, TG02, WLP^{+}14, YSZ22, ZZC^{+}14, ZM07]. Animations [PM22a, ASS22, CSP19, CMB17, JRZ^{+}23, MSRB17, PM20, PM22b, RPP20, dAU14]. AniNex [CWT^{+}23]. Anisomorphic [AA07]. Anisotropic [SHZ19, XXX^{+}23, YLT^{+}18, CZZ22, EWWL00, FB15b, LH14, PVM^{+}22, WXL^{+}23]. anisotropy [MF02]. ANNOR [KB15]. Annotated [MDM^{+}21, CLH^{+}16]. Annotation [PGS^{+}23, WDH09, WWWW22, AMPG22, KB15]. annotations [LDM^{+}11, LSE18, SGBP17, SLRP16, VTW23]. Announcement [Ano77, Ano95q, Ano95r, Ano95s, Ano01g, Ano01j, Ano01h, Ano01i, Ano01k, Ano01l, Ano01w, Ano01u, Ano01v, Ano01q, Ano01r, Ano01p, Ano01s, Ano01t, Ano01o, Ano01m, Ano01n, Ano01x, Ano01z, Ano01y, Ano01-27, Ano01-28, Ano01f, Ano01-29, Ano02f, Ano02e, Ano02g, Ano02h, Ano02i, Ano02j, Ano02l, Ano02k, Ano02o, Ano02p, Ano02m, Ano02n, Ano02q, Ano02r, Ano03-30, Ano05e, Ano05f, Ano05h, Ano05g, Ano05i, Ano05j, Ano05k, Ano05l, Ano05-27, Ano94p]. Announcements [Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano02s, Ano06d, Ano06e]. Annual [Ano95x, Ano95y, Ano95w, Ano01-59, Ano06-27, Ano04a, Ano04-58, Ano03h]. anomalies [SRA^{+}19]. Anomaly [JSMK14]. Anonymizing [Kur24]. anthropomorphism [DGC^{+}21]. anti [CL06, LLZ^{+}23, MA18, SG92]. anti-aliased [CL06]. anti-aliasing [MA18, SG92]. anti-spoofing [LLZ^{+}23]. Antialiased [HN85, CC01b, LK00, MX12]. antialiasing [GH98, VH02]. Antics [Kit77]. ANTS [RCD^{+}04]. ANU [LTC^{+}20]. ANU-Net [LTC^{+}20]. anxiety [JP10]. any [CNS^{+}06, II22, Seg88, YAKE23]. API [CS04, MGH22, NSG05, SK04, UKW23]. API-level [MGH22]. APIs [TBM^{+}04]. APL [GN80, GS87, Haz77, Pol83]. APL-Graphics [Haz77]. APM [WCLT21]. Apollonian [Bou06b]. App [NGAS23]. apparent [Baw97]. appear [Nij04]. appearance [AYA^{+}20, ABAA22, BDSP22, CLMA19, CYWM23, KB10, KK21, dSNJA22, WLL^{+}21a]. appearance-driven [dSNJA22]. Apple [Cou92b]. Appleton [HRGD88]. appliances [HK04, MSAR01]. applicable [BSAH^{+}23]. Application [Dai93, DW82, DTZ09, Gra85, GA88, MCM^{+}18, SGM97b, SKO83, Ste75, TNU^{+}01, Tes84, TBM^{+}04, Ano02-53, AG94, AAB92, CDPS06, CGR87, CFMS02, CMLR11, FGES96, GD00, GLT^{+}97, GB75, HAL^{+}21, HMHB08, IU09, IT11, JT86, JY98, KBL22, KKC94, LZLS18, Mar86, MRF06, Mil03, Mit87, MUH10, PPL91, PP90, RCB15, Rau06, RBF17, ST20, SGM97a, She12, SKSI95, SZW10, SPS96, SLK^{+}23, Van89b, VR16, ZW20]. Application-Oriented [Dai93, Gra85, Mar86]. Applications [All77, Ano94p, Ano01-54, Ano03-51, Ano06c, BGT16, Bjo85, FGMR22, GPS18, GMMP21, Kar92, LPS19, PBB21, ACP20, AJ94, Ano04p, Ano04-54, AFM93, BKR^{+}16, BBDM03, BCC^{+}22, BEFV94, CP19a, CTN^{+}17, CG87, CIK99, CUD06, DQF04, DL93, DHZL20, FGES96, FP87, FSS^{+}02, FN99, GBF14, Gin93, HMW91, Jac95, Kal04, KDS04, LZSG03, LJWcH07, LY08, LLS^{+}16, LPZ^{+}21, LWG^{+}23, MOS^{+}21, MGMB22, MGJ^{+}11, PP16, RE22, RFR02, SSV07, SdSR^{+}19, SDIM13, Sri02, VBP05, VCHR07, War76, WKS03, ZR12, ZTS02, ZGdDL^{+}96, dSJ23, Ano93b, Ano95q, Ano95r, Ano95s, Ano02-50, Ano12n]. Applied [SB77, BKV05, FSV17, Fou11, ST02, SS75a, TAF16, dCdLL14]. Applying [MMSS03, MA14, Tay87, MR17, WJD^{+}09]. Apports [Nan77]. apprentice [Eck90]. Approach [Bar77, Car84, Eas85, Ent77, Jer85, ME77, ASM12, ARL^{+}20, AM19, BBCG11, BBMGM22, BRdSOS17, BMT96, BCC10, dSC07, CPM19, CVL^{+}04, CS03, CSZ92, CCW13, CGWW16, CCM^{+}11, CSK97, Cho06, CR13, CK75, DVF06, DG06, Dor90, DKM^{+}20, EW99, FST97, For84, GWP00, GPTP10, GWBD17, GZW12, HWFQ09, HS14, Hor82, HL06, IA99, JZLP23, Kas87, KKJ^{+}23, KJ01, KCS24, LK18, LKL^{+}20, LWW08, LYZ15, LP92, LST96, LKC94, LWZ^{+}23b, Loh95, MKDM22, MQW09, Mar04, Mar09, MBA20, MPM15, MKPM17, MMS15, MCKS06, MCP^{+}22, MKHN01, MY16, MBV18, MTM22, MSO^{+}20, MKKM18, NS87, NMM09, Nov03, OFP^{+}11, PP16, PPM18, PMS87, PTL04, PRW^{+}22, PM95, PCPW15, PPSS96, PM13, RP22, RMD11, RR15, RAG05, RPP20, Sar03, Sch85a, SB86, Sen03, SWZZ23, SK16, STdKB11, SvL09, Smi75, SBS22]. approach [SPL14, Sta87b, SEKA19, SGZ10, SMM20, SS04b, TT12, TS87, UPTd92, VV89, WTM12, WMZW22, WSG10, WMW13, WZC^{+}24, XGZ11, XCZ^{+}16, XLQP12, YHW23, YAKE23, ZNGN16, Zay12, ZGW^{+}16, ZZH^{+}24]. Approaches [AB78, HF85, Neg77, NT00, BKR^{+}16, BSAH^{+}23, CACC24, DAHF04, GOdSC23, MSE17, RCBS10, SKSZ99, SF92]. Approaching [Gna84]. Approximate [Aum89, GO06, FH94, GLDK95, dGMW16, SL12b]. approximated [GTG06]. Approximating [dCNPdFS14, Mar09, Spr04]. Approximation [Ala85, DFNP84, DG96c, EH96, HY93b, BRSP15, BPS^{+}10, BMMZ23, LD09, LZ11, LJWcH07, LOdF02, Mar02b, MLP01, ML12, NAS18, OKBG08, PS12, Pat15, PC23, RB20, SXSY14, WG17]. approximations [BMMZ23, DZ91, KH00, KRRS12, PF09, SP23]. April [Ano20k, Ano21m, Ano22o, Ano23q, Ano24b]. AR-based [SBHC22]. AR/VR [MOS^{+}21]. arbitrarily [AA00]. Arbitrary [Cor82, EKP93, PQCT23, TFF^{+}20, ASC17, ASC18, BLNZ22, CDW11, CVB16, CM15b, CYM16, CDGC94, KC07, Kle86, LQ12, MW12, MS82, NAO13, Pet18, PR11, RLB^{+}02, SR15, Seg88, TPM14, XWF^{+}20, YJLZ21, dMTB^{+}21]. arc [CGH94, MW92, TT83, ZZY12]. archaeological [DSR11]. archaeology [IT11]. archeological [CLE13]. Archetype [MO90]. Archetype-oriented [MO90]. ARCHI [DZZ79]. Archimedean [ZEK^{+}17]. Architectural [DZZ79, Her85, PEVW15, ASKCK03, BBCG11, DVF06, DBS^{+}11, DA18, MBA20, ÖOK23, SS96, Shi06, SF92, YMYH12]. Architecture [AP88, Ano12p, GP86, Mar86, Rix84, AKPS00, AAK^{+}22b, BK89a, Ban85, BH91, BES00, BL96, CGR87, CVHM03, DH95b, EHSF17, EMB^{+}98, FB04, HIK05, HG21, HGJC21, JJL97, Kni95a, KS15, KNMP14, LXPP06, MPP98, MR95, MGH22, MNSJ99, OCR^{+}19, PWK95, Rog85, SEDT^{+}03, San93, SLS03, SR95, SDD95, ST02, SvLBF10, SBD^{+}94, WLYH19, WSHY22, WCL23, XWW^{+}18, dGHM97, PKK03]. Architectures [BOH97, MPR89, MB97a, TPB08, Ano94a]. archival [Kir93]. archive [LBLD11]. arcs [HE80]. Arctic [Sul85]. Area [BP93, DVND10, KK08, BR07, Bru75, DTZ09, DK92, GWEA^{+}11, GN94, Jas88, KJS17, KJS18, LZW^{+}21a, Ran87b, Str86, ZXL^{+}21]. Area-efficient [KK08]. areas [DSG21, WUH^{+}15]. ARF [Mit87]. arguments [RPÁM04, RPAM06]. Arithmetic [Ala85, ML79, FPC10, LP93, Mor76, dCNPdFS14, PdFS06, RHK^{+}20]. arm [Car92, MH21]. Aroundplot [JHPhR11]. arrangement [RFZ^{+}17]. Array [BY88, AC89, KK21, LP93, RR92, YC10]. array-based [YC10]. arrays [CMdL21]. Art [ACB12, CAS^{+}15, Her83a, Mar79, Mas92, MSRB17, AXG^{+}13, Ano83a, CCM^{+}07, CS01, CLT07, Ent89a, Fen88, Han97b, Hel95, HK15, Hol03, HH88, JCFN18, Kal04, Kel00, Ley02, MD99a, Mar07, McW91b, MPAC^{+}23, Nap95, Nol95, OC00, Rad96, Sac22, Sel89, Szy90a, WFC^{+}09, Wal93b, Wal93a, Wal94b, IIK12]. Art-directable [CAS^{+}15]. Art-directed [MSRB17]. ARtention [PAE^{+}21]. Articulated [MK85, ASS22, Ano06-35, BHTT94, BTC94, SKL^{+}13]. articulations [SKSI95]. articulatory [dJONM18]. artifact [AOL96]. artifacts [OFP^{+}11, SBS19]. Artificial [AS22, Ano07f, AMS22, BYC19, NK01, Spr98, CWGR01, FB12, MRS^{+}07, MKHN01]. Artist [Gin02, MHLB16, MMH^{+}21, Ley02, McW87, Pic92b]. Artist-oriented [MHLB16]. Artistic [Kei89, Car99a, GA07, GR09, KKH23, Loy91, Ye02, ZLLY06]. Artistry [Ano86a]. Artists [Cra02, CMA10, DMM23, MRS^{+}07]. artophile [CMA10]. arts [Ebe00, PE16]. artwork [DMM23]. ARVIKA [WKS03]. ary [ZGC15]. As-conformal-as-possible [PP12]. as-isometric-as-possible [ZLL^{+}15]. As-rigid-as-possible [CL18, DKV^{+}22, BCDD22]. Asia [Aok16]. ASIC [WWD^{+}95]. Aspect [JNR85]. Aspects [CC01a, NM85, BT94, Fra83, LAB^{+}14, Wol02a]. Assembling [KP09]. Assembly [LS90, DJC^{+}23, GZ99, KL02, LFP10, LUB^{+}13, TH90]. assessing [SLM^{+}22]. Assessment [KHTM17, Ng95b, dJONM18, DFF22, FWX^{+}18, LSS21, MSE17, PAE09, RMW^{+}17, WBJ^{+}21, WTM12, WLX^{+}23, ZK95]. assessments [HMA23]. assignment [CM14]. assignments [San06]. assimilation [KAV^{+}88]. assist [PP20, UL22]. assistance [HKS01, KS01b, LFP10]. assistant [MSAR01]. assistants [CIK99]. Assisted [CMDS17, PGS^{+}23, SPK19, Smi77, BMH99, Cor76, HBOS13, KWK23, LD11, NMM09, PSH^{+}09, SHS^{+}23, vLLSM09]. assisting [HWYL21, Spr75]. Associate [Ano11a, Ano13a, Bon03, Bon05, Ano24c, Bon95, Bon97, Pic92a]. associated [Ano06-33, Dra98, Med86, Mil93, Pic87a]. Association [Ano03h, LPPM07]. Associative [Wec79]. assumptions [ST02]. asteroid [DW89]. asymmetric [HG22, HZ23]. asymptotic [LLP^{+}21]. asynchronous [BBMK21, WIP08]. AT&T [You89]. Athens [KH86]. Atlas [MAS14]. ATLAS2000 [FMS98]. Atlases [FMS98, SSH16]. atmospheres [SGGC05]. atmospheric [GSMA06]. Atom [ST23]. atomistic [GNL^{+}15]. atoms [ZS14]. ATR [TMK94]. attached [KKHS03]. attacks [WLL^{+}24]. attempts [RBLB21]. Attention [HZLQ20, LTC^{+}20, LHH^{+}21, LCX^{+}23, PQCT23, PR23, CGWZ22, DZD^{+}23, JYC^{+}23, MMdOE^{+}22, MRWL23, MML12, NLdAL^{+}23, PWV^{+}18, SXL^{+}23, SLL^{+}23, VZP22, WCHM22, WRLZ23, WCL23, WB24, ZZF^{+}23, ZLL^{+}23b, ZHW^{+}21]. Attention-based [HZLQ20, LTC^{+}20, WRLZ23]. Attention-oriented [PR23]. attentive [ZPN^{+}21]. attraction [Joe92]. attractiveness [BEFV94, DGC^{+}21, XWW^{+}21]. attractor [Mar03, Mar09, Spr96]. attractor-wrapping [Mar09]. Attractors [Rei97a, Can94, CGR98, CC03, DHJ^{+}99, Goe95, HHL99, JR00, Mar01, Mar09, PdFS06, Pic88a, Rei97b, Rei02, Spr93, Spr98]. Attribute [Aon90, STW82, AM19, BSAH^{+}23, Bre01, TZvD^{+}21]. attribute-based [BSAH^{+}23, TZvD^{+}21]. attributes [Cad08, CWNA08, LXJL21, RCPB01]. attribution [DBS^{+}18]. audience [Cun00]. audiences [H^{+}00a]. Audio [AMPG22, AC77, Ano07-33, BCS^{+}99, CIT^{+}19, FCG^{+}21]. Audio-visual [AMPG22]. audiovisual [WLL^{+}24]. auditory [SZEG93]. augmentation [CHZ^{+}23, HSE10, RGE07, ZWWC23]. Augmented [AC77, ABK11, GANM21, HWEB22, HSTR20, KB12, MAM^{+}24, MTS^{+}22, NTTF21, PKK03, PS15b, RCD^{+}04, TKdJO22, ABAA22, Ano01-42, Ano03-36, Ano04-39, Ano12l, AP99, A^{+}99, BND^{+}17, BCS^{+}99, BSMG24, CIT^{+}19, DSG21, Dor99, ET18, FEBS07, FHFG99, FFP^{+}21, GWEA^{+}11, GPC^{+}17, HZ15, HEW^{+}18, HFT^{+}99, H^{+}01, JP10, KS03a, KSR99, LDM^{+}11, MGSC^{+}10, MML^{+}11, PvdSLJ99, PAE09, PAE^{+}21, PTY^{+}16, RPP21, RSP^{+}19, SSM11a, TBS^{+}23, TSY11, TOY^{+}14, WCS23, WKS03, WDH09, WTA11, YRD^{+}24, ZSM^{+}19, TPK13]. Augmenting [Sei99, RR01]. August [Ano20m, Ano21q, Ano22p, Ano23r, Hop86]. Australia [Hvi86a, Wil86]. authentication [SZL^{+}13]. Author [Ano95-42, Ano01-30, Ano02w, Ano03-61, Ano04-63, Ano80, Ano83c, Ano85e]. Authors [Ano11k, Ano15a]. authorship [Pea02]. auto [RFZ^{+}17]. auto-arrangement [RFZ^{+}17]. autocompletion [CDF14]. Autoencoder [MKM19, LZZ^{+}19, OBD^{+}23, LDLD22]. Autoencoder-based [MKM19]. autoencoders [JK21, LCL^{+}23]. AutoGraL [BDRV01]. AUTOLAY [Pra99]. automata [BW98, CR03, CK93, CV97, Fra94, Li89, Mas92, Rei10]. Automated [Bar77, CSL18, GBA15, MLC^{+}22, MPAC^{+}23, OGSSLM^{+}07, CLT07, CR13, Med86, WUH^{+}15, XGZ11, YY88]. Automatic [And85, ADT^{+}16, BRdSOS17, DFWW15, EPB^{+}19, GDDA13, HCV^{+}22, HZD^{+}19, KCH^{+}22, Kuo01, LRD07, LMC13, LYS^{+}19, LW88, LP83, MHW10, MMV^{+}14, OVWK16, Pic94, QWC14, RCB15, RPSP^{+}19, SAB12, SW21, SL12a, SSM11a, Spr93, Spr94, SP95, ÜT99, WLL^{+}12, WWH^{+}21, YZ17, YLS^{+}21, ZPP^{+}23, AJ94, AAK^{+}22b, ABJ90, CUD06, DSG21, FK82, GD95, HWFQ09, HS14, HAB75, LLL^{+}23, LUB^{+}13, LPV92, MSMP12, Pha89, PCGS15, SR02, TDR^{+}17, TH90, VFSL06, YF09]. Automatically [AWI^{+}09, ZZC^{+}20b, LLLC11, MD99b, SCH^{+}18]. automating [ESFGDZ97]. automaton [Ada97, McC02]. automotive [BDRV01, Gin93, OH83]. autonomous [FvdPT01, GS01a, GKT02, KCH^{+}22, PCS00, SFC01, ZTAP21]. autonomously [ZBP^{+}18]. autoregression [WM24]. AUTOSIGN [DKM^{+}20]. autostereoscopic [MBPF12]. Autumn [Loy91]. aux [Le 77]. auxetic [OMP^{+}18]. auxetic-inspired [OMP^{+}18]. auxiliary [QWC14]. avalanche [MMSS03]. avatar [DGC^{+}21]. avatars [D^{+}01, LCL^{+}21, NLG20, PCS00]. averaged [BF07, Rag09]. Averaging [Bar77, CMLR11]. aviation [For84, RCB15, SL02]. Avions [For84]. avoidance [LKC98]. Avoiding [PHO^{+}23, LMW06]. Award [Ano10a, Ano11a, Ano11b, Ano11c, Ano13a, Ano13b, Ano13c, Ano21j, Ano22k, Ano23k, DZ93, EDKS96, Ros92, Sou94, WS22, YQY90, Ano02t, Ano03k, Ano04m, Ano99c, Ano01-36, Ano06m, Ano09c, GLDK97, WF22]. Awards [Ano08b]. Aware [Ano12p, ASZ^{+}14, BWYZ24, CP21, CIK99, CB10, DYW^{+}22, DFF22, FFP^{+}21, GCYX23, GLS^{+}20, HDZR23, HCC13, JCFN18, KNC11, KJS17, LQ12, LZG^{+}23, LLM^{+}16, LZYQ22, LSWL13, LWZ^{+}23b, LSW12, MZ23, MRW^{+}21, MSHL22, MGJ^{+}11, PLJ^{+}13, PLJL15, RTB^{+}18, RGH^{+}19, RMSC11, Sac22, SM22, SSQL24, SBS13, WZ09, WHL^{+}09, WLL21b, YYG16, ZZL21, ZSL12, ZQL15, ZGWP16, ZZD^{+}19]. awareness [BSGT03, GN04, HJ03, KH03, KD03, LLP^{+}22]. away [CD23]. AWE [SL02]. axial [TH09, YAKE23]. axially [BR07]. axiom [CUD06]. Axis [MBN21, BAS^{+}20, CXT18, Egh83, ML12, SMM20, WKT21]. Axonometric [Sug84].

B [AFW^{+}18, BLNZ22, BS93, BW92, CY94, CS80b, CF13, DBG92, DTG15, FRWW14, KDG96, KKH23, Lam00, LP83, LYS^{+}16, MW12, Pha89, PS91, PGVACN06, SWZZ23, SGZ10, Wal94a, WZT97, WSHY22, WF11, Woo87, WWF^{+}18, XLM12, ZCC^{+}16]. B-Rep [KDG96, SGZ10]. B-spline [AFW^{+}18, BS93, BW92, CY94, CF13, DTG15, KKH23, Lam00, LP83, LYS^{+}16, MW12, PS91, PGVACN06, SWZZ23, Wal94a, WZT97, WF11, Woo87, WWF^{+}18, XLM12, ZCC^{+}16]. B-splines [CS80b, DBG92, Pha89, WSHY22]. B3 [Chu90]. B3-splines [Chu90]. Back [GD00, ENE11, KMGL99]. back-face [KMGL99]. back-tracking [ENE11]. backface [HIK05]. Background [Mac77, NZZ^{+}21, MFL11, SL12b]. baggage [WTW^{+}23]. baking [SSDS12]. balance [CFMS02, SCB^{+}24]. balanced [KFW16, XWW^{+}19]. Balancing [CD93, Dod09]. ball [Mar03, MPW^{+}12, RE22, VR06, WWF^{+}18]. balls [LCK16]. band [CZZ22, CK75, GPTB02, Hod91, RP18]. bands [HO94, PRÁM04]. bandwidth [CYKK09]. barcode [CZCG04, DHZL20]. Barycentric [Ska08, BPKB15]. base [Big86, WH96]. Based [Ano07-38, Ano12s, Bas77, Bax77, BP08, RBF20, Str85b, AALJ20, AMPG22, AK06, AKF^{+}20, AL98, AMR23, APS09, AC90, dSASP^{+}15, ARL^{+}20, ANGH11, Ano04-51, Ano06-49, ALD12, AVHT17, ABCO12, AVM05, ACGC22, ADT^{+}16, AÇÖ96, BDPR93, BSJC02, BKR^{+}16, BJP97, BX99, BF19, BDSP22, BSAH^{+}23, BWD13, BKCW14, BERW97, dLBRM^{+}12, BES00, Bim15, BBE14, BN03, BK23, BSL^{+}13, BP10, BCMM07, BJS01, BAS^{+}15, Bri95, BEKL00, Bru14, BRPC18, BAC14, BD17, BRO22, Bus97, BBP10, CZB^{+}22, CDW11, CGS^{+}21, CSP19, CIT^{+}19, CLM^{+}19, CTN^{+}17, CTS^{+}10, Cet23, CFH^{+}18, CK09, CSZ92, CHSB10, CS16, CZ22, CHZ^{+}23, CHC^{+}24, CWTL08, CGZZ15, CD15, CSLY01, CAL^{+}04, CPCS90, CYKK09, CG96, CN05, CMS22, CCCP04, CDGC94, CGW07, CLX^{+}19, CV97, DMM23, DLS^{+}11, DCV98, Dav95, DG96b, DVF06, DBS^{+}11, DMV06, DL09]. based [DW13, DMG20, DWX^{+}23, DFF22, DLZY14, DWH^{+}15, DGA02, DDPT98, Duc18, Eck90, ELC^{+}12, EBST14, EK15, EMB^{+}98, FBH^{+}01, Fat01, Fat02, FR98, Fen88, FWX^{+}18, FB14, FB15a, FIC21, FWW13, FR88, FJS11, FJ03, FM09, Fou11, FM17, FFV^{+}11, FK11, FLM^{+}15, Fun99, GWX^{+}18, GdMF03, GMP22, GV89, GI13, GBP^{+}17, GSV^{+}18, GKT02, GA12, GvK18, GA83b, GHFH08, GLDH97, GGW22, GWBD17, GLT^{+}97, GLZ^{+}21, GTFB23, GZL21, HO12, HZ15, HDZR23, HS03, HKL^{+}23, HG21, HLO16, HHLE17, HGJ16, HHZ^{+}22, HR07, HWR^{+}23, How88, HW16, HLL^{+}18, HZLQ20, HZLC22, HZD^{+}19, HHN^{+}23, HYP^{+}24, HMHB08, HY93a, IP03, IP23, IC96, ILLC01, II22, IKTS22, IKB00, IVCN20, JH89, JK15, JZLP23, JZX^{+}21, JXJ22, JLP00, JYLW14, JAS97, JY98, JHL^{+}12, JCFN18, JRJP^{+}22, KPA12, KBL22, KSM07, KPSN04, KS91, KCU^{+}22, KES22, KM14]. based [KSR99, KK22, KJ01, KB04, KPL08, KGK^{+}07, KDG96, KEVD18, KB15, KJ08, Lam00, LŽ03, LSW^{+}23, LNFC95, LD09, LNL^{+}23b, LNL^{+}23a, LMR^{+}19, LS90, LKS07, LdSP^{+}08, LGWM17, LCDN06, LW24, LV02, LW89, LS07, LJCW04, fLhLfT11, LLLC11, LL11, LLLZ16, LWW^{+}20, LTC^{+}20, LCL^{+}21, LMY^{+}21, LWWY22, LZY^{+}23, LWZ^{+}23a, LZG^{+}23, LYZ15, hLfTxDdZ09, LSE18, LP92, LTH^{+}19, LXT^{+}23, LZT^{+}24, LPP^{+}19, LD12, LS06, LLS^{+}16, LWD^{+}18, LPO20, LYZ^{+}22, LWZ^{+}23b, LSCJ23, LLL^{+}23, LHL23, LRY^{+}24, Liv18, LRG11, Loh95, LSS21, LMJH^{+}11, LUB^{+}13, LWP02, LPL^{+}05, LSR22, LPV92, LZL^{+}24, MWLZ22, MYL^{+}23, MA15, MTTL82, MC23, MAH00, MQW09, MJEG21, MPL02, MKC08, MMALRA01, MALI11, MXK^{+}19, MG09, MHCL15, MCTB11, McW91b, MRSS^{+}18, MCT08, MMdOE^{+}22, MSHL22, MFL11, MHYN23, MM10, MMK04, MRWL23, MMS15, MNI^{+}16, MLP01, MGH22, MBC^{+}23, MBV18, MVK^{+}22, MKM19, NP96]. based [NAS18, NUM24, NZL^{+}21, NFHS06, NRTT95, OT88, OSB07, OSSJ09, OK12, OPR92, PMZS97, PXH^{+}03, PP16, PPM18, PF97, PA07, PHLW15, PQCT23, PGL^{+}23, PLVT23, PACSG^{+}23, PPD22, PWV^{+}18, PTY^{+}16, PSM16, Pra99, PCPW15, Pum96, QYC^{+}22, QKS01, QL23, Rei98, RKG22, RSAF18, RM05, RNM^{+}19, RMG15, ROP11, Ros90, RFR02, RdMF08, SP04, SKKN10, STT^{+}18, SM22, Sch98, SD21, SM98, SD08, SPT18, SM92, SA17, SDD95, SJZW07, SWH^{+}17, SXL^{+}23, SBE20, ST02, SNS06a, Shi04, SG15, SLQ^{+}19, SVW23, SB84, SvL09, SJZ^{+}23, SPL14, Ste09, SF92, SZL^{+}13, SGZ10, SXW^{+}22, TBS^{+}23, TSK98, TNF14, TDR^{+}17, TMN^{+}00, TWBP03, TSD^{+}07, TF18, TH90, Tap06, TEC^{+}14, TKD16, TN02, TFF^{+}20, TPB08, TBDC20, TZvD^{+}21, TBM^{+}04, TMB^{+}05, TNFG14, TB18, TC00, UKL^{+}13, Van89b, VRV05, WBK98, WS12, WWY06, WW08]. based [WZ09, WHL^{+}09, WLDB11, WY11b, WWL^{+}12, WTM12, WLP^{+}14, WLZL18, WWH^{+}21, WKT21, WSK^{+}22, WSHY22, WRLZ23, WPL^{+}23, WXL^{+}23, WXC^{+}23, WLL^{+}24, WLQC18, WLJT19, WI00, WZC^{+}24, WM24, XD08, XZ00, XZY^{+}21, XL10, XSQ^{+}97, XWW^{+}18, XLL^{+}20, XXX^{+}23, XZY^{+}23, YRS^{+}18, YC10, YLYJ13, YWC22, YPSZ01, YHNC22, YHX10, YH21, ZTAP21, ZHC11, ZjLW^{+}14, ZY02, ZK07, ZO07, ZSW08, ZSL08, ZS14, ZLM^{+}15, ZBP^{+}18, ZZJ^{+}14, ZHP^{+}19, ZZCL14, ZGC^{+}19, ZSS20, ZYW23, ZGZS22, ZM07, ZXD^{+}14, ZMK18, ZZDZ10, ZSS^{+}18, ZCL^{+}22, ZYML23, ZR12, ZZD^{+}19, ZPL^{+}15, ZCZ^{+}18, ZHG^{+}21, ZP04, dAU14, dAPG18, dVTT18, vBT20a, vTCB^{+}21, vWS04, vdBB07, vRESH16, HZ18, SBHC22]. baselines [HWR^{+}23]. bases [GLDK95, Nan77, PDS21, She88]. Basic [Cas88, BL96, HB91, SHL^{+}24, Van10, ZR97, FD75]. basics [HGS23]. basins [QM96, SP95]. basis [ABMC^{+}15, BAD23, BC01, ET18, Fra86, KJS17, LZ88, Sar04]. Batch [WWL16]. Batlló [GPC^{+}17]. Batrachion [Pic95a]. battlefield [KH03]. Baustein [SKH83]. Bayesian [CLN^{+}16, KNC11, LZY^{+}23, WZL^{+}22, YIL09]. be [HH88, SS75b]. Beam [RA15a, GH98]. bearing [QMHH91]. beautification [FASS16]. beauty [Pic87a, Pic87b]. Beauty3DFaceNet [XWW^{+}21]. beginners [Hol03]. beginnings [Nol95]. behavior [CTJ^{+}14, Coo01a, GS01a, MGM^{+}23, Wal93b, WXZ^{+}18, YMYI11]. behavioral [CF96]. behaviors [GKT02]. behaviour [Coo00b, Ent89b, PLM^{+}05]. behavioural [MHK99, PHO^{+}23]. Behaviourally [GD04]. behaviours [TB18]. being [S^{+}01]. believability [LUMC04]. Beltrami [RBG^{+}09, XLHH21]. Bench [NSL16]. benchmark [CGG^{+}20, LLC^{+}22, MR17, NSL16, RRQ^{+}22, VBS^{+}15, YK23, vOMRI^{+}15]. Benchmarking [TMB^{+}05]. Bending [PJP23]. benefit [Cot75a]. Best [Ano99c, Ano01-36, Ano02t, Ano06m, Ano08b, Ano09c, Ano10a, Ano11a, Ano13a, Ano21j, Ano22k, CR07, DZ93, EDKS96, FH94, FT02, GLDK97, LWP02, MLPB02, NA02, PMPR^{+}16, Ros92, Shi02, Sou94, YQY90, ZS02, Ano03k, Ano04m, MS09b, PB23, SK13]. Best-effort [PMPR^{+}16]. beta [LH91, LL91]. beta-spline [LH91]. Beta-splines [LL91]. Better [Per02, ZSS^{+}18, DK97, KRP^{+}91]. Between [Reh85, Bro92, CCKW11, CZ98, CSM^{+}01, DMM23, DGA02, GBF14, Gro91, HAL^{+}21, HC95, IA91, JS08, KYKK19, KN88, MMS89, MR05, ME92, MAF19, Qui91, Sri02, SCT^{+}14, SK13, YY88]. bevel [SCH^{+}18]. Beyond [Ano12a, CH12, MCKS06, BES01, CP97, FWD21]. Bézier [CM15b, SR15, ASC17, ASC18, BS93, BFRA11, CCKW11, CYM16, DM00, LKL^{+}02b, MR90a, Pet18, SY23, SV18, SM07, ZZ15, GdMF03, ZFS03, ZFSY04]. BFGS [FRWW14]. Bi [YR96, ASC17, ASC18, BYQZ22, CV97, Pet18, YAKE23]. bi-axial [YAKE23]. bi-cubic [ASC17, ASC18, Pet18]. Bi-directional [YR96, BYQZ22]. bi-level [CV97]. Bianca [WF22]. biarc [SXSY14]. bias [KAVM23]. biased [Sar92b]. biases [MD99a]. bibliography [San06]. bicubic [Com01, MZ89]. bicycle [COSEV22]. Bidirectional [LZZ^{+}19, HE15, JC95, JL23, MMK04, STT^{+}18, OKBG08]. bifurcations [Dal00, FM00, Wal06]. big [GSF^{+}19, ITW^{+}20, KS15, TPRC18]. biharmonic [FS20]. Bijective [WHFL16]. bike [OST^{+}16]. bike-sharing [OST^{+}16]. bilateral [HZ23, YMZ^{+}15]. Bildschirmtext [FP87]. Binary [GWL^{+}20, Wir80, AF00, AS95, BK89b, LP93, LCD15, MLPB02, OM96, QB92, SJT20, SK06, WKT21, vBT21]. bincode [TC00]. bincode-based [TC00]. bincodes [HC95]. binding [GRF^{+}22, GD11]. Binet [NR07b]. binning [HGW^{+}24]. binocular [BBE14, RZY^{+}20]. binomial [ME92]. Bio [CLF^{+}06]. Bio-edutainment [CLF^{+}06]. biochemical [EBST14]. biographic [DPS10]. Biographies [Ano15a]. Biography [Ano02u]. Biological [CHMR78, CH96, GNL^{+}15, JH89]. biologically [SEKA19]. Biology [KPBR20, KKN^{+}21, PB11, SPK19]. biomechanic [LPV92]. Biomechanical [SH94]. biomedical [MGJ^{+}11]. biomolecular [DSJ19b]. Biomorphic [Ley02, Stu91]. biomorphs [OdlCA02]. biopsy [FSS^{+}02]. biplane [FR92a, K^{+}00a]. biquadratic [KP15]. bisection [GIZ95, Ran91]. Bit [Dur89, Fiu87, DMG20, MDJ^{+}95]. Bit-mapped [Fiu87]. bit-rate [DMG20]. bivariate [CLWQ09, ZDT07]. black [BD98, CC20, GR09, Lee01]. blades [RJKV12]. Blaschke [KKS93]. Blaubeuren [SIG02]. blended [WZT97]. Blending [LY15, OKT01, SSBT01, ACO01, BSPD10, BPKB15, Bru14, CCCS08, CW03, DGKK20, FS20, KP15, LW99, MS01, PP16, Pie89, SZ09, YCZ04, ZP07, ZQ12, ZCL^{+}22, dGWvdW09]. blends [GBF14, ZGC15]. Blind [LGZ^{+}21, JZY^{+}23, WLDB11]. blinkers [PR96]. blinks [PQ10]. blob [LJWcH07]. Blobtrees [dGWvdW09]. Block [ACG15, DS93, KJ08, PHLW15, SGS99, TTKA23]. block-based [PHLW15]. blocks [PR96]. blood [ASS^{+}19, KGK^{+}07, MVPL18, RHBS95, WCA^{+}11]. blood-coil [WCA^{+}11]. Bloodline [KGGP19]. Blowing [KSH^{+}19]. Blue [XHGL12, NSG05, ZGZ^{+}16]. blue-c [NSG05]. blue-noise [ZGZ^{+}16]. blur [BBHC15, RAF21]. blurred [DRFRD06, SK06]. Board [Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17h, Ano18e, Bro06a, GN94, GHFH08, SH96, You89, ZW88, Ano85d, Ano86f, Ano94k, Ano95t, Ano02-30, Ano02-31, Ano02-32, Ano03v, Ano03w, Ano03u, Ano03r, Ano03s, Ano03t, Ano04z, Ano04x, Ano04y, Ano04u, Ano04v, Ano04w, Ano05u, Ano05v, Ano05t, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i]. Board [Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano18a, Ano18b, Ano18c, Ano18d, Ano18f, Ano18g, Ano18h, Ano18i, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24d, Ano24e]. Board/Pub [Ano04x]. Board/Publication [Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17h, Ano03v, Ano03w, Ano03u, Ano04z, Ano04y, Ano05u, Ano05v, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c]. Boards [Per77, PGR83]. bodies [GW06, PF97]. Body [Wee21, BMdSVR18, dPCOO^{+}05, CSX^{+}19, CCM^{+}18, CS06, DGC^{+}21, FEVM10, GG14, HSR^{+}09, KG04, MQW09, MPS06, OH83, RZY^{+}20, WAM17, WWF^{+}23, XLL^{+}18, ZLLL21]. Bogota [OFP^{+}11]. boiling [PCPW15]. BOLD [ZLL^{+}20]. BOLD3D [ZLL^{+}20]. bones [DJH^{+}23, SH94]. book [MGSC^{+}10]. books [LTS96]. Boolean [FR88, MT88, PYD^{+}05, RF00, Wal89]. Booleans [XCZ^{+}16]. bootstrap [WLQC18]. borders [BGK04]. boring [AD94]. botanical [Car92, COM^{+}94]. bottle [Sac22]. bounce [SSH16]. bound [CGB13]. boundaries [BGPT18, CIPT15, GSSK^{+}13, MWDG13, PGVACN06, RKMP13, SK23, ZLZG12]. Boundary [Lin15, MAC19, SOC^{+}19, WBK98, BCC20, CCC97, FST97, HHL99, HL06, IMG22, KM14, LKL02a, LYX18, LST96, MG98, NKMI23, PMTK01, Sob89, SJ15, WFC14, Wan23, Wu89, XCZ^{+}16, Yin04, Zhu91]. Bounded [FS20, CFZC19, NS87, VL98]. Bounding [MCT08, SVCNM23, CC03, LAM06, LYXY19, Mar03, WG17]. Bounding-planes [MCT08]. bounds [BS98, CR99, SOG08, ZFSY04]. bowling [PXH^{+}03]. box [FEVM10, LYXY19]. box-represented [LYXY19]. Boxes [SVCNM23, CC03]. boxing [SWH^{+}17]. brain [ABG^{+}18, DBLC02, GSF^{+}19, GWW^{+}22, GSB^{+}21, H^{+}00b, McW91a, PRW^{+}22, SL18]. brain-computer [SL18]. BrainTrawler [GSF^{+}19]. branched [MD99b, ZZJ^{+}14]. branches [YYY19]. Branching [CZ22]. Brazil [Ano11m, Ber84, Oli08, SK12, Wu02]. Brazilian [MTB18, Aug84, Tes84]. BRDC [MLPB02]. BRDF [CPLB14, MPSB21]. BRDFs [JJPP^{+}22, dSNJA22]. bread [BPD15]. Breadth [HW89]. Breadth-first [HW89]. Break [Bar77]. Break-Up [Bar77]. breaking [BDM^{+}16, VJ06]. breast [GPTP10, RNM^{+}19]. Breguet [For84]. BRep [GT91]. bridges [SWS75]. brief [Bow95b]. brighten [WLY23]. BrightFormer [WLY23]. Bringing [BZYB21, OC00, Wol00]. Brisbane [Hvi86b]. broadcasting [BS04, BK02, LRMS92, Yam94]. broken [ÜT99]. Browser [CZR22]. browsing [HW89, LDM^{+}11]. brush [KRK^{+}06, SEMWC05, WI00]. brushing [HEK22]. BTF [KCL18, RPKLMG23]. buckling [dSASP^{+}15]. Bud [CZ98]. Bud-sequence [CZ98]. budget [VLD15]. Buffer [Bas77, CGW07, SR95, YS97, CHSD95, rKC93, YD00, ZP92, MCTB14]. buffers [CNC^{+}21]. Build [MBN21, FN08, Seg88, VHS12]. Builder [DCJH13]. builders [Gre96]. Building [AMC03, HYP^{+}20, SVCNM23, YSW^{+}96, ZI00, BKL^{+}95, HHK^{+}13, HHKF10, LNSW16, LBV14, LUB^{+}13, OVWK16, ÖOK23, TS95]. buildings [BBP13, DVF06, DKM^{+}20, FB12, JPCS18, MP19]. Built [MPS85]. bump [BERW97, CL06, EJRW96, KLP01, Zha98a]. bump-mapping [BERW97]. bunch [HJT^{+}13]. bundle [SKP99]. bundles [VBW12, VBTW13]. bundling [XZY^{+}23, ZZD^{+}19]. Buried [WFS^{+}82]. burning [MR92]. Bus [GP86, GL83]. business [Fra83, Mil75]. butterfly [LLC09]. BVH [BHH15, VHS12]. BVHs [WIP08].

c [NSG05, GSY94]. CG [IB22, HKS00, Ise21, RSK^{+}24]. C2F [JXW^{+}22]. C2F-3DToothSeg [JXW^{+}22]. C2SPoint [JDT^{+}23]. CA [Ano94-32, Ano94-31]. CA-techniques [Ano94-32, Ano94-31]. CAAD [Ano95j]. Cabbage [CWTL08]. cache [RLD^{+}12, TPRC18]. cache-oblivious [TPRC18]. caches [WGS99]. caching [FDA03, JK21]. CAD [Ano86d, Ano95k, Ano95l, Ano95m, Ano03i, TMH20, AC77, Ano95k, Ano95l, Ano95m, Ano01h, Ano03i, Ano21i, AAB92, Aug84, Bär90, Big86, Boa78, CR13, CDGA84, DZZ79, Eas85, Ela86, Fen88, FIC21, FIC23, FRC06, For84, Fra86, FT02, Goh84, GA83a, GA83b, GS83, GS87, HZ15, HLCF88, IKTS22, IA83, JZLP23, JT86, KN88, KKNT88, LQOW08, LWP02, MHW10, MDM^{+}21, MKDM22, MLX18, Mar76, Med86, MPOL96, Mes84, MBGK89, MLPB02, MKKM18, MKM19, NA02, OT88, OH83, PM84a, QGW08, Reh85, RRQ^{+}22, Sal85, Shi02, Smi78, SAK90, Tes84, TS87, WLL^{+}12, WR79, XLXG11, YC10, YY88, Yue86, ZSW08, ZPP^{+}23, ZS02]. CAD*I [BG88]. CAD-based [HZ15, OT88]. CAD-directed [LQOW08]. CAD-kinematics [Med86]. CAD-oriented [FRC06]. CAD-systems [Big86]. CAD-turnkey-systems [GA83a]. CAD/CAE [HLCF88]. CAD/CAM [CR13, CDGA84, For84, Fra86, JT86, OH83, Sal85]. CAD/Graphic [QGW08]. CAD/Graphics [TMH20, Ano95m, Ano95k, Ano95l, Ano21i, MLX18, Ano01h]. CADD [PP99]. CADDM’94 [Ano94t]. ‘CADSketchNet’ [MDM^{+}21]. CAe [Ano12s, HLCF88, MW83, NKA83]. Cage [CHSB10, CLM^{+}19, LLX^{+}15, RGdL^{+}18]. Cage-based [CHSB10]. CageR [LLX^{+}15]. Cahiers [CV77]. CAL [Lau77, Mad94, Mad95, PPL91, Smi78]. CAL/CGI [PPL91]. Calculating [Stü98]. Calculation [Lam87, Hor82, VBVS88, XLHH21]. calculation-minimized [Hor82]. calculations [AF16a]. calculator [MDJ^{+}95]. calibration [HT96, KSH17, LRG11, VKA^{+}23]. Call [Ano85a, Ano85b, Ano85c, Ano86b, Ano86c, Ano86d, Ano93c, Ano93a, Ano93e, Ano94d, Ano94e, Ano94h, Ano94f, Ano94g, Ano95q, Ano95r, Ano95s, Ano95z, Ano95-27, Ano95x, Ano95y, Ano95w, Ano95a, Ano95b, Ano95c, Ano95-41, Ano95-39, Ano95-40, Ano96a, Ano01-31, Ano01-32, Ano03j, Ano03z, Ano04g, Ano05m, Ano06f, Ano09a, Ano09b, Enc83d, Enc84b, Enc85a, Enc85b, Enc86a, Enc86b, Ano77, Ano86h, Ano94p, Ano01i]. calligraphic [IWM94]. calligraphy [WI00, YP05, ZCL^{+}18, SEMWC05]. Callosum [CJAR21]. CAM [Bjo85, CR13, CDGA84, For84, Fra86, JT86, MW83, OH83, Sal85, YY88]. CAM-X [MW83]. Camera [TOY^{+}14, AG13, CACC24, HT96, HZ18, KSF15, LRG11, MPQG18, RCB15, SSM11a, TSY11]. cameras [MBPF12, PSBD19, YK23]. camerawork [MY16]. campimetry [dMOHKO18]. Can [Bij85, YRD^{+}24]. cancer [ZBM00]. Canonical [ZLZG12]. Canopy [LCWZ14]. Canopy-frame [LCWZ14]. Cantor [PM90]. canvas [OCR^{+}19]. cao [YP05]. cap [Sac22]. capabilities [KAV^{+}88]. Capacity [XLGG11, ZGZ^{+}16, TT12, ZYW^{+}21]. Capacity-Constrained [XLGG11]. CAPD [SZL^{+}13]. CAPE [Ano94p]. capture [AOB17, CFB15, SSM87, You89, dBWK18]. captured [HQW14]. captures [PRBD22]. Capturing [SM07, ARM23, JY98, MS08, SR02, SWZZ23]. capturing-based [JY98]. car [JRJP^{+}22, KK22]. cardiac [ASS^{+}19, CCM^{+}11, KGGP19, MVPL18, PSK^{+}11]. cardinality [CCH94, HCC91]. cardiology [B^{+}00]. cardiovascular [SAKB75]. cards [TMB^{+}05]. care [GLL00]. caricature [LVM^{+}11]. Carlo [BSB^{+}02, JC95, PTW98, SB86, Shi92, WLZL18]. Carotid [Coo00b, Coo01b]. Carotid-Kundalini [Coo00b]. carpets [YS23]. Carrying [BF07]. cars [SGC^{+}19]. CARTE [WA75]. cartilage [LSCJ23]. Cartographic [Mir87, JH11]. Cartography [Haz79b, Sch85b]. Cartoonish [ZHP^{+}19]. cartoonization [WLL^{+}21a]. carved [AJAC23]. Casa [GPC^{+}17]. cascade [WSHY22, WRLZ23]. cascaded [KM21]. case [DQF04, GPC^{+}17, HWSW19, HGH^{+}18, MTS^{+}22, PVR87, PMBS14, SV06, SS82, SKM98, TB18]. case-based [TB18]. case-study [DQF04]. cases [KYKK19]. Cast [HN85]. castable [SJG19]. casting [AA07, BCG^{+}94, CC19, CSSC00, DGBNV18, GMP22, rKC93, WBK98, WF11, WGS99]. Catalonia [ACB12, Mir87]. Catalunya [BJAN^{+}95]. CATE [Ano04h]. categories [VHR^{+}18]. Categorizing [WFC^{+}09]. catheter [CCM^{+}11, LCDN06]. catheter-heart [CCM^{+}11]. Catmull [LAE^{+}19, NA02]. cats [Cra02]. cauliflowers [RPÁM04]. Causes [Bra95]. caustic [KM21]. caustics [KM21, WZ21]. caves [FM22]. Cayley [MR96]. CDA [HK93]. Cech [DI12]. CEIG [AL10, CJO19, OOC22, SPO22, SCG23]. Celebrating [Ano24c]. cell [CPG94, PVM^{+}22, PTL04, PM91, WBK98, XH88]. cell-based [WBK98]. Cellular [Ada97, RR92, ATAG^{+}21, BAD23, BTV83, DLR^{+}10, Fra94, Li89, Mas92, McC02, NR07a, Rei10, WHH21]. center [BMT96, Web87, AW93]. centered [BSAH^{+}23, BPKG07, FGES96, FEVM10, GTdS^{+}18]. centerline [LTR^{+}14]. central [And82]. centralised [How88]. Centralized [Dai93]. centre [Gob93a, GS93, Hvi86b]. centric [KPH^{+}05, WSJJ24, dF24]. centrifugal [KS20]. centroidal [RA15b, ZYW^{+}21]. century [Wol00]. Ceramic [ZLZ^{+}20, CGH97a, CGH97b]. ceramics [XZY^{+}21]. cerebral [MVPL18]. cerebrovascular [ZZJ^{+}14]. certain [GD87]. Certificate [Ano01v, Ano01-44, Ano01-45, Ano02s, Ano04-37, Ano04-32, Ano05z, Ano01u, Ano02-44, Ano02-45, Ano03-33, Ano04-36, Ano06-32]. Certification [Pfa83, Bro84, Tho84]. Certification/validation [Pfa83]. Certifying [BMP84]. CFD [Lar03]. CG [FT02, LWP02, MLPB02, NA02, Shi02, ZS02, Ano86e, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano04-59, PZH^{+}05]. cGAN [RM22]. CGE04 [Ano04i]. CGEMS [Ano06k, Ano07l, Ano07m, Ano05o, Ano05q, Ano05n, Ano06j, Ano06l, Ano06g]. CGI [Ano04j, Ano05p, CEM89, FF96, HTKRW88, PPL91]. CGI’05 [Ano05b]. CGI2002 [Ano02v]. CGIM [Ano03d, Ano04l, Ano04k]. CGV [TAS09]. CGV-An [TAS09]. chain [Bri03, PS02]. Chained [MHCL15]. chains [ASR^{+}22, FR98, GdMF03]. challenge [LPL^{+}20]. Challenges [Ano01-38, Gin93, SB97, dSJ23]. change [GYK^{+}23, KBL22, KGB^{+}21, LSR22, TOY^{+}14]. changes [BBHC15, CHSD95, HS00, LCCS04]. Channel [LHH^{+}21, CS04, HLO16, HWR^{+}23, JZY^{+}23, LXT^{+}23, QSLS23, RBFS10, VHON04]. channels [RFB23]. Chaos [Ano99a, Ano99b, BGR94, Bap99, CGR98, CZC02, CCW99, Coo02, Dra98, Fat02, FM99, Gin02, Har00, Jef92, Kru99b, McC02, MR90b, Nik98, OdlCA02, Rei02, Sen98, Sen99, VB99, Wal93a, BPS06, Bra95, CD93, DBLC02, HP91, LM89, OK02, VA96, Mar02b, Mar02a]. Chaotic [Coo00b, Coo01a, DHJ^{+}99, JR00, PRÁM04, RS08, Rei97b, Rei02, CGR98, Cha97, CC01a, Coo07, Kei89, PBN97, QV95, RC94]. chapters [Mad95]. character [CBU^{+}15, HHG97, JRZ^{+}23, MHLB16, RRC^{+}18, SZ95, SCSG18, ZBP^{+}18, dAU14]. characteristics [BHZ^{+}21, LJCW04, MAdS^{+}19]. characterization [MTS^{+}22, PK86, RLB^{+}02]. characterize [SMU22]. Characterizing [CA17, EK15, SKSI95]. Characters [GWW84, dPCOO^{+}05, CYJ^{+}13, FvdPT01, GD04, MZ23, MKHN01, PQ10, SFC01, VVCN12, ZMYH06]. chart [MVS14, ZZCY22]. Chebyshev [Pat15]. check [AG13]. Checking [PMK85]. cheese [PM90]. chemical [IE98, RFB^{+}21, RW87]. Chernoff [SZW10]. CHI [Ano01d, Ano06h, HGH^{+}18]. Chief [Enc83a, Enc93, Jor18c, Jor18d, Jor18e, Jor18f, Jor18g, Jor19d, Jor19e, Jor19f, Jor19g, Jor19h, Jor20e, Jor20f, Jor20g, Jor20h, Jor21e, Jor21f, Jor21g, Jor21h, Jor23c, Jor23d]. Children [SB77]. China [Ano12o]. Chinagraph [Liu21]. Chinese [CWTL08, GWW84, IWM94, JXY87, PMZS97, SEMWC05, WI00, YP05, ZCL^{+}18, ZM07]. chip [Ack96]. Choice [Ano23k, Kei89, Ye02]. choreograph [MO92]. choreographies [RDD^{+}18]. chromatic [HKPL98]. Chromodynamics [TBLH17]. CHuMI [JGA09]. CIM [Sal85, Sch85a]. circle [Car99b, FC00b]. Circles [Pic89, And94, PK91]. Circuit [Per77, CS88, PGR83]. Circular [BM08, FS88, TMSPB09, CGH94, WHH06]. circulation [AÇÖ96]. cities [RPP21]. City [RV01, CSM^{+}01, GYK^{+}23, PSP^{+}20, SV06]. CIVR [Ano07h, Ano07i]. CLAP [RR92]. Clark [LAE^{+}19, NA02]. class [ALC06, Ano04-56, BHZ^{+}21, HLCF88, MSHL22, SCFF16]. class- [ALC06]. class-aware [MSHL22]. Classes [Ran88, GD87, San06, ZXLF23]. Classification [Egh83, GCvdS^{+}20, HGH^{+}18, RNM^{+}19, Sch86b, ATAG^{+}21, BRdSOS17, CVL^{+}04, CR13, DWZ^{+}22, DHZL20, FGP^{+}10, GA12, GK96, GS11, GTFB23, HHN^{+}23, JDT^{+}23, KS01a, KGK^{+}07, OPR92, PPD22, RFB^{+}21, SPT18, SPL14, UPT97, WCLT21, WLZ^{+}23, YLH^{+}18, ZK07, dAPG18]. classification-based [SPL14, ZK07]. classification-to-saliency [JDT^{+}23]. classifications [McD06]. Classified [ZCZ^{+}18]. Classifier [HR07, IR06]. classifiers [RSB^{+}19]. classify [Cas96, KJ01]. Classifying [TPK13, LSGFRC^{+}13, WLW05]. classroom [BM03, ZWWC23]. clay [MDS^{+}21, MDS^{+}22]. cleaning [PPM18]. Clearance [OP15]. clicks [JXW^{+}22]. client [JHL^{+}12]. climate [KBL22, LSR22]. clinical [MSO^{+}20]. CLIP [WZC^{+}24]. CLIP-based [WZC^{+}24]. Clipmaps [DR15]. Clipping [CB78, DGR93, AS91, Aro89, CM15a, Day92, Dor90, KJ08, LL12, LWP02, MA15, SM92, SEC90, Ska93, Ska94, Ska96, Ska97, SPY87, TKOD24, WHH06]. Clodion [DSM^{+}99]. cloning [HLO16]. Closed [Cor82, Tar22, PK91, SMM20, VM15, WA02]. Closed-form [Tar22]. closest [GO06]. closings [ZQ12]. Cloth [ZY01, BD13, BWD13, DG07, FRC06, HSE10, KL07, LDG96, Lin97, LKC98, LSWL13, NGA95, ZSS20]. Clothed [CSL23]. clothing [II22, LLL^{+}23]. clothoid [HEWF13, MS09a, WM05]. Clothoidal [WM90]. cloud [BGLA18, Bou20, CXGL23, CZCG04, DWZ^{+}22, DFF22, ERWS12, GYK^{+}23, HDZR23, HG21, HZLQ20, HHN^{+}23, JDT^{+}23, KZ04, KS15, KVB^{+}20, KGB^{+}21, LCZ^{+}11, LZC^{+}15, MTSM10, MBST22, MRWL23, NOS15, NZL^{+}21, PECW22, QL23, RB20, RRQ^{+}22, SBY11, SPS12b, SWL^{+}16, TZT^{+}22, TWNL22, VBS^{+}15, VZP22, WCLT21, WZL^{+}22, WPL^{+}23, WLZ^{+}23, WCH^{+}24, WLX^{+}21, WM24, ZCL^{+}13, ZLL^{+}21, ZZH^{+}24, ZZC^{+}20b, ZMH^{+}23, ZDL^{+}24, vOMRI^{+}15]. clouds [AAK22a, AK21, BLW12, DFWW15, DG17, GYY^{+}23, GM04, GMd^{+}13, HLZ^{+}17, JWZ23, LSK^{+}10, LZG^{+}23, LHCL05, LJJPF22, LCXL20, OVWK16, ÖT21, RMP19, RBF17, RRB^{+}22, SMU22, SDT21, ST22, SSV07, SW11, SMK08, SHS^{+}23, TGG06, TCL15b, TKB07, VDOK19, WBB^{+}08, WSL^{+}19, WYC^{+}23, YZWJ24, ZGW^{+}16, ZDL22, ZWL^{+}22, ZZH^{+}23]. CloudWalker [MBST22]. cluster [PMBS14]. Clustered [AP22, GMP22]. clustering [BA09, BAC14, CMS22, MSE17, MKM19, NMM09, RVR04, SFS^{+}21, SZW10, TTKA23, WWS^{+}13, XLXG11, ZSW08, ZCL^{+}13, ZLG^{+}15]. clustering-based [BAC14]. clusters [TBM^{+}04]. clutter [HHLE17, vBT20b]. cluttered [MMV^{+}14, ZML^{+}18]. CMA [ZHW^{+}21]. CNN [BWZ^{+}18, US20, ZLLG18]. CNNs [CWL20, LYZ^{+}22]. CNSBS [WZT97]. Co [Ano15a, SJB^{+}21, WCY^{+}20, Ano05y, KES22, Loh95, PPV03, WXG^{+}14, WWS^{+}13, WSWL14]. Co-adaptive [SJB^{+}21]. Co-Authors [Ano15a]. co-located [KES22]. co-matting [WXG^{+}14]. co-occurrence-based [Loh95]. co-operation [PPV03]. co-segmentation [WWS^{+}13, WSWL14]. Co-skeletons [WCY^{+}20]. co-sponsored [Ano05y]. coaching [HGH^{+}18]. Coarse [JXW^{+}22, TZT^{+}22, JCFN18]. Coarse-to-fine [JXW^{+}22, TZT^{+}22]. coated [HKYM01]. Cocone [DDW11]. code [CVL^{+}04, MLPB02]. coded [TVS^{+}03, VW21]. coder [KNC11]. codes [SLGQ23, TH90, YPLL19]. Coding [NM85, BC88, Bri03, IWM94, KNC11, WW08, WVY16]. coevolution [CWGR01]. cognition [PBG^{+}14]. Cognitive [HWEB22, PS15b, MKHN01, PRM^{+}24, WJG^{+}19]. cognitive-state [PRM^{+}24]. Coherence [FT02, ZWS19, CMLR11, ST97, Wee21]. Coherence-sensitive [FT02]. coherency [GLC20]. Coherent [HHG97, ERWS12, GY19, HJW97, SCVCN16, TPN95, WBA16, YY14]. coil [WCA^{+}11]. Coimbra [TS94]. coin [Coo07]. Coincidence [CM15b, CYM16, SR15]. cold [Yue86]. Collaborating [S^{+}01]. collaboration [Bro00, MTS^{+}22, MST^{+}22, MWY^{+}10]. Collaborative [MML^{+}11, A^{+}01, Bou09, CSJ^{+}21, CSS^{+}24, CSF20, CGS^{+}96, CJJ99, DR09, DGC^{+}21, ET18, Elv96, FHFG99, H^{+}01, KS03a, KAFB18, MTS^{+}22, ÖOK23, RdCAM01, RSP^{+}19, TS95]. collage [GZSZ20]. collection [WZZZ18]. collections [BTS19, Le 77, LGWM17, ME17, PSMD14, PE16, TMP07, WCH^{+}11]. collective [BSC^{+}21, D^{+}01]. Collision [JS08, UOT83, BH91, CK09, GO06, GKLM07, JTT01, Kam93, LAM06, LKC98, MAF19, MO06, WTWT18]. Colloque [Cho77, Cib77]. colon [ZMKG11]. colonic [MMD^{+}20]. colonization [dLBRM^{+}12]. colonoscopy [MBC^{+}23]. Color [Ehl85, LLM^{+}16, McC84, Mic90, NYKN83, Tan80, AVHT17, CCC^{+}16, CCC00, CMDS17, CV97, Dix91, ER87, FO21, HGW^{+}24, HLO16, JX96, LS05, LZYQ22, MCMV22, MPAC^{+}23, PR11, RZF19, RCPB01, RNM^{+}19, SL12a, Shi93d, SW19, SSM11b, Smo03, TB19, TS75, VW21, WBRV16, XHL^{+}24, YL96, Zha96]. Color-aware [LLM^{+}16]. color-channel [HLO16]. color-coded [VW21]. Color-mapping [Tan80]. colorations [CP10]. colorectal [RNM^{+}19]. Coloring [WC89, McC02]. colorization [AVHT17, BF19]. colormap [WVY16]. Colour [Har83, KS84, RO87, CLN^{+}16, CCW99, Fle91, Fri03, LHC12, Nik98, RHS^{+}94, Tay87]. coloured [IA83]. colours [HEK22]. ColourVis [LHC12]. columns [DK97]. Combination [Van10, HR07, Xu08]. combinations [ZSH12]. combinatorial [CBG22, Wu89]. Combined [CBNJ^{+}15, FJS11, SGC00, GLZ^{+}21, HXH24, RWE05]. Combining [CSL23, DAHF04, DCLB19, GPC^{+}17, JXY87, MCT05, WCHM22, HEW^{+}18, KB15, LZ24, LTP19, ZPN^{+}21]. CoMEdiA [San93]. COMETT [BGK89]. command [SVV92]. Comment [CM15b, SR15]. Comments [Woo89]. commercial [AAB92, CS04]. commodity [BND^{+}17, MBPF12, NBE^{+}04, TBM^{+}04]. commodity-based [TBM^{+}04]. Common [CS80a, MHW10, Wan23]. communicate [MGS^{+}22]. Communicating [KA86, Vää93, FGES96, GCW23]. Communication [Ano94a, BNS24, Egl86a, SB77, SBOT78, GN94, GS83, HMdM^{+}95, OC21, SHOC23, SH96, SBD^{+}94]. communications [Cot75b, SM75]. community [LPP^{+}19, Tho84]. Comp [Ano04k]. Compact [LMY^{+}21, OK20, BSC^{+}21, CGB13, FWD21, KWK17, TKS11, YC10]. Compaction [BS82]. company [Goh84]. Comparative [MLCMGR23, AAB18, FRTT18, GSB^{+}21, HAL^{+}21, LSR22, MFL11, SC97]. compare [CMLR11]. Comparing [CPC^{+}18, WFC^{+}09, vTTK^{+}20, EHM84, HP03, vdLdFvdEV23]. Comparison [KCK17, QN98, Str85b, VK07, BFRA11, CMS98, HGW^{+}24, MLP19, SMS09a, SMS09b, UL22, vOHR20, JP10, YRD^{+}24]. compass [MP22]. compatibility [GANM21]. compatible [LZLS18]. compendium [WKO12]. compensating [LS08]. compensation [MPTA^{+}22]. competitive [RWD14]. competitiveness [ZXH^{+}12]. complement [RCG^{+}05]. complementarity [ANE17]. complementary [BWYZ24, VW21]. Complete [WA02, FDGM18]. completing [KKJ^{+}23]. completion [ALR23, AAB18, FJS11, HBM23, LYW^{+}11, MCK12, SJZW07, WM24, XCL^{+}19, ZZY12]. Complex [BX99, Li89, PK85, She93, SHLW89, SK16, AS95, AJAC23, Car96, CCL^{+}03, CGWW16, CXT18, CF99, Coo01b, Coo02, DI12, DCV98, DMG99, DKM^{+}20, Eng93, Ent89a, ES22, FHM98, GB91, GBV92, IFD15, JC95, JXM^{+}10, JS08, MC10, MMV^{+}14, PMTK01, Pic87b, RCM^{+}20, RPAM00, STM^{+}04, SKSI95, SS13, Szy90a, TSK98, TEC^{+}14, TB18, XPL90, Zom10]. complexes [CDI12, DGLRD18, FWD21, GLA23, PTL04]. complexity [AHR^{+}22, CM06, DG06, MRG^{+}19, MSMK19, RHS^{+}94, SBWS11, Shi92, Ska96]. Complicated [Cro80]. Component [LQ12, MLP01, DGA02, LXT^{+}23, LBB11, ZR97]. Component-aware [LQ12]. Component-based [MLP01, DGA02]. Components [PS86, LLW^{+}19, NGAS23, Pha95, PSR^{+}93, Pra99, RdCAM01, SvD03]. Composing [WWK11]. composite [PSR^{+}93, Pra99, SCVCN16, Shi93a]. Compositing [LMHRG10, BRV^{+}10, PP16, WX14]. composition [HW16, JWL12, PMZS97, SIPO17, ZP07]. Compound [DW05, Fra86]. Comprehending [PSMD14]. comprehension [RM05]. comprehensive [KD15, ZTAP21]. compressed [SMK08]. Compression [KG04, MGH13, VGP04, AF00, BKV05, BRSP15, BW98, CFB15, CCH94, CK93, CV97, ER87, GK96, HL97, HE15, HMHB08, KPA12, KCL18, Kra10, LJJPF22, MSE17, MCAH12, MPL02, Nah23, PA07, QL23, SPS12b, TMP07, ZO07]. compression-domain [BRSP15]. Compressive [GCW23, JGA09, JYLW14]. Comput [Cad08, KRA^{+}23a, MDS^{+}22, PM22a, SMS09a]. Computation [Ano95-29, Ano95-30, CCKW11, PAFL06, AF11, AZF13, BKR^{+}16, BK20, CPM19, CPLB14, DL09, GL83, GR93, KYKK19, Kar92, KLKE11, KMGL99, LCK16, MWDG13, Pat15, PM93, PPSS96, ST97, Ska08, SBS22, Sri02, SCT^{+}14, TyZfTM12, ZGC^{+}19, ZYW^{+}21]. Computational [AGM^{+}21, Ano85g, Ano86j, Ano12s, ABM^{+}06, CM18, KS20, LZY^{+}23, Mal87, MD99b, PJ21, PMPR^{+}16, SBD15b, ZWSW22, BT94, Gho93, Gro92, MWDG13, MCP^{+}22, Par93, SM22, SC97, CIW09]. computations [CC19, VBW12]. Computer [AK06, AD85, All77, AW93, AL10, Ano78, Ano83a, Ano85c, Ano93d, Ano94i, Ano94j, Ano94p, Ano95v, Ano95-29, Ano95-30, Ano01-38, Ano01-66, Ano03h, Ano03a, Ano05o, Ano05q, Ano06j, Ano06l, Ano06i, Ano07f, Ano07-28, Ano07-29, Ano07g, AL78, AN99, AA01, BWF18, Bax77, BSPR77, Ber84, Ber77, BGT16, BPS06, Bjo85, BG80, BSW78, BJAN^{+}95, BGK89, Car84, CE80, CGH97a, CGH97b, CHMR78, Com85, Coo79, CG93, Cor76, Cou92b, Cro80, CS85, Dan78, DW82, DSN75, Del80, DSM^{+}99, DIE78, Dro78, DS15, DB83, Ebe00, Egl86a, EK85, EF15, End85, Ent77, FF96, FS80, GPS18, GS84, Gom90, GSME04, Gra85, GA88, GHM^{+}96, GC86, HBJ82, Han95, HKPL98, Her83a, HO88, HRGD88, How79, Hun77, Hvi86a, IS05, IKM90, JM88, Jv95, JZR88, JMV90]. Computer [KA85, KAV^{+}88, Kit77, KKS86, Lam87, LD11, LS79, LD78, LS84, LPS19, LG89, Mac77, MTT84, Mag85, MD99a, Man99, Mar78, Mar79, ME83, MD99b, McC95, McW88b, Mei83, MHK99, Mit77, Mud99, MNSJ99, NYKN83, Neg77, Oli08, Paq05, Per84, Pic84, PK85, PCS00, Pos77a, Pot77a, Pot75, PP99, PSR^{+}93, Pra99, Rad96, Rag79, Rie78, Rob78, RW87, Rog85, Row86, San06, Sch88, SPK19, SGR^{+}99, SKO83, SGS99, SAKB75, SMG77, Smi77, Sou93, SD16, SO75, Str85b, Sug83, Szy89a, Szy90a, Tak77, Tan94, TS94, Tru86, Var92, VP77, VBVS88, VCHR07, Wal93b, Wal94b, WFS^{+}82, WF88, Wis87, Wu02, dos01, ASM12, Ano91a, Ano91b, Ano01-65, Ano04j, Ano04l, Ano04i, Ano04-59, Ano05b, Ano06-47, Ano12o, AMC03, Ban85, Ben79, BF02b, Bou02]. computer [BEKL00, BL82, Bru75, CGR87, CVHM03, CC05, CLT07, CK75, Cor84, Cot75a, Cun00, DKM^{+}20, EW75, Eck90, FEJ04, FVG15, FP75, GMM18, Gri88, Gro91, GS89, H^{+}00a, HS05, Han97a, Han97b, HGS23, HS00, HH88, HLS89, HvK87, HL02, IC87, JK90, JXY87, KD94a, KS98, KAFB18, Kel86, KF02, KS04, KY97, KB04, Kor87, LS18, LC23, Mar76, Mar86, MDSU88, McW87, McW90, McW91a, McW91b, MG08, MMV88, Mit87, MY16, MHM95, MY97, MBP14, MTB18, NT95, NK01, Nol95, Nov03, Nug91, OC00, OSZ00, PGB86, PGR83, Pau88, Per02, Pic92b, Pot77b, Pra92, PPV03, Qui91, Roj91, Rok93a, RWE05, She12, Shi06, SL12a, SL18, Smi93, SS75a, Spr75, SF91, Sul85, SS04b, TSD87, Tan80, Tax04, Tei96, TMK94, TMH20, Van94, Wal93a]. computer [Web87, Wol00, Wol02a, Wri75, ZP04, Ano01-34, Ano01j, Ano01-33, Ano01-35, Ano02v, Ano04m, Ano06k, Ano07l, Ano07m, Ano07n, Ano07o, Mil92a, Mil92b]. Computer-Aided [Dro78, GS84, Str85b, Coo79, HKPL98, JZR88, KAV^{+}88, VBVS88, WF88, Cor84, JK90, KAFB18, MMV88, PGB86, Tan80, TMH20]. Computer-Based [Bax77, Str85b, Eck90, McW91b]. Computer-Drawn [Pic84]. computer-generated [JM88]. computer-supported [Pau88]. Computer/Human [Ano07o]. Computerized [Ano77, MTC02, Dur89, Dur91, MSO^{+}20]. Computers [Ano96a, Ano97c, Ano03k, Ano05m, Ano08i, Bij85, CS80a, Coc79, Cum89, FB15a, GO85, Gra83, NKNN83, WGS^{+}18c, YS21, Ano03d, Ano04h, Ano24a, BBDM03, Hol03, Jor23f, KKŽ04, NBE^{+}04, Nij04, SVT86, WFC^{+}09, Ano85a, Ano86b, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97d, Ano99c, Ano01-36, Ano06m, Ano08b, Ano09c, Ano10a, Ano11a, Ano11b, Ano11c, Ano13a, Ano13b, Ano13c, CMS12, DZ93, EDKS96, Enc85a, Enc85b, Enc86a, Enc86b, GG96, GLDK97, Ros92, SOdSC18, Sou94, YQY90, vdPS08, Ano23a, Ano23j, Ano23l, Ano24f, Ano24g, BA23, Jor23e, SSS15]. Computing [Ano95-41, Ano95-39, Ano95-40, Ano01-50, Ano06-27, Ano06-38, ABTW77, Bay95, CDIM16, EF15, JS92, Kno87, KPBR20, KKN^{+}21, PF09, PDS21, PB11, SPK19, SK23, SJ94, Ste99, VP00, YCF18, Ano91d, Ano03-49, Ano04g, Ano04-58, Ano05-28, Bad96, BR96, CM15a, Elv96, FB14, FB15a, GRW00, GN04, HMdM^{+}95, KMS^{+}97, KS01b, KD03, KFH^{+}09, LMD96, LH83, MK89, MCS^{+}18, MZ89, MTM22, Rau06, RAA^{+}08, SG15, SH96, SMM20, VM15, YSD13, ZFSY04, dSJ23, Ano94e, Ano95-28, Ano02l]. COMRIS [dHT01]. concave [LL05, SV18]. concavity [JCFN18]. concavity-aware [JCFN18]. concealment [JPP01]. concentrated [RS08]. Concept [CH91, DLV06, GGW22, HRTK86, Mac85, Mac78, MUS83, Ano03-57, Aon90, EL04, EMB^{+}98, FF96, IB06, JRS21, JCFN18, KCS22, LSW^{+}23, MSMP12]. Concepts [Ano02f, Pic84, Tak77, Ano04d, Cas88, Che06, GGW22, HB91]. Conceptual [Eas85, Haz77, BCF06, IWM^{+}09, OCR^{+}19]. Conceptualisation [San93]. concrete [GBF14, VR16]. Concurrent [Pos77a]. condition [CM15b, CYM16, SR15]. conditioned [CYWM23, WLB^{+}23]. Conditions [YJLZ21, BM08, LYX18, PPG^{+}18, Yin04, ZZ15]. conductivity [TNF14]. cone [WTWT18]. cones [MT00]. conf [Ano04k]. CONFERENCE [Ano07f, Ano85c, Ano94p, Ano94r, Ano95k, Ano95l, Ano95m, Ano95q, Ano95r, Ano95s, Ano95v, Ano95-29, Ano95-30, Ano95-41, Ano95-39, Ano95-40, Ano01-28, Ano01-37, Ano01c, Ano01-46, Ano01-47, Ano01-58, Ano02l, Ano02-41, Ano02-38, Ano02-55, Ano03h, Ano03a, Ano04n, Ano04b, Ano04c, Ano05c, Ano05-41, Ano06c, Ano06-27, Ano06a, Ano06b, Ano06-29, Ano06-30, Ano06-31, Ano06-38, Ano06-50, Ano06-51, Ano07h, Ano07i, Ano07p, Ano07-28, Ano07-29, Ano07g, Ano07c, Ano07e, Ano07d, Ano07-30, Ano08a, AMS22, BYC19, DS15, FS17, Joh18, SD16, Tan94, TSC16, Ano91a, Ano91b, Ano94c, Ano94q, Ano94t, Ano94u, Ano94v, Ano01w, Ano01-65, Ano02-53, Ano03i, Ano03b, Ano03d, Ano03c, Ano04a, Ano04l, Ano04h, Ano05b, Ano05d, Ano05a, Ano05-40, Ano06-35, Ano06-33, Ano06-34, IK21]. conference [MCPW21, TMH20, Ano01e, Ano93b, Ano01-59, Ano01-64, Ano02g, Ano02h, Ano03a, Ano06n, Ano06-37, CL92, MCPW21]. Confidence [WXG^{+}14, LZW^{+}21b]. Confidence-driven [WXG^{+}14]. configurable [BDRV01, Pfa84]. configurations [CLWQ09, TPM14]. configuring [LLM^{+}90]. confocal [RPFC01]. conformal [Bru75, LX08, PP12, ZLZG12]. Conformance [PMK85, Sco84, Pfa84]. conforming [FSP15, MAC19, XCZ^{+}16]. conformity [Kru84]. confronts [SIPO17]. Congress [Ano85g]. congruence [LP93]. congruent [LL92, YAKE23]. Conic [Pie84, ATB98, Xu08]. conics [SZ09]. conjecture [CZ98]. Conjugate [MSE20, KL07]. conjunction [Ano07-39]. connect [OMW13]. Connected [Pha95, Jas88, LZT^{+}24, MP89, PR93, QL23, Str86, ZYW23]. connections [FS88]. Connectivity [GCDL22, SK06, And98, MAS14, QB92]. Connectivity-guaranteed [GCDL22]. connoted [CVL^{+}04]. conquer [HWFQ09]. conscientiousness [ADHC^{+}23]. consecutive [MRG^{+}19]. Consensus [SS12a, VT06]. conservative [FWCS97]. conserving [IY18]. considerations [BG01, CEPS13, HEW^{+}18, HWEB22]. consistency [FBTT^{+}22, LZ24, WYC^{+}23]. Consistent [GF09, WCY^{+}20, ASSF17, IFD15, LZW^{+}21b, MCMV22, SK13]. console [JD75]. consolidation [HLL^{+}18]. constant [RPM97]. Constrained [Bou06a, CTP^{+}21, Coo78, CSC10, SVV23, XLGG11, AZF13, BC13, ENE11, HHCM17, JLP00, LYL^{+}17, LSWL13, LLL^{+}15, MYF06, PHO^{+}23, Sah15, TWBP03, VP00, ZGZ^{+}16, ZZC^{+}14, ZYW^{+}21, BW98]. constraining [KRRS12, OSB07]. Constraint [JYL24, Mul01, YRS^{+}18, DZD^{+}23, Des00, Doh95, JAS97, JALS03, MMF03, MT00, TMN^{+}00, vdBB07]. Constraint-based [YRS^{+}18, vdBB07]. Constraints [HO94, AB03, BMB95, CGWW16, DLZY14, GDA^{+}13, NA02, PF09, PX06, RKMP13, RCLM18, SAK90, ZYX^{+}19]. Constructing [HWR^{+}23, Pre84a, Ang97, Lam00, SVV92, ZCT95]. Construction [AKW^{+}16, CCI12, DM79, GN89, KR92, MWY^{+}10, WG17, APS09, Ano04p, BHH15, BKL17, CFH^{+}18, CF77, DI12, DWH09, FST97, Fiu89, HWFQ09, KW14, KS09, LWW08, LLLZ16, LKHM19, LST96, LP83, LYS^{+}16, Mal87, ME17, NG05, PM90, SXG^{+}09, SA87, SNS06b, SKM98, TSK98, VH15, WIP08, YCF18, Zom10, ZJH87, ZWP^{+}93]. constructions [Har00, MP22, MGOH96]. Constructive [Con91, JALS03]. constructively [Tax04]. consumer [FRTT18]. consumption [JSMK14]. Contact [BWBM20, WLB^{+}23, ASR^{+}22, ES22, FM17, KK16, SKSI95, TPM14, VKA^{+}23]. Contact-conditioned [WLB^{+}23]. contagion [BGD18]. containing [Pie84]. Content [FH11a, FH11b, GLS^{+}20, KPL08, WZ09, BKM16, FB04, FJ03, MWA^{+}13, MVRB18, PdSP^{+}22, SFD06, ZPN^{+}21]. Content-aware [GLS^{+}20, WZ09]. Content-based [KPL08, FJ03]. content-style [ZPN^{+}21]. Contents [Ano95-42, Ano02w, Ano04-63, Ano80, Ano83c, Ano85e, Ano03-61, BWF18]. Contents/Author [Ano02w, Ano03-61].

Context [BTS19, Mac85, MGJ^{+}11, RMSC11, AAB18, CMD99, CSM^{+}01, DMV06, FST97, HZLQ20, JHPhR11, LKLW16, LLS^{+}16, MA15, RGE07, SWF^{+}20, SBG99, ZZL21, ZWWC23, dMTB^{+}21]. Context-adaptive [BTS19]. Context-aware [MGJ^{+}11, RMSC11, ZZL21]. context-dependent [DMV06]. context-free [FST97]. contexts [WL15]. Contextual [TKOD24, LRY^{+}24, LTP19, SLX^{+}16]. contextualized [MST^{+}22]. Contextually [BSB^{+}23]. continuation [Gom14]. continuity [YCZ04]. Continuous [Bou20, FBH^{+}01, LBLV16, BKCW14, CS06, GKLM07, JRSM17, KP09, PC23, RC08, RHC15, SZ09, Wil03, ZQ12, CAS^{+}15]. Continuum [Ano06c, GMMP21, JXM^{+}10]. Contour [NLSN11, RTB^{+}18, SSK87, BG91, CN05, Con91, FJW11, LSZQ21, LSW15, ML12, RMG15, SZ95, SHLW89, TC00, ZW88]. Contour-aware [RTB^{+}18]. contour-based [CN05, RMG15]. Contour-driven [NLSN11]. Contouring [Chr78]. contours [BKL15, BKL17, CT13, JX96, ZNA08, ZJH87]. contracting [KP18]. contraction [CHL^{+}11, DS18]. contrast [AEA13, ANGH11, LYX18, LPR^{+}14, WZC^{+}21, ZZJ18]. contrast-enhanced [ANGH11]. contrast-preserving [ZZJ18]. Contrastive [MJEG21, WCF^{+}23]. Contributions [Ano93a, Ano86b, Ano94d, Ano94e, Enc85a, Enc85b, Enc86a, Enc86b]. Control [AK13, Dai93, GS01a, Lea85, AYA^{+}20, ABCD93, BSPD10, BW94, BMT96, BDM^{+}16, CACC24, CVB16, DAG22, DS93, FMP96, GLS^{+}20, IP03, JPCS18, KFW16, Lam00, LLLC11, LS08, NC12, PF89, QV95, QM96, Rag80, RCB15, RJS98, Sar94a, SP00b, Shi04, SSM11a, SHBSS16, TDR01, YLYJ13, YMZ^{+}15, ZGC15, ZEK^{+}17, ZDT07, ZZC20a].

Controllable [ZCL^{+}22, HY23, LLX^{+}15]. Controlled [ML79, CF11, DN22, GD04, HEWF13, HY03, SFVP13, WM05]. controller [Mil03, Sta87a]. controllers [Par86]. controls [LL91]. convective [TKB07]. conventions [BES01]. Convergence [XLM12, Coo00a, Pat15]. converging [TMN^{+}00]. Conversation [WS22, D^{+}01, WF22]. Conversion [GP86, BTV83, Eas75, Lin79, MAC19, MvSE18, Shl83, WZT97, WJD^{+}09, YR98, ZEK^{+}17, ZYML23]. Convertible [ZC18]. Converting [BMW12, BW92, Kau88]. Convex [QKS01, KLKE11, LWW07, Mar09, MFOK94, Ska93, Ska97, Sri02, SGES12, TyZfTM12, TKZ^{+}13, ZM91]. Convexity [ZDT07, BB91]. Convolution [ZZH^{+}23, Fou11, HXH24, JYC^{+}23, JZY^{+}23, JT02, KYM12, LCGD24, NLdAL^{+}23, SHZ19, ZLL^{+}21, ZLL^{+}23a, ZDC^{+}23, ZWWC23, ZCL^{+}22]. Convolutional [JK21, BRPC18, FCG^{+}21, HZLC22, LCXL20, OBD^{+}23, SPT18, WGS^{+}18c, WGS^{+}18a, WGS^{+}18b, ZYW23]. convolutions [Bou20, FM09, PR23]. ConvPoint [Bou20]. CONVR [Ano04p, Ano04o]. Coons [Pie88, SNS06b, WZT97]. Coons-type [Pie88]. cooperation [Ano04-58, CSM^{+}01, SIG02]. cooperative [DYW^{+}22, GA07, San93, AHK03]. Coordinate [Ber77, IA91, MG86, EPS96, KK08, MMS89, MRG^{+}19, RM91, Suf88]. Coordinate-Free [Ber77]. Coordinated [WL15]. coordinates [Aro89, BPKB15, Ska08, SPS12a, SHCW22, SZL^{+}13, ZNGN16]. Coordination [NHR^{+}22, RdCAM01, CWGR01]. copy [BBP13]. Core [End83, GK04, LYC^{+}15, NIH08, SW11, TPB08, VP06, GN80]. core/shell [LYC^{+}15]. Corner [FS98, MS07, XL10, YHHS93]. corners [IA91]. Cornsweet [LPR^{+}14]. corotational [BD13, SZW^{+}14]. Corpus [CJAR21]. correct [BFSE03, CGMS00, JAS97, Koh96b, MA94]. Correction [DWL^{+}03, BHTT94, CIT^{+}19, KWK23, MKDM22, RZF19, SBE20]. correctness [CEPS13, Mar02a]. Correlation [GS84]. correlations [ZMH^{+}23]. correspondence [ATZM19, DLR^{+}20, FLV20, RGdL^{+}18, WYZ20, YYG16, ZZCY22]. correspondences [SK13]. Corrigendum [PM22a, SOdSC18, WGS^{+}18c, YS21]. corrupted [LH14]. COSAR [BA90]. Cosine [MPA^{+}10]. cosmetic [Bre01]. Cosserat [MBC^{+}23]. cost [Cot75a, DDQM98, KKNT88, Mey79, SM22, WG17]. cost-benefit [Cot75a]. count [vBT20b]. coupled [Coo01a]. coupling [MHYN23, TRLX22, WWCZ19]. Course [Mar78, Mit77, AM10, Ban97b, BHL^{+}94, Cun00, Pie90, RWE05, San00, San06, Wol00, Wol02a]. courses [BWF18, Bou02, KS04, Paq05, SSB^{+}08]. courseware [HS03]. coursework [SL16a]. Covariance [LQ15, PQCT23]. Covariance-Matching [PQCT23]. cover [MYL^{+}23, NAS18]. coverable [HXA^{+}12]. coverage [Bri03]. covering [AXG^{+}13, SIE14]. coverings [RdMF08]. COVID [AAK^{+}22b, ASR^{+}22]. COVID-19 [AAK^{+}22b, ASR^{+}22]. COVIR [AAK^{+}22b]. CPU [WJW^{+}23]. CQUATS [Gin02]. CR [DZD^{+}23]. CR-Net [DZD^{+}23]. crack [TRB^{+}22, WAM17]. cracks [YYY19]. Craniofacial [JZX^{+}21, DZD^{+}23, HZD^{+}19]. Crash [WP77]. Cray [GS89]. create [JCFN18, dSNJA22, Sch95]. created [CLT07]. Creating [Ban85, BBMK21, CCM^{+}07, JVS^{+}24, SSH16, ZSS20, BF19, BGD18, LaV07, MUH10]. Creation [GWW84, dSEM19, van89a, BSAH^{+}23, CS80b, JCT^{+}15, MD99a, Pic95b, SGBI02, Van09, YAKE23]. Creative [Tru86, Sel89]. Creativity [KPMT18]. creatures [HP01, TPM14]. Criteria [De 84, Pos77b, DKM^{+}20]. criterion [GTG06, LKL^{+}02b, TDR^{+}17]. Critical [FLV20, FR92b, LLC04, MTS^{+}22, MW14, TWNL22]. critically [Hol03]. critiques [Whi04]. crop [JWL12]. crop-and-warp [JWL12]. Cross [CMdL21, CWL20, LSS21, SDT21, SCFF16, Woo87, ZHW^{+}21, BKL17, HGJ16, KD11, KW14, PT16, Rei95, Roc89, SA17, VHR^{+}18, ZR97]. Cross-class [SCFF16]. Cross-domain [CWL20, LSS21]. Cross-modal [ZHW^{+}21]. cross-parameterization [KW14]. Cross-sectional [Woo87]. cross-sections [HGJ16, KD11, SA17]. cross-tabulating [VHR^{+}18]. Cross-time [SDT21]. Crossing [CM23, WJG^{+}19, EVRW23, WBJ^{+}21]. Crossing-free [CM23]. crossings [JYL17]. Crossroads [Bro92]. crosswalk [BRdSOS17]. Crowd [HYP^{+}20, BKM16, DSG21, G^{+}01, HBOS13, ICNV21, JXM^{+}10, LD12, OAYG10, RPP21, WLL^{+}24, ZZC^{+}14, vTP20, vTCB^{+}21]. crowded [GPC^{+}17]. crowds [BGD18, dLBRM^{+}12, KKPC23, WLJT19, vTCB^{+}21]. crowdsourcing [BRdSOS17]. crying [Pic95a]. cryo [GCvdS^{+}20]. cryo-electron [GCvdS^{+}20]. CryptoComparator [CEN^{+}23]. cryptocurrencies [CEN^{+}23]. cryptography [BD98]. Crystal [Pot77b, SA86]. crystallization [NR07a]. crystallographic [Ste75]. CSCW [Ano06o, Ano06p]. CSG [AJAC23, Ano95q, Ano95r, Ano95s, CCC97, Ger86, GT91, JV91, MS90, PVCM22, RVdF08, VCQ92]. CSG/BRep [GT91]. CSRBF [MBV18]. CSRBFs [KHS03]. CT [APA^{+}11, AAK^{+}22b, ST22, YLH^{+}18]. CT-scan [AAK^{+}22b]. CT-volumes [ST22]. CTA [KGK^{+}07]. CU [QL23]. cube [PWK95, KMS^{+}97]. Cube-4 [KMS^{+}97]. Cubes [CEPS13, KONS17, KAÖ98, NBE^{+}04, NY06]. Cubic [BB93b, Sar94a, ASC17, ASC18, BB91, CY94, DBG92, Dra98, DDPT98, FEVM10, LP83, LJH18, Mon87, Pet18, PS91, Sar92b, Sar92a, Sar94b, SAMA97, Sar00, SBH01, Sar02, Sar03, WMRA^{+}15, XCL^{+}19]. cubical [Mok88]. cubics [Sar93, SM07]. cubism [ABCO12]. cuboid [GZLW14]. cuboid-structured [GZLW14]. CUDA [YMYI11]. CudaHull [SGES12]. cues [WB24]. Culling [SBS04, BMH99, FJW11, GSF99, HIK05, RCBS10, YCL^{+}21]. Cultural [BL11, CMS11, ASM12, MD99a, MGMB22, SLL^{+}21, CMS12]. Cultures [Mar78]. cumulative [Mou13]. cumulus [DG17]. curls [Pic87b]. Current [CMSF11, Rie78, TMK94]. currents [XSQ^{+}97]. curricula [Ebe00, Han95, HS00, Paq05]. Curricular [Wol02b]. Curriculum [Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03l, Dom93, MC23, MWY^{+}10, TS94, Ano03m, Ano03n]. Curvature [HEWF13, SY23, ZR12, DTG15, GA12, HN20, HBM23, KP09, KKL02, LJH18, MSMK19, PMV06, TF18, Tsu17]. curvature-adaptive [TF18]. Curvature-based [ZR12]. Curvature-controlled [HEWF13]. Curve [AMT86, AB03, Fay85, Haz79a, HAL20, HY93b, IORM17, LL12, Moh77, NKNN83, SHBSS17, dGGDV11, BMU^{+}16, CJXZ23, Chu90, CY94, CZ06, CZCG04, GSY94, HEWF13, HY03, Kar92, Koh96b, LGRP14, LL04, LMW06, LZG^{+}23, LH91, LP83, LYS^{+}16, MPR89, MM10, MZ89, NA02, OK12, PMM18, Pha89, PS91, Pie83, SHBSS16, SWL^{+}16, WY11a, WYXM22, WA02, WCY^{+}20, Wüt98, Xu08, Zay12, ZJSB22, Zha98b, ZZY12]. curve-like [ZJSB22]. curve-polygon [LL04]. curve-skeletons [BMU^{+}16]. curved [AA00, BP94, BFSE03, Boe91, CT13, KB05, LMZ90, MSMK19, RdMF08, RF97, SM22, SGGC05, Wee21, Zhu91, dCdLL14]. curved-edge [RF97]. curved-space [Wee21]. Curves [ML79, Pie84, TY24, AS19, ALC06, ASWL11, AFW^{+}18, BGR94, Bri03, Bro08b, Cha90, CCKW11, CM15b, CYM16, DM00, DTG15, DW13, Elb22, Ger97, Gom14, GLT^{+}97, HEWF13, HZD^{+}19, IIK12, JOK^{+}07, JRSM17, Jut94, KLKE11, LLLC11, LT95, LOdF02, MW12, Mal87, MS07, MS09a, Mes00, Mok87, Mok88, NC12, dCNPdFS14, OSB07, PR97, PDS21, PSS04, RBF20, SY23, SR15, Sar93, Sar94a, SXSY14, SCT^{+}14, SMM20, VM15, WM89, WWF^{+}18, ZLS98, ZLS99, ZM92, ZFS03, ZFSY04]. curves/surfaces [CCKW11]. curvilinear [DTWT94, GZZS06, OK99, TMSPB09, YY96]. Custom [Jer95, KMS^{+}97]. customer [SSW^{+}23]. customization [LK18]. Customized [SF92]. cut [CF11, CFH^{+}18, JVS^{+}24, YL96]. cutaway [KMV^{+}18]. cuts [GDDA13]. Cutting [CP97, BS01b, Cho06, Cum89, MKKM18, WJW^{+}23]. CW [Ano04r, Ano04q]. cybersecurity [BRO22]. cybersickness [MVRB18]. CyberTape [Wan04]. Cyberworlds [Ano03o, Ano03a, Ano04r, Ano04s, FS17, Ano05a, IK21, Ano05r, Ano06a, Ano07c, Ano08a, IK21, Joh18, Kiy19]. cycle [AXG^{+}13, CWL20, MVPL18]. cycles [CM20]. Cyclic [ACC^{+}11, JR00, LL06, Pic87a, Rei10, ZLLY06]. cyclide [GBF14]. cylinder [BW92, CAL^{+}04, Roc89]. cylinders [FB11, FR92a, SH94, TCL15b, vEB98]. cylindrical [RGRG15].

D [Ano03-62, Ano05-41, Ano07-39, BF02a, BTD^{+}22, BPKG07, CB97, GVVJ99, LPD^{+}18, LSCJ23, Maj98, PCD^{+}15, RV01, STP^{+}20, TT19, ZPL^{+}15, dlLC99, AT08, AMHWW16, AK21, ASS22, ATAG^{+}21, AA07, ASS^{+}19, Ano03f, Ano05d, Ano13o, ATZM19, AA13, ABCO12, AG94, ALM19, ABM^{+}06, BM03, BK89a, BND^{+}17, BCS^{+}99, BDRV01, BBCG11, BY88, BWdBP13, BAPD23, BCC^{+}22, BSF13, BTS19, BFLP20, BDL^{+}22, BYQZ22, BPKG07, BN03, BGLA18, Bou09, BJS01, BG91, BD97, BS01b, BK89b, BHL^{+}15, BKL15, BAC14, BD17, CMLH21, CCCS08, CKS98, CTP^{+}21, CPC^{+}18, CGG^{+}20, CLH^{+}16, Car96, CMB17, CMSF11, CVB16, CW03, CC08, CSG^{+}17, CSX^{+}19, CYWM23, CCM^{+}18, CSK97, CBNJ^{+}15, CGH97a, CB10, CH12, CHSD95, CMDS17, CS18, DN22, DGV^{+}24, DCJH13, DEST95, DCLB19, DMS08, DG01, DSR11, Dur89, Dur91, DKY97]. D [DGR93, Elb22, EK22, EKP93, ET07, EBC^{+}15, ERB^{+}14, EME15, FAZ21, FH11a, FH11b, FDGM18, FHM98, FS98, FJW11, FWX^{+}18, FGZ^{+}22, FSS^{+}02, FCSB90, Fer01, FIC21, FIC23, FR92a, Fou11, FM22, FFP^{+}21, GM86, GWX^{+}18, GPTB02, GV89, GD95, Gin02, GF09, GMd^{+}13, Gom85, GA12, GZZS06, GvK18, GLC20, GBD88, GN89, GLA23, GYL^{+}13, HZ15, HLY^{+}19, HKHP11, HG21, HWSW19, HKBA17, HR88, HHCM17, HL93, HHKF10, HR07, HCLC16, HHN^{+}23, HYP^{+}24, HS99, IR06, IO91, JH89, JK15, JXW^{+}22, JTT01, JHPhR11, JJL97, JA84b, JRS21, JCFN18, dSJdML18, JRJP^{+}22, KS98, Kan85, KSM07, Kas87, Kau88, K^{+}00a, KD11, KRK^{+}06, KCK17, KP95, Kle86, KGGP19, Kor90, KYT^{+}17, KVB^{+}20, KGB^{+}21, KLL^{+}15, KEVD18, KD15, KCS22, KCS24, KYM12, Laf94, LDS^{+}21, LHS87, LBD17, LNL^{+}23b, LNL^{+}23a, LMC13, LK18, LDLD22]. D [LKL^{+}20, LBLV16, LVM^{+}11, LYW^{+}10, LLLC11, LLLZ16, LYL^{+}17, LZZ^{+}19, LWW^{+}20, LCL^{+}21, LSZQ21, LWWY22, LG94, LZT^{+}24, Lin97, LJWcH07, LXB^{+}15, LZL^{+}15, LM16, LXCW18, LWD^{+}18, LXJL21, LRY^{+}24, LSS21, LSGFRC^{+}13, LUB^{+}13, LYS^{+}16, LCXL20, LB19, LR90, LBTM15, LAE^{+}19, MG86, MAFL16, MHLB16, MDM^{+}21, MKDM22, MOS^{+}21, MLC^{+}22, Mar79, MG09, MAG^{+}12, Mar09, Mar10, MWA^{+}13, MS08, MMM^{+}20, MHYN23, MBST22, MFP11, MCKS06, MCP^{+}22, MY16, MSRB17, MS09b, NG88, NP96, NC12, NTAI20, NP88, NAK13, OK99, OK03, OP13, OCR^{+}19, ÖT21, OC21, PPS20, PSBD19, PRRR13, PK91, PS12, PS13b, PA07, PRW^{+}22, Pic88a, PPD22, PECW22, PPG^{+}18, PVCM22, PTY^{+}16, PCWD23, PPVT03, QYC^{+}22, RTB^{+}18, RJS01, RBB^{+}11, RAK^{+}15, RB06, RO13, RE22, RdCVL16, RMG15, RBF20, RPM97, RHBS95, RCG^{+}05, SMU22, SSB04, ST20, SDT21, ST22, SG92, Sak02, SBWS11]. D [SVNB99, SS22, ST97, SGC00, SM99, SJT20, SF98, SD15, SPT18, SA17, SLL^{+}17, SLX^{+}16, SKO83, Sin87, SBS13, SET^{+}88, Sla92, SJ09, SJZ^{+}23, SVP82, SHBSS17, SGES12, SK06, SS75b, SCFF16, SK13, SHS^{+}23, TBS^{+}23, TZT^{+}22, TRB^{+}22, TMK94, THL15, US20, ÜT99, VGP04, VZP22, VAGT08, VDOK19, WBRV16, WBA16, WS12, WBB^{+}08, WTM12, WZZZ18, WGS^{+}18a, WGS^{+}18b, WLYH19, WWH^{+}21, WRLZ23, WPB^{+}23, WUH^{+}15, WH96, WBL^{+}97, WR02, Wol02b, WWS^{+}13, WYZ20, WYC^{+}23, WLG04, XCL^{+}19, XWW^{+}21, XLL^{+}18, XWWK21, YF09, YZ17, YPLL19, YSZ22, YLT^{+}18, YNS94, YHNC22, YHX10, ZTAP21, ZIP^{+}19, ZLLL21, ZPP^{+}23, ZLLG18, ZLZ^{+}20, ZM07, ZMK18, ZLL^{+}20, ZDL22, ZZDZ10, ZSM^{+}19, ZHG^{+}21, dGGDV11, dlLC99, vBT20a, vBT20b, vBT21]. D-look [BKL15]. D-models [CKS98]. D-printing [EME15, LXJL21]. D-XCT [RAK^{+}15]. D/PHIGS [HR88, Sin87]. D/S3D [MY16]. DAGs [MTS21]. dairy [NFW^{+}24]. Dam [Mil92b]. damaged [LYW^{+}11]. d’aménagement [Dom77]. Dance [BSW78, FCG^{+}21, RDD^{+}18]. dans [CV77, Nan77]. Darmstadt [Ano02-55, KKMT06, EF15, Web87]. dart [Fat01]. dart-shaped [Fat01]. dashboards [BDRV01]. Dassault [For84]. Dassault-Breguet [For84]. Data [Ano93e, Ano01-58, Ano02g, Ano02h, Ano03-56, Ano04-60, Ano06n, Ano06-51, ARL78, BT78, BRO22, BKS23, C^{+}01, CS85, DJC^{+}23, EFP02, ER87, FZPM93, GM78, GB75, HLY^{+}19, KH00, MMGB17, Mit77, NHR^{+}22, RB20, Reh85, San85, Sch85b, SLX^{+}16, SBS13, TC93, WHW^{+}22, Wec79, AM91, AMM^{+}07, AKW^{+}16, ASS^{+}19, ANGH11, Ano01-28, Ano01-41, Ano03-48, Ano03-55, Ano04n, AKB22, AÇÖ96, BS98, BGD18, BFT23, BKM16, BRdSOS17, BK02, BPGW11, BLS15, BC88, BMB95, BSC^{+}21, BKS21, CJAR21, CCCS08, CSFG96, CTS^{+}10, CTJ^{+}14, CC01a, CR07, CHZ^{+}23, CFB15, CK02b, CZCG04, DH95a, DTWT94, DDM^{+}06, DGV^{+}24, DAG22, DGKK20, DKLP02, DH95b, E^{+}00, EDKS94, Fra86, Fru94, Gal88, GSF^{+}19, GWW^{+}22, Gia15, GPTP10, GSV^{+}18, GK96, GBKG04, GS11, GWBD17, HSD96, HWYL21, HW22, HKCL02, HN20, HGW^{+}24]. data [Her83b, HHLE17, HJDR95, HG02, HPD^{+}10, HS08, HY93a, IE98, IL97, ITW18, JSMK14, KPSN04, KCK17, KONS17, KF88, KGGP19, KHS03, KS15, Lar03, LZZ^{+}19, LLC04, LYS^{+}19, LYZ^{+}22, LWZ^{+}23b, MA15, MAM^{+}24, MMD^{+}20, MK83, MSD75, MTSM10, MPL21, MGS^{+}22, MA14, MCP^{+}22, NKMI23, NC07, NIH08, NLS07, dATNMC^{+}22, OL96, PSMD14, PSBD19, PRRR13, PL97, PLJL15, Par75a, PS12, PSH^{+}09, PHLW15, PRW^{+}22, PMBS14, PR93, Pol83, PTR^{+}92, PRM^{+}24, QWC14, RBF17, RPFC01, RAK^{+}15, RHS^{+}94, RBLB21, RA03, ROP11, RNM95, SKKN10, SK03, SA04, SVNB99, SAMA97, Sar00, SBH01, Sar02, Sar03, SSK87, SGC00, SBD15a, SPS12b, SDWE99, SVVS^{+}17, SJB^{+}21, SL02, Spr04, SL16a, SHD^{+}17, SWL^{+}16, TW24, THQ^{+}16, TC24, TBLH17, TRB^{+}22, TZvD^{+}21, TPRC18, THL15, Tsu17, TT12, TKB07, VT07, VP06, WK14, WHH21]. data [WZW97, WFC14, WWH^{+}21, Wel76, WMW13, WMDR08, WWF^{+}18, XZ00, XWW^{+}19, XH88, YWR03, YC10, YWC22, YPZ^{+}23, YHW23, ZNGN16, ZLM^{+}15, ZQL15, ZKS^{+}96, dSB04, vLvKV11, vOHR20, vOMRI^{+}15]. Data-aware [SBS13]. data-comparison [vOHR20]. data-dependent [CTS^{+}10]. Data-Driven [NHR^{+}22, BRO22, DJC^{+}23, HLY^{+}19, IE98, RB20, SLX^{+}16, WHW^{+}22, BGD18, BKM16, ZLM^{+}15]. data-hiding [TT12]. Data-sensitive [MMGB17]. data-users-tasks [MA14]. Database [AN99, CLH^{+}16, How88, PR82a, TSY11, ZR97]. Databases [BS82, CZR22, SS93]. dataflow [tHd90]. dataset [BHZ^{+}21, CLH^{+}16, LYL^{+}17, MDM^{+}21, RFB23, WZZ^{+}21]. Datasets [Ger02, HKCL02, HG02, LFL02, TWSH02, BSM^{+}22, BBP10, CGW07, FV13, LBV14, LWG^{+}23, MBV18, UL22, WPJP23]. DATICAM [CS85]. datums [PVCM22]. DaType [Sch98]. Davis [MS82]. DCE [GPTP10]. DCE-MRI [GPTP10]. DCT [JPP01]. DCU [ZYW23]. DCU-NET [ZYW23]. DDA [Fay85]. de-occlusion [LZKJ23]. dead [SFC01]. deblurring [JZY^{+}23, LH14, LGZ^{+}21, YSD13, ZZF^{+}23, ZLL^{+}23b]. December [Ano02-55, Ano03a, Ano19i, Ano20i, Ano21k, Ano22l, Ano23p]. Decentralized [FS86b]. deciding [MFOK94]. Decimation [KRA^{+}23a, KRA^{+}23b, Rei98]. decision [ADHC^{+}23, CFMS02, FMCM^{+}21, LWZ^{+}23a, Mil75, MSO^{+}20, PSMD14, QL23, RHFL14, Sch95, SK23]. decision-making [ADHC^{+}23, LWZ^{+}23a, RHFL14, Sch95]. declarative [Des00, STdKB11, Woo90]. decogons [VB99]. decolorization [ZZJ18]. decomposing [HL93]. Decomposition [LLGA12, CDIM16, DH95a, DRFRD06, FCM^{+}18, FWD21, FVG15, HLB^{+}06, JYL17, KS91, KM14, LWW07, LHG21, LXT^{+}23, MPL02, MRF06, RdCVL16, TWNL22, WHL^{+}09, WFC14, XGZ11, ZW20, ZTS02, WHZ^{+}18, ZSL08]. decompression [NIH08]. deconstruction [VB99]. decoration [CGH97a, CGH97b]. decoupled [GM05, SLGQ23, WCL23]. dedicated [Ano94u, Ano94v]. Deep [ASS22, BAPD23, CAAC20, FIC21, HBM23, JK21, LCXL20, MDM^{+}21, MCV18, PJVH^{+}24, PRBD22, VNMP23, WZZZ18, WWCZ19, WLL^{+}21a, XWW^{+}21, dMTB^{+}21, BRdSOS17, BGLA18, BHL^{+}15, DJH^{+}23, FAZ21, GTdS^{+}18, HG21, HGJC21, LSS21, LB19, MENS19, MKDM22, MCMV22, QSXT22, RMP19, SS22, VZP22, WLYH19, WZL^{+}22, WRLZ23, WPB^{+}23, WPJP23, WZZ^{+}18, XZPG21, YHW23, ZTAP21, ZDC^{+}23, ZGZS22, ZZH^{+}23]. deeper [JK21]. DeepPortraitDrawing [WWF^{+}23]. Deferred [ZP07, FBTT^{+}22]. Defined [Van85, BSF13, FPC10, LTV08, MMS89, Ran88, RH85, SJ09, UBW99]. Defining [GKT02, Pie84, Roc89, WJD^{+}09, Mar76]. Definition [Haz79a, Mar76]. Deformable [CC04, CSHZ04, GÖT93, JSV98, LST96, Ano06-35, ADOR02, Cho06, DKY97, GRPR08, GO10, GW06, GKLM07, GLT^{+}97, GÖT97, IMG22, KK16, KÇM00, LD05, LKLW16, LBB11, LBB12, MO06, OMP^{+}18, PMTK01, RBB^{+}11, TNFG14, WJW^{+}23]. Deformation [CF13, RdAMA21, WMRA^{+}15, ASPO15, AB97, Bec94, CMdL21, CBU^{+}15, CHSB10, CSC10, FTB12, FMP96, FVG15, GLS^{+}20, HLB^{+}06, hLfTxDdZ09, LSWL13, LLW^{+}19, LX08, MYF06, RGdL^{+}18, RSAF18, RHC15, RNM95, Sah15, SZW^{+}14, VR16, WZ09, ZZC20a, ZQL15, ZHP^{+}19, vRESH16]. deformation-aware [ZQL15]. deformation-based [vRESH16]. deformations [CVB16, DLR^{+}20, GO90, JLP00, KR92, Par93]. deghosting [KHTM17, SL12b]. degradation [HXC^{+}23, VPLL06]. degraded [SWF^{+}20]. degree [BSJC02, Boa78, CM15b, CYM16, IORM17, SR15]. degrees [MS82]. degrees-of-freedom [MS82]. dehazing [LYX18, LYZ^{+}22, LZL^{+}24, XZL^{+}22].

Dehydration [LYC^{+}15]. Delaunay [Ang97, CLWQ09, KKŽ04, KB02, PPM18, SWL^{+}16, SEKA19, SMM20, VP00, XLGG11, ZMYH06]. delay [AHK03, CDGC94, Sen99, Sen98]. Delft [Ano01-62, Jv95]. Delivering [Ano01-46, Ano01-47, Ano02-55, Ano04c, TC93, Ano01w]. Della [Fal89]. Delone [RCG^{+}05]. Demand [RS99]. Demand-driven [RS99]. demands [PSSP96]. demonstration [Gob93a, GS93]. DEMs [AJAC23]. Denoising [ALM19, JYLW14, PLJ^{+}13, WPL^{+}23, XSW23, YRS^{+}18, YBD^{+}24, ZLGH10]. Dense [Tap06, ZLLL21, ABG^{+}18, CCCS08, DSJ19b, DJH^{+}23, RGdL^{+}18, RBF17]. densely [HZLC22, ZYW23]. densities [vTCB^{+}21]. Density [FCG^{+}07, AMFH21, CTQ^{+}14, DS93, FJW11, KFW16, SG92, XXHM21, ZHG^{+}21]. dental [SKSI95, WCLZ14]. dependent [BKV05, BGK04, BFSE03, CTS^{+}10, CC08, DMV06, ESAH02, FV13, GAÖ02, GCRR11, GNL^{+}15, LFL02, MALI11, MWS04, MVG^{+}21, RGGB02, TWSH02]. depicting [AT11]. depiction [BMS^{+}11, JK15, MFP11]. deposition [KEVD18]. Depth [CVP^{+}16, LSE18, LR16, ALR23, APS09, AAB18, BBE14, CSL18, HAL^{+}21, ISPS17, KP95, KWK23, KCS22, LFY^{+}21, LMJH^{+}11, MBPF12, MENS19, MSL^{+}19, OK20, PSBD19, PCKB23, Rok93a, SPGR93, Sri02, SXW^{+}22, WPB^{+}23, ZYW23]. depth-assisted [KWK23]. depth-based [LSE18, LMJH^{+}11]. depth-map [CSL18]. depth-merge [SPGR93]. Deraining [LHH^{+}21, LXT^{+}23]. Derivative [Haz77]. derivatives [ASWL11, HG02]. derived [Lak89, Szy90b, ABG^{+}18]. descent [NAO13, RYNJ23]. Description [Reh85, Sug83, BYQZ22, Dom77, Lan88, LBLV16, Mar91, Sob89, TBDC20, ZWP^{+}93]. Descriptions [HN20, RS75]. descriptive [Pau88]. descriptor [CMdL21, CZCG04, TCL^{+}15a, ZLL^{+}20, vBT20a, vBT20b]. Descriptors [ATAG^{+}21, BWdBP13, CDF14, DFF22, EHBA10, TB19, vBT21]. Design [Ade85, AP88, All77, Ano12b, Ano12p, BERW97, CMD99, CE80, CFMS02, Coc79, DW82, DIE78, Dro78, GH91, Her85, HEW^{+}18, HYP^{+}20, JK84, KS98, Lau77, LS79, LWZ^{+}23a, LLHH94, LQOW08, Mag85, ME77, Mar79, ME83, MGSC^{+}10, MMS^{+}17, Nar15, PKRM21, PWJ^{+}18, RK84, SHL^{+}24, Str85b, Tan94, VKA^{+}23, War76, Wit84, Yue86, ZW88, vOMRI^{+}15, Ade86, AGM^{+}21, AC90, ACB12, Ano03-49, ALD12, AAB92, AFM93, BDPR93, Ban97a, BLNZ22, BG01, BCF06, BHTT94, CS88, CLN^{+}16, Cho06, CPCS90, CK75, Com85, Cor84, DLV06, DSM^{+}99, DKM^{+}20, Eck90, EL04, ESFGDZ97, FB12, FB15b, For84, FP75, GA83b, GS83, GSY94, Hel95, HLCF88, IC87, IWM^{+}09, IKM90, JZR88, JMV90, KAFB18, KAV^{+}88, KS20, Kni95b, KD00, KB20, KMWW^{+}14, KAAO75, LZY^{+}23, LZ88, MHW10, MSE20]. design [MMP18, MBA20, McW91b, Men85, MA14, MMV88, MHM95, MN90, OCR^{+}19, OK12, OPR92, Par93, PKK03, PGB86, PGR83, PPV07, PAE^{+}21, dMPF08a, dMPF08b, PSR^{+}93, Pra99, QMHH91, RJKV12, RCLM18, Sch98, SVT86, Sen98, Sen99, Sen03, SBHC22, SS96, SWvB95, SK04, SVP82, Sta87a, SJG19, SS75b, Tan80, Thi85, TS95, TMH20, UWC90, VV89, Wei84a, WWD^{+}95, Woo87, WMW13, XH88, ZjLW^{+}14, ZY02, ZZL21, ZC18, GA88]. design-Results [IWM^{+}09]. Designed [ML79, PDL^{+}21]. Designer [KW96]. Designing [Hvi86c, KA22, LSE18, LVVC06, NA02, Sar93, SS02, TKB07, BM03, OMP^{+}18]. designs [Ger97]. desktop [ABG^{+}18, HAL^{+}21, Koh96a, RBW01, TS95]. DeskVR [AMR23]. detail [ACC^{+}18, BP10, DDM^{+}06, GNL^{+}15, IWT13, LYL^{+}17, LHH^{+}21, PACSG^{+}23, RS99, RC08, SP00b, WHZ^{+}18, ZZDZ10, ZYML23]. detail-guided [LHH^{+}21]. detail-preserving [LYL^{+}17]. detailed [LS08, RLS^{+}12]. details [BRPC18, CGG19, GY19, RHC15, RPP20, ZSL12]. Detect [WTW^{+}23]. Detecting [AKB22, DLZY14, PCV16, CA17, CYW15, TFY00, ZJSB22, ZM92]. Detection [KRA^{+}23a, LXCW18, SP00a, AVM05, BH91, CGG^{+}20, CK09, ECG^{+}22, GYK^{+}23, GKLM07, HMA23, HL97, HZLC22, IVCN20, JSMK14, JXJ22, JDT^{+}23, JTT01, JS08, JPP01, Kam93, KGK^{+}07, KGB^{+}21, KRA^{+}23b, LDS^{+}21, LDM^{+}11, LAM06, LCCS04, LYZ15, LBB11, LLS^{+}16, LMHRG10, LTP19, MS07, MAF19, MO06, MKPM17, MMV^{+}14, SIPO17, TRB^{+}22, UOT83, WLC88, WTWT18, WLX^{+}23, WWCZ19, ZTAP21, ZDC^{+}23, ZWWC23, ZCT95, dMTB^{+}21]. Determination [MG86, DSG21, HHL99, JD99, MA94, TH90, YL23]. determined [MZPZ16]. Determining [QB92, BBMR89, HN20, SF98]. DEUNet [YBD^{+}24]. Developable [ZJSB22, CT13, FBR^{+}17, WF88]. developed [Egh83]. Developing [MR17, NFW^{+}24, Rix84, BBDM03, HFT^{+}99]. Development [AM10, CG93, CDGA84, Lau77, McW89, MMV88, RK84, Van89b, YNS94, Ano03p, Aum89, Beb75, CS04, Cor84, FGES96, GD87, JP10, LLHH94, LPZ^{+}21, MRS^{+}07, MGSC^{+}10, MBGK89, MWY^{+}10, OMF93, OSZ00, Pie90, Pra99, Rag80, She12, SET^{+}88]. Developmental [PM10]. developments [AA00, FCSB90]. Deviatoric [LM16]. Device [NM85, Ree85, BMR23, BCS^{+}99, BNS24, HTW^{+}19, MOS^{+}21, PS91, ZDL^{+}19]. device-agnostic [BNS24]. devices [AEA13, CL97, DGBNV18, Dix91, FRTT18, GJN^{+}21, KHK18, MYC15, RLT16, RLU^{+}19, RCLM18]. devised [Car92]. DF [HC95]. DF-expression [HC95]. DFSPH [RP18].

DGRN [YZC^{+}23]. DHOP [PCD^{+}15]. diagnosis [B^{+}00, GPTP10, KFH^{+}09]. diagnostic [Cha97]. diagnostics [BCS^{+}99]. diagram [MTM22, SMM20, ZGC^{+}19, ZYW^{+}21, dFP22]. Diagrams [Shi93b, BPGW11, CZ22, DG17, LKHM19, LW88, MD85, TKS11, Tes84, XCXC23, XZY^{+}23]. Dialogue [Han84, Web84, Web85]. diameter [BKR^{+}16]. diamond [dF24]. diamond-kite [dF24]. Dichotomous [Car92, Car93]. dichromats [WZC^{+}21]. dictionaries [DMG20]. dictionary [CBM23]. dies [JZR88]. diesel [BL82]. difference [CBU^{+}15, MK89, ZFSY04, WKO12]. difference-of-Gaussians [WKO12]. different [Frü91, GMM^{+}23, KP95, MMS89, Paq05, SRZK23, SSBT01]. Differentiable [BAD23]. Differential [HHZ^{+}22, BPS^{+}10, BW94, CYJ^{+}13, CLE13, Hod91, MRC15, NSS^{+}22, TMN^{+}00, VH02]. differentials [ZMKG11]. difficulties [VCHR07]. Difficulty [WJG^{+}19]. diffraction [CXXW20]. diffused [HL97]. Diffusion [Lee01, LBB11, AK15, Bou06a, CGM91, EKG06, Kur24, McG08, PS13a, PS13b, SVW23, WVY16, ZW20]. Diffusion-geometric [LBB11]. diffusion-limited [Bou06a]. Digital [APA^{+}11, Ano03p, Ano06-50, Ano07-33, AG94, CKS98, CMB17, Cor93, FHM98, FMS98, HEG98, HMHA98, IE98, PM84b, SGM97b, SK98, Spi02, SHG98, VP98, AF00, Ano05d, BPGW11, CS01, Cho06, CZ06, DTZ09, DBS^{+}11, DGA02, ER87, Fer01, FEJ04, FM09, HZC^{+}22, Jac95, Kon89, LCL15, LXPP06, LHC12, Mar07, MARI17, MH21, MF02, MPAC^{+}23, MTB18, NRTT95, OSZ00, SGM97a, SOdSC17, SOdSC18, SR97, SGBI02, SSW^{+}23, SLRP16, TR95, TBG^{+}20, VH02, WCH^{+}11, Yam94, YMYH12, ZLS98, ZK98, BSM^{+}22, MAAS15]. digitization [HAB75, SK06]. Digitized [CFMP84, EH96, Yin04]. Diglib [JM88]. dihedral [ZLLY06]. dilated [ZWWC23]. DiLight [CMB17]. Dimension [CDI12, IV93, Ano13n, BjOwKM12, Cas96, CNS^{+}06]. Dimension-independent [CDI12]. Dimensional [AR84, ARL78, BS77, BG80, CB78, FZPM93, Gab77, How79, Str85a, VP77, AS91, And98, Aro89, Baw97, Bec95, BFT23, CWTL08, CKCK09, CR99, DH07, DAG22, E^{+}00, EL22, Fal89, FJ03, GSF99, HTKRW88, HHLE17, IWM^{+}09, KP98, LW99, LPO20, Liv18, MB97a, MD85, Mor76, NMM09, NR07a, Oik98, OdlCA02, Osi05, PL97, Par75a, PTL04, PLM^{+}05, RFZ^{+}17, RO87, Roe00, RPM96, REG^{+}89, Sen03, SGS99, SA87, Sob89, SPY87, SCCS13, SO75, Ste75, SJ15, SF92, TZvD^{+}21, TWSH02, TS75, VJ06, WLYH19, WTF95, WC89, Wüt98, ZNGN16, ZMM^{+}90]. dimensionality [MJEG21, MCMT14, dATNMC^{+}22]. Dimensions [FN10, ABMC^{+}15, Bou06a, Bro07d, CDW11, DW89, DSB96, EPS96, FN08, GMNS93, NAO13]. DIMNet [ZLLL21]. Direct [BAG03, Gab77, LCZ^{+}11, PX06, SH03, WKE04, CCC^{+}16, FRTT18, JPCS18, KSR99, KJS17, MPA^{+}10, PJ21, QD03, RMD11, Shi04, YNS94, dHG^{+}97, dGHM97]. directable [CAS^{+}15]. Directed [CFMP84, AK13, CACC24, DWH^{+}15, LQOW08, MT88, MSRB17, PF89]. direction [Ano03p, CC08, HJ16, MF02, Seg88, ZZD^{+}19, ZSL08]. direction-aware [ZZD^{+}19]. direction-oriented [CC08]. Directional [LSY11, BYQZ22, CC19, ÇB22, PWJ^{+}18, VP00, YR96]. Directions [Ano86j, KCK17]. directly [PPZ^{+}10]. director [SGBI02]. DirectVoxGO [PLVT23]. DIS [CGR87]. disability [Ano06-33]. disambiguate [MM18]. disappear [Nij04]. disassembly [GYL^{+}13]. disaster [CTJ^{+}14]. disc [ZCZ^{+}18]. disciplines [BHL^{+}94, Han95]. Discovery [HLS89, GSV^{+}18, LAB^{+}14, MHW10, MMS04]. Discrete [And94, CPLB14, DLW23, HY93b, LLW^{+}19, RBG^{+}09, ACSW75, ASWL11, BD97, BK89b, CM06, CDIM16, DSJ19a, DS18, FLV20, GM05, IP23, IFD15, Jac95, KB06, KKL02, Liv18, PP12, PS13b, Pic87a, PCPW15, SG92, TGG06, ZZDZ10, ZXLF23]. discretisation [SGS99]. discretization [Mok87, Mok88, SGR^{+}99]. Discriminating [EK85, IR06]. discrimination [WCS23]. discriminative [RR15, ZTF^{+}22]. discriminators [MSHL22]. discs [SMM20]. discuss [AD94]. discussions [IWM^{+}09]. disease [CMLR11, MGS^{+}22]. Diseases [NGAS23]. disentangled [HY23, NZZ^{+}21]. disentanglement [CYWM23, ZGZS22]. dishwasher [AAB92]. disk [AFW^{+}18, CJT96, GYJZ15, KKS93, KJS18, KKH23, MKPM17]. disks [McW89, Mil93]. Disparity [FJ17, KWK23]. Displaced [LAL11]. displacement [KNMP14, MLPB02]. Display [ARL78, Bro79, BK89b, CHMR78, MWA^{+}13, McC84, Tin86, WP77, AEA13, Dix91, EHBM82, FS80, FP89, FBP96, GSA89, GYD75, Jar75, JJL97, Kei89, KGM75, KKO01, KP95, MSD75, Par86, PR82b, RO87, RM91, SB94, SA86, SA87, Sta87a, Wir80, ZW88]. display/printing [Dix91]. displaying [CL97, Ent89b, OK20, RH85]. Displays [Ano96a, Lea85, SMG77, BN07, BGMP08, BFSE03, Bor91, CPS^{+}22, CP19a, C^{+}01, CWC^{+}14, IKM^{+}20, MAG^{+}12, MWDG13, SDS89, SFD06, VSKG03, WCH^{+}11, WM89, YS23, YWR03]. dissection [ZSM^{+}19]. Dissections [Ano12a]. dissemination [APA^{+}11]. dissimilarity [vBT21]. Distance [LXY22, MMS89, CCKW11, DZD^{+}23, Fou11, KYKK19, LLHH94, MA18, PS13a, PC23, RMD11, SA17, SBHS10, SCT^{+}14, TF18, TMP07, WY11a, YLYJ13, ZR12]. distances [CPM19, FSP15]. Distinctive [WUH^{+}15]. Distinguished [Ano11b, Ano13b, Ano24c]. distorted [DBS^{+}18, Zha98a]. distortion [BFSE03, CJT96, JK15, KWK17, SB94, SDIM13, WHFL16]. distortion-free [KWK17]. distortions [JH11]. distracted [AKB22]. distraction [WB24]. Distributed [Ano95a, Ano95b, Ano95c, Bad96, ET18, Gra93, PSZ96, PS86, BCS^{+}99, CA15, CHSD95, DKZ14, LZSG03, LRMS92, LH83, LVVC06, MO92, RBP96, SWL^{+}16, SSA96, XWW^{+}19]. distributing [GM04]. Distribution [OKBG08, BJP97, C^{+}01, CCH94, HXC^{+}23, HO88, HCC91, SMU22, SWZZ23]. distributions [Bec95, HGW^{+}24, Wir80, XLGG11]. distributive [Gin02]. districts [PPS20]. Dither [PN83, Zha98b]. dithering [AM12, Elb22]. DIVE [CH93]. divergence [BGPT18]. divergence-free [BGPT18]. diverse [H^{+}00a]. Diversified [CHZ^{+}23]. diversity [KAVM23, OP15, PMBS14, TC24]. divide [HWFQ09]. divide-and-conquer [HWFQ09]. Dividing [BPGW11]. division [PM91]. DL [WMDR08]. DL-STDM [WMDR08]. DLP [CTP^{+}21]. DMS [Ano95a, Ano95b, Ano95c]. DNN [WWO^{+}23]. do [Hol03, MRG^{+}19, TKD16]. d’objets [Le 77]. document [CGR87, Gru87, TMP07]. Documentation [RK84, IKM90, ZM07]. documenting [OFP^{+}11]. documents [BDKK96, CKS98, FHM98, FS86b, HK93, Hor83, MO92, SHG98, SLRP16]. Dodecafoam [GS99]. Does [AMZ^{+}24, CPS^{+}22, FBH^{+}21]. dogs [Cra02]. doing [Wes94]. Domain [KG20, KW14, SJZ^{+}23, BRSP15, Bru14, CWL20, FB04, FJS11, HLB^{+}06, JPP01, LPL^{+}20, LDLD22, LSWZ23, Lod21, LSS21, LLZ^{+}23, QSLS23, Rei97c, SD90, WHL^{+}09, WZZ^{+}18]. domain-based [FJS11]. Domain-specific [SJZ^{+}23]. domains [CD15, LG03, PR93, SV18, SFS^{+}21, SGS99]. dominant [BMW12, KP19, LCZ^{+}11, WX14]. données [Le 77, Nan77]. dot [Mas92]. dot-patterns [Mas92]. Douady [RPÁM04]. doubly [LMZ90]. doubly-curved [LMZ90]. down [LL00, MTTL82, SS04b]. Downloaded [Ano11c, Ano13c]. Dr [Enc15b, EJ15]. Dr.-Ing [Enc15b, EJ15]. Drafting [Str85b, Coo79, LHS87]. drag [GBP^{+}17]. draping [KL07]. draw [Car92, HE80]. Drawing [GD00, uHRBK06, Hod91, RAA^{+}08, AB03, CC01b, CSG^{+}17, EPB^{+}19, FASS16, GIZ95, LHG21, MAdS^{+}19, OK03, SVP82, SL01, TL13, US20, UWC90, YZ17, Zhu91]. Drawings [AR84, CFMP84, JK84, BC13, EVRW23, FJ03, GZW12, KC07, OK99, PP99, SD08, WWL^{+}12, ZPP^{+}23]. Drawn [Pic84, AWI^{+}09, CDF14, EK15, KKJ^{+}23, LKS07, LYZ^{+}24]. dressed [HSR^{+}09]. dressing [FGLW03, XLL^{+}20]. DRIFT [PdSP^{+}22]. Driven [GM86, NHR^{+}22, BGD18, BKM16, BRO22, CSCF08, DJC^{+}23, GWW^{+}22, GZW12, HLY^{+}19, IE98, LL11, LLW^{+}19, LXJL21, MMP18, MBA20, MS09b, MA17, NLSN11, Nov03, dSNJA22, RB20, RS99, SLX^{+}16, SZW^{+}14, SVV92, SBH07, SKCP99, TMK94, VHS12, WXG^{+}14, WHW^{+}22, WCL23, WCF^{+}23, YWC22, YGS12, ZLM^{+}15, ZZC20a, dAPG18]. Driver [Rix84, Rix83]. driving [JXJ22, SGC^{+}19, ZTAP21]. DRM [LDS^{+}21]. droplet [YJC99]. drops [CCW13]. Droste [Ley07]. drum [McC08]. drusen [ZWS19]. DSPOBJ [ARL78]. DSV [Ano06q]. DSV-IS [Ano06q]. DT [CJAR21]. DT-MRI [CJAR21]. DTV [CS04]. Dual [BYQZ22, HXH24, LZL^{+}24, RHK^{+}20, YBD^{+}24, AA01, CBM^{+}22, Jut94, KONS17, KEVD18, LTV08, LXT^{+}23, QSLS23, SXL^{+}23, TP89, WTWT18, WZZ^{+}21, YZC^{+}23, ZMH^{+}23]. dual-channel [LXT^{+}23]. dual-cone [WTWT18]. dual-domain [QSLS23]. Dual-encoder [YBD^{+}24]. dual-exposure [CBM^{+}22]. dual-paradigm [TP89]. Dual-precision [RHK^{+}20]. Dual-resolution [HXH24]. duality [Aro89]. dubbing [PZM^{+}23]. dueling [Rei97a]. d’un [Cib77, Dom77]. d’une [Le 77]. duotone [AXG^{+}13]. Dupin [GBF14]. during [CIT^{+}19, MVPL18, MCG^{+}23, WCA^{+}11]. DVST [Gab77]. Dynamic [ADOR02, CAS^{+}15, CBU^{+}15, Dav90, FBTT^{+}22, GBP04, HKS00, IMG22, JA84a, KW96, RRC^{+}18, XZCOX09, ZZF^{+}23, AJAC23, Ano13j, ALM19, ACGC22, BCC20, CHSD95, DW05, EKG06, ESAH02, FvdPT01, FCG^{+}07, GLC20, GLS^{+}20, HOCN07, JZY^{+}23, Kam93, KHK18, KRRS12, LaV07, LAM06, LWW08, LMY^{+}21, LHCL05, Lib91, LWFZ23, MS82, MPA^{+}10, Mor75, PF89, PR11, SL12b, SK16, SDS89, SBD15b, TOY^{+}14, WPH^{+}14, WLG04, YD00, ZDC^{+}23, ZSL12, dGGV08, MCTB14]. Dynamical [Per84, G^{+}01, Osi05]. Dynamics [Cet23, Coo02, DM79, JKK02, CC01a, CDR01, CHC^{+}24, CCW97, CCW99, CCW01, DR09, DLR^{+}10, JS09, Kei89, KW96, Lop92, MP89, MBC^{+}23, NR07b, PF97, Rei92, Rei04b, RPAM00, STT^{+}18, TNF14, vTCB^{+}21, CCW98]. dynamiques [Dom77].

E-Learning [Ano07w, TSC16, Pan06]. E.h [Enc15b, EJ15]. early [HIK05, IWM^{+}09]. Earth [BSM^{+}22, MAAS15]. earthworm [Pra92]. easily [SOdSC17, SOdSC18]. easy [CL96, DGV^{+}24, WWWW22]. Easybowling [PXH^{+}03]. eat [Hol03]. EC [LLX^{+}15]. ECCS [Ano95-29, Ano95-30]. echocardiography [Roe00]. ecological [WMW13]. economical [MSD75]. Ed [Ano05s, Ano01-37, Ano02-29, Ano03q, Ano04t, Ano06r, Ano07p, Smi78]. Ed-media [Ano05s, Ano01-37, Ano02-29, Ano03q, Ano07p, Ano04t]. Edge [DS18, GCYX23, HL97, HHN^{+}23, LKHM19, AKW^{+}16, ASZ^{+}14, BX99, BWYZ24, Cum89, DLZY14, EVRW23, Kle86, LSY11, LZ24, Muk86, NLS07, OMW13, RF97, VA96, WEWL99, XGC18, XZY^{+}23, ZGWP16, ZZD^{+}19]. edge-angle [XGC18]. Edge-aware [GCYX23, BWYZ24, ZGWP16]. Edge-based [HHN^{+}23, DLZY14]. edge-consistency [LZ24]. edge-oriented [Kle86]. edge-preserved [HL97]. edge-preserving [BX99]. Edgebreaker [LRS^{+}03]. edgeLBP [TB19]. edges [WLW05, Zhu91]. Edible [OUZS18]. Edit [XHL^{+}24, MX12, MX14, YY14]. edited [Cyc93]. Editing [KRRS12, AJ94, BAS^{+}20, DN22, EB10, HEWF13, HPD^{+}10, HQW14, IY18, IKTS22, IMMS82, Lew75b, LZL^{+}19, LMJH^{+}11, Ros90, San93, SW11, SRZK23, TEC^{+}14, WBB^{+}08, YY14, Zay12, ZYX^{+}19, ZQL15, ZHP^{+}19, ZZXT18]. Edition [Mil92b, dPCOO^{+}05]. Editor [Ano11a, Ano13a, Sar02, BMU^{+}16, Bon03, Bon05, CS88, CÖ91, HH93, KD94b, MTTLL82, Mud93, PS03, SS04a, Wu02, Ano83b, Ano12q, Ano23k, Ber84, Bon95, Bon97, Bru92, Bus98, The15, Egl86b, Enc83a, Enc86c, Enc93, Enc15a, End85, Gob93b, Gra83, Gri93, Hvi86a, Jor18c, Jor18d, Jor18e, Jor18f, Jor18g, Jor19d, Jor19e, Jor19f, Jor19g, Jor19h, Jor20e, Jor20f, Jor20g, Jor20h, Jor21e, Jor21f, Jor21g, Jor21h, Jor23c, Jor23d, Kir96, Kje95, Kot90, Kro91, Kro96, Lis95, Mac95, Pic92a, Pre84b, Rob91, Rog85, Sch86c, Ska98, SLCN09, Str83a, Str95, Str96, Sul85, Shi02]. Editor-in-Chief [Enc83a, Enc93, Jor18c, Jor18d, Jor23c]. Editorial [ACO01, AC09, Ano99d, Ano99e, Ano99f, Ano99g, Ano00a, Ano11d, Ano12k, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17h, Ano18e, Ano22i, Ano23j, Ano24f, BS04, BES01, Bon01, Enc93, Enc07, Jor07, Jor10a, Jor10b, Jor11, Jor18a, Jor18b, Jor19a, Jor19b, Jor19c, Jor20a, Jor20b, Jor20c, Jor20d, Jor21a, Jor21b, Jor21c, Jor21d, Jor22a, Jor22b, Jor22c, Jor22d, Jor22e, Jor22f, Jor22g, Jor23a, Jor23b, KF02, KS01b, KT95, Mar07, MW05, NK01, Spi02, TFS06, VHE10, WW10, BP08, Enc83a, HBG14, Ano85d, Ano86f, Ano94k, Ano95t, Ano02-30, Ano02-31, Ano02-32, Ano03v, Ano03w, Ano03u, Ano03r, Ano03s, Ano03t, Ano04z, Ano04x, Ano04y, Ano04u, Ano04v, Ano04w, Ano05u, Ano05v, Ano05t, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07q, Ano07r, Ano07s, Ano07t]. Editorial [Ano07u, Ano07v, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano18a, Ano18b, Ano18c, Ano18d, Ano18f, Ano18g, Ano18h, Ano18i, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e]. Editorial [Ano20f, Ano20g, Ano20h, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano24d, Ano24e]. Editors’ [Mud99, Ska99, Ano24c, DG96a, GS96, HW85, HS97, JG00, Kli00, PZK96, SB96, SS97, SP97, TA90, BF02b, CK02a, EFP02]. edits [ZZLZ21]. EDP [Ban85]. Educating [CMA10]. Education [Ade85, Ano94m, Ano01-37, CK02a, CB01, FK04, HBJ82, KP05, KJ01, LS79, MB10, PB23, WR79, Wol02a, Ano04h, Ano04i, AMC03, Bro92, BGK89, CMA10, Dia94, Dom94, Gna83, GHFH08, Han97a, Han97b, Jv95, Kal04, KS03a, LPL^{+}05, MD99a, McC95, McC96, MBGK89, MWY^{+}10, MY16, MK03, NFW^{+}24, OC00, ÖOK23, Paq05, PPL91, PS18, RW87, Shi06, TN02, VCHR07, dos01, Aok16]. Educational [Ano05o, Ano06k, Ano07l, Ano07m, Ano07p, PM84a, Pic84, Ano05q, Ano06j, Ano06l, AKB22, Bro00, BBMK21, FEJ04, GHS06, HS05, KJTS96, MUH10, SvD03]. educators [HL02]. Edutainment [TSC16, BBDM03, CLF^{+}06, Pan06, Ano06y, Ano07x, Ano07w]. EEG [CMLR11, GYD75]. effect [FGP^{+}10, KYT^{+}17, Ley07, LGLK16]. Effective [KNMP14, BC88, CEG18, FS98, GV07, MENS19, NFW^{+}24, WL15, Whi04]. effectively [KP18]. Effectiveness [AC77, MNI^{+}16]. Effects [MVRB18, RHFL14, SBS19, AJ94, AHK03, BMdSVR18, Bra95, CIT^{+}19, CSF20, HGVV16, LSE18, MTN22, Rok93a, RC94, SSB^{+}08, SBHS10, VNMP23]. efficiency [Car93, Frü91]. Efficient [ATB98, And85, AZF13, BS93, BK23, Bro07a, CL06, CSH08, CK09, CSSC00, CL95, CCC97, DI12, DGR93, FEVM10, GMd^{+}13, GZLW14, GYJZ15, Hop98, HLZ^{+}17, HY93a, JYC^{+}23, JHL^{+}12, Kau88, KLKE11, KB15, LH14, LYX18, LHH^{+}21, LAE^{+}19, MX12, MX14, MKC08, Mar01, NKMI23, NUM24, PM93, SSB04, Sri02, SLF^{+}24, WEWL99, WTL^{+}11, WFC14, XCL^{+}19, XZPG21, Yu99, ZZLZ21, ZZJ18, AEW91, AF16a, And98, BGK04, BP94, BBCG11, BS09, BK93, CCI12, Eas75, FLR^{+}23, FRC06, FDA03, GM04, HG02, HWR^{+}23, HLB^{+}06, IKB00, JJPP^{+}22, KLP01, KK08, KHK18, KCL18, LKS07, LPV95, LJCW04, LCL^{+}23, LWP02, dGMW16, MS08, PAJ19, PS08, Roj91, RP18, SSDS12, SMK08, SM92, SEC90, Ska93, SPL14, SOG08, TPM14, TVS^{+}03, WMRA^{+}15, XLW^{+}24, ZWS19, ZLL^{+}23a, ZGC^{+}19, ZP92]. Efficiently [CNC^{+}21, DB83, MCS^{+}18, SKH^{+}05]. effort [Lew75a, PMPR^{+}16]. EFPG [Ano07e]. EG [KPBR20, KKN^{+}21, Ano94l, Ano04-27, Ano04-62, DKFC20, vdPS08]. EG-IEEE-TCVG [Ano04-62]. EGPGV’07 [Ano07y]. Egress [HYP^{+}20]. EGVE [AMS22, BYC19]. EI10 [Ano04n]. Eiffel [STN95]. Eigen [WLP^{+}14]. Eigen-based [WLP^{+}14]. eigenproblems [XLHH21]. Eigenvalue [LM16, WYZ20]. eight [PR96]. Eighth [Ano02-33]. eikonal [PSK^{+}11]. EKF [HZ18]. elaborate [ZCT95]. elastic [AALJ20, KD15, LM22, PKRM21]. elastically [GÖT97]. elasticity [TNF14]. eLearning [Ano04-38, Gra02]. Electrical [DIE78, Pot77a, DBLC02, Men85]. electro [GYD75]. electro-encephalogram [GYD75]. electromagnetic [GANM21, RGGB02]. electron [GCvdS^{+}20]. Electronic [Mar87, Boa78, Mar91, McW88a, SLRP16]. electrophysiology [PSK^{+}11]. electrostatic [GD11, SHK18]. electrotactile [VKA^{+}23]. elegant [Lym89]. Element [Bar77, NYKN83, Arb92, BA90, CSHZ04, DSJ19a, HL06, KPH^{+}05, LCCM02, MCK12, STW82]. Elementary [Tak77, BM03, Mar02a]. elements [BD13, CRT04, KS09, LWLT11, Sch75, WMRA^{+}15]. elimination [RM91]. Ellipse [TT83]. ellipses [FH94, YR98]. ellipsoid [AS19]. Ellipsoidal [DR15, LJWcH07]. Ellipsoidal-blob [LJWcH07]. Elliptic [AAK22a, BGR94, SCM94, TT83]. Elsevier [Ano05m]. EMBASSI [HKS01]. embedded [ACG15]. Embedding [BDK17, BSB^{+}23, CLH^{+}23, MMM^{+}20, YYG16]. Embodied [SL18, BCS08, BN07]. embodiment [BDPR93, FBH^{+}21]. embolization [WCA^{+}11]. Emergence [CWGR01]. Emergency [OAYG10, RCB15]. Emergent [Pea02]. emerging [LC23]. emotion [BGD18]. Emotional [GSME04, LYZ^{+}24, MML12, LCCS04, WLL^{+}24]. emotions [CF96]. empathy [HZC^{+}22]. Empirical [ZSL08, JK15, WHZ^{+}18]. Employing [IM07]. empty [KT17, MKPM17]. empty-region [KT17]. enabled [BSMG24, CUD06]. Enabling [Bro00, SL16a, BF02b]. encephalogram [GYD75]. enclosing [LCK16, Sch12]. encoded [JPP01]. encoder [YBD^{+}24]. Encoding [BK20, SPS12a, BW98, GC86, Her83b, HW89, KDG96, LMY^{+}21, Nah23, RPKLMG23, SPS12b, SPY87, US20, WCH^{+}24, XCW^{+}09]. Encodings [Sch86b]. encryption [QN98]. end [SLM^{+}22]. end-to-end [SLM^{+}22]. endoscope [MBC^{+}23]. Endoscopic [KÇM00, YSW^{+}96]. Endpoint [LZG^{+}23]. Endpoint-based [LZG^{+}23]. Energy [GR93, HCC13, KHK18, BSPD10, CWC^{+}14, CCC^{+}16, HKHP11, JRS21, KP98, LS08, SP23, WWY06]. Energy-aware [HCC13]. energy-based [WWY06]. energy-compensating [LS08]. Energy-efficient [KHK18]. energy-minimizing [HKHP11]. energy-saving [CWC^{+}14, CCC^{+}16]. engaging [BBMK21, NFW^{+}24]. engine [BL82, CCCP04, HIK05, SWvB95, Wei84b]. engine/propeller [BL82]. Engineering [All77, Ano95-29, Ano95-30, Ano03a, CFMP84, Dan78, DW82, HBJ82, Mar79, Pot77a, Pre84a, Smi77, Aug84, Boa78, DVF06, FBR^{+}17, GTdS^{+}18, Gra83, GZW12, JZLP23, JMV90, KKJ^{+}23, MGSC^{+}10, MRSS^{+}18, MPOL96, NKA83, OH83, PVCM22, Yin04, dos01, Ano94p]. engineers [Cra02, YRD^{+}24]. Enhance [All77, KKJ^{+}23, WCHM22]. Enhanced [CZZ22, KH03, MBPF12, PSBD19, ZNT^{+}18, ANGH11, AM19, KCL18, KCS24, LPP^{+}19, MBC^{+}23, RR01, WTWT18]. Enhancement [GCYX23, JM88, KCS22, NZL^{+}21, PLJL15, Row82, CS98, CH94, CST05, CRD10, Egh80, HLS89, ISPS17, KPMT18, LSY11, LPR^{+}14, LYZ^{+}22, MKDM22, MMdOE^{+}22, MFP11, PECW22, RYNJ23, SSQL24, SXL^{+}23, SL12a, TVS^{+}03, ZLL^{+}23a]. enhancement-and-restoration [LYZ^{+}22]. Enhancements [YT87, SR97]. Enhancing [ACGC22, CKS98, MHM95, SLK^{+}23, ZFG^{+}20, VKA^{+}23]. Enigma [AdBC^{+}04]. enriched [HHKF10, Wal89]. enseignements [Cib77]. Ensemble [SPT18, RSB^{+}19]. ensembles [EL22, KBL22, RdCVL16, Wir80, dSdCLBC^{+}22]. entertainment [MTB18, Ano06-50]. entire [TSD87]. entities [AA01]. entre [Dom77]. entropy [Bru14, JYL23]. Envelope [GD00, MTM22]. Environment [AW93, Del80, Egl86a, HYP^{+}20, MPS85, PKK03, ZPL88, Beb75, BB03, BJS01, BB98, CSJ^{+}21, COPR17, Cze90, DLV06, Elv96, FLM^{+}15, GWP00, Gia15, Hor83, HJL^{+}93, KDS04, LRMS92, LRD07, LPL^{+}05, LhCE97, MOS^{+}21, Mar91, MCMV22, MPOL96, Mey79, MCKS06, MHM95, PHO^{+}23, PH90, RBW01, SCB^{+}24, Shi04, SMS09a, SMS09b, SHOC23, SL18, SLM^{+}22, SL02, SD90, SSA96, TMSPB09, TKB07, YSW^{+}96, ZWQ^{+}06, ZGdDL^{+}96, AHK03]. Environmental [Gro91, RFS22, KSH^{+}19, RŞH^{+}22]. Environments [Ano96a, Ano12n, AMS22, BYC19, CM93, LC18, RCD^{+}04, Tei96, ABG^{+}18, AA13, AKB22, A^{+}99, Bad96, BP94, BMdSVR18, BES00, BOH97, CA15, CEG18, CH93, CAAC20, CSS^{+}24, CGS^{+}96, CUD06, CKK96, DCV98, DSJ19b, DAHF04, DHJ^{+}97, DR09, DGC^{+}21, ESAH02, FBT93, FHFG99, GHCH03, GPC^{+}17, Gre96, GSB^{+}21, HJL07, JXM^{+}10, JHPhR11, KJ01, KK16, LSW^{+}23, LBLD11, LCCS04, LMD96, LVVC06, MR95, MBA20, MPW^{+}12, MMV^{+}14, NC12, OAYG10, OP13, dJONM18, OMGGG^{+}19, PSZ96, PZH^{+}05, PCY^{+}06, PJVH^{+}24, RdCAM01, SKSZ99, SSM11a, SU93, SC97, SB97, TPM14, WL15, WCW^{+}24, XPL90, vLM09, Ano02-33].

EPCGI [GB18]. EPCGI’17 [GMM18]. EPCS [LZG^{+}23]. equal [RPM97]. equality [MYF06]. equality-constrained [MYF06]. equally [ITW18]. equation [GWX^{+}18, GLDK95, MRC15, PM93, QM96, SJZW07, SKCP99]. Equations [SMG77, BW94, CYJ^{+}13, Coo01a, JKK02, PS13b, RM91, TMN^{+}00]. equipped [RFB^{+}21]. equivalence [ZXLF23]. equivariant [CCC00]. eraser [DTG15]. Erlangen [PS86]. ERLNET [PS86]. erosion [CBS^{+}14, PM13]. Erratum [Cad08, FB15a, KRA^{+}23a, MDS^{+}22, SMS09a]. Error [AK15, Ber79, CF11, GRW00, LLX^{+}15, BS98, CFZC19, GR93, HL97, VL98, WZZ^{+}21, YRS^{+}18]. error- [HL97]. error-bounded [CFZC19]. Error-controlled [CF11]. Error-Prone [Ber79]. errors [Bor91, HGH^{+}18, JPP01, Mok87]. ESBD [Cet23]. escape [NG03]. essential [Cad08, CWNA08]. estimate [BAC14]. Estimating [Can94, HSR^{+}09, HAL^{+}21, VMAL16]. Estimation [WTF95, ZPIS23, ATHL14, ASWL11, ARL^{+}20, CJXZ23, FJ17, FM09, FGL23, FCG^{+}07, GS11, JL23, KCH^{+}22, KCS22, LSK^{+}10, LZL^{+}19, LFY^{+}21, LZC^{+}15, MLP19, MCMV22, PWK95, PCKB23, PRBD22, PR23, SK99, SdSR^{+}19, SBHS10, SYMW21, TSY11, TOY^{+}14, TCL15b, VDOK19, WXC^{+}23, WPJP23, XLQP12, XXHM21, YIL09, ZCL^{+}13, ZZH^{+}24, ZYW23, ZMK18, ZLL^{+}20, ZHW^{+}21]. estimator [MCV18, WZZZ18]. estimators [DTZ09]. ETC2 [Nah23]. ETGraph [GWBD17]. Euclidean [MA18, NdSV20a, SCMT91, Var92]. eugenic [Goe95]. Euler [ZZY12]. Eulerian [AAAN23, RP18, WZLQ19]. Eulerian-DFSPH [RP18]. EURIM [Cho77]. Eurographics [Ano86d, Ano93a, Ano94d, Ano94e, Ano94b, Ano94n, Ano94o, Ano02-33, AMS22, BYC19, SIG02, Ano02-34, Ano04-29, Ano05x, Ano85b, Ano86g, Ano86h, Ano91a, Ano93c, Ano94h, Ano94m, Ano95u, Ano01k, Ano01l, Ano01-38, Ano02-35, Ano03x, Ano03y, Ano04-28, Ano05w, Ano05y, Ano06z, Ano07y, Ano24c, Hop86, KPBR20, KKN^{+}21, LPD^{+}18, PB23, SPK19]. Eurographics’2007 [Ano07z]. Eurographics’86 [Ano85c]. Europe [EF15]. European [Ano03h, Ano85c, Tho84]. EuroVA [BA23, BV22]. EuroVis [BA23, Ise21, IB22]. evaluate [LPV95]. Evaluating [Bou09, Bow95a, KYT^{+}17, Lea87, LSGFRC^{+}13, RB08, COPR17, MR17, SvL09, YRD^{+}24]. Evaluation [AP88, Cad08, CWNA08, GP86, Pos77b, Pro85, Rob78, WPB^{+}23, ZFSY04, dHG^{+}97, AMFH21, AEA13, AGM^{+}21, BP94, BC01, CFMS02, EHBA10, FRTT18, JYL23, KSM07, LaV07, LTBZ13, MAM^{+}24, MTS^{+}22, MST^{+}22, MMS^{+}17, MFL11, MSL^{+}19, Muk86, SY23, SBHC22, SHL^{+}24, SLK^{+}23, THL15, VKA^{+}23, ZP92]. evaluations [HN20]. EvalViz [MSL^{+}19]. Event [KKMT06, GM05, vdLdFvdEV23]. events [CTJ^{+}14, dSdCLBC^{+}22]. éventuelle [Le 77]. Evolution [Lie85, Gly91, Sal85, CF77]. evolutionary [CCM^{+}07, MRS^{+}07]. Evolving [HP01, NG95a, DHJ^{+}99, WR02]. Exact [BPS03, AMGA12, EA19, MAFL16]. exaggeration [LVM^{+}11]. Example [BBP10, Dai93, Dan78, MM10, SD21, DSJ19a, GI13, GZW12, LVLD10, LL06, LYZ15, MALI11, QKS01, Smi75, SZW^{+}14]. Example-based [BBP10, MM10, SD21, GI13, LYZ15, MALI11]. example-driven [GZW12, SZW^{+}14]. example-palettes [DSJ19a]. examples [Fra94, SCFF16, XLL^{+}20, YS17, YS21]. Excellence [Ano24c]. Exchange [KN88, Lie85, BG88, BBMK21]. excited [PTR^{+}92]. excluded [GD11]. Exclusive [DLN^{+}18]. executable [Ano13o]. execution [vOMRI^{+}15]. Exemplar [LTH^{+}19, PBK13]. Exemplar-based [LTH^{+}19]. exercise [CIT^{+}19, PTR^{+}92]. Exercises [HBJ82]. exhaustive [RB06]. exhibited [SBS19]. exhibiting [Rei02]. Exhibition [Ano85c, ACB12, Ano91a, Ano03i]. exigences [Dom77]. Existing [App87, PPG^{+}18]. existing-conditions [PPG^{+}18]. EXOD [LBV14]. exoskeleton [LhCE97, dGGDV11]. expandable [EW75]. expansion [BCMD17, YBTB23]. expected [Dod09, TT12]. Experience [KS96, Wei84a, ACGC22, Boa78, Cor84, FBH^{+}21, KSM07, LZR22, MCG^{+}23, Tei96]. Experiences [GO85, TM75, BG01, Kre93]. experiencing [SBS19]. experiment [BCHM02, CMA10, Pie89]. Experimental [BKCS79, GM78, LYZ^{+}24, Mit77, AHK03, dSC07, CC20, Enc95b, Gal88, MR05, TM75]. experimentation [CV77]. Experiments [ABCD93, Pot77a, vDLS02, NGAS23]. expert [CLE13, Ela86, IWM^{+}09]. Explainable [AS22, MSO^{+}20]. explanations [TZvD^{+}21]. explanatory [HJ03]. explicatifs [CF77]. explicit [SWS10]. exploded [MM18]. exploit [LTC^{+}20, NT00]. Exploiting [HJW97, KPFT03, ST97]. Exploration [FGM^{+}20, GVTA10, GGW22, GLA23, MVPL18, PHO^{+}23, AMPG22, BRHB20, Bou09, BH15, CJAR21, DDM^{+}06, DMG20, GSF^{+}19, GVC^{+}20, HGW^{+}24, LGWM17, LV02, MMD^{+}20, MR05, MAG^{+}12, Nar15, PSMD14, dMPF08a, dMPF08b, PJVH^{+}24, PSM16, RSB^{+}19, RŞH^{+}22, SBKB23, SS12b, TPRC18, TKB07, VR16, YWR03, ZNGN16, ZC07, dDH87]. Explorative [OCR^{+}19]. Exploratory [EGL^{+}95, RCLM18, SBR^{+}22]. explore [BEFV94, THQ^{+}16, TZvD^{+}21]. Exploring [AK21, ADHC^{+}23, DBW^{+}12, HFT^{+}99, Kor87, LST96, LHC12, MTN22, MCG^{+}23, NSS^{+}22, PdSP^{+}22, PWV^{+}18, TBS^{+}23, BAS^{+}20, LBV14, TN02, WLX^{+}23, FST97]. explosion [TKS11]. Expo [Ano06b, Ano03b, Ano02-41, Ano02-38]. ExpoBlend [Bru14]. exponent [Sas04]. Exponential [Cet23, RPAM00]. exponentially [GSSK^{+}13]. exponents [MH89]. exposure [Bru14, CBM^{+}22, KHTM17, MVRB18]. exposures [CBM^{+}22]. express [MAO^{+}12]. Expressing [Des00]. Expression [HL06, LYZ^{+}24, tHV10, CGZZ15, Dom77, HMW91, HC95, OBD^{+}23, VR16, YGS12, YBTB23, ZM07]. expression-invariant [tHV10]. Expressive [BH15, DKFC20, Mou15, Ano13m, LW89, GW17, Sez16, SA19]. Extended [DVG^{+}18, Nav89, TY24, CHC^{+}24, FPC10, GdMF03, LZ11, TPN95, Woo89, WAM17, WKO12]. Extending [SS93, CS04]. extensible [CR13]. Extension [Del80, Kan85, Sob89, CEM89, HK93, LS06, PK91]. Extensions [HE15, Lam00, MGOH96, BF15, MTT82]. extent [Can94]. External [RPÁM04, RPAM06, BDM^{+}16, RAA^{+}08, TWNL22]. extracted [LLLC11]. Extracting [CBM23, KS09, LWS15, Par88, PVCM22, VR16, WLX^{+}21, LBLD11, PSS04, ZS14]. Extraction [GY19, RBF17, TCL15b, BKR^{+}16, DA18, DW13, GWX^{+}18, Ger02, HJW^{+}08, HHN^{+}23, II22, KB05, Kuo01, LCL15, LZG^{+}23, LTR^{+}14, NBE^{+}04, PMTK01, PS08, RPSP^{+}19, Sch12, TPG99, XD08, ZMYH06, ZGW^{+}16]. extrema [DTG15]. extreme [SB94, dSdCLBC^{+}22, vTCB^{+}21]. extremity [SCB^{+}24]. extrinsic [MMM^{+}20, SDIM13, TSY11]. Extrinsically [HJ16]. extrusion [JZR88, KEVD18, SLYY97, WY03, XZY^{+}21]. extrusion-based [XZY^{+}21]. eye [AKB22, GWBD17, Kur24, PRM^{+}24, WSJJ24]. eye-movement [PRM^{+}24]. eye-tracking [GWBD17, Kur24].

F [Sou93]. FA [VW21]. fabric

[AKF^{+}20, CZB^{+}22, CLMA19, GCC23].

Fabrication

[Ano12p, CM18, FCM^{+}18, ZWSW22, AGM^{+}21, MNS^{+}19, PKRM21, ZPIS23]. Fabrication-Aware [Ano12p]. fabrics [PGB86, PACSG^{+}23]. facade [WS12]. Face [MSAR01, OBD^{+}23, BWdBP13, CYWM23, DGLRD18, GK96, Güd97, HWYL21, HHCM17, IC96, JD99, KMGL99, LLLC11, LZKJ23, LRY^{+}24, LLZ^{+}23, MMH^{+}21, PSBD19, PSS04, QD03, QWC14, VBP05, WSJJ24, ZHP^{+}19, ZSS^{+}18, tHV10]. Face-to-face [MSAR01]. Faces [DB83, Pic84, CGZZ15, DMM23, ESFGDZ97, KP22a, KD15, Nav89, Par75b, Sch95, SZW10, Zhu91]. faceted [Roc89]. FaceTuneGAN [OBD^{+}23]. Facial [DMV06, fLhLfT11, dPCOO^{+}05, DGC^{+}21, HEG98, HLS89, HY03, IC96, KFN23, KAVM23, LWG^{+}23, MZCD21, MAO^{+}12, NUM24, PSS04, WWCZ19, XWW^{+}21, YSZ22, YGS12, ZM07]. facilitate [BL82]. facilitates [SL18]. facilities [IC87]. facility [LRMS92]. Facing [ABMC^{+}15]. Factor [JNR85, BP94, PTW98, ZP92]. Factoring [AEW91]. factorization [AF16a, PHLW15]. Factors [Mag85, MHYN23, WJG^{+}19, WBP92]. factory [CP10]. Failure [Thi85, Ade86]. failures [XhKKL16]. fair [SZ09]. Falcidieno [WF22]. Falcon [SHD^{+}17]. Fall [Ano01b, Ano02a]. false [CUD06, JX96]. Families [AMT86, WCY^{+}20]. family [Dra98, Eng93, FM00, GBG04, RBP^{+}01]. fan [FDA03]. far [FJW11]. far-field [FJW11]. farm [NFW^{+}24]. fashio [KAAO75]. Fashion [GSME04, DWX^{+}23]. Fast [AMT86, BWD13, Bro79, CY94, EKP93, FV13, FPC10, GKLM07, HHL99, IO91, KYKK19, KS01a, KB05, KMV^{+}18, LSZQ21, LHCL05, LS05, LWFZ23, MAFL16, MAF19, Mor76, dATNMC^{+}22, PT16, PCGS15, RA03, Rok93a, RAF21, SHK18, SMK08, SSB^{+}08, SZW^{+}14, SGES12, TNF14, VDOK19, WIP08, WWY06, WZC^{+}21, WGS99, XLHH21, ZLL^{+}15, Zom10, ANE17, BWZ^{+}18, CPM19, GRPR08, KLP01, Kni95b, LŽ03, LHG21, LCD15, LDT02, Mey79, PLVT23, Pie83, QL23, SSV07, SL12b, Ska97, SPY87, WTM12, XWF^{+}20]. Faster [Baw97, IL97]. Fat [BGK04, Mes00]. fatigue [CMLR11, JXJ22]. faultlines [PMBS14]. FAUST [SGBP17]. FCC [CM20]. FDDAA [VH02]. feasibility [GHM^{+}96, LWZ^{+}23a]. feathers [BDSP22]. FEATS [Ano01-39]. Feature [BDPR93, CZL14, CPCS90, FFV^{+}11, GMP22, KM14, LH00, LXW^{+}10, LSW12, PLJ^{+}13, SJT20, SHS^{+}23, TEC^{+}14, WLL22, ZMYH06, ZLL^{+}21, ZML^{+}18, AC90, BHL^{+}15, CYW15, DW13, DYW^{+}22, Egh83, EBST14, EHBA10, GY19, GJN^{+}21, GO06, GT91, HJW^{+}08, HHN^{+}23, JZLP23, JPP01, KBL22, LD09, LCL15, LLP^{+}22, LWWY22, MCS^{+}18, MR89, NMM09, OPR92, RAK^{+}15, RBF20, Ros90, SK03, SLL^{+}17, TSY11, TH90, WY11b, WLL^{+}12, WPL^{+}23, WCH^{+}24, WWK11, XZL^{+}22, YF09, YPZ^{+}23, YLH^{+}18, ZjLW^{+}14, ZGW^{+}16, ZZH^{+}24, ZS02]. Feature-aligned [LXW^{+}10, MCS^{+}18]. Feature-assisted [SHS^{+}23]. Feature-aware [LSW12, PLJ^{+}13]. Feature-based [BDPR93, CPCS90, GMP22, KM14, TEC^{+}14, AC90, EBST14, OPR92, Ros90, TH90]. feature-free [DYW^{+}22]. Feature-preserving [CZL14, WLL22, WLL^{+}12]. featured [BZYB21]. Features [EK85, LYZ^{+}24, TA90, AOL96, BWZ^{+}18, CA17, DEW75, GTFB23, HBM23, HZLC22, JRZ^{+}23, KB15, LSK^{+}10, LWS15, LTC^{+}20, LBB12, MN90, MKKM18, NZL^{+}21, PCWD23, PM13, RB06, SJ09, TKD16, TAF16, WUH^{+}15, ZJSB22, ZMH^{+}23]. February [Ano20o, Ano21n, Ano22r, Ano23o, Ano24h].

Feedback [AC77, Bai07, BSMG24, CIT^{+}19, HGH^{+}18, KES22, LD05, SCB^{+}24]. Feeling [ABTW77]. Feiner [Mil92b]. FEM [AALJ20, SS82, XZ00]. FEM-based [XZ00]. Fermi [Tin86]. Few [GOdSC23, FK11, YS17, YS21, YK23]. Few-shot [GOdSC23]. fiber [EKG06, MMS15]. fibers [MMS15]. Fibonacci [Rei04b]. fidelity [CS16, KAFB18, LSWZ23, MR05, RSP^{+}19, SS13, VLD15, YSZ22]. fiducial [AMZ^{+}24, OUZS18]. Fiedler [BNPS10]. Field [NYKN83, Phi92, PF89, ATZM19, BGMP08, BB15, BKL17, BBP10, CPS^{+}22, CGB13, CNC^{+}21, CG96, FBH^{+}01, FWCS97, FJW11, Fou11, HJW^{+}08, HW22, HS03, ID17, JWZ23, JMC^{+}04, KBL22, LWWY22, hLfTxDdZ09, LSHL18, LR16, LCGD24, MLP19, MPL02, MAG^{+}12, MMK04, POBB09, RdCVL16, Rok93a, SHK18, TF18, WWL16, YLYJ13, YJC99, ZNT^{+}18, ZZF^{+}23, ZZJ^{+}14, ST23]. Field-directed [PF89]. fields [AMFH21, CLM^{+}19, CJT96, DWL^{+}03, DW89, DS18, EKG06, FS80, HJ16, fLhLfT11, LYW^{+}13, LM16, LWFZ23, LSW12, MVG^{+}21, MW14, PLVT23, PC23, RMP19, RPKLMG23, REG^{+}89, SMMS01, SPMA13, SLF^{+}24, Tap06, VJ06, WCF^{+}23, XZCOX09, XLGS16]. Fifth [Ano91b, Ano94p, Ano05d, Ano94d, Ano94e, Ano94n, Ano94o]. figure [BHTT94, BTC94]. figures [ASS22, Car92, Mes00, SB86]. filament [ZPIS23]. files [BPR90, KPL08]. fill [FS98, Kno87, SZ95, YCF18]. fill-lines [Kno87]. Filling [Nov03, PMV06, AKF^{+}20, AS95, AG94, AAB18, BGK04, BK93, CN05, MMALRA01, Oik98, Ran87b, SEKA19, TC00, WLL^{+}12, YHX10, Zha98b]. film [AJ94]. filmic [dSC07]. films [HKYM01]. Filter [TJ85, CC08, EWWL00, PLJL15, WLZL18]. filtered [MYL^{+}23, MTS21]. Filtering [JYL24, AnD19, ALC06, BM08, BWYZ24, CZL14, Fou11, GP91, GSSK^{+}13, JYL23, LCW^{+}18, Mou13, PLJ^{+}13, RC94, SD15, SLKD16, ZGWP16, Egh80]. filters [LCXL20, WWL^{+}12]. Finding [SK13, Ada97, FN10, IORM17, LL92, PF16, RP20, RP22, Sam89, XL10]. Fine [ZLS99, DWX^{+}23, GY19, GvK18, JXW^{+}22, LNL^{+}23b, LNL^{+}23a, TZT^{+}22, XZPG21, XLW^{+}24]. fine-grained [DWX^{+}23, GvK18, LNL^{+}23b, LNL^{+}23a]. fine-texture [XZPG21]. Fine-tuning [ZLS99, XLW^{+}24]. Finger [FSM94, KY97, AK13, CGG^{+}20, MA17]. Finger-Pointer [FSM94]. Finite [Bar77, CK96, CV97, NYKN83, Arb92, BW98, BD13, BA90, CBU^{+}15, CSHZ04, CRT04, CK93, KP18, KPH^{+}05, LCCM02, Mar02b, STW82, WMRA^{+}15, YHNC22, ZCL^{+}22]. finite-time [Mar02b]. fire [MHYN23, YHNC22]. firm [Aug84]. First [Ano94p, Ano94q, Ano03z, Kel86, Koh97, EVRW23, HW89, Lar03, Ano86d, Ano94r]. First-hit [Koh97]. First-time [Kel86]. fish [MPQG18]. Fisher [GTG06]. fisheye [RJS01]. fit [IY18, LK18, XLL^{+}20, GBG04]. FIT-recommending [GBG04]. Fit4CAD [RRQ^{+}22]. fitness [FGL23, NC07]. Fitting [Coc79, LYW^{+}10, RRB^{+}22, WWF^{+}18, BLW12, Chu90, CY94, FT02, GG14, IORM17, LD09, LMC13, Pha89, RRQ^{+}22, SS22, SWZZ23, WY11a, XCL^{+}19, ZCC^{+}16]. Fitts [MCG^{+}23]. five [GSF99, Ng95b]. five-dimensional [GSF99]. Fixed [VH02, Dra98, LB75, RHK^{+}20]. Fixed-point [VH02, RHK^{+}20]. flame [CGM91]. flames [ZZQW11]. flare [BK23]. flat [MLC^{+}22, QMHH91, VBTW13]. flatting [LZ24]. Flavor [Ela86]. flexibility [BRHB20]. Flexible [FGES96, LWW^{+}20, ACA07, CR13, FS20, ICNV21, Lar03, MSMP12, dMPF08a, dMPF08b, QWC14, RPP20, SR95, SKL^{+}13, ZEK^{+}17, dGGV08]. flight [Ano91d, BH91]. flip [Coo07]. Floating [Kra10]. flood [BK93]. flood-filling [BK93]. floor [MAM^{+}24, TMSPB09]. floorplan [CLXJ22]. Flow [Gra93, Lam87, AT11, ASS^{+}19, BH15, CBC19, ELC^{+}12, FWCS97, JZX^{+}21, LB75, LWS15, Lin15, LW88, Liv18, MRG^{+}19, MVPL18, PL97, RAHA88, WWL^{+}12, dLvL00]. flow-based [WWL^{+}12]. flowing [ZWQ^{+}06]. flows [CCW13, TWSH02]. fluid [ANE17, LF22, Lin15, PL97, PVM^{+}22, TRLX22, WZLQ19, XXX^{+}23, ZDD23, dVTT18, vTCB^{+}21]. fluid-hair [Lin15]. fluid-solid [TRLX22]. fluids [dSASP^{+}15, BGPT18, IP23, JS09, MP89, TNF14]. fluorescence [WHH21]. Flux [HHZ^{+}22]. FluxomicsExplorer [HHZ^{+}22]. fly [SBR23, STBG19]. FlyAway [HS99]. Focus [JHPhR11, MA15]. focused [BKR^{+}16]. Fold [Ano12a, IY18, MNS^{+}19]. folded [GCC23, Ste99, ZSS20]. folding [HHK^{+}13]. folds [SM22]. Foley [Mil92b]. foliation [CBC19]. foliations [WMZW22]. Font [LH91, PMZS97, Par93, SR02]. Fonts [Mar87, IWM94, SR97]. food [ZWR^{+}18]. fooled [SOdSC17, SOdSC18]. Footpoint [SCT^{+}14]. footprints [LBLD11]. Footstep [BPKB15]. Force [LhCE97, BDM^{+}16, CSHZ04, Cho06, DWH^{+}15, GBP^{+}17, GÖT97]. force-directed [DWH^{+}15]. forces [Rei95]. forcing [MH89]. foreground [MFL11, NZZ^{+}21]. foreground/background [MFL11]. forest [BLZD12, Edg00, MHYN23, YHNC22]. Foreword [ACP20, AS16, AF16b, AEJZ18, Ano12l, Ano12o, Ano12m, Ano12n, Ano13j, Ano13k, Ano13n, Ano13l, Ano13m, Ano15k, Ano21i, Ano22j, Aok16, AdSMD23, AMS22, ALP17, BK19, BV22, BTD^{+}22, BGT16, BFLP20, BDL^{+}22, BLZ22, BMP22, BPH20, BLM23, BYC19, BKS21, BKS23, CHPS20, CJO19, CP19b, CWT^{+}23, CH15, CM18, DKFC20, DK24, DS15, The15, EK18, EZ22, Enc15a, FT24, FGMR22, FS17, GPS18, GW17, GMMP21, GHK18, GB18, GMM18, GL10, GSKZ22, HSTR20, HJL07, IK21, JZZ16, JZ17, Joh18, Kiy19, KPBR20, KKN^{+}21, Kry14, KNDT20, LS19, LPD^{+}18, LC18, LC23, Lin10, Liu21, LPS19, MRT10, MLX18, MS16, MB14, MB10, Mou15, MTB18, MCPW21, NTB18, NOJS^{+}22, ND23, NTTF21, OP19, OOC22, PB22, PB23, PB10, PDK16, PBB21, PK17, RSK^{+}24]. Foreword [RMSB22, RS22, RGJdQ18, SK22, SH12, STP^{+}20, SPK19, SPO22, SCG23, Sez16, SC12, SNB17, SMMG22, SD16, SUF^{+}18, SSS15, SA19, TK20, TT19, TSC16, TMH20, TP17, TKdJO22, VB17, VT22, Zac15, ZT16, ZWSW22, vLM09, vLLSM09]. forgery [SIPO17]. Form [Ano12b, CF13, GT91, BP94, EB10, FBH^{+}01, FMP96, GCDL22, GdMF03, GO90, JZLP23, PK91, PTW98, RNM95, SGBP17, SE01, Tar22, VJ06, WBP92, ZZC20a, ZM92, ZP92, dMF99]. form-factor [BP94, ZP92]. form-factors [WBP92]. Formal [Car84, Gna84, QKS01, RS75]. formalization [UPT97]. Forman [ID17]. Format [HB91, CR13, LZ11, LZT^{+}24]. formation [PBN97, Wei99, ZZC^{+}14]. formats [DK97]. formerly [Ano01-66]. forming [MMV88, Yue86]. forms [HHG97, Ste09, TGG06]. formulation [GÖT97]. Forward [AZ23, NN15, PF97, AN99, GP91, Mil93]. Foundations [dSJ23]. four [DSB96, EPS96, FN08, Osi05]. four-dimensional [Osi05]. Fourier [SKS09]. Fourteenth [Ano06-27]. Fourth [Ano91c, Ano95k, Ano95l, Ano95m, SIG02]. Foveated [YCL^{+}21, GMd^{+}13, MIGS22]. FOX [MAG^{+}12]. FPAN [LZKJ23]. FPGA [CDGC94]. Fractal [Bow95b, Bro06a, CCKN01, Coo00a, DBG93, Fat01, Fat02, FN08, GBV92, IWM94, Lan97, McD06, MR96, PK86, QM96, Bou06b, Cas96, CZ98, Edg00, FP89, FC00b, Gon98, Kel00, Loy91, OdlCA02, Pe04, Pic95a, Pum96, SLL^{+}23, VJ06, Van09, Van10, WW08, Ye02, YHHS93]. fractal-based [Pum96]. fractality [KRP^{+}91]. Fractals [BGV93, GB91, Sas04, Shi93a, Shi93b, ADR01, Bou09, Bro07c, Car96, Car99a, Car99b, CB97, Dav95, Dav90, Ent89a, Ley05, Maj98, Mar04, Mar10, NG95a, NG03, Pic94, SS89a, Shi93c, SCNT03]. fraction [KONS17]. Fracture [MBP14, BDP^{+}17, DJH^{+}23]. fractured [DJC^{+}23]. fractures [SLQ^{+}19, YYY19]. Fragment [MCTB14, KJS17]. fragments [WLG04]. frame [CHSD95, GZZS06, Kuo01, LCWZ14, PMPR^{+}16, WQL^{+}17]. framebuffer [KKHS03]. frames [EPS96, SGC^{+}19]. framework [ACA07, BAS^{+}20, BC01, BHL^{+}15, CH91, DWZ^{+}22, DWX^{+}23, DGA02, ET18, Enc95b, FIC21, GSF^{+}19, GW06, Gho93, Gia15, Gna84, GD00, GV07, GCCZ14, HCC13, IWT13, JYLW14, KD94a, KB20, KD15, LWZ^{+}23a, LB19, LSR22, MZPZ16, MRW^{+}21, MMD^{+}20, MUH10, PKK03, PBG^{+}14, PP20, SGBP17, SM99, SLM^{+}22, WJW^{+}23, WXL^{+}23, WYC^{+}23, YWR03, ZLL^{+}23a, ZQL15, ZLLG18, dGGV08]. framework-based [LWZ^{+}23a]. Francesca [Fal89]. Francis [Mil92a]. Francisco [KGM75]. Fraunhofer [KKMT06]. freckle [LTH^{+}19]. Free [Ano12b, Ber77, CF13, GdMF03, GO90, JZLP23, SGBP17, BGPT18, CMM16, CM23, DYW^{+}22, EB10, FBH^{+}01, FMP96, FST97, FGP^{+}10, FGLW03, GCDL22, GS01b, JRS21, KWK17, MCG^{+}23, OK03, RNM95, SLQ^{+}19, SE01, WQL^{+}17, WLG04, YCF18, ZZC20a, ZM92, dGHM97, dMF99]. Free-Form [Ano12b, CF13, GdMF03, GO90, JZLP23, SGBP17, EB10, FBH^{+}01, FMP96, RNM95, SE01, ZZC20a, ZM92, dMF99]. free-hand [OK03]. free-sketch [FGP^{+}10]. free-viewpoint [WLG04]. freedom [MS82, YT83]. Freeform [Ano12p, SLRP16, WY03, BF02a, EK22, FTB12, KLKE11, LPV95, vdBB07]. Freehand [JK84, GPTB02, GD00, RO13, WWF^{+}23]. Freestyle [SCC11]. Frenét [SGC^{+}19]. frequencies [DTZ09]. Frequency [SSQL24, FJS11, JRS21, KHK18, LSWZ23, PSK^{+}11, Sen99, XZL^{+}22]. Frequency-aware [SSQL24]. frequency-modulated [Sen99]. Fresh [CS16]. Freshman [HBJ82]. friction [CZB^{+}22]. Frieze [CGR98]. from-polygon [AMGA12]. front [PSK^{+}11, PWJ^{+}18]. frontal [KC07]. froth [IA99]. fruits [LYC^{+}15]. ftc [Kra10]. full [BZYB21, KMWW^{+}14, LTC^{+}20]. full-featured [BZYB21]. full-surround [KMWW^{+}14]. fully [QL23, WGS^{+}18c, WGS^{+}18a, WGS^{+}18b, WYZ20]. Function [BG79, CSL23, LS06, AJRV00, BKR^{+}16, Can94, CDR01, CC03, Coo00b, Coo01b, DDPT98, FA94, GR93, Gro94, HE15, KJS17, KLL^{+}15, LW90, LS07, MG09, Nik98, Nik06, dMPF08a, dMPF08b, Spr94, TACS22, Van09, WEWL99, ZLLL21, ZSH12, vWS04]. Function-based [LS06, LS07]. Functional [Pfa84, CBM23, LLL^{+}15, MMM^{+}20, WYZ20]. functionalities [KJTS96]. functionality [CLH^{+}23]. functionally [LAE^{+}19]. Functions [Ala85, NKNN83, ABMC^{+}15, AZ23, BAD23, BPS^{+}10, BSF13, BDK17, CP21, CTQ^{+}14, CGB13, Car93, Cas87, CVP^{+}16, CEM89, Dra98, Fay85, HKHP11, IP23, KL07, Koh96b, Koh97, LLLC11, Mel19, MMK04, PPP88, PF09, PS12, Pie89, RMD11, SMU22, SY23, SAMA97, Sar04, Sas04, Shi93a, SPS96, Suf88, XLSW22, OKBG08]. Fundamental [Pau88, Kon89]. furniture [GCLZ16, YLS^{+}21, ZZC^{+}20b, ZC18]. Further [Ree92]. fused [KEVD18, LYZ^{+}22, ZPIS23]. fused-filament-fabrication [ZPIS23]. Fusion [ZML^{+}18, BS09, CKM^{+}23, HG22, HXH24, JXJ22, JYC^{+}23, LWWY22, LHL23, MRWL23, STBG19, XWW^{+}21, XZL^{+}22]. Future [Ano77, Ano86j, Ano95-38, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99j, Ano01-51, Ano02o, Ano02p, Ano02n, Ano02-47, Ano02-49, Ano03-43, Ano03-41, Ano04-47, Ano06-40, Ano06-39, Ano07-36, McC84, Str85b, Ano94z, Ano94-27, Ano94-28, Ano95-33, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano01x, Ano01z, Ano01y, Ano01-52, Ano01-53, Ano02-48, Ano03-42, Ano03-44, Ano03-45, Ano03-46, Ano04-45, Ano04-46, Ano04-48, Ano04-49, Ano05-34, Ano05-35, Ano05-36, Ano06-41, Ano06-42, Ano06-43, Ano07-34, Ano07-35, BCHM02, FMS98, HK15, KCK17, Roe00, Sal85]. Futures [Ano95j, Ano01-49]. Fuzzy [CR03, RAK^{+}15, CMS22, Des00, IKTS22, SP00b].

G [CNC^{+}21, IB06]. G&VC [Ise21].

G-strokes [IB06]. Gabriel [AAK22a]. Gain [MPTA^{+}22, MTN22]. Gaining [dSB04]. GALATEA [FP75]. game [EL04, Jef92, MWY^{+}10, MNI^{+}16, MMH^{+}21, Pan06, SBHC22, Mar02b, Mar02a, MNI^{+}16, Rei97c]. gameplay [WK14]. Games [LC18, OP19, SNB17, TSC16, Ano03p, Bro06a, CHPS20, DPS10, GSKZ22, KS98, MCV18, MTB18, ND23, Per02, Wee21, Ano07x, Ano07w]. gaming [CLF^{+}06, MLCMGR23, SSB^{+}08]. GAMORRA [MGH22]. gamuts [RHS^{+}94]. GAN [ALR23, CZB^{+}22, CYWM23, LCL^{+}21, LZKJ23, LSWZ23, LHL23, MSHL22]. GAN-based [CZB^{+}22, LCL^{+}21, MSHL22]. gap [BGK04, TC24]. gaps [Nov03]. garment [ABJ90, ADT^{+}16, LMC13, LK18, LYW^{+}10, RMSC11, XLL^{+}20]. gas [VH15]. gasket [Lak89, Szy90b]. gated [RFB23]. Gaudí’s [Bro08a]. Gauguin [HLS89]. Gauss [LM89, ZJSB22]. Gaussian [GSSK^{+}13, ME92]. Gaussians [WKO12]. Gaze [BSMG24, CACC24, Duc18, LPO20, AKB22, DGV^{+}24, PQ10, PAE^{+}21, PRM^{+}24, SBE20, WSJJ24]. gaze-adaptive [PAE^{+}21]. Gaze-based [Duc18, LPO20]. gaze-centric [WSJJ24]. Gaze-directed [CACC24]. Gaze-enabled [BSMG24]. gazing [Pot77b]. GCG [CH96]. GCN [HWR^{+}23]. GD3 [How79]. GDDT [RK84]. GDR [KUMW90]. GEANS [IMMS82]. gear [MSE20]. GEMO [LR90]. gems [GHS06]. gender [MVRB18]. General [CCC00, JLP00, JYLW14, RJS01, ZTF^{+}22, AB03, EW75, FT97, FH94, HH91, Lam00, LRMS92, LKHM19, LP83, MTTLL82, Nav89, Par93, PYD^{+}05, RF00, SB86, SP95, TDR01, YWR03]. general-purpose [MTTLL82]. Generalization [SNS06b, LWJ^{+}22]. generalizations [Mar02a]. Generalized [AKF^{+}20, CBC19, CBS^{+}14, CS85, CB78, GBP^{+}17, JPCS18, McG08, NR07b, Sar94b, SDIM13, BW92, CG87, CDR01, CAL^{+}04, FR92a, FR92b, JLP00, KS20, LAM06, PMM18, Roc89, SH94, YHW23, vEB98, BCC^{+}22]. generate [FCG^{+}21, GD11, HP01, WWWW22, Ye02]. Generated [PK85, AKF^{+}20, BW94, DS18, HL96, JM88, LP93, Li89, Mic90, SL12a, Shi93c, SPS12b, SCH^{+}18, Szy89a, Szy90a, Szy91, TC24]. Generating [HW22, Shi93b, SZ09, WWF^{+}23, CGH94, Jas88, PBH19, SS89a, Str86, WMZW22, ZLLY06]. Generation [AL78, BGV93, BT78, DSB96, Gab77, LJJPF22, NKNN83, NG03, Rag79, SGS99, SSW^{+}23, ATB98, ASS22, ASSF17, Ban97a, BR89, Bap99, BKM16, CSL18, CZB^{+}22, CYW15, CXGL23, Cas96, CLM^{+}19, CP98, CHZ^{+}23, CKCK09, CL95, CEM89, CD93, DKZ14, E^{+}00, Fay85, FNM20, FM22, GD95, Goe95, GBD88, HAL20, HY23, KKS93, Krö98, LD03, LRD07, LLLC11, LHG21, LLP^{+}22, LH91, LYS^{+}19, LWW10, LCCM02, LW88, LYXY19, LSW12, MLC^{+}22, MPR89, MPQG18, MMH^{+}21, NK01, ÖT21, OGSSLM^{+}07, PS91, Pic94, PJP23, RS75, RPP21, Rok93a, RBP^{+}01, SFVP13, SBR23, SGC^{+}19, SAB12, SZ95, SGR^{+}99, Spr93, Spr94, SP95, TT82, THL15, VJ06, WWH^{+}21, WCL23, XCZ^{+}16, YHHS93, YK18, Zha96, ZMK18, ZW88, dSMBG23]. Generative [BAPD23, CS16, HY23, HYP^{+}24, KKJ^{+}23, KFN23, NZZ^{+}21, OBD^{+}23, ZZF^{+}23, ZLL^{+}23b, ZDL^{+}24, dSMBG23]. generator [OP13]. Generators [AMT86, Bow95a, PAFL06]. generic [AFM93, SBS13, ZK98]. Genetic [JALS03, BC13, Goe95, HR04, ZSW08]. Genie [OO04]. genres [MGM^{+}23]. Gentropy [WR02]. genus [PT16, TGG06, ZLZG12]. genus-1 [TGG06]. genus-zero [PT16]. geo [KPSN04, MCP^{+}22, SA04, RFB23]. geo-data [MCP^{+}22]. GEO-Nav [RFB23]. geo-referenced [SA04]. geo-spatial [KPSN04]. Geodesic [Chr78, LSW15, CPM19, Har00, JZX^{+}21, Mou13, PKRM21, RBB^{+}11]. Geographic [BS82, BC88, DMM23, Jun94, VV89]. Geographical [Ano93e]. Geological [LNP^{+}13]. Geometric [ASWL11, AC90, Ano03-49, Ano04-30, AB11, BAPD23, BCC^{+}22, BT94, CL97, CTLG94, FR98, KF02, LT95, PPP88, QLCV96, STM^{+}04, SIE14, SAK90, AA01, BMMZ23, Bre01, CA17, CLXJ22, CGB13, CCM^{+}11, DZD^{+}23, DMG99, Dod09, Doh95, JH11, JAS97, JALS03, JCFN18, KZ04, LS90, LM22, LTS88, LSZQ21, LBB11, LSWL13, LB19, MVCNI21, MPS06, NS87, PMZS97, PSBM10, Pie90, PX06, RFB23, Ran87a, RRQ^{+}22, RRB^{+}22, Sar94b, SM92, SBS22, SM75, TBG^{+}20, TA90, XLM12, ZD04, ZYX^{+}19, Ano02-36, Ano02-37]. Geometrical [BG80, BD97, RLU^{+}19, Big86, FS86a, Fra86, IA99, IKTS22, SAK90, ZZDZ10]. GEOMETRIE [SKH83]. geometries [CXT18, ES22, NdSV20b]. Geometry [HIK05, LY08, Sou93, WHL^{+}09, ZZC^{+}14, Ano03-54, Ano06-48, BKV05, BK02, Bro08a, CP21, CLE13, Con91, DA18, DYW^{+}22, DFF22, EHSF17, Fun99, GMP22, HKBA17, KC07, KS03a, KG20, LJCW04, LZW^{+}21b, LCGD24, LR90, MRSS^{+}18, NSS^{+}22, Pau88, PEVW15, RBB^{+}11, SP00a, SCM94, SHS^{+}23, Var92, WW06, WHZ^{+}18, XWW^{+}21, ZO07]. geometry- [Fun99]. Geometry-aware [WHL^{+}09, DFF22]. geometry-consistent [LZW^{+}21b]. Geometry-constrained [ZZC^{+}14]. Geometry-optimized [LY08]. geometry-perceptive [LCGD24]. geometry/topology [MRSS^{+}18]. geophysical [Coo00a]. geoscientific [HKCL02]. geospatial [BLS15, BSM^{+}22, Gia15, KS15]. German [Han97a, Sch86a]. Germany [EF15, KKMT06, Ano02-55, Han97b, SIG02]. GESIM [LCGN92]. gestural [SKL^{+}13]. Gesture [SS96, CPC^{+}18, CGS^{+}21, FNM20, FGP^{+}10, IC96, KY97, LWW^{+}20, LMY^{+}21, RO13, TVL16, ZWWC23]. gestures [CGG^{+}20, CH12, ECG^{+}22, KSF15]. Getti [BMR23]. getting [BMR23]. GFLOPS [TC93]. GGR [JZX^{+}21]. ghosts [BK23]. giant [CL96]. Gierer [Coo01a]. gigantic [GM04]. Gingerbread [Gly91]. girth [XCZP14]. GIS [CSFG96, GVVJ99, MPOL96, NAS18, WWH^{+}21, dlLC99]. GIS-based [NAS18]. GKS [BBMR89, Bro84, BMP84, Cul84, Cze90, GO85, HTKV84, HRTK86, HR88, Kan85, LZ88, Mac85, Mag84, NP88, PVR87, PMK85, Pro85, Rix83, Rix84, Sin87, XZL88]. GKS- [HR88, NP88, Sin87]. GKS-standard [LZ88]. GKSGRAL [Cul84]. GLAM [ABG^{+}18]. glass [CCW13]. glasses [KSH17]. Glassner [Cyc93]. GLERIAM [Con77]. Global [NdSV20a, SKP99, AP22, BDK17, CHL^{+}11, CXCH23, CMS22, DLZY14, DM01, GLC20, GBP08, GLZ^{+}21, KHK18, KB15, LCZ^{+}11, LLZ^{+}11, LG23, LB19, Mok87, PAJ19, PM93, Stü98, TDR^{+}17, UPT97, XWW^{+}18, ZMH^{+}23]. globally [XXT18]. Globular [MP89]. gloss [JK15]. Glossary [Pot78, Enc83b]. glossy [Stü98]. Glycogen [ABG^{+}18]. Glycogen-derived [ABG^{+}18]. glyph [ROP11, SVW23]. glyph-based [ROP11]. glyphs [VW21]. GMDN [ZHG^{+}21]. GMFIM [KFN23]. GML [HHKF10]. GMM [LZT^{+}24]. GMM-ICQ [LZT^{+}24]. GMP [Ano02-36, Ano02-37]. GMT [YHW23]. go [YCO23, ZDL^{+}24]. Goal [AK13, BHTT94, SFC01]. Goal-oriented [BHTT94, SFC01]. Gogh [HLS89]. gold [Jac93, OFP^{+}11]. gons [KP19]. good [Bro07a, FF96, Hvi86c, S^{+}01, VFSL06]. Gordon [BF15]. Gouraud [Zha96]. GPU [ANE17, BKR^{+}16, CTN^{+}17, CTS^{+}10, CMM16, CMS22, CF13, FRWW14, GSV^{+}18, HO12, HBOS13, HZ18, KJ08, LSH^{+}12, LTR^{+}14, MG09, RC08, RVdF08, RdMF08, SG15, SS13, SGES12, SHCW22, TyZfTM12, WJW^{+}23, YK18, ZGWP16, ZGC^{+}19]. GPU-accelerated [YK18]. GPU-assisted [HBOS13]. GPU-based [HZ18, BKR^{+}16, CTN^{+}17, CTS^{+}10, HO12, MG09, RdMF08, SG15, ZGC^{+}19]. GPUs [HCC13, LCD15, NSL16]. graded [LAE^{+}19]. Gradient [JYL24, hLfTxDdZ09, PWK95, SJZW07, CZL14, CFZL16, CDIM16, KJS17, MLP19, MDJ^{+}95, RYNJ23, YZC^{+}23]. Gradual [SBHS10]. GRAFEDIT [MTTLL82]. GRAFIELDS [PF89]. grained [DWX^{+}23, GvK18, LNL^{+}23b, LNL^{+}23a]. Grammar [CSP19, WS12, DA18, FST97, HHKF10]. Grammar-based [CSP19, WS12]. grammars [TKZ^{+}13, TAF16]. granular [IWT13, ZK24]. grape [HJT^{+}13]. Graph [Cad08, CMLR11, Del80, KRA^{+}23a, LF22, MDS^{+}22, PSM16, PM22a, Abe04, CM14, DWX^{+}23, DG06, DWH^{+}15, EVRW23, FCG^{+}21, GDDA13, GWBD17, GTFB23, LKS07, LLLZ16, LBV14, LCXL20, MCS13, PR23, RVR04, SvLBF10, SLG97, SMM20, TSK98, TAF16, TAS09, TTKA23, ZHG^{+}21, BDP82, SB84]. Graph-based [PSM16, DWX^{+}23, GWBD17, GTFB23, LKS07, TSK98, ZHG^{+}21]. Graph-Theoretic [Del80]. GRAPHEX68 [DEW75]. Graphic [AMT86, GS84, JA84a, JK90, KS84, Kit77, RPM96, Sch85b, Sen99, Wec79, ACSW75, Ano06i, Ano06-49, BBH90, CS88, FP87, HTKRW88, KS96, Lan88, LZ88, OT88, Rei75, Roj91, Sen98, Sen03, TS95, QGW08]. Graphical [BA90, Bij85, Dai93, DEW75, HP91, IMMS82, JNR85, KS86, LB75, Lea85, Mag85, Pro85, Ric89, RK84, SB84, Szy92, AA92, BDKK96, CG96, DL93, DSB96, Enc95b, FM99, Fiu89, Jar75, LPV92, MTT82, Ner75, Pol83, PP90, Sch88, SD08, SSS90, SS75b, XZL88, van89a]. graphical-mathematical [SS75b]. graphically [KD94a]. graphically-oriented [KD94a]. GraphiCon [Ano95v]. ’Graphics [AD94, Ano02m, Ade85, All77, AB78, AW93, Ano85c, Ano93d, Ano94m, Ano94b, Ano94t, Ano95u, Ano95v, Ano96a, Ano97c, Ano99c, Ano99a, Ano99b, Ano01-36, Ano01-35, Ano01-61, Ano01-66, Ano02-46, Ano03-27, Ano04k, Ano04-31, Ano05o, Ano05m, Ano06m, Ano06k, Ano07l, Ano07m, Ano07n, Ano07f, Ano07y, Ano07-38, Ano08b, Ano09c, Ano10a, Ano11a, Ano11b, Ano11c, Ano13a, Ano13b, Ano13c, AN99, BWF18, Bap99, Bas77, Bax77, BSPR77, Ber77, Ber79, BGT16, BS77, BG79, Bjo85, BCMM07, BP08, BHL^{+}94, BKCS79, BKS23, CKS98, Car84, CGR98, CG87, CS80a, CZC02, CE80, CCW99, Coo02, Dan78, DW82, De 84, DZ93, Del80, DSM^{+}99, Dra98, DS15, DB83, Egl86a, Egl86b, EDKS96, EZ22, EK85, Enc83b, Ent77, Fat02, FHM98, FM99, FB15a, FCSB90, FMS98]. Graphics [FGMR22, Gab77, GPS18, Gin02, Gna83, Gra85, GLDK97, HBJ82, Har00, Her83a, HMHA98, How79, IE98, IS05, JV84, Jer85, JA84c, KA85, KD86, KUMW90, Kru99b, Kry14, KA86, Lam87, LD78, Lie85, LPS19, Mac78, Mac77, Man99, Mar78, ME77, ME83, MD99b, MS82, McC02, McC84, Mei83, MHK99, Mil75, Mil92a, Mil92b, Mud99, MNSJ99, MCPW21, NYKN83, Neg77, Nic84, Nik98, OKK83, OdlCA02, Per84, Per77, PK85, PBB21, Pos77a, Pos77b, Pot78, PP99, Pra99, PS86, Rag79, Ram89, Rei02, Rie78, Rob78, RMSB22, Ros92, Row86, SIG02, SRF08, SB77, Sch86a, SK98, Sch86b, Sen98, Sen99, SGR^{+}99, SGS99, SMS09a, Sou93, Sou94, SD16, SM75, Str85a, Str83a, SHG98, TK20, Tak77, Tho84, VP77, VB99, Wal93a, WGS^{+}18c, WFS^{+}82, Web84, Web85]. Graphics [Wel76, WWD^{+}95, Wis86, Wu02, YS21, YQY90, ZSW08, vJ84, AEMT88, AEW91, ACP20, AD85, AT11, AL10, Ano78, Ano83a, Ano91a, Ano94i, Ano94j, Ano01-41, Ano02v, Ano03f, Ano03k, Ano03d, Ano03-40, Ano03-48, Ano04j, Ano04l, Ano04i, Ano04-27, Ano04-51, Ano04-59, Ano05q, Ano05b, Ano05-29, Ano06j, Ano06l, Ano12o, Ano13m, Ano24a, App87, AM10, AMC03, AA01, Ban97b, Ban85, BF19, BDP82, BDRV01, Ben79, Ber84, BPS06, BF02b, Bou02, Bre01, BEKL00, Bru75, BJAN^{+}95, BGK89, BKS21, CGR87, CSP19, Cas87, CB97, CC05, CSK97, CT75, CK75, Coo79, Cor76, CÖ91, Cot75a, Cot75b, Cun00, DSN75, DPS10, Dia94, EW75, Ebe00, EK15, EF15, End85, ERB^{+}14, FF96, FEJ04, FWW13, Fiu87, Fra83, FVG15, FEJM75, FP75, Gna84, Gom90, GMM18, GHS06]. graphics [GKLM07, Gra83, Gri88, Gro91, GN80, GB75, H^{+}00a, HS05, Han95, Han97a, Han97b, HGS23, HO88, HS00, HRGD88, How88, HH88, HLS89, HvK87, HL02, HJL^{+}93, Hvi86a, IC87, IKM90, JM88, Jv95, JH89, JD75, JXY87, JMV90, Jor23f, JA84b, Kel86, KF02, KS04, KKS86, KB04, Kor87, Laf94, Lam83, LNFC95, LS84, Lew75a, LC23, LLHH94, LRR87, MTTLL82, MTTL82, MD99a, MP22, Mar76, Mar86, Mar10, MDSU88, MUS83, McC95, McC96, MSD75, McW88b, MB97b, MG08, Mey79, MY16, MGH22, MHM95, Mor76, MY97, MBP14, NT95, Nik06, Nov03, Oli08, OC00, OSZ00, PSZ96, Paq05, Par86, PMS87, PR82a, PPL91, Pfa83, Pfa84, Pic92b, Pol83, PSSP96, Pot75, Pot77b, PSR^{+}93, Pra92, PPV03, Pur87a, Qui91, Rag80, RG93, Ran88]. graphics [RAHA88, RJS01, RO87, RW87, Rog85, Rok93a, RWE05, SR95, SM99, Séq13, She88, SSB^{+}08, SAKB75, SS82, Sla92, Smi75, Smi93, SVV92, SS75a, Spr75, Sri02, SVP82, SO75, Ste75, SF91, Sul85, SS04b, Szy89a, TSD87, Tax04, Tay87, TS94, Tei96, TMK94, TMB^{+}05, TMH20, TS75, TM75, UKW23, Van94, Van89b, Var92, VCHR07, War76, Web87, WMFR89, WBL^{+}97, Wis87, Wol00, Wol02a, Wri75, XXC94, ZP04, ZFJ90, dDH87, dos01, tHd90, Ano95m, Ano05y, Ano86b, Ano95k, Ano95l, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97d, Ano02i, Ano03h, Ano03-28, Ano04m, Ano05j, Ano05-30, Ano07-27, Ano07-31, Ano08i, Ano12m, Ano21i, BK19, CMS12, EK18, Enc85a, Enc85b, Enc86a, Enc86b, GG96, Haz77, MLX18, PK17, SK22]. Graphics [SOdSC18, Tan94, ZT16, vdPS08, Ano85a, Ano01h, Ano01-65, Ano23a, Ano23j, Ano23l, Ano24f, Ano24g, BA23, Jor23e, SSS15]. graphics-based [Van89b]. Graphite [Ano05f]. Graphs [LD78, Moh77, AMPG22, ACG15, Bou02, CB01, Hvi86c, ITW^{+}20, KT17, KZ04, PK86, SS12b, SWH^{+}17, SJ15, MR96]. Graphtracker [SvRvL07]. Grapics [Ano01j]. grasp [VR16, JA84b, JA84c]. grasping [AN99, ES22, OMGGG^{+}19]. gratings [Rag09]. gravitational [YJC99]. Gravity [FFP^{+}21, CC01a, KB12]. Gravity-aware [FFP^{+}21]. Gray [PN83, WW08]. gray-level [WW08]. grayscale [JVS^{+}24, MALI11]. Greed [Bro07a]. Green [BDK17, TACS22]. Grenoble [Ano07e]. grey [PPS20]. grey-levels [PPS20]. Grid [Bar77, PLVT23, CMM16, CM20, CM23, DS93, JZX^{+}21, JWZ23, Koh96b, LTV08, LS84, Liv23, MAFL16, PKRM21, SGR^{+}99, XMD^{+}12, XWY15]. Grid-based [PLVT23]. grid-free [CMM16]. grid-particle [XMD^{+}12]. grids [BF02a, BLNZ22, BAG03, CZ06, FTB12, FH94, GRW00, Mok87, Mok88, PM95, WWL16]. groove [QMHH91]. grooves [QMHH91]. group [CCW97, CCC00, CCW01, CDGA84, DR09, Kon89, McC96, SFVP13]. groupe [CV77]. grouped [DLN^{+}18]. Grouping [BSF13, SPL14]. groups [CCW01, ICNV21, PAFL06]. groupwork [MUH10]. growing [FDA03, SHS^{+}23, VH15, XD08]. Growth [RPAM00, COM^{+}94, Düc90]. growth-regulations [COM^{+}94]. GRSI [Ano21j, Ano22k]. GSM [KPB96]. GsNeRF [LWFZ23]. GSS [OT88]. Guaranteed [GS01b, GCDL22]. Guest [BP08, HBG14, HH93, KD94b, Mud93, PS03, SS04a, Wu02, Ano83b, Ber84, BF02b, Bru92, Bus98, CK02a, DG96a, EFP02, The15, Egl86b, Enc15a, End85, Gob93b, GS96, Gra83, Gri93, HW85, HS97, Hvi86a, JG00, Kir96, Kje95, Kli00, Kot90, Kro91, Kro96, Lis95, Mac95, Mud99, PZK96, Pre84b, Rob91, Rog85, SB96, SS97, Sch86c, SP97, Shi02, Ska98, Ska99, Str83a, Str95, Str96, Sul85, TA90]. GUI [Pra99]. GUI-based [Pra99]. guidance [BB15, GKW^{+}24, HIS83, LSWZ23, MMT^{+}23, NZL^{+}21, SJB^{+}21, TBS^{+}23, YZC^{+}23, YRD^{+}24]. GUIDE [CMD99]. Guided [DA18, JYL24, AnD19, AN99, CLM^{+}19, CACC24, FSS^{+}02, HAL20, JYL23, KFN23, KGGP19, LYW^{+}11, LHH^{+}21, LQOW08, LZC^{+}15, MFP11, MMH^{+}21, PPV07, QSXT22, QSLS23, SCB^{+}24, SWZZ23, SHL^{+}24, SCFF16, TFF^{+}20, WH96, WZC^{+}24, YPZ^{+}23, YK23, YLH^{+}18, ZZH^{+}24]. Guidelines [BBMR89, Dom93, Mag85, CMB17]. guidewire [LCDN06]. guiding [AMPG22, ÇB22, LZL^{+}15, WLYH19, WB24]. GUIs [AA07]. Gulliver [OO04]. guyed [LRR87]. GVE [Ano94m]. Gyeongju [PKK03].

H [Kre93]. Hadamard [Egh80, Egh83]. hair [Lin15, SPS96]. Hairstyle [SLGQ23]. half [SO75]. half-tone [SO75]. halfedge [HY93a]. Halftone [BY88, Fri03]. Halftoning [IU09, SGM97b, KEVD18, SGM97a]. Halley [Ree91, Ree92]. halos [DVND10]. Hamel [LB75]. Hamiltonian [AXG^{+}13, CM20]. Hand [LYZ^{+}24, ZWWC23, AWI^{+}09, CGG^{+}20, CGS^{+}21, EK15, FRTT18, ILLC01, IA83, KES22, KSF15, LKS07, LMY^{+}21, MENS19, MAdS^{+}19, OK03, VR16, WLB^{+}23, WPJP23, ZSL12]. Hand-drawn [LYZ^{+}24, AWI^{+}09, EK15, LKS07]. hand-held [WLB^{+}23]. hand-made [MAdS^{+}19]. hand-object [WPJP23]. Hand-raising [ZWWC23]. hand-written [IA83]. handheld [DBW^{+}12, FFP^{+}21, PTY^{+}16, KB12]. handle [DL09]. handles [LRS^{+}03]. Handling [DGV^{+}24, Mar87, SGC^{+}19, FM17, Lin15, TH90]. Hands [Bai07, VBP05, MCG^{+}23]. hands-free [MCG^{+}23]. Hands-on [Bai07]. Handsketching [JNR85]. handwritten [SL16a, ZMYH06]. Haptic [AHK03, Ano96a, BCF06, GO10, PR97, AALJ20, A^{+}01, BMR23, BOH97, HR97, HTW^{+}19, IYH97, KK16, LFP10, LD05, LS08, PRRR13, PSBM10, DHJ^{+}97]. Haptics [SB97]. HARDWARE [Ano05y, AMT86, Ano07-27, De 84, EJRW96, FC00a, KSKS96, Ack96, Ano02i, Ano03-27, Ano04-31, BERW97, DH95b, HIK05, KNMP14, LLM^{+}90, MLM03, SLS03, Ano94b]. hardware-accelerated [MLM03]. hardware-based [BERW97]. Hardware-supported [EJRW96]. hardwood [OT88]. harm [FBH^{+}21]. Harmonic [SBY11, CLM^{+}19, KL07, LVM^{+}11, LXW^{+}10, fLhLfT11, PS13b, WMZW22, XZCOX09, ZMKG11]. harmonics [MCV18]. Harmonograms [Bro07b]. harmony [LZYQ22]. hashing [DLN^{+}18, XZPG21]. Hatching [KEVD18, BKL15, GI13]. Hausdorff [KYKK19]. having [COM^{+}94, PSBM10]. hazard [dGHM97]. hazard-free [dGHM97]. haze [LYZ^{+}22]. HCI [Ano91b, Ano07-28, Ano07-29]. HCPS [LWZ^{+}23a]. HDR [Cad08, CWNA08, CKM^{+}23, CJXZ23, CBM^{+}22, KHTM17, LMR^{+}19, UKL^{+}13, YBD^{+}24]. HDR-LFNet [CKM^{+}23]. HDR-video [UKL^{+}13]. head [CPS^{+}22, CP19a, HLY^{+}19, IKM^{+}20, KYT^{+}17, LY08, PQ10]. head-mounted [CPS^{+}22, CP19a, IKM^{+}20]. head-tracking [KYT^{+}17]. health [GLL00]. heart [CCM^{+}11, K^{+}00a, PiP00]. Heat [PS13a, STT^{+}18, GWX^{+}18, JZX^{+}21, Liv18, Pat15]. Heat-based [STT^{+}18]. Height [BKL17, FS80, SPMA13, XLGS16]. Height-field [BKL17]. held [WLB^{+}23]. helical [MR90a, ZHC11]. helixes [XCZP14]. Helmholtz [RdCVL16]. help [Ehl85]. hemianopia [ZFG^{+}20]. hemispherical [RBL95]. hence [Pot75]. Henon [MR89, RBP^{+}01]. herein [Car92]. Heritage [CMS11, PCD^{+}15, ASM12, DBS^{+}11, JH11, LXPP06, MGMB22, SLL^{+}21, SEMWC05, YMYH12, CMS12]. Hermite [LJH18]. Hessian [EA19, LYL^{+}17]. Hessian-constrained [LYL^{+}17]. Heterogeneous [QD03, CGG^{+}20, ECG^{+}22, GSF^{+}19, JY98, LW24, MCP^{+}22, WFC14]. Heuristic [LK18, MMP18, CCW13, CCI12, IORM17, MZ89]. heuristical [SB86]. heuristically [LW89]. heuristically-based [LW89]. hex [JDGS88]. hex-tree [JDGS88]. hexagonal [CR03, LS84]. Hexahedral [XGC18, GLA23, fLhLfT11]. HF [JZX^{+}21]. HF-GGR [JZX^{+}21]. Hidden [AP88, HF85, Koh96b, LD78, WLC88, YD00, AC89, Eas75, HE80, Hor82, HH91, HLS89, JD99, KS09, PR82b, PF80, Ran87a, RM91, SPGR93, YS97]. Hidden-curve [Koh96b]. Hidden-line [YD00, AC89, PF80, Ran87a, YS97]. Hidden-Surface [HF85]. hiding [TT12]. Hierarchical [Abe04, BCDD22, DGLRD18, DFNP84, FZPM93, KMGL99, PF16, PHLW15, AF16a, BD97, CS98, CB01, HKL^{+}23, How88, ILLC01, JDGS88, LXCW18, Mar02a, NP96, PBH19, RP20, SSM87, VHR^{+}18, YWR03, YF09, ZLG^{+}15, ZQL15, ID17]. hierarchically [BPR90]. hierarchies [KFW16, VHR^{+}18, WFG03, WG17]. Hierarchy [HHLE17, dSEM19, LAM06, XH88]. Hierarchy-based [HHLE17]. High [Ano93b, Ano94a, Ano13j, ASZ^{+}14, BBE14, HGVV16, HH88, KS84, KSH17, LZW^{+}21b, LZLS18, LSWZ23, Mac85, NBE^{+}04, PL97, RZY^{+}20, RLS^{+}12, She12, SL12b, WCH^{+}24, YSZ22, AEA13, ABMC^{+}15, ANE17, Ano91d, BDP82, Bec95, BFT23, BM08, CCCS08, C^{+}01, DAG22, DG96c, DL93, Egh80, FJ03, FCW^{+}10, GP91, GLC20, GBP04, GS04, HHLE17, HOCN07, HY23, IWT13, IC87, JJL97, KK12, LLC^{+}22, MZPZ16, MAF19, MMdOE^{+}22, MCS^{+}18, MMK04, Nah23, NMM09, NUM24, Pur87a, Rok93b, RSP^{+}19, SS13, SVW23, TZvD^{+}21, TT12, VLD15, WZLQ19, XZL^{+}22, YWH^{+}16, ZNGN16, ZLZG12, ZGdDL^{+}96]. high-capacity [TT12]. High-dimensional [PL97, BFT23, DAG22, FJ03, HHLE17, NMM09, TZvD^{+}21, ZNGN16]. high-dimensions [ABMC^{+}15]. High-fidelity [LSWZ23, YSZ22]. high-frequency [XZL^{+}22]. high-intensity [MMdOE^{+}22]. High-level [She12, BDP82, IC87, KK12, Pur87a]. high-noise-variance [BM08]. High-order [ASZ^{+}14]. high-performance [Ano91d, JJL97, ZGdDL^{+}96]. High-precision [RZY^{+}20]. High-quality [LZLS18, GP91, GBP04, GS04, HY23, MCS^{+}18, MMK04, Nah23]. High-Resolution [KS84, BBE14, CCCS08, C^{+}01, LLC^{+}22, NUM24]. High-Speed [Ano93b, Ano94a, FCW^{+}10, WZLQ19]. High-tech [HH88]. higher [CMA10]. Highlight [PH90, JK15]. highly [BH91, HS03, LG94, XLL^{+}20]. highway [ALD12]. Hilbert [SWL^{+}16]. Hill [Mil92a]. Hinged [Ano12a]. hints [LPD13]. hip [BL11]. hip-hop [BL11]. HIST [Jun94]. histogram [BWdBP13, WLZL18]. histograms [CDF14]. historian [MPAC^{+}23]. historic [Jun94, LBLD11]. historical [DDM^{+}06, PACSG^{+}23]. hit [Koh97]. HLODs [GK04]. HNA [PF16]. HNA* [RP20]. Hodge [RdCVL16]. Hole [MKPM17, AAB18, WLL^{+}12]. hole-filling [WLL^{+}12]. holes [GCC23, PMV06, QB92]. hollow [YCF18]. hollow-to-fill [YCF18]. holography [Fer01]. home [GLL00, MSAR01]. homme [CV77, Nan77]. Homogeneous [Aro89, Van94, AEW91, LYZ^{+}22, Mok88, Ska08]. homogenization [LLP^{+}21]. homology [GCDL22, PAFL06, YL23, ZDL22]. Honors [Ano24c]. Hookean [TTKA23]. hop [BL11]. horns [Ste09]. hosted [Ano07e]. hot [GSV^{+}18]. Hotel [Ano03a]. hotwire [JVS^{+}24]. Hough [ZDC^{+}23, ZMK18]. Hough-space-based [ZMK18]. HoughLaneNet [ZDC^{+}23]. HPNet [WCH^{+}24]. HQ [Nah23]. HSHTF [Egh80]. HT [RBF20]. HT-Based [RBF20]. Hub [EF15]. huge [GK04, JGA09, SW11, WBB^{+}08]. Hughes [Mil92b]. hull [CYCL09, KLKE11, SGES12, SMM20, TyZfTM12, VM15]. hulls [Mar09]. Humaines [Cib77, Le 77]. Human [Ano07-28, Ano07-29, CSL23, FGL23, GKW^{+}24, JNR85, LBLD97, LTP19, Mag85, MT00, San06, WP77, ASM12, Ano91b, AN99, BSAH^{+}23, BNS24, BPKG07, CSX^{+}19, CCM^{+}18, DFWW15, DJG^{+}04, GGW22, Güd97, HKL^{+}23, HLY^{+}19, JL23, KD94a, KSH^{+}19, KY97, KCS22, KCS24, LDLD23, LL00, LYW^{+}10, LLLZ16, LCL^{+}21, LY08, LPV92, Mag84, MQW09, MSHL22, Mit87, MPS06, Par75b, PBG^{+}14, PHLW15, PPD22, PS18, PR23, RP22, RZY^{+}20, RB08, RWE05, SH94, SLM^{+}22, Spr75, SCSG18, SSW^{+}23, TC24, WPB^{+}23, WXC^{+}23, Wil75, WWF^{+}23, XLL^{+}18, YSDG24, ZLLL21, ZZCL14, ZHG^{+}21, VTW23, Ano07o]. human-centered [BSAH^{+}23, BPKG07]. Human-Computer [Ano07-28, Ano07-29, Mag85, San06, ASM12, Ano91b, Mit87]. human-human [MSHL22]. human-like [RP22]. Humanistic [CLT07]. humanoid [MZ23]. Humans [SOdSC17, SOdSC18, FGLW03, HSR^{+}09, HZC^{+}22, MZCD21, Nij04, PHO^{+}23, WFC^{+}09]. Hybrid [AKPS00, BRV^{+}10, MRSS^{+}18, SGM97a, SGM97b, WZLQ19, YLH^{+}18, ASM12, BN03, CXT18, DVF06, FEBS07, Fun99, GT91, HBOS13, HCC13, KGK^{+}07, LRY^{+}24, LL12, RP18, Wal94a, WL15, WPB^{+}23, WCH^{+}24, ZYW^{+}21]. Hybrid-feature-guided [YLH^{+}18]. hybrid-level [LRY^{+}24]. hydraulic [PM13]. hydro [RŞH^{+}22]. hydro-meteorological [RŞH^{+}22]. hydroelectric [RJKV12]. HyOctane [RBP96]. hyper [AALJ20, KAÖ98, Liv18]. hyper-cubes [KAÖ98]. hyper-elastic [AALJ20]. hyperbolic [AJRV00, CCW01, OCCZ12, Rei97c]. hyperchaos [BPS06]. hypercube [SA87]. hypercubes [CBG22]. hyperdimensional [MB97b]. hyperdocument [RBP96]. hyperelastic [WXL^{+}23]. hyperelasticity [CHC^{+}24]. HyperGraph [HKL^{+}23]. Hypermedia [Ano07p, GN94, Kir93, MO92, OMF93, San93, Vää93]. HyperNet [MO92]. hyperparameter [FLR^{+}23]. hyperscanning [GSB^{+}21]. Hyperspace [SS89b]. hyperspectral [LJJPF22]. hypervideo [FB04]. hypotheses [ME92]. hypothesis [ZMK18].

i* [ZNGN16]. i-Media [ZR97]. I-patches [SVSV20]. IA* [CDW11]. Iamascope [FM99]. IARIGAI [Ano07e, Ano07d]. IASTED [Ano04l, Ano04h, Ano04k, Ano03d]. Iberian [Ano04b, Ano05c]. Ibero [PPV03]. IBM [BGV93, Pic94]. ICA [Ano94-32, Ano94-31]. ICAT [AMS22, BYC19]. ICAT-EGVE [AMS22, BYC19]. ICB [You89]. ICHIM [Ano01m, Ano01n]. ICIMA’05 [Ano05g]. ICIP [Ano01q, Ano01r]. ICL [LH83]. ICL/Three [LH83]. ICME [Ano03b, Ano06b, Ano02-38]. Iconic [GH91]. icons [KPL08, Spr96]. ICPNM [Ano01u, Ano01v, Ano01-44, Ano01-45, Ano02s, Ano02-44, Ano02-45, Ano01o, Ano01-40, Ano02j, Ano02-39, Ano02-40, Ano03-29, Ano03-30, Ano03-33, Ano04-32, Ano04-33, Ano05-27, Ano05z, Ano06-28]. ICQ [LZT^{+}24]. IDC [Sta87a]. ideas [Pau88]. Identification [Pur87a, AALJ20, HP91, PSBD19, Pop93, RHFL14, RBF20, ZWL^{+}22]. Identifying [DMM23, DB83, ITW18, vLvKV11, FIC23, MKKM18]. identity [PZM^{+}23, ZHP^{+}19]. IEC [HB91]. IEE [Ano03-31]. IEEE [Ano94r, Ano95-41, Ano95-39, Ano95-40, Ano02-38, Ano04-62, Ano07-39, Ano94q, Ano94s, Ano95z, Ano95-27, Ano95x, Ano95y, Ano95w, Ano01q, Ano01r, Ano01p, Ano01s, Ano01t, Ano01-42, Ano01-41, Ano01-43, Ano02k, Ano02-41, Ano02-42, Ano02-43, Ano03-32, Ano04-39, Ano04-40, Ano04-34, Ano04-35, Ano05h, Ano06-29, Ano06-30, Ano06-31, Ano07-39, Ano07-30, Lin10]. IFS [CB97, HHL99, Mar01, Mar02a, Mar03, Mar04, Mar09, Mar10, Ste09]. IFS-based [Ste09]. IFT [CGW07]. IGD [KKMT06].

IGES [Wei84a]. IHEP [KKS86]. II [CGH97b, Cib77, McW91a, YT83]. IIF [HB91]. IKS [CH91]. Illumination [CZR22, LZL^{+}19, MWS04, YK23, AVM05, AP22, CMS22, DM01, FV13, GBP08, HWYL21, HKYM01, HOCN07, KHK18, KJS17, LG23, MR05, NdSV20a, PAJ19, PM93, RR15, SSH16, SBD15b, Stü98, SKR15, TDR^{+}17, UPT97, VRV05, XLQP12, XWW^{+}18, ZXL^{+}21, ZZQW11, ZLL^{+}23a]. Illumination-dependent [MWS04]. Illumination-guided [YK23]. illuminations [Lan88]. illusion [AYA^{+}20, Hod91]. Illusory [AYA^{+}20]. Illustrating [GYL^{+}13]. illustration [APS09, BBP10, IA83]. illustration-inspired [APS09]. illustrations [LaV07]. illustrative [BRV^{+}10, CSC10, GRIG12, GVTA10, HGVV16, LLH17, RBFS10, VHE10]. Image [Ano94r, Ano04b, Ano05c, Ano07h, Ano07i, Ano07-37, BB93a, CG85, CWTL08, CCH94, CK93, Egh80, EKP93, Ent77, EAAY23, GS84, Gro92, GS89, HCC91, HF85, II22, JYL24, JZY^{+}23, KKS93, KPH^{+}05, LW24, LHH^{+}21, LXT^{+}23, LPD13, LG03, LYXY19, LHL23, LZ24, LCGN92, MYL^{+}23, ME92, Mou13, SBOT78, SIPO17, SPK19, SLKD16, SLL^{+}23, SBD^{+}94, WLQC18, WX14, YZC^{+}23, YPZ^{+}23, ZP07, ZLL^{+}23b, ZXD^{+}14, ZPL88, vWS04, AEA13, Ano94q, APB07, AAB18, Ban97a, BW98, BX99, Bap99, BND^{+}17, BYQZ22, BFSE03, Bim15, BS09, BBE14, BRPC18, CZB^{+}22, CH91, CIT^{+}19, CZ98, CR07, CWC^{+}14, CJXZ23, CTLG94, CÖ91, CLX^{+}19, CD93, DCV98, Day92, DLZY14, DSR11, EK15, EHBA10, FSS^{+}02, FS20, FLR^{+}23, FJS11, FLM^{+}15, FSM94, Fun99, GCLZ16, GS05, GN94, GDA^{+}13, GKW^{+}24, HWYL21, HXC^{+}23, HCX^{+}23, HKL^{+}23]. image [HKHP11, HEK22, HGJC21, Hor83, JK15, JYL23, JYL17, JYLW14, JJL97, JVS^{+}24, KHTM17, KFN23, Kir93, KK22, KB15, KCS22, KCS24, LMR^{+}19, Les01, Ley07, LSH^{+}12, LTC^{+}20, LCL^{+}21, LLP^{+}22, LSE18, LYX18, LS05, LWD^{+}18, LYZ^{+}22, LLL^{+}23, LRY^{+}24, LMJH^{+}11, LZL^{+}24, MENS19, MKC08, MXK^{+}19, MSD75, Mol96, MR17, NZZ^{+}21, NAK13, Owe94, PR97, PP16, PM84b, Pic88b, PJVH^{+}24, PdSP^{+}22, PCWD23, QSLS23, RYNJ23, RS99, RO87, SGC^{+}19, SB94, SD15, SW21, SSQL24, SZ95, SDD95, SJZW07, SXL^{+}23, She12, SWZZ23, SL12b, SvLBF10, SJZ^{+}23, SF92, SHG98, TL02, TDR01, UKL^{+}13, UPTd92, VP98, WW08, WXG^{+}14, WZC^{+}21, WSK^{+}22, WLY23, WCL23, WH96, WGS99, WKO12, XZPG21, YS23, YL96, YBD^{+}24, YHHS93, You89, YS97, YSD13, ZLM^{+}15, ZCL^{+}18, ZPN^{+}21, ZLL^{+}23a, ZS94, ZZJ18, ZSS20, ZZXT18, dSMBG23, vBT20b, Frü91, HB91]. Image-based [CWTL08, II22, LW24, WLQC18, ZXD^{+}14, BBE14, DCV98, EK15, FLM^{+}15, Fun99, KK22, LMR^{+}19, SJZ^{+}23, ZLM^{+}15]. image-guided [FSS^{+}02, WH96]. image-rendering [SDD95]. image-space [BND^{+}17, CWC^{+}14, YS97]. image-to-friction [CZB^{+}22]. Image-to-Image [YPZ^{+}23]. image-warping [SvLBF10]. image-warping-based [JK15]. image/video [LSH^{+}12]. Imagery [Cro80, San85, Kru99b, Pum96, SL18]. Images [BY88, DBG93, McC08, MCPW21, PN83, Szy89b, AF00, AS95, AM12, ALR23, And98, AOL96, ACC^{+}18, Bus97, CPG94, CH94, CEM89, CK96, CV97, Dix91, Elb22, FSV17, FP89, GV89, GCW23, GBV92, GZLW14, HL97, HZ15, HMHA98, HLS89, IU09, ILLC01, ISPS17, Jac95, JCT^{+}15, JPP01, JX96, dSJdML18, JRJP^{+}22, KG20, LH14, LNSW16, LSCJ23, Lod21, LMHRG10, Loy91, LHC12, MX12, MF02, MML12, NLdAL^{+}23, Nik98, OdlCA02, OK20, PMTK01, Par88, PPG^{+}18, Pop93, PPZ^{+}10, PR11, QB92, RNM^{+}19, RA03, SM07, SOdSC17, SOdSC18, SL12a, Shl83, Smo03, STBG19, SO75, SK06, TRB^{+}22, TVS^{+}03, TKM^{+}24, WKT21, WCHM22, WLB^{+}23, WPB^{+}23, WPJP23, WZZ^{+}18, WWF^{+}23, Ye02, YLH^{+}18, ZJSB22, ZLS98, ZGS17, ZWR^{+}18, ZXD^{+}14, ZLLY06, ZPL^{+}15, dMTB^{+}21]. ImageSwitcher [Ban97a]. imagination [NH83]. Imaging [Ano03-35, Ano04k, BK91, SGM97b, Tru86, Ano03d, Ano03-34, Ano04l, Ano05d, Ano13j, BK89a, FSS^{+}02, GTTC03, HMW91, KK21, KWK23, MDJ^{+}95, Sak02, SGM97a, SVW23, WVY16]. IMC [Ano95-28]. immediate [Bai07]. immersion [ZXH^{+}12]. Immersive [HSB^{+}10, KHS^{+}10, KMWW^{+}14, OC21, WBJ^{+}21, YH21, vDLS02, BB03, CDPS06, CEG18, CP19a, CAAC20, CSF20, DBS^{+}18, DGC^{+}21, GWP00, GA07, GMM^{+}23, HWSW19, KA22, LSW^{+}23, LPZ^{+}21, MVRB18, MYC15, NFW^{+}24, dJONM18, PRW^{+}22, S^{+}01, SCB^{+}24, SW19, SHOC23, SHL^{+}24, SU93, Wol02b, ZZL21]. Impact [DGC^{+}21, FBH^{+}21, LYZ^{+}24, MPS85, Ano03p, AMZ^{+}24, ZPIS23]. implement [LRMS92, LS18]. Implementation [Cul84, DM00, HBJ82, HTKV84, PR82a, PGR83, Pos77b, PF80, Sch75, SGGC05, Aon90, CH91, CDGA84, GN80, HR88, Hop98, LTS88, LRS^{+}03, NP88, PVR87, PKK03, SM22, SCMT91, SCM94, Sin87, STN95, TMB^{+}05, TC00, WTL^{+}11, Wei84a, Wri75, XZL88, Yue86, vOMRI^{+}15]. Implementations [BMP84, KMS^{+}97, ZR12, UPTd92]. Implementing [EWWL00, GO85, Sam89, SH94, TM75]. Implications [OC21, Bre01, Cra02, Gre96]. Implicit [CSL23, CD23, NKNN83, ASSF17, AZ23, BS01a, BSL^{+}13, BMS^{+}11, CIPT15, DG96b, Fay85, FPC10, FFV^{+}11, GM07, GVPN09, Gom14, GCRN23, HCV^{+}22, HKHP11, IORM17, JSP03, JOK^{+}07, LL11, LYL^{+}17, LOdF02, dCNPdFS14, RMD11, RGE07, RRC^{+}18, SVSV20, SWS10, TFY00, TMN^{+}00, TNU^{+}01, TW24, WYXM22, WCW^{+}24, YL23, ZHC11, ZGC15, ZLLL21, ZQ12]. implicitly [LTV08, LZT^{+}24, SJ09]. implicitly-connected [LZT^{+}24]. implicits [NSS^{+}22]. Importance [SKCP99, YPZ^{+}23, ZZQW11, LZ14, UWC90, YWC22]. importance-driven [YWC22]. imposed [HO94]. Impossible [Bro07c, Elb11]. impostor [DW05]. impostors [MCT05]. impressions [MR05]. improve [DR09, HR88, PCWD23]. Improved [GIZ95, IVCN20, JK15, MA18, Nah23, PTW98, SKS17, WP77, Ang97, GZW12, KP22a, LLLC11, MTS^{+}22, TPG99, WW08, WJW^{+}23, XZY^{+}23]. Improvement [SKKN10, MVK^{+}22, XGC18]. Improvements [Ger86, KB02]. improves [HKBA17]. Improving [BMMZ23, EVRW23, GR09, LLH17, LFL02, PRM^{+}24, VZP22, YMYH12, HR97, LWJ^{+}22]. Impulse [SMG77]. impulsive [LH14]. in/for [CKS98, FHM98, FMS98, HMHA98, IE98, SK98, SHG98]. inaccessible [LUMC04]. Inbetweening [CMB17]. Inc. [Pot78]. inCCsight [CJAR21]. Inception [HG21]. incidence [FB15b]. incidents [BGD18]. include [HS05]. Including [Ban97b, CKS98, DMG99, MT00, WKO12]. Inclusion [FT97, RF97, dMF99, AZ23, FTU95, MRF06, WGLS00]. incomplete [ASM12, LKL^{+}20, MBV18, PHLW15]. incompressible [CIPT15]. inconsistencies [WTF95]. Incorporating [LD05, KAVM23]. increase [RHS^{+}94, RPP20]. Increasing [KSH^{+}19]. increasingly [RCM^{+}20]. Incremental [BHH15, NKNN83, Ang97, Bap99, KB02, dATNMC^{+}22, RLU^{+}19, SLYY97, SHS^{+}23, YR96, MG98]. Independent [HLO16, NHR^{+}22, CDI12, FJW11, LMC13, Lod21, MCTB14, ST20, WVY16]. Index [Ano95-42, Ano99h, Ano00b, Ano01-30, Ano02w, Ano03-61, Ano04-63, Ano11k, Ano80, Ano83c, Ano85e]. Indexing [IL97, FJ03, vBT20a, vBT21]. India [AN99, DSM^{+}99, Man99, MD99b, MHK99, Mud99, MNSJ99, PP99, Pra99, SGR^{+}99, SGS99]. Indic [MNSJ99]. indicator [IP23, XLSW22]. indicators [GRW00]. Indirect [AMR23, LXB^{+}15, WTA11, GPC^{+}17, LTP19, SSH16, ZXL^{+}21]. individual [LYW^{+}10, McW90]. individuals [XLL^{+}20]. indoor [CLN^{+}16, GPC^{+}17, HFT^{+}99, KCU^{+}22, LWW^{+}20, LWD^{+}18, MOS^{+}21, MBA20, MMV^{+}14, OVWK16, PPG^{+}18, PRBD22, RGGB02, WZZZ18, ZLM^{+}15]. induced [AYA^{+}20, RB08]. induction [IKM^{+}20]. Industrial [WR79, Ano91a, AFM93, Com85, LPZ^{+}21, MAM^{+}24, RAK^{+}15, SET^{+}88, WKS03]. industries [Ano03p, Dro78]. Industry [PM84a, ABJ90, Ano02-53, MWY^{+}10, NFW^{+}24, SSW^{+}23, Tes84]. industry-university [MWY^{+}10]. inertial [BS09, PR23]. INESC [FCSB90]. INFADS [KAAO75]. infections [ASR^{+}22]. inference [LCL^{+}21, SXW^{+}22]. Inferring [BBHC15]. Infinite [YHHS93, FC00b, Sbe98, SKP99]. Infinite-corner-point [YHHS93]. Infinitex [HPD^{+}10]. inflation [CGWW16]. Influence [HZC^{+}22, POBB09, ALD12, RB08, dFP22, vOHR20]. influences [DMM23]. Info [Ano04x, Ano04y, Ano05u, Ano03u, RV01]. informatics [dos01]. Information [Ano93e, Ano95-28, Ano03v, Ano03w, Ano04z, Ano05v, Ano06t, Ano06v, Ano06w, Ano06x, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17h, Bax77, Bru14, Ehl85, Haz79b, HFP06, Mac77, ZML^{+}18, Ano94-29, Ano94-30]. information [Ano06u, ADT^{+}16, AMZ^{+}24, BL96, BF02b, Car92, CSCF08, CLE13, Fer01, GN94, HIS83, HAL^{+}21, HMHA98, Jun94, KR96, KN88, KPB96, LRY^{+}24, LTP19, Mar80, Mar82, PBG^{+}14, RGE07, RBFS10, TG02, TPK13, VV89, VKA^{+}23]. Informatique [Cib77]. informatiques [CV77]. infotainment [HKS01]. infrared [ADT^{+}16, YSZ22]. infrastructure [KKC94]. infrastructures [HKS01]. Ing [Enc15b, EJ15]. ingredients [AA92]. inhabited [LCCS04]. Inhibition [Row82]. inhomogeneous [hLfTxDdZ09, PBK13, SGGC05]. initial [ABCD93, LaV07, Lie85]. initialization [Liv23, ZSS^{+}18]. injective [Liv23]. Injectivity [ZZ15, YJLZ21]. ink [BPGW11, Lee01, SLRP16, XMD^{+}12]. Inlier [ZWL^{+}22]. Inlier/outlier [ZWL^{+}22]. inner [CGB13]. Innovative [CVHM03]. INP [Ano07e]. Inpainting [GCC23, FDGM18, FLR^{+}23, LLP^{+}22]. Input [Bij85, BKCS79, IA83, Mac85, LRHS14, MPAC^{+}23, Ner75, WY03]. inputs [CH12]. inputting [WMW13]. inquiry [JMC^{+}04]. INR [TW24]. inscribed [BR07]. insects [GCCZ14, CWTL08]. insertion [BG91, KB02, RBF20]. InShaDe [ATAG^{+}21]. Insight [TC93, BRO22]. insights [Ree92]. inspection [LQOW08, WTW^{+}23]. inspired [APS09, MSR^{+}13, OMP^{+}18, SEKA19, ZIP^{+}19]. instability [BjOwKM12, SYMW21]. installations [GA07]. Instance [FIC21, WSHY22]. instances [FIC23]. instant [PAJ19, WZS19]. institute [Mir87, GA88]. instruction [Cor76]. instructions [TMK94]. Instrument [Pot77a, FM99]. Instruments [Gra85].

Integer [ML79, ATB98, FH94, IKB00, Liv23, ML12, SWMdF21]. integer-based [IKB00]. Integral [Haz77]. integratable [Sch85a]. Integrated [Eas85, HMHA98, Sch86a, SL16b, For84, GDA^{+}13, Gru87, IYH97, LRHS14, SZEG93, SMS09a, SMS09b, SC97, Sta87a, Yam94, ZR97]. Integrating [ACV03, EL04, ESAH02, App87, CSFG96, FB12, LLW^{+}23]. Integration [GA83b, Nik06, OH83, SVT86, Ano94-32, Ano94-31, AP99, Ban85, DEST95, LPV95, MCP^{+}22, NSG05, PTW98, PCWD23, RAG05, REG^{+}89]. integrative [MIGS22]. Integrity [TJ85]. Intel [XWW^{+}18]. Intelligence [AS22, Ano04g, LWW^{+}20, Ano07f]. Intelligent [Ano86d, Ano94-29, Ano94-30, Ano02f, KS01b, PP99, ZWS19, Ano03-37, Ano03c, Ano04d, Ano04-40, HEW^{+}18, JD75, KD94a, LVVC06, ZI00]. intensity [BJP97, MMdOE^{+}22, MKG00, Rok93b]. intent [GR09, VR16]. intention [Nap95]. intentions [LWW^{+}20]. inter [ESFGDZ97, GSB^{+}21]. inter-brain [GSB^{+}21]. inter-faces [ESFGDZ97]. Interact [Ano05i]. Interacting [ACSW75, VSKG03, AA07, GD00, Pla00, ZBP^{+}18]. Interaction [Ano07o, Ano07-28, Dai93, Eas85, FGLW03, JNR85, Pro85, RMSB22, San06, Ano91b, AM10, BZYB21, BN07, B^{+}00, BES00, BNS24, CCM^{+}11, CJJ99, CS18, Duc18, FGES96, FBT93, GMM18, GSKZ22, HWSW19, KSH^{+}19, KS96, LSW^{+}23, LLW^{+}23, MK83, MLCMGR23, MR95, MSHL22, MCKS06, MCG^{+}23, ND23, OMGGG^{+}19, PR97, PSBM10, QSLS23, RHN03, RHS^{+}94, RWE05, Shi04, Sla92, SSS15, SL16b, SKL^{+}13, Vää93, WL15, Wes94, WPJP23, ZWS19, van89a, Ano07-29, Ano12m, Kry14]. Interaction-free [FGLW03]. interactions [CHPS20, FRTT18, Gro91, KD94a, LCWZ14, Lin15, LhCE97, MYC15, SLM^{+}22]. Interactive [ASC17, ANGH11, Ano06-50, APB07, BS77, BG79, BL82, BS01b, CC19, CS88, CIK99, CE80, CF99, Cho06, CMS22, DQF04, DSJ19a, DM79, Del80, DMG20, ERWS12, EB10, Fru94, GM78, GVC^{+}20, GI13, GYD75, GHCH03, GSY94, GWW84, Haz77, Haz79b, HKHP11, HJDR95, JH89, KJS17, KJS18, KD86, KK12, LFP10, LZSG03, LMR^{+}19, LCDN06, LS07, LVM^{+}11, LSH^{+}12, LZ88, LRR87, Mac78, MMD^{+}20, ME77, Mar79, ME83, MYF06, Met85, PSMD14, PZH^{+}05, Per77, Rag79, Rau06, RMD11, RS75, SB77, SBOT78, Sch85b, SRZK23, SD15, SS12b, SWS75, SS13, SB84, SS75a, Spr75, SJG19, SS75b, TMK94, VP06, WHH21, WSK^{+}22, WWO^{+}23, WCA^{+}11, WSWL14, YWR03, ASPO15, Ano03f, Bai07, BS04, BMU^{+}16, Beb75, BCC20, BH15, Bru75, BEFV94, BBP10, dSC07, Cas87, CACC24, Che06]. interactive [CS04, Chu90, CST05, CB10, CT75, CK75, Cra02, CS80b, DSM^{+}99, DG06, DGA02, EHBM82, EL04, ENE11, FP75, GCLZ16, GSV^{+}18, GBP08, GNL^{+}15, GB75, HS03, HIS83, HPD^{+}10, How75, IE98, IC87, IKTS22, JGA09, Jer95, KS98, KS01a, KLW12, Kir93, KS01b, Kor87, KB20, KAAO75, Lar03, Lew75b, LG94, LKC94, LZLS18, LWD^{+}18, LG23, LH83, MTTLL82, Mar76, Mar86, MUS83, MSD75, McW87, McW89, MPA^{+}10, Mey79, MBC^{+}23, Nar15, PGB86, PS13b, PMS87, PR82a, Pea02, PSK^{+}11, PF89, Pol83, PSM16, PJJSH16, PP90, Rag80, RAHA88, Rei75, RLD^{+}12, RMW^{+}17, RSN^{+}07, SEDT^{+}03, SWF^{+}20, SLKD16, SLL^{+}17, SdSR^{+}19, SVW23, SS82, SS02, Smi75, SDWE99, SGBI02, SVP82, Ste75, SHS^{+}23, Tay87, TAS09, TBM^{+}04, TS75, TM75, VHON04, VW21, WBB^{+}08, Wan04, WJW^{+}23, ZNGN16]. interactive [ZSS20, dDH87]. Interactive/Graphics [BG79]. Interactively [WHW^{+}22, UBW99, XLGS16]. interactivities [HS05]. Interchange [HB91, CG96]. Interchanging [Hor83]. intercommunication [dJONM18]. interest [BWdBP13, RS99]. Interface [Ano03-28, BK19, EK18, PK17, Ree85, Reh85, SK22, TK20, ZT16, ACB12, BG88, BKP01, CS04, CH96, DL93, FF96, FSM94, GVVJ99, HWSW19, H^{+}01, IC87, IKTS22, IYH97, JHPhR11, KY97, KPMT18, Lea87, Mag84, MAO^{+}12, Mit87, MSAR01, NFHS06, NGAS23, dJONM18, SVT86, SL18, SEMWC05, THQ^{+}16, Thi85, TS95, WMW13, YY88, YNS94]. Interfaces [Ano07-39, Ano12s, KA86, Mag85, Vää93, Web85, ABCD93, ANE17, DHJ^{+}97, Enc95b, Fiu89, GHFH08, HR97, HFT^{+}99, KYT^{+}17, LVVC06, MAH00, MO90, MCM^{+}18, NGAS23, PAE^{+}21, Sif99, SSS90, SKL^{+}13, VKA^{+}23, XL10]. interferometry [HKPL98]. interior [Com85, FB12, FSP15, WKW16, XWY15, YCF18, YLS^{+}21]. Interlocked [ZEK^{+}17, AKF^{+}20]. intermediate [SVCNM23]. intermittency [Coo07]. internal [Hoo91, TWNL22]. International [Ano86e, Ano91a, Ano94p, Ano94r, Ano94u, Ano94v, Ano94a, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano95v, Ano95x, Ano95y, Ano95w, Ano95-41, Ano95-39, Ano95-40, Ano01j, Ano01w, Ano01-35, Ano01-42, Ano01c, Ano02-41, Ano02-38, Ano02-55, Ano03a, Ano03b, Ano03-33, Ano03-52, Ano04-36, Ano04c, Ano04-32, Ano05z, Ano05-29, Ano05-41, Ano06c, Ano06a, Ano06b, Ano06-32, Ano06-33, Ano06-37, Ano06-50, Ano07h, Ano07i, Ano07f, Ano07-28, Ano07-29, Ano07g, Ano07c, Ano07e, Ano07d, Ano07-33, Ano08a, Ano13k, Ano15k, ALP17, BA23, BLZ22, BMP22, BPH20, BLM23, CP19b, FS17, GHK18, Joh18, SH12, TSC16, Ano91b, Ano91c, Ano94c, Ano94q, Ano01-65, Ano02v, Ano02-53, Ano03i, Ano03d, Ano03c, Ano04j, Ano04-39, Ano04-55, Ano05b, Ano05d, Ano05a, Ano05-28, Ano05-40]. international [Ano06i, Ano06-34, IK21, MCPW21, TMH20, Ano93d, Ano94t, Ano95-29, Ano95-30, Ano95-31, Ano01u, Ano01v, Ano01-44, Ano01-45, Ano01-46, Ano01-47, Ano02s, Ano02l, Ano02-44, Ano02-45, Ano04-37, Ano04-38, Ano05-30, Ano07n, Ano07-31, AMS22, BYC19]. Internet [Ano01-62, Ano03-34, Ano03-35, FB04, FMS98, IE98]. interoperability [HS03]. interplay [Qui91]. interpolant [RB20]. interpolants [BFRA11]. interpolated [GMP22]. Interpolating [MW12, Com01, JSP03, KR92, ZDT07]. Interpolation [BB93b, Cor82, GBD88, MR90a, Wal94a, ACGC22, BLL15, BF15, BB91, CP96, FSP15, KK08, KONS17, KJS17, LYW^{+}13, LW99, Mon87, NA02, OM96, PWK95, PB96, PC23, Pum96, RNM95, Sar94b, SBH01, WM95, Wil03, XLM12, Xu08, YF09, YR96, ZLL^{+}15, ZLL^{+}21, vRESH16]. interpolations [LJH18]. interpolative [SBH07]. interpolator [BSPD10]. Interpolatory [Sar92b]. INTERPRESS [Mar91]. interpretable [GGW22]. Interpretation [AR84, JK84, Sug84, TC93, Arb92, dPCOO^{+}05, PP20, PP99, ST02]. Interpretive [CC05]. interrelationships [TS87]. interrogation [FPC10, GOZ95, Ros90]. Interruptible [MO06]. intersecting [HH91, PMTK01]. Intersection [Bro76, AFW^{+}18, AO91, Boe91, CK09, DM00, GS01b, JOK^{+}07, JS92, LJCW04, LT95, LL12, MK89, MZ89, PS02, SF98, SJ94, TFY00, WA02, ZM91, Zub88, vBT20b]. intersection-free [GS01b]. intersections [FFV^{+}11, HPKE19, RHK^{+}20]. interspersed [SD08]. Interval [Ala85, SF91, CM15a, Lam99, Sar92b]. Intervention [SPK19, CCY^{+}03, CCM^{+}11]. interweaving [CLH^{+}23]. intra [CCM^{+}11]. intra-cardiac [CCM^{+}11]. Intrinsic [MMM^{+}20, RZF19, BBHC15, CW03, CLE13, LW99, LLL^{+}15, LXCW18, TOY^{+}14]. Intrinsic/extrinsic [MMM^{+}20].

Introducing [GHS06, DZD^{+}23]. Introduction [All84, ADL06, Bon95, Bon97, Bus98, CSD09, Cyc93, DG96a, Enc83c, Enc84a, FJA08, FR04, GM06, GS96, Gri93, GN04, HS97, JG00, Kir96, KD03, KS06, Kje95, KS03b, KKE07, Kro96, Lis95, Mac95, Mud99, Pic92a, PZK96, SB96, SS97, SP97, Ska98, Ska99, SS07, Str95, Str96, dA06, Ano83b, Ber84, BF02b, Bon03, Bon05, Bou06b, Bru92, CGR87, CK02a, EFP02, Egl86b, Enc86c, End85, Gob93b, Goh84, Gra83, HW85, HH93, Hvi86a, Kot90, Kro91, KD94b, Mud93, PS03, PLM^{+}05, Pre84b, Rob91, Rog85, Sch86c, Shi02, SS04a, SLCN09, Str83a, Sul85, TA90, Wu02, Kli00]. introductory [BWF18, Bou02, Cun00, RWE05, San00, San06, SS04b, Wol00, Wol02a]. intuitive [CZR22, JXW^{+}22]. invariance [ZGZS22]. Invariances [dCdLL14]. Invariant [ATAG^{+}21, BAPD23, CMdL21, HZ18, LRG11, Osi05, OCCZ12, RBB^{+}11, SvRvL07, VMAL16, WSX12, XCL^{+}19, tHV10]. invasive [CCY^{+}03, WH96]. inventing [Cun00]. Inverse [ABMC^{+}15, CKM^{+}23, FB12, FB15b, KRRS12, AN99, BAD23, BMT96, ENE11, EA19, FB14, FB15a, KNMP14, MMP18, PPV07, TKM^{+}24, YK23]. inversion [CC01a, Coo00a, FC00b, Ley05, LSWZ23]. inverted [CBC19]. Investigating [IWM^{+}09, MGM^{+}23]. investigation [HSD96, HKCL02, Shi93c, TS87]. invoking [Sch75]. involving [KB12, SSV07]. ionospheric [SVVS^{+}17]. Iota [Mac78]. IP [PR23]. iQmulus [VBS^{+}15]. iris [WSG10, Nic84]. irradiance [DMG99, MRC15]. Irregular [AM12, ABJ90, CD15, CZ06, SKKN10, WZW97, XM12]. irregularly [SHD^{+}17]. ISDB [Yam94]. ISMAR [Ano03-36, Ano04-39]. iso [FJW11, TPG99, XD08, CGR87, HB91]. iso-surface [TPG99, XD08]. iso-surfaces [FJW11]. ISO/IEC [HB91]. IsoBAS [LCD15]. isochromatics [RS08]. isoluminance [Con91]. Isometric [BLL15, DLR^{+}20, KLL^{+}15, LSW15, ZLL^{+}15]. isophotes [XLGS16]. isosurface [CP98, CSSC00, LST96, LFL02, LCD15, NBE^{+}04, NG05, PS08, VP06, Wan23]. isosurfaces [CL95, CGMS00, Ger02, ITW18, ZCT95]. isothetic [CZ06]. isotropic [LVLD10, LH14]. isovalue [MSE17]. Israel [Dro78]. Issue [AS16, Ano96a, Ano23j, Ano23l, ALP17, BLZ22, FS17, HKS00, HJL07, IS05, Jor23c, Jor23d, Jor23e, KP05, MLX18, Rei04a, ACO01, Ano04g, Ano04-56, Ano05m, Ano13o, Ano24f, Ano24g, BDL^{+}22, BMP22, DPS10, FJA08, FR04, FK04, Jor23f, KF02, Mar07, MK03, Pan06, QGW08, vLM09, vLLSM09, vdPS08]. Issues [Ano95-38, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano99a, Ano99j, Ano01-51, Ano02o, Ano02p, Ano02n, Ano02-47, Ano02-49, Ano03-46, Ano04-47, Ano06-40, Ano06-39, Ano07-36, McC95, Ros90, Ano94z, Ano94-27, Ano94-28, Ano95-33, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano99b, Ano01x, Ano01z, Ano01y, Ano01-52, Ano01-53, Ano02-48, Ano03-42, Ano03-43, Ano03-44, Ano03-45, Ano03-41, Ano04-45, Ano04-46, Ano04-48, Ano04-49, Ano05-34, Ano05-35, Ano05-36, Ano06-41, Ano06-42, Ano06-43, Ano07-34, Ano07-35, CMSF11, Cou92a, MSAR01, TN02]. Issues/Future [Ano02n, Ano06-39]. ISSV [Ano95-31]. iStar [ZNGN16]. ISU [LhCE97]. ISVC05 [Ano05-28, Ano05-31, Ano05-32]. IT-based [EMB^{+}98, KJ01]. ITAU [CDGA84]. items [WTW^{+}23]. Iterated [AJRV00, Can94, CC03, FA94, Gro94, JSG04, LW90, Nik98, Nik06, Spr94, Van09, vWS04]. Iteration [PK85, Rei92, Sel89, ASKCK03, Dra98, LZ11, Mic90, Roj94, Szy89a, Van10]. Iterative [Pop93, AnD19, BMMZ23, EVRW23, GSF^{+}19, KM14, LCK16, LWJ^{+}22, LZ11, MMS^{+}17, MNI^{+}16, VR16]. ITV [FB04, IA83]. Iv [Ano06-35]. IV’04 [Ano04-40]. IVA [Ano03-37]. iVRoad [WBJ^{+}21]. IWACA [Ano94a]. IWVR [Ano04-41, Ano04-42, Ano04-43].

Jacobi [Bri95]. Jadeite [CWTL08]. Jam [MNI^{+}16]. Jam-based [MNI^{+}16]. January [Ano12r]. Japan [Ano02-37, MWY^{+}10, MNI^{+}16]. Java [BDRV01]. jaw [Güd97]. JDCAD [LG94]. Jeffery [LB75]. jet [dSASP^{+}15, Ran88]. jewelry [ZjLW^{+}14]. jigsaw [CL96, EK22]. Joaquim [Ano07-40]. John [Mil92b]. Joint [Ano12s, RGdL^{+}18, WZZ^{+}18, BCMM07, LH14, MW12, MT00]. jointly [KKMT06]. joints [HCV^{+}22, LMY^{+}21, SCH^{+}18]. Jorge [Ano07-40]. journal [Ano05m, CMS12]. Journey [Bag93, Pic92b]. JPEG [CPG94, TVS^{+}03]. JPEG-coded [TVS^{+}03]. Julia [CCC00, Coo01b, DSB96, DMT03, Eng93, Ent89a, Gin02, Lak91, MR92, Mil93, Nor89, OdlCA02, She93]. July [Ano02-37]. June [KKMT06, Ano20l, Ano21p, Ano22m, Ano23m]. Justification [Mes84].

KaraKter [ZBP^{+}18]. Karate [ZBP^{+}18]. karst [FM22]. Keplerian [Dav95]. Kernel [GH91, JYL24, Pat15, SB84, CZZ22, Fou11, JK90, PS13a, SBS22, XZL88, XXHM21, ZZD^{+}19, BK91]. kernels [JT02, LLLZ16, ZCL^{+}22]. key [BPS06, RJS98]. kind [HWSW19]. kinematic [KRRS12, MR90a]. Kinematics [LCG19, AN99, ENE11, EA19, Med86]. Kinetic [CS06]. kinetics [BMT96]. KinetiX [OMP^{+}18]. kinodynamics [KRRS12]. kiss [Pic89]. kite [Fat01, dF24]. kite- [Fat01]. knee [LSCJ23]. knee-cartilage [LSCJ23]. knitted [EW99]. knitting [TKM^{+}24]. knitwear [DG07]. knot [ZCC^{+}16]. knots [Bro06b]. know [Fra86]. know-how [Fra86]. Knowledge [PSH^{+}09, Big86, DEST95, DWX^{+}23, Ela86, MMS04, MTN22, NMM09, She88, SD90, WJD^{+}09, vLLSM09]. Knowledge-assisted [PSH^{+}09, NMM09]. knowledges [Ano03-59]. Koch [HY93b, McC08]. Konstanz [KKMT06]. KRT [How88]. Kruger [Enc95a]. Kumite [ZBP^{+}18]. Kundalini [Coo00b, Coo01b]. Kunii [WF22].

L [APS09, FRWW14, LWW10, MRF06, NSL16]. L-Bench [NSL16]. L-BFGS-B [FRWW14]. L-systems [APS09, LWW10]. lab [Bro92]. label [CM14, GJN^{+}21, KAVM23, NKMI23, SLL^{+}17, WSWL14]. labeled [BFT23, FK11, GTG06, HHLE17]. labeling [AOB17, BCC20, BGLA18, CB10, LZL^{+}15]. labelling [BC13, Pha95]. labels [LYXY19, PAE09]. Laboratory [Mit77, Pot77a, Gal88, HRGD88]. labyrinths [Ada97]. Lactate [ABG^{+}18]. lag [Ban97a]. Lagrangian [IWT13, LF22, TRLX22, WZLQ19, ZNT^{+}18]. laminate [PSR^{+}93]. laminated [Pra99]. land [TS87]. land-use [TS87]. land-use/transportation [TS87]. Landfill [Rob78]. Landmark [MCS13, TSY11, WWCZ19]. Landmarking [HK99]. landmarks [NUM24, RKG22]. landscapes [KB10]. Lane [ZDC^{+}23]. Langevin [WLZL18]. Language [BKCS79, Del80, ME77, XLW^{+}24, BDP82, CG87, DEW75, Des00, DWX^{+}23, GV07, HMW91, LZ88, Mar76, Mar91, Pur87a, ZWP^{+}93, LS06]. Languages [Pos77b, dPCOO^{+}05, McW89]. laparoscopic [YRD^{+}24]. Laplace [AMHWW16, RBG^{+}09, WSX12, XLHH21]. Laplacian [HLO16, SZL^{+}13, TACS22]. Laplacian-membrane [HLO16]. Laplacians [VBTW13]. LARFNet [HG22]. Large [Ano12b, BLZD12, CA15, Ger02, HKCL02, HG02, KM21, KRA^{+}23a, LFL02, PJJSH16, TWSH02, AF00, AEA13, ACA07, Ano01-41, Ano03-48, BN07, BMH99, BRHB20, BRdSOS17, BLS15, CHZ^{+}23, CGW07, DR09, EHBA10, FTB12, Gia15, HPD^{+}10, HLB^{+}06, HK99, KRA^{+}23b, LBV14, MAG^{+}12, NAS18, PSMD14, PCGS15, RSP^{+}19, SSV07, SKH^{+}05, SKS09, SBSL04, SHD^{+}17, SXW^{+}22, TMSPB09, UL22, Van89b, WHH21, WWH^{+}21, WUH^{+}15, WLJT19, YWR03, YWC22, ZMH^{+}23, dVTT18]. large-data [Ano01-41, Ano03-48]. Large-Scale [Ano12b, KRA^{+}23a, BLZD12, KM21, PJJSH16, BRdSOS17, DR09, EHBA10, KRA^{+}23b, MAG^{+}12, SKS09, SXW^{+}22, WWH^{+}21, WUH^{+}15, YWC22, ZMH^{+}23]. large-screen [TMSPB09]. large-size [CHZ^{+}23]. Laser [SR89, Fer01, LQOW08, RBF17, RPFC01, vLvKV11]. laser-guided [LQOW08]. Latency [dGHM97, SvLBF10]. Latency- [dGHM97]. Latent [ÖT21, CLH^{+}23, GGW22, SLGQ23, ZGZS22, ZC07]. lattice [FEVM10, Oik98, Seg88, ZZC20a]. law [MCM^{+}18]. layer [CXXW20, HXH24, JCT^{+}15, WTL^{+}11, ZWS19]. Layered [GM04, LX08, SJ15, BES00, MW99, OP13, Yam94]. Layering [Kel00, ACO12]. layers [CKCK09, NLdAL^{+}23, WLW05]. Layout [BDKK96, CB10, CYW15, IKM90, RE22, XH88, YLS^{+}21]. Layout-aware [CB10]. layouting [GWW^{+}22]. layouts [DKM^{+}20, MMV^{+}14, PSP^{+}20]. LBARNet [HZ23]. LDR [LMR^{+}19]. leading [EF15]. learned [DMG20, KKPC23, WKS03]. Learning [Ano07x, Ano07w, BAPD23, CLF^{+}06, Che06, CBM^{+}22, CLX^{+}19, FCG^{+}21, FS78, GYY^{+}23, GCRN23, GTFB23, HXC^{+}23, JMC^{+}04, MC23, Smi77, TSC16, WWWW22, XhKKL16, XLSW22, XWWK21, YK23, ZDL22, ZK24, AKB22, BKV05, BKM16, BRdSOS17, BHL^{+}15, Cum89, FAZ21, FIC21, GOdSC23, GZSZ20, GTdS^{+}18, HG21, HBM23, HS14, KJ01, LDLD22, LDLD23, LLHH94, LSS21, LPL^{+}05, LB19, MKDM22, MH21, McC96, MMdOE^{+}22, MHM95, NRTT95, Nov03, PZH^{+}05, Pan06, PCY^{+}06, QLF^{+}09, RMP19, Roj94, SHL^{+}24, VZP22, VCHR07, WLYH19, WPB^{+}23, WCF^{+}23, WLX^{+}23, Wes94, Wol02b, WPJP23, YS15, YHW23, ZTAP21, ZXH^{+}12, ZLLG18, ZZH^{+}23, VTW23, EMB^{+}98, PP20]. learning-based [FIC21, LSS21]. Least [Coo78, Zha98a, CC01a, LZC^{+}15, MYF06, RGRG15]. Least-Squares [Coo78, CC01a]. Leather [MHZ94]. Lectures [Mei83]. left [FN08, PiP00]. legibility [DK97]. length [OMW13, Sbe98, XCZP14, YR98]. lens [AMPG22, BK23, RLS^{+}12]. lens-based [AMPG22]. lens-shifted [RLS^{+}12]. lenses [RGE07]. Leonardo [Sch95]. lesions [AAK^{+}22b, SBR^{+}22]. Less [LS79, CCY^{+}03, LSZQ21]. Lessons [CB01, WKS03, EL04]. Let [YCO23]. Letter [Jor18c, Sar02]. Level [DDM^{+}06, Mac85, PN83, ANE17, Ban97b, BDP82, CRD10, CV97, DL93, EB10, GN80, GNL^{+}15, HJW^{+}08, HE15, IC87, KK12, LRY^{+}24, PACSG^{+}23, Pur87a, RC08, SD15, She12, SK99, Tho84, TT12, UKW23, WW08, YK18, dLvL00, MGH22]. level-of-abstraction [SD15]. Level-of-detail [DDM^{+}06, GNL^{+}15]. level-set [EB10]. levels [JP10, PPS20, PACSG^{+}23, RS99, SP00b, ZZDZ10]. LFNet [CKM^{+}23]. LFPeers [SBKB23]. LGCPNet [GLZ^{+}21]. LGV [SD90]. LHX [IYH97]. Liberation [SDS89]. Libraries [CKS98, FHM98, FMS98, HMHA98, IE98, SK98, SHG98, AF00, ANE17]. library [BDRV01, OSZ00]. license [NLdAL^{+}23]. LIDAR [MPTA^{+}22, ZTAP21, HLZ^{+}17, LLC^{+}22]. LIDAR-based [ZTAP21]. liés [Dom77]. life [BGD18, CLF^{+}06, DSN75, MKHN01, NK01, PR96, VA96, ZGS17, Bou09, Rei97c]. LifeBrush [DSJ19b]. Lifelong [EMB^{+}98]. lifting [KPA12]. lifting-based [KPA12]. ligand [GRF^{+}22]. Light [APSS01, BP93, KD00, LG23, SPMA13, AYA^{+}20, BGMP08, BB15, CMB17, CNC^{+}21, CPLB14, DWL^{+}03, GMM^{+}23, LW24, LWW08, LMHRG10, MCV18, MAG^{+}12, OK20, PA91, PJP23, RLS^{+}12, SSDS12, SSH16, SSQL24, SKS17, WCHM22, YCL^{+}21, YSZ22, dMTB^{+}21]. light-field [BB15]. light-G-buffers [CNC^{+}21]. light-scattering [LW24]. Lighting [GCYX23, PH90, BCMD17, CDPS06, FB12, FB15b, KK22, KD00, MCMV22, MPA^{+}10, UKL^{+}13]. lightness [GLC20]. lightning [Kru99b, VR06]. lights [KJS17, KJS18, Rok93b, ZXL^{+}21]. Lightweight [HG22, HZ23, LHY23, PCWD23, QL23, ZLLG18, ZHG^{+}21]. like [IU09, RP22, RBP^{+}01, ZJSB22]. likelihood [CTQ^{+}14, WLZL18, ZZJ^{+}14]. limb [LLW^{+}23]. limit [FC00b]. limited [Bou06a, Dix91, YPZ^{+}23]. Lindenmayer [CG93]. Line [AR84, AB78, Bor91, CSG^{+}17, DBG92, EKP93, KFW16, KEVD18, REG^{+}89, SM92, Ska96, ACSW75, AC89, AS91, AA01, BR89, BPS03, BC13, Cas88, CC01b, CZ06, CS80b, Dor90, EVRW23, ENE11, GIZ95, Hor82, HH91, IB06, LWWY22, LWP02, MKC08, MLPB02, OK03, PR82b, PLM^{+}05, PF80, Ran87a, Ran91, RGMJ22, RM91, SMMS01, SH94, SLL^{+}17, SEC90, Ska93, Ska94, Ska97, SPY87, SL01, SLK^{+}23, TT82, US20, VBW12, VBTW13, WLC88, WWL^{+}12, WVY16, YZ17, YS97, YD00, Zhu91, ZCL^{+}22, SL01]. Line-based [KEVD18, MKC08]. line-coding [WVY16]. line-drawing [US20]. line-feature [SLL^{+}17]. line-of-action [SH94]. Line-of-sight [REG^{+}89]. line-search [ENE11]. Linear [BP93, BG80, BWYZ24, LW90, OKBG08, ZM91, ANE17, BFSE03, DRFRD06, DJC^{+}23, Fal89, GRPR08, GSA89, IA91, KH00, Lam99, MZ89, PS15a, PA91, TT82, TMSPB09, Van09, WEWL99, WWL16, XCL^{+}19, XLHH21, YR96, ZM92]. Lines [LYZ^{+}24, SBOT78, AWI^{+}09, APS09, BC13, Day92, EHBA10, GY19, uHRBK06, IIK12, Kno87, LD09, PK91, Phi92, SHK18, Tsu17, ZK95, ZGW^{+}16]. linguistic [Kas87]. linkage [WLL^{+}24]. linkages [EK15]. linked [CMLH21]. LinkNet [CMLH21]. lip [PZM^{+}23]. lip-sync [PZM^{+}23]. Liquid [CIPT15, RP18, SOC^{+}19]. liquids [RPP20, SLQ^{+}19]. Lisa [Sch95]. List [Ano80, Ano83c, Ano85e, Ano94w, Ano95-32, Ano95-42, Ano97e, Ano99i, Ano01-48, Ano03-38, Ano04-44, Ano05-33, Ano06-36, FD75, Ano07-32]. LiSurveying [LLC^{+}22]. literature [KCU^{+}22, PdSP^{+}22]. lithic [APA^{+}11]. live [GHS06, APB07]. LivePaper [RR01]. liver [CSJ^{+}21, CSS^{+}24, MTC02]. living [GZL21]. load [HWEB22]. lobes [MPA^{+}10]. Local [AVHT17, BW94, BHL^{+}15, CGG19, FSV17, GLZ^{+}21, LCGD24, MG86, PS12, RM91, BWdBP13, BYQZ22, BMMZ23, BFRA11, DK92, DLZY14, GS01a, GKT02, JZLP23, KB15, LSZQ21, LYX18, ODZS11, Sar04, SWZZ23, SXSY14, ZMH^{+}23, vBT20a, vBT21]. Local-global [GLZ^{+}21]. Localised [CP21]. locality [WLZL18]. localization [HZ18, ZCZ^{+}18]. Localized [DDW11, KP22b, LLW^{+}19, PZM^{+}23]. Locally [SBD15a, dGWvdW09, KP19, ZEK^{+}17]. located [KES22]. Location [LMD96, LWZ^{+}23b, MAFL16, SBG99]. location-based [LWZ^{+}23b]. locations [JRSM17]. locomotion [CP19a, GCCZ14]. locus [DAG22]. LOD [BGK04, CCL^{+}03, MCT08, RCBS10]. Lofts [SEKA19]. log [Bru14, TY24]. log-aesthetic [TY24]. log-domain [Bru14]. logic [IKTS22, OGSSLM^{+}07, PP90, RR92, SSM87]. logical [Szy91]. logico [Dom77]. logico-sémantiques [Dom77]. Logistic [SLG97, BF07, MH89, QM96, Szy90b]. LOGO [SCMT91, SCM94]. LogP [DG17]. logs [OT88]. Long [PCKB23]. Long-range [PCKB23]. longest [PMM18]. Longitudinal [SBR^{+}22, JZY^{+}23, SRA^{+}19]. Look [Lan97, BKL15, EVRW23, H^{+}00a, SK99]. loom [Pic95b]. loop [MW12]. loops [DL09, FB11, MT88, Mic90]. Lorenz [OK02]. loss [CPG94, WSJJ24]. losses [PMPR^{+}16]. Lossless [AF00, BW98]. Lot [CE80]. Low [DDQM98, KAFB18, LZL^{+}15, SvLBF10, VLD15, ALM19, AMZ^{+}24, CM06, CS16, KKNT88, LZG^{+}23, Mey79, NSL16, RHK^{+}20, SM22, SSQL24, Sla92, UKW23, WHFL16, WCHM22, WZZ^{+}21, YZWJ24, ZCL^{+}13]. low-complexity [CM06]. low-cost [KKNT88]. Low-fidelity [KAFB18]. low-level [UKW23]. low-light [SSQL24, WCHM22]. low-power [NSL16, RHK^{+}20]. low-quality [LZG^{+}23]. Low-rank [LZL^{+}15, ALM19, ZCL^{+}13]. low-time [AMZ^{+}24]. Lowell [EGL^{+}95, Gri88]. lower [MTM22, SCB^{+}24]. LPMNet [ÖT21]. LS [LHY23]. LS-MVSNet [LHY23]. LSPIHT [CFB15]. lumigraph [ARL^{+}20]. Lunda [Ger97]. Lunda-designs [Ger97]. lung [AAK^{+}22b, GWL^{+}20, YLH^{+}18]. lust [BEFV94]. LUT [CDGC94]. LUT-based [CDGC94]. Lyapunov [MH89].

M [SR95]. M-Buffer [SR95]. MA [RP20]. MA-HNA* [RP20]. Mach [Hod91, HO94]. Machine [HRTK86, Mac78, Web85, Ano94u, Ano94v, AKB22, HR97, HIS83, HS14, KMS^{+}97, LPD13, MK83, MVK^{+}22, MN90, PXH^{+}03, TR95, YHNC22]. machines [JS09]. Machining [MBN21, Bär90, CXT18, PPVT03]. macromodular [FEJM75]. Made [EF15, AAB92, CL96, DGV^{+}24, GvK18, MAdS^{+}19, OUZS18, XLL^{+}20, ZYX^{+}19, BBDM03]. Magic [AdBC^{+}04]. MagicBook [BKP01]. Magnetic [NYKN83, AP99]. magnification [AAAN23, FO21]. Mahalanobis [RA15b]. maintainability [MMF03]. maintenance [GZ99]. major [KAV^{+}88]. majors [She12]. make [KL02, QMHH91, RBLB21]. maker [GR09]. Making [HCLC16, MTS^{+}22, Séq13, ADHC^{+}23, ABJ90, BPD15, LWZ^{+}23a, Mil75, RHFL14, Sch95]. malleable [OK12]. Man [HIS83, Web85, GvK18, HR97, MK83, ZYX^{+}19, Gly91]. Man-Machine [Web85, HIS83, HR97]. man-made [GvK18, ZYX^{+}19]. man/machine [MK83]. manageable [BWYZ24]. Management [BT78, Haz79b, AD85, Ban85, DG06, ET07, GB75, H^{+}01, LCGN92, RC08, RJS98, SVT86, vOMRI^{+}15, Ano94-29, Ano94-30]. manager [MMF03]. Managing [SP00b, SBSL04]. mandala [CS01]. Mandarin [Ano03a]. Mandelbrot [CCC00, DSB96, DMT03, FR92b, FPR92, Gin02, Hoo91, Jon90, OdlCA02, PRÁM04, Phi92, PFR94, Rad96, Roj91, RPÁM04, RPAM06, RAA^{+}08]. Manent [CUD06]. manifold [ASC17, CDW11, CLX^{+}19, MX14, MCAH12, MG98, WHL^{+}09, YY96]. Manifolds [QV95, ACG15, EK22, GTFB23, Mel19, Osi05, SXG^{+}09, SJ15, VBTW13, dCdLL14]. manipulating [CH12, SLGQ23]. Manipulation [CHMR78, UBW99, AMR23, AK13, AA07, CXCH23, CEG18, CEM89, FFP^{+}21, GZLW14, KFN23, KSR99, LPO20, OMGGG^{+}19, RS75, SH03, Shi04, SHOC23, SWS10, WSK^{+}22, YNS94]. manipulations [PX06, RHM^{+}12]. manufacture [GA83b, PSR^{+}93]. manufacturing [BT94, For84, HBA13, IR06, LCA19, SHD^{+}17]. many [JYL17]. Map [BT78, BS82, BSGT03, Bri95, CSL18, Car99b, CZR22, Coo00b, DWH09, Dur89, FJ17, HWFQ09, HAB75, MH89, MVCNI21, MR89, SLL^{+}17, SP95, SLG97, VW21, WWK11, Zha96, ABG^{+}18]. MAPEDIT [HAB75]. mapped [Fiu87]. Mapping [CCCS08, CM93, GPC84, Jer85, LJJPF22, RHS^{+}94, Ack96, Aon90, ASZ^{+}14, AA00, BERW97, Bru75, Cad08, CWNA08, CL06, CKM^{+}23, CZC02, CDGC94, DG96c, DS93, EKG06, ERWS12, Ent89b, EJRW96, FCM^{+}18, GBA15, GP91, GLC20, HXA^{+}12, IMG22, KDS04, KLP01, KNMP14, LXW^{+}10, fLhLfT11, LX08, MA18, OCCZ12, PPS20, PP12, PP16, PT16, RLT16, SEDT^{+}03, SHK18, SLRP16, TWBP03, Tan80, WYZ^{+}11, WA75, XWW^{+}19, Yu99, ZBM00, Zha98a, ZLZG12, dMM19]. mappings [CCC00, WC89]. MapReduce [Gia15]. Maps [OCR^{+}19, Abe04, BF07, BTV83, CA17, Car99b, CBS^{+}14, Eng93, FJW11, GJN^{+}21, GO06, HOCN07, JC95, KM21, KCH^{+}22, LBLD11, LYS^{+}19, LLL^{+}15, Liv23, MAS14, Mic90, MML^{+}11, NP96, PMS87, PCGS15, RPM96, RPM97, RBP^{+}01, SK03, SK23, WBRV16, WCdA98, ZI00, SFS^{+}21]. Marathon [ZXH^{+}12]. marbling [AT11]. Marcel [For84]. marching [CBG22, CK75, KONS17, NBE^{+}04, NY06, TPG99, WA02, CEPS13]. MARF [ST23]. Marina [Ano03a]. Mark [Hol94, dPCOO^{+}05]. mark-up [dPCOO^{+}05]. Marker [ABJ90, ARM23, AOB17, AMZ^{+}24, OUZS18]. Marker-making [ABJ90]. markerless [MA15]. markers [CGW07, Dor99]. Markov [LW90]. markup [CG87]. MARS [HFT^{+}99, Pic92b]. mask [AC89, KFN23, LLP^{+}22, WWWW22, LZKJ23]. Mask-FPAN [LZKJ23]. mask-guided [KFN23]. Masked [CCCS08]. masking [BRV^{+}10, HKBA17]. MaskNet [ZWL^{+}22]. Mass [KES22, BMT96, FWCS97, LNSW16, McD06, TH09, WKW16]. mass-spring [McD06, TH09]. Massachusetts [EGL^{+}95]. Massive [vOMRI^{+}15, BGMP08, DMG20, MAG^{+}12, SBR23]. Massively [RBL95, GPR^{+}95]. Master’s [Ban97b]. Master’s-level [Ban97b]. match [MLM^{+}17]. Matching [BWdBP13, PQCT23, SE01, BWD13, BL82, BDK17, CSLY01, CBM23, DJH^{+}23, IT11, LD11, LH00, LSW15, MTM22, PGS^{+}23, RZF19, VDOK19, YF09, ZS02, ZML^{+}18]. Material [LSWL13, AALJ20, BBMK21, CZB^{+}22, FTB16, GHS06, IWT13, LAE^{+}19, NGA95, OMP^{+}18, PSBM10, PRBD22, SK16, WAM17, dBWK18]. Material-aware [LSWL13]. Materials [Ano05o, Ano06k, Ano07l, Ano07m, Lau77, ABAA22, Ano05q, Ano06j, Ano06l, BKCW14, FEJ04, IMG22, KJS18, LM22, LLHH94, MC10, SS13, WXL^{+}23]. matériels [Le 77]. Mathematical [LTS96, HMW91, LaV07, NKMI23, SS75b]. Mathematics [KS03a, Pic87a, Pic87b, Qui91]. MathPad [LaV07]. matrices [EA19, HL93, Van94]. Matrix [DH95a, FB14, FB15a, Fos87, LQ15, PPL91, PHLW15, TPN95, YY14]. matter [MA14]. matting [LS05, WXG^{+}14]. maxillofacial [hLfTxDdZ09]. maximal [GYJZ15]. maximally [LBB11]. Maximizing [BR07]. Maximum [MKG00, TT12]. Maxine [BCS08]. May [Ano03i, Ano20p, Ano21r, Ano22n, Ano23n]. maze [IU09]. maze-like [IU09]. MBKD [PDL^{+}21]. McKim [Kre93]. Meaningful [ZFJ90]. Means [NYKN83, CMLR11, HR88, WAM17]. measure [CVP^{+}16, MXK^{+}19, Sri02, SCT^{+}14, ZZJ18]. measured [WMZW22]. measurement [AALJ20, GG14, Wan04]. measurement-based [AALJ20]. measurements [KB12, PR23, ZFJ90, dBWK18]. Measures [BAPD23, BHZ^{+}21, MS01, PHO^{+}23]. Measuring [BFT23, Par75a, XCZP14, LQOW08]. Mechanical [JK84, LLP^{+}21, CPCS90, EK15, Gra83, JMV90, NKA83, PMV06, RCLM18, UWC90]. Mechanics [Tak77, Ano85g, Ano86j]. mechanism [DZD^{+}23, DSM^{+}99, MMdOE^{+}22, WCHM22].

Mechanisms [JA84a, GS01a, PCS00, VZP22]. medallions [RPAM06, RAA^{+}08]. Media [Ano01u, Ano01v, Ano01-44, Ano01-45, Ano01-49, Ano02s, Ano02-44, Ano06-37, Ano07e, Ano07d, FH11a, FH11b, Ano95-39, Ano95-40, Ano03-33, Ano04t, Ano04-36, Ano05s, Ano06-32, Ano06-34, BD17, CMSF11, CH15, ERDS14, Hvi86b, LW24, RWD14, WCH^{+}11, ZR97, ZK24, Ano01-37, Ano02-29, Ano03q, Ano06r, Ano07p, Ano04-37, Ano04-32, Ano05z, ZR97, Ano02-45]. Medial [RJKV12, BAS^{+}20, GYY^{+}23, ML12, SMM20, WKT21, ST23]. median [ZGWP16]. mediated [MTS^{+}22]. mediation [MAH00]. Medical [HKS00, PM22a, SGM97b, SPK19, BK89a, CGW07, DEST95, FGES96, GV89, HXC^{+}23, HCX^{+}23, HKHP11, KFH^{+}09, LTC^{+}20, LPL^{+}05, MA15, MGS^{+}22, MGM^{+}23, PM20, PM22b, RPHL14, ROP11, Sak02, SGM97a, SDWE99, SF92, ZKS^{+}96]. Medicine [Ano01-59, KPBR20, KKN^{+}21, PB11, SPK19, Ano04a, GPR^{+}95, PSSP96]. Medicube [BK89a]. Meditation [Web85]. Meditations [Mad00]. medium [CS03, COSEV22, Mok87]. medium-global [Mok87]. Medley [Rei10]. meeting [GMM18]. Meets [Ano01-59, Ano04a, BPS06, H^{+}00a, HH88, RCB15, RBUB94]. Meinhardt [Coo01a]. Melting [ZWQ^{+}06]. members [GBG04]. membrane [HLO16]. memoir [Nol95]. Memoriam [Adz22, Enc95a, Enc15b]. memorization [CPS^{+}22]. Memory [MCTB14, SSDS12, CYKK09, KKŽ04, NRTT95, SLS03, dGHM97, TPRC18]. Memory-optimized [MCTB14]. mental [CMLR11, Sif99, Wes94]. menus [DH07]. merge [HE80, SPGR93]. Merging [Žal01, Lym89, Que93]. Mesh [BAPD23, BA09, Fou11, Gab77, KNC11, KRA^{+}23a, LKL02a, PGL^{+}23, RCPB01, TCL^{+}15a, ZLGH10, AMHWW16, ASSF17, AF11, BjOwKM12, BCDD22, CA15, CKCK09, CMS98, DN22, FWX^{+}18, GYY^{+}23, GAÖ02, GTFB23, HKL^{+}23, HEK22, JGA09, KPA12, KYKK19, KRA^{+}23b, LY15, LBD17, LDLD23, LZP^{+}04, fLhLfT11, LWLT11, hLfTxDdZ09, LZT^{+}24, LCCM02, LTBZ13, LZL^{+}15, LLS^{+}16, LCW^{+}18, LCGD24, MAC19, MYF06, MVS14, MRSS^{+}18, MFL11, MGJ^{+}11, MKKM18, PPD22, SMS09a, SMS09b, SA86, SA87, TSD^{+}07, TAF16, WLDB11, WY11b, WTM12, WLL22, XhKKL16, XGZ11, XLXG11, XCZ^{+}16, XCW^{+}09, XGC18, YPSZ01, YWH^{+}16, ZD04, ZJSB22, ZYX^{+}19, ZQL15, dAU14, dCdLL14]. Mesh-aware [KNC11]. mesh-based [dAU14]. Mesh-Invariant [BAPD23]. mesh-to-points [BCDD22]. meshed [DFWW15, ZSK95]. Meshes [AYZ12, PF16, TJ85, ACG15, ASC17, AXC22, AK15, ALM19, BMW12, BGMP08, BFRA11, BAS^{+}15, BS01b, BAC14, CPM19, CCL^{+}03, CC08, FJ17, GY19, GCC23, GLA23, Hop98, HXA^{+}12, KP19, KNC11, KCS24, KYM12, LGRP14, LAL11, LDLD22, LBLV16, LLX^{+}15, MCAH12, MGH13, MLP19, MMS15, NC12, NAO13, OP13, OP15, PS13a, PT16, PMV06, PB96, RP20, Rei98, RA15a, RMG15, RBF20, SXG^{+}09, SBSL04, SCC11, TB19, TT12, VGP04, VMAL16, VH15, WFG03, WCLZ14, ZY01, ZZCY22, ZSL12, ZCT95, dF24, dGGV08, vTTK^{+}20]. Meshing [TGG06, DLS^{+}11, GVPN09, LCZ^{+}11, LCA19, Liv23, MG98, SS12a]. Meshless [LLZ^{+}11, XCXC23, SOG08, ZLGH10]. message [Ela86]. Meta [GOdSC23, LM22, Sif99]. Meta-learning [GOdSC23]. meta-materials [LM22]. meta-model [Sif99]. metaball [YJC99]. metaballs [JLP00]. metabolite [GVC^{+}20]. Metabolomics [HHZ^{+}22]. Metafile [NM85, Ree85, Sch86b]. Metafiles [Row86]. metal [MMV88]. metal-forming [MMV88]. metallic [KK22]. metamorphosis [Bow95b, KP98, SFVP13]. metaphor [GS05, Shi04]. meteorological [RŞH^{+}22, UL22]. Meteorology [H^{+}00a]. Method [Bar77, CGH94, CS85, Lam87, NYKN83, Uga06, Wal93b, AF16a, ASPO15, And82, Bap99, BCG^{+}94, BW92, BAC14, CXGL23, Car99b, Car93, CM06, CP96, CM15a, CYWM23, Cla90, CPLB14, CS80b, DWH^{+}15, FWCS97, FWX^{+}18, FB14, FB15a, FIC23, FV13, FEBS07, GH98, Gil94, GLDK95, HJW^{+}08, HLB^{+}06, HY93a, IP23, JKK02, JK15, JYL17, JZY^{+}23, Jon91, JJPP^{+}22, KB06, KCL18, KGK^{+}07, Lam00, Les01, LW89, LUB^{+}13, MYL^{+}23, MPL02, MF02, MR90b, OCCZ12, PMZS97, PVM^{+}22, PTY^{+}16, PLFT21, PCPW15, Pur87b, QGGW97, Ree91, Ree92, RH85, Rei98, SGGC05, SKS09, SDS89, Spr04, Ste09, Szy90a, TNU^{+}01, TRLX22, TDR01, VRV05, VH15, WW08, WLP^{+}14, WTWT18, WZLQ19, WPB^{+}23, Wan23, WLL^{+}24, WAM17, XPL90, XM12, XMD^{+}12, XWW^{+}19, YLYJ13, YHHS93, ZK95, ZK07, ZDL^{+}24, ZM91, ZM92, ZYW^{+}21, ZCT95, ZLLY06]. méthodes [CV77]. Methodologies [Cot75a]. methodology [HLCF88, JCFN18, LCCM02, MTTL82, VJ06]. Methods [Ala85, Ent89b, AMFH21, AAAN23, AS22, ABM^{+}06, BMMZ23, Cad08, CWNA08, Car96, Car99a, CRT04, DMT03, FAZ21, GLDH97, GANM21, GOZ95, HN20, JT02, KHTM17, KP95, LB75, LSW12, MLP19, MLCMGR23, MCG^{+}23, Per02, RCBS10, RCG^{+}05, RNM95, ST20, SS89a, SF91, UPT97, Wal94b, WMRA^{+}15, WPJP23, XLW^{+}24, ZTAP21]. metric [LCG19, Lod21, ZZH^{+}24, ZZDZ10]. metrics [LTBZ13, TSD^{+}07, YRS^{+}18, vTTK^{+}20]. Metropolis [APSS01, SKS17]. MFS [LXW^{+}10]. MHz [MDJ^{+}95]. MIC [XWW^{+}18]. MICCAI [SPK19]. Micro [Ano85f, APA^{+}11, KK21]. micro-CT [APA^{+}11]. micro-mirror [KK21]. microblog [CTJ^{+}14]. Microcomputer [Ade85, ML79, Ano91c]. microcontrollers [Nar15]. microcosm [BSB^{+}23]. microfacet [CXXW20, SP23]. Micrographics [Ade86]. Microprocessor [Bas77, PR82b]. Microprocessor-Based [Bas77]. microscope [RPFC01]. microscopy [WHH21]. Microstructure [BDSP22]. Microstructure-based [BDSP22]. microstructures [LLP^{+}21]. microvascular [RHBS95]. Mid [CS18, CPC^{+}18, KK21, OK20, VR16]. Mid-air [CS18, CPC^{+}18, KK21, OK20, VR16]. MIDAS [War76]. midgets [PFR94]. MIG [RS22]. Mill [MNS^{+}19]. mimicking [AWI^{+}09]. Mind [LAB^{+}14]. mine [dDH87]. minimal [BSPD10, EME15, KS09, Las91, MPAC^{+}23, MCS13]. minimalistic [CPM19]. minimax [MW92]. minimisation [GR93]. Minimization [BG79, CZL14, CFZL16, KP98, LH14, LJH18, PMV06, WY11a]. minimized [Hor82]. Minimizing [OMW13, HKHP11]. minimum [CCKW11, IU09]. Mining [SA04, BPGW11, BPKG07, DW89, HMHA98, KPSN04, LPPM07, PSMD14, SL16a, VT07]. MiniVR [FBT04]. Minkowski [Gho93, SM98]. MIP [KJ08]. mipmap [SK99]. Mira [MTT82]. Mirage [Ano03-39]. mirror [Ger97, KK21, OK20, PCV16]. mis [GTG06]. mis-labeled [GTG06]. mis-segmented [GTG06]. misconceptions [RBLB21]. MISD [SR95]. misfit [LK18]. missing [CR07, CK02b, Spr04]. Mitigating [KAVM23, WSJJ24]. mitosis [Stu91]. Mixed [BES01, BNS24, MCK12, PGS^{+}23, Ano03-36, Ano04-39, Hor83, KSH^{+}19, KTMW12, MCV18, MCMV22, Mor76, PCY^{+}06, PMBS14, QLF^{+}09]. Mixed-element [MCK12]. mixed-point [Mor76]. mixing [VP98]. mixture [ZHG^{+}21]. mixup [LLZ^{+}23]. MLS [BGPT18]. MMCN [Ano04-58, Ano04-57].

MMCN02 [Ano01-50]. MMCN’07 [Ano06-27, Ano06-38]. MMM [Ano06-37]. MMV [HK93]. Mobile [Ano95-28, ABK11, BSGT03, dMOHKO18, RBP96, ABAA22, Bad96, BZYB21, BR96, BBDM03, BG01, BL96, CYKK09, DGBNV18, DR09, FN99, HZ15, HFT^{+}99, H^{+}01, HCC13, JHL^{+}12, KHK18, KR96, KS01b, LMD96, MLCMGR23, MCKS06, MML^{+}11, NSL16, NGAS23, PvdSLJ99, RV01, Rau06, RLT16, RLU^{+}19, RJS98, SKSZ99, SH96, ZDL^{+}19]. mobility [OO04]. Möbius [FS20]. mocap [PHLW15]. Mockup [DCJH13]. Modal [KLL^{+}15, Bus97, CMLH21, CK02b, CLH^{+}23, PRW^{+}22, ZPN^{+}21, ZPIS23, ZHW^{+}21]. modalities [SHOC23]. Mode [WHZ^{+}18, ZSL08, BRHB20, WLX^{+}23]. Model [CVB16, GM86, MDM^{+}21, MZCD21, Nav89, NFHS06, PACSG^{+}23, WP77, AALJ20, BG88, BTS19, BDM^{+}16, CC04, CGM91, CXXW20, CB97, CWTL08, CMD99, COM^{+}94, CSHZ04, CSLY01, CZ06, Con91, DBLC02, DJC^{+}23, DFWW15, DLR^{+}10, Fri03, GM02, GCDL22, GTdS^{+}18, GK96, GA83b, Gro92, Güd97, HBOS13, HMdM^{+}95, HKYM01, HY23, HYP^{+}24, IC96, ILLC01, ID17, ICNV21, IA83, JZLP23, JZX^{+}21, JL23, JHL^{+}12, KFN23, KB05, KR96, KN88, KL02, LM22, LST96, LDG96, LSCJ23, LHL23, LPV92, LX08, MKDM22, MLC^{+}22, MBA20, MGH22, MBC^{+}23, MKKM18, NP96, NR07a, Par75b, PSK^{+}11, PMV06, PQ10, Rok93b, SPT18, Shi93d, Sif99, SJZ^{+}23, SGZ10, SYMW21, TH90, THL15, VA96, VCQ92, Wal06, WZ09, WLZL18, WXZ^{+}18, WCHM22, WJGG15, Wil75, WI00, Wu89, Wüt98, XLXG11, XLGS16, Yam94, YJC99, ZIP^{+}19, ZK98, ZZJ^{+}14, ZCZ^{+}18]. Model-based [NFHS06, WI00]. modeled [AJAC23, BBP13]. Modeler [CS16]. modèles [CF77]. Modeling [Ano01-60, Ano02a, Ano02-36, Ano02-37, Ano02-50, Ano03-52, Ano03-51, Ano06-37, Ano06-52, Ano12s, Ano13k, Ano15k, ALP17, BLZ22, BMP22, BPH20, BLM23, Elb11, EBC^{+}15, FHM98, GCLZ16, GHK18, Gro94, Her85, Lib91, MW99, MPSB21, MY97, Now86, SH12, SKSI95, SJ09, Wit84, XLL^{+}18, YYY19, ZD04, ZBM00, ZCL^{+}18, ZXLF23, AKPS00, ACG15, ASC17, Ano03-60, Ano04-30, Ano04-55, Ano04-54, Ano04-61, Ano05d, Ano06-34, Ano13l, BDPR93, BAD23, BX99, BBCG11, BBE14, BAS^{+}15, CGB13, CYJ^{+}13, CT13, CSX^{+}19, CCM^{+}18, CCM^{+}11, CPCS90, CS18, CLX^{+}19, CB01, DCJH13, DVF06, Des00, Doh95, EHSF17, EHBM82, Ela86, FBH^{+}01, FST97, Fun99, GWP00, GDDA13, GBF14, GA12, GO90, HOCN07, HQW14, HLZ^{+}17, HZLQ20, HJT^{+}13, HL06, JCFN18, KSM07, KRK^{+}06, KF02, KLW12, KKNT88, KK16, KK12, KYT^{+}17, KKH23, KEVD18]. modeling [LLLC11, LG94, LCCM02, LXPP06, LTPN96, MT88, MZPZ16, MWLZ22, MRR98, MP19, MT00, MYC15, MCP^{+}22, MBP14, NS87, NC12, OSSJ09, PPP88, PTL04, Pie90, PWV^{+}18, PP02, PM13, QD03, RMD11, RLD^{+}12, RFZ^{+}17, RLB^{+}02, RLU^{+}19, RMSC11, SVV23, SH94, SKS09, SLX^{+}16, SS96, Shi06, SGR^{+}99, SWS75, SET^{+}88, STdKB11, SJZ^{+}23, SEKA19, TKM^{+}24, VPBY02, WBA16, WWY06, WHL^{+}09, WZLQ19, WMFR89, Wei84b, WKW16, XZ00, YMYI11, YLT^{+}18, YMYH12, YH21, ZjLW^{+}14, ZLM^{+}15, ZZC^{+}20b, ZSS20, ZGZS22, ZPL^{+}15, dlLC99, tHV10, vRESH16, Ano07a, Ano07b]. modélisation [Nan77]. Modelling [All84, Ano95q, Ano95r, Ano95s, Ano01-54, Ano01-57, B^{+}00, CP19b, FHP89, Kle86, MPS85, PMPR^{+}16, Wen84, vEB98, Ano03-53, Ano05-40, BBH90, CSFG96, CSZ92, CCCP04, DAHF04, DMS08, FS86a, Fal89, FR98, GdMF03, GT91, Her83b, HY03, JK90, JDGS88, K^{+}00a, KH86, LTS96, LS07, LS06, LR90, MQW09, NGA95, PCS00, Pla00, PM10, RBUB94, STM^{+}04, SL12b, SPS96, Ste09, WM95, YT83, Ano93e]. Models [BP93, BKL^{+}95, DM79, GÖT93, IO91, SKO83, Web84, XLW^{+}24, ASPO15, ASS22, AGM^{+}21, AVM05, AB11, BMH99, Bec94, BA90, CA17, CCCS08, CKS98, CC20, CB10, CDGC94, CMDS17, DA18, DGBNV18, Dol95, DLZY14, EHBM82, Elb11, FIC21, FIC23, GM04, GF09, GKLM07, GÖT97, GYL^{+}13, GK04, HMA23, HAL^{+}21, HFP06, HHKF10, HCLC16, HK99, IVCN20, JCFN18, KPFT03, KM14, Kuo01, LLH17, LD03, Lee01, LYW^{+}10, LNSW16, LZYQ22, LCCM02, LJWcH07, LXCW18, LUB^{+}13, LVVC06, MHW10, MAG^{+}12, ME92, MCT05, MAF19, MPS06, MKM19, OVWK16, PA07, PSBM10, PVCM22, PSS04, RC08, RVdF08, Ros90, SBR23, SK98, SGC00, SD08, SP23, STW82, SH03, Sif99, SC97, STBG19, SZL^{+}13, TNFG14, US20, VZP22, WLL^{+}12, WWO^{+}23, WXL^{+}23, Wes94, XM12, YK18, ZK07, ZYX^{+}19, ZPP^{+}23, van89a]. modern [APA^{+}11, FWW13, Mad00, UKW23]. Modification [GWW84, ÖT21, SDS89, van89a]. Modified [AnD19, CKCK09, SZEG93, WW08, XLSW22]. modular [CCW97, CCC00, CCW01, KS15, LP93, Sch85a, SEKA19, YAKE23]. modulated [Sen99]. modulation [APS09, HLO16]. module [HHN^{+}23, II22, JXJ22, LDS^{+}21, WCLT21, WZZ^{+}21]. modules [NLdAL^{+}23, Wol02b]. Moebius [FSV17]. MOGRA [RG93]. Moiré [Sen00, Rag09, YT87]. molds [HLCF88, JVS^{+}24, SJG19]. Molecular [BKS23, HJW^{+}08, HvK87, RG93, WMFR89, BKS21, EHBM82, HGVV16, SLK^{+}23, VFSL06, ZS14, BKS23, SFS^{+}21]. molecule [HSD96]. Molecules [CHMR78]. Moliverse [BSB^{+}23]. mollusc [vEB98]. MolVA [BKS21, BKS23]. Moment [TSD^{+}07]. Moment-based [TSD^{+}07]. moments [WLDB11]. Mona [Sch95]. Monastery [ACB12]. monitor [FP89]. monitoring [Han84, MAM^{+}24]. monocular [ET18, HZ15, MPQG18, PCKB23, WWK11, ZYW23]. monoid [PDS21]. monopole [SHK18]. monotonic [DBG92, SAMA97, Sar00, Sar02, Sar03]. monotonicity [SY23]. montage [TR95]. Monte [BSB^{+}02, JC95, PTW98, SB86, Shi92, WLZL18]. Morphing [Sch95, CJT96, FBH^{+}01, GS01b, LBTM15, Mar04, PT16, YF09]. morphologic [WCLZ14]. Morphological [SFVP13, WSK^{+}22]. Morphology [CG93, NKMI23, YHNC22]. Morphosis [Eng93]. Morse [CDI12, CDIM16, DS18, FLV20, IFD15]. mosaics [BX99]. Most [Ano11c, Ano13c, GL83]. MoSuMo [GD11]. Motif [Jer95]. motifs [CGH94, MRG^{+}19]. Motif^{T M} [STN95]. Motion [LYZ^{+}24, MK85, OP19, SNB17, YMZ^{+}15, ARM23, AAAN23, AOB17, Ano94u, Ano94v, Ano06-35, BBHC15, BWBM20, BPKB15, BHTT94, CHPS20, CSP19, CFB15, CPG94, CS06, ESAH02, GG14, GSKZ22, HHCM17, ILLC01, JZY^{+}23, JL23, JRZ^{+}23, KRRS12, LLLZ16, LZZ^{+}19, LGZ^{+}21, LWJ^{+}22, LDG96, Lin97, LKC94, LB12, LPV92, MTS21, MSHL22, MA17, ND23, PF89, POBB09, RAF21, RCLM18, SRA^{+}19, SYMW21, TH09, UOT83, VR16, WLP^{+}14, Wil75, WMDR08, YSDG24, ZLL^{+}23b]. motion-JPEG [CPG94]. motions [FCG^{+}21, FO21]. motivated [CFB15]. Motor [BMdSVR18, CIT^{+}19, FvdPT01, HGH^{+}18, RHM^{+}12, SL18]. motorcycle [GLA23]. mounted [CPS^{+}22, CP19a, IKM^{+}20]. mouse [CH12, KY97]. movable [Güd97, OK20]. move [Car92]. movement [BTC94, BSC^{+}21, PQ10, PRM^{+}24]. movements [DBS^{+}18, FGL23]. movie [CLH12, FP75]. movies [JM88]. Moving [EPS96, Gon98, Jut94, KB05, PDL^{+}21, RCM^{+}20, RGRG15, ZMM^{+}90]. MP [GS89]. MPEG [DJG^{+}04, HEG98, QN98]. MPEG-4 [HEG98]. mpLBP [TBDC20]. MPVConv [ZZH^{+}23]. MR [EKG06, GVC^{+}20, PPN^{+}23]. MR-Net [PPN^{+}23]. MRF [SXW^{+}22]. MRF-based [SXW^{+}22]. MRI [CJAR21, GPTP10, KGGP19, LSCJ23]. MSOKS [JK90]. MSRS [GRPR08]. Multi [Ano95-39, Ano95-40, Ano06-37, DS93, EL22, FTB16, GCYX23, HWR^{+}23, JCT^{+}15, JD75, JYC^{+}23, JWZ23, LPL^{+}20, LZW^{+}21a, LLW^{+}23, LRG11, LLZ^{+}23, MDSU88, MBN21, Moh77, PBK13, RP20, RVR04, SV18, SVSV20, SXW^{+}22, TF18, TKS11, VHON04, WHZ^{+}18, XZL^{+}22, ZPN^{+}21, ZZJ^{+}14, dLvL00, ALC06, AK13, Ano06-34, ASZ^{+}14, AP99, BWD13, BYQZ22, BES00, BRPC18, Bus97, BBP10, CMLH21, CH93, CF11, CXT18, CDI12, CS06, CLX^{+}19, CLH^{+}23, DCLB19, DG06, DKM^{+}20, E^{+}00, GJN^{+}21, GVVJ99, GZL21, GS04, HKCL02, HE15, HXH24, JXJ22, KHTM17, KP15, KP22a, KK16, KL02, KMWW^{+}14, LBD17, LWW^{+}20, LZW^{+}21b, LFY^{+}21, LCCM02, LHY23, LCGD24, MRF06, MVS14, NKMI23, OP13, PP16, PSK^{+}11, PRW^{+}22, PKRM21, PRBD22, QSXT22, RZY^{+}20, Ran87b, SVV23, SSH16, SGS99, SK16, SLL^{+}23, SCCS13, SIE14, Tap06, TPB08, WTL^{+}11, WRLZ23]. multi [WCL23, WLX^{+}23, WAM17, XZPG21, YS23, YK18, ZZCY22, FS78, ZZH^{+}23]. Multi-agent [RP20]. Multi-Axis [MBN21, CXT18]. Multi-block [DS93, SGS99]. multi-body [CS06, WAM17]. multi-bounce [SSH16]. Multi-branched [ZZJ^{+}14]. Multi-camera [LRG11]. Multi-channel [HWR^{+}23, VHON04]. multi-chart [MVS14, ZZCY22]. Multi-console [JD75]. multi-core [TPB08]. multi-covering [SIE14]. multi-criteria [DKM^{+}20]. Multi-Curve [Moh77]. Multi-dimensional [EL22, E^{+}00, SCCS13]. Multi-domain [LPL^{+}20, LLZ^{+}23]. Multi-exemplar [PBK13]. multi-exposure [KHTM17]. multi-field [BBP10]. multi-finger [AK13]. multi-front [PSK^{+}11]. multi-graph [DG06]. Multi-grid [JWZ23]. multi-image [YS23]. multi-L-REP [MRF06]. multi-label [NKMI23]. Multi-layer [JCT^{+}15, HXH24, WTL^{+}11]. multi-layered [BES00, OP13]. Multi-level [dLvL00, HE15, YK18]. Multi-material [FTB16, SK16]. Multi-media [Ano95-39, Ano95-40]. Multi-modal [ZPN^{+}21, Bus97, CMLH21, CLH^{+}23, PRW^{+}22]. multi-mode [WLX^{+}23]. multi-module [JXJ22]. multi-object [LWW^{+}20]. multi-page [GJN^{+}21]. multi-patch [PKRM21]. Multi-perspective [TKS11]. Multi-Pierre [FS78]. multi-point [KK16]. multi-polygon [Ran87b]. Multi-purpose [MDSU88]. multi-rate [KK16]. multi-region [LCCM02]. Multi-Resolution [GCYX23, LZW^{+}21a, BWD13, CF11, CDI12, GS04, KL02]. multi-ring [LCGD24]. Multi-scale [TF18, WHZ^{+}18, XZL^{+}22, ALC06, ASZ^{+}14, BYQZ22, PP16, SLL^{+}23, WCL23]. multi-sensory [HKCL02]. Multi-sided [SV18, SVSV20, KP15, KP22a, SVV23]. multi-space [GZL21]. multi-stage [BRPC18]. Multi-threading [RVR04]. Multi-user [LLW^{+}23, AP99, CH93, KMWW^{+}14]. multi-variate [Tap06]. Multi-View [JYC^{+}23, SXW^{+}22, CLX^{+}19, DCLB19, GVVJ99, LBD17, LZW^{+}21b, LFY^{+}21, LHY23, PRBD22, QSXT22, RZY^{+}20, WRLZ23, XZPG21]. multiagent [CWGR01]. multibody [KW96]. multichannel [CMLR11]. multicolor [RPFC01]. multidimensional [AOL96, CVP^{+}16, PTR^{+}92, Sah15, TZvD^{+}21, dSB04]. multidisplay [BB03]. multigrid [GW06, WMZW22, WMRA^{+}15]. multilayer [HKYM01, HZ18]. Multilevel [JA84c, PN83, GRW00, MTTL82, ZY01]. MULTIMEDIA [Ano02b, Ano93b, Ano94c, Ano94y, Ano94x, Ano95u, Ano95a, Ano95b, Ano95c, Ano95-41, Ano02c, Ano02-41, Ano02-38, Ano03e, Ano04f, Ano06-27, Ano06b, Ano07k, Ano07p, NSG05, ABCD93, Ano94-29, Ano94-30, Ano01g, Ano01c, Ano03b, Ano04e, Ano04-58, BEFV94, CYKK09, Gir93, GV07, HK04, HK93, KS01b, LZSG03, LSS97, PP16, SKSZ99, VHON04, Ano07j, Ano01-50, Ano04-58, Ano06-38]. multimedia-appliances [HK04]. Multimodal [AT08, ST22, SKR15, VW21, HKS01, LSW^{+}23, RKMP13, RHN03, ST20, SLM^{+}22, SL16b, ZZJ18]. Multimodality [ZKS^{+}96]. multiparametric [Dal00]. Multipass [dSEM19]. multiphase [ZWQ^{+}06]. Multiple [ASKCK03, ARL78, Lea85, BMR23, BH15, Bro07d, CP10, DKY97, Elb22, GL83, Gil94, IKTS22, KLW12, LW24, LWW10, MC10, McC95, MHYN23, MVG^{+}21, PSS04, RCBS10, RPAM06, RAA^{+}08, SSM11a, Sta87b, SBR^{+}22, TZvD^{+}21, ZQ12]. multiple-spiral [RPAM06, RAA^{+}08]. multiplexed [dBWK18]. multiply [PR93]. multiprocessor [ZW88]. Multiresolution [Ger02, Lam99, OSB07, PPN^{+}23, Rei98, RLB^{+}02, AOL96, GM04, HS99, KPFT03, KNC11, MVS14, MCT05, MRS17, PR97, SK98, WSG10, YPSZ01]. Multiresolutions [BS11]. Multiscale [AS19, HCX^{+}23, LKLW16, BNPS10, ID17, WJGG15]. multiselection [JALS03]. multisensory [MGMB22]. Multitask [HTKV84, ZLLG18]. Multivariate [YBTB23, PMBS14, ROP11, SS12b, SVVS^{+}17, SHD^{+}17, TW24, THQ^{+}16, YWR03, YHW23, dSB04]. Multiview [DSR11]. Mumford [WLL22]. mundane [CS01]. mural [MPAC^{+}23]. murals [RM22]. Murray [Pha95]. muscle [DGKK20]. muscles [RRC^{+}18, SH94]. Musculoskeletal [NGAS23]. Museum [ACB12, GPC^{+}17, McW88a, OFP^{+}11]. Music [Ano02-55, Ano04c, GS84, Ano01w, KPL08, Ano01-46, Ano01-47]. Musical [CS80a, FM99, GC86]. Mutual [QSLS23, CSCF08, KKO01]. MVE [FLM^{+}15]. MVSNet [LHY23]. MWFormer [PGL^{+}23]. myocardial [SRA^{+}19].

n [FB11]. n-loops [FB11]. nanometric [BAS^{+}20]. Nanyang [Ano03a]. Narrative [MGS^{+}22, dSC07, MGM^{+}23, SS02]. Narrow [GPTB02, CZZ22]. Narrow-band [GPTB02]. Narrowband [SBOT78]. National [ACB12]. Natural [BN06, BKCS79, GM86, MAG^{+}12, MPS85, Ano05m, Des00, FB12, GBF14, HL96, JS92, LS05, Ste09, VRV05, dMTB^{+}21]. naturalness [WZC^{+}21]. nature [COM^{+}94, KRP^{+}91]. Nav [RFB23]. NAVER [PKK03]. navigable [DSG21]. Navigating [dSEM19]. Navigation [NRTT95, PF16, SW91, BTS19, BG01, BJS01, DBW^{+}12, FHM98, HK99, KA22, KCU^{+}22, MAG^{+}12, MMGB17, MCM^{+}18, OP13, OP15, RP20, SVNB99, SAB12, SDWE99, dAPG18, vTTK^{+}20, vTP20]. navigational [ZI00]. NBS [ZSW08]. Near [OP13, BW98, HPKE19, KYKK19, LSW15, RKG22]. near-isometric [LSW15]. near-lossless [BW98]. near-optimal [RKG22]. near-singular [HPKE19]. near-zero [KYKK19]. nearest [ATZM19, Li89, dGMW16, SSV07]. nearest-neighbor [ATZM19]. nearly [ITW18]. NeatSankey [XZY^{+}23]. NEDLAN [Mar76]. needs [GHFH08]. Negative [Pie89, QMHH91]. neglect [WBJ^{+}21]. neighbor [ATZM19, dGMW16, Sam89, SSV07]. neighborhood [ANGH11, LCXL20, MX14, QD03, RA03]. neighbors [Li89]. Neo [TTKA23]. Neo-Hookean [TTKA23]. NEOGEN [OP13]. Nerve [SMG77]. NeRVI [GCW23]. ness [WSJJ24]. Nested [OCR^{+}19, AG94, GRW00, KB20, LTC^{+}20, SCCS13]. nesting [SB86, Seg88]. NET [Ano01-63, CDGC94, Spr98, Wil03, ZYW23, AAK^{+}22b, CGWZ22, DZD^{+}23, HG21, LTC^{+}20, LSCJ23, LCX^{+}23, PPN^{+}23]. net-delay [CDGC94]. Nets [KS86, MRR98]. Network [Ano07-33, GCYX23, LLR93, LHH^{+}21, LCX^{+}23, ME77, ZXH^{+}12, AHK03, BF02a, CMLH21, CKM^{+}23, CLN^{+}16, CSK97, CT75, Cot75b, DZD^{+}23, DK92, FP87, FCG^{+}21, FK11, GWW^{+}22, GLZ^{+}21, GK04, HCX^{+}23, HDZR23, HWEB22, HWR^{+}23, HG22, HZ23, HXH24, HZLC22, HJL^{+}93, IMMS82, JDT^{+}23, KKHS03, LBD17, LZW^{+}21b, LF22, LPP^{+}19, LZL^{+}24, MYL^{+}23, Mar76, Mar80, Mar82, NZZ^{+}21, PBH19, PRW^{+}22, PLFT21, QSLS23, QL23, SSQL24, SLL^{+}23, WZL^{+}22, WRLZ23, WLZ^{+}23, WXC^{+}23, WWCZ19, WZZ^{+}21, WM24, XWWK21, XZL^{+}22, ZDL^{+}19, ZLLL21, ZZF^{+}23, ZLL^{+}23a, ZLL^{+}23b, ZMH^{+}23, ZHG^{+}21, dGGDV11, Sch86a]. network-accelerated [LF22]. network-attached [KKHS03]. network-based [FK11, QL23]. Network-Oriented [ME77]. networked [HW89, PZH^{+}05, S^{+}01, PKK03]. Networking [Ano93b, Ano01-50, Ano06-27, Ano06-38, Ano04-58]. Networks [MDM^{+}21, Mar80, Now86, BGLA18, BC01, BRPC18, CYW15, CMLR11, Egl86b, HIS83, KMV^{+}18, LCXL20, Mar76, Mar82, MCMV22, OBD^{+}23, OK12, PPN^{+}23, RM22, Roj94, RHBS95, SM99, SPT18, SHBSS16, SHBSS17, WGS^{+}18c, WGS^{+}18a, WGS^{+}18b, WWH^{+}21, YSDG24, ZYW23, dSMBG23]. NeuBTF [RPKLMG23]. Neural [CYWM23, CD23, FK11, HGJC21, LXY22, MDM^{+}21, RPKLMG23, TACS22, TKM^{+}24, WCW^{+}24, BF02a, BAS^{+}20, BC01, BRPC18, GCW23, GCRN23, HZLC22, LBD17, LF22, MZ23, MCMV22, MVK^{+}22, NSS^{+}22, OBD^{+}23, PPN^{+}23, Roj94, SPT18, Spr98, TW24, VH15, WZL^{+}22, WCF^{+}23, WWCZ19, XWWK21, ZLL^{+}23a, ZYW23]. neurally [TFF^{+}20]. neurally-guided [TFF^{+}20]. neurodegenerative [CMLR11]. neurosurgery [H^{+}00b]. neuroticism [DAG22]. Newton [Car99a, Car99b, CZC02, Gil94, JKK02, LdSP^{+}08, MR90b, Wal93b, Wal94b, YHHS93]. Next [PJP23, Li89, Mad95, Pea02]. Nibble [MG98]. Nil [NdSV20b]. NN [AAK^{+}22b]. No [WLX^{+}23, Lod21, QWC14, WW08]. No-reference [WLX^{+}23, Lod21]. no-search [WW08]. Node [Bar77, LPP^{+}19, XZY^{+}23]. nodes [IA91]. nodule [YLH^{+}18]. noise [BM08, CGG19, LH14, LYZ15, SJT20, Smo03, TDR^{+}17, XHGL12, XXHM21, YIL09, ZGZ^{+}16]. noisy [DRFRD06, GS11, MTSM10]. Nomadic [KPB96]. Non [Ano12s, HQW14, LYZ^{+}22, LMJH^{+}11, Mac78, MVCNI21, RGMJ22, SB77, Van09, WH96, YSD13, AOB17, AVM05, BLNZ22, Bay95, BFSE03, BCC20, CDW11, CLT07, CST05, CBM23, DS93, DYW^{+}22, DLR^{+}20, GBF14, Gin02, GS11, LKHM19, MG98, MML12, NdSV20a, PMTK01, She12, SNS06a, TSK98, WA02, ZLL^{+}23a, ZSL12, CI11, SCMT91, Var92]. non-computer [She12]. non-cooperative [DYW^{+}22]. non-distributive [Gin02]. non-Euclidean [NdSV20a, SCMT91, Var92]. non-general [LKHM19]. Non-homogeneous [LYZ^{+}22]. Non-Interactive [Mac78, BCC20]. Non-invasive [WH96]. non-isometric [DLR^{+}20]. Non-line-of-sight [RGMJ22]. Non-linear [Van09, BFSE03]. non-manifold [CDW11, MG98]. non-natural [GBF14]. Non-overlapping [MVCNI21, WA02]. Non-parametric [HQW14, SNS06a]. Non-Photorealistic [Ano12s, LMJH^{+}11, AVM05, CLT07, CST05, MML12, CI11]. non-planar [Bay95]. non-rectangular [DS93]. non-rigid [AOB17, CBM23, ZSL12]. non-self-intersecting [PMTK01]. non-sequential [TSK98]. Non-Speaking [SB77]. non-stationary [BLNZ22]. Non-uniform [YSD13, ZLL^{+}23a]. non-uniformly [GS11]. nonconvex [SA86, WLL22]. noniterative [Jas88]. nonlinear [FA94, G^{+}01, Gro94, McD06, Zay12]. nonlinearity [PSBM10, Szy89b]. nonnegative [PHLW15]. nonsmooth [WLL22]. nontrivial [VBW12]. nonuniform [PLFT21]. norm [KKL02]. normal [BRHB20, CST05, DCLB19, GS11, LSK^{+}10, LZC^{+}15, LCW^{+}18, MFP11, NZL^{+}21, SPS12a, SHBSS16, WM95, XLGS16, XXT18, YRS^{+}18, ZCL^{+}13, ZQ12]. normalization [HWYL21, Kni95b, ZPN^{+}21]. Normalized [TMP07, Bru14, RNM^{+}19]. Normals [FM09]. Norte [FCSB90]. NOSSDAV [Ano07-33]. Notation [BSW78, CS80a]. Note [Ano10i, Ano22i, Ano23j, Ano23l, Ano24g, Chr78, CN05, FP89, Hoo91, Jor18a, Jor18b, Jor18d, Jor18f, Jor18g, Jor19a, Jor19b, Jor19c, Jor19d, Jor19e, Jor19f, Jor19g, Jor19h, Jor20a, Jor20b, Jor20c, Jor20d, Jor20f, Jor20g, Jor20h, Jor21b, Jor21c, Jor21d, Jor21e, Jor21f, Jor21h, Jor22a, Jor22b, Jor22c, Jor22e, Jor22f, Jor22g, Jor23a, Jor23b, Jor23c, Jor23d, Jor23f, Jor23e, Ano12k, Ano24f, BTV83, Coo07, Dal00, Jor18e, Jor20e, Jor21g, Jor22d, MW92, Pic88a, Pic89, Pra92, Ran87b, Ree91, VHE10, Ano06-45]. Notes [Nug91, DPS10, ZW88]. notice [Ano07-40]. Notions [LS18]. NOVAction23 [TC24]. novel [AAK^{+}22b, AM19, CXGL23, CGG^{+}20, CTN^{+}17, FWX^{+}18, FBTT^{+}22, GCCZ14, HY23, KONS17, KCS22, KCS24, LWFZ23, MMS^{+}17, MBV18, OFP^{+}11, PP16, PHO^{+}23, Shi04, Wan23, WCL23, XD08, ZYW^{+}21, ZZXT18]. novelty [Dia94]. November [Ano19j, Ano20j, Ano21l, Ano22s, Ano23s]. novice [AMZ^{+}24, RBLB21]. NPAR [Ano12s, CI11]. NPSNET [FZPM93]. NPSOFF [ZWP^{+}93]. nteractive [Mon87]. Nu [Sar04]. nuclear [ATAG^{+}21]. Number [Ano03j, Bow95a, MW12, SKP99]. numbers [FN10, Lak89, Szy92]. numeric [SK03]. Numerical [SP23, CGM91, Cor76, LB75, SLQ^{+}19, Szy89a]. numerically [BS11]. nurbs [Sar92a, AS19, BGK04, HY03, IC96, MCK12, PL97, PX06, QGGW97, WZT97, Yin04, ZFS03]. NURCC [MAC19].

O [AAK^{+}22b, CGW07]. O-Buffer [CGW07]. O-Net [AAK^{+}22b]. OBB [CK09]. OBB-based [CK09]. Object [BTD^{+}22, DSM^{+}99, FGZ^{+}22, KAÖ98, LPD^{+}18, STP^{+}20, She88, STN95, TT19, WLZL18, Wis86, ABAA22, AMR23, Ano13o, AA13, BFLP20, BDL^{+}22, CSFG96, CEG18, CH91, CMD99, CSLY01, Cze90, Day92, DJC^{+}23, FDGM18, FVG15, FFP^{+}21, GMd^{+}13, GBD88, HMA23, HWSW19, HOCN07, HJL^{+}93, IR06, JSV98, JRS21, dSJdML18, Kas87, KLW12, KP98, LDS^{+}21, Les02, LL04, LL05, LL06, LTS88, LWW^{+}20, LSZQ21, LPO20, OMGGG^{+}19, PLVT23, PPSS96, QD03, QKS01, SBS19, SS93, SS02, SJ09, SSS90, SO75, SCFF16, ST23, UPTd92, VV89, WZZZ18, WLB^{+}23, WPJP23, YZ17, ZTAP21, ZS94, ZLLG18, ZMK18, ZHW^{+}21, ZWP^{+}93, VTW23, vBT20a, vBT21]. object-orientation [SS02]. Object-Oriented [Wis86, She88, STN95, CSFG96, CH91, Cze90, HJL^{+}93, PPSS96, SS93, SSS90, UPTd92, VV89]. Object-space [KAÖ98]. Objective [PAE09]. Objects [Haa94, Van85, ACV03, ACC^{+}11, AF89, Ano06-35, ADOR02, AG94, BF02a, BD97, BK89b, CL97, Cho06, CKCK09, CP10, CH12, CR99, DG96b, DZ91, EME15, FGES96, Fiu89, GRPR08, GO10, GTG06, GKT02, GvK18, GN89, HZ15, HKYM01, IKTS22, JC95, JHPhR11, Jut94, KD11, KP95, KDG96, LY15, LKL^{+}20, Lin97, LWD^{+}18, LAE^{+}19, MW99, MMS89, MA94, MD99b, MC10, MG09, MS08, MO06, MLP01, ME17, NG88, OK99, OK03, PdSMdML^{+}12, PLM^{+}05, RRQ^{+}22, RBP96, SCVCN16, SKSI95, SLL^{+}21, SDS89, Sri02, SPL14, SK13, ÜT99, WZ21, WJW^{+}23, WTF95, YK23, YLT^{+}18, YY96, Zhu91, ZPL^{+}15]. Oblique [dSEM19]. oblivious [TPRC18]. observations [SM92]. obstacle [LMW06]. obstacle-avoiding [LMW06]. occluded [WTW^{+}23]. Occlusion [FHFG99, SBS04, TKOD24, AM12, AMGA12, BMH99, CC19, CMdL21, DVND10, ET07, GSF99, KKO01, KJ08, LZKJ23, SPMA13]. occlusions [HSE10]. Occupant [Wil75, WP77]. occurrence [FBH^{+}21, Loh95]. Ocean [TG02]. Oct [SW91]. Oct-Tree [SW91]. October [Ano20n, Ano21o, Ano22q, Ano23t]. Octpaths [Her85]. Octree [KDG96, San85, ZO07, AF89, IA91, KONS17, LTV08, SVNB99, MCT08]. Octree-based [ZO07]. OctreeBTFs [KWK17]. octrees [GSA89, Lib91, Sam89, MCT05]. octtree [Nav89, Wal89]. octtrees [MMS89, Woo89]. odd [FN10]. ODE [FRC06]. ODF [SVW23]. ODFNet [SMU22]. ODOPs [SVCNM23]. off [Bor91, Coo07, JHPhR11]. off-screen [JHPhR11]. office [CGR87, Hor83, Str83a]. offset [HPKE19, Kar92, MCS^{+}18, WJ91, ZR12]. offsetting [MHCL15]. oil [CC20, Rei75, SLKD16]. oilfield [KAV^{+}88]. OLED [CWC^{+}14]. OMiCroN [dSEM19]. Omni [Par86]. Omnidirectional [dSJ23, FSV17, FS20, PJVH^{+}24]. on-board [SH96]. On-Line [AB78, PLM^{+}05]. on-off [Coo07]. on-patient [MA15]. on-site [LZR22]. on-the-fly [STBG19]. One [PR96, BWBM20, Dra98, DK97, FN08, HH88, RPM96, TGG06, ZCL^{+}18]. one-dimensional [RPM96]. one-forms [TGG06]. one-parameter [Dra98]. Online [BKM16, DWH09, PCD^{+}15, CMLH21, CGG^{+}20, ECG^{+}22, PLFT21, VH15, WHW^{+}22, ZMYH06]. onto [LYW^{+}10, SXSY14, Zha98a]. OOPS [MAH00]. Op [IIK12]. opaque [SLYY97]. Open [Egl86a, FGZ^{+}22, Ano05k, CLH^{+}16, FP87, LBV14, LTPN96, NLOdS23, SBD^{+}94, TN02, ŽJŽ03]. Open-Set [FGZ^{+}22, NLOdS23]. OpenCL [HE15]. OpenGL [Bai07, BMH99, CC05]. OpenGL-assisted [BMH99]. opening [FB15b]. OpenSG [Rei04a, Rei04a, SBS04, WKE04]. opera [ZM07]. Operating [Ano07-33]. operation [GHM^{+}96, KL02, LTS88, PPV03]. Operations [BB93a, DN22, DM00, FR88, Gho93, MT88, PYD^{+}05, RF00, RA03, Wal89]. operator [AMHWW16, KD94a, WSX12]. Operators [Haz77, Row82, EB10, RBG^{+}09, SM98, Szy91]. optic [ZCZ^{+}18]. optical [AP99, Bim15, CS18, Hod91, IA99, IKM^{+}20, KKO01, KSH17, OK20, Sei99, SvRvL07, SvL09, YSD13, ZWS19]. Optically [KSR99]. optics [MWDG13]. Optimal [BSB^{+}02, DRFRD06, HLCF88, MG09, MB97a, PBN97, SB86, BCC^{+}22, CTQ^{+}14, GR93, JWL12, LCK16, LS18, OP13, Rag80, RKG22, SKM98, XXT18, ZK24]. optimisation [CCCP04, GCCZ14, Mul01]. optimisation-based [CCCP04]. Optimization [HYP^{+}20, BCC10, CXCH23, CSHZ04, CBNJ^{+}15, CB10, DKM^{+}20, JK90, JL23, dSJdML18, KPA12, LWLT11, LWW^{+}20, LZY^{+}23, LZT^{+}24, PJJSH16, RYNJ23, SWZZ23, SS82, WYZ^{+}11, WPB^{+}23, WKW16, WWF^{+}18, XCW^{+}09, XGC18, ZCC^{+}16, ZTF^{+}22, ZDL^{+}24]. optimizations [SEDT^{+}03]. Optimized [CXT18, GSF99, RE22, vRESH16, CGWW16, LY08, MCTB14, SPS12a].

Optimizing [ARM23, DA18, MBN21, RC08, KH00, WHW^{+}22]. Optimum [IORM17]. Options [WFS^{+}82]. Orbit [ZLLY06, Car99a, Ye02]. Orbits [She93, Mil93]. Order [Ano15a, ASZ^{+}14, CD93, HG02, LM22, LM16, MCTB14, RB06]. order-independent [MCTB14]. ordered [Egh80, SK03, Zha98b]. ordering [LPP^{+}19]. ordinary [CYJ^{+}13]. ordinate [CPLB14]. Ordinates [ERDS14]. Organ [FGM^{+}20]. organic [Bro08a, EPB^{+}19]. organisational [PBN97]. organization [SHLW89, WZZZ18]. Organized [Ano01d, KKMT06, PM13]. organizing [HK04, SK03]. organometallic [WMFR89]. Oriental [CLT07]. Orientation [And85, EME15, FTU95, MBN21, And82, SMU22, SBY11, SS02]. orientations [SHBSS17, XXT18]. Oriented [Dai93, Gra85, ME77, SVCNM23, Wis86, AMM^{+}07, BHTT94, CSFG96, CH91, CC08, CL18, Cze90, DSM^{+}99, FCM^{+}18, FRC06, GA83b, GS87, GB75, HJL^{+}93, KD94a, KS96, Kle86, KKŽ04, MGH13, MHLB16, Mar80, Mar82, Mar86, MO90, MA14, Muk86, PR23, PPSS96, SMMS01, She88, SS93, STN95, SSS90, SFC01, UPTd92, VV89]. Orienting [CHL^{+}11, LW10]. origami [Van10]. originally [Car92]. orthant [LCK16]. orthodontics [RSN^{+}07]. OrthoGaze [LPO20]. Orthogonal [HBA13, DLW23, HL93, LPO20, LYS^{+}16, WLX^{+}21]. orthographic [GZZS06, GN89, JCFN18, NG88, SLYY97, WTF95, YY96, ZPP^{+}23]. orthonormal [LLL^{+}15]. orzanized [Ano03a]. Oscillating [Roj94]. oscillators [Cha97]. osculatory [Pic89]. OSF [STN95]. OSF/Motif^{T M} [STN95]. Other [GO85, NM85, BMB95, Pic88b, PR96, Ste09]. Our [Web85, NH83]. Out-of-core [SW11, GK04, NIH08, VP06]. out-of-reach [MMS^{+}17]. outdoor [CSL18, DLZY14, DBW^{+}12, HFT^{+}99, LDM^{+}11, LZYQ22, OAYG10, WDH09, XLQP12, ZXD^{+}14]. outlier [ZWL^{+}22]. outliers [XXHM21]. outline [SR97]. outlined [SZ95]. outlines [SR02, SM07, SWZZ23]. output [Dix91, HR88, MBA20]. output-driven [MBA20]. overhead [ZK95]. overlap [TDR01, ZZH^{+}24, vOHR20]. overlapping [BD97, LYS^{+}16, MVCNI21, WA02, YZWJ24]. overlay [CPS^{+}22]. overlaying [PvdSLJ99, dMTB^{+}21]. Overview [Par86, AAAN23, Bro84, BK91, DDM^{+}06, Den90, DMT03, GOdSC23, MTC02, ZFG^{+}20]. own [HH88]. ownership [DGC^{+}21].

P [YD00]. P-buffer [YD00]. PA [LCX^{+}23]. PA-Net [LCX^{+}23]. Pacific [Ano94t, Ano95a, Ano95b, Ano95c, Ano02m, Ano02-46, Ano03-40, EZ22]. Package [KD86, CO88, PGB86, Rag80, Van89b, WSL92]. Packaging [Per77]. packed [ZK24, TPRC18]. Packed-Memory [TPRC18]. packing [BvW06, uHRBK06, Pic89]. Padé [Pat15]. Page [Mar91, Ano11o, GJN^{+}21]. PageRank [DWH^{+}15]. Pages [Ano19i, Ano19j, Ano20k, Ano20m, Ano20i, Ano20o, Ano20l, Ano20p, Ano20j, Ano20n, Ano21m, Ano21q, Ano21k, Ano21n, Ano21p, Ano21r, Ano21l, Ano21o, Ano22r, Ano22o, Ano22n, Ano22m, Ano22p, Ano22q, Ano22s, Ano22l, Ano23o, Ano23q, Ano23n, Ano23m, Ano23r, Ano23t, Ano23s, Ano23p, Ano24h, RGE07]. paint [SLKD16]. painterly [KPMT18]. Painting [DSJ19b, JXY87, LPD13, PJJSH16, ZYML23]. paintings [CHZ^{+}23, HLS89, Lee01, MPAC^{+}23]. paints [KK22]. Pair [LDS^{+}21]. Pair-wise [LDS^{+}21]. Pairwise [WLX^{+}21]. palettes [CCKN01, DSJ19a, XHL^{+}24]. Palindrome [RS96]. palindromization [Krö98]. palpation [LBLD97]. panels [Jar75, JVS^{+}24, ZEK^{+}17]. panes [CCW13]. panorama [DWH09, SPT18]. PANORAMA-based [SPT18]. panoramas [ZZLZ21]. Panoramic [GTTC03, PJVH^{+}24]. Paper [Ano99c, Ano01-36, Ano06m, Ano08b, Ano09c, Ano10a, Ano11c, Ano13c, Ano21j, Ano22k, DZ93, EDKS96, GLDK97, Ros92, Sou94, YQY90, Ano02t, Ano03k, Ano04m, Enc84b, Lee01, MML^{+}11, RR01, ST02]. Papers [Ano93e, Ano95q, Ano95r, Ano95s, Ano96a, Ano03z, Ano09a, Ano09b, FT02, LWP02, MLPB02, NA02, Shi02, ZS02, AL10, Ano77, Ano85a, Ano85b, Ano85c, Ano86c, Ano86d, Ano94p, Ano03j, Ano04g, Ano05m, Ano06f, Ano13o, Enc83d, Oli08, PB23, Ano01-31, Ano01-32]. parabola [Szy90b, WHH06]. Parabolic [RGRG15]. Parabolic-cylindrical [RGRG15]. paradigm [Jac93, MM18, SPGR93, TP89, dAPG18]. paradigms [PJ21]. paradox [GMNS93]. Parallax [SPMA13]. Parallel [Ano07y, AF11, CTS^{+}10, DTWT94, FRWW14, GP86, Haz79a, HS08, HF85, LCK16, LWW10, LTR^{+}14, MP93, ME17, PS91, RKG22, SIG02, SRF08, SPGR93, Smi93, TPB08, TTKA23, VHR^{+}18, WPH^{+}14, ASPO15, Ano01-41, Ano03-48, Ano04-27, AZF13, AÇÖ96, Ban97a, Baw97, BH91, CSK97, DMT03, Fle91, GPR^{+}95, Kla91, KKŽ04, KAÖ98, LL92, MR95, MB97a, MTM22, PMPR^{+}16, Pic94, RP20, RBL95, SDD95, SG15, SGES12, SHCW22, TP89, TBM^{+}04, WIP08, WJW^{+}23, WWO^{+}23]. parallelism [LSH^{+}12, NT00]. parameter [BKM16, Dra98, Eng93, EL22, GBA15, HBG14, PSMD14, RYNJ23, TSY11, TOY^{+}14, TY24]. parameter-space [EL22]. parameterisation [UBW99]. Parameterised [vdBB07]. parameterization [APS09, KW14, LKL02a, LZP^{+}04, LCZ^{+}11, LLZ^{+}11, LZL^{+}24, MC10, PT16, RGdL^{+}18, ZLGH10]. parameterizations [AF11, AZF13, CFH^{+}18, GCRN23]. Parameterized [XWWK21, BPKB15, MCMT14, WM89]. Parameterized-View-Learning [XWWK21]. parameters [BBHC15, Com01, COPR17, GR93, HEG98, LRHS14, Muk86, TCL15b]. Parametric [AO91, BSJC02, CSX^{+}19, CCM^{+}18, ML79, OdlCA02, Xu08, BLW12, BFRA11, Cha90, Dol95, Dor90, FMP96, HQW14, KCH^{+}22, LY15, LCCM02, MK83, OVWK16, Pie83, SNS06a, SXG^{+}09, SXSY14, VCQ92, ZQ12]. Parametric-surface [BSJC02]. parametrically [PTR^{+}92, RH85]. parametrization [LGWM17, WHFL16]. parametrized [van89a]. parent [LP93]. Parking [CE80]. parsing [LZKJ23]. ParSys [PLM^{+}05]. Part [Ano01-51, CLH^{+}23, HR07, IA83, KL02, LZG^{+}23, LTP19, MKM19, ÖT21, CGH97a, CGH97b]. part-aware [LZG^{+}23]. part-in-whole [MKM19]. Part/Future [Ano01-51]. Parthenon [CDPS06]. Partial [ATZM19, WYZ20, vBT21, BW94, FDGM18, IT11, LSW15, MRC15]. partially [CK02b, CDF14, KKJ^{+}23]. participating [CS03, ERDS14]. Participation [Ano93c, Ano95z, Ano95y, Ano95w, Ano95a, Ano95b, Ano95c, Ano95-41, Ano95-39, Ano95-40, Ano86h, Ano94h, Ano94f, Ano94g, Ano01i, Ano95-27, Ano95x]. Particle [dSASP^{+}15, CD15, dSJdML18, RPP20, CYCL09, DG96b, EW99, GBP^{+}17, GBP08, IP23, JSV98, LWW^{+}20, MP89, PVM^{+}22, PLM^{+}05, Pla00, PCPW15, SKKN10, SOC^{+}19, SKH^{+}05, WLZL18, WXL^{+}23, WWF^{+}18, XMD^{+}12, XXX^{+}23]. Particle-based [dSASP^{+}15, CD15, GBP^{+}17, IP23, SKKN10, WXL^{+}23, XXX^{+}23]. particle-in-cell [PVM^{+}22]. particles [CL18]. partitioning [BPS03, DBG92, EL22, SBS13]. parts [GO10, GvK18, JRJP^{+}22, RGdL^{+}18]. PASCAL [BDP82, JA84b, FN10, MTT82, Nug91]. PASCAL/GRAPH [BDP82]. Pasko [Adz22]. pass [LK00, PZM^{+}23, ZP07]. passenger [HIS83]. passes [MRG^{+}19]. passing [Ela86, MRG^{+}19]. passive [RB08]. Past [Ano94z, Ano94-27, Ano94-28, Ano95-33, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano95-38, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99j, Ano01x, Ano01z, Ano01y, Ano01-52, Ano01-53, Ano02o, Ano02p, Ano02n, Ano02-47, Ano02-48, Ano02-49, Ano03-42, Ano03-43, Ano03-44, Ano03-45, Ano03-41, Ano03-46, Ano04-45, Ano04-46, Ano04-47, Ano04-48, Ano04-49, Ano05-34, Ano05-35, Ano05-36, Ano06-40, Ano06-41, Ano06-42, Ano06-43, Ano06-39, Ano07-34, Ano07-35, Ano07-36, NT95]. Past/Future [Ano95-38, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k, Ano97f, Ano97g, Ano97h, Ano97i, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99j, Ano02o, Ano02p, Ano02-47, Ano02-49, Ano03-43, Ano04-47, Ano06-40, Ano07-36, Ano94z, Ano94-27, Ano94-28, Ano95-33, Ano95-34, Ano95-35, Ano95-36, Ano95-37, Ano01x, Ano01z, Ano01y, Ano01-52, Ano01-53, Ano02-48, Ano03-42, Ano03-44, Ano03-45, Ano03-46, Ano04-45, Ano04-46, Ano04-48, Ano04-49, Ano05-34, Ano05-35, Ano05-36, Ano06-41, Ano06-42, Ano06-43, Ano07-34, Ano07-35]. paste [BBP13]. Patch [CYW15, AÇÖ96, FJS11, LKL^{+}02b, MXK^{+}19, PKRM21, TFF^{+}20, WX14, WM24, YWC22]. patch-based [MXK^{+}19, TFF^{+}20, YWC22]. patches [ASC17, ASC18, CT13, GdMF03, KP09, MR90a, MZ89, Pet18, Pie88, PGVACN06, SVV23, SVSV20, TCL^{+}15a, Tar22, Zha98a]. PatchSwapper [ZZXT18]. path [BP10, ÇB22, LFP10, MTS21, PF16, RP20, RP22, SGC^{+}19, Sbe98, SAB12, SK23, STBG19]. path-based [BP10]. path-finding [PF16]. path-tracing [MTS21, STBG19]. pathfinding [RKG22]. paths [Car92, Car93, CSH08, CM23, KCR02, NAO13, Roj94]. patient [MA15, WH96]. patients [NGAS23, ZFG^{+}20]. Pattern [Ano04b, Ano05c, CFMP84, HXA^{+}12, Wei99, Zha96, DHZL20, Fen88, LAB^{+}14, LRG11, QMHH91, RS08, RBP^{+}01, TBDC20, XXC94, YBTB23]. Pattern-Directed [CFMP84]. Patterns [FBR^{+}17, MCPW21, RBLB21, Szy91, DTZ09, Dod09, GTFB23, Kar92, LP93, Li89, LSHL18, MLC^{+}22, Mas92, MVPL18, OCCZ12, Rag09, RPSP^{+}19, Sen00, Spr04, TB19, Van09]. Paul [HLS89, TL13]. pavement [TRB^{+}22]. PC [GV89, KGGP19]. PC-based [GV89]. PC-MRI [KGGP19]. PConv [RM22]. PDA [BJS01, CG96]. PDA-based [CG96]. PDE [UBW99, Uga06, YCZ04, ZY02]. Peak [Mok87]. peaks [YT87]. pearlescent [KK22]. pedestrians [KCH^{+}22, SGC^{+}19]. Peeling [PMM18, WTL^{+}11]. Pelvic [FGM^{+}20]. pen [HS11, LdSP^{+}08, SPL14, LdSP^{+}08]. pen-based [LdSP^{+}08]. Pencil [KYM12, AWI^{+}09, LHG21]. pending [Ano05y, SIG02]. pendula [PTR^{+}92]. Penny [TVL16]. Penrose [McC02]. People [Bij85, VCHR07]. per-pixel [LYZ15]. percentile [FWX^{+}18]. Perception [TDR^{+}17, YGS12, BX99, ESAH02, GS01a, GKT02, GMM^{+}23, Gro92, KP95, KS04, LLH17, LHL23, MWDG13, MZCD21, MTN22, MSL^{+}19, Nan77, POBB09, SM75, YSDG24, ZDD23]. Perception-driven [YGS12]. perceptive [LCGD24]. Perceptual [AYA^{+}20, ABCO12, CRD10, SLK^{+}23, WCS23, BCC10, Cad08, CWNA08, IP03, LS05, MXK^{+}19, ZZDZ10, vOHR20]. Perceptually [CFB15, MSR^{+}13]. percolation [RNM^{+}19]. PerEL’05 [Ano04-38]. perfect [BD98]. perforated [YPLL19]. perform [AA00, TKD16]. Performance [AP88, BS82, GP86, Ano91d, AMZ^{+}24, DAG22, DG96c, EHM84, EK15, JJL97, KKJ^{+}23, KSKS96, KYT^{+}17, LBLD97, Lea87, MA17, NC07, NBE^{+}04, Sla92, VZP22, ZFG^{+}20, ZZCL14, ZPIS23, ZGdDL^{+}96, ZFJ90]. performance-driven [MA17]. perfusion [HO12]. peridynamics [BDP^{+}17]. Periodic [XXC94, MH89, SWMdF21, WSHY22]. periodicities [WC89]. peripheral [SU93, Var92]. PERIS [ZPL88]. permutation [WCLT21]. perpendicular [Baw97]. PERQ [GO85, LH83]. Persistence [DL09, DS18, DHZL20]. Persistence-based [DL09]. persistence-generated [DS18]. persistent [GCDL22, YL23, ZDL22]. person [PSBD19, WCL23]. Personal [BG01, Mad94, Mad95, Str85b, Web85, CIK99, HvK87, MBPF12, McW87, Nug91, Shi04]. Personality [YSDG24, IM07]. personality-rich [IM07]. Personalized [WBRV16, ZWR^{+}18]. persons [IM07]. persons-New [IM07]. Perspective [GPC84, Kor90, San85, Suf88, Fal89, HT96, HK15, RR15, Sch88, TKS11, YZ17, ZS94]. perspectives [Bay95, CP21, CMSF11, DDM^{+}06, McC95]. perturbation [KLP01]. Pervasive [Ano02l, Ano04-50, Ano05-37, MAM^{+}24, Ano04g, Ano04-38, Ano06-44]. Petri [KS86, MRR98]. PEX-environment [PH90]. PGV [Ano04-27]. Phase [PTR^{+}92, Rag79, Pic87a, STT^{+}18, ZZJ^{+}14]. phase-field [ZZJ^{+}14]. phases [BCF06]. phasing [FNM20]. Phenomena [BN06, Ano05m, GSMA06, MG08, SAKB75]. phenomenon [VR06]. PHI [NP88]. PHI-GKS [NP88]. PHIGS [AEMT88, BBMR89, HR88, How88, Laf94, NP88, PH90, Sin87]. Philosophy [Wri75]. phone [BZYB21, KPB96]. Phong [BERW97, Cla90, OM96]. photo [CCCS08, GZSZ20, JWL12]. photogrammetric [BKL^{+}95, Ano07-37]. photogrammetry [ST02]. photogrammetry-based [ST02]. photograph [FBR^{+}17]. photographic [CCCS08, CS01]. photographs [GR09, LMR^{+}19, PJJSH16, WWL^{+}12]. photography [BB15, GS05, LGLK16]. photometric [LYZ15, MRC15]. Photon [JC95, ERWS12, LZ14, XWW^{+}19]. Photorealism [NT95]. Photorealistic [Ano12s, AVM05, CLT07, CST05, HZC^{+}22, LMJH^{+}11, MML12, PdSMdML^{+}12, CI11]. photos [NLSN11]. PHSColograms [MSF95]. PHSColograms^{REG} [MSF95]. phyllotaxis [FHP89]. Physical [Ano07a, Ano07b, AK06, AKW^{+}16, AM10, D^{+}01, MR05, Owe94, PLM^{+}05, RFB^{+}21, Shi04]. Physicalizing [ASS^{+}19]. Physically [XSQ^{+}97, CS16, CSLY01, DG96b, IP03, KES22, LCDN06, MMS15, TPB08]. physically-based [DG96b, KES22, LCDN06]. Physics [RSAF18, AJAC23, Gro91, GLT^{+}97, Mad00, NUM24]. Physics-based [RSAF18, GLT^{+}97, NUM24]. physiology [Gro91]. PIA07 [Ano07-37]. PIA07-Photogrammetric [Ano07-37]. Picture [BPR90, JA84c, NM85, Row82, ACA16, HW89, LZL^{+}19, OGSSLM^{+}07, RS75]. picture-logic [OGSSLM^{+}07]. picturebook [AA92]. Pictures [AL78, BG80, BPR90, ER87, RS96, Sch88]. piece [SAMA97, SJG19]. piece-wise [SAMA97]. pieces [ASM12]. Piecewise [BB93b, AO91, BJP97, BB91, Com01, GLDK95, HEWF13, LZ24, MS09a, PC23, WWL16, ZZC^{+}20b, ZPL^{+}15]. Piero [Fal89]. Pierre [FS78]. PIG [HG21]. PIG-Net [HG21]. PIK [BK91]. pile [AF89]. Piles [MB97b]. pilots [AMZ^{+}24]. Pin [CEG18]. Pincher [TVL16]. PinMesh [MAFL16]. pipe [FS88]. pipe-connections [FS88]. pipeline [AEW91, Bre01, FWW13, HR88, Liv23, MGH22, NIH08, PZM^{+}23, Sin87, TZT^{+}22]. pipes [DVG^{+}18]. PIPS [KR92]. pit [AM12, SD21]. Pittsburgh [AW93]. pivoting [MZ89, ZM92]. PIX [HJL^{+}93]. Pixel [CSL23, KPSN04, KK08, KSH17, KGK^{+}07, LYZ15, NLdAL^{+}23, SNS06a]. Pixel-Aligned [CSL23]. pixel-based [KGK^{+}07, SNS06a]. pixel-wise [KSH17]. Pixelated [GDA^{+}13]. PixelFlow [Mol96]. pixelization [SW21]. pixels [BD98, Bou02]. PL [FLV20]. Place [Dia94, BWZ^{+}18]. placement [AM12, ABJ90, GJN^{+}21, PGR83, SLL^{+}17]. Planar [GPC84, JA84a, LYW^{+}13, LJH18, SCH^{+}18, AXC22, AA00, Bay95, CFH^{+}18, CMdL21, CM15a, DGLRD18, DBG92, DHJ^{+}99, EK15, FTU95, FR88, Gom14, KLKE11, LDT02, LSW12, MSE20, MS07, MW92, Mes00, MPM15, MKPM17, NP96, RGdL^{+}18, RF00, SM07, SA17, SXSY14, SMM20, VM15, WSL92, ŽC99, ZEK^{+}17, ZXL^{+}21, ZZC^{+}20b, ZM92, ZXLF23, ZPL^{+}15, Zub88, ZJH87]. Plane [DBG93, LCX^{+}23, PK85, Rag79, She93, AJRV00, Baw97, Car96, Ent89a, GB91, GBV92, JS92, LŽ03, Lop92, Mal87, MKKM18, dCNPdFS14, Nav89, SB86, SJ94, SWMdF21, Szy90a, VRV05, ZP92]. planes [Bro76, CZ98, CP97, DTZ09, LPO20, MCT08, Pic87a, WLX^{+}21]. planet [DR15, LZY^{+}23]. planet-sized [DR15]. Planning [Dai93, GA88, AC90, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, B^{+}00, CXT18, CSJ^{+}21, CJJ99, GG96, GA83b, GB75, HO88, LFP10, LS90, LKC94, LQOW08, MTC02, MMT^{+}23, PBH19, RSN^{+}07, SGBP17, SVT86, Tay87, TB18, dDH87]. plans [How75]. plant [OMF93]. plants [VRV05]. plasma [Jar75]. plate [NLdAL^{+}23]. plates [KK21]. Platform [NHR^{+}22, BCS08, CH93, CS04, CJJ99, DBS^{+}11, GA07, SGPC20, ZMM^{+}90]. platforms [AZF13]. PLATO [WK14]. plausible [ACA16, ACC^{+}18, AAB18]. playback [EHM84]. players [MLM^{+}17]. playing [Che06, SGBI02]. Plot [Wen84]. plots [SHCW22]. Plotters [ML79]. plotting [Pie83, SSK87, TSD87]. plus [MA15, SBS04, Rei04a]. PLX [LRMS92]. pockets [GD11]. poetry [CHZ^{+}23]. Poetry4painting [CHZ^{+}23]. Poincaré [WCdA98]. Point [Ano04-51, Ano07-38, BSL^{+}13, BCMM07, BP08, DFF22, GYK^{+}23, GJN^{+}21, KZ04, KF88, LWW07, LWWY22, ODZS11, SP04, TWNL22, WM95, ZCL^{+}13, ZZH^{+}23, ZP04, AAK22a, AK21, ACA07, ADR01, Ano06-49, BAD23, Bec95, BLW12, BGLA18, Bou20, CHL^{+}11, CXGL23, CRT04, CZCG04, COPR17, DWZ^{+}22, DFWW15, GYY^{+}23, GM04, GMd^{+}13, GS11, GLZ^{+}21, HDZR23, HG21, HLZ^{+}17, HZLQ20, HHN^{+}23, HMHB08, JDT^{+}23, JWZ23, KD11, KYL15, KK16, KB04, KVB^{+}20, KGB^{+}21, Lam00, LSK^{+}10, LCZ^{+}11, LZG^{+}23, Lin90, LZC^{+}15, LJJPF22, LCXL20, MAFL16, MKC08, MRF06, MTSM10, MBST22, MPM15, MKPM17, MRWL23, Mor76, NOS15, NZL^{+}21, OVWK16, OMW13, ÖT21, PLJ^{+}13, PS13a, PECW22, QL23, RB20, RMP19, RHK^{+}20, RBF17, RRQ^{+}22, RRB^{+}22, SMU22, SDT21, ST22, SSV07, Sar92b, SW11, SMK08, SBY11, SG15, SPS12b, SWL^{+}16, SHS^{+}23, TZT^{+}22, TGG06, TBDC20, TCL15b]. point [VBS^{+}15, VZP22, VH02, VDOK19, WBB^{+}08, WLW05, WW06, WTL^{+}11, WSL^{+}19, WCLT21, WZL^{+}22, WPL^{+}23, WLZ^{+}23, WCH^{+}24, WSL92, WAM17, WGLS00, WLX^{+}21, WYC^{+}23, WM24, WLG04, XSW23, XLGG11, XWW^{+}18, YRS^{+}18, YZWJ24, YIL09, YHHS93, ZK07, ZP07, ZSW08, ZGW^{+}16, ZLL^{+}21, ZZH^{+}24, ZZC^{+}20b, ZTF^{+}22, ZMH^{+}23, ZDL^{+}24, ZS02, ZDL22, ZWL^{+}22, vOMRI^{+}15, YCO23]. Point-Based [Ano07-38, BP08, Ano04-51, BSL^{+}13, BCMM07, SP04, ZP04, Ano06-49, GLZ^{+}21, KB04, MKC08, TBDC20]. point-clouds [SSV07, SMK08]. Point-in-polygon [LWW07, MRF06, WLW05]. point-in-polyhedron [Lin90]. Point-line [LWWY22]. point-sampled [GM04, WW06]. point-set [PLJ^{+}13]. point-sets [WSL92]. Point-Voxel [ZZH^{+}23]. Point2MM [GYY^{+}23]. Pointer [FSM94]. Pointing [FSM94]. PointMatch [WYC^{+}23]. PointNGCNN [LCXL20]. Points [Cor82, AM91, BWdBP13, BCDD22, CVB16, Dra98, EH96, FR92b, LW10, MW14, PK91, SS22, SXSY14, TWNL22, Wil03, WWF^{+}18, Yin04]. Poisson [GYJZ15, KJS18, MBV18, SJZW07, XXHM21]. Poisson-disk [GYJZ15]. polar [CGR98, CDF14, Lam00, Suf88, VRV05]. polar-plane-based [VRV05]. Polarization [dMM19]. polarized [CPLB14]. policies [AK13]. Politècnica [BJAN^{+}95]. polycube [FCM^{+}18, HWFQ09, SP16, WYZ^{+}11]. Polygon [GP86, TP89, ŽJŽ03, AMGA12, BR07, CC08, GS01b, GK04, HLL^{+}18, Kno87, KAÖ98, LŽ03, LL04, LL05, LWW07, MCAH12, MRF06, Ran87b, SF98, Shl83, Ska93, WLW05, ZCL^{+}22]. Polygonal [PS02, SA86, BMH99, BTV83, BC88, DAHF04, DZ91, LOdF02, MAF19, MCS^{+}18, MLP01, MS01, PR93, RLB^{+}02, SV18, SGC00, SMS09a, SMS09b, SSBT01, TT12, ZD04]. polygonization [CCC97]. polygons [CW03, FTU95, Kau88, LDT02, LLGA12, MK89, PYD^{+}05, Pha95, RF00, RFR02, RdMF08, RF97, SG15, SWMdF21, Van09, ŽC99, Žal01, Zub88]. Polyhedra [Chr78, Bro76, FT97, JS08, MFOK94, Ran88, TKZ^{+}13]. Polyhedral [ZTS02, FR88, GH98, KP22b, KB05, PCV16, PVCM22, VL98, Wan04, XhKKL16, YZ17]. Polyhedron [SBS22, Sug84, CSZ92, KS09, Lin90, Ska97, ZM91]. PolyMeCo [SMS09a, SMS09b]. Polynomial [PS15a, BJP97, CM15a, JT02, Shi93a]. Polynomials [Ala85, Dra98, GLDK95, IORM17, IPV96, Lop92]. Polynomiography [Kal04]. polytopal [SA87]. polytopes [Sob89, VT06]. polytree [WLC88]. POMAR [MGH13]. pooling [FWX^{+}18]. pop [LTS96, WLX^{+}21]. pop-up [LTS96]. popularity [BD17]. populating [BM03]. population [DLR^{+}10, LRD07, LQ15]. porous [DHZL20, GCDL22, GVTA10, Lin15, VAGT08, YL23, YLT^{+}18]. Portable [Cul84, FBT04, Liv23, Smi75]. Portrait [TL13, SW21, ZDL^{+}19, ZWR^{+}18]. PortraitNet [ZDL^{+}19]. Portraits [Rag79, PTR^{+}92]. Portugal [Gom90, LG89]. Portuguese [GMM18, DQF04]. pose [DYW^{+}22, FGL23, KCS22, LMC13, LTP19, MENS19, PR23, SRZK23, SdSR^{+}19, TOY^{+}14, VDOK19, WPB^{+}23, WXC^{+}23, WPJP23, ZMK18, ZLL^{+}20, ZHW^{+}21, ZHG^{+}21]. pose-independent [LMC13]. poses [RB08]. posing [DFWW15, MHLB16]. Position [BKCW14, FM17, WLJT19, BMdSVR18, CHC^{+}24, LKHM19, MBC^{+}23, NZL^{+}21, STT^{+}18, TNF14, WCH^{+}24, ZCC^{+}16]. Position-based [BKCW14, WLJT19, MBC^{+}23, STT^{+}18, TNF14]. positional [KYT^{+}17]. positioning [DWH^{+}15, PTY^{+}16, XZY^{+}23]. positions [MLM^{+}17]. positive [VM15, YJLZ21]. Positivity [BB93b, AB03, BMB95]. Possibilities [Sal85, Vää93]. possible [BCDD22, CL18, DKV^{+}22, PP12, ZLL^{+}15]. post [RPP20]. post-processing [RPP20]. postgraduates [BHL^{+}94]. postprocessing [CTLG94]. PostScript [Laf94]. Posture [SWH^{+}17, WSX12, ILLC01, II22]. Posture-based [SWH^{+}17]. Posture-invariant [WSX12]. Potential [Lam87, CAL^{+}04, PM93, SSB^{+}08]. potential-based [CAL^{+}04]. potentials [GD11]. Pothole [TRB^{+}22]. pottery [VR16]. Power [Pic94, WFS^{+}82, JSMK14, NSL16, RHK^{+}20, XCXC23, ZGC^{+}19, ZYW^{+}21]. powerful [REG^{+}89]. Practical [BMP84, CFZC19, CEPS13, GBP08, Wri75, FP89, MWY^{+}10, San06, XLQP12, vTTK^{+}20]. Practically [KKŽ04]. Practicals [Mei83]. Practice [Mil92b, Ano94-32, Ano94-31, Ano04-59]. practice/ICA [Ano94-32]. pragmatic [Smi75]. praxis [Nap95]. pre [DWH^{+}15]. pre-positioning [DWH^{+}15]. Precise [MvSE18, GH98, MSE20]. precision [HCC13, Kra10, RHK^{+}20, RZY^{+}20, VGP04, WHH21, WCH^{+}24]. Predicting [HYP^{+}20, XLL^{+}20, MZCD21]. prediction [BBMGM22, ES22, KCS22, LZW^{+}21b, SLM^{+}22, XWW^{+}21, YCO23]. predictions [DCLB19]. Predictive [FMCM^{+}21, HJ03, PSMD14]. Preface [Ano10h, Ano11l, BLS15, DPS10, FLMY16, NP12, YSWP09]. preferable [HWSW19]. preferences [SW19]. Preliminary [GS83]. PREMO [RBUB94]. Preparing [BSW78]. Presence [MVRB18, A^{+}01, BMdSVR18, D^{+}01, DGC^{+}21, HZC^{+}22, JP10, KSH^{+}19, LW24, SBHS10, LUMC04]. Present [McC84, NT95]. Presentation [Con77, Pic84, HK93, JK90, KR96, KP95, KN88, RBUB94, TPK13]. presentations [VHON04]. Presenter [PCD^{+}15]. preservation [APA^{+}11, KFN23, ZYML23]. preserve [BMB95]. preserved [HL97]. Preserving [BB91, BB93b, SAMA97, AB97, BS98, BX99, BWBM20, Bru14, CZL14, ITW^{+}20, LYL^{+}17, MX14, Mon87, PB96, SJT20, SP23, SK06, VL98, WYZ^{+}11, WLL^{+}12, WZC^{+}21, WLL22, WLL21b, ZQL15, ZZJ18, ZFS03]. Press [CS16]. pressure [BGPT18]. Prevention [Moh77]. Previewing [JV84, WZS19]. Previous [TMK94]. PREVIS [FMCM^{+}21]. Primal [CHC^{+}24]. primary [BM03]. prime [Lak89]. primitive [HL93, SF92, WEWL99]. primitives [DKY97, FD75, KK12, MTT82, PDL^{+}21, RRQ^{+}22, RRB^{+}22, SS22, ZHC11]. Principal [ERDS14]. Principal-Ordinates [ERDS14]. principle [Aro89]. Principles [Kru84, Mil92b, MYC15, ABCO12, BERW97, HS99, Mit87]. print [Fri03, GR09]. Printed [Per77, LAE^{+}19, PGR83, YPLL19, YLT^{+}18, ZLZ^{+}20]. printer [TBS^{+}23]. printers [Fer01, SR89]. Printing [Ano07e, Ano07d, SB83, ASS^{+}19, BCC^{+}22, CTP^{+}21, CSG^{+}17, Dix91, EME15, LXJL21, XZY^{+}21, ZWR^{+}18]. prints [HKBA17, KEVD18]. Prior [LSCJ23, HHCM17, HGJC21, JL23, QSXT22, QSLS23]. Prior-based [LSCJ23]. priori [WRLZ23]. prioritization [SKH^{+}12]. priors [CLXJ22, CLE13]. prism [WFC14]. privacy [KFN23, LWZ^{+}23b]. privacy-aware [LWZ^{+}23b]. Prize [Ano93d, Ano94i]. proactive [VSKG03]. Probabilistic [KYL15, YWC22, ATHL14, Kru99b]. probability [FN10, TMB^{+}05]. probability-based [TMB^{+}05]. probe [MCV18]. probes [CNC^{+}21]. Problem [DW82, ANE17, ARL^{+}20, Bec95, BjOwKM12, How75, PK91, SS93, SS75a]. Problèmes [Dom77]. Problems [SvD03, BPS06, Des00, Kor87, LW99, Mal87, MN90, Rag80, Sal85]. ProbNet [WZL^{+}22]. Procedural [BPD15, BTC94, CP96, DG07, E^{+}00, EHSF17, FM22, HR04, HJT^{+}13, LVLD10, PP02, RPP21, TAF16, Ano13l, BAD23, CSP19, CGG19, HY23, JPCS18, KK12, MP19, Per02, RLD^{+}12, SBR23, STdKB11, TKM^{+}24, XM12, YMYH12]. proceduralization [DA18]. procedurally [BBP13, FPC10]. procedure [FK82, Med86, MZ89, Mon87]. procedures [SS75b]. proceedings [Ano08i, Tan94]. process [ABCD93, AC90, BAD23, BBH90, GA83b, LW88, MTS^{+}22, ÖOK23, Sel89, TBS^{+}23, Thi85, UWC90, VCHR07, Whi04, WWK11]. Processes [JK84, BT94, GZ99, IR06, IKM90, JMV90, Lak91, LCA19, MMV88, Mor75, SJB^{+}21, WJD^{+}09]. Processing [Ano02-36, Ano02-37, Ent77, GM78, Haz79a, HF85, Pos77a, Row82, Sch86a, Sch85b, WBB^{+}08, Ano03-54, Ano04-30, Ano06-48, AMZ^{+}24, BS09, BLS15, Bou20, Bre01, CP21, CH91, CLT07, CRT04, DA18, FS86b, FSM94, FD75, Gal88, Gia15, GS83, GBKG04, HHG97, JVS^{+}24, KHS03, KS15, Les01, MD85, Pic88b, RR92, Ran87a, RPP20, She12, Ska96, Smi93, WLL22, XZCOX09, YPSZ01, Yu99, Ano94r, Ano94q]. Processor [BY88, CYKK09, EW75, GPR^{+}95, LK00, LZ88, Lin79, TP89, TPB08, WBL^{+}97]. Processors [OKK83, GKLM07]. produced [Rag09]. producing [Wal93a]. Product [ÇB22, RBUB94, BCF06, Lam00, LQ12, LPZ^{+}21, MvSE18, Sen99, Sen98]. product-delay [Sen99, Sen98]. Production [Ano94p, DK92, CH15, HPD^{+}10, JK90, MNI^{+}16, MMH^{+}21]. products [FBR^{+}17, KKS93]. Professor [Adz22, Enc15b, EJ15]. Profile [Big86]. profiled [MA94]. profiles [LLP^{+}21]. Program [Ano01v, Ano01-44, Ano01-45, Ano02s, Ano04-37, Ano04-32, Ano05z, Pie90, SB84, Ano01u, Ano02-44, Ano02-45, Ano03-33, Ano04-36, Ano06-32, Goh84, MNI^{+}16, RM05, WA75, PB23]. programmable [LK18, TMB^{+}05]. Programme [San06, BL82]. programmed [Fra86]. Programmer [BK91]. programmers [Spr75]. Programming [Haz77, Web84, Web85, ZPL88, BDP82, HR04, Lam99, Lew75a, Ner75, PP90, Pur87a, SBHC22, TM75, ZM91, ZGdDL^{+}96]. Programs [PMK85, WP77, And82, GBG04, GD00, JSG04, SVV92]. progression [BDP^{+}17]. Progressive [DGBNV18, GM07, KWK23, LLP^{+}22, MCAH12, OK03, PR11, SPS12b, WZS19, AÇÖ96, BKV05, BJP97, BLZD12, HCX^{+}23, Hop98, JL23, rKC93, LZ11, LFL02, MGH13, RBL95, SBSL04, VGP04, WBP92, WFG03, ZP92]. progressive-iteration [LZ11]. Project [Smi77, BM03, BSGT03, BGK89, Dur91, VPLL06]. projected [ENE11, Muk86]. Projecting [SXSY14]. Projection [PPD22, Sug84, AYA^{+}20, ABMC^{+}15, And82, Baw97, CTQ^{+}14, ET07, HT96, HHLE17, IMG22, LYS^{+}16, LBTM15, MKG00, PWJ^{+}18, RBL95, SvRvL07, TMSPB09, YZ17, BCC^{+}22]. Projection-based [PPD22]. projection-induced [AYA^{+}20]. Projections [Pre84a, BFT23, CVP^{+}16, FR92a, JCFN18, MCMT14, Roc89, TZvD^{+}21]. projective [CMdL21, Yu99]. projects [KS96]. promising [GL83]. Prompt [XLW^{+}24, HCX^{+}23]. PROMS [Ano01c]. Prone [Ber79]. proof [Mar02a, QKS01]. Propagated [LCW^{+}18]. propagation [CSHZ04, Cho06, FJS11, KD00, LFL02, MX12, MX14, RGGB02, WAM17, WSWL14, XHL^{+}24, YY14, ZZLZ21, ERDS14]. propeller [BL82]. propellers [RJKV12]. propensity [PQ10]. Properly [LLL^{+}15]. properties [ASWL11, ATZM19, BPS^{+}10, Coo00a, RFB^{+}21]. property [LLP^{+}21, US20, WKW16]. propos [Le 77]. proposal [DM01, LRMS92, OPR92, Paq05]. proposed [BK91]. props [ARM23, GHFH08]. ProSeCo [BHZ^{+}21]. prospects [HK15]. prostate [CC04, ZBM00]. Protein [GRF^{+}22, AGM^{+}21, BRHB20, FEJM75, GD11, LPL^{+}20, RFB^{+}21, SFS^{+}21]. Protein-ligand [GRF^{+}22]. proteins [MRW^{+}21]. protocols [Ano01c]. proton [MMT^{+}23]. prototiles [Fat01, Fat02]. prototype [LLM^{+}90]. prototypes [Nar15]. Prototyping [PRRR13, CS16, KAFB18]. provide [HGH^{+}18]. Providing [Gra02, Cun00]. proximal [MAG^{+}12]. proximity [DMM23, KZ04]. Proxy [ZZC20a]. Proxy-driven [ZZC20a]. prune [CS06]. pruning [GCRR11, ML12]. Pseudo [Car96, DCL07, HLXL23, Mac78, Bow95a, SXG^{+}09, Wir80]. Pseudo- [Car96]. Pseudo-Machine [Mac78]. pseudo-manifolds [SXG^{+}09]. pseudo-random [Bow95a, Wir80]. Pseudo-training [HLXL23]. pseudocolor [Pic88b]. pseudorandom [Ric89, Szy92]. PSMs [KCH^{+}22]. Pub [Ano04x]. Public [CTJ^{+}14, NFW^{+}24, SFD06, VSKG03]. Publication [Ano03v, Ano03w, Ano03u, Ano04z, Ano04y, Ano05u, Ano05v, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano17h, FEJ04]. publications [Ben79]. Publisher [Ano10i]. PUBLISHER’S [Ano06-45]. Publishing [Mar87, Mar91, VHON04, ZR97]. PUPs [MRS17]. Purpose [KAV^{+}88, EW75, MTTLL82, MDSU88]. purposes [Gro92, KJTS96, SET^{+}88]. pushes [HGH^{+}18]. Pushing [MWDG13]. puzzle [BRO22, Edg00, LYS^{+}19]. puzzle-based [BRO22]. puzzles [EK22, KS20, OGSSLM^{+}07]. PVG [Ano01p, Ano01-41, Ano03-47, Ano03-48]. PVLNet [XWWK21]. pyramid [LFY^{+}21, Nug91, WCH^{+}24]. pyramidal [KNMP14]. pyramids [FO21, LSGFRC^{+}13]. Pythagorean [Bro07a].

Q [Haz79a]. Q-Splines [Haz79a]. QCD [TBLH17]. QCDVis [TBLH17]. QR [YPLL19]. quad [AKW^{+}16, BMW12, Gin02, HXA^{+}12, HW89, KYKK19, KP19, LCZ^{+}11, LWLT11, Liv23, SA86]. quad-dominant [KP19, LCZ^{+}11]. quad-edge [AKW^{+}16]. quad-meshes [HXA^{+}12]. quad-pattern-coverable [HXA^{+}12]. quad-tree [SA86]. quadrangulation [LLZ^{+}11, Tar22]. Quadratic [Pha89, WY11a, CM15a, Eng93, Mil93, PB96, RPM96, RPM97, SP95, YRS^{+}18]. quadric [HL96, JS92, Kuo01, SJ94]. quadrics [GBF14, LNFC95]. quadrilateral [AXC22, DLS^{+}11, LZT^{+}24]. quadrilaterals [Jas88, Str86]. Quadruple [BR89]. Quadruple-step [BR89]. quads [ZXLF23]. quadtree [HY93a, ME17, TPRC18]. Qualitative [GZSZ20, NFLYCO99, SA04]. Quality [LZC^{+}15, MSE17, SR97, XCW^{+}09, BFT23, CVP^{+}16, DLV06, DFF22, FWX^{+}18, Frü91, GP91, GBP04, GS04, HGVV16, HY23, LPV95, LZG^{+}23, LZLS18, Lod21, MZPZ16, MCMT14, MCS^{+}18, MMK04, MMH^{+}21, Nah23, PSMD14, SvLBF10, WTM12, WLX^{+}23, XGC18, YWH^{+}16]. Quantification [MSMK19, SP00a, VAGT08]. Quantifying [AHR^{+}22, MF02]. Quantitative [THL15, ZZDZ10, NFLYCO99]. Quantization [Dix91, Fle91, HE15, SMK08, YL96]. quantized [CH94, JX96]. Quantum [Tak77, TBLH17]. quartet [GZSZ20]. quartic [FM00]. Quasi [CT13, She93, MR92, RB20, SKCP99, XCZP14]. Quasi-developable [CT13]. quasi-helixes [XCZP14]. quasi-interpolant [RB20]. quasi-Julia [MR92]. quasi-random [SKCP99]. Quasi-Stable [She93]. quasicrystalline [Rei02]. quasicrystals [ADR01]. quaternion [Jut94]. quaternions [Nor89]. queries [MAFL16, RGE07, SS12b]. Query [DN22, KKJ^{+}23, KS91, KJ08, LWZ^{+}23b, MKDM22, WGLS00, YBTB23]. Query-replace [DN22]. querying [vBT20a]. questions [McW91b, RBLB21]. queueing [Wal06]. QUICCI [vBT21]. quick [Lin90]. QuickETC2 [Nah23].

QuickETC2-HQ [Nah23]. quintic [LJH18]. Quotient [Rei92].

Radial [vBT20b, ABMC^{+}15, HZD^{+}19, KJS17]. radiance [CZR22, JK21, LWFZ23, PLVT23, SLF^{+}24, WCF^{+}23]. radii [BYQZ22]. radio [PSK^{+}11]. radio-frequency [PSK^{+}11]. radiofrequency [RGH^{+}19]. Radiological [KHS^{+}10]. radiometric [AG13]. Radiosity [DMG99, AF16a, AÇÖ96, BJP97, BSB^{+}02, CSH08, CS98, CM06, FB14, FB15a, Ng95b, PAJ19, RBL95, RGGB02, Sbe98, SDD95, SZEG93, Shi92, WBP92, XPL90, ZP92]. Radiotherapy [FGM^{+}20, FMCM^{+}21]. Radon [Bim15]. RAGI [SMMG22]. railway [HIS83]. rain [CP13, RJG06, WJGG15]. raindrop [QSLS23]. rainfall [VPLL06]. raising [ZWWC23]. Random [MBST22, MKKM18, Sbe98, WPL^{+}23, Bow95a, CSH08, QLCV96, SKCP99, Wir80]. randomized [KB02, WYXM22]. randomly [MGH13, SK03, SWL^{+}16]. Range [Ala85, Ano13j, H^{+}00a, Han95, HOCN07, Mou13, PLJL15, PCGS15, PCKB23, PR11, SL12b, TSD87, vLvKV11, CAS^{+}15]. rank [ALM19, LZL^{+}15, ZCL^{+}13, PP20]. ranking [BSAH^{+}23, LKLW16, TCL^{+}15a]. RankViz [PP20]. RANSAC [WLX^{+}21]. Raphael [KH86]. Raphson [Wal93b]. Rapid [Dol95, Goe95, LD03, RFZ^{+}17, SWL^{+}16, dBWK18, AS95, CGH97b, CHSD95, WXC^{+}23, YHX10]. Rapidly [KP18]. Raster [JV84, Jer85, Neg77, Rix84, Shl83, vJ84, ADR01, Bor91, CEM89, GL83, Lin79, MCTB11, Par88, Rix83, SDS89, WBL^{+}97, WM89]. raster-based [MCTB11]. rasterization [Ack96, KK08, Kla91, LK00, MB97a, MGH22, Wüt98]. rasterization-based [MGH22]. rasterized [DGLRD18]. rate [DMG20, DH95b, KK16]. rates [LB75]. ratio [CMdL21]. Rational [Ala85, DDPT98, FK82, GSY94, LL91, MR90a, PS15a, SY23, Sar92b, Sar92a, Sar93, Sar94b, SAMA97, Sar00, SBH01, Sar02, Sar03, Sas04, ZDT07, ZZ15, ZFS03, ZFSY04]. rationalisation [Le 77]. ratios [GVC^{+}20, MRC15]. raw [CHL^{+}11, LYL^{+}17, NTAI20]. raw-scanned [NTAI20]. Ray [Cyc93, GMP22, Ger86, IO91, MS90, SW91, SKH^{+}12, ST23, SSA96, Van85, dVTT18, vWdlHFK23, AA07, AZ23, BS93, BCG^{+}94, BHH15, BSB^{+}02, CC19, CF99, CSSC00, DGBNV18, HP03, Hod91, HMHB08, JC95, JK21, KPFT03, rKC93, K^{+}00a, Koh97, KM21, LG23, MHCL15, Mar01, PM91, QGGW97, RHK^{+}20, Sam89, SEDT^{+}03, SLS03, SGGC05, SDD95, SSB^{+}08, Ste75, SLF^{+}24, SKM98, SKP99, UPTd92, VHS12, WIP08, WBK98, WZ21, WTW^{+}23, WF11, WGS99, WKW16, ZM91]. ray-based [MHCL15]. ray-bundle [SKP99]. Ray-casting [GMP22, AA07, rKC93]. ray-convex [ZM91]. ray-reps [WKW16]. ray-traced [LG23]. ray-tracing [AZ23]. ray-triangle [RHK^{+}20]. raycasting [DEST95, GBKG04]. RayFilling [BCG^{+}94]. Rayleigh [Rei92]. rays [AF89, RAA^{+}08, SKP99]. RBF [ACGC22, TWBP03]. RBF-based [TWBP03]. RBW [Shi93d]. Re [Cun00, LKLW16, LZL^{+}24, PSBD19]. re-identification [PSBD19]. Re-inventing [Cun00]. re-parameterization [LZL^{+}24]. re-ranking [LKLW16]. reach [MMS^{+}17]. reaching [DBS^{+}18]. reaction [McG08]. reaction-diffusion [McG08]. reactive [MSHL22]. reactivity [WMFR89]. readability [EVRW23, XZY^{+}23]. Reader [BCHM02]. reading [BCHM02]. ready [WWK11, XLL^{+}20]. ready-made [XLL^{+}20]. ready-to-use [WWK11]. Real [AOB17, AJAC23, BLZD12, BFSE03, BP10, CJXZ23, CF13, DVND10, DG17, ES22, FZPM93, FCW^{+}10, GSV^{+}18, GG14, JXJ22, KLW12, KK22, LCCS04, LSHL18, MWLZ22, MPQG18, Mor75, MA17, NAS18, PdSMdML^{+}12, PSBM10, RHC15, SdSR^{+}19, SCSG18, TSY11, WF11, XMD^{+}12, YHNC22, ZXL^{+}21, ZDL^{+}19, ZYX^{+}19, ZGWP16, ACV03, AALJ20, AJ94, ABAA22, ADOR02, AP22, BGD18, Bay95, BSF13, BH91, CP19a, CF11, CAAC20, CTN^{+}17, CB97, CVHM03, DVG^{+}18, DWZ^{+}22, EW75, EHM84, ERDS14, Eng93, FP75, GRPR08, GW06, GYD75, GCRR11, GSB^{+}21, GPR^{+}95, HBOS13, HG22, HZ23, HXH24, HGH^{+}18, JJPP^{+}22, KL07, KKO01, KM21, LRD07, LWWY22, Lin79, LD12, LM16, LYZ^{+}22, LCX^{+}23, LSS21, Lop92, MAM^{+}24, MDJ^{+}95, Mar10, MMK04, MYC15, MSRB17, Nah23, NLSN11, PGS^{+}23, RJG06, SBR23, SLS03, SV06, SWvB95, SPS96, STBG19, TSD87]. real [TC24, TG02, TKOD24, WLP^{+}14, WZ21, WJW^{+}23, WJGG15, WCA^{+}11, WLJT19, WLG04, XLQP12, YSZ22, YK23, ZGS17, ZZH^{+}24, ZLLG18, ZZXT18, ZW88, ZMM^{+}90, dHG^{+}97]. real- [PGS^{+}23]. Real-Time [FZPM93, AOB17, AJAC23, BLZD12, BFSE03, BP10, CJXZ23, CF13, DVND10, DG17, FCW^{+}10, GSV^{+}18, GG14, JXJ22, KLW12, KK22, LCCS04, LSHL18, MWLZ22, MPQG18, MA17, NAS18, PdSMdML^{+}12, PSBM10, RHC15, SdSR^{+}19, SCSG18, TSY11, WF11, XMD^{+}12, YHNC22, ZXL^{+}21, ZDL^{+}19, ZYX^{+}19, ZGWP16, AALJ20, AJ94, AP22, CF11, CAAC20, CTN^{+}17, CB97, CVHM03, DVG^{+}18, EW75, EHM84, ERDS14, FP75, GRPR08, GW06, GCRR11, GPR^{+}95, HBOS13, HG22, HZ23, HXH24, HGH^{+}18, KL07, KKO01, KM21, LRD07, LWWY22, LD12, LCX^{+}23, LSS21, MAM^{+}24, MDJ^{+}95, Mar10, MMK04, MSRB17, Nah23, RJG06, SBR23, SLS03, SWvB95, TKOD24, WLP^{+}14, WZ21, WJW^{+}23, WJGG15, WCA^{+}11, WLJT19, WLG04, XLQP12, YSZ22, ZLLG18, ZZXT18, ZW88, ZMM^{+}90, dHG^{+}97]. real-world [DWZ^{+}22, JJPP^{+}22, LYZ^{+}22, MYC15, TC24]. realisation [San93]. Realism [vJ84, GMM^{+}23, SL12a].

Realistic [AALJ20, AL78, CG85, CP13, HSE10, LLLC11, Mar10, MRC15, Pla00, RJG06, YMYI11, YP05, ZPL88, BLZD12, BGD18, CDPS06, FCG^{+}21, FNM20, Maj98, OMGGG^{+}19, PF97, ZDD23, ZSS20]. réalité [Nan77]. realities [SW19]. Reality [Ano95x, Ano95y, Ano95w, Ano01-59, Ano01d, Ano06c, Ano06-29, Ano06-30, Ano06-31, Ano07-39, Ano07-30, Ano11m, AMS22, ABK11, BYC19, CSF20, CL92, Gob93b, GMM^{+}23, HWEB22, HSTR20, KNDT20, LS06, MAM^{+}24, MTS^{+}22, NTTF21, PS15b, PGS^{+}23, TKdJO22, vDLS02, AG13, ABAA22, Ano01f, Ano01-42, Ano02-53, Ano03-36, Ano04a, Ano04p, Ano04-39, Ano06-33, Ano12l, AP99, A^{+}99, BZYB21, BND^{+}17, BCS^{+}99, BSMG24, BES01, BNS24, BAS^{+}20, Bro00, BBMK21, Bry93, CEG18, CP19a, CIT^{+}19, CACC24, CSJ^{+}21, DBS^{+}18, Dor99, ET18, FBT04, FEBS07, FFP^{+}21, GWEA^{+}11, GPC^{+}17, Gob93a, GS93, GZ99, GLL00, HZ15, HGS23, HAL^{+}21, HWSW19, HEW^{+}18, HFT^{+}99, H^{+}01, HGH^{+}18, Jac93, JP10, KS03a, KKC94, KSH^{+}19, KSR99, KTMW12, Koh96a, KÇM00, LDM^{+}11, LBLD97, MCV18, MCMV22, MCM^{+}18, MML^{+}11, OMGGG^{+}19]. reality [ÖOK23, PCY^{+}06, PvdSLJ99, PAE09, PAE^{+}21, PTY^{+}16, QLF^{+}09, Que93, RSP^{+}19, SWF^{+}20, SCB^{+}24, Sei99, SBE20, SK12, SUF^{+}18, TBS^{+}23, TSY11, TOY^{+}14, VKA^{+}23, WL15, WHW^{+}22, WCS23, Wee21, WKS03, WDH09, WTA11, YRD^{+}24, YH21, Ano95u, Ano01a, GMMP21, KB12, TPK13]. reality-based [TBS^{+}23]. Reality-Virtuality [GMMP21]. Realization [Ano12b, MSAR01]. Realizing [BN07, MT88, IKTS22]. Realtime [VFSL06, BSM^{+}22, CDPS06, CP13, YWH^{+}16]. Reasonable [YL23]. reasoning [AC90, FV06, GTG06, LS90, LL06, SA04, SAK90, TA90]. Rebuttal [CM15b]. receptive [LCGD24, ZZF^{+}23]. recipe [Gon98, Loy91]. recipes [Dol95]. Reciprocal [KTMW12]. reckoning [SFC01]. Recognition [Ano04b, Ano05c, BG80, CDF14, BWZ^{+}18, BSMG24, BPGW11, CPC^{+}18, CGG^{+}20, CGS^{+}21, CMdL21, CSLY01, CZ06, FGP^{+}10, FK11, FGL23, GRF^{+}22, GMd^{+}13, GZW12, HZ15, HWR^{+}23, JRS21, JHL^{+}12, Kas87, KY97, LH00, LMY^{+}21, MAH00, RRB^{+}22, SD08, TVL16, TC24, XWWK21, YS15, YS17, YS21, ZLLG18, ZMK18, tHV10]. recognition-based [MAH00]. recognizer [EK15, LKS07]. recolouring [WZC^{+}21]. recommending [GBG04]. reconfigurable [SEDT^{+}03]. Reconstruct [CGWZ22, CNC^{+}21, PVCM22]. reconstructed [FDGM18]. Reconstructing [ABAA22, DKY97, KC07, KH86, LNSW16, RKMP13, VH15, XSW23]. Reconstruction [CSL23, CGMS00, GZZS06, KD11, Tsu17, VP77, XXHM21, YY96, ZS94, AM91, ACA07, ACA16, ASZ^{+}14, BF02a, BD98, BKL^{+}95, CLXJ22, CLWQ09, CZZ22, CWTL08, CK02b, CZ06, CBM^{+}22, CLE13, CCCP04, DDW11, DYW^{+}22, DSR11, FAZ21, FV13, FR92a, FLM^{+}15, GDDA13, GV89, GHCH03, GS89, HKL^{+}23, HLY^{+}19, HHCM17, HGJ16, IP23, JSV98, JH89, JZX^{+}21, JWZ23, KCS24, LW24, LYL^{+}17, LW10, LCX^{+}23, LLL^{+}23, LRY^{+}24, LSS21, LBTM15, MOS^{+}21, MKC08, MPM15, MBV18, MMV^{+}14, NKMI23, NOS15, OVWK16, OK99, OK03, ODZS11, PMM18, PLVT23, PECW22, RTB^{+}18, RBF17, RCG^{+}05, SOC^{+}19, SA17, SLYY97, SHLW89, SJ09, SXW^{+}22, SLF^{+}24, SHS^{+}23, TZT^{+}22, TF18, ÜT99, VAGT08, WS12, WTL^{+}11, WFC14, WSL^{+}19, WYXM22, WRLZ23, WLB^{+}23, WPB^{+}23, XLSW22, YZ17, YBD^{+}24, ZLGH10, ZLLL21, ZPP^{+}23, ZZCL14, ZSM^{+}19, dFP22, vLvKV11]. reconstructions [ABG^{+}18, WHW^{+}22]. Recovering [PPG^{+}18, ASM12, GR09]. recovery [AG13, PHLW15, WHZ^{+}18]. rectangles [MS07, SDS89, vLvKV11]. rectangular [DS93, RJS01]. rectification [DWL^{+}03, FS20]. rectilinear [DMG20]. recurrent [HZLC22, LBD17, LZZ^{+}19, Mar02a, PR23, SH03]. Recursive [JSG04, Ran91, TJ85, Cha90, GIZ95, NG05, OdlCA02, SS22]. Redesigning [ÖOK23]. reduce [HHLE17, RHS^{+}94]. reduced [LM22, MLC^{+}22]. reduced-order [LM22]. Reducing [JX96, Lew75a, Ban97a, Cla90, DJH^{+}23]. reduction [BSJC02, CHC^{+}24, DH95a, LK18, MJEG21, MCMT14, dATNMC^{+}22, Smo03, WLYH19]. Reeb [SJ15, TSK98]. Reel [Cou92b]. Refereed [FEJ04]. reference [FB04, GMNS93, HMdM^{+}95, Lod21, Pfa84, SC97, SCFF16, WLX^{+}23, ZSS20]. referenced [SA04]. Refinable [KP19, SP16]. refined [GBKG04, SBD15a]. refinement [CFZC19, GM07, GAÖ02, GBP04, IORM17, KP22a, rKC93, KWK23, LZZ^{+}19, LWD^{+}18, MMS^{+}17, NTAI20, RCPB01, WFG03, WYZ20]. refining [HG22]. Reflectance [HOCN07, MMK04, OKBG08, KWK17, WPH^{+}14]. reflecting [LhCE97]. Reflection [KJTS96]. reflective [HS08, LRD07]. reflector [MMP18, PPV07]. Reflexive [BMdSVR18]. refraction [GM02]. refresh [Coo79]. Refreshing [Bas77]. regard [COPR17]. Region [TC00, dFP22, CTQ^{+}14, CGZZ15, CN05, KT17, LSZQ21, LCCM02, LZ24, SGZ10, SHS^{+}23, XD08, ZZXT18]. region-based [CGZZ15]. Region-filling [TC00]. region-growing [XD08]. region-piecewise [LZ24]. region-swapping [ZZXT18]. regional [AS95, BPS03, RŞH^{+}22, SRA^{+}19, SHL^{+}24]. Regions [Gab77, SB94, ALD12, Fri03, HPKE19, LL92, LP92, MMdOE^{+}22, RS08, RS99, RMW^{+}17]. Registration [FIC21, H^{+}00b, BCDD22, DZD^{+}23, HZD^{+}19, LLM^{+}16, ME92, PM84b, ST20, SDT21, ST22, SKS09, WCH^{+}24, WH96, YZWJ24, ZZH^{+}24, ZSL12, ZTF^{+}22, ZDL^{+}24]. regolith [LZY^{+}23]. regression [BRPC18, JZX^{+}21, LTP19, MENS19, YZC^{+}23, ZSS^{+}18, ZHG^{+}21]. regression-based [ZSS^{+}18]. regular [KPA12, RPSP^{+}19, SWMdF21, Van09, Wil03, WA02, YBTB23]. Regularised [TPG99]. regularization [JWZ23, WLL22, WLL^{+}21a]. regularized [LTS88]. regulations [COM^{+}94]. Rehabilitation [Ano04-43, Ano04-42, CIT^{+}19, LLW^{+}23, RHM^{+}12, SCB^{+}24, WJG^{+}19]. Reinforcement [MC23, KKPC23, MH21]. reinterpretation [BL11]. rejection [LWP02]. relatifs [Le 77]. Relation [LGWM17, CDF14, HZLQ20, LDS^{+}21]. Relation-based [LGWM17]. Relational [MD85, PR82a]. relations [Dom77, LSS97]. relationship [KS91, MR05, ME92]. Relationships [HYP^{+}20, IA91]. relative [LYW^{+}13]. relatively [Sla92]. relatives [FN10]. relaxation [Liv18]. relaxed [VLD15, WYXM22]. release [YCO23]. Relevance [RYNJ23]. reliable [FPC10, dATNMC^{+}22, XLL^{+}20]. relief [GTFB23]. reliefs [TBG^{+}20, ZCL^{+}18]. relighting [WLQC18, ZXD^{+}14]. remapping [MWA^{+}13]. remarks [Sar02]. remesher [FTB16]. Remeshing [AYZ12, CFZC19, GYJZ15, KP22b, MMM^{+}20, PJ21, VLD15]. remote [BNS24, CSFG96, Gia15, MTS^{+}22, MST^{+}22, Mil03, THL15, TKB07]. Removal [AP88, HF85, MBN21, Eas75, HH91, QSLS23, SJT20, SPGR93, SD21, SLQ^{+}19, WZZ^{+}21]. removers [LCL^{+}23]. rendered [CAAC20, DZ91]. renderer [FWW13]. Rendering [Ano03y, Ano12s, CCL^{+}03, CI11, PMPR^{+}16, Pic88b, PPZ^{+}10, SK98, WZ21, AALJ20, AT08, AJAC23, AAK^{+}22b, ACA16, ABCO12, ACGC22, BGK04, BLZD12, BDSP22, BAG03, BGMP08, BK23, BSL^{+}13, BC01, BB98, BRV^{+}10, BCC10, Car96, Car99a, CF11, Cha90, CVHM03, CCC^{+}16, CYWM23, CLT07, CSSC00, CL95, CRD10, CP13, DTWT94, DAHF04, DW05, DK92, Den90, DR15, DM01, ESAH02, ERWS12, ERDS14, FEVM10, FBP96, Fun99, GM07, GO10, GPTB02, Ger02, GM04, Gro94, GLDK95, GLDH97, GBP04, GCRR11, GZL21, GS04, HO12, HGVV16, HJDR95, Hol94, IP03, IL97, IIK12, IVCN20, JK15, rKC93, KS01a, Kni95a, KK16, KJS17, KJS18, KPMT18, KAÖ98, KYM12, LMR^{+}19, Lee01, LWW08, LZW^{+}21a, LLC04, LHCL05, LCD15, LJJPF22, LS08, MLM03, MKC08, MPP98, Mar10, MHZ94, MMK04, MCKS06, MIGS22, NIH08, NAS18, NT00]. rendering [NLSN11, Oik98, PdSMdML^{+}12, Pic88a, PM95, Rok93b, RVdF08, RJG06, RGMJ22, RdMF08, RBFS10, SP04, SKKN10, SR89, SMK08, SDD95, SZEG93, SVW23, SWvB95, SF92, SKCP99, TP89, TKOD24, WZW97, WPH^{+}14, WJGG15, WKE04, XXX^{+}23, YK23, YK18, YD00, YWH^{+}16, ZK07, ZP07, ZLLY06, dHG^{+}97, dGHM97, dAPG18, dVTT18, vWS04, Ano93a, Ano94d]. renderings [CST05, Gin02]. rep [SGZ10, MRF06, KDG96]. repair [CG96, TBS^{+}23]. RepDehazeNet [LZL^{+}24]. Repeat [Ano04-33, Ano06g, Ano06-28, Ano04-57, Ano04-60, Ano06h, Ano06p, Ano06y, Ano06-44]. repeatability [FBH^{+}21]. repeatable [RPSP^{+}19]. repertoire [FvdPT01]. replace [DN22]. replacement [QWC14]. replacing [CR07]. replicas [GMM^{+}23]. replicating [KSF15, RHC15]. replication [PCS00]. report [CL96, KW96, NGAS23, OL96, RG93]. repositories [VT07]. repository [HS05]. represent [Car92, LMZ90, YRD^{+}24]. Representation [BS82, LDLD22, Pre84a, AA01, BA90, Bri03, CDW11, CZR22, CTLG94, CAL^{+}04, CD23, CBM23, DGLRD18, DJG^{+}04, DMG20, DWX^{+}23, EPS96, Ela86, FWD21, FR88, GCW23, HKL^{+}23, HG02, ICNV21, JWZ23, KM14, Koh96b, KWK17, KLL^{+}15, Lam99, LGZ^{+}21, LZC^{+}15, LP83, MVS14, Mel19, Mes00, MLPB02, MS90, NLG20, Nav89, PL97, PMS87, Ric89, RFR02, SSB04, SBS19, SBD15a, Sob89, SWMdF21, SJ15, ST23, Szy92, TSK98, TW24, TBDC20, WKT21, WCW^{+}24, YBTB23, ZY02, ZSW08, ZSK95, Zhu91, dF24]. representational [JDGS88]. representations [BAS^{+}20, CW03, GCRN23, Jon90, MJEG21, MHCL15, MR96, Rei98, Roj91, SL16b, WLC88, Wu89, YK18]. represented [LYXY19, MMS89]. Representing [AF89, DE92, TPK13, dAU14, Wal93b, XH88]. Reproducing [VVCN12, MAdS^{+}19]. reproduction [AEA13, CLMA19]. reps [WKW16]. requantization [HCC91]. required [IM07]. Requirements [Dur91, PM84a, GS83]. resampling [ASPO15, Lar03, SOC^{+}19]. rescue [MMSS03]. rescued [Bor91]. Research [All77, Ano07e, Ano07d, Sch86a, Ban97b, Bro92, EGL^{+}95, FF96, FCSB90, FEJM75, Jun94, Nov03, RW87, SB97, ZBP^{+}18]. research-driven [Nov03]. Researcher [WS22]. réseaux [Le 77]. reservoir [CC20, LG23]. residual [CHC^{+}24, HZ23, HZLC22, MYL^{+}23, SLL^{+}23]. residuals [SWZZ23]. residuals-distribution-guided [SWZZ23]. resistance [XXHM21]. resistant [vBT20b]. resizing [WZ09]. Resolution [GCYX23, KS84, BWD13, BBE14, CCCS08, CF11, C^{+}01, CDI12, EVRW23, GS04, HW22, HXH24, KL02, LTC^{+}20, LZW^{+}21a, LLC^{+}22, MAF19, Mul01, NLdAL^{+}23, NUM24, RLS^{+}12, SVW23, SLL^{+}23, TW24, YZC^{+}23]. resolved [PAJ19]. resolving [MAH00]. resonance [BL11]. Resource [FN99]. resources [HL02]. respecting [MC10]. Response [ASC18, AG13, CTJ^{+}14, FBH^{+}21, GO06, HLCF88, Kam93, LW89, MML12]. restitution [MPAC^{+}23]. Restoration [CFMP84, RM22, LYZ^{+}22, Pop93]. restoring [WB24]. restricted [Ang97, MTM22, YRS^{+}18, dGWvdW09]. result [SBKB23]. Results [IWM^{+}09, ABCD93, BA90, GCW23, Qui91, YRD^{+}24]. Retargeting [BWBM20, LWJ^{+}22]. retexturing [ADT^{+}16, LSH^{+}12]. Reticulated [Har00]. retiming [YMZ^{+}15]. retinal [ZWS19]. Retinex [LHG21, WCHM22]. retouching [LTH^{+}19]. Retrace [Moh77]. Retrieval [AOL96, BTD^{+}22, FGZ^{+}22, LPD^{+}18, MDM^{+}21, RFB^{+}21, STP^{+}20, SLL^{+}21, TT19, TBG^{+}20, Ano94-29, Ano94-30, Ano13o, BFLP20, BDL^{+}22, BRPC18, CUD06, EHBA10, FJ03, HR07, JZLP23, KKJ^{+}23, LPL^{+}20, LNL^{+}23b, LNL^{+}23a, LKL^{+}20, LLLZ16, LKLW16, LG03, LXB^{+}15, LB19, MKDM22, MKM19, PPD22, QYC^{+}22, RFZ^{+}17, SPT18, SBS13, SJZ^{+}23, TKD16, WLP^{+}14, WLYH19, WWK11, XZPG21, VTW23, vBT20a, vBT21, Ano07h, Ano07i]. Retrieving [LWD^{+}18, TB19, CWL20]. retroreflective [ARM23, Dor99]. retroreflector [OUZS18]. retrospective [Duc18, WKS03]. return [BF07, Wes94]. reuse [CSH08, JZLP23]. reveal [RBLB21]. Revealing [CVL^{+}04]. Reverse [DVF06, FBR^{+}17, FM00, Wal06, Yin04, LLX^{+}15, MRSS^{+}18, SS09]. Reverse-engineering [FBR^{+}17]. reversible [KB06]. Review [Mag85, AAB18, Ben79, Cou92a, GTTC03, KCU^{+}22, MPW^{+}12, MHK99, NLOdS23, NGAS23, ST20, dSMBG23]. Reviewer [Ano11b, Ano13b, Ano23a]. ReviewerNet [SGPC20]. Reviewers [Ano97e, Ano99i, Ano06-36, Ano09j, Ano10j, Ano11n, Ano15l, Ano16i, Ano17i, Ano19k, Ano20q, Ano21s, Ano22t, Ano94w, Ano95-32, Ano01-48, Ano03-38, Ano04-44, Ano05-33, Ano15m, Ano24a, SGPC20, Ano07-32, Ano13p]. revised [FPC10]. revisit [BLW12]. revisited [DL09, JR00, SM92]. revolute [SJ94]. revolution [EH96, FS86a, GBF14, Pea02]. rewriting [BBP13, MP19]. RGB [CMLH21, dSJdML18, LWWY22, LLL^{+}23, TRB^{+}22]. RGB-D [CMLH21, dSJdML18, LWWY22, TRB^{+}22]. Rheumatic [NGAS23]. RIAO [Ano94-29, Ano94-30]. rich [GD04, HFP06, HL02, IM07, ZSL12]. Richardson [Wal94b]. ridge [NZL^{+}21]. ridge-valley [NZL^{+}21]. Riesz [FO21]. rigging [MZ23, NTAI20]. right [McW91a]. Rigid [GÖT93, AOB17, ADOR02, BCDD22, CL18, CBM23, DKV^{+}22, GO10, LD05, PF97, SJG19, ZEK^{+}17, ZSL12]. rim [SBD15b]. ring [LCGD24]. ringed [ZQ12]. rings [And94]. Ripoll [ACB12]. Rips [Zom10]. RIST [KHS^{+}10]. Rivers [LH83]. road [TRB^{+}22, WJG^{+}19, WBJ^{+}21, WWH^{+}21]. Robert [Kre93]. Robot [FS78, Per84, Car92, LQOW08, MH21, TL13]. Robotic [Dai93, PGS^{+}23, GA07, JVS^{+}24, MDS^{+}21, MDS^{+}22]. Robotic-Assisted [PGS^{+}23]. robotics [Ano94u, Ano94v]. robots [SBD15b]. Robust [DYW^{+}22, Dor99, ISPS17, JL23, LDM^{+}11, LSK^{+}10, LHG21, LLL^{+}23, LOdF02, PdFS06, Sug83, WLDB11, WMDR08, XXHM21, YPSZ01, ZSL12, ACG15, BWZ^{+}18, BK02, BMT96, CMdL21, DZD^{+}23, KYKK19, LGRP14, Liv23, Smo03, WWK11, WWCZ19, YZWJ24, ZS02, VTW23]. rodent [ABG^{+}18]. rods [MBC^{+}23]. ROI [LWD^{+}18, VW21]. ROI-based [LWD^{+}18]. Role [Web85, ADHC^{+}23, CSM^{+}01, FR92b, GMM^{+}23]. roll [Yue86]. room [MMV^{+}14]. root [CM15a, JALS03]. roots [Gil94]. rotated [MSF95]. rotating [OK20, SB94]. rotating-screen [SB94]. Rotation [EKP93, Baw97, MG09, XCL^{+}19]. rotation-invariant [XCL^{+}19]. rotational [K^{+}00a, MA94]. rotations [HL93]. rotoscopy [LPV92]. rough [PPM18]. roughness [WTM12]. roughness-based [WTM12]. round [Bor91, EHSF17]. round-off [Bor91]. routes [GSV^{+}18]. routines [CK75]. routing [PGR83]. RP [Shi06]. RP-aided [Shi06]. RPR [PdSMdML^{+}12]. RPR-SORS [PdSMdML^{+}12]. rubber [AGM^{+}21]. Rule [Fen88, LUB^{+}13, BPKG07, JAS97, KP22a]. Rule-based [Fen88, LUB^{+}13, JAS97]. ruled [Aum89, GD87, PR82b]. ruler [LHS87]. rules [BDKK96, LPPM07, Wei99]. Ruleset [MP19, BBP13]. Ruleset-rewriting [MP19]. ruling [SM22]. ruling-aware [SM22]. Run [YR98]. Run-length [YR98]. Running [SL01, LRMS92, MLCMGR23]. runs [SL01]. runtime [RLD^{+}12]. RUS [LLR93]. rush [Jac93]. Russia [Ano95v]. Rutherford [HRGD88]. RVCA [GMMP21]. RWM [YL96]. RWM-cut [YL96].

S [Cyc93, Mil92a, Ano99d, Ano99e, Ano99f, Ano99g, Ano00a]. S3D [MY16]. safety [RCB15]. SAH [WG17]. Sal [BBMGM22]. Saliency [KRA^{+}23a, KRA^{+}23b, LZYQ22, SBE20, BBMGM22, BAC14, CACC24, CAAC20, DSNW13, GBA15, HZLC22, JDT^{+}23, KCH^{+}22, LLS^{+}16, MFP11, PP16, PWV^{+}18, SKH^{+}12, TCL^{+}15a, TKD16]. Saliency-aware [LZYQ22]. Saliency-based [SBE20, TKD16]. saliency-guided [CACC24]. Salient360 [DGV^{+}24]. SalientGaze [SBE20]. SAM [MK85]. same [ACO12]. Sample [rKC93, FB14, FB15a]. sample-based [FB14, FB15a]. sampled [CCCS08, DW13, GM04, GS11, OMW13, SHD^{+}17, WW06]. sampler [LZY^{+}23]. samples [FK11, KJS18, WLG04, XCL^{+}19]. Sampling [DG96b, SG92, TMN^{+}00, BA09, GYJZ15, Pur87b, SNS06a, SLF^{+}24, TNU^{+}01, WPH^{+}14, XHGL12, YWC22, ZZQW11, ZGZ^{+}16, HHZ^{+}22]. San [BL11, KGM75]. Sand [ZYML23]. sandy [KB10]. Sanitary [Rob78]. Sankey [XZY^{+}23]. Satellite [Ber79, JD75, Kir93, Mey79]. satisfaction [Des00, Doh95]. save [CYKK09]. saving [CWC^{+}14, CCC^{+}16]. sawing [OT88]. SAXRegEx [YBTB23]. SBC [Ano01a, Ano01f]. SBGames [MTB18]. SBIM [Ano12s, vdPS08]. scaffold [LUB^{+}13]. Scaffolded [QLF^{+}09]. Scalable [BGMP08, Gir93, G^{+}01, KKHS03, RVdF08, IWM94, JHL^{+}12, KLW12, Kni95a, LPP^{+}19, SHCW22]. scalar [BPS^{+}10, DMG20, DKLP02, ID17, KBL22, MVG^{+}21, MW14, PF09, PS12, REG^{+}89]. Scale [Ano12b, KRA^{+}23a, LCL15, MALI11, ALC06, ASZ^{+}14, BLZD12, BRdSOS17, BYQZ22, BAS^{+}20, DVG^{+}18, DR09, EHBA10, FTB12, Gia15, KM21, KRA^{+}23b, MAG^{+}12, PP16, PRBD22, PJJSH16, RSP^{+}19, SKS09, SLL^{+}23, SXW^{+}22, TF18, WHZ^{+}18, WWH^{+}21, WCL23, WUH^{+}15, XZL^{+}22, YWC22, ZMH^{+}23, dVTT18]. Scale-dependent [MALI11]. Scale-space [LCL15]. scaling [FPR92, KHK18, RPM97, Sah15]. Scan [GP86, Neg77, AAK^{+}22b, Eas75, GL83, Kau88, LCK16, Lin79, YR98, YHX10, ZM07]. scan-based [ZM07]. scan-conversion [YR98]. scan-converting [Kau88]. Scanline [BB93a]. scanned [NTAI20, RBF17, Tsu17]. scanners [WBB^{+}08]. Scanning [Bri03]. Scanning-curves [Bri03]. scans [BWdBP13, GR09, MPTA^{+}22]. SCAPE [ZZCL14]. SCAPE-based [ZZCL14]. scapulo [MT00]. scapulo-thoracic [MT00]. scattered [AM91, KHS03, RNM95, SSK87, WWF^{+}18, XZ00]. scattering [LW24, ZK24]. scatterplot [MJEG21]. scatterplot-based [MJEG21]. scatterplots [MM18, vOHR20]. SCCG [SD16, VB17]. SCCG’2015 [DS15]. scenario [CKK96, LLZ^{+}23, RŞH^{+}22, RHFL14]. scenario-universal [LLZ^{+}23]. scenarios [JP10, JJPP^{+}22, MAM^{+}24, SEMWC05]. scene [AMPG22, AAB18, Bou02, BG91, CB01, FBTT^{+}22, KD00, LD05, LWW^{+}20, LZL^{+}19, LCX^{+}23, MCKS06, NZZ^{+}21, PRBD22, RFZ^{+}17, RO87, RVR04, SLX^{+}16, SvLBF10, SGBI02, SXW^{+}22, WZZZ18, WUH^{+}15, XWY15, ZZF^{+}23, vLvKV11]. scene-scale [PRBD22]. Scenes [GM86, MTT84, NH83, Sug83, ACV03, AEA13, BCC20, CCL^{+}03, CCI12, CHSD95, GYK^{+}23, GH98, GLC20, KVB^{+}20, KGB^{+}21, LD12, MPA^{+}10, NFLYCO99, PdSMdML^{+}12, RCBS10, SSB04, SD15, WIP08, WPH^{+}14, ZLM^{+}15, ZGS17, ZML^{+}18]. scheduling [SFD06]. schematic [Men85, SSM87]. Scheme [PMPR^{+}16, CP98, CWC^{+}14, CCC^{+}16, GT91, GBKG04, KD11, KONS17, Liv18, MPL02, MCT08, MS90, NG05, RP18, SEDT^{+}03, vBT20a]. schemes [BLNZ22, BF15, BD98, LH14, SK04]. School [Ano01-66, Ano03a, DZZ79, KH86, Ano91c]. Schröder [Dra98]. Science [GA88, Lau77, CLF^{+}06, JMC^{+}04, Kal04, Owe94, She12, Smi77]. Sciences [Cib77, Le 77, RFS22, CV77, Nan77]. Scientific [Ano94e, Ano95-31, CM93, Dom93, LLR93, WCdA98, vDLS02, Beb75, Bry93, Gin93, PdSP^{+}22, Wri75, YWC22, YHW23]. scientificalness [ZXH^{+}12]. scintillation [SVVS^{+}17]. sclerosis [SBR^{+}22]. scope [dos01]. Scores [BSW78, GC86]. Scouts [SS89b]. Screen [CXCH23, AEA13, DK97, GJN^{+}21, JHPhR11, KFW16, MB97a, SB94, TMSPB09, XXX^{+}23]. screen-space [KFW16]. Screened [XXHM21]. screening [WPL^{+}23]. screening-based [WPL^{+}23]. screens [PWJ^{+}18]. screenspace [KJS17]. Scribble [BF19]. Scribble-based [BF19]. Script [IC96]. Script-based [IC96]. Scripta [CUD06]. scrollytelling [MGM^{+}23]. ScrutinAI [HMA23]. Sculpting [SCC11, SCCS13, Kru99a, MDS^{+}21, MDS^{+}22, SM98]. sculpture [SCVCN16]. sculptured [NS87]. Sculptures [Ano12b, AXG^{+}13, PJP23]. SDS [PM10]. sea [CS03]. Seamless [RGE07, WL15, ZZCY22, HLO16, KWK17]. search [ENE11, HGJC21, RM91, SBKB23, TKB07, WW08, YBTB23]. Searching [KSF15, Rei75, BWZ^{+}18, Las91, dGMW16]. Seashells [Cor93, Ste09]. Seaweed [CG93]. Secant [Wal94b]. secants [Jon91, Szy90a]. Second [Mil92b, RB06, Ano94c, HG02, Hod91, LM16, Bou09]. second-order [HG02, LM16]. secondary [VCQ92]. Secret [TL02]. SECT [BC88]. Section [ATB98, AF16b, AEJZ18, Ano11m, Ano21i, Ano22j, Aok16, AMS22, ABK11, BX99, BK19, BV22, BTD^{+}22, BGT16, BPH20, BYC19, BSB^{+}02, CJO19, CP19b, CMS11, CM18, CR99, DKFC20, DS15, EK18, FH11a, FH11b, FR98, FGMR22, GM02, GPTB02, GPS18, GW17, GMMP21, GHK18, GZ99, GB18, GMM18, Gra02, GAÖ02, HSTR20, IA99, JZ17, JD99, JKK02, JT02, Joh18, Kiy19, KB02, KPBR20, Kry14, KNDT20, LPD^{+}18, LV02, LC18, LPS19, MW99, MRR98, MD99a, MPP98, Mar02a, MB10, MF02, Mou15, MTB18, NTB18, NTTF21, OP19, OK99, Oik98, OOC22, PS02, PTW98, PBB21, PK17, PB11, RLB^{+}02, RFS22, RS22, RGJdQ18, SK22, SF98, SPO22, SCG23, SM98, Sez16, SNB17, Sif99, SD16, SUF^{+}18, SA19, TK20, TT19, TL02, TSC16, TP17, TKdJO22, VH02]. Section [VL98, VB17, VT22, WEWL99, WBK98, Yu99, ŽC99, ZLS99, ZWSW22, dMF99, ACP20, Ano83a, Ano86b, Ano86i, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, Ano12l, Ano12o, Ano12m, Ano12n, Ano12s, Ano13n, Ano13l, Ano13m, AdSMD23, BA23, BFLP20, BLM23, BCMM07, BP08, BKS21, BKS23, CHPS20, CMS12, CI11, CH15, DK24, EZ22, Enc85a, Enc85b, Enc86a, Enc86b, FT24, GG96, GSKZ22, IK21, JZZ16, KKN^{+}21, LS19, LC23, Lin10, Liu21, MS16, Mar02b, MB14, MCPW21, ND23, NN15, PB22, PB23, PDK16, RSK^{+}24, Rei95, RMSB22, SR02, STP^{+}20, SPK19, SC12, SK12, SSS15, TMH20, VHE10, ZT16, ZM92]. sectional [Woo87]. Sections [VP77, ATB98, HGJ16, KD11, Roc89, SA17, Xu08]. security [Gra02, HWEB22]. sedimentation [CBS^{+}14]. see [IKM^{+}20, KKO01, KSH17]. see-through [IKM^{+}20, KKO01, KSH17]. seed [FS98, Oik98, YHX10]. seed-filling [YHX10]. seeding [BH15, VW21]. seemingly [Elb11]. segment [MHCL15, SF98, TT82, XhKKL16]. segment-polygon [SF98]. Segmentation [ACC^{+}18, GOdSC23, JA84c, AMFH21, ATHL14, AAK^{+}22b, APB07, BGLA18, Bus97, CMLH21, CFZL16, DEST95, GF09, GLZ^{+}21, GWL^{+}20, HCX^{+}23, HKHP11, HG21, HS14, HZLQ20, HG22, HZ23, HXH24, HLXL23, JZLP23, JXW^{+}22, JRJP^{+}22, KGK^{+}07, KVB^{+}20, LBD17, LWS15, LTC^{+}20, LTBZ13, LZL^{+}15, LXJL21, LSCJ23, MRSS^{+}18, MFL11, MRWL23, NLOdS23, PCWD23, RBG^{+}09, RMG15, VZP22, WS12, WY11b, WGS^{+}18c, WGS^{+}18a, WGS^{+}18b, WSHY22, Wan23, WLQC18, WWS^{+}13, WSWL14, WCLZ14, WLL21b, WYC^{+}23, XLXG11, ZWS19, ZJSB22, ZLG^{+}15, ZDL^{+}19, ZZJ^{+}14, ZMH^{+}23, dCdLL14]. segmented [GTG06, PPS20, RGdL^{+}18]. segmenting [DJH^{+}23, HS11]. Segments [Pie84, DRFRD06, LYS^{+}16, WHH06]. seismic [Arb92, PSH^{+}09]. Selected [Ano22j, BF02b, Wir80]. Selection [De 84, Row82, AL10, AA13, AM19, HCC13, MYL^{+}23, MMS^{+}17, Oli08, RO13, SGPC20, SW11, VFSL06, WSL^{+}19]. selections [JPCS18]. Selective [WFG03, DTG15, LPR^{+}14, SMSS13]. Selectively [ZSK95]. Self [DBS^{+}18, GMNS93, HMHB08, LM89, NGAS23, PM13, ZYW23, VTW23, AM12, BB98, CJXZ23, HK04, HHK^{+}13, HPKE19, JWZ23, Lan97, LKC98, LHY23, LYS^{+}16, MMdOE^{+}22, MRWL23, PMTK01, PBN97, RHC15, SK03, SGC^{+}19, SCC11, WTWT18]. self- [PBN97]. self-adaptive [SCC11]. self-affine [Lan97]. self-attention [MRWL23]. Self-attribution [DBS^{+}18]. self-collision [LKC98, WTWT18]. self-driving [SGC^{+}19]. self-folding [HHK^{+}13]. self-intersections [HPKE19]. self-learning [MMdOE^{+}22]. self-occlusion [AM12]. Self-organized [PM13]. self-organizing [HK04, SK03]. self-overlapping [LYS^{+}16]. Self-reference [GMNS93]. self-replicating [RHC15]. Self-report [NGAS23]. Self-similar [LM89]. Self-similarity [HMHB08]. Self-supervised [ZYW23, VTW23, CJXZ23, JWZ23, LHY23]. self-visualizing [BB98]. Semantic [CLH12, EAAY23, FH11a, FH11b, JRJP^{+}22, LXPP06, RGE07, ZjLW^{+}14, BGLA18, CMLH21, CGWZ22, DBS^{+}11, DWX^{+}23, GKW^{+}24, HMA23, HZLQ20, HG22, HZ23, HXH24, HLXL23, KVB^{+}20, LBLV16, LPD13, LYXY19, MK83, MRWL23, MPS06, RLU^{+}19, SJZ^{+}23, VZP22, WCL23, WCF^{+}23, WYC^{+}23, ZMH^{+}23, ZC07, GD11]. semantic-based [DBS^{+}11]. semantically [HHKF10, RPP21]. Semantics [CMSF11, GRIG12, MS09b, CUD06, Fiu87, Mar80, Mar82, YMYH12]. Semantics-driven [MS09b]. sémantiques [Dom77]. Semi [LD09, LZKJ23, ÇB22, CUD06, GWP00, HLXL23, KPA12]. semi-adaptive [ÇB22]. semi-automatic [CUD06]. semi-immersive [GWP00]. semi-regular [KPA12]. Semi-sharp [LD09]. Semi-supervised [LZKJ23, HLXL23]. Semiautomatic [AR84]. sense [BMdSVR18, LUMC04, RBLB21]. sensing [CSFG96, Gia15, JYLW14, THL15, TKB07, YSZ22]. sensitive [CMD99, FT02, MMGB17, WLZL18]. Sensitivity [Fra94]. sensor [Bim15, BS09]. sensors [CBM^{+}22, ISPS17, MMSS03, Nar15, SHBSS17]. Sensory [AM10, HKCL02]. sentence [LHL23]. sentence-word [LHL23]. separable [HXH24]. separated [PAE09]. Separation [Fri03, BHZ^{+}21, PMBS14, SS75a, DW89]. September [SIG02]. Sequence [BWZ^{+}18, CZ98, CLH12, CXT18, LDLD23, PPD22, WM24]. sequences [AMHWW16, CH96, GHCH03, Jef92, KG04, K^{+}00a, LM89, MRG^{+}19, Ric89, TC24, vdLdFvdEV23]. sequency [Egh80]. sequential [KM14, ME92, PM84b, TSK98]. serendipity [Wal93a]. Serial [VP77]. series [Ano06-33, BSM^{+}22, Cha97, DSR11, Lan97, SHD^{+}17, TKB07, YBTB23]. Serious [LC18, DPS10, SBHC22]. server [JHL^{+}12, RBP96, ZR97]. server-client [JHL^{+}12]. service [CG96, GD11, HKS01, LMD96, SKSZ99, SSW^{+}23]. services [KPB96, Yam94]. session [BLS15, Ise21, IB22]. Set [Ano95q, Ano95r, Ano95s, FGZ^{+}22, Hoo91, ARM23, CCCS08, Car99a, DM00, Eng93, Ent89a, EB10, FR92b, FPR92, FC00b, GBP04, HJW^{+}08, HMHB08, Jon90, KYL15, Les01, LTS88, MPM15, MKPM17, MTM22, NSL16, NLOdS23, ODZS11, OdlCA02, PLJ^{+}13, PRÁM04, Phi92, PFR94, Rad96, Roj91, RPÁM04, RPAM06, RAA^{+}08, STN95, SLG97, VM15, WTL^{+}11, YRS^{+}18, Žal01, ZTF^{+}22, ZXD^{+}14]. Set-theoretic [Ano95q, Ano95r, Ano95s]. Sets [ARL78, She93, AOB17, ACA07, ADR01, CHL^{+}11, Cas96, CZC02, CCC00, CRT04, Coo01b, Des00, DSB96, DMT03, FLV20, Gin02, HJDR95, HPD^{+}10, ITW18, KD11, Lak91, MR92, Mil93, Nor89, OMW13, OdlCA02, PS13a, PCGS15, RA03, SKKN10, SK03, VT06, VP06, WSL92, XSW23, YWR03, YIL09, ŽJŽ03]. Setting [SGBI02, LQ15]. settings [GMM^{+}23].

Seventh [Ano02-50]. Several [CG93]. severe [CMdL21]. SFINGE [CGG^{+}20]. SGML [CG87]. shaded [BF19, Kle86]. Shader [CYKK09, FWW13, SMSS13]. Shader-based [CYKK09, FWW13]. shaders [Bai07]. Shading [BP93, BK89b, PA91, BERW97, BKL15, Cla90, DH95b, HT96, HO94, KTMW12, LP92, OM96, PH90, RR15, SDD95, SMSS13, Zha96]. Shadow [IKM^{+}20, WXZ^{+}18, WZZ^{+}21, CC19, GSSK^{+}13, JV91, MA18, MVCNI21, WLX^{+}23, WWWW22]. shadowing [PA91]. Shadows [HN85, BC13, DLZY14, XLGS16]. shaft [LWW08]. Shah [WLL22]. shallow [CD15, DVG^{+}18, GM02]. Shape [ATAG^{+}21, ALR23, Ano01-54, Ano03-52, BPS^{+}10, BMS^{+}11, CYJ^{+}13, CK02b, CSLY01, DLR^{+}20, FV06, FIC21, FSP15, GTG06, GHK18, GS11, HT96, Les02, LL91, LWLT11, LLS^{+}16, MNS^{+}19, NC12, NTAI20, SHBSS17, TKZ^{+}13, VL98, WKW16, ZLG^{+}15, AK21, ACG15, ACO01, Ano05-40, AHR^{+}22, BLL15, BKR^{+}16, BWBM20, BWD13, BYQZ22, BHL^{+}15, BDK17, CP21, CXCH23, CLH^{+}16, CMdL21, CW03, CWL20, CCM^{+}18, CBNJ^{+}15, CZCG04, CBM23, Com01, E^{+}00, FCM^{+}18, FIC23, GBF14, GR93, GLZ^{+}21, HCV^{+}22, HSR^{+}09, HK15, HBM23, HO94, HHKF10, HHN^{+}23, HYP^{+}24, IY18, Kle86, KLL^{+}15, KD15, LPL^{+}20, LGWM17, Les01, LL05, LL06, LS07, LLLC11, LYW^{+}13, LQ15, LKLW16, LW99, LS06, LXB^{+}15, LB19, LBTM15, MSE20, MENS19, MSMK19, MBST22, MSL^{+}19, MFP11, Mon87, Par93, QYC^{+}22, RB06, RKMP13, RE22, RBG^{+}09, RR15]. shape [RHC15, RCG^{+}05, Sah15, SFVP13, Sar94a, SAMA97, SRZK23, SS82, SBS13, TKD16, TY24, VPBY02, VR16, WJ91, WGS^{+}18c, WGS^{+}18a, WGS^{+}18b, WLYH19, WPB^{+}23, WSWL14, WCY^{+}20, WYZ20, WSX12, XCL^{+}19, XCZP14, XWWK21, YF09, YYG16, ZLL^{+}15, ZW20, ZQL15, vBT20a, vBT20b, vOHR20, vRESH16, Ano13k, Ano15k, ALP17, BMR23, BLZ22, BMP22, BPH20, BLM23, CP19b, SH12]. Shape-A-Getti [BMR23]. shape-aware [CP21]. shape-preserving [ZQL15]. shaped [Fat01, Fat02, SBH01, ZYW23]. ShapeGraMM [SBR23]. shapes [AKF^{+}20, ABJ90, ATZM19, ABM^{+}06, BMR23, BPGW11, BDK17, CDW11, DRFRD06, EPB^{+}19, KKJ^{+}23, KKH23, LBB11, LBB12, LSGFRC^{+}13, MMM^{+}20, MS01, MS09b, OKT01, PMZS97, PS12, PS13b, PCV16, PVCM22, RBF17, RBB^{+}11, SBWS11, SJG19, SSBT01, TSK98, WWS^{+}13, WZZ^{+}18, ZZC^{+}20b, dGGDV11, vdBB07]. shaping [Ano04-55, FCW^{+}10, MNSJ99]. Share [Pot78]. shareable [How88]. Shared [AHK03, GP86, Gre96, BM03, Bro00, BOH97, CSM^{+}01, KKŽ04, RSAF18, SHOC23, SvLBF10]. shared-scene-graph [SvLBF10]. sharing [CSF20, COSEV22, JY98, Mey79, OST^{+}16, TL02]. sharp [LD09, LSK^{+}10, SJ09]. Sharpness [RMP19, CC08]. Shear [SGC00]. Shear-Warp [SGC00]. sheared [PWK95]. sheet [CGWW16, YL23]. Shell [XZY^{+}21, LYC^{+}15, LLW^{+}19, ZIP^{+}19]. shells [vEB98]. shift [H^{+}00b]. shifted [RLS^{+}12]. shifting [STT^{+}18]. shifts [PQ10]. ship [Cor84, MR92, RJKV12]. ShipShape [FASS16]. Shock [PB96]. shooting [BSB^{+}02, HP03, LZ14, SKM98]. Shop [MAM^{+}24]. short [WB24]. shortest [SK23]. ShortStraw [XL10]. ShortStraw-based [XL10]. shot [GOdSC23]. shoulder [MT00]. Showcase [BFSE03]. SHREC [CGS^{+}21, ECG^{+}22, GRF^{+}22, GYK^{+}23, GCvdS^{+}20, KVB^{+}20, KGB^{+}21, LPL^{+}20, RFB^{+}21, RRB^{+}22, SLL^{+}21, TBG^{+}20, TRB^{+}22]. SHREC’20 [DLR^{+}20]. SHREC’21 [AHR^{+}22]. SHREC’22 [FGZ^{+}22, QYC^{+}22]. Shrinking [PSP^{+}20]. Shroud [Pic88b]. Shrubbery [ZIP^{+}19]. Shrubbery-shell [ZIP^{+}19]. shu [YP05]. SIAM [Ano03-49].

Siamese [HDZR23]. SIBGRAPI [AF16b, Ano01-27, Ano01-55, Ano01-56, Ano02q, AdSMD23, FT24, LS19, MS16, MCPW21, NN15, Oli08, PB22, RGJdQ18, TP17]. SIBGRAPI-Conference [MCPW21]. side [EBC^{+}15, McW91a, RTB^{+}18]. side-view [EBC^{+}15, RTB^{+}18]. sided [KP15, KP22a, PGVACN06, SV18, SVV23, SVSV20, Tar22]. Siegel [Mil93]. Sierpinski [Maj98]. Sifting [PE16]. SIG [Ano04-58, Ano01d]. SIG-CHI [Ano01d]. SIGGRAD [PDK16]. SIGGRAPH [Ano01d, Ano05y, Ano06c, Ano05-38, Cou92b, Ano02-51, Ano02-52, Ano03f, Ano03-50, Ano04-52, Ano04-53, Ano05-39, Ano06-46, Aok16, Cou92a]. sight [RGMJ22, REG^{+}89]. SIGMA [DZZ79]. SIGMA-ARCHI [DZZ79]. sign [NLG20]. signage [DKM^{+}20]. signal [GCRN23]. signals [RC94]. signboard [LZYQ22]. Signed [SA17, PC23, RMD11, TF18, YLYJ13]. signing [NLG20]. Silhouette [AVM05, IVCN20]. silhouettes [KB05, ODZS11]. Silicon [Nic84]. SIMD [Fle91, Kla91, NBE^{+}04]. similar [LM89, TBG^{+}20]. similarities [MFOK94]. Similarity [MS01, HMHB08, MXK^{+}19, MG09, RCG^{+}05, SBKB23, SFS^{+}21]. Simple [DGR93, MENS19, SM22, ADOR02, BMR23, BR07, BMB95, CPC^{+}18, CST05, CO88, CV97, GM04, Joe92, KAFB18, Lak89, LW89, LKC98, LRS^{+}03, PMM18, PCGS15, RRQ^{+}22, RRB^{+}22, SRZK23, SM92, Shi93a, SG15, Szy90b, dGGV08]. simple-deformable [ADOR02]. simplex [CLWQ09, SA87]. simplicial [CDW11, FR98, FWD21, GdMF03, MLP19, NAO13, RdMF08, WWL16]. simplicity [FTU95]. Simplification [DZ91, AXC22, AL98, BS98, BA09, CA15, CSCF08, CMS98, DCV98, IFD15, KKL02, KM14, MSE20, MNS^{+}19, SGZ10, TSD^{+}07, VL98, ZZCY22]. Simplified [YWH^{+}16]. simplifying [IB06]. simulate [Car93, PMV06]. Simulated [Mit77, CCY^{+}03, WPB^{+}23]. Simulating [dLBRM^{+}12, KB10, KK21, LL00, MK85, SU93, TH09, TB18, WXL^{+}23, COM^{+}94, DSJ19b, IWT13]. Simulation [Ano95-29, Ano95-30, BN06, BDM^{+}16, CG93, CKK96, FZPM93, GSMA06, JA84a, KS86, Per84, RGGB02, VR06, WFS^{+}82, WP77, AALJ20, AK06, ACSW75, Ano05m, BD13, BWD13, BKCW14, BKM16, BH91, CZB^{+}22, CC04, CS03, CGM91, CCW13, CCM^{+}11, CD15, CSHZ04, CL18, Coo07, CBS^{+}14, DVG^{+}18, DSM^{+}99, DSG21, DG17, Düc90, EL22, FRC06, FM17, GM05, GRPR08, GW06, G^{+}01, HSB^{+}10, IA99, ICNV21, JK90, JXM^{+}10, Kam93, KS98, KL07, KES22, KPH^{+}05, KÇM00, LBLD97, Lar03, LD05, LTS96, LM22, LCDN06, LCWZ14, LF22, hLfTxDdZ09, LSWL13, LGLK16, LCA19, LRHS14, MLC^{+}22, McW87, MMS15, MBC^{+}23, NFW^{+}24, OAYG10, PXH^{+}03, PVM^{+}22, PSK^{+}11, PCPW15, PPVT03, RSN^{+}07, RB08, RP18, STT^{+}18, SM22, SAKB75, SZW^{+}14, Sou92, STBG19, TNF14, TRLX22, TPB08, UOT83, VPLL06, WXZ^{+}18, WZLQ19, WJW^{+}23]. simulation [WLL^{+}24, WMRA^{+}15, WCA^{+}11, WLJT19, XSQ^{+}97, XMD^{+}12, XXX^{+}23, YSW^{+}96, ZY01, ZDD23, dSdCLBC^{+}22, vTP20, vTCB^{+}21, Ano94n, Ano94o]. Simulations [Gra93, DGA02, EBST14, GBP^{+}17, HBOS13, LVVC06, MC23, RCB15, RGH^{+}19, SK16, TMB^{+}05]. simulator [CTN^{+}17, OT88, SHL^{+}24, SvL09]. simulator-based [SvL09]. simulators [MTN22]. Simultaneous [Elb22, FO21, KHS^{+}10, YBD^{+}24]. Singapore [Ano03a]. Singe [ZLM^{+}15]. Single [Ack96, ACA16, CSL23, FAZ21, LHH^{+}21, ALR23, CLX^{+}19, DLZY14, EPB^{+}19, GCLZ16, HKL^{+}23, HSE10, HYP^{+}24, JXW^{+}22, KCS22, KCS24, KS09, LK00, LCL^{+}21, LSE18, LZL^{+}19, LWD^{+}18, LLL^{+}23, LRY^{+}24, MENS19, OK03, PECW22, PJVH^{+}24, PCV16, QSLS23, WLB^{+}23, YZ17, YBD^{+}24, ZP07, ZGZS22, ZZXT18, ZPL^{+}15, ZW88, dCdLL14]. single-image [ZZXT18]. single-pass [ZP07]. Single-picture [ACA16]. Single-View [FAZ21, CSL23, ALR23, KS09, PECW22, PCV16, WLB^{+}23]. singular [DW13, HPKE19]. singularities [Gom14]. sinus [MT00, YSW^{+}96]. sinusoidal [PPN^{+}23, Pic88b]. sinusoids [Ram89]. SIR [MPP98]. SIT6 [AMR23]. site [Dom77, DSR11, GRF^{+}22, LZR22]. Sites [Rob78]. Situated [RHN03, GANM21, LZR22, SFD06]. Situation [KD03, BSGT03, CIK99, GN04, HW16, KH03]. Situation- [KD03]. situation-aware [CIK99]. situation-awareness [BSGT03]. Six [Bar77]. Sixth [Ano01-57]. size [CHZ^{+}23, MX14, POBB09, QL23, WCS23]. sized [DR15]. sizes [OP15]. Sizing [BSPR77]. skeletal [BMW12, HHG97]. Skeleton [CGS^{+}21, CLM^{+}19, LL11, ML12, BKR^{+}16, CD23, GWX^{+}18, HWR^{+}23, LZZ^{+}19, LMY^{+}21, LWJ^{+}22, LZG^{+}23, NTAI20, PSBD19, SKL^{+}13, WCLZ14, ZCL^{+}22]. Skeleton-based [CGS^{+}21, HWR^{+}23, LMY^{+}21, ZCL^{+}22]. skeletonization [IKB00, MWLZ22]. SkeletonLab [BMU^{+}16]. skeletons [BMU^{+}16, SJT20, WSK^{+}22, WCY^{+}20]. Sketch [APS09, Ano12s, BAS^{+}15, BRPC18, GA12, LNL^{+}23b, MDM^{+}21, MAO^{+}12, OSSJ09, OK12, QYC^{+}22, SD08, YS17, YS21, ZPL^{+}15, ALD12, BES00, CWL20, EBC^{+}15, FGP^{+}10, HR07, IKTS22, KSM07, KS09, LKL^{+}20, LV02, MWLZ22, MQW09, MFL11, RFZ^{+}17, XL10, YS15, YH21, ZjLW^{+}14, ZHP^{+}19, ZSS20, ZGZS22]. sketch-aided [LKL^{+}20]. Sketch-Based [Ano12s, APS09, BAS^{+}15, BRPC18, LNL^{+}23b, OSSJ09, OK12, QYC^{+}22, ZPL^{+}15, ALD12, BES00, HR07, IKTS22, KSM07, LV02, MWLZ22, MQW09, MFL11, XL10, YH21, ZjLW^{+}14, ZHP^{+}19, ZSS20, ZGZS22]. Sketch2Jewelry [ZjLW^{+}14]. SketchANIMAR [LNL^{+}23b]. SketchCADGAN [KKJ^{+}23]. SketchCleanNet [MKDM22]. sketched [BPGW11, EHBA10, WY03]. sketches [EK15, KKJ^{+}23, MKDM22, PPM18, PCV16, PVCM22, RTB^{+}18, RMSC11, SRZK23, TEC^{+}14, WWF^{+}23]. Sketching [BRPC18, MS09a, TCP00, ZNA08, CCCP04, DCLB19, IWM^{+}09, MAO^{+}12, OC21, VD98, WZS19]. skew [Aum89]. SkewT [DG17]. SkewT/LogP [DG17]. skiing [LWZ^{+}23a]. skill [SWH^{+}17]. skills [FvdPT01, JMC^{+}04, MY16, SHOC23, SHL^{+}24]. skin [CBU^{+}15, LTH^{+}19]. skinning [HCV^{+}22, HAL20, KKH23, LL11, MD99b, RRC^{+}18]. Skouras [WS22]. skull [WH96]. skulls [LYW^{+}11]. Skyscraper [MVG^{+}21]. SLAM [ET18, LWWY22, PTY^{+}16, WCW^{+}24]. SLAM-based [PTY^{+}16]. slice [YR98, YHX10]. slice2mesh [LCA19]. Sliced [LXY22]. slicing [HBA13, LAE^{+}19, RM05]. slicing-based [RM05]. SlidAR [FFP^{+}21, PTY^{+}16]. slideshows [MGM^{+}23]. sliding [MS07]. slopes [LL00]. SM [Ano04-54]. SM’04 [Ano03-51]. Small [HWYL21, AEA13, ANGH11, Car92, COSEV22, DVG^{+}18, GJN^{+}21, PXH^{+}03, YT87]. small-scale [DVG^{+}18]. small-screen [GJN^{+}21]. Smaller [CCW98]. Smallest [HG02]. Smart [GPS18, ACP20, GKT02, HHK^{+}13, XLL^{+}18, SW91, Ano05j, BGT16, CEG18, FGMR22, LPS19, PBB21]. SmartKcom [RHN03]. smartphone [MOS^{+}21]. smartphones [BB15]. smartwatch [LSW^{+}23]. smartwatch-based [LSW^{+}23]. SMI [AS16, Ano04-55, Ano05-40, Ano11k, WF22, WS22]. SMI’04 [Ano03-52]. SMI’09 [Ano08i]. SMI’10 [Ano09b]. SMI2021 [BLZ22]. SMI2022 [BMP22]. SMI2023 [BLM23]. SMOG [MDSU88]. SMOG-85 [MDSU88]. smoke [TRLX22, YLYJ13]. Smooth [AM91, AL98, GSSK^{+}13, KP15, SS09, ASC17, BF19, DH95b, GY19, HJ16, KP19, LMW06, PF09, SXG^{+}09]. smooth-shaded [BF19]. smoothing [AAK22a, GG14, GWL^{+}20, KB06, LGRP14, LZ24, MGJ^{+}11, ZZD^{+}19]. sn [Bri95]. Snail [RBP^{+}01]. Snail-like [RBP^{+}01]. snake [LKC94]. Snap2cad [MOS^{+}21]. SnapNet [BGLA18]. SNet [HDZR23]. snow [LCL^{+}23, NAS18, NR07a]. Snowed [LCL^{+}23]. snowflakes [CR03]. soap [IA99]. soccer [MLM^{+}17, MRG^{+}19, YMYI11]. Social [BD17, RWD14, DGC^{+}21, HZC^{+}22, ICNV21, KSH^{+}19, LLW^{+}23, TB18]. society [BF02b]. SoDA [VPLL06]. sodium [RFB23]. soft [DLZY14, GZL21, KG04, MLP01, PSBM10]. soft-body [KG04]. Software [Ano86i, Ano01d, BG79, CL96, Fra83, HK04, KW96, KHS03, KD86, KNDT20, MK85, OL96, RG93, RK84, Str85b, Ano86b, CJAR21, CH96, Enc85a, Enc85b, Enc86a, Enc86b, FWW13, Gna84, IYH97, JM88, Koh96a, KS15, LLM^{+}90, McW88b, Pfa83, Pfa84, Pra99, RCM^{+}20, Sco84, SET^{+}88, SvD03, SUF^{+}18, VT07, You89]. soil [VPLL06]. Sol [NdSV20b]. Solid [All84, Ano95q, Ano95r, Ano95s, Ano02-50, Ano03-53, Ano03-51, Ano07a, Ano07b, CSZ92, DB83, Her83b, Her85, KR92, Pre84a, SLYY97, Wit84, Ano04-54, BG88, CYCL09, CL18, Con91, DMG99, DKY97, Ela86, FT02, GWP00, GdMF03, GO90, GN89, HAL20, JDGS88, KKNT88, LX08, MT88, MG98, MHCL15, PP02, Ros90, SET^{+}88, TRLX22, YT83, YK18, van89a, Ano01-57]. Solid-interpolating [KR92]. solid-space [PP02]. solids [CCC97, Dav95, FM17, HH91, Kle86, LAE^{+}19, NS87, Nav89, Wei84b, Woo90, XCZ^{+}16, dMF99]. sols [CF77]. Solution [DW82, How75, Ada97, CPLB14, FRC06, HH91, LB75, LWJ^{+}22, PK91, Rag80, SKCP99, dVTT18].

Solutions [SMG77, CBU^{+}15, HK04, LW99, LSS21, PM93, ZK24]. solvent [GD11]. solver [ANE17, GRPR08, JAS97, WAM17]. Solving [DGKK20, EA19, SMSS13, How75, JKK02, JALS03, SJZW07, SS93, YHHS93]. SOM [BF02a]. Some [All77, ABCD93, EK85, Fra94, FA94, MTT82, McW91b, Mei83, Pos77b, Qui91, Sar02, WFS^{+}82, BCC^{+}22, DDM^{+}06, DTZ09, Hoo91, JKK02, BPS06]. sonar [CTN^{+}17]. SORS [PdSMdML^{+}12]. sort [Las91]. sorting [dGMW16]. sound [ZPIS23]. soup [HLL^{+}18]. soups [SHK18]. Source [Ano06k, Ano07l, Ano07m, Mac77, Ano05q, Ano06j, Ano06l, LMHRG10, OK20, Ano05o]. Sources [BP93, FB12, PA91]. southern [AMC03]. SP [MCT05]. SP-Octrees [MCT05]. Space [And98, Bec94, Har83, HF85, IY18, PS08, Wen84, Zha98b, AKF^{+}20, AB97, Bag93, BND^{+}17, BjOwKM12, dLBRM^{+}12, BPS03, BSC^{+}21, BEFV94, CXCH23, CWC^{+}14, CK02b, CCW01, CLH^{+}23, Day92, EL22, GZL21, HBG14, KFW16, KAÖ98, LCG19, LCL15, LS05, Med86, dSNJA22, PSMD14, PAE^{+}21, PP02, RMG15, SEKA19, SYMW21, TCL^{+}15a, TPK13, TCP00, Wee21, Wüt98, XCL^{+}19, XXX^{+}23, YS97, ZGZS22, ZMK18]. Space-filling [Zha98b, AKF^{+}20]. Space-Plot-Modelling [Wen84]. Spaced [Cor82, ITW18]. SpacePicture [Kir93]. spaces [AK21, Bec95, BRHB20, GGW22, NdSV20a, S^{+}01, ZC07]. SPACESKETCH [NC12]. spacetime [CGZZ15]. Spain [AL10, IS05]. Span [DGR93]. spanning [IU09]. SPARP [LK00]. SPARQ [KS91]. Sparse [Mel19, QSXT22, WLZ^{+}23, ABG^{+}18, CD15, DMG20, FV13, HXC^{+}23, LGZ^{+}21, LR16, LLW^{+}19, NUM24, PR23, SSM11a, WRLZ23, XCL^{+}19, ZZLZ21, ZXD^{+}14, dBWK18]. SparseFormer [WLZ^{+}23]. sparsely [OMW13]. SparseSoftDECA [NUM24]. sparsification [ITW^{+}20]. Sparsity [KPA12, ZGZS22]. Sparsity-based [KPA12]. Spatial [Ano93e, GWW^{+}22, KS91, dGMW16, RNM95, AF00, BM08, CPS^{+}22, Can94, CCL^{+}03, CWGR01, CDF14, DDQM98, DLR^{+}10, DLN^{+}18, FWX^{+}18, FJS11, HDZR23, HHL99, KPSN04, KCK17, KSH17, LLH17, LYS^{+}19, LW99, LSGFRC^{+}13, MGSC^{+}10, Mok88, ROP11, SA04, Spr04, SSS15, WBJ^{+}21, ZWWC23]. spatial-aware [HDZR23]. Spatial-data-driven [GWW^{+}22]. spatial-temporal [LYS^{+}19]. Spatially [MCMV22, UKL^{+}13, WPH^{+}14]. SpatialRugs [BSC^{+}21]. Spatio [MTS21, MPSB21, RDD^{+}18, ASR^{+}22, BBMGM22, ÇB22, DMG20, EL22, FBTT^{+}22, HP91]. spatio-directional [ÇB22]. Spatio-temporal [MTS21, MPSB21, RDD^{+}18, ASR^{+}22, BBMGM22, DMG20, EL22, FBTT^{+}22, HP91]. Spatiotemporal [TW24, JL23, MLM^{+}17, TPRC18]. Speaking [SB77]. Special [AS16, AF16b, AEJZ18, Ano96a, Ano04-56, Ano11m, Ano12s, Ano13o, Ano21i, Ano22j, Aok16, AMS22, ALP17, ABK11, BK19, BV22, BTD^{+}22, BGT16, BLZ22, BPH20, BCMM07, BP08, BYC19, BKS23, CJO19, CP19b, CMS11, CI11, CH15, CM18, DKFC20, DS15, EK18, FH11a, FH11b, FK04, FGMR22, FS17, GW17, GMMP21, GB18, GMM18, HKS00, HSTR20, HJL07, Ise21, IB22, IS05, JZ17, Joh18, KF02, Kiy19, KP05, KPBR20, KKN^{+}21, KNDT20, LPD^{+}18, LC18, LPS19, MLX18, MB10, MK03, MTB18, NTB18, NTTF21, OP19, OOC22, Pan06, PBB21, PK17, PB11, QGW08, Rei04a, RFS22, RS22, RGJdQ18, SK22, SPO22, SCG23, Sez16, SNB17, SK12, SD16, SUF^{+}18, SA19, TK20, TT19, TSC16, TP17, TKdJO22, VB17, VT22, ZWSW22, vLM09, AEMT88, ACP20]. special [Ano04g, Ano05m, Ano12l, Ano12o, Ano12m, Ano12n, Ano13n, Ano13l, Ano13m, AdSMD23, BA23, BFLP20, BDL^{+}22, BMP22, BLS15, BLM23, BKS21, CHPS20, CMS12, DPS10, DK24, EZ22, FJA08, FT24, FR04, GHFH08, GSKZ22, IK21, JZZ16, LS19, LL00, LC23, Lin10, Liu21, Mar07, MS16, MB14, MCPW21, ND23, NN15, PB22, PB23, PDK16, RSK^{+}24, RMSB22, STP^{+}20, SPK19, SC12, SSS15, TMH20, VHE10, ZT16, vLLSM09, vdPS08, ACO01, GPS18, GHK18, Kry14, Mou15]. Species [CG93]. specific [BC01, SJZ^{+}23]. Specification [Car84, Lam00, STW82, SSS90, Lie85]. specified [ALD12, Jas88, NTAI20, Str86, YLS^{+}21]. spectator [LZR22, YMYI11]. Spectral [GD95, LDLD23, VBW12, ALM19, CP21, IP03, JJPP^{+}22, LDLD22, WWS^{+}13, WYZ20]. SpecTrHuMS [LDLD23]. spectroscopy [GVC^{+}20]. Spectrum [ITW^{+}20]. Spectrum-preserving [ITW^{+}20]. specular [KCL18]. specular-enhanced [KCL18]. speech [CUD06, HWSW19, IC96, Ner75, Par75b, PK86]. speech-enabled [CUD06]. Speed [Ano93b, Ano94a, DL09, FCW^{+}10, HS11, Koh97, MKG00, WZLQ19]. Speeder [BCHM02]. speeding [GL83, QGGW97]. SpeedSeg [HS11]. speleothems [FM22]. SPH [BGPT18, CIPT15, SLQ^{+}19, vTCB^{+}21]. SPH-based [SLQ^{+}19]. Sphere [CFH^{+}18, Ley05, Van85, AS19, Coo02, HY93a, dFP22, PACSG^{+}23]. Sphere-based [CFH^{+}18]. sphere-of-influence [dFP22]. Spheres [Chr78, AJAC23, And94, LKHM19, MTM22]. spherical [AF11, BBMGM22, CZ22, DWH09, KG20, LZP^{+}04, MCV18, PT16, PPG^{+}18, QMHH91, RM91, SK03, SPS12a, WHFL16]. Sphinx [AdBC^{+}04]. SPIE [Ano01-28, Ano01-58, Ano04-57, Ano04-58]. spike [FPR92]. spin [BYQZ22, CGR98, LSW15]. spin-image [BYQZ22]. spinning [DYW^{+}22]. Spiral [BvW06, CCW99, FHP89, QMHH91, RPAM06, RAA^{+}08]. spirals [Rei10, ZEK^{+}17]. splashing [RPP20]. splat [IVCN20, ZLG^{+}15]. splat-based [IVCN20]. splatters [GGW22]. Spline [ML79, WKT21, ALC06, AFW^{+}18, BS93, BjOwKM12, BW92, CY94, CF13, DTG15, DDPT98, HAL20, KH00, KKH23, Lam00, LQ15, LH91, LP83, LYS^{+}16, MW12, PS15a, PS91, PGVACN06, RB20, Sar92b, Sar92a, Sar94a, Sar00, SBH01, Sar02, Sar03, SWZZ23, SBD15a, TT83, Wal94a, WM05, WZT97, WHL^{+}09, WSK^{+}22, WF11, Wil03, Woo87, WWF^{+}18, XCL^{+}19, XLM12, ZDT07, ZCC^{+}16]. Spline-based [WKT21, WSK^{+}22]. spline-interpolation [Wil03]. Splines [Haz79a, BLNZ22, CLWQ09, Chu90, CS80b, DBG92, FEVM10, FK82, GSY94, KP15, KP22b, LQ12, MW92, Pha89, PF89, Sar94b, Sar04, SP16, SZ09, WM90, WSHY22, ZZY12, LL91]. splitting [EVRW23]. Sponsored [Ano01d, Ano05y]. spoofing [LLZ^{+}23]. Sports [LZR22, HGH^{+}18, MPW^{+}12]. spot [CGG19]. spread [MHYN23, YHNC22]. spring [GÖT97, LM22, McD06, TH09, CSFG96, DS15, SD16]. square [CM23, Mok87]. squared [WY11a]. Squares [Coo78, CC01a, LZC^{+}15, MYF06, RGRG15]. squats [HGH^{+}18]. SSI [Las91]. sspline [Mon87]. SST [BBMGM22]. SST-Sal [BBMGM22]. stability [MW14]. Stable [LBB12, LS08, She93, BWD13, BCC20, DVG^{+}18, FRC06, Kur24, LBB11]. stack [LP92]. stack-based [LP92]. stacking [CTP^{+}21]. STAG [ACP20, FGMR22, LPS19, PBB21]. stage [BRPC18, DS93]. Standard [Cul84, KA86, PMK85, BK91, CG87, LZ88]. Standardization [Ree85, Str85a]. Standards [Cot75b, NM85, Wis86, AD94, End85, GN94, Gna84, HTKRW88, KUMW90, Sta87b]. Star [LGLK16, ZNGN16]. Star-effect [LGLK16]. StarLogo [KCR02]. State [HK15, CSK97, CK96, PRM^{+}24, YHNC22]. Stateless [DKZ14]. statement [Gin02, Ley02]. States [Nol95]. static [CCI12, EVRW23, KK16, MB97a, PSBM10, Zay12, ZQL15]. statics [LdSP^{+}08]. station [Men85]. stationary [BLNZ22]. Statistical [PMS87, Pic84, Pol83, KB06, KD15, Pur87b, RLU^{+}19, WSX12, ZGW^{+}16]. Statistics [ZML^{+}18, ANGH11]. Stats [LZR22]. status [SB97]. STDM [WMDR08]. Steady [JS09]. Steelcase [Ano86a]. Steepest [NAO13]. Steering [JYL24, CVP^{+}16, G^{+}01, MCM^{+}18]. steganalysis [MYL^{+}23]. Steganography [WW06]. Stellar [FWD21]. stencil [JCT^{+}15]. stenotic [RMW^{+}17]. step [BR89, Mel19, ZDL^{+}24, RBUB94]. steps [Sch95]. Stereo [JYC^{+}23, CACC24, JK15, LZW^{+}21b, LYZ15, LFY^{+}21, LHY23, MRC15, PGS^{+}23, QSXT22, RZF19, SXW^{+}22, WRLZ23]. stereolithography [WM95]. stereopsis [RZY^{+}20]. Stereoscopic [NAK13, ACGC22, BCC10, CL97, COPR17, FBP96, NC12, PJVH^{+}24, Sou92]. stereoscopically [PAE09]. stereovision [KKO01]. Steven [Mil92b]. stick [ASS22]. sticks [MS90]. stiff [MMS15]. stiffness [WXL^{+}23, ZEK^{+}17]. stimuli [Kur24]. stimulus [LW89]. stimulus-response [LW89]. stippling [MALI11, MARI17, MAdS^{+}19]. stitched [ASC17, ASC18, Pet18]. Stochastic [JOK^{+}07, TFY00, ASKCK03, BAD23, BA09, LVLD10, Lin97, McC02, Pur87b, Rag09, TMN^{+}00, TNU^{+}01]. stopping [TDR^{+}17]. Storage [Bas77, Gab77, GC86, LZT^{+}24, MD85]. store [SS75a]. story [EL04]. storyline [PBH19]. Storylines [ZC07, CLH12]. Storytelling [Ano06-50, Ano03-58, Cra02, DGA02, LNP^{+}13, Spi02, SGBI02]. Strafor [Ano86a]. straight [CZ06, EVRW23]. straight-line [EVRW23]. straightedge [MP22]. Strain [Cet23, PiP00, BSPD10]. Strain-based [Cet23]. Strange [Spr96, Goe95, PdFS06, Pic88a, Spr93]. strange-attractors [Pic88a]. Straßer [EJ15, Enc15b]. strategies [AAB18, C^{+}01, CWGR01, Sal85, SA04, SWF^{+}20, THL15, WWO^{+}23, vTP20]. Strategy [BMP84, ASKCK03, FWX^{+}18, LPP^{+}19, dGMW16, MMdOE^{+}22, WSL^{+}19, WLQC18]. stratification [PTL04, SKS17]. stratigraphic [STM^{+}04]. Stream [LD12, BH15, SMMS01]. Stream-based [LD12]. stream-line [SMMS01]. Streaming [GK04, dATNMC^{+}22, SSB04, SWF^{+}20, TPRC18]. streamline [LWS15, LSW12, MPL02, SP00a]. streamline-based [MPL02]. Streamlined [TVL16]. streams [FCW^{+}10, GSV^{+}18, LZSG03]. street [KVB^{+}20, KGB^{+}21]. strength [VBVS88]. stresses [RS08]. stretchable [IMG22]. string [K^{+}00b]. strip [DW89]. stripe [LSHL18]. strips [VK07]. stroke [HS14, Sch98]. stroke-based [Sch98]. strokes [BKL15, HS11, IB06, MHLB16, SPL14]. Structerf [WCW^{+}24]. Structerf-SLAM [WCW^{+}24]. Structural [Ade85, Ano77, BS77, ACO12, CXCH23, CTLG94, LZL^{+}24, PPL91, WCW^{+}24, ZPIS23]. Structure [AXC22, DFNP84, FST97, Krö98, Sac22, Wec79, CLH^{+}16, DKY97, FEJM75, GM04, Her83b, HHCM17, HY93a, JYL17, LZL^{+}19, LPP^{+}19, LCD15, LZL^{+}15, LXCW18, Mar80, Mar82, Nav89, NLS07, OK02, PLJL15, PBN97, PMBS14, PDL^{+}21, PR93, Rei02, SVNB99, TPRC18, Vää93, WLL21b, XWY15, XH88, YL23, YC10, YCF18, YAKE23, YLT^{+}18]. Structure-aware [Sac22, PLJL15]. structure-preserving [WLL21b]. structure-texture [JYL17]. structured [BPR90, GZLW14, Lew75a, RLS^{+}12, YSZ22]. Structures [BSPR77, FZPM93, Gru87, San85, Ade86, ABG^{+}18, AKW^{+}16, BMW12, BAD23, BAS^{+}20, BK20, BWYZ24, CBNJ^{+}15, DEST95, DLW23, DHZL20, GK96, GVTA10, GNL^{+}15, HHK^{+}13, Hoo91, IL97, IV93, JH89, KH00, Lew75b, LCGN92, MHW10, Mar76, Mar80, Mar82, Mor76, OMP^{+}18, PKRM21, PPG^{+}18, RMW^{+}17, SGBP17, STM^{+}04, SA86, SCCS13, SCH^{+}18, SEKA19, VAGT08, VMAL16, Wal89, WBA16, WQL^{+}17]. Structuring [Eas85, Mac85, Sif99, BD97, EPB^{+}19, Pol83]. STSR [TW24]. STSR-INR [TW24]. student [Bow95b, BBMK21, GS87, KS96]. Students [Dan78, AKB22, Boa78, MGSC^{+}10]. studies [GVC^{+}20, IWM^{+}09, SRA^{+}19, YRD^{+}24]. Studio [ZSM^{+}19]. Study [AC77, GLT^{+}97, LYZ^{+}24, SHOC23, AHK03, BRHB20, CC20, Cas87, CMLR11, DQF04, GPC^{+}17, GSB^{+}21, HAL^{+}21, HWSW19, HGH^{+}18, KP95, LWZ^{+}23a, MAM^{+}24, MLCMGR23, PVR87, PMBS14, RB06, RLT16, SLRP16, ZGZS22]. Studying [JH11]. stuntman [FvdPT01]. Style [LXY22, BCMD17, CSF20, PQCT23, TFF^{+}20, WZZZ18, WZC^{+}24, XWF^{+}20, ZPN^{+}21, ZGZS22]. StyleGAN2 [SLGQ23]. StyleTerrain [HY23]. Stylistic [CLN^{+}16]. stylization [IB06, LPD13, MSRB17, MR17, NAK13, SLKD16, SKH^{+}12, WKO12, ZIP^{+}19]. Stylized [MDS^{+}21, MDS^{+}22, US20, WCH^{+}11, WCF^{+}23]. sub [KMV^{+}18, NLdAL^{+}23]. sub-pixel [NLdAL^{+}23]. sub-terrain [KMV^{+}18]. Subdivision [JSP03, LDD07, MRS17, TJ85, ASSF17, BLNZ22, CCL^{+}03, DAHF04, GSF99, KP18, KD11, KJS17, Lam99, LŽ03, LD09, LAL11, LKL^{+}02b, LLC09, MB97a, NA02, NG05, OKT01, RSAF18, SS09, SMFF04, SK04, SBH07, TAF16, VPBY02, XLHH21]. subdivisions [BS11]. subject [AB03, Yin04]. subjected [PX06]. Subjective [Lod21, MR05, PAE09]. subjects [SBD15b]. subnets [LZL^{+}24]. subpath [LG23]. subspace [XLHH21, ZCL^{+}13]. subspaces [vRESH16]. substitution [GS99, QLCV96]. subsurface [CPLB14, ZK24]. subsystem [Mol96]. subtle [FO21]. subtrees [WG17]. Successes [Lie85]. successful [CMA10]. successive [Bro76]. suggestive [KPMT18]. SUI [SSS15]. Sumi [NLSN11]. Sumi-e [NLSN11]. summarization [RDD^{+}18, YWC22]. Summary [KKMT06]. summed [LZW^{+}21a, DVND10]. summed-area [LZW^{+}21a]. sums [LM89]. Super [NLdAL^{+}23, HW22, SLL^{+}23, TW24, YZC^{+}23, ZP92]. super-plane [ZP92]. Super-resolution [NLdAL^{+}23, HW22, SLL^{+}23, TW24, YZC^{+}23]. Supercomputer [LLR93]. supercomputers [Gin93]. Supercomputing [AW93]. Supercover [CZ06]. superimposition [ADOR02]. superior [KSKS96]. superpixel [WLL21b]. superresolution [HXC^{+}23]. supersampling [SMSS13]. supervised [CJXZ23, DWZ^{+}22, GOdSC23, HLXL23, JWZ23, LWS15, LZKJ23, LHY23, MRWL23, WYC^{+}23, ZYW23, VTW23]. supervision [LRY^{+}24]. Support [Ano07-33, HYP^{+}20, MBN21, WQL^{+}17, Ack96, BDPR93, BR96, BPKG07, BB98, CFMS02, CKK96, EME15, FMCM^{+}21, GS01a, Jun94, KD94a, Lew75a, MAM^{+}24, MMF03, MSO^{+}20, NMM09, NP88, OSZ00, PSZ96, Pol83, RM05, Sar04, SHOC23, Sla92, YCF18, ZCL^{+}22]. Support-free [WQL^{+}17, YCF18]. supported [ADHC^{+}23, EJRW96, MST^{+}22, Pau88]. Supporting [CEN^{+}23, LRHS14, SFD06, CSM^{+}01, FB04, HMA23, LLH17, MAH00, ZWP^{+}93]. supports [TKOD24]. suppression [SGZ10]. Surface [AP88, AXG^{+}13, CLWQ09, CRT04, DFNP84, ELC^{+}12, GOZ95, Haz79a, HF85, IP23, IT11, KKL02, Lam87, LXJL21, MSL^{+}19, MBV18, Reh85, SFS^{+}21, SGR^{+}99, Tin86, VH15, WY11b, WSL^{+}19, XXHM21, ZJH87, AM91, ABG^{+}18, ACA07, BSJC02, BNPS10, BP10, BMB95, BW92, BS01b, CGG19, CP96, CT13, CZZ22, CMM16, Chu90, CD23, Com01, DAHF04, DCJH13, DDW11, Dol95, DCL07, Eas75, ES22, EB10, FBH^{+}01, FST97, GDDA13, GVPN09, GCRN23, GWL^{+}20, HN20, HLCF88, HGJ16, HPKE19, HAL20, IORM17, JWZ23, Koh96b, LY15, LDD07, LD09, LGRP14, LJCW04, LST96, LLM^{+}16, LW10, MGOH96, MAC19, MSMK19, MZ89, Muk86, NKMI23, NOS15, dCNPdFS14, OK12, PJ21, PMTK01, PLJ^{+}13, PS13a, PR82b, Pha89, PS91, RAG05, RP18, SOC^{+}19, Sch12, SJT20, SPGR93, SA17, SKS09, SLQ^{+}19, SHLW89, SOG08, TF18, TB19]. surface [TBDC20, TPG99, ÜT99, VPLL06, WZT97, Wan04, WWY06, WZ09, WY11a, WFC14, WYXM22, XD08, XLSW22, XZCOX09, Yin04, Zha98a, ZY02, ZS14, ZCC^{+}16, ZSS20, ZSK95, ZTS02, ZW88, APB07]. Surface-based [ELC^{+}12]. surface-editing [EB10]. Surface-reconstructing [VH15]. surface-texture [DCL07]. surface/volume [RAG05]. Surfaces [Bro79, Cor82, LD78, MG86, AXG^{+}13, Ano04-56, AO91, Aum89, AZ23, AA00, BS01a, BP94, BS93, BF07, BSF13, BW94, BLW12, Boe91, BSL^{+}13, BH15, BMS^{+}11, Bri03, Bro08b, CCKW11, CZ22, DW13, DMG99, DKV^{+}22, DE92, EH96, FS86a, FB11, FMP96, FJW11, FM09, FPC10, GM07, GD11, GLT^{+}97, GYJZ15, GD87, HL96, HMHB08, JT02, JSP03, JOK^{+}07, KC07, KP09, KP18, KP19, KB06, KYL15, KZ04, Koh97, Kuo01, Lam99, Lam00, LMZ90, Lan88, LAL11, LPV95, LCL15, Lib91, LPR^{+}14, LT95, LCCM02, Liv18, LRS^{+}03, MW12, MZPZ16, MP22, MCK12, MvSE18, MCS^{+}18, MVG^{+}21, NS87, NA02, OKT01, OSB07, Par75a, PDS21, PS15a, PPD22, PX06, QGGW97, RB20, RFB^{+}21, RH85, RMD11, Rei98, RLB^{+}02, RGRG15, RSAF18, RLS^{+}12, SV18, SR89, Sar93, SMFF04, SXG^{+}09, SBD15a]. surfaces [SPS12b, SJ09, SHBSS17, Ste99, Stü98, SHZ19, SWS10, SCT^{+}14, SE01, TFY00, TMN^{+}00, TNU^{+}01, TAF16, TBG^{+}20, Tsu17, UBW99, Uga06, VPBY02, VD98, Wal94a, WHZ^{+}18, WF11, WF88, Woo87, XHGL12, XXT18, YMZ^{+}15, YCZ04, ZD04, ZDT07, ZSW08, ZLZG12, ZS14, ZGZ^{+}16, ZZ15, ZM92, ZLZ^{+}20, ZFSY04, ZQ12, Zhu91, ZCL^{+}22]. Surfacing [SHBSS16, SVSV20]. surfel [GBP04]. surgeons [YRD^{+}24]. Surgery [PGS^{+}23, BEKL00, CC04, CSJ^{+}21, CSS^{+}24, GHM^{+}96, KÇM00, hLfTxDdZ09, MTC02, WH96, YSW^{+}96, YRD^{+}24]. surgical [FEBS07]. surprising [Dod09]. surround [KMWW^{+}14, PJVH^{+}24]. surround-view [PJVH^{+}24]. Survey [FAZ21, Hun77, KHS^{+}10, MAAS15, PM22a, ROP11, AS22, Ano86b, Ano86i, AA13, BWF18, Bec94, Boe91, CP21, CP19a, CCM^{+}18, DH07, DG01, Doh95, Enc85a, Enc85b, Enc86a, Enc86b, ERB^{+}14, FVG15, GTdS^{+}18, Han97a, Han97b, JTT01, Kam93, KCK17, KB04, LLS^{+}16, LPZ^{+}21, LWG^{+}23, MARI17, MGMB22, MWDG13, MCTB11, MG08, NLG20, NG88, NY06, OSSJ09, ÖOK23, PCS00, PS18, PM20, PM22b, RdAMA21, SSM11b, WPJP23, XLW^{+}24, ZTAP21, vdLdFvdEV23]. sutras [Mad00]. SVBRDF [ARL^{+}20, HQW14]. SVG [Ano05k]. SVM [TDR^{+}17]. SVR [HSTR20, NTTF21, TKdJO22, NTB18]. swapping [ZZXT18]. swaps [WSJJ24]. swarm [dSJdML18, LWW^{+}20, WWF^{+}18]. swarming [MBA20]. swaying [AK06]. sweep [CS06, MA94, MGOH96]. Sweeping [AA92, Van85, ZLZ^{+}20]. sweeps [MSE20]. Swept [HL96]. SWF [LHL23]. SWF-GAN [LHL23]. swin [HXC^{+}23]. Swinging [Ano12a]. switchable [YS23]. switched [Lak91]. Symbol [SB77, FK11, GZW12]. symbolic [SV06, YBTB23]. symbols [CDF14, LKS07]. symmetric [BR07, CSX^{+}19, LM16, Spr96, SK13, TT82]. symmetrical [ZS94]. symmetries [AKF^{+}20, CCW01, LLL^{+}15]. symmetrized [Mas92]. Symmetry [LYW^{+}11, VJ06, YYG16, AJRV00, CCW99, CCC00, DHJ^{+}99, Egh83, HS08, JR00, LXCW18, PCV16, Rei97a, Rei97b, ZLLY06]. Sympathy [HZC^{+}22]. Symposium [Ano91d, Ano94f, Ano94g, Ano95-31, Ano95x, Ano95y, Ano95w, Ano01d, Ano01-41, Ano01a, Ano02-50, Ano03y, Ano03-54, Ano06-47, Ano06-48, Ano06-49, Ano07a, Ano07b, Ano07y, Ano07-39, Ano07-31, Ano12s, Aok16, AMS22, BTD^{+}22, BYC19, KNDT20, Rei04a, TKdJO22, Ano01f, Ano01-42, Ano03f, Ano04-39, Ano04-40, Ano04-51, Ano04-62, Ano05-28, BF02b, DKFC20, MTB18, SUF^{+}18, Ano77, Ano86j, Ano01-57, Ano03-62, Ano07-38, BCMM07, BP08, HSTR20, NTTF21]. sync [PZM^{+}23]. synchronization [HK93]. synchronized [Par75b]. Synchronizing [HK93, vTP20]. synchrony [GSB^{+}21]. Synergies [DGA02]. Synergistically [DJH^{+}23]. Synthesis [BY88, CG85, EK22, EAAY23, GM86, KA85, MG86, RAHA88, ZPL88, ACA16, ACC^{+}18, CM14, CGG19, DSJ19a, DCL07, FJ17, FBTT^{+}22, GCDL22, GS05, GvK18, GCCZ14, HR04, LWW^{+}20, LR16, LYZ^{+}22, LWFZ23, Loh95, MXK^{+}19, MS82, MHZ94, MPSB21, MSHL22, MM10, MSR^{+}13, MA17, NLG20, NZZ^{+}21, Owe94, PBK13, PJVH^{+}24, RO87, RPSP^{+}19, SD21, SNS06a, SCFF16, TFF^{+}20, TG02, UPTd92, WZZZ18, WSG10, WI00, XWY15, YP05, YGS12, ZSL08, ZPN^{+}21, Ano77]. Synthesizing [ACO12, CYCL09]. synthetic [ACV03, ADOR02, HZC^{+}22, Kru99b, LhCE97, NRTT95, PdSMdML^{+}12, PLM^{+}05, PRM^{+}24, TC24, dMTB^{+}21]. syringe [HTW^{+}19]. System [AC77, ARL78, Bax77, BR96, Ber79, BG79, BKCS79, DZZ79, DM79, FS78, GM78, Gom85, GÖT93, How79, JA84c, LMZ90, Mac78, SBOT78, SB84, VP77, Wec79, Wen84, Wit84, AEMT88, AKPS00, ALD12, AVM05, AAB92, AFM93, BBH90, BCS^{+}99, BA90, BHTT94, BJS01, Cas87, CS16, CC03, CFMS02, CT75, DSM^{+}99, DSG21, Dur91, EW99, EHBM82, Ela86, EMB^{+}98, FBT04, FRC06, FEJM75, FP75, GWEA^{+}11, GN94, GA83b, GAÖ02, GN80, GS87, GB75, HR97, HKCL02, HIS83, HO88, HPD^{+}10, HFT^{+}99, HAB75, How88, HS99, II22, IMMS82, IA83, Jar75, JK90, JXJ22, JA84b, Jun94, KSM07, KKJ^{+}23, KAV^{+}88, Kir93, KKO01, KKNT88, KY97, KGGP19, KMWW^{+}14, KAAO75, LNFC95, LdSP^{+}08, Les02, LL05, LL06, Lew75b, LWZ^{+}23a, LG94, LLW^{+}23, LTPN96, LR90, LPV92, LCGN92, MTTL82, MBPF12]. system [MRG^{+}19, MKDM22, MDSU88, McD06, MPOL96, MPL21, MP89, MW83, NKA83, Ner75, NGA95, Nic84, OMGGG^{+}19, OK20, OK02, OMF93, PBH19, PMZS97, PMPR^{+}16, Pfa84, Pic87a, PLM^{+}05, Pol83, RŞH^{+}22, RR01, RSN^{+}07, Sch98, SWS75, SZW10, SVP82, SH96, SS75b, TH09, TN02, TAS09, TS75, TM75, VBVS88, WK14, WJD^{+}09, War76, XZL88, YT83, YSZ22, YNS94, YSD13, Yue86, ZR97, ZZL21, dHG^{+}97, dHT01, GH91, KGM75, Pic94, PS86]. Systematic [OK99, AMM^{+}07, KCU^{+}22, NLOdS23, RLT16, SLRP16]. systèmes [Le 77]. Systemised [Wal93a]. Systems [AB78, Ano95a, Ano95b, Ano95c, Ano07-33, Car84, CG93, DM79, DIE78, Eas85, Egl86a, Haz79b, Lea85, MG86, Mar87, Reh85, SL02, Str85a, WFS^{+}82, AJRV00, ABG^{+}18, APS09, Ano94-29, Ano94-30, Ano01c, Ano04d, App87, AAB92, A^{+}99, Ban85, BG01, Big86, Can94, Cor84, COSEV22, Cot75a, DG96b, EHM84, EBST14, End83, FA94, Gag95, G^{+}01, GA83a, Gro94, HP01, JSV98, JT86, JY98, KD94a, KAFB18, KN88, KKNT88, KAVM23, LHS87, LW90, LSS97, LZ88, LWW10, MMS89, Mar86, MS82, Mes84, MMV88, MR90b, Nik98, Nik06, OH83, OST^{+}16, OdlCA02, Osi05, Pla00, PS18, QV95, QLCV96, RŞH^{+}22, RVR04, SB94, SS02, Sla92, Spr94, SAK90, TACS22, Tan80, TR95, Van09, VV89, VP98, vWS04, Ano86d, Ano93e, Ano95-41, Ano95-39, Ano95-40, Ano02f, Ano07g]. Systolic [AP88, MPR89, PS91].

T [BjOwKM12, HAL20, MAC19, WHL^{+}09]. T-mesh [BjOwKM12]. T-NURCC [MAC19]. T-spline [HAL20]. T-spline-based [WHL^{+}09]. table [BC13, CMB17, Lan97, CJJ99]. table-lines [BC13]. tables [LZW^{+}21a, DVND10]. tablet [Par75a]. tabletop [KAFB18, RR01]. tableware [CGH97a, CGH97b]. tabulating [VHR^{+}18]. tactile [CZB^{+}22, dJONM18]. Tai [HGH^{+}18]. Taiji [Bro07d]. tailor [AAB92]. tailor-made [AAB92]. tale [LQ15]. Tangent [SBH07]. tangential [PK91]. Tangible [WBA16, ZGS17, RBW01, Séq13, SEMWC05]. target [ILLC01, RCM^{+}20, RHFL14]. targets [DYW^{+}22, SYMW21]. TARig [MZ23]. task [AMZ^{+}24, BC01, FGP^{+}10, GTdS^{+}18, GN04, KD03, KYT^{+}17, MCG^{+}23, PBH19, PWV^{+}18, TPM14, XWW^{+}19]. task-and-technique [GTdS^{+}18]. task-awareness [KD03]. task-based [PWV^{+}18]. Tasks [Dai93, DGC^{+}21, LFP10, MSL^{+}19, MA14]. TAssembly [DJC^{+}23]. Taubin [AAK22a]. taught [SS75b]. taxonomy [DH07, RPHL14, SB97]. Taylor [Sou93]. TAYRA [WBL^{+}97]. TCVG [Ano04-62]. teach [BEKL00, CB01, SBHC22]. teacher [BBMK21]. Teaching [Bai07, Bre01, DZZ79, DMS08, FWW13, HGS23, Owe94, Tak77, Tax04, Wol02a, Boa78, FF96, GS05, KJ01, OSZ00, PPSS96, SHOC23, SS04b, Tei96, Hol03]. Teachware [MBGK89]. teams [MRG^{+}19, PMBS14]. teamwork [DR09]. tech [HH88]. Technical [ATB98, BX99, BSB^{+}02, CR99, FR98, GM02, GPTB02, GZ99, Gra02, GAÖ02, IA99, JD99, JKK02, JT02, KB02, LV02, MW99, MRR98, MD99a, MPP98, Mar02b, Mar02a, MF02, OK99, Oik98, PS02, PTW98, RLB^{+}02, SR02, SF98, SM98, Sif99, TL02, VH02, VL98, WEWL99, WBK98, Yu99, ŽC99, ZLS99, dMF99, FJ03, Thi85]. Technique [Pro85, AS95, BC88, Cha90, CST05, GTdS^{+}18, GL83, HS11, JDGS88, JRS21, KKH23, Kru99b, LZ14, LDT02, MT88, MAG^{+}12, MS08, MMS^{+}17, MFP11, Par93, Sob89, TKOD24, UL22, WYXM22, YT87, ZZXT18]. Techniques [Ano95q, Ano95r, Ano95s, BS82, JV84, KA85, Mag85, PN83, ADR01, Ano94-32, Ano94-31, Arb92, AA13, B^{+}00, BC01, BHL^{+}94, CP19a, DG96c, Den90, Doh95, DM01, Fay85, FGP^{+}10, FBT93, Frü91, GBD88, GS04, HSD96, H^{+}00b, HLCF88, Hol94, H^{+}01, HLS89, KA22, Kam93, Kel00, KCK17, KB04, MAH00, MD99b, MCTB11, McC96, Mul01, Nah23, NT00, Ng95b, Pic88b, PP02, RM05, RBUB94, ROP11, SP04, SK99, Tan80, TPB08, WZW97, ZM07, dAPG18, vdLdFvdEV23]. Technological [Ano03a]. Technologies [Ano06-50, McC84, Ano06-33, BCF06, LPZ^{+}21, SB83]. technologists [CMA10]. Technology [Ano01d, Ano07e, Mar07, McW88b, APA^{+}11, Ano86a, Ano04h, BS04, BF02b, CDGC94, CDGA84, GA83b, Hol03, JHL^{+}12, McW87, Nap95, SUF^{+}18, Ano05-41, Ano07d, Jv95, KNDT20]. technophile [CMA10]. tele [A^{+}01, D^{+}01]. tele-conversation [D^{+}01]. tele-presence [A^{+}01, D^{+}01]. Telebuddy^{T M} [D^{+}01]. telecommunications [MPOL96, Que93, Ano07p]. TeleInVivo^{T M} [CGS^{+}96]. telemedicine [GLL00]. teleoperation [BNS24]. telepresence [Gag95, MBPF12]. telerobotics [KKC94].

Teleservices [Ano94a, Ano94y, Ano94x]. Teletel [Lam83]. Televirtuality [Que93]. television [CS04, DQF04, EL04]. Telexistence [AMS22, BYC19]. temperature [SW19]. Template [DLS^{+}11, HGJ16, DJC^{+}23, LH00, LYW^{+}11, MZ23, VDOK19]. template-aware [MZ23]. Template-based [DLS^{+}11, HGJ16]. temple [RM22]. Temporal [LSS97, PA07, SBKB23, AMHWW16, ASR^{+}22, BBMGM22, DMG20, EL22, FCW^{+}10, FBTT^{+}22, HW22, HFP06, HP91, LAB^{+}14, LYS^{+}19, LPP^{+}19, MTS21, MPSB21, RDD^{+}18]. Temporally [BCC20, SCVCN16, YY14, CBM^{+}22, ERWS12, FJ17, HJW97]. temporally-alternating [CBM^{+}22]. temporary [WG17]. Tencon [Ano94s]. tendon [TH09]. tension [MZPZ16, Sar92b]. tension-determined [MZPZ16]. tensor [BRSP15, EKG06, Lam00, LQ12, LM16, MvSE18, VW21, WVY16, YRS^{+}18]. tensor-product [LQ12, MvSE18]. Tensorpose [SdSR^{+}19]. Tenth [Ano07f]. teragons [VB99]. Teramac [KMS^{+}97]. Terminal [Gab77, KD86, Rei75]. Terminals [Bas77]. Termination [LKL^{+}02b]. terms [WSJJ24]. Ternary [LLC09]. terracotta [JVS^{+}24]. terrain [CP96, DR15, HY23, HS99, KMV^{+}18, LV02, LL00, LZW^{+}21a, Pum96, SD21, SAB12, SBD15a, TEC^{+}14]. terrains [AJAC23, NAS18]. TerraMobilita [VBS^{+}15]. TerraMobilita/iQmulus [VBS^{+}15]. terrorist [WLL^{+}24]. tesselation [CP98, SMFF04]. Tessellated [Bro79, YWH^{+}16]. tessellation [BSJC02, CYKK09, HY93a]. Tessellations [CCW97, CCW01, RA15b]. test [And98, CK09, FTU95, FT97, JCFN18, Lin90, MRF06, RF97, WLW05, dMF99]. testing [GANM21, KA22, Kru84, Mag84, MF02, Pfa84, Sco84]. tests [LWW07, ME92, Wan23]. tetrahedra [FN08, TPG99]. tetrahedral [CP98, SGR^{+}99, XCZ^{+}16, XCW^{+}09, ZCT95]. tetrahedron [Rei97b]. Text [LNL^{+}23a, LHL23, Sch86a, WCF^{+}23, App87, BPGW11, CIT^{+}19, CÖ91, FP87, HCX^{+}23, Hor83, LPPM07, SHG98, WZC^{+}24]. Text-based [LNL^{+}23a]. Text-driven [WCF^{+}23]. text-guided [WZC^{+}24]. Text-to-Image [LHL23]. TextANIMAR [LNL^{+}23a]. textbook [EK15]. textile [BK20]. textiles [EW99, Pla00]. texts [HMHA98, MNSJ99]. TextStyler [WZC^{+}24]. textual [PdSP^{+}22]. Texture [CM14, CM93, KA85, MG86, WLL21b, ZSL08, Ack96, AnD19, AL98, AVHT17, AA00, BAD23, BN03, Bus97, CGWZ22, DSJ19a, DG96c, DFF22, DCL07, EWWL00, GP91, GD95, GLS^{+}20, HE15, HR04, HPD^{+}10, JYL17, KK08, KSKS96, Kra10, LSY11, LCL^{+}21, LSZQ21, LBTM15, MXK^{+}19, MHZ94, MWS04, MMK04, Mol96, Nah23, PBK13, SHK18, SK98, SNS06a, SL16b, TWBP03, Tap06, WLL^{+}21a, XZPG21, XWW^{+}21, Yu99]. Texture-aware [WLL21b]. texture-based [AL98, AVHT17, BN03, Tap06]. texture-less [LSZQ21]. textured [CCL^{+}03, HOCN07, JCFN18, VH15, ZZCY22]. Textures [CG85, GM86, GPC84, AT08, CYCL09, DW05, DMG99, LVLD10, LB12, Loh95, McG08, MMH^{+}21, SG92, WR02]. Texturing [FDGM18, CCKN01, DG01, LRD07, SD15]. TG [CGWZ22]. TG-Net [CGWZ22]. Thank [Ano09j, Ano10j, Ano11n, Ano13p]. Thanks [Ano15l, Ano15m, Ano16i, Ano17i, Ano19k, Ano20q, Ano21s, Ano22t]. Theater [PKK03]. theatre [GHM^{+}96]. Their [SGM97b, Bär90, BF15, LZLS18, McC96, RBF20, SGM97a, SBHS10, TCL15b]. them [HE80, LS18]. Thematic [BT78, WA75]. themed [WBA16]. Theoretic [Del80, Ano95q, Ano95r, Ano95s]. Theoretical [vTTK^{+}20, SS75a]. Theories [OKK83]. theory [Ano94-32, Ano94-31, Ano04-59, Cas96, Fiu87, FLV20, JYLW14, K^{+}00b, LS18]. therapy [FSS^{+}02, MMT^{+}23]. There [SBG99, S^{+}01]. thermal [TNF14]. THESEUS [DL93]. thick [SMMS01]. thickening [CA17, XZY^{+}21]. thickness [CKCK09, YL23]. thin [CA17, CKCK09, Les02, XXT18]. think [Hol03]. Thinking [Kre93, MHM95, OC21]. Third [Ano95-41, Ano95-39, Ano95-40, Ano06-50, Pot77a, Ano02-53]. Third-Year [Pot77a]. thoracic [MT00]. Thoughts [Eck90]. threading [RVR04]. threat [FBH^{+}21]. Three [AR84, Ano24c, ARL78, BS77, CB78, DH07, FZPM93, Fal89, Ger86, HTKRW88, Jon90, KP98, LW99, Roe00, Str85a, VP77, Baw97, CWTL08, CKCK09, CR99, DK97, GZZS06, GMNS93, IWM^{+}09, LPO20, McC02, Mor76, NT00, NR07a, Par75a, PLM^{+}05, RFZ^{+}17, RO87, REG^{+}89, Sch95, Sen03, SGS99, SO75, Ste75, SJ15, SF92, TS75, VJ06, WTF95, ZMM^{+}90, LH83]. three-coloring [McC02]. Three-Dimensional [AR84, ARL78, BS77, CB78, FZPM93, Str85a, VP77, DH07, Fal89, HTKRW88, KP98, LW99, Roe00, CWTL08, CKCK09, CR99, IWM^{+}09, LPO20, Mor76, NR07a, Par75a, PLM^{+}05, RFZ^{+}17, RO87, REG^{+}89, Sen03, SGS99, SO75, Ste75, SJ15, SF92, TS75, VJ06, WTF95, ZMM^{+}90]. thresholds [WCS23]. throwing [YCO23]. TIDSE [Ano06-50]. TIFF [LCGN92]. Tight [Mar03, SOG08, WSL92]. tightly [ZK24]. tightly-packed [ZK24]. tile [BK23, LYS^{+}19]. tile-based [BK23]. tileable [MLC^{+}22]. tiles [AKF^{+}20, McC02, YAKE23]. Tiling [FB11, MD99b, SDS89, Wei99]. tilings [CCW99, Fat01, Fat02, GS99, SWMdF21]. Time [FZPM93, LLC04, Pic87a, Rag09, Shi92, Sug83, AALJ20, AMM^{+}07, AJ94, AOB17, AJAC23, AMZ^{+}24, AP22, BKV05, BLZD12, Bay95, BH91, BFSE03, BP10, BWYZ24, BSC^{+}21, BSM^{+}22, CF11, CAAC20, CTN^{+}17, Cha97, CB97, CVHM03, CJXZ23, CF13, DVG^{+}18, DRFRD06, DVND10, DMG20, DR09, DG17, DKV^{+}22, EW75, EHM84, ERDS14, ES22, FCW^{+}10, FP75, GRPR08, GW06, GYD75, GSV^{+}18, GG14, GCRR11, GPR^{+}95, HBOS13, HG22, HZ23, HXH24, HH88, HGH^{+}18, JXJ22, KL07, Kel86, KLW12, KKO01, KK22, KM21, LAB^{+}14, LRD07, LCCS04, LWWY22, LSHL18, Lin79, LD12, LCX^{+}23, LSS21, MWLZ22, MAM^{+}24, MDJ^{+}95, Mar02b, Mar10, MVRB18, MPQG18, MMK04, MVG^{+}21, Mey79, MA14, MSRB17, Mor75, MA17, NIH08, Nah23, NAS18, NG03, PAJ19, PdSMdML^{+}12, PSBM10, PS08, PGS^{+}23, RB08, RHC15, RGGB02, RJG06, SDT21]. time [SBR23, SLS03, SD08, SV06, SdSR^{+}19, SWvB95, SVVS^{+}17, SHD^{+}17, SCSG18, TSY11, TW24, TKOD24, TWSH02, TKB07, VP06, WLP^{+}14, WZ21, WJW^{+}23, WJGG15, WCA^{+}11, WF11, WLJT19, WLG04, XLQP12, XMD^{+}12, YSZ22, YMZ^{+}15, YHNC22, YBTB23, ZXL^{+}21, ZDL^{+}19, ZYX^{+}19, ZQL15, ZGWP16, ZLLG18, ZZXT18, ZW88, ZMM^{+}90, dHG^{+}97]. Time-averaged [Rag09]. time-based [SD08]. Time-critical [LLC04]. time-dependent [BKV05, MVG^{+}21, RGGB02, TWSH02]. Time-discrete [Pic87a]. time-oriented [AMM^{+}07, MA14]. time-resolved [PAJ19]. time-sharing [Mey79]. Time-Varying [Sug83, DMG20, DKV^{+}22, LLC04, NIH08, RB08, SVVS^{+}17, TW24, VP06, ZQL15]. timeslicing [PLFT21]. TIPS [OKK83, You89]. TIPS-1 [OKK83]. tissue [GZL21, KÇM00, PSBM10]. Title [Ano11o]. TLS [LLC^{+}22]. TLS-LiDAR [LLC^{+}22]. today [Roe00, BK96]. together [S^{+}01]. tomograms [GCvdS^{+}20]. Tomographic [NOS15, Bus97]. tomography [GS89, ZWS19]. Tomorrow [Bag93, BK96]. tonal [MSD75]. Tone [BCMD17, GLC20, ASZ^{+}14, BMS^{+}11, Cad08, CWNA08, CKM^{+}23, CJXZ23, GBA15, LTH^{+}19, SO75].

Tool [Haz79b, PBB21, RK84, ACP20, Ban85, BGK89, CEG18, Cha97, CUD06, COSEV22, ESFGDZ97, FF96, GZ99, Gre96, HMA23, IYH97, KD00, LaV07, Lar03, LD05, LBV14, LCA19, MO92, OL96, PPVT03, REG^{+}89, SHOC23, SDWE99, SEMWC05, SSS90, SWS10, TBLH17, WBJ^{+}21, Wan04, ZSM^{+}19, ZMM^{+}90]. toolbox [DGV^{+}24]. tooling [CGH97b]. toolkit [DL93, MAH00, SVV92, SKL^{+}13]. Tools [BGT16, FGMR22, Gag95, GPS18, Haz77, LPS19, NHR^{+}22, Bär90, BHL^{+}94, Cun00, IM07, KHS03, LSR22, MUS83, OC21, RPM96, Wol02b]. ‘TOON’ [Dur91]. Tooth [WCLZ14, JXW^{+}22]. top [MTTL82, SS04b]. top-down [MTTL82, SS04b]. topics [BF02b]. Topological [AMFH21, DSNW13, KS09, Las91, LM16, Liv23, AA92, AG94, BA09, BD97, BK20, CEPS13, IKTS22, LSW12, vTP20, dlLC99]. Topologically [ACG15, AKF^{+}20, IFD15, CGMS00, DN22, TSK98]. Topologically-consistent [IFD15]. topologies [SSBT01]. Topology [BS98, KBL22, SZL^{+}13, TWSH02, ZFS03, ASC17, ASC18, AG94, AB11, BLNZ22, CBNJ^{+}15, DS93, LQ12, LKHM19, MMALRA01, MRSS^{+}18, MTM22, Pet18, SGS99, SvRvL07, SCC11, WYZ^{+}11, ZGC15, ZZC20a, dLvL00]. topology-adjustable [ZZC20a]. Topology-based [KBL22, MMALRA01]. topology-preserving [WYZ^{+}11]. topometric [GWEA^{+}11]. toric [YJLZ21]. torques [RB08]. torus [Bri95, Coo02]. Tosiyasu [WF22]. TOSL [SMMS01]. Total [ZW20, LH14, MSMK19]. touch [AMR23, DHJ^{+}97]. touch-based [AMR23]. touching [Ano13n, SMM20]. touchscreen [CH12]. tourist [CMD99]. tower [LRR87]. toy [VA96]. TP [Ano04-59]. traced [KM21, LG23]. Tracer [dVTT18, vWdlHFK23]. traces [ASR^{+}22]. Tracing [Cyc93, Ger86, IO91, Van85, AZ23, BS93, BHH15, CF99, GH98, HMHB08, JC95, JK21, KPFT03, Koh97, MTS21, Mar01, MS90, PM91, QGGW97, Sam89, SEDT^{+}03, SLS03, SGGC05, SDD95, SSB^{+}08, STBG19, SSA96, SKP99, UPTd92, VHS12, WIP08, WZ21, ZM92, PACSG^{+}23]. track [ECG^{+}22, FGZ^{+}22, QYC^{+}22, BHTT94]. tracker [SvRvL07]. trackers [SvL09]. Tracking [A^{+}99, RLT16, WM89, AP99, BND^{+}17, BS09, CMM16, CGZZ15, CCC97, CS18, DYW^{+}22, Dor99, DKV^{+}22, EKG06, ENE11, FEBS07, GWBD17, IMG22, JHL^{+}12, dSJdML18, KLW12, KYT^{+}17, Kur24, LDM^{+}11, LSZQ21, PGS^{+}23, RAK^{+}15, SP00a, SGC^{+}19, Sei99, TWSH02, VBP05, WLZL18, WWK11]. tractography [VW21]. tradition [HH88]. traditional [GPC^{+}17, JXY87, MAdS^{+}19, MCKS06]. Traffic [HW16, KGM75, MC23, Tay87, WXZ^{+}18, WHW^{+}22, dMTB^{+}21]. Training [Gra85, WR79, BGK89, BRO22, CCY^{+}03, CSS^{+}24, DGA02, Goh84, GV07, HWEB22, HTW^{+}19, HLXL23, JRS21, KÇM00, LWZ^{+}23a, LLW^{+}23, MTN22, MRWL23, MPW^{+}12, OMF93, RSN^{+}07, SCB^{+}24, SL18, WJG^{+}19, WWO^{+}23, WYC^{+}23, XWF^{+}20, ZBP^{+}18, KHS^{+}10]. trajectories [CPC^{+}18, CGG^{+}20, LRHS14, McD06, SKH^{+}05]. trajectory [GSV^{+}18, KB05, SLM^{+}22]. transcendental [CZC02]. Transfer [LXY22, Reh85, ZSH12, AVHT17, BWBM20, CHSB10, LSY11, MSMP12, MCS13, OBD^{+}23, PZM^{+}23, PQCT23, dMPF08a, dMPF08b, PR11, RdAMA21, RPKLMG23, SKSZ99, SLGQ23, TFF^{+}20, WCL23, WZC^{+}24, XWF^{+}20, YSDG24, ZHP^{+}19]. transferring [fLhLfT11]. transfinite [BF15]. Transform [Egh80, Bim15, Egh83, JPP01, KG20, Lam99, Lan97, MA18, ML12, Sch75, WW08, WKT21, ZDC^{+}23]. transformable [HCLC16]. Transformation [Jac95, AEW91, AA01, CZC02, CK96, HS08, KLL^{+}15, Ley07, Mor76, Van94, VT06].

Transformations [HC95, PK85, Sou92, AA01, CL97, CS01, Dur89, FSV17, FS20, JSG04, Kor90, Les01, WX14]. Transformer [CSL23, PGL^{+}23, WXC^{+}23, FLR^{+}23, HXC^{+}23, HXH24, LDLD23, SXL^{+}23, WLY23, WLZ^{+}23]. Transformer-based [WXC^{+}23]. transformers [HDZR23]. transforms [KPA12]. Transient [PAJ19, RGMJ22]. transition [AJ94, CSF20]. transitional [BKP01]. Transitioning [UKW23]. transitions [AL98, PS15a, SBHS10]. translation [MVK^{+}22, YPZ^{+}23, YHW23]. translator [Wei84a]. translucent [KJS18]. Transmission [PN83, WFS^{+}82, CPG94, RS99, SYMW21]. Transparency [dSC07, MCTB11, MCTB14, RAF21]. transparent [Ger02, WZ21]. TransPIFu [CSL23]. Transport [Ano94y, Ano94x, CPLB14, LS18, SAKB75, SKS17, ZK24, APSS01]. Transportation [IC87, Kor87, TS87]. transposition [GC86]. trap [Car99a, ZLLY06]. trapezoid [Kla91]. trapezoidation [LDT02, ŽC99, ŽJŽ03]. trapezoids [ŽJŽ03]. traps [Ye02]. travail [CV77]. traversal [KNMP14, WEWL99, Zhu91]. treadmill [MLCMGR23]. treasure [CWTL08]. treatment [B^{+}00, CPG94, KFH^{+}09, OMF93, RSN^{+}07, SGBP17, WF88, Woo90]. treatments [Sch88]. Tree [SW91, AK21, BWYZ24, ÇB22, HLZ^{+}17, JD99, JDGS88, Lew75b, PSMD14, RBF17, SA86, TT12, XM12, YYY19, YHNC22, YH21, vBT21, Lew75b]. Trees [Ger86, AK06, AJAC23, ACA16, BNPS10, Bro07a, Car92, COM^{+}94, CCI12, Gon98, GCRR11, IU09, Kru99a, LD11, MS90, RE22, VRV05, WSK^{+}22]. Trends [Ano77, Rie78, Sak02, Str85b, COSEV22]. tri [WFC14]. tri-prism [WFC14]. trial [MWY^{+}10]. triangle [BGMP08, CK09, FN10, KYKK19, MA14, RHK^{+}20, RMG15, RFR02, SHK18, SVW23, SXG^{+}09, VK07, VMAL16, VH15, WSL^{+}19, YC10, ZSL12]. triangle-based [RFR02, SVW23]. triangle-triangle [CK09]. triangles [KYL15]. triangular [BFRA11, CPM19, CT13, FJ17, GdMF03, KP09, LKL^{+}02b, PJ21, PB96, VGP04, Wil03]. triangulated [XCZ^{+}16]. triangulating [PR93]. Triangulation [Chr78, XLGG11, YC10, ZS14, CNS^{+}06, KKŽ04, LŽ03, MG98, PPM18, SM22, SWL^{+}16, ZMYH06, ID17]. Triangulations [Ang97, CTS^{+}10, LZLS18, dCNPdFS14, VP00]. Tribox [CR99]. trigonometric [Mic90]. trilinear [CGMS00]. trimmed [MvSE18]. trimmed-surfaces [MvSE18]. Trimming [HPKE19, Uga06]. trivariate [LQ12, PL97, PPP88]. Trots [VP77]. Truchet [Bro08b]. True [Nik98]. True-colour [Nik98]. truncated [Ram89]. try [II22]. try-on [II22]. TSNeRF [WCF^{+}23]. TSR [HW22]. TSR-VFD [HW22]. tub [PR96]. Tube [Bas77, Gab77]. Tubelets [SKH^{+}05]. tubular [CZ22, FDGM18, KMV^{+}18, RBF17, XCZP14]. Tumor [Düc90, LBLD97]. tumors [RNM^{+}19]. tuning [ACGC22, GBA15, XLW^{+}24, ZLS99]. tunnel [DL09]. turbines [RJKV12]. Turbulent [Pic87b]. Turin [Pic88b]. Turncae [Ano85f]. Turning [YS23]. turnkey [GA83a, MW83]. TURP [CC04]. turtle [JSG04]. tutorial [Ano94j, Bow95b, Cha97, Gon98, Maj98, Mil93, Roj91]. tutorials [FT24, Hop86]. tutoring [LdSP^{+}08]. tutors [HEW^{+}18]. TV [GBG04, Lin79, Rau06]. Twelth [Ano04-58]. twelve [PR96]. twill [ACC^{+}11]. twill-woven [ACC^{+}11]. Twin [HLXL23, MH21]. Twist [Ano12a]. twisted [MA94]. twisted-profiled [MA94]. TWIXT [Gom85]. Two [AB78, BG80, Car99a, CXXW20, CB78, Dav95, Gab77, GLDH97, How79, Mar78, McW90, Osi05, SKSZ99, SS89a, SM92, AC89, AS91, Aro89, CCKW11, CM15b, CYM16, CRD10, DS93, EHM84, GMNS93, KPA12, KHS^{+}10, Koh96b, Koh97, MK89, MD85, MZ89, OK99, Par75a, PZM^{+}23, PTL04, PF80, SR15, SKSI95, SPY87, SJG19, TRLX22, TWSH02, WC89, Xu08, ZDL^{+}24, ZWL^{+}22, Zub88]. Two- [CB78]. Two-Dimensional [BG80, Gab77, How79, Osi05, AS91, MD85, Par75a, PTL04, SPY87, TWSH02, WC89]. Two-layer [CXXW20]. two-level [CRD10]. two-pass [PZM^{+}23]. two-piece [SJG19]. two-step [ZDL^{+}24]. two-view [OK99]. two-way [TRLX22]. type [MVRB18, Pie88, WZT97, YLH^{+}18]. typeface [Sch98].

U [LTC^{+}20, LSCJ23, ZYW23]. U-Net [LTC^{+}20, LSCJ23]. U-shaped [ZYW23]. U.K. [Smi77]. UAV [JRJP^{+}22, LNSW16]. UAV-based [JRJP^{+}22]. UDAformer [SXL^{+}23]. UFormer [PCKB23]. UGRAF3 [BBH90]. Ultra [SV06]. Ultra-real [SV06]. UltRASim [SHL^{+}24]. Ultrasound [STBG19, ANGH11, BNS24, FSS^{+}02, GPTB02, SHL^{+}24]. ultrasound-guided [SHL^{+}24]. umbrella [LCWZ14]. un-distortion [JK15]. uncalibrated [RR15, Sei99]. uncann [WSJJ24]. uncertain [MW14]. Uncertainty [ATHL14, HBG14, MMT^{+}23, RPHL14, RMW^{+}17, RGH^{+}19, DYW^{+}22, DKLP02, EKG06, MRW^{+}21, RHFL14, SLK^{+}23, VT06, ZWS19]. Uncertainty-aware [RGH^{+}19, DYW^{+}22, MRW^{+}21]. uncontrolled [HOCN07]. unconventional [Jon90]. Uncovering [ASR^{+}22]. Undergraduate [Lau77, Boa78]. underlying [Tsu17]. Understand [Bij85]. Understanding [CH12, KKPC23, LL04, LL05, PGL^{+}23, PRW^{+}22, ABM^{+}06, CUD06, FV06, KRP^{+}91, Les01, Les02, LL06, SLX^{+}16, dSB04]. Underwater [MMdOE^{+}22, SXL^{+}23, ZLL^{+}23a]. UNet [YBD^{+}24]. unfold [XhKKL16]. unfolding [SGBP17, ZMKG11]. unified [Gho93, LZ88, MPM15, SMM20, WXZ^{+}18, YLYJ13]. uniform [uHRBK06, LŽ03, MAFL16, YSD13, ZLL^{+}23a]. uniformly [GS11]. Unifying [EK85]. unilateral [WBJ^{+}21]. unimodal [Wal06]. union [MK89]. unique [AWI^{+}09]. uniquely [TC24]. unit [AA01, KKS93, MZCD21]. United [Nol95]. unity [RPM97]. univariate [HGW^{+}24, Pie89]. Universal [SCNT03, LHS87, LLZ^{+}23, ŽC99]. universe [BSB^{+}23, IA91]. Universitat [BJAN^{+}95]. universities [Han97a]. University [Ano03a, Jv95, KKMT06, MWY^{+}10, EGL^{+}95, Gri88, Pie90, TS94]. Unleashing [LAB^{+}14]. unorganized [LW10, SA17, ZZC^{+}20b]. unoriented [WSL^{+}19, XSW23]. unprepared [A^{+}99]. unrestricted [PCS00]. unsalient [TCL^{+}15a]. Unsharp [HKBA17]. unsmoothed [RS08]. unsteady [HW22, Tap06]. unstructured [ARL^{+}20, BAG03, Fru94, LMR^{+}19, OMW13, PM95]. Unsupervised [BKV05, EAAY23, FIC23, SS22, WWS^{+}13, HDZR23, LWJ^{+}22, WLL^{+}21a]. Untrimming [MvSE18]. unusual [MTT82]. update [BG91, McW88b]. updated [ZZCY22]. updates [CF11]. upper [LLW^{+}23]. upsampling [HDZR23, RPP20, WYZ20]. UPU [HDZR23]. UPU-SNet [HDZR23]. Urban [Bax77, BBCG11, CJJ99, DAHF04, GDDA13, HIS83, LD03, LRD07, LBLD11, LCX^{+}23, NFLYCO99, THL15, VBS^{+}15, ZZL21, vLvKV11]. UrbanVR [ZZL21]. Usability [CIT^{+}19, NGAS23, RHS^{+}94, YRD^{+}24, dHT01]. usage [COSEV22, DQF04]. Use [Bax77, Ehl85, Gra85, LS79, Mit77, ML79, Pic84, Rob78, BBMR89, Bru75, DBW^{+}12, Fal89, FP89, Gin93, IA83, Mar76, McC96, MML^{+}11, Par93, PACSG^{+}23, Per02, SVCNM23, SWF^{+}20, SVT86, WBP92, WWK11, dDH87]. use/transportation [TS87]. used [Car92, LB75, Sch95]. useful [SS02]. User [ACB12, Ano07-39, CS04, FGES96, H^{+}01, PPV07, SW19, ALD12, AP99, BNS24, CLH^{+}16, CH93, CH12, DAG22, DL93, Enc95b, Fiu89, GDA^{+}13, GD04, HAL^{+}21, HWSW19, HFT^{+}99, IWM^{+}09, KSM07, KYT^{+}17, KMWW^{+}14, Lea87, LWW^{+}20, LLW^{+}23, LVVC06, MAM^{+}24, MO90, MR95, MGM^{+}23, MCG^{+}23, MSAR01, NC07, NFHS06, NTAI20, NGAS23, PAE^{+}21, SFVP13, Sif99, SEMWC05, SSS15, SKL^{+}13, Thi85, TS95, YLS^{+}21]. User-centered [FGES96]. user-controlled [GD04, SFVP13]. User-guided [PPV07]. User-interface [ACB12, Thi85]. user-specified [ALD12, NTAI20, YLS^{+}21]. Users [Ber79, JA84b, KHS^{+}10, MA14, Nug91, RV01, Vää93, Wri75]. Using [BGD18, BS77, BPGW11, BEKL00, BB93b, Cas96, CJJ99, CG93, Dai93, DR09, FGP^{+}10, Gab77, GS05, Haz77, JP10, Mac78, ME77, ME83, MXK^{+}19, MN90, Oik98, RHS^{+}94, SMU22, SBOT78, SSM11b, SEMWC05, SPS96, TH90, TZvD^{+}21, TPM14, WMW13, XLGS16, AYA^{+}20, ACSW75, ABAA22, ADR01, ABCD93, ABJ90, ASWL11, APS09, ANE17, AOL96, AMGA12, ATZM19, AKB22, AN99, AA00, AA01, BW98, BX99, BMR23, Baw97, BPKB15, BD13, BWdBP13, BYQZ22, BCMD17, BPS03, BW94, BMMZ23, BC01, Bri03, BB91, BDM^{+}16, BBMK21, BKL17, BRPC18, Cad08, CWNA08, CLXJ22, CLWQ09, CM14, CMB17, CSCF08, Cet23, CKM^{+}23, CYJ^{+}13, CBU^{+}15, CJT96, CCL^{+}03, CWL20, CLT07, CSHZ04, Cho06, CCI12, CSLY01, CCC97, CRT04, CLE13, CMS22, Con91, CDF14, CS80b, CK93, CB01, DCV98, DJC^{+}23, Des00, DFWW15, DS93, DMS08]. using [DGKK20, DMG99, DSG21, Dor99, Egh80, Egh83, ERWS12, EA19, FCM^{+}18, FB11, Fay85, Fer01, FJ17, FWW13, FEVM10, FO21, Fou11, FPC10, FK11, GBA15, GDDA13, GdMF03, GP91, GO06, GTG06, GMd^{+}13, GSV^{+}18, GSF99, GKLM07, GSY94, GLDK95, GO90, GAÖ02, GSB^{+}21, HEWF13, HGW^{+}24, Her83b, Her85, HLCF88, HS11, HFP06, HEK22, HR04, HLS89, HLB^{+}06, HS99, IMG22, IP23, IC96, ILLC01, IC87, JS09, Jer95, JK15, JL23, JRS21, JRZ^{+}23, JHL^{+}12, JRJP^{+}22, Jut94, KL07, KES22, KKL02, KM14, KHK18, KKHS03, Kle86, KZ04, KJS17, KJS18, KPMT18, KHS03, Kor87, KM21, KÇM00, KCS22, KYM12, KJ08, Lam99, Lan97, LBLD97, LLH17, Lee01, LDLD22, LXW^{+}10, LSH^{+}12, LYW^{+}13, LLLZ16, LZW^{+}21a, LYX18, LZL^{+}19, LPD13, LM16, LWD^{+}18, LPO20, LPPM07, LL12, LB19, LBTM15, LhCE97]. using [LX08, LVVC06, MSE17, MLM^{+}17, MAFL16, MBPF12, MRG^{+}19, MP22, MOS^{+}21, Mar76, Mar86, MCMV22, MG09, MAG^{+}12, MYF06, MSMK19, MZCD21, MRC15, MMS^{+}17, MBN21, MS90, Mou13, MTM22, Mul01, MPAC^{+}23, Nar15, NLdAL^{+}23, OKT01, OBD^{+}23, OK12, ÖOK23, PBH19, PdFS06, PLJL15, PBN97, PR82b, PHLW15, PLVT23, PMV06, Pfa84, Pha95, Pic88b, Pla00, PdSP^{+}22, PCPW15, Pro85, Pum96, QLCV96, RMP19, RM22, RMD11, RdMF08, SP00a, STT^{+}18, SK03, SGC^{+}19, Sar93, SAMA97, SM07, SHK18, SGC00, SJT20, Sei99, SH94, SD08, She12, SLYY97, SSM11a, SXG^{+}09, SPS12a, SBS22, SvD03, SBD15b, SJG19, SKH^{+}12, SL01, SL16b, SWL^{+}16, SMM20, SHS^{+}23, TBS^{+}23, TNF14, TSY11, TGG06, TFF^{+}20, TAF16, TB19, TRB^{+}22, TG02, TBM^{+}04, TTKA23, TPN95, Tsu17, TT12, US20, UBW99, UKL^{+}13, VP98]. using [WEWL99, WFG03, WWL^{+}12, WTWT18, WLZL18, WZ21, WSK^{+}22, WSHY22, Wan23, WMRA^{+}15, WLQC18, WCH^{+}24, WWD^{+}95, Wir80, WX14, WCLZ14, WKW16, WWF^{+}18, WLX^{+}21, WSX12, XZPG21, XMD^{+}12, XXHM21, YRS^{+}18, YL23, YC10, YY14, YSZ22, YMZ^{+}15, YJC99, YSD13, YT87, ZD04, ZMYH06, ZMKG11, ZLS98, ZLS99, ZY01, ZLL^{+}23a, ZHP^{+}19, ZFG^{+}20, ZSS20, ZLLG18, ZK24, vBT21, vOHR20, vTCB^{+}21]. USIS [EAAY23]. utilisations [CF77]. utilising [MH21]. utilized [KCK17]. Utilizing [Wec79, BCS^{+}99, CTJ^{+}14, Jar75]. utterance [NLG20]. UV [LZKJ23].

v

[Fat02, Ano03-34, Ano03-35, Pic87b, SZW10]. v-shaped [Fat02]. V-system [SZW10]. vacation [Pic92b]. valence [KNC11]. Validation [DM01, BCF06, MGSC^{+}10, Pfa83, SD90]. valley [NZL^{+}21]. valuable [FF96]. value [BSPD10, Lam00]. values [DDPT98]. VAPOR [FGM^{+}20]. Variability [FGM^{+}20, FMCM^{+}21]. Variable [FLR^{+}23, CKCK09, DMG20, TY24]. Variable-hyperparameter [FLR^{+}23]. variables [Koh96b, Koh97, SLK^{+}23]. variance [APSS01, BM08]. variant [BYQZ22]. variate [Tap06]. variation [LH14, LJH18, PMV06, ZW20]. Variational [GVPN09, YIL09]. Variations [Woo90, Bro07d, FO21]. variety [ACO12]. various [JT02, LG03, TNU^{+}01, Wei99]. Varying [Sug83, DMG20, DKV^{+}22, LLC04, NIH08, RB08, SVVS^{+}17, TW24, UKL^{+}13, VP06, WPH^{+}14, ZQL15]. vascular [CCY^{+}03, HAL^{+}21, LLH17, RMW^{+}17, SGBP17, THQ^{+}16]. vasculature [EDKS94]. vase [ZY02]. VCBM [KPBR20, KKN^{+}21, RSK^{+}24, SPK19]. vcps [CZC02]. VDA [Ano02g, Ano02h, Ano04-60, Ano06-51]. VDA2004 [Ano03-55, Ano03-56]. VDAFS [Reh85]. vection [TMSPB09]. Vector [AT11, MP22, BF19, DS18, Egh83, Fou11, Fru94, GS89, HW22, HE15, Kni95b, Lan88, LSW12, MPL02, NT00, Nik06, RdCVL16, SMMS01, SMK08, Tap06, WWL16, ZNT^{+}18]. vector-parallelism [NT00]. vectorisation [HEK22]. Vectorization [BGV93, CFMP84, DHZL20, ZLS98, ZLS99]. Vectorized [WSHY22, XWW^{+}18]. vectors [AA01, Par88, SPS12a]. vegetation [ACA16]. vehicle [SH96]. vehicles [Ano91d, Ano04-40]. Velocity [TNFG14, LYW^{+}13]. Velocity-based [TNFG14]. ventricle [PiP00]. Verba [CUD06]. verbal [TMK94]. Verification [Bar77, GZ99, ST02, Tan80, ZMK18, ZCZ^{+}18]. Verifier [PMK85]. Verifying [TR95]. verisimilitude [Hel95]. Versatile [SLQ^{+}19]. version [AKW^{+}16, GS89]. versus [AAB92, SvLBF10]. vertex [EVRW23, FDA03, LZT^{+}24, Rei98, WFG03, dF24, BM03]. vertex-centric [dF24]. vertex-optimization-based [LZT^{+}24]. vertices [IA91]. Very [Ger02, HKCL02, HG02, TWSH02, LFL02]. Vessel [KGM75, KGK^{+}07, LD11, RHBS95]. VesselMap [THQ^{+}16]. vessels [CLE13]. vest [XLL^{+}18]. VFD [HW22]. VFire [HSB^{+}10]. VGA [FP89]. VHDL [WWD^{+}95]. via [AALJ20, ASS^{+}19, ALM19, Ban97a, Ban85, BRdSOS17, CTQ^{+}14, CYCL09, CLN^{+}16, CYWM23, CF13, CLH^{+}23, DJH^{+}23, FBTT^{+}22, GKW^{+}24, HXC^{+}23, HHCM17, HZLC22, JXW^{+}22, KFW16, KSH^{+}19, KA86, LCZ^{+}11, LWS15, LGZ^{+}21, LHG21, LHH^{+}21, LSWL13, LLP^{+}21, LJH18, LSW12, MT88, MSE20, Mor76, ODZS11, RZY^{+}20, SVT86, SLGQ23, SPY87, STBG19, TSD87, TW24, TCL^{+}15a, VZP22, WY11a, WFC14, WGS^{+}18c, WHZ^{+}18, WGS^{+}18a, WGS^{+}18b, WLL22, WCF^{+}23, WLX^{+}23, WSJJ24, WWS^{+}13, WSWL14, WZZ^{+}18, WYZ20, WLL^{+}21a, WZZ^{+}21, XCL^{+}19, XHL^{+}24, XGC18, XLHH21, XZL^{+}22, YF09, YCF18, YWC22, ZCL^{+}13, ZNT^{+}18, ZZF^{+}23, ZLL^{+}23b, ZMH^{+}23, vTP20]. vibrating [McC08]. vibration [JRS21, MS82]. vibro [ZPIS23]. vibro-acoustic [ZPIS23]. vibrotactile [KES22]. Video [Ano07h, Ano07i, CFZL16, AAAN23, BS04, BEFV94, CVL^{+}04, CCM^{+}07, CAS^{+}15, CPG94, CBM^{+}22, Cou92a, DH95b, FCW^{+}10, FP75, GHCH03, Gir93, HSE10, HW16, JM88, KKHS03, LSH^{+}12, McW89, Met85, Mou13, MMH^{+}21, QL23, SHG98, UKL^{+}13, WLG04, YY14, dAPG18, Ano07-33]. video-based [QL23, dAPG18]. Videodisc [Str83b]. Videogames [GM05, YMYI11]. VideoHandles [KSF15]. videos [BBMGM22, CMLH21, FO21, HMHA98, KSF15, Lod21, MX12, PGS^{+}23, SSM11a, XLQP12, YSDG24, ZHP^{+}19]. Videotape [CL92]. videotex [FP87]. Vietoris [Zom10]. View [ET07, FAZ21, FJW11, Gab77, GCRR11, GNL^{+}15, JYC^{+}23, XWWK21, ZGZS22, AMM^{+}07, ALR23, Ano01-62, BGK04, BFSE03, CPS^{+}22, CSL23, CSF20, CLX^{+}19, DCLB19, ESAH02, EBC^{+}15, FJ17, FV13, FBTT^{+}22, GVVJ99, GAÖ02, GZLW14, HSE10, HY93a, KS09, LBD17, LZW^{+}21b, LFY^{+}21, LFL02, LWFZ23, LHY23, MM18, MRWL23, MIGS22, MS09b, OK99, Oik98, PECW22, PJVH^{+}24, PCV16, PRBD22, POBB09, QSXT22, RTB^{+}18, RZY^{+}20, SXW^{+}22, WRLZ23, WLB^{+}23, WVY16, XZPG21]. view-based [MRWL23]. View-dependent [GCRR11, GNL^{+}15, BGK04, BFSE03, ESAH02, GAÖ02, LFL02]. View-independent [FJW11, WVY16]. View-projection [ET07]. View/style [ZGZS22]. Viewer [And85, GK04, JGA09, BSM^{+}22]. Viewing [MPS85, Rog83, FJW11, SRZK23, Sin87, Zhu91]. Viewpoint [CSCF08, RBFS10, WR79, KB05, WLG04]. Viewpoint-driven [CSCF08]. Views [Haa94, Rei04b, FDGM18, GZZS06, GN89, LR16, LYS^{+}16, NG88, PR96, RJS01, SLYY97, Suf88, Sul85, VFSL06, WTF95, YY96]. Vincent [HLS89]. VIRIM [GPR^{+}95]. Virtual [AHK03, Ano95u, Ano95x, Ano95y, Ano95w, Ano96a, Ano01-59, Ano01d, Ano01a, Ano02-33, Ano03-57, Ano03-58, Ano04-43, Ano06c, Ano06-29, Ano06-30, Ano06-31, Ano07-39, Ano07-30, Ano11m, AMS22, AFM93, BAS^{+}20, BYC19, Bry93, dPCOO^{+}05, CSF20, CLE13, CL92, Gob93a, Gob93b, GZ99, GLL00, GHFH08, HSTR20, KKC94, KNDT20, LC18, LPL^{+}05, NM85, NTTF21, PCY^{+}06, PKK03, Ree85, SK22, TKdJO22, WI00, Zhu91, vDLS02, ABG^{+}18, ABAA22, AAK^{+}22b, Ano01f, Ano02-53, Ano03-37, Ano03c, Ano04a, Ano04p, Ano04-42, Ano06-33, AA13, Bag93, BZYB21, BM03, BGD18, BMdSVR18, BES00, BBMK21, BOH97, CEG18, CP19a, CH93, CL97, CACC24, CSJ^{+}21, CS04, CUD06, COPR17, CKK96, DVG^{+}18, DKZ14, DBS^{+}18, DJG^{+}04, DLV06, DHJ^{+}97, DMG99, DR09, DGC^{+}21, FBT04, FvdPT01, FBT93, FBH^{+}21, FGLW03, GWP00]. virtual [GHCH03, GS93, GS01a, GMM^{+}23, Gre96, GG14, GSB^{+}21, HGS23, HAL^{+}21, HWSW19, HFP06, HZC^{+}22, HP01, HTW^{+}19, HJL07, HGH^{+}18, II22, IM07, Jac93, JMC^{+}04, JP10, KES22, KSH^{+}19, Koh96a, KÇM00, LBLD97, LKL02a, LY08, MMF03, MZCD21, MPW^{+}12, MHM95, MCM^{+}18, NFW^{+}24, Nij04, dJONM18, OMGGG^{+}19, ÖOK23, PSZ96, PXH^{+}03, PZH^{+}05, PHO^{+}23, PQ10, PSSP96, PS18, Que93, RdCAM01, RSAF18, RCBS10, RPP21, SWF^{+}20, SBS19, SLS03, SCB^{+}24, SD15, SBE20, SW19, SSM11a, SU93, STdKB11, SK12, SEMWC05, SB97, SUF^{+}18, SSW^{+}23, TB18, TMSPB09, VVCN12, VR16, VKA^{+}23, WJG^{+}19, WBJ^{+}21, WL15, WHW^{+}22, WCS23, Wee21, YSW^{+}96, YCO23, YH21, ZXL^{+}21, ZSM^{+}19, ZWP^{+}93, vLM09, Ano12n, BFSE03, CJJ99, LS06, SKH^{+}05, ZXH^{+}12, vWdlHFK23]. Virtuality [GMMP21]. Virtues [Hel95]. Visception [KB20]. viscosity [ZDD23]. viscous [dSASP^{+}15, BGPT18, CBC19, MP89, TNF14]. visemes [DMV06]. Visibility [MTSM10, VHS12, AMGA12, AMZ^{+}24, BPS03, BRV^{+}10, CHL^{+}11, CF99, CCI12, Fun99, Gro91, HJW97, JRSM17, Kar92, KT17, NFLYCO99, PPSS96, RH85, SVNB99, ST97, SG15, Ste99, WPH^{+}14, Yu99]. visio [PSBM10]. visio-haptic [PSBM10]. Vision [Ano02f, Ano02a, Ano04-61, Ano07g, EK85, GHM^{+}96, SKO83, XLW^{+}24, Ano86a, Ano94u, Ano94v, Ano04d, AM10, DWX^{+}23, KCU^{+}22, MST^{+}22, NRTT95, SU93, TR95, Var92, Ano01-60, Ano06-52, PCKB23]. vision-based [KCU^{+}22]. vision-Development [AM10]. Vision-Language [XLW^{+}24]. visitors [CSM^{+}01]. vison [Ano03-60]. VisSym [Ano04-62]. Visual [Ano03-59, A^{+}01, BBP13, BRHB20, BHZ^{+}21, BA23, BD98, BKS23, CLH12, CEN^{+}23, EF15, FZPM93, FGM^{+}20, HGW^{+}24, HZLC22, KKMT06, KS04, KFH^{+}09, KPBR20, KKN^{+}21, Kre93, LPPM07, MLM^{+}17, Mag84, MMS04, MCMT14, MPS85, MFP11, NC07, NHR^{+}22, OST^{+}16, PZM^{+}23, PECW22, PB11, RHM^{+}12, RAK^{+}15, RSB^{+}19, SPK19, SCB^{+}24, SRA^{+}19, SVVS^{+}17, SHD^{+}17, VT07, VT22, WB24, ZDD23, ZC07, ABG^{+}18, AMPG22, ATAG^{+}21, AS22, ANGH11, Ano05-28, ASR^{+}22, AM19, BWZ^{+}18, BN07, BX99, BS09, BD17, BKS21, CVL^{+}04, CAAC20, CTJ^{+}14, CYCL09, CGWZ22, COM^{+}94, CH15, COSEV22, Ebe00, ET18, EBST14, ESFGDZ97, FLR^{+}23, FMCM^{+}21, GSF^{+}19, GBA15, GTdS^{+}18, GVC^{+}20, GPTP10, Gro92, GWBD17, HMA23, HK15, HHLE17, HMdM^{+}95, HHZ^{+}22, JSMK14, JRZ^{+}23, KPSN04, KB20, LL06, LWZ^{+}23b, LRHS14, MRW^{+}21, MMD^{+}20, MRG^{+}19, McW89]. visual [MPL21, MA14, MMGB17, MHM95, MML12, MSO^{+}20, OL96, OC21, PE16, PWV^{+}18, RŞH^{+}22, RGH^{+}19, SBWS11, Sel89, STM^{+}04, SS12b, SZEG93, Sou92, SJB^{+}21, SC97, SKH^{+}12, TMP07, TPRC18, TKB07, WK14, WJD^{+}09, WTM12, WWO^{+}23, WCS23, Wol02a, ZK95, ZFG^{+}20, dSJ23]. visual-inertial [BS09]. Visualisation [Ano93e, Ano04-60, BS01a, HKS00, Dia94, DMT03, LLHH94, MCT05, PLFT21, PTR^{+}92, TBLH17, VP06, dlLC99]. visualisations [H^{+}00a]. Visualising [BSM^{+}22, DBLC02]. Visualization [Ano94e, Ano94m, Ano95v, Ano95-28, Ano95-31, Ano95z, Ano95-27, Ano01t, Ano01-43, Ano01-58, Ano01-60, Ano02g, Ano02h, Ano02k, Ano02a, Ano02-42, Ano02-43, Ano03-32, Ano03-55, Ano03-56, Ano04-35, Ano06n, Ano06-51, Ano06-52, Ano07y, Arb92, BMB95, BDP^{+}17, CM93, Dom93, Dom94, EDKS94, Fer01, Ger02, Gra93, HO12, HKCL02, HG02, IPV96, Jut94, K^{+}00b, LLR93, LFL02, MP93, MHYN23, MVK^{+}22, NdSV20b, PMBS14, Pic94, RFS22, SIG02, SK03, SRF08, SBH01, TWSH02, vDLS02, ATHL14, ADHC^{+}23, Ano01s, Ano01-28, Ano01-41, Ano03-48, Ano03-60, Ano04n, Ano04-27, Ano04-34, Ano04-62, Ano04-61, Ano05h, BRSP15, Ban97b, B^{+}00, BPKG07, BB03, BHL^{+}94, Bro92, Bry93, BSC^{+}21, CJAR21, CC20, CLH12, CWC^{+}14, CGS^{+}96, Dal00, DAG22, DMS08, DG06, DWH^{+}15, DG07, E^{+}00, EFP02, ELC^{+}12, EPS96, Elv96]. visualization [EGL^{+}95, FGES96, FWCS97, Frü91, Fru94, GWW^{+}22, GRIG12, GRW00, Gin93, GSV^{+}18, GBP08, GCW23, GANM21, GAÖ02, GPR^{+}95, GNL^{+}15, GK04, HBG14, HSB^{+}10, HW16, HS99, ITW^{+}20, JS09, Jef92, JH11, Jer95, JOK^{+}07, KH03, KKHS03, KR96, KMV^{+}18, KD00, KB20, LLH17, LYS^{+}19, LPP^{+}19, LTPN96, LSR22, MA15, MW99, MAM^{+}24, Maj98, MR95, MSL^{+}19, MVG^{+}21, MGS^{+}22, Mil93, MGM^{+}23, MMT^{+}23, NMM09, NT00, NGA95, OGSSLM^{+}07, OCCZ12, OSZ00, PdFS06, PiP00, PSH^{+}09, PBG^{+}14, PP20, PACSG^{+}23, PRW^{+}22, dMPF08a, dMPF08b, PB96, QKS01, RPFC01, RCM^{+}20, RMD11, RM05, RPHL14, RMW^{+}17, ROP11, RHBS95, REG^{+}89, SGPC20, SBR23, Sar00, Sar02, Sar03, SGC00, SWH^{+}17, SS93, Shi04, SV06, SSM11b, SS13, SDWE99, SL02, SLK^{+}23, SBR^{+}22, Tap06, TAS09, TBM^{+}04, TPN95, UL22, VHE10, VW21]. visualization [VHR^{+}18, Wol02b, WMW13, WCdA98, XZ00, YWR03, YC10, YWC22, YHW23, YHNC22, ZWS19, ZD04, ZNT^{+}18, ZGdDL^{+}96, ZKS^{+}96, ZMM^{+}90, dLvL00, dSB04, vLLSM09, vdLdFvdEV23, Ano94f, Ano94g, San00]. Visualization-based [MVK^{+}22]. visualization-supported [ADHC^{+}23]. Visualizations [CGM91, BN03, GVTA10, KCK17, LZR22, MPL02, PBH19, PRRR13, RBLB21]. Visualize [Ehl85]. Visualizer [GAÖ02]. Visualizing [AMM^{+}07, ADR01, CTQ^{+}14, CDR01, DKLP02, EKG06, IE98, MW14, NLS07, OK02, PM90, Rei95, Rei92, RC94, SMMS01, dSdCLBC^{+}22, Bec95, BB98, DSJ19b, PL97, RHFL14, SKKN10, SKH^{+}05]. visually [OMGGG^{+}19]. visuals [HS05]. Vitae [Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03l, Ano03m, Ano03n]. VLSI [Kni95b, Sta87a, Tan80]. VLSI-design [Kni95b, Sta87a]. VMV [Ano06-52, Ano03-60]. VMV01 [Ano01-60]. VMV2002 [Ano02a]. voice [HZC^{+}22]. Vol [Ano02w, Ano03-61, Ano04-63]. Volant [CUD06]. VolDmi [LTPN96]. Volkswagen [JT86]. voltage [KHK18, RFB23]. voltage-gated [RFB23]. Volume [Ano94f, Ano94g, Ano95-42, Ano01-30, Ano03j, Ano03-61, Ano04-63, Ano05-30, Ano07-31, AB97, BCMM07, CJT96, Elv96, Frü91, Har83, LTV08, MP93, dMPF08a, dMPF08b, PM95, RPFC01, Sch12, SF92, Wen84, AT08, AKPS00, Ano05-29, BRSP15, BAG03, BC01, CC19, CF11, CVHM03, CCC^{+}16, CL95, CGS^{+}96, CRD10, CSC10, DTWT94, DH95b, DKV^{+}22, EDKS94, EME15, FTB16, FEVM10, Fou11, GPTB02, GRIG12, GA83b, GBKG04, GLDK95, GLDH97, GCW23, GPR^{+}95, GS04, HO12, Her83b, HG02, HS08, IL97, rKC93, KS01a, KM14, KONS17, Kni95a, LAM06, LLC04, LXJL21, LTPN96, MLM03, MSE17, MA15, Man99, MDJ^{+}95, MCK12, MCT08, NIH08, NT00, NTAI20, PSM16, QKS01, RA03, RAG05, RBFS10, SKKN10, SG92, SGC00, SGR^{+}99, SWvB95, SDWE99, SKR15, TPN95, WBK98, WHH21, WZW97, WLDB11, WKE04, WG17, ZGdDL^{+}96, ZKS^{+}96]. volume [dHG^{+}97, dGHM97, Ano01-61]. volume-based [Fou11, MCT08]. volume-data [HG02]. Volume-enclosing [Sch12]. Volume-preserving [AB97]. Volume-primitive [SF92]. volume-rendering [GLDK95]. VolumeEVM [RAG05]. volumes [APB07, DMG20, Fru94, FFV^{+}11, FCG^{+}07, HL96, SJT20, YJLZ21, ST22]. Volumetric [ASSF17, MC10, Pat15, ZMKG11, ASPO15, AVM05, CC20, CSSC00, CBC19, DGBNV18, DKLP02, FSP15, GDDA13, HN20, HJDR95, ITW18, KW14, LXW^{+}10, fLhLfT11, LYL^{+}17, hLfTxDdZ09, LST96, LBB12, MKKM18, NKMI23, NG03, Oik98, PP12, PMTK01, PS12, SB94, TW24, TKOD24, WFC14, XGZ11, ZZQW11, ZLLG18]. Voronoi [DW13, CZ22, LKHM19, MTM22, RA15b, RCG^{+}05, Shi93b, SMM20]. Voronoi-based [DW13]. vortex [TRLX22]. vortices [SP00a]. Vorticity [Lam87]. Voting [WM24, CHL^{+}11, YRS^{+}18]. Voting-based [WM24]. Voxel [ZZH^{+}23, ZZD^{+}19, AP22, CCI12, DCLB19, FTB12, JH89, SJT20, SM98, YK18]. Voxel-based [ZZD^{+}19, JH89]. voxel-visibility [CCI12]. voxelisation [ZFS03]. voxelization [FC00a, PM95]. voxels [DE92, KJS18]. VoxMorph [FTB12]. Voyaging [Wil86]. VPL [CMS22]. VPL-based [CMS22]. VR [PKK03, Ano03-31, AKB22, AFM93, ACGC22, Ban97a, CCY^{+}03, CAAC20, CSS^{+}24, CCM^{+}11, CUD06, FBH^{+}21, GVVJ99, HSD96, KDS04, KES22, LWZ^{+}23a, Lin10, MOS^{+}21, MGMB22, MMS^{+}17, MTN22, MUH10, MCG^{+}23, PKK03, SL18, SvLBF10, TN02, VBP05, WJG^{+}19, WB24, ZBP^{+}18]. VR-based [TN02, ZBP^{+}18]. VRAI’2002 [Ano02-53]. VRAIS [Ano95x, Ano95y]. VRArchEducation [ÖOK23]. VRCAI [Ano04-64]. VRCIA [Ano06c]. VRdeo [BBMK21]. VRIPHYS [AEJZ18, JZZ16, JZ17]. VRIPHYS’05 [Ano05l]. VRML [BK96, LPL^{+}05]. VRST [Ano02e, Ano02d]. vs [McW87, WB24]. VSI [RSK^{+}24]. VTK [DMS08, LPL^{+}05]. Vulkan [UKW23].

W2GAN [YPZ^{+}23]. WAALSURF [HvK87]. walk [CSH08, Sbe98, SKCP99]. Walking [Ano01-62, LL00, MCM^{+}18]. walks [MBST22]. Walkthroughs [DCV98, ASKCK03, HJW97, WGS99]. Wall [XWY15, ZMKG11]. wallpaper [CGR98, CCW01]. Walls [SPMA13]. Walter [Sou93]. wand [AdBC^{+}04, FRTT18]. Warashi [YNS94]. Warp [ZHC11, JWL12, KRK^{+}06, MKG00, SGC00]. Warp-based [ZHC11]. WarpCurves [SWS10]. Warped [PFR94, GSSK^{+}13]. Warping [BB93a, JRZ^{+}23, BFSE03, JK15, SvLBF10, TDR01]. Wasserstein [DZD^{+}23, LXY22]. water [CCW13, CD15, CFMS02, DVG^{+}18, GM02, KM21, MY97, OMF93, XSQ^{+}97, YJC99]. watercolor [MSRB17]. Watermarking [LZP^{+}04, DDQM98, HEG98, LDD07, SHG98, VP98, WLDB11, WMDR08, YPSZ01, ZK98]. watershed [CDIM16, CGW07]. Watt [Hol94, Hol94]. wave [GM02, RGGB02]. waveform [PK86]. waveforms [Sen99]. Wavelet [Bus97, HLL^{+}18, YPZ^{+}23, ASZ^{+}14, BX99, GLDK95, GLDH97, HYP^{+}24, JZY^{+}23, KPA12, Lam99, MPL02, PA07]. Wavelet-based [HLL^{+}18, BX99, GLDH97, HYP^{+}24, PA07]. wavelets [OSB07]. waves [BDM^{+}16, PP02, TG02, XSQ^{+}97]. WAVEVISIONS [Koh96a]. way [DSN75, EF15, TRLX22, UKW23, WKS03]. weak [LRY^{+}24]. Weakly [GOdSC23, MRWL23, DWZ^{+}22, WYC^{+}23]. Weakly-supervised [GOdSC23]. wearable [MMSS03, VKA^{+}23]. wearables [dHT01]. weather [H^{+}00a, MHYN23, SL02, dSdCLBC^{+}22]. weathering [MG08, MB14, VNMP23]. Web [Ano01-46, Ano01-47, Ano02-55, Ano04c, Ano05-41, ERB^{+}14, KA22, LSR22, THQ^{+}16, Ano01-63, Ano01w, Ano03-62, BDKK96, BEKL00, CH96, GD11, HS03, RGE07]. web-based [LSR22, BEKL00, HS03]. WEB3D [Ano01-64, Ano05-41, Ano22j]. WEDELMUSIC [Ano04c, Ano01-29, Ano02r, Ano02-54, Ano02-55]. Weight [YPZ^{+}23, TPN95]. Weighted [JYL23, JYL24, Sar04, AB11, BW98, CK93, JT02, ZGWP16]. weighting [FWX^{+}18, JYL23]. weights [YJLZ21]. Weitek [Wei84b]. Welcome [Ano07-40]. Well [Bro79, AG94, TKD16]. well-nested [AG94]. Well-Tessellated [Bro79]. Weymouth [DSR11]. Wheel [LM16]. whence [Pot75]. Where [Nij04]. which [HWSW19]. while [MLCMGR23, dSEM19]. white [GR09]. whole [DJH^{+}23, MKM19]. whole-to-whole [DJH^{+}23]. wide [GWEA^{+}11, H^{+}00a, Han95]. wider [Cun00]. widget [STN95]. widgets [Jer95]. Wild [Bro06b, CGS^{+}21, LZKJ23, QYC^{+}22]. wildfire [HSB^{+}10]. Will [All77]. wind [KSH^{+}19]. Window [JY98, PGL^{+}23, ZFG^{+}20]. Window-based [PGL^{+}23]. windows [LRMS92, WHH06]. winged [NLS07]. winged-edge [NLS07]. winner [WF22, WS22]. Winter [Ano01-66].

WINVIZ [OL96]. wire

[AXG^{+}13, GZZS06, Kuo01, PSS04].

wire-frame [GZZS06, Kuo01]. wireframe [CXGL23, KS09, PCV16, TZT^{+}22]. WireframeNet [CXGL23]. wiring [Tes84]. wise [KSH17, LDS^{+}21, SAMA97]. within [BSGT03, BES00, BEFV94, CUD06, Sch86a, Wis87, ZCT95]. without [VTW23]. Wixom [BF15]. wizard [MSL^{+}19]. Wolfgang [Enc15b, EJ15, Enc95a]. wood [CGWZ22]. word [LHL23]. work [PMBS14]. workbook [Mad94]. workbooks [Mad95]. workers [CG96]. working [Ano03c]. workload [MGH22]. Workshop [Ano86d, Ano93a, Ano94d, Ano94e, Ano94m, Ano94b, Ano94n, Ano94o, Ano94a, Ano95a, Ano95b, Ano95c, Ano02-33, Ano02a, Ano05-30, BA23, KPBR20, KKN^{+}21, LPD^{+}18, SIG02, SPK19, Ano04i, Ano04-38, Ano05-29, Ano07-33, BDL^{+}22, CWT^{+}23, STP^{+}20, Ano94-32, Ano94-31, KKMT06, Ano01b]. Workstation [KS84, Gru87, Mit87, ZFJ90, tHd90]. Workstations [MP93, Rix84, Frü91, Gir93, McW90, Sta87a]. World [Ano01-37, Ano07p, DWZ^{+}22, GHFH08, JJPP^{+}22, LYZ^{+}22, MYC15, Pic92b, RCM^{+}20, TC24, TG02, ZWP^{+}93, Ano85g]. worlds [Bag93, BM03, DKZ14, Jac93, LUMC04, RSAF18, STdKB11]. worldwide [MO92]. worms [Bär90]. worst [SKM98]. worst-case [SKM98]. woven [ACC^{+}11, DG07, PGB86, YAKE23]. Wrap [KL02]. Wrap-around [KL02]. wrapping [Mar09, WSL92, vdBB07]. wrinkle [LLLC11, WWY06, ZSL12]. wrinkle-aware [ZSL12]. write [GL83]. written [IA83]. WSCG [Ano01-32, Ano03z, Ano03-63, SK99, Ska99, SDWE99, SKP99, SKCP99, YJC99]. WSCG’2002 [Ano01-66, Ano01-65]. WWW [FN99, KJTS96]. WWW2GCG [CH96]. x [Pe04, CLF^{+}06, GS89, Jer95, K^{+}00a, LRMS92, MW83, Ste75, WTW^{+}23]. X-MP [GS89]. X-ray [K^{+}00a, Ste75, WTW^{+}23]. X3D [HFP06]. XCT [RAK^{+}15]. XDoG [WKO12]. XGKS [HTKV84]. XPBD [Cet23, TTKA23]. XPLG [TM75]. XVII [MTB18].

Yang [Bro07d]. yarn [PACSG^{+}23]. yarn-level [PACSG^{+}23]. yarns [TKM^{+}24]. Year [Pot77a, Duc18]. Years [Mei83, KS96, Wil86]. yields [KP18]. Yin [Bro07d]. Young [WS22].

Zashiki [YNS94]. Zashiki-Warashi [YNS94]. zero [JYL17, KYKK19, PT16]. Zeus [SGR^{+}99]. ZGDV [Web87]. zigzag [LM22]. zigzagspring [LM22]. Zooming [Pum96, VR84].

References

Azuma:1999:TUE

[A^{+}99] Ronald Azuma et al. Tracking in unprepared environments for augmented reality systems. Computers and Graphics, 23(6): 787–793, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/29/article.
pdf.

Ansar:2001:VHC

[A^{+}01] Adnan Ansar et al. Visual and haptic collaborative tele-presence. Computers and Graphics, 25(5):789– 798, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/33/abstract.
html.

Akman:1992:SAG

[AA92] Varol Akman and Ahmet Arslan. Sweeping with all graphical ingredients in a topological picturebook. Computers and Graphics, 16(3):273–281, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Azariadis:2000:UPD

[AA00] Phillip N. Azariadis and Nikos A. Aspragathos. On using planar developments to perform texture mapping on arbitrarily curved surfaces. Computers and Graphics, 24(4):539–554, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/32/article.
pdf.

Azariadis:2001:CGR

[AA01] Phillip Azariadis and Nikos Aspragathos. Computer graphics representation and transformation of geometric entities using dual unit vectors and line transformations. Computers and Graphics, 25(2):195– 209, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/28/article.
pdf.

Andujar:2007:ARC

[AA07] Carlos Andujar and Ferran Argelaguet. Anisomorphic ray-casting manipulation for interacting with 2D GUIs. Computers and Graphics, 31(1):15–25, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001816.

Argelaguet:2013:SOS

[AA13] Ferran Argelaguet and Carlos Andujar. A survey of 3D object selection techniques for virtual environments. Computers and Graphics, 37(3):121–136, May 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001793.

Ahmed:2023:OEV

[AAAN23] Ahmed Mohamed Ahmed, Mohamed Abdelrazek, Sunil Aryal, and Thanh Thi Nguyen. An overview of Eulerian video motion magnification methods. Computers and Graphics, 117(??): 145–163, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002522.

Asteasu:1992:DDS

[AAB92] C. Asteasu, J. C. Astiazaran, and J. M. Besga. A dishwasher design system: An application of tailor-made CAD systems versus commercial systems. Computers and Graphics, 16 (4):395–399, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Atapour-Abarghouei:2018:CRP

[AAB18] Amir Atapour-Abarghouei and Toby P. Breckon. A comparative review of plausible hole filling strategies in the context of scene depth image completion. Computers and Graphics, 72(??): 39–58, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300219.

Agathos:2022:EGT

[AAK22a] Alexander Agathos, Philip Azariadis, and Sofia Kyratzi.

Elliptic Gabriel taubin smoothing of point clouds. Computers and Graphics, 106 (??):20–32, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000814.

Amara:2022:CVR

[AAK^{+}22b] Kahina Amara, Ali Aouf, Hoceine Kennouche, A. Oualid Djekoune, Nadia Zenati, Oussama Kerdjidj, and Farid Ferguene. COVIR: a virtual rendering of a novel NN architecture O-Net for COVID-19 CTscan automatic lung lesions segmentation. Computers and Graphics, 104 (??):11–23, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000358.

Abdulali:2020:RHR

[AALJ20] Arsen Abdulali, Ibragim R.

Atadjanov, Seungkyu Lee, and Seokhee Jeon. Realistic haptic rendering of hyper-elastic material via measurement-based FEM model identification and real-time simulation. Computers and Graphics, 89(??): 38–49, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932030042X.

Anderson:1978:TAL

[AB78] David C. Anderson and Robert L. Belleville. Two approaches to on-line graphics systems. Computers and Graphics, 3(1):49–61, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Aubert:1997:VPS

[AB97] Fabrice Aubert and Dominique Bechmann. Volumepreserving space deformation. Computers and Graphics, 21(5):625–639, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=5&aid=9700040.

Asim:2003:CDS

[AB03] M. R. Asim and K. W. Brodlie. Curve drawing subject to positivity and more general constraints. Computers and Graphics, 27 (4):469–485, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Attene:2011:GMW

[AB11] M. Attene and S. Biasotti. Geometric models with weighted topology. Computers and Graphics, 35(3): 542–548, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000525.

Alhejri:2022:RRO

[ABAA22] Aisha Alhejri, Naizheng Bian, Entesar Alyafeai, and Mousa Alsharabi. Reconstructing real object appearance with virtual materials using mobile augmented reality. Computers and Graphics, 108(??): 1–10, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001480.

Alty:1993:EUM

[ABCD93] James L. Alty, Marius Bergan, Penny Craufurd, and Ciaran Dolphin. Experiments using multimedia interfaces in process control: Some initial results. Computers and Graphics, 17(3): 205–218, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arpa:2012:PRB

[ABCO12] Sami Arpa, Abdullah Bulbul, Tolga Capin, and Bulent Ozguc. Perceptual 3D rendering based on principles of analytical cubism. Computers and Graphics, 36(8): 991–1004, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001197.

Abello:2004:HGM

[Abe04] James Abello. Hierarchical graph maps. Computers and Graphics, 28(3):345– 359, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Agus:2018:GGD

[ABG^{+}18] Marco Agus, Daniya Boges, Nicolas Gagnon, Pierre J. Magistretti, Markus Hadwiger, and Corrado Calí. GLAM: Glycogen-derived Lactate Absorption Map for visual analysis of dense and sparse surface reconstructions of rodent brain structures on desktop systems and virtual environments. Computers and Graphics, 74(??):85–98, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830058X.

Amaral:1990:MMU

[ABJ90] Carlos Amaral, Joao Bernardo, and Joaquim Jorge. Markermaking using automatic placement of irregular shapes for the garment industry. Computers and Graphics, 14 (1):41–46, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Azuma:2011:SSM

[ABK11] Ronald Azuma, Mark Billinghurst, and Gudrun Klinker. Special section on mobile augmented reality. Computers and Graphics, 35(4):vii– viii, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001208.

Attene:2006:CMU

[ABM^{+}06] M. Attene, S. Biasotti, M. Mortara, G. Patanè, M. Spagnuolo, and B. Falcidieno. Computational methods for understanding 3D shapes. Computers and Graphics, 30(3): 323–333, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000549.

Amorim:2015:FHD

[ABMC^{+}15] Elisa Amorim, Emilio Vital Brazil, Jesús Mena-Chalco, Luiz Velho, Luis Gustavo Nonato, Faramarz Samavati, and Mario Costa Sousa. Facing the high-dimensions: Inverse projection with radial basis functions. Computers and Graphics, 48(??): 35–47, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000230.

Atkinson:1977:CF

[ABTW77] William D. Atkinson, Karen E. Bond, Guy L. Tribble III, and Kent R. Wilson. Computing with feeling. Computers and Graphics, 2(2): 97–103, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Allan:1977:ESC

[AC77] J. J. Allan III and A. M. Chiu. An effectiveness study of a CAD system augmented by audio feedback. Computers and Graphics, 2(4): 231–233, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alvisi:1989:TAM

[AC89] L. Alvisi and G. Casciola. On the two array mask hiddenline algorithm. Computers and Graphics, 13(2): 193–206, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anderson:1990:GRF

[AC90] D. C. Anderson and T. C.

Chang. Geometric reasoning in feature-based design and process planning. Computers and Graphics, 14(2): 225–235, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alvarado:2009:E

[AC09] Christine Alvarado and Marie-Paule Cani. Editorial. Computers and Graphics, 33(4):439, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000831.

Allegre:2007:FFS

[ACA07] Rémi Allègre, Raphaëlle Chaine, and Samir Akkouche. A flexible framework for surface reconstruction from large point sets. Computers and Graphics, 31(2): 190–204, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002391.

Argudo:2016:SPR

[ACA16] Oscar Argudo, Antonio Chica, and Carlos Andujar. Single-picture reconstruction and rendering of trees for plausible vegetation synthesis. Computers and Graphics, 57(??):55–67, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300188.

Andujar:2012:UID

[ACB12] C. Andujar, A. Chica, and P. Brunet. Userinterface design for the Ripoll Monastery exhibition at the National Art Museum of Catalonia. Computers and Graphics, 36 (1):28–37, February 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001518.

Akleman:2011:CTW

[ACC^{+}11] Ergun Akleman, Jianer Chen, YenLin Chen, Qing Xing, and Jonathan L. Gross. Cyclic twill-woven objects. Computers and Graphics, 35(3):623–631, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000422.

Argudo:2018:SAI

[ACC^{+}18] Oscar Argudo, Marc Comino, Antonio Chica, Carlos Andújar, and Felipe Lumbreras. Segmentation of aerial images for plausible detail synthesis. Computers and Graphics, 71(??):23–34, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301887.

Akleman:2015:BMT

[ACG15] Ergun Akleman, Jianer Chen, and Jonathan L. Gross. Block meshes: Topologically robust shape modeling with graphs embedded on 3-manifolds. Computers and Graphics, 46 (??):306–326, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001083.

Avan:2022:EVE

[ACGC22] Emre Avan, Tolga K. Capin, Hasmet Gurcay, and Ufuk Celikcan. Enhancing VR experience with RBF interpolation based dynamic tuning of stereoscopic rendering. Computers and Graphics, 102(??):390–401, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002107.

Ackermann:1996:SCH

[Ack96] Hans-Josef Ackermann. Single chip hardware support for rasterization and texture mapping. Computers and Graphics, 20(4): 503–514, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600022.

Aykanat:1996:PPR

[AÇÖ96] Cevdet Aykanat, Tolga K. Çapin, and Bülent Özgüç. A parallel progressive radiosity algorithm based on patch data circulation. Computers and Graphics, 20(2): 307–324, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500132.

Alexa:2001:ESI

[ACO01] Marc Alexa and Daniel Cohen-Or. Editorial — Special issue on shape blending. Computers and Graphics, 25(1):1–2, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/26/article.
pdf.

Assa:2012:MSS

[ACO12] Jackie Assa and Daniel Cohen-Or. More of the same: Synthesizing a variety by structural layering. Computers and Graphics, 36(4):250–256, June 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000180.

Agus:2020:FSS

[ACP20] Marco Agus, Massimiliano Corsini, and Ruggero Pintus. Foreword to the special section on smart tool and applications for graphics (STAG 2019). Computers and Graphics, 91 (??):A3–A4, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030087X.

Alemparte:1975:IDS

[ACSW75] M. Alemparte, D. Chheda, D. Seeley, and W. Walker. Interacting with discrete simulation using on line graphic animation. Computers and Graphics, 1(4): 309–318, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Abad:2003:ISO

[ACV03] Francisco Abad, Emilio Camahort, and Roberto Vivó. Integrating synthetic objects into real scenes. Computers and Graphics, 27(1):5–17, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alem:1985:CGA

[AD85] W. K. Alem and S. S. Deshmukh. Computer graphics in airport management. Computers and Graphics, 9 (4):423–433, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arnold:1994:GSB

[AD94] D. B. Arnold and A. M. Day. ’graphics standards are boring’ — discuss. Computers and Graphics, 18(3): 287–293, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Adamatzky:1997:CAL

[Ada97] Andrew Adamatzky. Cellular automaton labyrinths and solution finding. Computers and Graphics, 21(4): 519–522, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700027.

Abaci:2004:MWE

[AdBC^{+}04] Tolga Abacı, Rachel de Bondeli, Ján Cíger, Mireille Clavien, Fatih Erol, Mario Gutiérrez, Stéphanie Noverraz, Olivier Renault, Frédéric Vexo, and Daniel Thalmann. Magic wand and the Enigma of the Sphinx. Computers and Graphics, 28 (4):477–484, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Adeli:1985:MGS

[Ade85] H. Adeli. Microcomputer graphics in structural design education. Computers and Graphics, 9(3):299–301, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Adeli:1986:MFA

[Ade86] Hojjat Adeli. Micrographics in failure analysis and design of structures. Computers and Graphics, 10(1):71–74, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alves:2023:ERC

[ADHC^{+}23] Tomás Alves, Tiago Delgado, Joana Henriques-Calado, Daniel Gonçalves, and Sandra Gama. Exploring the role of conscientiousness on visualizationsupported decision-making. Computers and Graphics, 111(??):47–62, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000110.

Andres:2006:I

[ADL06] Eric Andres, Guillaume Damiand, and Pascal Lienhardt. Introduction. Computers and Graphics, 30(1): 29, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002050.

Argotti:2002:DSS

[ADOR02] Yann Argotti, Larry Davis, Valerie Outters, and Jannick P. Rolland. Dynamic superimposition of synthetic objects on rigid and simpledeformable real objects. Computers and Graphics, 26 (6):919–930, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Allis:2001:VPS

[ADR01] Ned W. Allis, Jeffrey P. Dumont, and Clifford A. Reiter. Visualizing point sets, fractals, and quasicrystals using raster techniques. Computers and Graphics, 25(3):519–527, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/40/article.
pdf.

Apolinario:2023:FSS

[AdSMD23] Antonio L. Apolinário, Jefersson A. dos Santos, Fabio Miranda, and Cosimo Distante. Foreword to special section on SIBGRAPI 2022. Computers and Graphics, 114(??):A4–A6, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001711.

Avots:2016:AGR

[ADT^{+}16] Egils Avots, Morteza Daneshmand, Andres Traumann, Sergio Escalera, and Gholamreza Anbarjafari. Automatic garment retexturing based on infrared information. Computers and Graphics, 59(??):28– 38, October 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300450.

Adzhiev:2022:MPA

[Adz22] Valery Adzhiev. In memoriam: Professor Alexander Pasko (8.12.1960–15.11.2022). Computers and Graphics, 109(??):A3–A4, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322002096.

Akyuz:2013:EIR

[AEA13] Ahmet Oguz Akyüz, M. Levent Eksert, and M. Selin Aydin. An evaluation of image reproduction algorithms for high contrast scenes on large and small screen display devices. Computers and Graphics, 37(7): 885–895, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001143.

Andrews:2018:FSS

[AEJZ18] Sheldon Andrews, Kenny Erleben, Fabrice Jaillet, and Gabriel Zachmann. Foreword to the special section on VRIPHYS 2018. Computers and Graphics, 76 (??):A3–A4, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301493.

Abbi-Ezzi:1988:SGS

[AEMT88] Salim S. Abbi-Ezzi, Jorge F. Molina, and Michael A. Toelle. A special graphics system for PHIGS. Computers and Graphics, 12(2):

155–162, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Abi-Ezzi:1991:FHT

[AEW91] Salim S. Abi-Ezzi and Michael J. Wozny. Factoring a homogeneous transformation for a more efficient graphics pipeline. Computers and Graphics, 15(2): 249–258, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Akman:1989:ROR

[AF89] Varol Akman and Wm. Randolph Franklin. Representing objects as rays, or how to pile up an octree. Computers and Graphics, 13(3): 373–379, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ageenko:2000:LCL

[AF00] Eugene Ageenko and Pasi Fränti. Lossless compression of large binary images in digital spatial libraries. Computers and Graphics, 24 (1):91–98, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/13/20/47/27/35/article.
pdf.

Athanasiadis:2011:PCS

[AF11] Theodoros Athanasiadis and Ioannis Fudos. Parallel computation of spherical parameterizations for mesh analysis. Computers and Graphics, 35(3): 569–579, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000616.

Aguerre:2016:HFM

[AF16a] José Pedro Aguerre and Eduardo Fernández. A hierarchical factorization method for efficient radiosity calculations. Computers and Graphics, 60(??):46– 54, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300978.

Aliaga:2016:FSS

[AF16b] Daniel G. Aliaga and Leandro A. F. Fernandes. Foreword to the special section on SIBGRAPI 2016. Computers and Graphics, 60 (??):A1–A2, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301108.

Astheimer:1993:VDG

[AFM93] Peter Astheimer, Wolfgang Felger, and Stefan Müller. Virtual design: a generic VR system for industrial applications. Computers and Graphics, 17(6):671–677, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ao:2018:IAD

[AFW^{+}18] Xuefeng Ao, Qian Fu, Zhongke Wu, Xingce Wang, Mingquan Zhou, Quan Chen, and Hock Soon Seah. An intersection algorithm for disk B-spline curves. Computers and Graphics, 70 (??):99–107, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301140.

Arques:1994:DTD

[AG94] Didier Arques and Olivier Grange. Digital topology in 2.5 D: an application to topological filling of wellnested objects. Computers and Graphics, 18(3): 373–393, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Akyuz:2013:RCR

[AG13] Ahmet Oguz Akyüz and Asli Gençtav. A reality check for radiometric camera response recovery algorithms. Computers and Graphics, 37 (7):935–943, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001015.

Alderighi:2021:CDF

[AGM^{+}21] Thomas Alderighi, Daniela Giorgi, Luigi Malomo, Paolo Cignoni, and Monica Zoppè. Computational design, fabrication and evaluation of rubber protein models. Computers and Graphics, 98 (??):177–187, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001096.

Alhalabi:2003:ESE

[AHK03] M. Osama Alhalabi, Susumu Horiguchi, and Susumu Kunifuji. An experimental study on the effects of network delay in Cooperative Shared Haptic Virtual Environment. Computers and Graphics, 27(2): 205–213, April 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arslan:2022:PSQ

[AHR^{+}22] Mazlum Ferhat Arslan, Alexandros Haridis, Paul L. Rosin, Sibel Tari, Charlotte Brassey, James D. Gardiner, Asli Genctav, and Murat Genctav. SHREC'21: Quantifying shape complexity. Computers and Graphics, 102(??):144–153, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001916.

Aigrain:1994:ART

[AJ94] Philippe Aigrain and Philippe Joly. The automatic realtime analysis of film editing and transition effects and its applications. Computers and Graphics, 18(1): 93–103, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alonso:2023:RTR

[AJAC23] Jesús Alonso, Robert Joan-Arinyo, and Antoni Chica. Real-time rendering and physics of complex dynamic terrains modeled as CSG trees of DEMs carved with spheres. Computers and Graphics, 114(??):306– 315, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001152.

Adcock:2000:IFS

[AJRV00] Bruce M. Adcock, Kevin C.

Jones, Clifford A. Reiter, and Lisa M. Vislocky. Iterated function systems with symmetry in the hyperbolic plane. Computers and Graphics, 24(5):791– 796, October 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/37/article.
pdf.

Akagi:2006:CAS

[AK06] Y. Akagi and K. Kitajima. Computer animation of swaying trees based on physical simulation. Computers and Graphics, 30 (4):529–539, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000835.

Andrews:2013:GDM

[AK13] S. Andrews and P. G. Kry. Goal directed multi-finger manipulation: Control policies and analysis. Computers and Graphics, 37(7): 830–839, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000599.

Alexa:2015:EDM

[AK15] Marc Alexa and Jan Eric Kyprianidis. Error diffusion on meshes. Computers and Graphics, 46(??): 336–344, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000983.

Aiteanu:2021:ESS

[AK21] Fabian Aiteanu and Reinhard Klein. Exploring shape spaces of 3D tree point clouds. Computers and Graphics, 100(??):21– 31, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001448.

Asish:2022:DDS

[AKB22] Sarker Monojit Asish, Arun K.

Kulshreshth, and Christoph W. Borst. Detecting distracted students in educational VR environments using machine learning on eye gaze data. Computers and Graphics, 109(??):75– 87, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001856.

Akleman:2020:GAT

[AKF^{+}20] Ergun Akleman, Vinayak R. Krishnamurthy, Chia-An Fu, Sai Ganesh Subramanian, Matthew Ebert, Matthew Eng, Courtney Starrett, and Haard Panchal. Generalized abeille tiles: Topologically interlocked space-filling shapes generated based on fabric symmetries. Computers and Graphics, 89 (??):156–166, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300674.

Adzhiev:2000:HSA

[AKPS00] Valery Adzhiev, Maxim Kazakov, Alexander Pasko, and Vladimir Savchenko. Hybrid system architecture for volume modeling. Computers and Graphics, 24 (1):67–78, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/33/article.
pdf.

Akleman:2016:CPV

[AKW^{+}16] Ergun Akleman, Shenyao Ke, You Wu, Negar Kalantar, AliReza Borhani, and Jianer Chen. Construction with physical version of quad-edge data structures. Computers and Graphics, 58 (??):172–183, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300516.

Aoki:1978:CGR

[AL78] Masayoshi Aoki and Martin D. Levine. Computer generation of realistic pictures. Computers and Graphics, 3(4):149–161, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Aliaga:1998:STT

[AL98] D. G. Aliaga and A. A.

Lastra. Smooth transitions in texture-based simplification. Computers and Graphics, 22(1):71–81, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
1/515.pdf.

Andujar:2010:CGS

[AL10] Carlos Andújar and Javier Lluch. Computer graphics in Spain: a selection of papers from CEIG 2009. Computers and Graphics, 34(6):688, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001305.

Alander:1985:IAR

[Ala85] J. Alander. On interval arithmetic range approximation methods of polynomials and rational functions. Computers and Graphics, 9 (4):365–372, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Amati:2006:SCC

[ALC06] Giancarlo Amati, Alfredo Liverani, and Gianni Caligiana. From spline to class-A curves through multi-scale analysis filtering. Computers and Graphics, 30(3): 345–352, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000562.

Applegate:2012:SBS

[ALD12] C. S. Applegate, S. D. Laycock, and A. M. Day. A sketch-based system for highway design with userspecified regions of influence. Computers and Graphics, 36 (6):685–695, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000726.

Allan:1977:SRA

[All77] John J. Allan III. Some research advances in computer graphics that will enhance applications to engineering design. Computers and Graphics, 2(2):55–58, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Allen:1984:ISM

[All84] G. Allen. An introduction to solid modelling. Computers and Graphics, 8(4): 439–447, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arvanitis:2019:DDM

[ALM19] Gerasimos Arvanitis, Aris S.

Lalos, and Konstantinos Moustakas. Denoising of dynamic 3D meshes via lowrank spectral analysis. Computers and Graphics, 82 (??):140–151, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300780.

Attene:2017:FSI

[ALP17] Marco Attene, Sylvain Lefebvre, and Daniele Panozzo. Foreword to the special issue on Shape Modeling International 2017. Computers and Graphics, 66(??): 1–3, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300754.

Alhamazani:2023:CGS

[ALR23] Fahd Alhamazani, Yu-Kun Lai, and Paul L. Rosin.

3DCascade-GAN: Shape com-

pletion from single-view depth images. Computers and Graphics, 115(??):412– 422, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001620.

Agishtein:1991:SSR

[AM91] Michael E. Agishtein and Alexander A. Migdal. Smooth surface reconstruction from scattered data points. Computers and Graphics, 15 (1):29–39, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ariga:2010:SVD

[AM10] Taeko Ariga and Koichi Mori. Sensory visiondevelopment of a course for physical interaction and graphics. Computers and Graphics, 34(6):800–810, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001469.

Alexa:2012:IPP

[AM12] Marc Alexa and Wojciech Matusik. Irregular pit placement for dithering images by self-occlusion. Computers and Graphics, 36 (6):635–641, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000258.

Artur:2019:NVA

[AM19] Erasmo Artur and Rosane Minghim. A novel visual approach for enhanced attribute analysis and selection. Computers and Graphics, 84(??):160–172, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301438.

Asare:2003:BCG

[AMC03] Sampson D. Asare, Petros M. Mashwama, and Steve Cunningham. Building computer graphics education in southern Africa. Computers and Graphics, 27(3):369–372, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Abrikosov:2021:TAD

[AMFH21] Alexei I. Abrikosov, Talha Bin Masood, Martin Falk, and Ingrid Hotz. Topological analysis of density fields: an evaluation of segmentation methods. Computers and Graphics, 98(??):231– 241, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100114X.

Apostu:2012:AAO

[AMGA12] Oana Apostu, Frédéric Mora, Djamchid Ghazanfarpour, and Lilian Aveneau. Analytic ambient occlusion using exact frompolygon visibility. Computers and Graphics, 36 (6):727–739, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001094.

Abrevaya:2016:LOT

[AMHWW16] Victoria Fernández Abrevaya, Sandeep Manandhar, Franck Hétroy-Wheeler, and Stefanie Wuhrer. A 3D + t Laplace operator for temporal mesh sequences. Computers and Graphics, 58(??):12–22, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300619.

Aigner:2007:VTO

[AMM^{+}07] Wolfgang Aigner, Silvia Miksch, Wolfgang Müller, Heidrun Schumann, and Christian Tominski. Visualizing time-oriented data — a systematic view. Computers and Graphics, 31(3): 401–409, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000611.

Ahsan:2022:AVA

[AMPG22] Moonisa Ahsan, Fabio Marton, Ruggero Pintus, and Enrico Gobbetti. Audiovisual annotation graphs for guiding lens-based scene exploration. Computers and Graphics, 105(??):131– 145, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200070X.

Almeida:2023:SIT

[AMR23] Diogo Almeida, Daniel Mendes, and Rui Rodrigues. SIT6: Indirect touch-based object manipulation for DeskVR. Computers and Graphics, 117 (??):51–60, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002509.

Argelaguet:2022:FSS

[AMS22] Ferran Argelaguet, Ryan P. McMahan, and Maki Sugimoto. Foreword to the special section on the International Conference on Artificial Reality and Telexistence and Eurographics Symposium on Virtual Environments (ICAT-EGVE 2020). Computers and Graphics, 103(??):A5–A6, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000346.

Alia:1986:FHG

[AMT86] G. Alia, E. Martinelli, and N. Tani. Fast hardware graphic generators of curve families. Computers and Graphics, 10(1):27–36, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ayala:2024:DFM

[AMZ^{+}24] Naila Ayala, Diako Mardanbegi, Abdullah Zafar, Ewa Niechwiej-Szwedo, Shi Cao, Suzanne Kearns, Elizabeth Irving, and Andrew T. Duchowski. Does fiducial marker visibility impact task performance and information processing in novice and low-time pilots? Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000165.

Aydin:1999:CGI

[AN99] Yahya Aydin and Masayuki Nakajima. Computer graphics in India — database guided computer animation of human grasping using forward and inverse kinematics. Computers and Graphics, 23(1):145–154, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/656.pdf.

Anderson:1982:OMC

[And82] D. P. Anderson. An orientation method for central projection programs. Computers and Graphics, 6 (1):35–37, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anderson:1985:EAA

[And85] D. P. Anderson. Efficient algorithms for automatic viewer orientation. Computers and Graphics, 9(4): 407–413, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Andres:1994:DCR

[And94] Eric Andres. Discrete circles, rings and spheres. Computers and Graphics, 18(5):695–706, September– October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Andujar:1998:SEC

[And98] Carlos Andújar. Space efficient connectivity test for n-dimensional images. Computers and Graphics, 22(4): 557–558, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/4/576.pdf.

Al-nasrawi:2019:MIG

[AnD19] Mukhalad Al-nasrawi and Guang Deng. Modified iterative guided texture filtering algorithm. Computers and Graphics, 79(??): 81–100, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300093.

Andersen:2017:FLC

[ANE17] Michael Andersen, Sarah Niebe, and Kenny Erleben. A fast linear complementarity problem solver for fluid animation using high level algebra interfaces for GPU libraries. Computers and Graphics, 69(??):36– 48, December 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301589.

Anglada:1997:IIA

[Ang97] Marc Vigo Anglada. An improved incremental algorithm for constructing restricted Delaunay triangulations. Computers and Graphics, 21(2):215–223, March–April 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600085.

Angelelli:2011:IVA

[ANGH11] Paolo Angelelli, Kim Nylund, Odd Helge Gilja, and Helwig Hauser. Interactive visual analysis of contrastenhanced ultrasound data based on small neighborhood statistics. Computers and Graphics, 35(2): 218–226, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001913.

Anonymous:1977:ACP

[Ano77] Anonymous. Announcement and call for papers for Symposium on Future Trends on Computerized Structural Analysis and Synthesis. Computers and Graphics, 2(3):195, ???? 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849377900425.

Anonymous:1978:CG

[Ano78]  Anonymous.  Computer

graphics. Computers and Graphics, 3(2–3):113, ???? 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        009784937890016X.

Anonymous:1980:LCA

[Ano80] Anonymous. List of contents and author index. Computers and Graphics, 5(2– 4):i–ii, ???? 1980. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849380900175.

Anonymous:1983:CGA

[Ano83a]  Anonymous.  Computer

graphics art section. Computers and Graphics, 7(1): 103–104, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900559.

Anonymous:1983:GEI

[Ano83b] Anonymous. Guest Editor’s introduction. Computers and Graphics, 7(3– 4):213–214, ???? 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900158.

Anonymous:1983:LCA

[Ano83c] Anonymous. List of contents and author index. Computers and Graphics, 7(3– 4):i–v, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900365.

Anonymous:1985:CPC

[Ano85a] Anonymous. Call for papers — Computers & graphics. Computers and Graphics, 9(3):331–332, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900615.

Anonymous:1985:CPEa

[Ano85b] Anonymous. Call for papers — Eurographics ’86. Computers and Graphics, 9 (3):333–335, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900627.

Anonymous:1985:CPEb

[Ano85c] Anonymous. Call for papers: European Computer Graphics Conference and Exhibition Eurographics’86. Computers and Graphics, 9 (4):457–459, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900421.

Anonymous:1985:EB

[Ano85d]  Anonymous.  Editorial

Board. Computers and Graphics, 9(1):ifc, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0097849385900019.

Anonymous:1985:LCA

[Ano85e] Anonymous. List of contents and author index. Computers and Graphics, 9(4): iii–vii, ???? 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900433.

Anonymous:1985:TM

[Ano85f] Anonymous. Turncae on micro. Computers and Graphics, 9(1):79, ???? 1985.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900147.

Anonymous:1985:WCC

[Ano85g] Anonymous. World Congress on Computational Mechanics. Computers and Graphics, 9(4):455, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784938590041X.

Anonymous:1986:ATS

[Ano86a] Anonymous. Artistry and technology — the Steelcase Strafor vision. Computers and Graphics, 10 (3):263, 265, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900117.

Anonymous:1986:CCN

[Ano86b] Anonymous. Call for contributions to the new software survey section to Computers & Graphics. Computers and Graphics, 10(3): i–iii, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900166.

Anonymous:1986:CP

[Ano86c] Anonymous. Call for papers. Computers and Graphics, 10(2):199, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900476.

Anonymous:1986:CPF

[Ano86d] Anonymous. Call for papers: First Eurographics Workshop on Intelligent CAD Systems. Computers and Graphics, 10 (3):271–272, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900142.

Anonymous:1986:CI

[Ano86e]  Anonymous.  CG Inter-

national ’87. Computers and Graphics, 10(3): 273, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900154.

Anonymous:1986:EB

[Ano86f]  Anonymous.  Editorial

Board. Computers and Graphics, 10(1):ifc, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0097849386900622.

Anonymous:1986:E

[Ano86g]  Anonymous.  Eurograph-

ics ’86. Computers and Graphics, 10(2):201–203, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900488.

Anonymous:1986:EAC

[Ano86h]  Anonymous.  Eurograph-

ics ’87 — Amsterdam call for participation. Computers and Graphics, 10(3): 269, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900130.

Anonymous:1986:SSS

[Ano86i] Anonymous. Software survey section. Computers and Graphics, 10(2): i–iv, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784938690049X.

Anonymous:1986:SFD

[Ano86j] Anonymous. Symposium on Future Directions of Computational Mechanics. Computers and Graphics, 10(3): 267, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900129.

Anonymous:1991:EIC

[Ano91a] Anonymous. Eurographics ’91 — international conference on computer graphics and industrial exhibition. Computers and Graphics, 15 (3):457–??, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1991:FIC

[Ano91b] Anonymous. Fifth international conference on humancomputer interaction (HCI international ’93). Computers and Graphics, 15(3): 459–??, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1991:FIS

[Ano91c] Anonymous. Fourth international school — Microcomputer ’91. Computers and Graphics, 15(3):458, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1991:SHP

[Ano91d] Anonymous. Symposium on high-performance computing for flight vehicles. Computers and Graphics, 15(3):458, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1993:EWR

[Ano93a] Anonymous. 5th Eurographics Workshop on Rendering — call for contributions. Computers and Graphics, 17(6):717–??, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1993:CHS

[Ano93b] Anonymous. Conference on High-Speed Networking and Multimedia Applications. Computers and Graphics, 17(6):716–??, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1993:ECP

[Ano93c] Anonymous. Eurographics ’94 — call for participation. Computers and Graphics, 17(6):719–??, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1993:ICG

[Ano93d]  Anonymous.  Interna-

tional Computer Graphics Prize 1994. Computers and Graphics, 17(6):718–??, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1993:MVS

[Ano93e]  Anonymous.  Modelling

and Visualisation of Spatial Data in Geographical Information Systems — call for papers. Computers and Graphics, 17(6):715–??, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1994:IIW

[Ano94a] Anonymous. 2nd IWACA 2nd International Workshop on Advanced Teleservices and High-Speed Communication Architectures. Computers and Graphics, 18(4): 605–606, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900795.

Anonymous:1994:EWGb

[Ano94b] Anonymous. The 9th Eurographics Workshop on Graphics Hardware. Computers and Graphics, 18 (4):604, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900787.

Anonymous:1994:AMS

[Ano94c] Anonymous. ACM Multimedia ’94: The second ACM international conference on multimedia. Computers and Graphics, 18(4): 611–612, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900833.

Anonymous:1994:CCFa

[Ano94d] Anonymous. Call for contributions: Fifth Eurographics Workshop on Rendering. Computers and Graphics, 18(1):139, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901287.

Anonymous:1994:CCFb

[Ano94e] Anonymous. Call for contributions: Fifth Eurographics Workshop on Visualization in Scientific Computing. Computers and Graphics, 18(1):141, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901309.

Anonymous:1994:CPSa

[Ano94f] Anonymous. Call for participation 1994 Symposium on Volume Visualization. Computers and Graphics, 18(3):435, May/June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900493.

Anonymous:1994:CPSb

[Ano94g] Anonymous. Call for participation 1994 Symposium on Volume Visualization. Computers and Graphics, 18 (4):608, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900817.

Anonymous:1994:CPE

[Ano94h]  Anonymous.  Call for

participation Eurographics ’94. Computers and Graphics, 18(1):137–138, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901279.

Anonymous:1994:CGP

[Ano94i]  Anonymous.  Computer

graphics prize. Computers and Graphics, 18 (1):142, January/February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
0097849394901317.

Anonymous:1994:CGT

[Ano94j]  Anonymous.  Computer

graphics tutorial. Computers and Graphics, 18 (1):140, January/February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0097849394901295.

Anonymous:1994:EB

[Ano94k]  Anonymous.  Editorial

Board. Computers and Graphics, 18(1):ifc, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901082.

Anonymous:1994:E

[Ano94l] Anonymous. EG ’94. Computers and Graphics, 18 (4):602, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900760.

Anonymous:1994:EWGa

[Ano94m]  Anonymous.  Eurograph-

ics Workshop on Graphics and Visualization Education (GVE). Computers and Graphics, 18(4):603, July/August 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900779.

Anonymous:1994:FEWa

[Ano94n] Anonymous. Fifth Eurographics Workshop on Animation and Simulation. Computers and Graphics, 18(3):431, May/June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900469.

Anonymous:1994:FEWb

[Ano94o] Anonymous. Fifth Eurographics Workshop on Animation and Simulation. Computers and Graphics, 18 (4):601, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900752.

Anonymous:1994:FAC

[Ano94p] Anonymous. First announcement and call for papers CAPE ’95: the Fifth International Conference on Computer Applications in Production and Engineering. Computers and Graphics, 18(3):440, May/June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900531.

Anonymous:1994:FIIa

[Ano94q] Anonymous. First IEEE international conference on image processing. Computers and Graphics, 18(3): 439, May/June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900523.

Anonymous:1994:FIIb

[Ano94r] Anonymous. First IEEE International Conference on Image Processing. Computers and Graphics, 18 (4):614, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784939490085X.

Anonymous:1994:IT

[Ano94s] Anonymous. IEEE Tencon ’94. Computers and Graphics, 18(3):429, May/ June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900442.

Anonymous:1994:ICP

[Ano94t] Anonymous. International conference Pacific Graphics ’94/CADDM’94. Computers and Graphics, 18(3): 430, May/June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900450.

Anonymous:1994:IDCa

[Ano94u] Anonymous. International dedicated conference on robotics, motion and machine vision. Computers and Graphics, 18(3): 436, May/June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900507.

Anonymous:1994:IDCb

[Ano94v] Anonymous. International dedicated conference on robotics, motion and machine vision. Computers and Graphics, 18(4):613, July/August 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900841.

Anonymous:1994:LR

[Ano94w] Anonymous. List of 1993 reviewers. Computers and Graphics, 18(1):135–136, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901260.

Anonymous:1994:MTT

[Ano94x]  Anonymous.  Multime-

dia transport and teleservices. Computers and Graphics, 18(3):437–438, May/June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900515.

Anonymous:1994:MT

[Ano94y]  Anonymous.  Multime-

dia transport and teleservices. Computers and Graphics, 18(4):615–616, July/ August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900868.

Anonymous:1994:PFIa

[Ano94z] Anonymous. Past/future issues. Computers and Graphics, 18(1):133, January/February 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394901252.

Anonymous:1994:PFIb

[Ano94-27] Anonymous. Past/future issues. Computers and Graphics, 18(3):427, May/ June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849394900434.

Anonymous:1994:PFIc

[Ano94-28] Anonymous. Past/future issues. Computers and Graphics, 18(4):599, July/ August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849394900744.

Anonymous:1994:RIMa

[Ano94-29]  Anonymous.  RIAO ’94

Intelligent multimedia information retrieval systems and management. Computers and Graphics, 18(3): 433–434, May/June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849394900485.

Anonymous:1994:RIMb

[Ano94-30]  Anonymous.  RIAO ’94

Intelligent multimedia information retrieval systems and management. Computers and Graphics, 18(4): 609–610, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849394900825.

Anonymous:1994:WICb

[Ano94-31]  Anonymous.  Workshop

on integration of CAtechniques in theory and practice ICA ’94. Computers and Graphics, 18 (4):607, July/August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900809.

Anonymous:1994:WICa

[Ano94-32]  Anonymous.  Workshop

on integration of CAtechniques in theory and practice/ICA ’94. Computers and Graphics, 18(3): 432, May/June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849394900477.

Anonymous:1995:PWDa

[Ano95a] Anonymous. 1996 Pacific Workshop on Distributed Multimedia Systems (DMS ’96) — call for participation. Computers and Graphics, 19(4):649–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PWDb

[Ano95b] Anonymous. 1996 Pacific Workshop on Distributed Multimedia Systems (DMS ’96) — call for participation.

Computers and Graphics, 19(5):791–??, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PWDc

[Ano95c] Anonymous. 1996 Pacific Workshop on Distributed Multimedia Systems (DMS ’96) — call for participation. Computers and Graphics, 19(6):902–??, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:Aa

[Ano95d]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):169, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849395900330.

Anonymous:1995:Ab

[Ano95e]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):170, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849395900349.

Anonymous:1995:Ac

[Ano95f]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):171, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900357.

Anonymous:1995:Ad

[Ano95g]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):172, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900365.

Anonymous:1995:Ae

[Ano95h]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):173, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900373.

Anonymous:1995:Af

[Ano95i]  Anonymous.  Announce-

ments. Computers and Graphics, 19(1):174–175, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900381.

Anonymous:1995:CF

[Ano95j] Anonymous. CAAD Futures ’95. Computers and Graphics, 19(2):340–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CGFa

[Ano95k] Anonymous. CAD/Graphics ’95: The Fourth International Conference on CAD and CG. Computers and Graphics, 19(2):341–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CGFb

[Ano95l] Anonymous. CAD/Graphics ’95: The Fourth International Conference on CAD and CG. Computers and Graphics, 19(3):484–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CGFc

[Ano95m] Anonymous. CAD/GRAPHICS ’95: The Fourth International Conference on CAD and CG. Computers and Graphics, 19(4):639–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CIa

[Ano95n]  Anonymous.  CG Inter-

national ’96. Computers and Graphics, 19(4):648–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CIb

[Ano95o]  Anonymous.  CG Inter-

national ’96. Computers and Graphics, 19(5):790–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CIc

[Ano95p]  Anonymous.  CG Inter-

national ’96. Computers and Graphics, 19(6):901–??, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CSTa

[Ano95q] Anonymous. CSG ’96 — Settheoretic Solid Modelling: Techniques and Applications — conference announcement and call for papers. Computers and Graphics, 19(4): 645–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CSTb

[Ano95r] Anonymous. CSG ’96 — Settheoretic Solid Modelling:

Techniques and Applications

— conference announcement and call for papers. Computers and Graphics, 19(5): 787–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:CSTc

[Ano95s] Anonymous. CSG ’96 — Settheoretic Solid Modelling: Techniques and Applications

— conference announcement and call for papers. Computers and Graphics, 19(6):898–??, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:EB

[Ano95t]  Anonymous.  Editorial

Board. Computers and Graphics, 19(1):ifc, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900284.

Anonymous:1995:EGM

[Ano95u]  Anonymous.  Eurograph-

ics ’95: Graphics. Multimedia. Virtual Reality. Computers and Graphics, 19(2): 339–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:GIC

[Ano95v]  Anonymous.  GraphiCon

’95: The 5th International Conference on Computer Graphics and Visualization in Russia. Computers and Graphics, 19(2): 337–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IVRc

[Ano95w] Anonymous. IEEE Virtual Reality Annual International Symposium 1996 — call for participation. Computers and Graphics, 19(6):897–??, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IVRa

[Ano95x]  Anonymous.  IEEE Vir-

tual Reality Annual International Symposium 1996 (VRAIS ’96) — Call for Participation. Computers and Graphics, 19(4):644–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IVRb

[Ano95y]  Anonymous.  IEEE Vir-

tual Reality Annual International Symposium 1996 (VRAIS ’96) — call for participation. Computers and Graphics, 19(5):786–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IVCa

[Ano95z] Anonymous. IEEE Visualization ’95 — call for participation. Computers and Graphics, 19(3):486–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IVCb

[Ano95-27] Anonymous. IEEE Visualization ’95 — Call for Participation. Computers and Graphics, 19(4):641–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:IIV

[Ano95-28] Anonymous. IMC ’96 — Information Visualization and Mobile Computing. Computers and Graphics, 19(5): 785–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:ICEa

[Ano95-29] Anonymous. International Conference on Engineering Computation and Computer Simulation (ECCS ’95). Computers and Graphics, 19 (3):488–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:ICEb

[Ano95-30] Anonymous. International Conference on Engineering Computation and Computer Simulation (ECCS ’95). Computers and Graphics, 19(4):643–??, July– August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:ISS

[Ano95-31] Anonymous. International Symposium on Scientific Visualization (ISSV). Computers and Graphics, 19 (3):483–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:LR

[Ano95-32] Anonymous. List of 1994 reviewers. Computers and Graphics, 19(1):167–168, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849395900322.

Anonymous:1995:PFIa

[Ano95-33] Anonymous. Past/future issues. Computers and Graphics, 19(1):165, January/February 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849395900314.

Anonymous:1995:PFIb

[Ano95-34] Anonymous. Past/future issues. Computers and Graphics, 19(2):335–??, March– April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PFIc

[Ano95-35] Anonymous. Past/future issues. Computers and Graphics, 19(3):481–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PFId

[Ano95-36] Anonymous. Past/future issues. Computers and Graphics, 19(4):637–??, July– August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PFIe

[Ano95-37] Anonymous. Past/future issues. Computers and Graphics, 19(5):783–??, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:PFIf

[Ano95-38] Anonymous. Past/future issues. Computers and Graphics, 19(6):895–??, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:TIIb

[Ano95-39] Anonymous. The Third IEEE International Conference on Multi-media Computing and Systems — call for participation. Computers and Graphics, 19(5):789–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:TIIc

[Ano95-40] Anonymous. The Third IEEE International Conference on Multi-media Computing and Systems — call for participation. Computers and Graphics, 19(6):900–??, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:TIIa

[Ano95-41] Anonymous. The Third IEEE International Conference on Multimedia Computing and Systems — call for participation. Computers and Graphics, 19(4): 647–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1995:VLC

[Ano95-42] Anonymous. Volume 19 list of contents and author index. Computers and Graphics, 19(6):I–??, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CPC

[Ano96a] Anonymous. Call for papers: 1997 computers and graphics special issue on haptic displays in virtual environments. Computers and Graphics, 20(6):IV–??, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CGPa

[Ano96b]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 20(2): III–??, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CGPb

[Ano96c]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 20(3):III– ??, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CGPc

[Ano96d]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 20(4): III–??, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CGPd

[Ano96e]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 20(5):III–??, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:CGPe

[Ano96f]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 20(6):III–??, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:PFIa

[Ano96g] Anonymous. Past/future issues. Computers and Graphics, 20(2):I–??, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:PFIb

[Ano96h] Anonymous. Past/future issues. Computers and Graphics, 20(3):I–??, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:PFIc

[Ano96i] Anonymous. Past/future issues. Computers and Graphics, 20(4):I–??, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:PFId

[Ano96j] Anonymous. Past/future issues. Computers and Graphics, 20(5):I–??, September– October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1996:PFIe

[Ano96k] Anonymous. Past/future issues. Computers and Graphics, 20(6):I–??, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:CGPa

[Ano97a]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 21(1):III– ??, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:CGPb

[Ano97b]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 21(3):III– ??, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:CGPc

[Ano97c]  Anonymous.  Computers

and graphics is planning a new section. Computers and Graphics, 21(4): iii–??, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:CGPd

[Ano97d]  Anonymous.  Computers

and Graphics is planning a new section. Computers and Graphics, 21(6):III–??, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:LR

[Ano97e] Anonymous. List of 1996 reviewers. Computers and Graphics, 21(1):III– ??, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:PFIa

[Ano97f] Anonymous. Past/future issues. Computers and Graphics, 21(1):I–??, January– February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:PFIb

[Ano97g] Anonymous. Past/future issues. Computers and Graphics, 21(3):I–??, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:PFIc

[Ano97h] Anonymous. Past/future issues. Computers and Graphics, 21(4):i–??, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1997:PFId

[Ano97i] Anonymous. Past/future issues. Computers and Graphics, 21(6):I–??, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1998:PFIa

[Ano98a] Anonymous. Past/future issues. Computers and Graphics, 22(4):I–??, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1998:PFIb

[Ano98b] Anonymous. Past/future issues. Computers and Graphics, 22(5):i–??, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1998:PFIc

[Ano98c] Anonymous. Past/future issues. Computers and Graphics, 22(6):783–??, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1999:CGPa

[Ano99a] Anonymous. Chaos and graphics — past/future issues. Computers and Graphics, 23(2):295–??, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1999:CGPb

[Ano99b] Anonymous. Chaos and graphics — past/future issues. Computers and Graphics, 23(3):455–??, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1999:CGB

[Ano99c] Anonymous. Computers and Graphics Best Paper Award. Computers and Graphics, 23 (1):1–??, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1999:Ea

[Ano99d] Anonymous. Editorial(s).

Computers and Graphics, 23 (4):467–468, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/27/article.
pdf.

Anonymous:1999:Eb

[Ano99e] Anonymous. Editorial(s). Computers and Graphics, 23 (5):633–634, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
27/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/27/abstract.
html.

Anonymous:1999:Ec

[Ano99f] Anonymous. Editorial(s). Computers and Graphics, 23 (6):777–778, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/27/article.
pdf.

Anonymous:1999:Ed

[Ano99g] Anonymous. Editorial(s). Computers and Graphics, 23(6):839, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/37/article.
pdf.

Anonymous:1999:I

[Ano99h] Anonymous. Index. Computers and Graphics, 23 (6):ix–x, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
48/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/48/article.
pdf.

Anonymous:1999:LR

[Ano99i] Anonymous. List of 1998 reviewers. Computers and Graphics, 23(1):177–??, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1999:PFI

[Ano99j] Anonymous. Past/future issues. Computers and Graphics, 23(1):175–??, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2000:E

[Ano00a] Anonymous. Editorial(s).

Computers and Graphics, 24(3):321–324, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/27/article.
pdf.

Anonymous:2000:I

[Ano00b] Anonymous. Index. Computers and Graphics, 24 (6):XI–XII, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/36/article.
pdf.

Anonymous:2001:SSV

[Ano01a] Anonymous. 4th SBC Symposium on Virtual Reality. Computers and Graphics, 25(4):715, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/33/47/abstract.
html.

Anonymous:2001:FW

[Ano01b]  Anonymous.  6th Fall

workshop. Computers and Graphics, 25(4):722, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
13/20/57/33/54/abstract.
html.

Anonymous:2001:ICPb

[Ano01c] Anonymous. 6th International Conference on protocols for multimedia systems — PROMS 2001. Computers and Graphics, 25(4): 716, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/48/abstract.
html.

Anonymous:2001:ASV

[Ano01d] Anonymous. ACM Symposium on Virtual Reality Software & Technology 2001 Organized and Sponsored by ACM, SIGGRAPH, and SIG-CHI. Computers and Graphics, 25(5):920, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/47/abstract.
html.

Anonymous:2001:AC

[Ano01e] Anonymous. AFRIGRAPH 2001 CONFERENCE. Computers and Graphics, 25(5): 918, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        13/20/57/34/45/abstract.
                                        html.

Anonymous:2001:ASS

[Ano01f]  Anonymous.  Announce-

ment: 4th SBC symposium on virtual reality. Computers and Graphics, 25 (3):546, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        49/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/49/article.
                                        pdf.

Anonymous:2001:AAM

[Ano01g]  Anonymous.  Announce-

ment: ACM multimedia 2001. Computers and Graphics, 25(3):543, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        46/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/46/article.
                                        pdf.

Anonymous:2001:ACGb

[Ano01h]  Anonymous.  Announce-

ment: CAD/graphics’ 2001. Computers and Graphics, 25(2):353–354, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/13/20/57/29/
43/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/43/article.
pdf.

Anonymous:2001:ACP

[Ano01i]  Anonymous.  Announce-

ment: call for participation. Computers and Graphics, 25(3):540, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
43/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/43/article.
pdf.

Anonymous:2001:ACGa

[Ano01j]  Anonymous.  Announce-

ment: Computer Grapics International 2001. Computers and Graphics, 25 (2):352, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/42/article.
pdf.

Anonymous:2001:AEa

[Ano01k]  Anonymous.  Announce-

ment: Eurographics 2001. Computers and Graphics, 25 (2):356, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        45/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/45/article.
                                        pdf.

Anonymous:2001:AEb

[Ano01l]  Anonymous.  Announce-

ment: Eurographics 2001. Computers and Graphics, 25 (3):542, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        45/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/45/article.
                                        pdf.

Anonymous:2001:AIb

[Ano01m]  Anonymous.  Announce-

ment: ICHIM 2001. Computers and Graphics, 25 (2):355, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        44/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/44/article.
                                        pdf.

Anonymous:2001:AId

[Ano01n]  Anonymous.  Announce-

ment: ICHIM 2001. Computers and Graphics, 25 (3):541, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
44/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/44/article.
pdf.

Anonymous:2001:AIa

[Ano01o]  Anonymous.  Announce-

ment: ICPNM. Computers and Graphics, 25(2): 351, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/41/article.
pdf.

Anonymous:2001:AIP

[Ano01p]  Anonymous.  Announce-

ment: IEEE 2001 (PVG 2001). Computers and Graphics, 25(3):548, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
51/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/51/article.
pdf.

Anonymous:2001:AIIa

[Ano01q]  Anonymous.  Announce-

ment: IEEE ICIP 2001. Computers and Graphics, 25 (2):357, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
46/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/46/article.
pdf.

Anonymous:2001:AIIb

[Ano01r]  Anonymous.  Announce-

ment: IEEE ICIP 2001. Computers and Graphics, 25 (3):544, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
47/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/47/article.
pdf.

Anonymous:2001:AIVa

[Ano01s]  Anonymous.  Announce-

ment: IEEE visualization 2001. Computers and Graphics, 25(2):358, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
47/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/47/article.
pdf.

Anonymous:2001:AIVb

[Ano01t]  Anonymous.  Announce-

ment: IEEE Visualization 2001. Computers and Graphics, 25(3):547, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
50/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/50/article.
pdf.

Anonymous:2001:AICb

[Ano01u]  Anonymous.  Announce-

ment: International certificate program for New Media (ICPNM). Computers and Graphics, 25(3):539, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/42/article.
pdf.

Anonymous:2001:AICc

[Ano01v]  Anonymous.  Announce-

ment: International Certificate Program for New Media (ICPNM). Computers and Graphics, 25(6):1049, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        13/20/57/35/37/abstract.
                                        html.

Anonymous:2001:AICa

[Ano01w]  Anonymous.  Announce-

ment: International conference on Web delivering of music. Computers and Graphics, 25(2):359, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        48/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/48/article.
                                        pdf.

Anonymous:2001:APFa

[Ano01x]  Anonymous.  Announce-

ment: Past/future issues. Computers and Graphics, 25 (2):360, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        49/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/49/article.
                                        pdf.

Anonymous:2001:APFc

[Ano01y]  Anonymous.  Announce-

ment: Past/future issues. Computers and Graphics, 25 (6):1052, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-

ng/10/13/20/57/35/40/abstract.
html.

Anonymous:2001:APFb

[Ano01z]  Anonymous.  Announce-

ment: Past/future issues (25/3). Computers and Graphics, 25(3):550, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
53/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/53/article.
pdf.

Anonymous:2001:AS

[Ano01-27]  Anonymous.  Announce-

ment: SIBGRAPI 2001. Computers and Graphics, 25 (3):545, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
48/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/48/article.
pdf.

Anonymous:2001:ASC

[Ano01-28]  Anonymous.  Announce-

ment: SPIE Conference on visualization and data analysis 2002. Computers and Graphics, 25(6):1050, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
13/20/57/35/38/abstract.
html.

Anonymous:2001:AW

[Ano01-29]  Anonymous.  Announce-

ment: WEDELMUSIC 2001. Computers and Graphics, 25 (3):549, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
52/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/52/article.
pdf.

Anonymous:2001:AIVc

[Ano01-30] Anonymous. Author index: Volume 25. Computers and Graphics, 25(6):XI–XII, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/41/abstract.
html.

Anonymous:2001:CP

[Ano01-31] Anonymous. Call for Papers. Computers and Graphics, 25(1):184, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
51/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/51/article.
pdf.

Anonymous:2001:CPW

[Ano01-32] Anonymous. Call for Papers:

WSCG 2002. Computers and Graphics, 25(2):361, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
50/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/50/article.
pdf.

Anonymous:2001:CAa

[Ano01-33] Anonymous. Computer animation 2001. Computers and Graphics, 25(4): 721, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/53/abstract.
html.

Anonymous:2001:CAb

[Ano01-34] Anonymous. COMPUTER ANIMATION 2001. Computers and Graphics, 25(5): 919, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/46/abstract.
html.

Anonymous:2001:CGI

[Ano01-35]  Anonymous.  Computer

Graphics International 2001. Computers and Graphics, 25(1):180, February 2001.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
47/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/47/article.
pdf.

Anonymous:2001:CGB

[Ano01-36]  Anonymous.  Computers

& Graphics Best Paper Award. Computers and Graphics, 25(6):iii–vi, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/27/abstract.
html.

Anonymous:2001:EMW

[Ano01-37] Anonymous. ED-MEDIA 2001, world conference on education. Computers and Graphics, 25(1): 179, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
46/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/46/article.
pdf.

Anonymous:2001:ECC

[Ano01-38]  Anonymous.  Eurograph-

ics 2001, challenges in computer. Computers and Graphics, 25(1):181, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
48/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/48/article.
pdf.

Anonymous:2001:F

[Ano01-39] Anonymous. FEATS 2001.

Computers and Graphics, 25(1):178, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
45/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/45/article.
pdf.

Anonymous:2001:I

[Ano01-40] Anonymous. ICPNM. Computers and Graphics, 25(1): 174, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/41/article.
pdf.

Anonymous:2001:ISP

[Ano01-41] Anonymous. IEEE 2001 Symposium on parallel and large-data visualization and graphics (PVG 2001). Computers and Graphics, 25(4): 718, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/50/abstract.
html.

Anonymous:2001:IAI

[Ano01-42]  Anonymous.  IEEE and

ACM International symposium on augmented reality 2001. Computers and Graphics, 25(4):720, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/52/abstract.
html.

Anonymous:2001:IV

[Ano01-43] Anonymous. IEEE Visualization 2001. Computers and Graphics, 25(4): 717, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/49/abstract.
html.

Anonymous:2001:ICPa

[Ano01-44] Anonymous. International Certificate Program for New Media (ICPNM). Computers and Graphics, 25(4): 713, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/45/abstract.
html.

Anonymous:2001:ICPc

[Ano01-45] Anonymous. International Certificate Program for New Media (ICPNM). Computers and Graphics, 25(5): 917, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/44/abstract.
html.

Anonymous:2001:ICWa

[Ano01-46] Anonymous. International Conference on Web Delivering of Music. Computers and Graphics, 25(4): 723, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/55/abstract.
html.

Anonymous:2001:ICWb

[Ano01-47] Anonymous. International Conference on Web Delivering of Music. Computers and Graphics, 25(5):922, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/49/abstract.
html.

Anonymous:2001:LR

[Ano01-48] Anonymous. List of reviewers in 19991/2000. Computers and Graphics, 25 (1):171–173, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/26/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/40/article.
                                        pdf.

Anonymous:2001:MF

[Ano01-49] Anonymous. Media futures. Computers and Graphics, 25(1):176, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/26/
                                        43/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/43/article.
                                        pdf.

Anonymous:2001:MCN

[Ano01-50]  Anonymous.  Multimedia

Computing and Networking 2002 (MMCN02). Computers and Graphics, 25(5): 923, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/34/50/abstract.
                                        html.

Anonymous:2001:PFIa

[Ano01-51] Anonymous. Part/future issues to: 25/1. Computers and Graphics, 25(1): 185, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
52/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/52/article.
pdf.

Anonymous:2001:PFIb

[Ano01-52] Anonymous. Past/future issues. Computers and Graphics, 25(4):725, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/33/57/abstract.
html.

Anonymous:2001:PFIc

[Ano01-53] Anonymous. Past/future issues. Computers and Graphics, 25(5):928, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/55/abstract.
html.

Anonymous:2001:SMA

[Ano01-54] Anonymous. Shape modelling and applications. Computers and Graphics, 25(1):175, February 2001.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/42/article.
pdf.

Anonymous:2001:Sa

[Ano01-55]  Anonymous.  SIBGRAPI

2001. Computers and Graphics, 25(1):182, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
49/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/49/article.
pdf.

Anonymous:2001:Sb

[Ano01-56]  Anonymous.  SIBGRAPI

2001. Computers and Graphics, 25(4):714, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/33/46/abstract.
html.

Anonymous:2001:SMS

[Ano01-57] Anonymous. Solid Modelling 2001, Sixth ACM Symposium. Computers and Graphics, 25(1):177, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
44/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/44/article.
pdf.

Anonymous:2001:SCV

[Ano01-58] Anonymous. SPIE Conference on Visualization and Data Analysis 2002. Computers and Graphics, 25(5): 924, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/51/abstract.
html.

Anonymous:2001:AMM

[Ano01-59]  Anonymous.  The 10th

Annual Medicine Meets Virtual Reality Conference. Computers and Graphics, 25(6):1051, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/39/abstract.
html.

Anonymous:2001:VMV

[Ano01-60] Anonymous. Vision, Modeling and Visualization 2001 (VMV01). Computers and Graphics, 25(5):921, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        13/20/57/34/48/abstract.
                                        html.

Anonymous:2001:VG

[Ano01-61] Anonymous. Volume Graphics 2001. Computers and Graphics, 25(1):183, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/26/
                                        50/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/50/article.
                                        pdf.

Anonymous:2001:WTV

[Ano01-62]  Anonymous.  Walking

through the view of Delft

— on Internet. Computers and Graphics, 25(5):927, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/34/54/abstract.
                                        html.

Anonymous:2001:WN

[Ano01-63] Anonymous. WEB NET 2001. Computers and Graphics, 25(4):719, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/57/33/51/abstract.
                                        html.

Anonymous:2001:WC

[Ano01-64] Anonymous. WEB3D 2002 Conference. Computers and Graphics, 25(5):926, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/53/abstract.
html.

Anonymous:2001:WIC

[Ano01-65] Anonymous. WSCG’2002 — the 10th international conference on computer graphics. Computers and Graphics, 25(4):724, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/33/56/abstract.
html.

Anonymous:2001:WFW

[Ano01-66] Anonymous. WSCG’2002 (formerly the Winter School of Computer Graphics). Computers and Graphics, 25(5):925, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/34/52/abstract.
html.

Anonymous:2002:FWV

[Ano02a] Anonymous. 7th Fall Workshop on Vision, Modeling, and Visualization 2002 (VMV2002). Computers and Graphics, 26(4):648, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/51/abstract.
html.

Anonymous:2002:AMa

[Ano02b] Anonymous. ACM MUL-TIMEDIA 2002. Computers and Graphics, 26(4):648, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/52/abstract.
html.

Anonymous:2002:AMb

[Ano02c] Anonymous. ACM Multimedia 2002. Computers and Graphics, 26(5):826, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/46/abstract.
html.

Anonymous:2002:AV

[Ano02d] Anonymous. ACM VRST 2002. Computers and Graphics, 26(2):388, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
gej-ng/10/13/20/68/41/
51/abstract.html.

Anonymous:2002:AAV

[Ano02e]  Anonymous.  Announce-

ment: ACM VRST 2002. Computers and Graphics, 26 (4):647, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/55/50/abstract.
html.

Anonymous:2002:AAC

[Ano02f]  Anonymous.  Announce-

ment: Advanced Concepts for Intelligent Vision Systems. Computers and Graphics, 26(3):528, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/45/abstract.
html.

Anonymous:2002:ACVa

[Ano02g]  Anonymous.  Announce-

ment: Conference on Visualization and Data Analysis (VDA 2003). Computers and Graphics, 26(5):827, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/48/abstract.
html.

Anonymous:2002:ACVb

[Ano02h]  Anonymous.  Announce-

ment: Conference on Visualization and Data Analysis (VDA 2003). Computers and Graphics, 26(6): 982, December 2002. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:AGH

[Ano02i]  Anonymous.  Announce-

ment: Graphics hardware 2002. Computers and Graphics, 26(3):527, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/43/abstract.
html.

Anonymous:2002:AI

[Ano02j]  Anonymous.  Announce-

ment: ICPNM. Computers and Graphics, 26(4):645, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/46/abstract.
html.

Anonymous:2002:AIV

[Ano02k]  Anonymous.  Announce-

ment: IEEE Visualization 2002. Computers and Graphics, 26(4):647, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

13/20/68/55/49/abstract.
html.

Anonymous:2002:AICb

[Ano02l]  Anonymous.  Announce-

ment: International Conference on Pervasive Computing. Computers and Graphics, 26(3):526, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/54/42/abstract.
html.

Anonymous:2002:APG

[Ano02m]  Anonymous.  Announce-

ment: PACIFIC GRAPH-ICS 2002. Computers and Graphics, 26(4):646, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/48/abstract.
html.

Anonymous:2002:API

[Ano02n] Anonymous. Announcement past issues/future issues. Computers and Graphics, 26(6):983, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:APFa

[Ano02o]  Anonymous.  Announce-

ment: Past/future issues. Computers and Graphics, 26(3):531, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/54/47/abstract.
html.

Anonymous:2002:APFb

[Ano02p]  Anonymous.  Announce-

ment: Past/future issues. Computers and Graphics, 26 (5):829, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/56/49/abstract.
html.

Anonymous:2002:AS

[Ano02q]  Anonymous.  Announce-

ment: SIBGRAPI 2002. Computers and Graphics, 26 (4):646, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/55/47/abstract.
html.

Anonymous:2002:AW

[Ano02r]  Anonymous.  Announce-

ment: WEDELMUSIC 2002. Computers and Graphics, 26(4):649, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/53/abstract.
html.

Anonymous:2002:AICa

[Ano02s]  Anonymous.  Announce-

ments: International Certificate Program for New Media (ICPNM). Computers and Graphics, 26(3):525, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/41/abstract.
html.

Anonymous:2002:BPA

[Ano02t] Anonymous. Best paper award 2001. Computers and Graphics, 26(6):831– 836, December 2002. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:B

[Ano02u]  Anonymous.  Biography.

Computers and Graphics, 26 (4):534, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/55/30/abstract.
html.

Anonymous:2002:CGI

[Ano02v]  Anonymous.  Computer

graphics international 2002 (CGI2002). Computers and Graphics, 26(2):386, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.

                                        elsevier.com/gej-ng/10/
                                        13/20/68/41/45/abstract.
                                        html.

Anonymous:2002:CAI

[Ano02w]  Anonymous.  Contents/

author index for vol 26. Computers and Graphics, 26(6):??, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:CVa

[Ano02x]  Anonymous.  Curriculum

vitae. Computers and Graphics, 26(4):534, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/29/abstract.
                                        html.

Anonymous:2002:CVb

[Ano02y]  Anonymous.  Curriculum

vitae. Computers and Graphics, 26(4):590, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/36/abstract.
                                        html.

Anonymous:2002:CVc

[Ano02z]  Anonymous.  Curriculum

vitae. Computers and Graphics, 26(4):591, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/37/abstract.
html.

Anonymous:2002:CVd

[Ano02-27]  Anonymous.  Curriculum

vitae. Computers and Graphics, 26(5):655, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/27/abstract.
html.

Anonymous:2002:CVe

[Ano02-28]  Anonymous.  Curriculum

vitae. Computers and Graphics, 26(5):656, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/28/abstract.
html.

Anonymous:2002:EM

[Ano02-29] Anonymous. ED-MEDIA 2002. Computers and Graphics, 26(1):201–202, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/49/abstract.
html.

Anonymous:2002:EBa

[Ano02-30]  Anonymous.  Editorial

Board. Computers and Graphics, 26(4):CO2, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/27/abstract.
                                        html.

Anonymous:2002:EBb

[Ano02-31]  Anonymous.  Editorial

Board. Computers and Graphics, 26(5):CO2, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/56/25/abstract.
                                        html.

Anonymous:2002:EBc

[Ano02-32]  Anonymous.  Editorial

Board. Computers and Graphics, 26(6):CO2, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:EEW

[Ano02-33] Anonymous. Eighth Eurographics Workshop on Virtual Environments. Computers and Graphics, 26(1):200– 201, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

13/20/68/27/46/abstract.
html.

Anonymous:2002:Ea

[Ano02-34] Anonymous. EUROGRAPH-ICS 2002. Computers and Graphics, 26(1):204, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/54/abstract.
html.

Anonymous:2002:EG

[Ano02-35]  Anonymous.  Eurograph-

ics 2002. Computers and Graphics, 26(2):387, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/48/abstract.
html.

Anonymous:2002:GMPa

[Ano02-36]  Anonymous.  Geometric

Modeling and Processing 2002 (GMP 2002). Computers and Graphics, 26(1): 202, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/50/abstract.
html.

Anonymous:2002:GMPb

[Ano02-37]  Anonymous.  Geometric

Modeling and Processing 2002 (GMP 2002) — Japan, July 10–12, 2002. Computers and Graphics, 26(2):386, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/41/46/abstract.
                                        html.

Anonymous:2002:III

[Ano02-38] Anonymous. ICME 2002 IEEE International Conference on Multimedia and Expo. Computers and Graphics, 26(2):387, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/13/20/68/41/
                                        50/abstract.html.

Anonymous:2002:Ia

[Ano02-39] Anonymous. ICPNM. Computers and Graphics, 26 (5):825, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/56/45/abstract.
                                        html.

Anonymous:2002:Ib

[Ano02-40] Anonymous. ICPNM. Computers and Graphics, 26(6): 981, December 2002. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2002:IIC

[Ano02-41] Anonymous. IEEE International Conference on Multimedia and Expo. Computers and Graphics, 26(1): 203, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/52/abstract.
html.

Anonymous:2002:IVa

[Ano02-42] Anonymous. IEEE Visualization 2002. Computers and Graphics, 26(1):204, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/53/abstract.
html.

Anonymous:2002:IVb

[Ano02-43] Anonymous. IEEE Visualization 2002. Computers and Graphics, 26(2):388, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/52/abstract.
html.

Anonymous:2002:ICPa

[Ano02-44] Anonymous. International certificate program for New Media (ICPNM). Computers and Graphics, 26(1):199, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/45/abstract.
html.

Anonymous:2002:ICPb

[Ano02-45] Anonymous. International certificate program for new media (ICPNM). Computers and Graphics, 26(2):385, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/44/abstract.
html.

Anonymous:2002:PG

[Ano02-46] Anonymous. Pacific Graphics 2002. Computers and Graphics, 26(2):387–388, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/49/abstract.
html.

Anonymous:2002:PFIa

[Ano02-47] Anonymous. Past/future issues. Computers and Graphics, 26(1):205, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
gej-ng/10/13/20/68/27/
55/abstract.html.

Anonymous:2002:PFIb

[Ano02-48] Anonymous. Past/future issues. Computers and Graphics, 26(2):389, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
gej-ng/10/13/20/68/41/
53/abstract.html.

Anonymous:2002:PFIc

[Ano02-49] Anonymous. Past/future issues. Computers and Graphics, 26(4):651, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/54/abstract.
html.

Anonymous:2002:SAS

[Ano02-50] Anonymous. Seventh ACM Symposium on Solid Modeling and Applications. Computers and Graphics, 26(1): 201, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/48/abstract.
html.

Anonymous:2002:Sa

[Ano02-51] Anonymous. SIGGRAPH 2002. Computers and Graphics, 26(1):203, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/13/20/68/27/
                                        51/abstract.html.

Anonymous:2002:Sb

[Ano02-52] Anonymous. SIGGRAPH 2002. Computers and Graphics, 26(2):386, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/13/20/68/41/
                                        47/abstract.html.

Anonymous:2002:TIC

[Ano02-53] Anonymous. The third international conference on virtual reality and its application in industry (VRAI’2002). Computers and Graphics, 26(1):200, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/27/47/abstract.
                                        html.

Anonymous:2002:W

[Ano02-54] Anonymous. WEDELMU-SIC 2002. Computers and Graphics, 26(5):826, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/56/47/abstract.
                                        html.

Anonymous:2002:WIC

[Ano02-55] Anonymous. WEDELMU-SIC 2002 — 2nd International Conference on Web Delivering of Music December 9–11, 2002, Darmstadt, Germany. Computers and Graphics, 26(3):530, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/46/abstract.
html;   http://www.wedelmusic.
org/wedelmusic2002.

Anonymous:2003:ICC

[Ano03a] Anonymous. 2003 International Conference on Cyberworlds: Marina Mandarin Hotel, Singapore 3-5 December 2003, orzanized by School of Computer Engineering, Nanyang Technological University Conference. Computers and Graphics, 27 (4):663–664, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:ICM

[Ano03b] Anonymous. 2003 International conference on multimedia and expo (ICME). Computers and Graphics, 27(1):159, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IWC

[Ano03c] Anonymous. 4th international working conference on intelligent virtual agents.

Computers and Graphics, 27 (3):458, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IIC

[Ano03d] Anonymous. 6th IASTED international conference on computers, graphics, and imaging — CGIM 2003. Computers and Graphics, 27 (3):457, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:AM

[Ano03e] Anonymous. ACM Multimedia 2003. Computers and Graphics, 27(3): 461, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:ASS

[Ano03f] Anonymous. ACM SIG-GRAPH 2003 symposium on interactive 3D graphics. Computers and Graphics, 27 (1):157–158, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:A

[Ano03g]  Anonymous.  AIM2003.

Computers and Graphics, 27 (3):456, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:ACE

[Ano03h] Anonymous. Annual Conference of the European Association for Computer Graphics. Computers and Graphics, 27(2):323, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CIC

[Ano03i] Anonymous. CAD 2004:

international CAD conference and exhibition May 24– 28. Computers and Graphics, 27(4):664, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CPV

[Ano03j] Anonymous. Call for papers for volume 27 number 3. Computers and Graphics, 27(3):465, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CGB

[Ano03k]  Anonymous.  Comput-

ers & graphics best paper award 2002. Computers and Graphics, 27(6): 829–834, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CVa

[Ano03l] Anonymous. Curriculum vitae. Computers and Graphics, 27(4):468, August 2003.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CVb

[Ano03m] Anonymous. Curriculum Vitae. Computers and Graphics, 27(6):837, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:CVc

[Ano03n] Anonymous. Curriculum Vitae. Computers and Graphics, 27(6):838, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:C

[Ano03o] Anonymous. Cyberworlds 2003. Computers and Graphics, 27(1):160–161, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:DGI

[Ano03p] Anonymous. Digital games industries: development, impact and direction. Computers and Graphics, 27(3): 459, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EM

[Ano03q] Anonymous. ED-MEDIA 2003. Computers and Graphics, 27(2):322, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBa

[Ano03r]  Anonymous.  Editorial

Board. Computers and Graphics, 27(4):CO2, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBb

[Ano03s]  Anonymous.  Editorial

Board. Computers and Graphics, 27(5):CO2, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBc

[Ano03t]  Anonymous.  Editorial

Board. Computers and Graphics, 27(6):CO2, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBPc

[Ano03u]  Anonymous.  Editorial

Board/publication info. Computers and Graphics, 27(3): CO2, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBPa

[Ano03v]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 27(1):CO2, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:EBPb

[Ano03w]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 27(2):CO2, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:E

[Ano03x]  Anonymous.  Eurograph-

ics 2004. Computers and Graphics, 27(6):987, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:ESR

[Ano03y]  Anonymous.  Eurograph-

ics Symposium on Rendering 2003. Computers and Graphics, 27(2):322, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:FWC

[Ano03z] Anonymous. First WSCG 2004 call for papers. Computers and Graphics, 27 (5):827–828, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:GH

[Ano03-27] Anonymous. Graphics hardware 2003. Computers and Graphics, 27(3):456, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:GI

[Ano03-28] Anonymous. Graphics Interface 2003. Computers and Graphics, 27(1):158, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:I

[Ano03-29] Anonymous. ICPNM. Computers and Graphics, 27(1): 156, February 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IA

[Ano03-30] Anonymous. ICPNM announcement. Computers and Graphics, 27(6):983, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IVb

[Ano03-31] Anonymous. IEE VR 2004.

Computers and Graphics, 27(6):986, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IVa

[Ano03-32] Anonymous. IEEE Visualization 2003. Computers and Graphics, 27(2):323– 324, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:ICP

[Ano03-33] Anonymous. International certificate program for new media (ICPNM). Computers and Graphics, 27(2): 321, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IIVb

[Ano03-34] Anonymous. Internet imaging V. Computers and Graphics, 27(5):825, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IIVc

[Ano03-35] Anonymous. Internet Imaging V. Computers and Graphics, 27(6):984, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IMA

[Ano03-36] Anonymous. ISMAR 03 (mixed and augmented reality). Computers and Graphics, 27(4):660, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:IIVa

[Ano03-37]  Anonymous.  IVA 2003

(intelligent virtual agents). Computers and Graphics, 27 (4):658, August 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:LR

[Ano03-38] Anonymous. List of reviewers in 2000/2001. Computers and Graphics, 27 (1):153–155, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:M

[Ano03-39] Anonymous. Mirage 2003.

Computers and Graphics, 27(1):157, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PG

[Ano03-40] Anonymous. Pacific graphics 2003. Computers and Graphics, 27(4):661, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFIe

[Ano03-41] Anonymous. Past /future issues. Computers and Graphics, 27(5):826, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFIa

[Ano03-42] Anonymous. Past/future issues. Computers and Graphics, 27(1):162, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFIb

[Ano03-43] Anonymous. Past/future issues. Computers and Graphics, 27(2):325, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFIc

[Ano03-44] Anonymous. Past/future issues. Computers and Graphics, 27(3):463, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFId

[Ano03-45] Anonymous. Past/future issues. Computers and Graphics, 27(4):664, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PFIf

[Ano03-46] Anonymous. Past/future issues. Computers and Graphics, 27(6):988, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:P

[Ano03-47] Anonymous. PVG 2003.

Computers and Graphics, 27 (3):460, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:PPL

[Ano03-48]  Anonymous.  PVG 2003

(parallel and large-data visualization and graphics). Computers and Graphics, 27 (4):662, August 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:SGD

[Ano03-49] Anonymous. SIAM (geometric design computing). Computers and Graphics, 27(5):823, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:S

[Ano03-50] Anonymous. SIGGRAPH 2003. Computers and Graphics, 27(1):160, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:SSMb

[Ano03-51] Anonymous. SM’04 (Solid Modeling and Applications). Computers and Graphics, 27(6):987, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:SSMa

[Ano03-52] Anonymous. SMI’04 (Shape Modeling International). Computers and Graphics, 27(6):986, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:SM

[Ano03-53] Anonymous. Solid modelling 2003. Computers and Graphics, 27(1):158– 159, February 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:SGP

[Ano03-54] Anonymous. Symposium on geometry processing. Computers and Graphics, 27(1): 159, February 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VVDa

[Ano03-55] Anonymous. VDA2004 (visualization and data analysis). Computers and Graphics, 27(5):825, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VVDb

[Ano03-56] Anonymous. VDA2004 (Visualization and Data Analysis). Computers and Graphics, 27(6):984, December 2003. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VC

[Ano03-57] Anonymous. Virtual concept 2003. Computers and Graphics, 27(3):461, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VS

[Ano03-58] Anonymous. Virtual storytelling 03. Computers and Graphics, 27(5):824, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VK

[Ano03-59] Anonymous. Visual knowledges. Computers and Graphics, 27(3):459, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VVM

[Ano03-60] Anonymous. VMV 2003 (vison, modeling and visualization). Computers and Graphics, 27(4):663, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:VCA

[Ano03-61] Anonymous. Volume contents/author index for vol 27. Computers and Graphics, 27(6):??, December 2003. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:WS

[Ano03-62] Anonymous. Web 3D 2003 Symposium. Computers and Graphics, 27(1):157, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2003:W

[Ano03-63] Anonymous. WSCG 2004. Computers and Graphics, 27 (6):985–986, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:AMM

[Ano04a] Anonymous. The 13th annual medicine meets virtual reality conference. Computers and Graphics, 28(6): 1007–1008, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:ICPc

[Ano04b] Anonymous. 2nd Iberian Conference on Pattern Recognition and Image Analysis. Computers and Graphics, 28 (6):1008, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:ICW

[Ano04c]  Anonymous.  4th In-

ternational Conference on Web Delivering of Music WEDELMUSIC 2004. Computers and Graphics, 28 (4):618–619, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:AAC

[Ano04d] Anonymous. Acivs 2004 (advanced concepts for intelligent vision systems). Computers and Graphics, 28(3): 460, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:AMa

[Ano04e] Anonymous. ACM multimedia 2004. Computers and Graphics, 28(3): 461, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:AMb

[Ano04f] Anonymous. ACM Multimedia 2004. Computers and Graphics, 28(4): 620, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CPS

[Ano04g] Anonymous. Call for papers: special issue on pervasive computing and ambient intelligence. Computers and Graphics, 28(1):

135, February 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CICb

[Ano04h] Anonymous. CATE 2004 (IASTED conference on computers and advanced technology in education). Computers and Graphics, 28 (3):459, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CWC

[Ano04i] Anonymous. CGE04 workshop on computer graphics education. Computers and Graphics, 28(1):132, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CCG

[Ano04j]  Anonymous.  CGI 2004

(computer graphics international). Computers and Graphics, 28(1):133, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CICc

[Ano04k] Anonymous. CGIM 2004 (IASTED conf. on Comp Graphics; Imaging). Computers and Graphics, 28(3): 459, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CICa

[Ano04l] Anonymous. CGIM 2004 (IASTED conference on computer graphics and imaging). Computers and Graphics, 28(1):133, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CGB

[Ano04m]  Anonymous.  Computer

Graphics best paper award (2003). Computers and Graphics, 28(6):791–798, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CVD

[Ano04n] Anonymous. Conference on visualization and data analysis 2005 (EI10). Computers and Graphics, 28(5): 788, October 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Cb

[Ano04o] Anonymous. CONVR 2004.

Computers and Graphics, 28 (4):619, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CCA

[Ano04p] Anonymous. CONVR 2004 (construction applications of virtual reality). Computers and Graphics, 28(3):460– 461, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Cd

[Ano04q]  Anonymous.  CW 2004.

Computers and Graphics, 28 (4):621, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:CC

[Ano04r] Anonymous. CW 2004 (cyberworlds). Computers and Graphics, 28(3):462, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Ca

[Ano04s] Anonymous. Cyberworlds 2004. Computers and Graphics, 28(2):306, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EM

[Ano04t]  Anonymous.  ED-media

2004. Computers and Graphics, 28(1):133, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBa

[Ano04u]  Anonymous.  Editorial

Board. Computers and Graphics, 28(1):CO2, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBb

[Ano04v]  Anonymous.  Editorial

Board. Computers and Graphics, 28(3):CO2, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBc

[Ano04w]  Anonymous.  Editorial

Board. Computers and Graphics, 28(5):CO2, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBPb

[Ano04x]  Anonymous.  Editorial

Board/pub info. Computers and Graphics, 28(4): CO2, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBPc

[Ano04y]  Anonymous.  Editorial

Board/publication info. Computers and Graphics, 28 (6):CO2, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EBPa

[Ano04z]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 28(2):CO2, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:EPP

[Ano04-27] Anonymous. EG PGV 2004 (parallel graphics and visualization). Computers and Graphics, 28(2):303, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Ea

[Ano04-28] Anonymous. Eurographics 2004. Computers and Graphics, 28(1):134, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:E

[Ano04-29] Anonymous. EUROGRAPH-ICS 2004. Computers and Graphics, 28(3):460, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:GMP

[Ano04-30]  Anonymous.  Geometric

modeling and processing 2004. Computers and Graphics, 28(1):130, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:GH

[Ano04-31] Anonymous. Graphics hardware 2004. Computers and Graphics, 28(3):459– 460, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IIC

[Ano04-32] Anonymous. ICPNM — International Certificate Program for New Media. Computers and Graphics, 28(3): 457–458, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IR

[Ano04-33] Anonymous. ICPNM — repeat. Computers and Graphics, 28(6):1005–1006, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IVa

[Ano04-34] Anonymous. IEEE visualization 2004. Computers and Graphics, 28(3): 462, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IVb

[Ano04-35] Anonymous. IEEE Visualization 2004. Computers and Graphics, 28(4):

620, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:ICPa

[Ano04-36] Anonymous. International certificate program for new media. Computers and Graphics, 28(1):129–130, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:ICPb

[Ano04-37] Anonymous. International Certificate Program for New Media. Computers and Graphics, 28(4):617–618, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IWP

[Ano04-38]  Anonymous.  Interna-

tional workshop on pervasive eLearning (PerEL’05). Computers and Graphics, 28 (6):1008, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IIA

[Ano04-39] Anonymous. ISMAR 04: IEEE and ACM international symposium on mixed augmented reality. Computers and Graphics, 28 (5):786–787, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:III

[Ano04-40] Anonymous. IV’04 (IEEE intelligent vehicles symposium). Computers and Graphics, 28(2):304, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:I

[Ano04-41] Anonymous. IWVR 2004.

Computers and Graphics, 28 (4):619, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IVRa

[Ano04-42] Anonymous. IWVR 2004 (virtual rehabilitation). Computers and Graphics, 28(2): 306, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:IVRb

[Ano04-43] Anonymous. IWVR 2004 (Virtual Rehabilitation). Computers and Graphics, 28 (3):461, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:LR

[Ano04-44] Anonymous. List of reviewers in 2002/2003. Computers and Graphics, 28 (1):125–127, February 2004.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PFIa

[Ano04-45] Anonymous. Past/future issues. Computers and Graphics, 28(1):136, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PFIb

[Ano04-46] Anonymous. Past/future issues. Computers and Graphics, 28(3):463, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PFIc

[Ano04-47] Anonymous. Past/future issues. Computers and Graphics, 28(4):622, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PFId

[Ano04-48] Anonymous. Past/future issues. Computers and Graphics, 28(5):789, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PFIe

[Ano04-49] Anonymous. Past/future issues. Computers and Graphics, 28(6):1009, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:P

[Ano04-50] Anonymous. Pervasive 2004.

Computers and Graphics, 28(1):131, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:PBG

[Ano04-51] Anonymous. Point-based graphics symposium. Computers and Graphics, 28 (1):131–132, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Sa

[Ano04-52] Anonymous. SIGGRAPH 2004. Computers and Graphics, 28(2):305, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:Sb

[Ano04-53] Anonymous. SIGGRAPH 2004. Computers and Graphics, 28(3):458, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:SSMb

[Ano04-54] Anonymous. SM 2004 (solid modeling and applications). Computers and Graphics, 28(1):132, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:SSMa

[Ano04-55]  Anonymous.  SMI 2004

(shaping modeling international). Computers and Graphics, 28(1):132, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:SIC

[Ano04-56] Anonymous. Special issue on class A surfaces. Computers and Graphics, 28(3): 464, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:SMR

[Ano04-57] Anonymous. SPIE MMCN 2005 — repeat. Computers and Graphics, 28(6):1007, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:SMT

[Ano04-58] Anonymous. SPIE MMCN 2005: Twelth annual multimedia computing and networking (MMCN ’05) in cooperation with: ACM SIG Multimedia. Computers and Graphics, 28(5): 788, October 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:TCT

[Ano04-59] Anonymous. TP CG ’04 (theory and practice of computer graphics). Computers and Graphics, 28(2): 303, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:VVD

[Ano04-60] Anonymous. VDA 2005 (Visualisation and Data Analysis) — repeat. Computers and Graphics, 28(6):1007, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:VMV

[Ano04-61] Anonymous. Vision, modeling and visualization 2004. Computers and Graphics, 28(5):787, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:VEI

[Ano04-62] Anonymous. VisSym 2004 (EG-IEEE-TCVG symposium on visualization). Computers and Graphics, 28(1): 131, February 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:VCA

[Ano04-63] Anonymous. Volume contents and author index, vol 28. Computers and Graphics, 28(6):??, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2004:V

[Ano04-64] Anonymous. VRCAI 2004. Computers and Graphics, 28 (2):304, April 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:ICC

[Ano05a] Anonymous. 2005 international conference on cyberworlds. Computers and Graphics, 29(2):304, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:CGI

[Ano05b] Anonymous. The 23rd computer graphics international conference (CGI’05). Computers and Graphics, 29(1): 173, February 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:ICP

[Ano05c] Anonymous. 2nd Iberian Conference on Pattern Recognition and Image Analysis. Computers and Graphics, 29(1):171, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:FIC

[Ano05d] Anonymous. 3DIM 2005:

the fifth international conference on 3-D digital imaging and modeling. Computers and Graphics, 29(1): 172, February 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AA

[Ano05e]  Anonymous.  Announce-

ment: Acivs 2005. Computers and Graphics, 29(3): 472, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AG

[Ano05f]  Anonymous.  Announce-

ment: Graphite 2005. Computers and Graphics, 29(3): 474, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AIa

[Ano05g]  Anonymous.  Announce-

ment: ICIMA’05. Computers and Graphics, 29(3): 471, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AIV

[Ano05h]  Anonymous.  Announce-

ment: IEEE visualization 2005. Computers and Graphics, 29(3):472, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AIb

[Ano05i]  Anonymous.  Announce-

ment: Interact 2005. Computers and Graphics, 29(3): 472, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:ASG

[Ano05j]  Anonymous.  Announce-

ment: Smart Graphics. Computers and Graphics, 29 (3):471, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:ASO

[Ano05k]  Anonymous.  Announce-

ment: SVG open 2005. Computers and Graphics, 29(3): 470, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:AV

[Ano05l]  Anonymous.  Announce-

ment: VRIPHYS’05. Computers and Graphics, 29(3): 473, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:CPS

[Ano05m] Anonymous. Call for papers: special issue of the Elsevier journal “Computers and Graphics”, natural phenomena simulation. Computers and Graphics, 29(2):306–307, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Ca

[Ano05n] Anonymous. CGEMS. Computers and Graphics, 29(2): 300, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:CCG

[Ano05o] Anonymous. CGEMS — Computer Graphics Educational Materials Source. Computers and Graphics, 29(3):468–469, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Cb

[Ano05p] Anonymous. CGI 05. Computers and Graphics, 29(2): 301, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:CGE

[Ano05q]  Anonymous.  Computer

graphics educational materials source (CGEMS). Computers and Graphics, 29 (1):169–170, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Cc

[Ano05r]  Anonymous.  CYBER-

WORLDS 2005. Computers and Graphics, 29(3): 473, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:EM

[Ano05s] Anonymous. Ed-media 2005. Computers and Graphics, 29 (2):301, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:EBa

[Ano05t]  Anonymous.  Editorial

Board. Computers and Graphics, 29(3):CO2, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:EBPa

[Ano05u]  Anonymous.  Editorial

Board/publication info. Computers and Graphics, 29(1): CO2, February 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:EBPb

[Ano05v]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 29(2):CO2, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Ea

[Ano05w] Anonymous. Eurographics 2005. Computers and Graphics, 29(2):303, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Eb

[Ano05x] Anonymous. EUROGRAPH-ICS 2005. Computers and Graphics, 29(3):471, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:GHC

[Ano05y]  Anonymous.  GRAPH-

ICS HARDWARE 2005 cosponsored by ACM SIG-GRAPH and Eurographics (pending). Computers and Graphics, 29(2): 303, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:IIC

[Ano05z] Anonymous. ICPNM — International Certificate Program for New Media. Computers and Graphics, 29(3): 467–468, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:IA

[Ano05-27] Anonymous. ICPNM announcement. Computers and Graphics, 29(2):299– 300, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:ISV

[Ano05-28] Anonymous. international symposium on visual computing ISVC05. Computers and Graphics, 29(1):173, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:IWVa

[Ano05-29] Anonymous. International workshop on volume graphics. Computers and Graphics, 29(1):172, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:IWVb

[Ano05-30]  Anonymous.  Interna-

tional Workshop on Volume Graphics. Computers and Graphics, 29(2): 301, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Ib

[Ano05-31] Anonymous. ISVC05. Computers and Graphics, 29(2): 304, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Ic

[Ano05-32] Anonymous. ISVC05. Computers and Graphics, 29(3): 474, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:LR

[Ano05-33] Anonymous. List of reviewers in 2003/2004. Computers and Graphics, 29 (1):165–167, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:PFIa

[Ano05-34] Anonymous. Past/future issues. Computers and Graphics, 29(1):175, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:PFIb

[Ano05-35] Anonymous. Past/future issues. Computers and Graphics, 29(2):305, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:PFIc

[Ano05-36] Anonymous. Past/future issues. Computers and Graphics, 29(3):475–476, June 2005. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:P

[Ano05-37] Anonymous. Pervasive 2005. Computers and Graphics, 29(1):171, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Sa

[Ano05-38] Anonymous. Siggraph 2005. Computers and Graphics, 29 (2):303, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:Sb

[Ano05-39] Anonymous. SIGGRAPH 2005. Computers and Graphics, 29(3):469–470, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:SIC

[Ano05-40]  Anonymous.  SMI 2005:

international conference on shape modelling. Computers and Graphics, 29(1):172, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2005:WIC

[Ano05-41] Anonymous. WEB3D 2005: the 10th International Conference on 3D Web Technology. Computers and Graphics, 29(1):170–171, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2006:ICC

[Ano06a] Anonymous. 2006 International Conference on CY-BERWORLDS. Computers and Graphics, 30(5): 881, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001907.

Anonymous:2006:ICM

[Ano06b] Anonymous. 2006 International Conference on Multimedia & Expo (ICME). Computers and Graphics, 30 (2):315, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000136.

Anonymous:2006:ASI

[Ano06c] Anonymous. ACM SIG-GRAPH International Conference on Virtual Reality Continuum and Its Applications VRCIA 2006. Computers and Graphics, 30 (2):313, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000094.

Anonymous:2006:Aa

[Ano06d]  Anonymous.  Announce-

ments. Computers and Graphics, 30(3):474–481, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000719.

Anonymous:2006:Ab

[Ano06e]  Anonymous.  Announce-

ments 30/4. Computers and Graphics, 30(4):669– 675, August 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000720.

Anonymous:2006:CP

[Ano06f] Anonymous. Call for papers. Computers and Graphics, 30(1):149, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002207.

Anonymous:2006:CRa

[Ano06g] Anonymous. CGEMS — repeat from 29/6. Computers and Graphics, 30 (1):151–152, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002578.

Anonymous:2006:CRb

[Ano06h] Anonymous. CHI 2006 — repeat from 29/6. Computers and Graphics, 30(1): 153, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002591.

Anonymous:2006:CGI

[Ano06i]  Anonymous.  Computer

graphic international 2006. Computers and Graphics, 30 (2):314, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000100.

Anonymous:2006:CGEa

[Ano06j]  Anonymous.  Computer

graphics educational materials source (CGEMS). Computers and Graphics, 30(2): 312–313, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000082.

Anonymous:2006:CGEb

[Ano06k]  Anonymous.  Computer

Graphics Educational Materials Source (CGEMS). Computers and Graphics, 30(5):880, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001889.

Anonymous:2006:CGEc

[Ano06l]  Anonymous.  Computer

graphics educational materials source (CGEMS). Computers and Graphics, 30 (6):1042, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002160.

Anonymous:2006:CGB

[Ano06m] Anonymous. Computers & Graphics Best Paper Award. Computers and Graphics, 30 (6):885–890, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001658.

Anonymous:2006:CVD

[Ano06n]  Anonymous.  Conference

on Visualization and Data Analysis 2007. Computers and Graphics, 30(6): 1043, December 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002196.

Anonymous:2006:C

[Ano06o] Anonymous. CSCW 2006. Computers and Graphics, 30 (2):319, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000069.

Anonymous:2006:CRc

[Ano06p] Anonymous. CSCW 2006 — repeat from 29/6. Computers and Graphics, 30(1): 155, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930500261X.

Anonymous:2006:D

[Ano06q] Anonymous. DSV-IS 2006.

Computers and Graphics, 30 (2):316, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600015X.

Anonymous:2006:EM

[Ano06r]  Anonymous.  ED ME-

DIA 2066. Computers and Graphics, 30(2):314– 315, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000124.

Anonymous:2006:EB

[Ano06s]  Anonymous.  Editorial

Board. Computers and Graphics, 30(1):??, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009784930500244X.

Anonymous:2006:EBPa

[Ano06t]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 30(2):??, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000392.

Anonymous:2006:EBPb

[Ano06u]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 30(3):??, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000914.

Anonymous:2006:EBPc

[Ano06v]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 30(4):??, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930600104X.

Anonymous:2006:EBPd

[Ano06w]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 30(5):??, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001671.

Anonymous:2006:EBPe

[Ano06x]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 30(6):??, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001993.

Anonymous:2006:ER

[Ano06y] Anonymous. Edutainment 2006 — repeat from 29/6. Computers and Graphics, 30 (1):152–153, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930500258X.

Anonymous:2006:E

[Ano06z]  Anonymous.  Eurograph-

ics 2006. Computers and Graphics, 30(1):154, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002219.

Anonymous:2006:FAM

[Ano06-27] Anonymous. Fourteenth Annual Multimedia Computing and Networking Conference (MMCN’07). Computers and Graphics, 30(5): 882, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001932.

Anonymous:2006:IR

[Ano06-28] Anonymous. ICPNM — repeat from 29/6. Computers and Graphics, 30 (1):150–151, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002566.

Anonymous:2006:IVRa

[Ano06-29] Anonymous. IEEE Virtual Reality Conference 2007. Computers and Graphics, 30(5):883, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001956.

Anonymous:2006:IVRb

[Ano06-30] Anonymous. IEEE Virtual Reality Conference 2007. Computers and Graphics, 30 (6):1044, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002202.

Anonymous:2006:IVRc

[Ano06-31] Anonymous. IEEE Virtual Reality Conference 2007. Computers and Graphics, 30 (6):1044, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002214.

Anonymous:2006:ICP

[Ano06-32]  Anonymous.  Interna-

tional certificate program for new media. Computers and Graphics, 30(2): 311–312, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000070.

Anonymous:2006:ICS

[Ano06-33] Anonymous. International conference series on disability, virtual reality and associated technologies. Computers and Graphics, 30 (2):319, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000203.

Anonymous:2006:IMM

[Ano06-34] Anonymous. The international multi media modeling conference. Computers and Graphics, 30(6): 1043, December 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002172.

Anonymous:2006:ICA

[Ano06-35]  Anonymous.  Iv confer-

ence on articulated motion and deformable objects (AMDO-e 2006). Computers and Graphics, 30(2): 315–316, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000148.

Anonymous:2006:LR

[Ano06-36] Anonymous. List of reviewers in 2004/2005. Computers and Graphics, 30 (1):145–148, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002190.

Anonymous:2006:MIM

[Ano06-37] Anonymous. MMM 2007: The International Multi Media Modeling Conference. Computers and Graphics, 30(5):882, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001920.

Anonymous:2006:MCN

[Ano06-38]  Anonymous.  Multimedia

Computing and Networking Conference (MMCN’07). Computers and Graphics, 30 (6):1043, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002184.

Anonymous:2006:PIF

[Ano06-39] Anonymous. Past issues/ future issues. Computers and Graphics, 30(5): 884, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001439.

Anonymous:2006:PFIa

[Ano06-40] Anonymous. Past/future issues. Computers and Graphics, 30(1):156, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849305002220.

Anonymous:2006:PFIb

[Ano06-41] Anonymous. Past/future issues. Computers and Graphics, 30(2):320, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000057.

Anonymous:2006:PFIc

[Ano06-42] Anonymous. Past/future issues. Computers and Graphics, 30(4):676–677, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001014.

Anonymous:2006:PFId

[Ano06-43] Anonymous. Past/future issues. Computers and Graphics, 30(6):1045, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001646.

Anonymous:2006:PR

[Ano06-44] Anonymous. PERVASIVE 2006 — repeat from 29/6. Computers and Graphics, 30(1):153, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002608.

Anonymous:2006:PN

[Ano06-45] Anonymous. PUBLISHER’S NOTE. Computers and Graphics, 30(2):157, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849306000239.

Anonymous:2006:S

[Ano06-46] Anonymous. SIGGRAPH 2006. Computers and Graphics, 30(2):317, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000173.

Anonymous:2006:SCA

[Ano06-47] Anonymous. Symposium on computer animation 2006. Computers and Graphics, 30 (2):317, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000185.

Anonymous:2006:SGP

[Ano06-48] Anonymous. Symposium on geometry processing. Computers and Graphics, 30 (2):314, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000112.

Anonymous:2006:SPB

[Ano06-49] Anonymous. Symposium on point-based graphic 2006. Computers and Graphics, 30 (2):316, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000161.

Anonymous:2006:TTI

[Ano06-50] Anonymous. TIDSE 2006: Third International Conference on Technologies for Interactive Digital Storytelling and Entertainment. Computers and Graphics, 30(5): 881, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001919.

Anonymous:2006:VCV

[Ano06-51] Anonymous. VDA 2007: Conference on Visualization and Data Analysis 2007. Computers and Graphics, 30(5):882, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001944.

Anonymous:2006:VMV

[Ano06-52] Anonymous. Vision, Modeling, and Visualization 2006 (VMV 2006). Computers and Graphics, 30(5):881– 883, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001890.

Anonymous:2007:ASSa

[Ano07a] Anonymous. 12th ACM Symposium on Solid and Physical Modeling. Computers and Graphics, 31(1): 152, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002342.

Anonymous:2007:ASSb

[Ano07b] Anonymous. 12th ACM Symposium on Solid and Physical Modeling. Computers and Graphics, 31 (2):309, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000246.

Anonymous:2007:ICCb

[Ano07c]  Anonymous.  2007 In-

ternational Conference On Cyberworlds. Computers and Graphics, 31(3): 535, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001203.

Anonymous:2007:IIRb

[Ano07d] Anonymous. 34th IARI-GAI International Research Conference: Advances in Printing and Media Technology. Computers and Graphics, 31(3):534, June 2007.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000751.

Anonymous:2007:IIRa

[Ano07e] Anonymous. 34th IARIGAI International Research Conference Advances in Printing and Media Technology (hosted by INP Grenoble, EFPG). Computers and Graphics, 31(2): 311, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000295.

Anonymous:2007:CTI

[Ano07f]  Anonymous.  3IA ‘2007

CONFERENCE: The Tenth International Conference On Computer Graphics And Artificial Intelligence. Computers and Graphics, 31 (2):308, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000210.

Anonymous:2007:ICCa

[Ano07g] Anonymous. 5th International Conference on Computer Vision Systems. Computers and Graphics, 31(1): 151, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002329.

Anonymous:2007:ACAa

[Ano07h] Anonymous. ACM — CIVR 2007: ACM International Conference on Image and Video Retrieval. Computers and Graphics, 31(2): 310, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000271.

Anonymous:2007:ACAb

[Ano07i] Anonymous. ACM-CIVR 2007: ACM International Conference on Image and Video Retrieval. Computers and Graphics, 31(3): 532–533, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000696.

Anonymous:2007:AMa

[Ano07j] Anonymous. ACM MUL-TIMEDIA 2007. Computers and Graphics, 31 (2):311, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000301.

Anonymous:2007:AMb

[Ano07k] Anonymous. ACM Multimedia 2007. Computers and Graphics, 31(3): 535, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000775.

Anonymous:2007:CGEa

[Ano07l]  Anonymous.  Computer

Graphics Educational Materials Source (CGEMS). Computers and Graphics, 31 (1):150–152, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002366.

Anonymous:2007:CGEb

[Ano07m]  Anonymous.  Computer

Graphics Educational Materials Source (CGEMS). Computers and Graphics, 31 (3):532, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000684.

Anonymous:2007:CGI

[Ano07n]  Anonymous.  Computer

Graphics International. Computers and Graphics, 31 (2):308, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000222.

Anonymous:2007:CHI

[Ano07o]  Anonymous.  Computer/

Human Interaction ‘07. Computers and Graphics, 31(1):152, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002330.

Anonymous:2007:EMW

[Ano07p] Anonymous. ED-MEDIA 2007: World Conference on Educational Multimedia, Hypermedia & Telecommunications. Computers and Graphics, 31(2): 310, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700026X.

Anonymous:2007:EBPa

[Ano07q]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(1):??, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000106.

Anonymous:2007:EBPb

[Ano07r]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(2):??, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000829.

Anonymous:2007:EBPc

[Ano07s]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(3):??, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001033.

Anonymous:2007:EBPd

[Ano07t]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(4):??, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001409.

Anonymous:2007:EBPe

[Ano07u]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(5):??, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001732.

Anonymous:2007:EBPf

[Ano07v]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 31(6):??, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001860.

Anonymous:2007:ELGb

[Ano07w]  Anonymous.  Edutain-

ment 2007: E-Learning and Games. Computers and Graphics, 31(2): 309, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000258.

Anonymous:2007:ELGa

[Ano07x] Anonymous. Edutainment 2007E-Learning and Games. Computers and Graphics, 31(1):152, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002354.

Anonymous:2007:ESP

[Ano07y]  Anonymous.  Eurograph-

ics Symposium on Parallel Graphics and Visualization (EGPGV’07). Computers and Graphics, 31 (2):308, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000209.

Anonymous:2007:E

[Ano07z]  Anonymous.  Eurograph-

ics’2007. Computers and Graphics, 31(3):534, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009784930700074X.

Anonymous:2007:GH

[Ano07-27]  Anonymous.  Graphics

Hardware 2007. Computers and Graphics, 31(3): 533, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000714.

Anonymous:2007:HIIa

[Ano07-28] Anonymous. HCI international 2007: 12th international conference on humancomputer interaction. Computers and Graphics, 31 (2):310, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000283.

Anonymous:2007:HIIb

[Ano07-29] Anonymous. HCI International 2007: 12TH International Conference on Human-Computer Interaction. Computers and Graphics, 31(3):533, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000702.

Anonymous:2007:IVR

[Ano07-30] Anonymous. IEEE Virtual Reality Conference 2007. Computers and Graphics, 31(1):151, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002317.

Anonymous:2007:ISV

[Ano07-31]  Anonymous.  Interna-

tional Symposium on Volume Graphics. Computers and Graphics, 31(3): 534, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000738.

Anonymous:2007:LR

[Ano07-32]  Anonymous.  LIST OF

REVIEWERS. Computers and Graphics, 31(1):147– 149, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002287.

Anonymous:2007:NIW

[Ano07-33]  Anonymous.  NOSSDAV

2007: 17th International workshop on Network and Operating Systems Support for Digital Audio & Video. Computers and Graphics, 31 (2):309, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000234.

Anonymous:2007:PFIa

[Ano07-34] Anonymous. Past/future issues. Computers and Graphics, 31(1):153, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002275.

Anonymous:2007:PFIb

[Ano07-35] Anonymous. Past/future issues. Computers and Graphics, 31(2):312, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000313.

Anonymous:2007:PFIc

[Ano07-36] Anonymous. Past/future issues. Computers and Graphics, 31(3):536, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000660.

Anonymous:2007:PPI

[Ano07-37] Anonymous. PIA07-Photogrammetric Image Analysis. Computers and Graphics, 31 (3):535, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000763.

Anonymous:2007:SPB

[Ano07-38] Anonymous. Symposium on Point-Based Graphics 2007. Computers and Graphics, 31(3):533–534, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000726.

Anonymous:2007:ISU

[Ano07-39] Anonymous. The 2nd IEEE Symposium on 3D User Interfaces 2007: (in conjunction with IEEE Virtual Reality 2007). Computers and Graphics, 31(1): 151, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002305.

Anonymous:2007:WNJ

[Ano07-40] Anonymous. Welcome notice for Joaquim Jorge. Computers and Graphics, 31 (1):1, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002421.

Anonymous:2008:ICC

[Ano08a]  Anonymous.  2008 In-

ternational Conference on Cyberworlds. Computers and Graphics, 32(2): 282, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000046.

Anonymous:2008:CGB

[Ano08b]  Anonymous.  Comput-

ers & Graphics Best Paper Awards. Computers and Graphics, 32(6):613– 614, December 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001428.

Anonymous:2008:EBPa

[Ano08c]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(1):??, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000101.

Anonymous:2008:EBPb

[Ano08d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(2):??, April 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000344.

Anonymous:2008:EBPc

[Ano08e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(3):??, June 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000605.

Anonymous:2008:EBPd

[Ano08f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(4):??, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000794.

Anonymous:2008:EBPe

[Ano08g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(5):??, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001027.

Anonymous:2008:EBPf

[Ano08h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 32(6):??, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001313.

Anonymous:2008:SPC

[Ano08i] Anonymous. SMI’09 proceedings on Computers and Graphics. Computers and Graphics, 32(5): 611, October 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000848.

Anonymous:2009:CP

[Ano09a] Anonymous. Call for papers. Computers and Graphics, 33(6):??, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001484.

Anonymous:2009:CPS

[Ano09b] Anonymous. Call for papers: SMI’10. Computers and Graphics, 33(4): ??, August 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001095.

Anonymous:2009:CGB

[Ano09c] Anonymous. Computers & Graphics Best Paper Award. Computers and Graphics, 33(6):659, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001472.

Anonymous:2009:EBPa

[Ano09d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(1):??, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900003X.

Anonymous:2009:EBPb

[Ano09e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(2):??, April 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000478.

Anonymous:2009:EBPc

[Ano09f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(3):??, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000703.

Anonymous:2009:EBPd

[Ano09g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(4):??, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000958.

Anonymous:2009:EBPe

[Ano09h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(5):??, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001198.

Anonymous:2009:EBPf

[Ano09i]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 33(6):??, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930900137X.

Anonymous:2009:TYR

[Ano09j] Anonymous. Thank you to reviewers. Computers and Graphics, 33(6):

??, December 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001460.

Anonymous:2010:CGB

[Ano10a]  Anonymous.  Computers

& Graphics Best Paper Award 2009. Computers and Graphics, 34(6): 639, December 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001640.

Anonymous:2010:EBPa

[Ano10b]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(1):??, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000038.

Anonymous:2010:EBPb

[Ano10c]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(2):??, April 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000270.

Anonymous:2010:EBPc

[Ano10d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(3):??, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000592.

Anonymous:2010:EBPd

[Ano10e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(4):??, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000968.

Anonymous:2010:EBPe

[Ano10f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(5):??, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001160.

Anonymous:2010:EBPf

[Ano10g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 34(6):??, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001524.

Anonymous:2010:P

[Ano10h] Anonymous. Preface. Computers and Graphics, 34(3): 183–186, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000518.

Anonymous:2010:PN

[Ano10i]  Anonymous.  Publisher’s

note. Computers and Graphics, 34(4):479–480, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000701.

Anonymous:2010:TYR

[Ano10j] Anonymous. Thank you to reviewers. Computers and Graphics, 34(6): ??, December 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001421.

Anonymous:2011:CGB

[Ano11a] Anonymous. Computers & Graphics Best Associate Editor Award 2010. Computers and Graphics, 35 (6):1072, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001592.

Anonymous:2011:CGD

[Ano11b] Anonymous. Computers & Graphics Distinguished Reviewer Award 2008–2010. Computers and Graphics, 35 (6):1073, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001610.

Anonymous:2011:CGM

[Ano11c] Anonymous. Computers & Graphics Most Downloaded Paper Award 2010. Computers and Graphics, 35 (6):1071, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001609.

Anonymous:2011:E

[Ano11d]  Anonymous.  Editorial.

Computers and Graphics, 35(1):??, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310002086.

Anonymous:2011:EBPa

[Ano11e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(1):??, February 2011.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000094.

Anonymous:2011:EBPb

[Ano11f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(2):??, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000264.

Anonymous:2011:EBPc

[Ano11g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(3):??, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000835.

Anonymous:2011:EBPd

[Ano11h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(4):??, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001233.

Anonymous:2011:EBPe

[Ano11i]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(5):??, October 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001397.

Anonymous:2011:EBPf

[Ano11j]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 35(6):??, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001543.

Anonymous:2011:ISA

[Ano11k]  Anonymous.  Index of

SMI 2011 authors. Computers and Graphics, 35 (3):??, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000999.

Anonymous:2011:P

[Ano11l] Anonymous. Preface. Computers and Graphics, 35(3): 441–443, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001051.

Anonymous:2011:SSV

[Ano11m] Anonymous. Special section on virtual reality in Brazil.

Computers and Graphics, 35 (2):??, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000227.

Anonymous:2011:TYR

[Ano11n] Anonymous. Thank you to reviewers. Computers and Graphics, 35(6):1074– 1077, December 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001373.

Anonymous:2011:TP

[Ano11o] Anonymous. Title page.

Computers and Graphics, 35 (3):??, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000847.

Anonymous:2012:BSH

[Ano12a] Anonymous. Beyond swinging: Hinged dissections that twist or fold. Computers and Graphics, 36(5): 307, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000921.

Anonymous:2012:DRL

[Ano12b] Anonymous. The design and realization of large-scale free-form sculptures. Computers and Graphics, 36(5): 308, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000933.

Anonymous:2012:EB

[Ano12c]  Anonymous.  Editorial

Board. Computers and Graphics, 36(7):??, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849312001501.

Anonymous:2012:EBPa

[Ano12d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(1):??, February 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001762.

Anonymous:2012:EBPb

[Ano12e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(2):??, April 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000106.

Anonymous:2012:EBPc

[Ano12f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(3):??, May 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000283.

Anonymous:2012:EBPd

[Ano12g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(4):??, June 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000775.

Anonymous:2012:EBPe

[Ano12h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(5):??, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000970.

Anonymous:2012:EBPf

[Ano12i]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(6):??, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001240.

Anonymous:2012:EBPg

[Ano12j]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 36(8):??, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001653.

Anonymous:2012:EN

[Ano12k] Anonymous. Editorial note. Computers and Graphics, 36(7):??, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001434.

Anonymous:2012:FSSa

[Ano12l]  Anonymous.  Foreword

to special section on augmented reality. Computers and Graphics, 36(7): ??, November 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001161.

Anonymous:2012:FSSc

[Ano12m] Anonymous. Foreword to special section on Graphics Interaction. Computers and Graphics, 36(8):

??, December 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001604.

Anonymous:2012:FSSd

[Ano12n] Anonymous. Foreword to special section on Virtual Environments and Applications. Computers and Graphics, 36(8):??, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001458.

Anonymous:2012:FSSb

[Ano12o] Anonymous. Foreword to the special section on computer graphics in China. Computers and Graphics, 36(7):??, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001409.

Anonymous:2012:FAF

[Ano12p] Anonymous. Freeform architecture and fabricationaware design. Computers and Graphics, 36(5): 306, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200091X.

Anonymous:2012:E

[Ano12q] Anonymous. From the Editor. Computers and Graphics, 36(6):??, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001227.

Anonymous:2012:J

[Ano12r] Anonymous. January 2012. Computers and Graphics, 36(1):??, February 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000027.

Anonymous:2012:SSJ

[Ano12s] Anonymous. Special section on the 2011 Joint Symposium on Computational Aesthetics (CAe), Non-Photorealistic Animation and Rendering (NPAR), and Sketch-Based Interfaces and Modeling (SBIM). Computers and Graphics, 36(6): ??, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001112.

Anonymous:2013:CGB

[Ano13a] Anonymous. Computers & Graphics Best Associate Editor Award 2012. Computers and Graphics, 37(1– 2):94, February/April 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000241.

Anonymous:2013:CGD

[Ano13b] Anonymous. Computers & Graphics Distinguished Reviewer Award 2012. Computers and Graphics, 37(1– 2):95, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000253.

Anonymous:2013:CGM

[Ano13c] Anonymous. Computers & Graphics Most Downloaded Paper Award 2011. Computers and Graphics, 37(1– 2):93, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300023X.

Anonymous:2013:EBPa

[Ano13d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(1–2):??, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000186.

Anonymous:2013:EBPb

[Ano13e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(3):??, May 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931300037X.

Anonymous:2013:EBPc

[Ano13f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(4):??, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000502.

Anonymous:2013:EBPd

[Ano13g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(5):??, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931300068X.

Anonymous:2013:EBPe

[Ano13h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(6):??, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001167.

Anonymous:2013:EBPf

[Ano13i]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 37(7):??, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849313001349.

Anonymous:2013:FHD

[Ano13j] Anonymous. Foreword — high dynamic range imaging. Computers and Graphics, 37(7):??, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300126X.

Anonymous:2013:FSM

[Ano13k]  Anonymous.  Foreword

to Shape Modeling International 2013. Computers and Graphics, 37(6): ??, October 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001003.

Anonymous:2013:FSSb

[Ano13l]  Anonymous.  Foreword

to special section on advances in procedural modeling. Computers and Graphics, 37(4):??, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000162.

Anonymous:2013:FSSc

[Ano13m] Anonymous. Foreword to the special section on expressive graphics. Computers and Graphics, 37(5): ??, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000307.

Anonymous:2013:FSSa

[Ano13n] Anonymous. Foreword to the special section on touching the 3rd dimension. Computers and Graphics, 37 (3):??, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000290.

Anonymous:2013:SIE

[Ano13o] Anonymous. Special issue on executable papers for 3D object retrieval. Computers and Graphics, 37(5): ??, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000587.

Anonymous:2013:TYR

[Ano13p] Anonymous. Thank you to Reviewers. Computers and Graphics, 37(1–2):96– 99, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000265.

Anonymous:2014:EBPa

[Ano14a]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 41(??):IFC, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000430.

Anonymous:2014:EBPb

[Ano14b]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 42(??):IFC, August 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849314000557.

Anonymous:2014:EBPc

[Ano14c]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 43(??):IFC, October 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000685.

Anonymous:2014:EBPd

[Ano14d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 44(??):ifc, November 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849314000855.

Anonymous:2014:EBPe

[Ano14e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 45(??):IFC, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001320.

Anonymous:2015:BCA

[Ano15a] Anonymous. Biographies of co-authors (alphabetic order). Computers and Graphics, 53 (part A)(??):92– 93, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931500151X.

Anonymous:2015:EBPa

[Ano15b]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 46(??):IFC, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849314001460.

Anonymous:2015:EBPb

[Ano15c]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 47(??):IFC, April 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000187.

Anonymous:2015:EBPc

[Ano15d]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 48(??):ifc, May 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000333.

Anonymous:2015:EBPd

[Ano15e]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 49(??):ifc, June 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500076X.

Anonymous:2015:EBPe

[Ano15f]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 50(??):ifc, August 2015.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000953.

Anonymous:2015:EBPf

[Ano15g]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 51(??):ifc, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001284.

Anonymous:2015:EBPg

[Ano15h]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 52(??):ifc, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849315001557.

Anonymous:2015:EBPh

[Ano15i]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 53 (part A)(??):ifc, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001739.

Anonymous:2015:EBPi

[Ano15j]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 53 (part B)(??):ifc, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001855.

Anonymous:2015:FSM

[Ano15k] Anonymous. Foreword to Shape Modeling International 2014. Computers and Graphics, 46(??):A1– A3, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001290.

Anonymous:2015:TRa

[Ano15l] Anonymous. Thanks to reviewers. Computers and Graphics, 46(??):358–360, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001514.

Anonymous:2015:TRb

[Ano15m]  Anonymous.  Thanks to

reviewers. Computers and Graphics, 50(??):71– 76, August 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000928.

Anonymous:2016:EBPa

[Ano16a]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 54(??):ifc, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849315002034.

Anonymous:2016:EBPb

[Ano16b]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 55(??):ifc, April 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300048.

Anonymous:2016:EBPc

[Ano16c]  Anonymous.  Editorial

Board/publication information. Computers and Graphics, 56(??):ifc, May 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300255.

Anonymous:2016:EBPd

[Ano16d]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 57(??):ifc, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300371.

Anonymous:2016:EBPe

[Ano16e]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 58(??):ifc, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300723.

Anonymous:2016:EBPf

[Ano16f]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 59(??):ifc, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300929.

Anonymous:2016:EBPg

[Ano16g]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 60(??):ifc, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849316301145.

Anonymous:2016:EBPh

[Ano16h]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 61(??):ifc, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301273.

Anonymous:2016:TR

[Ano16i]  Anonymous.  Thanks to

reviewers. Computers and Graphics, 56(??):I– IV, May 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300310.

Anonymous:2017:EBa

[Ano17a]  Anonymous.  Editorial

Board. Computers and Graphics, 63(??):ifc, April 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849317300237.

Anonymous:2017:EBb

[Ano17b]  Anonymous.  Editorial

Board. Computers and Graphics, 64(??):ifc, May 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009784931730033X.

Anonymous:2017:EBc

[Ano17c]  Anonymous.  Editorial

Board. Computers and Graphics, 65(??):ifc, June 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784931730050X.

Anonymous:2017:EBd

[Ano17d]  Anonymous.  Editorial

Board. Computers and Graphics, 66(??):ifc, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317300882.

Anonymous:2017:EBe

[Ano17e]  Anonymous.  Editorial

Board. Computers and Graphics, 67(??):ifc, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301401.

Anonymous:2017:EBf

[Ano17f]  Anonymous.  Editorial

Board. Computers and Graphics, 68(??):ifc, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317301620.

Anonymous:2017:EBg

[Ano17g]  Anonymous.  Editorial

Board. Computers and Graphics, 69(??):ifc, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301796.

Anonymous:2017:EBP

[Ano17h]  Anonymous.  Editorial

board/publication information. Computers and Graphics, 62(??):ifc, ???? 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730002X.

Anonymous:2017:TR

[Ano17i]  Anonymous.  Thanks to

reviewers. Computers and Graphics, 62(??):I– IV, ???? 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300079.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Computers and Graphics, 70(??):ii, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317302054.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Computers and Graphics, 70(??):328–331, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317302091.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Computers and Graphics, 71(??):ii, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0097849318300177.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Computers and Graphics, 72(??):ii, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0097849318300414.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

board. Computers and Graphics, 73(??):ii, June 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

https://www.sciencedirect.
com/science/article/pii/
S0097849318300736.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Computers and Graphics, 74(??):ii, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849318300967.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Computers and Graphics, 75(??):1–82, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301262.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Computers and Graphics, 76(??):ii, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849318301638.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Computers and Graphics, 77(??):ii, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301870.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Computers and Graphics, 78(??):ii, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849319300044.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. Computers and Graphics, 79(??):ii, April 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849319300196.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

Board. Computers and Graphics, 80(??):ii, May 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849319300494.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Editorial

Board. Computers and Graphics, 81(??):ii, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849319300688.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Computers and Graphics, 82(??):ii, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S009784931930113X.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Computers and Graphics, 83(??):ii, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301517.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Computers and Graphics, 84(??):ii, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849319301669.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Computers and Graphics, 85(??):ii, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301797.

Anonymous:2019:PD

[Ano19i] Anonymous. Pages 1–110 (December 2019). Computers and Graphics, 85 (??):1–110, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2019:PN

[Ano19j] Anonymous. Pages 1–212 (November 2019). Computers and Graphics, 84 (??):1–212, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2019:TR

[Ano19k] Anonymous. Thanks to reviewers 2018. Computers and Graphics, 78(??):116– 119, February 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300123.

Anonymous:2020:EBa

[Ano20a]  Anonymous.  Editorial

Board. Computers and Graphics, 86(??):ii, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849320300157.

Anonymous:2020:EBb

[Ano20b]  Anonymous.  Editorial

Board. Computers and Graphics, 87(??):ii, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849320300340.

Anonymous:2020:EBc

[Ano20c]  Anonymous.  Editorial

Board. Computers and Graphics, 88(??):ii, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849320300510.

Anonymous:2020:EBd

[Ano20d]  Anonymous.  Editorial

Board. Computers and Graphics, 89(??):ii, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849320300844.

Anonymous:2020:EBe

[Ano20e]  Anonymous.  Editorial

Board. Computers and Graphics, 90(??):ii, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849320301060.

Anonymous:2020:EBf

[Ano20f]  Anonymous.  Editorial

Board. Computers and Graphics, 91(??):ii, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301333.

Anonymous:2020:EBg

[Ano20g]  Anonymous.  Editorial

Board. Computers and Graphics, 92(??):ii, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849320301618.

Anonymous:2020:EBh

[Ano20h]  Anonymous.  Editorial

Board. Computers and Graphics, 93(??):ii, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301722.

Anonymous:2020:PD

[Ano20i] Anonymous. Pages 1–108 (December 2020). Computers and Graphics, 93 (??):1–108, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PN

[Ano20j] Anonymous. Pages 1–114 (November 2020). Computers and Graphics, 92 (??):1–114, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PAa

[Ano20k] Anonymous. Pages 1–122 (April 2020). Computers and Graphics, 87(??):1– 122, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PJ

[Ano20l]  Anonymous.  Pages 1–

178 (June 2020). Computers and Graphics, 89(??):1– 178, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PAb

[Ano20m] Anonymous. Pages 1–192 (August 2020). Computers and Graphics, 90(??): 1–192, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PO

[Ano20n] Anonymous. Pages 1–300 (October 2020). Computers and Graphics, 91(??):1– 300, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PF

[Ano20o] Anonymous. Pages 1–96 (February 2020). Computers and Graphics, 86 (??):1–96, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:PM

[Ano20p]  Anonymous.  Pages 1–

96 (May 2020). Computers and Graphics, 88(??): 1–96, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2020:TR

[Ano20q]  Anonymous.  Thanks to

reviewers 2019. Computers and Graphics, 86(??): 93–96, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300091.

Anonymous:2021:EBa

[Ano21a]  Anonymous.  Editorial

Board. Computers and Graphics, 94(??):ii, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321000200.

Anonymous:2021:EBb

[Ano21b]  Anonymous.  Editorial

Board. Computers and Graphics, 95(??):ii, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321000364.

Anonymous:2021:EBc

[Ano21c]  Anonymous.  Editorial

Board. Computers and Graphics, 96(??):ii, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784932100087X.

Anonymous:2021:EBd

[Ano21d]  Anonymous.  Editorial

Board. Computers and Graphics, 97(??):iv, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784932100100X.

Anonymous:2021:EBe

[Ano21e]  Anonymous.  Editorial

Board. Computers and Graphics, 98(??):ii, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321001539.

Anonymous:2021:EBf

[Ano21f]  Anonymous.  Editorial

Board. Computers and Graphics, 99(??):ii, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001990.

Anonymous:2021:EBg

[Ano21g]  Anonymous.  Editorial

Board. Computers and Graphics, 100(??):ii, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321002363.

Anonymous:2021:EBh

[Ano21h]  Anonymous.  Editorial

Board. Computers and Graphics, 101(??):ii, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002545.

Anonymous:2021:FSS

[Ano21i] Anonymous. Foreword to the special section on CAD/ Graphics 2021. Computers and Graphics, 99(??): A9–A12, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001217.

Anonymous:2021:GBP

[Ano21j] Anonymous. GRSI best paper award. Computers and Graphics, 99(??):A5– A6, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001953.

Anonymous:2021:PD

[Ano21k] Anonymous. Pages 1–106 (December 2021). Computers and Graphics, 101 (??):1–106, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PN

[Ano21l] Anonymous. Pages 1–152 (November 2021). Computers and Graphics, 100 (??):1–152, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PAa

[Ano21m] Anonymous. Pages 1–164 (April 2021). Computers and Graphics, 95(??):1– 164, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PF

[Ano21n] Anonymous. Pages 1–196 (February 2021). Computers and Graphics, 94 (??):1–196, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PO

[Ano21o] Anonymous. Pages 1–272 (October 2021). Computers and Graphics, 99(??):1– 272, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PJ

[Ano21p]  Anonymous.  Pages 1–

292 (June 2021). Computers and Graphics, 97(??):1– 292, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PAb

[Ano21q] Anonymous. Pages 1–344 (August 2021). Computers and Graphics, 98(??): 1–344, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:PM

[Ano21r]  Anonymous.  Pages 1–

82 (May 2021). Computers and Graphics, 96(??): 1–82, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2021:TR

[Ano21s] Anonymous. Thanks to reviewers 2020. Computers and Graphics, 94(??):193– 196, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000169.

Anonymous:2022:EBa

[Ano22a]  Anonymous.  Editorial

Board. Computers and Graphics, 102(??):ii, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322000152.

Anonymous:2022:EBb

[Ano22b]  Anonymous.  Editorial

Board. Computers and Graphics, 103(??):ii, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322000449.

Anonymous:2022:EBc

[Ano22c]  Anonymous.  Editorial

Board. Computers and Graphics, 104(??):ii, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322000759.

Anonymous:2022:EBd

[Ano22d]  Anonymous.  Editorial

Board. Computers and Graphics, 105(??):ii, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784932200098X.

Anonymous:2022:EBe

[Ano22e]  Anonymous.  Editorial

Board. Computers and Graphics, 106(??):ii, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322001340.

Anonymous:2022:EBf

[Ano22f]  Anonymous.  Editorial

Board. Computers and Graphics, 107(??):ii, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001728.

Anonymous:2022:EBg

[Ano22g]  Anonymous.  Editorial

Board. Computers and Graphics, 108(??):ii, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322001959.

Anonymous:2022:EBh

[Ano22h]  Anonymous.  Editorial

Board. Computers and Graphics, 109(??):ii, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322002199.

Anonymous:2022:EN

[Ano22i] Anonymous. Editorial note. Computers and Graphics, 104(??):A1–A3, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000711.

Anonymous:2022:FSS

[Ano22j] Anonymous. Foreword to special section on WEB3D 2021 selected. Computers and Graphics, 106(??): A6, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001303.

Anonymous:2022:GBP

[Ano22k] Anonymous. GRSI Best Paper Award 2021. Computers and Graphics, 108 (??):A6, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001911.

Anonymous:2022:PAAh

[Ano22l] Anonymous. Pages 1–120, A1–A6 (December 2022). Computers and Graphics, 109(??):1–120, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAd

[Ano22m] Anonymous. Pages 1–152, A1–A14 (June 2022). Computers and Graphics, 105 (??):1–152, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAc

[Ano22n] Anonymous. Pages 1–172, A1–A6 (May 2022). Computers and Graphics, 104 (??):1–172, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAb

[Ano22o] Anonymous. Pages 1–222, A1–A12 (April 2022). Computers and Graphics, 103 (??):1–222, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAe

[Ano22p] Anonymous. Pages 1–298, A1–A6 (August 2022). Computers and Graphics, 106 (??):1–298, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAf

[Ano22q]  Anonymous.  Pages 1–

328, A1–A12 (October 2022). Computers and Graphics, 107(??):1–328, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAa

[Ano22r] Anonymous. Pages 1–646, A1–A16 (February 2022). Computers and Graphics, 102(??):1–646, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:PAAg

[Ano22s] Anonymous. Pages 1–86, A1–A6 (November 2022). Computers and Graphics, 108(??):1–86, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2022:TR

[Ano22t]  Anonymous.  Thanks to

reviewers 2021. Computers and Graphics, 102(??): i–vii, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200022X.

Anonymous:2023:CGR

[Ano23a]  Anonymous.  Computers

and Graphics reviewer acknowledgments 2022. Computers and Graphics, 110 (??):I–XI, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300016X.

Anonymous:2023:EBa

[Ano23b]  Anonymous.  Editorial

Board. Computers and Graphics, 110(??):ii, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323000201.

Anonymous:2023:EBb

[Ano23c]  Anonymous.  Editorial

Board. Computers and Graphics, 111(??):ii, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323000535.

Anonymous:2023:EBc

[Ano23d]  Anonymous.  Editorial

Board. Computers and Graphics, 112(??):ii, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323000870.

Anonymous:2023:EBd

[Ano23e]  Anonymous.  Editorial

Board. Computers and Graphics, 113(??):ii, June 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323001012.

Anonymous:2023:EBe

[Ano23f]  Anonymous.  Editorial

Board. Computers and Graphics, 114(??):ii, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323002145.

Anonymous:2023:EBf

[Ano23g]  Anonymous.  Editorial

Board. Computers and Graphics, 115(??):ii, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002650.

Anonymous:2023:EBg

[Ano23h]  Anonymous.  Editorial

Board. Computers and Graphics, 116(??):ii, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323002789.

Anonymous:2023:EBh

[Ano23i]  Anonymous.  Editorial

Board. Computers and Graphics, 117(??):ii, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002972.

Anonymous:2023:ENC

[Ano23j] Anonymous. Editorial note Computers and Graphics issue 114. Computers and Graphics, 114(??):A1– A3, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002078.

Anonymous:2023:ECA

[Ano23k] Anonymous. Editor’s choice award 2022. Computers and Graphics, 117(??): A3, December 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002935.

Anonymous:2023:NCG

[Ano23l] Anonymous. Note Computers and Graphics issue 116. Computers and Graphics, 116(??):A1–A3, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002741.

Anonymous:2023:PAAd

[Ano23m] Anonymous. Pages 1–112, A1–A2 (June 2023). Computers and Graphics, 113 (??):1–112, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAc

[Ano23n] Anonymous. Pages 1–142, A1–A8 (May 2023). Computers and Graphics, 112 (??):1–142, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAa

[Ano23o] Anonymous. Pages 1–172, A1–A6 (February 2023). Computers and Graphics, 110(??):1–172, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAh

[Ano23p] Anonymous. Pages 1–222, A1–A4 (December 2023). Computers and Graphics, 117(??):1–222, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAb

[Ano23q] Anonymous. Pages 1–224, A1–A6 (April 2023). Computers and Graphics, 111 (??):1–224, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAe

[Ano23r] Anonymous. Pages 1–414, A1–A6 (August 2023). Computers and Graphics, 114 (??):1–414, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAg

[Ano23s] Anonymous. Pages 1–512, A1–A6 (November 2023). Computers and Graphics, 116(??):1–512, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2023:PAAf

[Ano23t]  Anonymous.  Pages 1–

542, A1–A12 (October 2023). Computers and Graphics, 115(??):1–542, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2024:ACG

[Ano24a] Anonymous. Acknowledging computers and graphics reviewers. Computers and Graphics, 118(??):A3– A11, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932400027X.

Anonymous:2024:A

[Ano24b] Anonymous. April 2024. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:2024:CEE

[Ano24c] Anonymous. Celebrating excellence: Eurographics honors three distinguished Associate Editors. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000633.

Anonymous:2024:EBa

[Ano24d]  Anonymous.  Editorial

Board. Computers and Graphics, 118(??):ii, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849324000323.

Anonymous:2024:EBb

[Ano24e]  Anonymous.  Editorial

Board. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849324000670.

Anonymous:2024:ENC

[Ano24f]  Anonymous.  Editorial

note Computers and Graphics issue 119. Computers and Graphics, 119(??): ??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000621.

Anonymous:2024:NCG

[Ano24g] Anonymous. Note Computers and Graphics issue 118. Computers and Graphics, 118(??):A1–A2, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000268.

Anonymous:2024:PAA

[Ano24h] Anonymous. Pages 1–232, A1–A12 (February 2024). Computers and Graphics, 118(??):1–232, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Asteasu:1991:PPS

[AO91] C. Asteasu and A. Orbegozo. Parametric piecewise surfaces intersection. Computers and Graphics, 15 (1):9–13, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Alexanderson:2017:RTL

[AOB17] Simon Alexanderson, Carol O’Sullivan, and Jonas Beskow. Real-time labeling of nonrigid motion capture marker sets. Computers and Graphics, 69(??):59–67, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730167X.

Aoki:2016:FSS

[Aok16] Miho Aoki. Foreword to the special section on SIG-GRAPH Asia 2015 Symposium on Education. Computers and Graphics, 61 (??):A1, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630084X.

Ang:1996:RAI

[AOL96] Y. H. Ang, S. H. Ong, and Zhao Li. Retrieval of artifact images using multidimensional multiresolution features. Computers and Graphics, 20(1):51– 59, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-

bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500092.

Aono:1990:AMC

[Aon90] Masaki Aono. Attribute mapping — concept and implementation. Computers and Graphics, 14(3–4): 465–475, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ajjanagadde:1988:DPE

[AP88] Venkatramana G. Ajjanagadde and L. M. Patnaik. Design and performance evaluation of a systolic architecture for hidden surface removal. Computers and Graphics, 12(1):71–74, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Auer:1999:IOM

[AP99] Thomas Auer and Axel Pinz. The integration of optical and magnetic tracking for multi-user augmented reality. Computers and Graphics, 23(6): 805–808, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/32/article.
pdf.

Ayerbe:2022:CVR

[AP22] Alejandro Cosin Ayerbe and Gustavo Patow. Clustered voxel real-time global illumination. Computers and Graphics, 103(??):75– 89, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200005X.

Abel:2011:DPD

[APA^{+}11] R. L. Abel, S. Parfitt, N. Ashton, Simon G. Lewis, Beccy Scott, and C. Stringer. Digital preservation and dissemination of ancient lithic technology with modern micro-CT. Computers and Graphics, 35(4):878– 884, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000409.

Armstrong:2007:ISI

[APB07] Christopher J. Armstrong, Brian L. Price, and William A. Barrett. Interactive segmentation of image volumes with Live Surface. Computers and Graphics, 31(2): 212–229, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930600241X.

Appelt:1987:ESI

[App87] Wolfgang Appelt. Existing systems for integrating text and graphics. Computers and Graphics, 11(4): 369–375, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anastacio:2009:SBP

[APS09] Fabricio Anastacio, Przemysław Prusinkiewicz, and Mario Costa Sousa. Sketchbased parameterization of Lsystems using illustrationinspired construction lines and depth modulation. Computers and Graphics, 33 (4):440–451, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000806.

Ashikhmin:2001:VAM

[APSS01] Michael Ashikhmin, Simon Premože, Peter Shirley, and Brian Smits. A variance analysis of the Metropolis Light Transport algorithm. Computers and Graphics, 25(2):287–294, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/35/article.
pdf.

Aldefeld:1984:STD

[AR84] B. Aldefeld and H. Richter. Semiautomatic three-dimensional interpretation of line drawings. Computers and Graphics, 8(4):371–380, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arbeloa:1992:VFE

[Arb92] Francisco Jose Seron Arbeloa. Visualization and finite element techniques for seismic interpretation. Computers and Graphics, 16 (4):383–394, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ashbaugh:1978:DSD

[ARL78] J. B. Ashbaugh, D. P. Roland, and L. F. Laird. DSPOBJ — system for display of multiple sets of threedimensional data. Computers and Graphics, 3(2– 3):63–70, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Andrade:2020:ULB

[ARL^{+}20] Beatriz Trinchão Andrade, Benjamin Resch, Hendrik P. A. Lensch, Olga Regina Pereira Bellon, and Luciano Silva. An unstructured lumigraph based approach to the SVBRDF estimation problem. Computers and Graphics, 93(??):95–107, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301473.

Acevedo:2023:ORM

[ARM23] Pedro Acevedo, Banafsheh Rekabdar, and Christos Mousas. Optimizing retroreflective marker set for motion capturing props. Computers and Graphics, 115 (??):181–190, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001504.

Arokiasamy:1989:HCP

[Aro89] A. Arokiasamy. Homogeneous coordinates and the principle of duality in two dimensional clipping. Computers and Graphics, 13(1): 99–100, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Andreev:1991:NAT

[AS91] Rumen Andreev and Elena Sofianska. New algorithm for two-dimensional line clipping. Computers and Graphics, 15(4):519–526, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Albert:1995:RRF

[AS95] Thomas A. Albert and Dick W. Slaaf. A rapid regional filling technique for complex binary images. Computers and Graphics, 19(4):541–549, July– August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=4&aid=9500032.

Alexa:2016:FSI

[AS16] Marc Alexa and Michela Spagnuolo. Foreword to the special issue on SMI 2016. Computers and Graphics, 58 (??):A1–A2, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300693.

Alderson:2019:MNC

[AS19] Troy Alderson and Faramarz Samavati. Multiscale NURBS curves on the sphere and ellipsoid. Computers and Graphics, 82 (??):243–249, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300913.

Alicioglu:2022:SVA

[AS22] Gulsum Alicioglu and Bo Sun.

A survey of visual analytics for explainable artificial intelligence methods. Computers and Graphics, 102 (??):502–520, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001886.

Akleman:2017:IMS

[ASC17] Ergun Akleman, Vinod Srinivasan, and Jianer Chen. Interactive modeling of smooth manifold meshes with arbitrary topology: G^{1} stitched bi-cubic Bézier patches. Computers and Graphics, 66(??):64–73, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317300584.

Akleman:2018:RT

[ASC18] Ergun Akleman, Vinod Srinivasan, and Jianer Chen. Response to “On G^{1} stitched bi-cubic Bézier patches with arbitrary topology”. Computers and Graphics, 71 (??):157–158, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301954.

See [Pet18].

Antal:2003:MSS

[ASKCK03] György Antal, László Szirmay-Kalos, Ferenc Csonka, and Csaba Kelemen. Multiple strategy stochastic iteration for architectural walkthroughs. Computers and Graphics, 27(2):285– 292, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Adan:2012:HHC

[ASM12] Antonio Adán, Santiago Salamanca, and Pilar Merchán. A hybrid human-computer approach for recovering incomplete cultural heritage pieces. Computers and Graphics, 36(1):1– 15, February 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931100149X.

Aguilera:2015:PRM

[ASPO15] Alejandro Rodríguez Aguilera, Alejandro León Salas, Domingo Martín Perandrés, and Miguel A. Otaduy. A parallel resampling method for interactive deformation of volumetric models. Computers and Graphics, 53 (part B)(??):147– 155, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001636.

Antweiler:2022:UCI

[ASR^{+}22] Dario Antweiler, David Sessler, Maxim Rossknecht, Benjamin Abb, Sebastian Ginzel, and Jörn Kohlhammer. Uncovering chains of infections through spatiotemporal and visual analysis of COVID-19 contact traces. Computers and Graphics, 106(??):1–8, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000851.

Ang:2019:PCB

[ASS^{+}19] Kathleen D. Ang, Faramarz F. Samavati, Samin Sabokrohiyeh, Julio Garcia, and Mohammed S. Elbaz. Physicalizing cardiac blood flow data via 3D printing. Computers and Graphics, 85 (??):42–54, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301554.

Akman:2022:DGA

[ASS22] Alican Akman, Yusuf Sahillioglu, and T. Metin Sezgin. Deep generation of 3D articulated models and animations from 2D stick figures. Computers and Graphics, 109(??):65– 74, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001820.

Altenhofen:2017:VSC

[ASSF17] Christian Altenhofen, Felix Schuwirth, André Stork, and Dieter Fellner. Volumetric subdivision for consistent implicit mesh generation. Computers and Graphics, 69 (??):68–79, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301577.

An:2011:GPE

[ASWL11] Yi An, Cheng Shao, Xiaoliang Wang, and Zhuohan Li. Geometric properties estimation from discrete curves using discrete derivatives. Computers and Graphics, 35 (4):916–930, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000343.

Artusi:2014:HOW

[ASZ^{+}14] Alessandro Artusi, Zhuo Su, Zongwei Zhang, Dimitris Drikakis, and Xiaonan Luo. High-order wavelet reconstruction for multi-scale edge aware tone mapping. Computers and Graphics, 45 (??):51–63, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000764.

Abellan:2008:MVR

[AT08] Pascual Abellán and Dani Tost. Multimodal volume rendering with 3D textures. Computers and Graphics, 32 (4):412–419, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000514.

Ando:2011:VGD

[AT11] Ryoichi Ando and Reiji Tsuruno. Vector graphics depicting marbling flow. Computers and Graphics, 35 (1):148–159, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001718.

Al-Thelaya:2021:PII

[ATAG^{+}21] Khaled Al-Thelaya, Marco Agus, Nauman Ullah Gilal, Yin Yang, Giovanni Pintore, Enrico Gobbetti, Corrado Calí, Pierre J. Magistretti, William Mifsud, and Jens Schneider. InShaDe: Invariant shape descriptors for visual 2D and 3D cellular and nuclear shape analysis and classification. Computers and Graphics, 98 (??):105–125, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000790.

Agathos:1998:TSE

[ATB98] A. Agathos, T. Theoharis, and A. Boehm. Technical section — efficient integer algorithms for the generation of conic sections. Computers and Graphics, 22(5):621–628, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        5/603.pdf.

Al-Taie:2014:UEV

[ATHL14] Ahmed Al-Taie, Horst K. Hahn, and Lars Linsen. Uncertainty estimation and visualization in probabilistic segmentation. Computers and Graphics, 39(??): 48–59, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001532.

Arbel:2019:PCS

[ATZM19] Nadav Yehonatan Arbel, Ayellet Tal, and Lihi Zelnik-Manor. Partial correspondence of 3D shapes using properties of the nearestneighbor field. Computers and Graphics, 82(??):183– 192, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930072X.

Augusto:1984:CBE

[Aug84] Alberto Augusto, Jr. CAD in a Brazilian engineering firm. Computers and Graphics, 8(3):227–230, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Aumann:1989:ADS

[Aum89] Günter Aumann. Approximate development of skew ruled surfaces. Computers and Graphics, 13(3): 361–366, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Arbelot:2017:LTB

[AVHT17] B. Arbelot, R. Vergne, T. Hurtut, and J. Thollot. Local texture-based color transfer and colorization. Computers and Graphics, 62(??):15–27, ???? 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630139X.

Arroyo:2005:SDV

[AVM05] Germán Arroyo, Francisco Velasco, and Domingo Martín. Silhouette detection in volumetric models based on a non-photorealistic illumination system. Computers and Graphics, 29(2):209– 216, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Andrews:1993:PSC

[AW93] Phil Andrews and Joel Welling. The Pittsburgh Supercomputing Center’s computer graphics environment. Computers and Graphics, 17 (1):5–8, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

AlMeraj:2009:AMU

[AWI^{+}09] Zainab AlMeraj, Brian Wyvill, Tobias Isenberg, Amy A. Gooch, and Richard Guy. Automatically mimicking unique hand-drawn pencil lines. Computers and Graphics, 33(4):496– 508, August 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000636.

Akram:2022:SSP

[AXC22] Muhammad Naeem Akram, Kaoji Xu, and Guoning Chen. Structure simplification of planar quadrilateral meshes. Computers and Graphics, 109(??): 1–14, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001790.

Akleman:2013:HCA

[AXG^{+}13] Ergun Akleman, Qing Xing, Pradeep Garigipati, Gabriel Taubin, Jianer Chen, and Shiyu Hu. Hamiltonian cycle art: Surface covering wire sculptures and duotone surfaces. Computers and Graphics, 37(5):316– 332, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300006X.

Akiyama:2020:ILP

[AYA^{+}20] Ryo Akiyama, Goshiro Yamamoto, Toshiyuki Amano, Takafumi Taketomi, Alexander Plopski, Yuichiro Fujimoto, Masayuki Kanbara, Christian Sandor, and Hirokazu Kato. Illusory light: Perceptual appearance control using a projectioninduced illusion. Computers and Graphics, 91(??): 129–140, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301102.

Aghdaii:2012:MRA

[AYZ12] Nima Aghdaii, Hamid Younesy, and Hao Zhang. 5-6-7 meshes: Remeshing and analysis. Computers and Graphics, 36(8):1072–1083, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001574.

Aydinlilar:2023:FIF

[AZ23] Melike Aydinlilar and Cédric Zanni. Forward inclusion functions for ray-tracing implicit surfaces. Computers and Graphics, 114(??):190– 200, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300081X.

Athanasiadis:2013:ECC

[AZF13] Theodoros Athanasiadis, Georgios Zioupos, and Ioannis Fudos. Efficient computation of constrained parameterizations on parallel platforms. Computers and Graphics, 37(6):596– 607, October 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000988.

Behr:2000:MVI

[B^{+}00] Johannes Behr et al. Modelling, visualization, and interaction techniques for diagnosis and treatment planning in cardiology. Computers and Graphics, 24 (5):741–753, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/34/article.
pdf.

Berger:1990:GRM

[BA90] H. Berger and J. Altenbach. Graphical representation of models and results in the finite element system COSAR. Computers and Graphics, 14(3–4): 395–404, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Boubekeur:2009:MSS

[BA09] Tamy Boubekeur and Marc Alexa. Mesh simplification by stochastic sampling and topological clustering. Computers and Graphics, 33(3):241–249, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000417.

Bernard:2023:CGS

[BA23] Jürgen Bernard and Marco Angelini. Computers and Graphics special section on the 13th International EuroVis Workshop on Visual Analytics (EuroVA) 2022. Computers and Graphics, 115(??):A4–A5, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001358.

Bulbul:2014:CBM

[BAC14] Abdullah Bulbul, Sami Arpa, and Tolga Capin. A clustering-based method to estimate saliency in 3D animated meshes. Computers and Graphics, 43(??): 11–20, October 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000521.

Badrinath:1996:DCM

[Bad96] B. R. Badrinath. Distributed computing in mobile environments. Computers and Graphics, 20(5):615–617, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.

cgi?year=1996&volume=20&
issue=5&aid=9600035.

Baldi:2023:DPP

[BAD23] Guillaume Baldi, Rémi Allègre, and Jean-Michel Dischler. Differentiable point process texture basis functions for inverse procedural modeling of cellular stochastic structures. Computers and Graphics, 112(??): 116–131, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000419.

Bagiana:1993:TSJ

[Bag93] Frank Bagiana. Tomorrow’s space: Journey to the virtual worlds. Computers and Graphics, 17(6):687–690, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Berk:2003:DVR

[BAG03] Hakan Berk, Cevdet Aykanat, and Uğur Güdükbay. Direct volume rendering of unstructured grids. Computers and Graphics, 27(3):387– 406, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bailey:2007:TOS

[Bai07]  Mike Bailey.  Teaching

OpenGL shaders: Handson, interactive, and immediate feedback. Computers and Graphics, 31(3): 524–531, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000647.

Bannon:1985:CEA

[Ban85] Robert T. Bannon. Creating and EDP architecture via computer graphics — a tool for systems integration management. Computers and Graphics, 9(1):51–56, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Banks:1997:IDR

[Ban97a] David C. Banks. The ImageSwitcher: a design for reducing VR lag via parallel image generation. Computers and Graphics, 21(2): 171–177, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600080.

Banks:1997:IGV

[Ban97b] David C. Banks. Including graphics and visualization research in a Master’slevel course. Computers and Graphics, 21(3):379– 387, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700016.

Baptista:1999:CGM

[Bap99] Humberto Rossetti Baptista.

Chaos and graphics — a method for incremental image generation. Computers and Graphics, 23(3):449– 454, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/694.pdf.

Besnier:2023:TMI

[BAPD23] Thomas Besnier, Sylvain Arguillère, Emery Pierson, and Mohamed Daoudi. Toward mesh-invariant 3D generative deep learning with geometric measures. Computers and Graphics, 115 (??):309–320, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300122X.

Barten:1977:AFE

[Bar77] H. J. Barten. Automated finite element grid breakup method — a verification of the six node averaging approach. Computers and Graphics, 2(2):75–79, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bar:1990:CWT

[Bär90] Gert Bär. CAD of worms and their machining tools. Computers and Graphics, 14 (3–4):405–411, 1990. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Basil:1977:MBR

[Bas77] M. Basil. A microprocessorbased refreshing buffer for storage tube graphics terminals. Computers and Graphics, 2(4):205–208, 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brazil:2015:SBM

[BAS^{+}15] Emilio Vital Brazil, Ronan Amorim, Mario Costa Sousa, Luiz Velho, and Luiz Henrique de Figueiredo. Sketchbased modeling and adaptive meshes. Computers and Graphics, 52(??):116– 128, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001235.

Boges:2020:VRF

[BAS^{+}20] Daniya Boges, Marco Agus, Ronell Sicat, Pierre J. Magistretti, Markus Hadwiger, and Corrado Calì. Virtual reality framework for editing and exploring medial axis representations of nanometric scale neural structures. Computers and Graphics, 91 (??):12–24, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300789.

Bawa:1997:FAT

[Baw97] Bhalinder S. Bawa. Faster apparent three dimensional rotation using parallel projection perpendicular to the xy plane. Computers and Graphics, 21(1):15– 21, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600066.

Baxter:1977:UGC

[Bax77] R. S. Baxter. Use of graphics in a computer-based urban information system. Computers and Graphics, 2(3): 155–161, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bayarri:1995:CNP

[Bay95] Salvador Bayarri. Computing non-planar perspectives in real time. Computers and Graphics, 19(3): 431–440, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500013.

Brodlie:1991:PCU

[BB91] K. W. Brodlie and S. Butt. Preserving convexity using piecewise cubic interpolation. Computers and Graphics, 15(1):15–23, 1991. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Breene:1993:IWS

[BB93a] L. Anne Breene and Jack Bryant. Image warping by scanline operations. Computers and Graphics, 17(2): 127–130, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Butt:1993:PPU

[BB93b] S. Butt and K. W. Brodlie. Preserving positivity using piecewise cubic interpolation. Computers and Graphics, 17(1):55–64, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Briggs:1998:SVR

[BB98] Edward S. Briggs, Jr. and R. Daniel Bergeron. A selfvisualizing rendering support environment. Computers and Graphics, 22(4): 547–555, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/4/574.pdf.

Blanke:2003:AVM

[BB03] William J. Blanke and Chandrajit Bajaj. Active visualization in a multidisplay immersive environment. Computers and Graphics, 27 (5):681–691, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Birklbauer:2015:AGL

[BB15] C. Birklbauer and O. Bimber. Active guidance for light-field photography on smartphones. Computers and Graphics, 53 (part B)(??):127–135, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001533.

Bellotti:2011:AAE

[BBCG11] Francesco Bellotti, Riccardo Berta, Rosario Cardona, and Alessandro De Gloria. An architectural approach to efficient 3D urban modeling. Computers and Graphics, 35(5): 1001–1012, October 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001336.

Bellotti:2003:MDE

[BBDM03] F. Bellotti, R. Berta, A. De Gloria, and M. Margarone. MADE: developing edutainment applications on mobile computers. Computers and Graphics, 27(4):617– 634, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Blumenthal-Barby:2014:HRD

[BBE14] David C. Blumenthal-Barby and Peter Eisert. Highresolution depth for binocular image-based modeling. Computers and Graphics, 39(??):89–100, April 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313002045.

Barth:1990:UGS

[BBH90] R. Barth, R. Bose, and I. Heilemann. UGRAF3.

A graphic system for process and modelling. Computers and Graphics, 14(3– 4):429–433, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Barber:2015:ICI

[BBHC15] Alastair Barber, Matthew Brown, Paul Hogbin, and Darren Cosker. Inferring changes in intrinsic parameters from motion blur. Computers and Graphics, 52(??): 155–170, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000692.

Bernal-Berdun:2022:SSS

[BBMGM22] Edurne Bernal-Berdun, Daniel Martin, Diego Gutierrez, and Belen Masia. SST-Sal: a spherical spatio-temporal approach for saliency prediction in 360^{}˚ videos. Computers and Graphics, 106 (??):200–209, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001042.

Bruza:2021:PVC

[BBMK21] Vojtech Bruza, Jan Byska, Jan Mican, and Barbora Kozlíková. VRdeo: Creating engaging educational material for asynchronous student–teacher exchange using virtual reality. Computers and Graphics, 98 (??):280–292, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001242.

Bettels:1989:GDW

[BBMR89] Jurgen Bettels, Peter R. Bono, Eileen McGinnis, and Joachim Rix. Guidelines for determining when to use GKS and when to use PHIGS. Computers and Graphics, 13(1):91–98, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Busking:2010:EBI

[BBP10] Stef Busking, Charl P. Botha, and Frits H. Post. Example-based interactive illustration of multi-field datasets. Computers and Graphics, 34(6):719–728, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000110X.

Barroso:2013:VCP

[BBP13] Santiago Barroso, Gonzalo Besuievsky, and Gustavo Patow. Visual copy & paste for procedurally modeled buildings by ruleset rewriting. Computers and Graphics, 37(4):238–246, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000058.

Bonfatti:1988:SEC

[BC88] Flavio Bonfatti and Lorenza Cavazza. SECT: An effective coding technique for polygonal geographic data. Computers and Graphics, 12 (3–4):503–513, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Braines:2001:FEV

[BC01] Simon A. Braines and Richard J. Cant. A framework for the evaluation of volume rendering techniques on a task specific basis using neural networks. Computers and Graphics, 25(4):643– 663, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/38/abstract.
html.

Bonnici:2013:CGA

[BC13] Alexandra Bonnici and Kenneth P. Camilleri. A constrained genetic algorithm for line labelling of line drawings with shadows and table-lines. Computers and Graphics, 37 (5):302–315, August 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000101.

Bulbul:2010:PAS

[BCC10] Abdullah Bulbul, Zeynep Cipiloglu, and Tolga Capin. A perceptual approach for stereoscopic rendering optimization. Computers and Graphics, 34(2):145– 157, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001526.

Bobak:2020:TSB

[BCC20] Petr Bobák, Ladislav Cmolík, and Martin Cadík. Temporally stable boundary labeling for interactive and noninteractive dynamic scenes. Computers and Graphics, 91 (??):265–278, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301230.

Bhat:2022:SAG

[BCC^{+}22] Krishnaraj Vilasraj Bhat, Gabriele Capasso, Simone Coniglio, Joseph Morlier, and Christian Gogu. On some applications of Generalized Geometric Projection to optimal 3D printing. Computers and Graphics, 102(??):199–212, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100217X.

Bourquat:2022:HMP

[BCDD22] Pierre Bourquat, David Coeurjolly, Guillaume Damiand, and Florent Dupont. Hierarchical mesh-to-points as-rigid-as-possible registration. Computers and Graphics, 102(??):320–328, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002260.

Bordegoni:2006:HTC

[BCF06] Monica Bordegoni, Giorgio Colombo, and Luca Formentini. Haptic technologies for the conceptual and validation phases of product design. Computers and Graphics, 30(3): 377–390, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000604.

Belloli:1994:RNM

[BCG^{+}94] V. Belloli, S. Callegari, C. Gatti, M. Della Monica, and D. Marini. Ray-Filling: a new method to accelerate ray casting. Computers and Graphics, 18(5):723–732, September– October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Back:2002:SRE

[BCHM02] Maribeth Back, Jonathan Cohen, Steve Harrison, and Scott Minneman. Speeder Reader: an experiment in the future of reading. Computers and Graphics, 26(4):623–627, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/43/abstract.
                                        html.

Bist:2017:TEU

[BCMD17] Cambodge Bist, Rémi Cozot, Gérard Madec, and Xavier Ducloux. Tone expansion using lighting style aesthetics. Computers and Graphics, 62(??): 77–86, ???? 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301406.

Botsch:2007:SSJ

[BCMM07] Mario Botsch, Baoquan Chen, Raghu Machiraju, and Torsten Möller. Special section on the joint Symposium on Point-based Graphics and Volume Graphics 2006. Computers and Graphics, 31(2):155–156, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000337.

Behringer:1999:DDD

[BCS^{+}99] Reinhold Behringer, Steven Chen, Venkataraman Sundareswaran, Kenneth Wang, and Marius Vassiliou. A distributed device diagnostics system utilizing augmented reality and 3D audio. Computers and Graphics, 23 (6):821–825, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/34/00000751.
pdf.

Baldassarri:2008:MPE

[BCS08] Sandra Baldassarri, Eva Cerezo, and Francisco J. Seron. Maxine: a platform for embodied animated agents. Computers and Graphics, 32(4):430– 437, August 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000472.

Braquelaire:1997:GTH

[BD97] Jean-Pierre Braquelaire and Jean-Philippe Domenger. Geometrical, topological, and hierarchical structuring of overlapping 2-D discrete objects. Computers and Graphics, 21(5):587–597, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=5&aid=9700037.

Blundo:1998:VCS

[BD98] Carlo Blundo and Alfredo De Santis. Visual cryptography schemes with perfect reconstruction of black pixels. Computers and Graphics, 22(4):449–455, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        4/568.pdf.

Bender:2013:ACS

[BD13] Jan Bender and Crispin Deul. Adaptive cloth simulation using corotational finite elements. Computers and Graphics, 37(7): 820–829, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000605.

Bulbul:2017:SMB

[BD17] Abdullah Bulbul and Rozenn Dahyot. Social media based 3D visual popularity. Computers and Graphics, 63(??): 28–36, April 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300146.

Burghard:2017:ESG

[BDK17] Oliver Burghard, Alexander Dieckmann, and Reinhard Klein. Embedding shapes with Green’s functions for global shape matching. Computers and Graphics, 68 (??):1–10, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730078X.

Borchers:1996:LRG

[BDKK96] Jan Borchers, Oliver Deussen, Arnold Klingert, and Clemens Knörzer. Layout rules for graphical Web documents. Computers and Graphics, 20(3):415–426, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=20&
                                        issue=3&aid=9600011.

Biasotti:2022:FSIa

[BDL^{+}22] Silvia Biasotti, Roberto M. Dyke, Yu-Kun Lai, Paul L. Rosin, and Remco Veltkamp. Foreword to the special issue on 3D object retrieval 2021 workshop (3DOR2021). Computers and Graphics, 102(??):A14–A16, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849321002168.

Brousset:2016:SCB

[BDM^{+}16] Mathias Brousset, Emmanuelle Darles, Daniel Meneveaux, Pierre Poulin, and Benoît Crespin. Simulation and control of breaking waves using an external force model. Computers and Graphics, 57(??): 102–111, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300164.

Barth:1982:HLG

[BDP82] W. Barth, J. Dirnberger, and W. Purgathofer. The high-level graphics programming language PAS-CAL/GRAPH. Computers and Graphics, 6(3):109–119, 1982. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Bussler:2017:VFP

[BDP^{+}17] Michael Bußler, Patrick Diehl, Dirk Pflüger, Steffen Frey, Filip Sadlo, Thomas Ertl, and Marc Alexander Schweitzer. Visualization of fracture progression in peridynamics. Computers and Graphics, 67 (??):45–57, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300456.

Bachmann:1993:FBM

[BDPR93] Tobias Bachmann, Martin Daniel, Gerhard Pahl, and Joachim Rix. Feature-based modeling in support of embodiment design. Computers and Graphics, 17(3): 285–294, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bellotti:2001:AJG

[BDRV01] F. Bellotti, A. De Gloria, M. Risso, and A. Villamaina. AutoGraL: a Java 2D graphics library for configurable automotive dashboards. Computers and Graphics, 25(2): 259–268, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/33/article.
                                        pdf.

Baron:2022:MBA

[BDSP22] Jessica Baron, Daljit Singh Dhillon, N. Adam Smith, and Eric Patterson. Microstructurebased appearance rendering for feathers. Computers and Graphics, 102 (??):452–459, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002028.

Bebb:1975:IES

[Beb75] Joan Bebb. An interactive environment for scientific development. Computers and Graphics, 1(4): 303–307, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bechmann:1994:SDM

[Bec94] Dominique Bechmann. Space deformation models survey. Computers and Graphics, 18(4):571–586, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Beckmann:1995:PVP

[Bec95] Peter E. Beckmann. On the problem of visualizing point distributions in high dimensional spaces. Computers and Graphics, 19(4): 617–629, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500040.

Burrill:1994:LES

[BEFV94] Victoria Burrill, Gillie Evans, Dirk Fokken, and Kaisa Väänänen. The lust to explore space: The attractiveness of interactive video within multimedia applications. Computers and Graphics, 18(5):675–683, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brodlie:2000:UWB

[BEKL00] Ken Brodlie, Nuha El-Khalili, and Ying Li. Using Web-based computer graphics to teach surgery. Computers and Graphics, 24 (1):157–161, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/41/article.
pdf.

Benest:1979:RCG

[Ben79] I. D. Benest. A review of computer graphics publications. Computers and Graphics, 4(2):95–136, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Berk:1977:CFC

[Ber77] Toby S. Berk. Coordinatefree computer graphics. Computers and Graphics, 2 (2):105–109, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Berk:1979:SGS

[Ber79] T. S. Berk. A satellite graphics system for errorprone users. Computers and Graphics, 4(3–4):185–187, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Beraldo:1984:GEI

[Ber84] A. T. M. Beraldo. Guest Editor’s introduction: Computer graphics in Brazil. Computers and Graphics, 8 (3):225–226, ???? 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849384900013.

Bennebroek:1997:DPH

[BERW97] K. Bennebroek, I. Ernst, H. Rüsseler, and O. Wittig. Design principles of hardware-based Phong shading and bump-mapping. Computers and Graphics, 21(2):143–149, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600077.

Bimber:2000:MLA

[BES00] Oliver Bimber, L. Miguel Encarnação, and André Stork. A multi-layered architecture for sketch-based interaction within virtual environments. Computers and Graphics, 24(6): 851–867, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/29/article.
pdf.

Bimber:2001:EMR

[BES01] O. Bimber, L. M. Encarnação, and A. Stork. Editorial: Mixed reality — beyond conventions. Computers and Graphics, 25(5):727– 730, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        13/20/57/34/27/abstract.
                                        html.

Barhak:2002:ARF

[BF02a] J. Barhak and A. Fischer. Adaptive reconstruction of freeform objects with 3D SOM neural network grids. Computers and Graphics, 26(5):745–751, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/56/37/abstract.
                                        html.

Bono:2002:GEI

[BF02b] Peter R. Bono and Dieter W. Fellner. Guest editors’ introduction: Selected topics from the symposium: computer graphics—enabling technology for the information society. Computers and Graphics, 26(4):533, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/28/abstract.
                                        html.

Bedient:2007:CSR

[BF07] Richard Bedient and Michael Frame. Carrying surfaces for return maps of averaged logistic maps. Computers and Graphics, 31(6): 887–895, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001343.

Belyaev:2015:TGW

[BF15] Alexander G. Belyaev and Pierre-Alain Fayolle. On transfinite Gordon–Wixom interpolation schemes and their extensions. Computers and Graphics, 51(??): 74–80, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000576.

Bao:2019:SBC

[BF19] Bin Bao and Hongbo Fu.

Scribble-based colorization for creating smooth-shaded vector graphics. Computers and Graphics, 81(??): 73–81, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300445.

Biasotti:2020:FSS

[BFLP20] Silvia Biasotti, Bianca Falcidieno, Guillaume Lavoué, and Ioannis Pratikakis. Foreword to the special section on 3D object retrieval 2019. Computers and Graphics, 88(??):A4–A5, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300303.

Boschiroli:2011:CLP

[BFRA11] Maria Boschiroli, Christoph Fünfzig, Lucia Romani, and Gudrun Albrecht. A comparison of local parametric C^{0} Bézier interpolants for triangular meshes. Computers and Graphics, 35 (1):20–34, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001445.

Bimber:2003:RTV

[BFSE03] Oliver Bimber, Bernd Fröhlich, Dieter Schmalstieg, and L. Miguel Encarnação. Realtime view-dependent image warping to correct non-linear distortion for curved Virtual Showcase displays. Computers and Graphics, 27 (4):515–528, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Benato:2023:MQP

[BFT23] Bárbara C. Benato, Alexandre X. Falcão, and Alexandru C. Telea. Measuring the quality of projections of high-dimensional labeled data. Computers and Graphics, 116(??): 287–297, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001929.

Birta:1979:IGS

[BG79] L. G. Birta and B. Guirguis.

An interactive/graphics software system for function minimization. Computers and Graphics, 4(1):29–41, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Borovsky:1980:CAR

[BG80] B. Borovsky and G. Gotchev.

Computer analysis and recognition of two-dimensional linear geometrical pictures. Computers and Graphics, 5 (2–4):83–86, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bey:1988:CIS

[BG88] Ingward Bey and Ulrich Gengenbach. The CAD*I interface for solid model exchange. Computers and Graphics, 12(2):181–190, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Braquelaire:1991:DSU

[BG91] J. P. Braquelaire and P. Guitton. 21/2D scene update by insertion of contour. Computers and Graphics, 15(1):41–48, 1991. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bieber:2001:PMN

[BG01] Gerald Bieber and Martin Giersich. Personal mobile navigation systems — design considerations and experiences. Computers and Graphics, 25(4):563– 570, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/29/abstract.
html.

Basak:2018:URL

[BGD18] Ahmet Eren Basak, Ugur Güdükbay, and Funda Durupinar. Using real life incidents for creating realistic virtual crowds with data-driven emotion contagion. Computers and Graphics, 72(??):70–81, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300244.

Buhmann:1989:CGT

[BGK89] A. Buhmann, M. Günther, and G. Koberle. Computer graphics as a tool in training and education: a COMETT project. Computers and Graphics, 13(4): 523–527, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Balazs:2004:FBG

[BGK04] Ákos Balázs, Michael Guthe, and Reinhard Klein. Fat borders: gap filling for efficient view-dependent LOD NURBS rendering. Computers and Graphics, 28 (1):79–85, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Boulch:2018:SPC

[BGLA18] Alexandre Boulch, Joris Guerry, Bertrand Le Saux, and Nicolas Audebert. Snap-Net: 3D point cloud semantic labeling with 2D deep segmentation networks. Computers and Graphics, 71(??):189–198, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301942.

Bettio:2008:SRM

[BGMP08] Fabio Bettio, Enrico Gobbetti, Fabio Marton, and Giovanni Pintore. Scalable rendering of massive triangle meshes on light field displays. Computers and Graphics, 32(1):55– 64, February 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001999.

Band:2018:MPB

[BGPT18] Stefan Band, Christoph Gissler, Andreas Peer, and Matthias Teschner. MLS pressure boundaries for divergence-free and viscous SPH fluids. Computers and Graphics, 76(??):37– 46, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830116X.

Balkin:1994:CEC

[BGR94] Sandy D. Balkin, Elizabeth L. Golebiewski, and Clifford A. Reiter. Chaos and elliptic curves. Computers and Graphics, 18(1): 113–117, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Biasotti:2016:FSS

[BGT16] Silvia Biasotti, Andrea Giachetti, and Marco Tarini. Foreword to the special section on Smart Tools and Applications in Computer Graphics 2015. Computers and Graphics, 60(??): A3–A4, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300851.

Bhavsar:1993:VGF

[BGV93] Virendra C. Bhavsar, Uday G.

Gujar, and Nagarjuna Vangala. Vectorization of generation of fractals from z ← z^{2} + c on IBM 3090/180VF. Computers and Graphics, 17(2):169–174, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bickerstaff:1991:HPA

[BH91] M. A. Bickerstaff and G. R.

Hellestrand. A highly parallel architecture for real time collision detection in flight simulation. Computers and Graphics, 15(3): 355–363, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brambilla:2015:ESM

[BH15] Andrea Brambilla and Helwig Hauser. Expressive seeding of multiple stream surfaces for interactive flow exploration. Computers and Graphics, 47(??):123– 134, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000035.

Bittner:2015:IBC

[BHH15] Jirí Bittner, Michal Hapala, and Vlastimil Havran. Incremental BVH construction for ray tracing. Computers and Graphics, 47 (??):135–144, April 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001435.

Brodlie:1994:GVT

[BHL^{+}94] Ken Brodlie, Terry Hewitt, Steve Larkin, Phil Willis, and Julian Gallop. Graphics and visualization

— techniques and tools: a course for postgraduates of all disciplines. Computers and Graphics, 18(3): 263–268, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bu:2015:LDF

[BHL^{+}15] Shuhui Bu, Pengcheng Han, Zhenbao Liu, Junwei Han, and Hongwei Lin. Local deep feature learning framework for 3D shape. Computers and Graphics, 46 (??):117–129, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000958.

Boulic:1994:GOD

[BHTT94] Ronan Boulic, Zhiyong Huang, Nadia Magnenat Thalmann, and Daniel Thalmann. Goal-oriented design and correction of articulated figure motion with the TRACK system. Computers and Graphics, 18(4): 443–452, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bernard:2021:PPV

[BHZ^{+}21] Jürgen Bernard, Marco Hutter, Matthias Zeppelzauer, Michael Sedlmair, and Tamara Munzner. ProSeCo: Visual analysis of class separation measures and dataset characteristics. Computers and Graphics, 96(??): 48–60, May 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000406.

Bigelmaier:1986:PGK

[Big86] Anton Bigelmaier. Profile of a geometrical knowledge base for CAD-systems. Computers and Graphics, 10(4): 297–305, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bijl:1985:GIC

[Bij85] A. Bijl. Graphical input:

Can computers understand people. Computers and Graphics, 9(2):85–95, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bimber:2015:ISB

[Bim15] Oliver Bimber. An image sensor based on optical Radon transform. Computers and Graphics, 53 (part A)(??):37–43, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001351.

Brunet:1995:CGU

[BJAN^{+}95] P. Brunet, R. Juan-Arinyo, I. Navazo, D. Tost, and S. Villa. Computer graphics at the Universitat Politècnica de Catalunya. Computers and Graphics, 19(3): 467–474, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500018.

Bjorke:1985:CGC

[Bjo85] O. Bjorke. Computer graphics in CAM applications. Computers and Graphics, 9 (2):165–169, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Berdinsky:2012:PID

[BjOwKM12] Dmitry Berdinsky, Min jae Oh, Tae wan Kim, and Bernard Mourrain. On the problem of instability in the dimension of a spline space over a T-mesh. Computers and Graphics, 36 (5):507–513, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000441.

Bao:1997:PRA

[BJP97] Hujun Bao, Xiaogang Jin, and Qunsheng Peng. A progressive radiosity algorithm based on piecewise polynomial intensity distribution. Computers and Graphics, 21(3):281– 288, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700004.

Brachtl:2001:PBN

[BJS01] M. Brachtl, J.Šlajs, and P. Slavík. PDA based navigation system for a 3D environment. Computers and Graphics, 25(4):627– 634, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/36/abstract.
html.

Bakalash:1989:MMI

[BK89a] Reuven Bakalash and Arie Kaufman. Medicube: a 3D medical imaging architecture. Computers and Graphics, 13(2):151–157, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bryant:1989:DDB

[BK89b] Jack Bryant and Clifford Krumvieda. Display of discrete 3D binary objects. I. shading. Computers and Graphics, 13(4): 441–444, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Butler:1991:OPI

[BK91] Timothy Butler and Pat Krolak. An overview of the Programmer’s Imaging Kernel (PIK) proposed standard. Computers and Graphics, 15(4):465–472, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Burtsev:1993:EFF

[BK93] S. V. Burtsev and Ye. P. Kuzmin. An efficient floodfilling algorithm. Computers and Graphics, 17 (5):549–561, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Broll:1996:VTT

[BK96] Wolfgang Broll and Tanja Koop. VRML: Today and tomorrow. Computers and Graphics, 20(3): 427–434, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9500012.

Bischoff:2002:TRB

[BK02] Stephan Bischoff and Leif Kobbelt. Towards robust broadcasting of geometry data. Computers and Graphics, 26(5):665–675, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/30/abstract.
html.

Batty:2019:FSS

[BK19] Christopher Batty and Paul G.

Kry. Foreword to the special section on Graphics Interface 2018. Computers and Graphics, 78(??):A3– A4, February 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301948.

Bright:2020:ETC

[BK20] Matt Bright and Vitaliy Kurlin. Encoding and topological computation on textile structures. Computers and Graphics, 90(??): 51–61, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300650.

Bodonyi:2023:ETB

[BK23] Andrea Bodonyi and Roland Kunkli. Efficient tilebased rendering of lens flare ghosts. Computers and Graphics, 115(??):472– 483, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001486.

Brown:1979:EGS

[BKCS79] David C. Brown, Stan C. Kwasny, B. Chandrasekaran, and Norman K. Sondheimer. Experimental graphics system with natural language input. Computers and Graphics, 4(1):13–22, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bender:2014:PBS

[BKCW14] Jan Bender, Dan Koschier, Patrick Charrier, and Daniel Weber. Position-based simulation of continuous materials. Computers and Graphics, 44(??):1– 10, November 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400065X.

Braun:1995:MPB

[BKL^{+}95] C. Braun, T. H. Kolbe, F. Lang, W. Schickler, V. Steinhage, A. B. Cremers, W. Förstner, and L. Plümer. Models for photogrammetric building reconstruction. Computers and Graphics, 19(1):109–118, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400126.

Bui:2015:LSC

[BKL15] Minh Tuan Bui, Junho Kim, and Yunjin Lee. 3D-look shading from contours and hatching strokes. Computers and Graphics, 51 (??):167–176, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000734.

Bui:2017:HFC

[BKL17] Tuan Minh Bui, Junho Kim, and Yunjin Lee. Heightfield construction using cross contours. Computers and Graphics, 66(??): 53–63, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300699.

Bera:2016:OPL

[BKM16] Aniket Bera, Sujeong Kim, and Dinesh Manocha. Online parameter learning for data-driven crowd simulation and content generation. Computers and Graphics, 55(??):68–79, April 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001703.

Billinghurst:2001:MTA

[BKP01] Mark Billinghurst, Hirokazu Kato, and Ivan Poupyrev. The MagicBook: a transitional AR interface. Computers and Graphics, 25 (5):745–753, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/57/34/29/abstract.
                                        html.

Baldacci:2016:GBA

[BKR^{+}16] Andrea Baldacci, Rastislav Kamenický, Adam Riecický, Paolo Cignoni, Roman Durikovic, Roberto Scopigno, and Martin Madaras. GPUbased approaches for shape diameter function computation and its applications focused on skeleton extraction. Computers and Graphics, 59 (??):151–159, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300796.

Byska:2021:FSS

[BKS21] Jan Byska, Michael Krone, and Bjorn Sommer. Foreword to the special section on molecular graphics and visual analysis of molecular data (MolVA 2021). Computers and Graphics, 99 (??):A7–A8, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001473.

Byska:2023:FSS

[BKS23] Jan Byska, Michael Krone, and Bjorn Sommer. Foreword: Special section on Molecular Graphics and Visual Analysis of Molecular Data (MolVA 2023). Computers and Graphics, 112 (??):A5, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000821.

Baby:2005:ULA

[BKV05] Thomas Baby, Youngmin Kim, and Amitabh Varshney. Unsupervised learning applied to progressive compression of timedependent geometry. Computers and Graphics, 29(3): 451–461, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Broome:1982:ICP

[BL82] D. R. Broome and T. H. Lambert. Interactive computer programme to facilitate diesel engine/propeller matching. Computers and Graphics, 6(3):101–108, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bonigk:1996:BAM

[BL96] Jörg Bönigk and Astrid Lubinski. A basic architecture for mobile information access. Computers and Graphics, 20(5):683–691, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=5&aid=9600042.

Blake:2011:CRR

[BL11] Edwin Blake and Ilda Ladeira. Cultural reinterpretation and resonance: The San and hip-hop. Computers and Graphics, 35(2): 383–391, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000045.

Baek:2015:ISI

[BLL15] Seung-Yeob Baek, Jeonghun Lim, and Kunwoo Lee. Isometric shape interpolation. Computers and Graphics, 46 (??):257–263, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001137.

Bonneau:2023:FSS

[BLM23] Georges-Pierre Bonneau, Ligang Liu, and Michela Mortara. Foreword to the special section on Shape Modeling International 2023 (SMI2023). Computers and Graphics, 115(??):A6– A9, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002455.

Barrera:2022:BSN

[BLNZ22] D. Barrera, A. Lamnii, M.- Y. Nour, and A. Zidna. α-B-splines non-stationary subdivision schemes for grids of arbitrary topology design. Computers and Graphics, 108(??):34– 48, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001650.

Boehm:2015:PSS

[BLS15] Jan Boehm, Roderik Lindenbergh, and Michela Spagnuolo. Preface to the special session on processing of large geospatial data. Computers and Graphics, 49(??):A1–A2, June 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931500045X.

Bo:2012:RFP

[BLW12] Pengbo Bo, Ruotian Ling, and Wenping Wang. A revisit to fitting parametric surfaces to point clouds. Computers and Graphics, 36 (5):534–540, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000751.

Biasotti:2022:FSIb

[BLZ22] Silvia Biasotti, Yang Liu, and Bo Zhu. Foreword to the special issue on Shape Modeling International 2021 (SMI2021). Computers and Graphics, 103(??):A7– A9, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000334.

Bao:2012:LSF

[BLZD12] Guanbo Bao, Hongjun Li, Xiaopeng Zhang, and Weiming Dong. Large-scale forest rendering: Real-time, realistic, and progressive. Computers and Graphics, 36(3):140–151, May 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000064.

Bailey:2003:VPD

[BM03] Fiona Bailey and Magnus Moar. The VERTEX project: designing and populating shared 3D virtual worlds in the primary (elementary) classroom. Computers and Graphics, 27(3): 353–359, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bhoi:2008:CSF

[BM08] Nilamani Bhoi and Sukadev Meher. Circular spatial filtering under high-noise-variance conditions. Computers and Graphics, 32 (5):568–580, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000903.

Brodlie:1995:VSD

[BMB95] Ken Brodlie, Petros Mashwama, and Sohail Butt. Visualization of surface data to preserve positivity and other simple constraints. Computers and Graphics, 19(4): 585–594, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=4&aid=9500036.

Bessa:2018:EBP

[BMdSVR18] Maximino Bessa, Miguel Melo, A. Augusto de Sousa, and José Vasconcelos-Raposo. The effects of body position on Reflexive Motor Acts and the sense of presence in virtual environments. Computers and Graphics, 71(??):35–41, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301863.

Bartz:1999:OAO

[BMH99] Dirk Bartz, Michael Meißner, and Tobias Hüttner. OpenGLassisted occlusion culling for large polygonal models. Computers and Graphics, 23 (5):667–679, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
31/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/31/abstract.
html.

Bo:2023:IGI

[BMMZ23] Pengbo Bo, Xiangjian Mai, Wenlong Meng, and Caiming Zhang. Improving geometric iterative approximation methods using local approximations. Computers and Graphics, 116 (??):33–45, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300170X.

Brodlie:1984:PSC

[BMP84] K. W. Brodlie, M. C.

Maguire, and G. E. Pfaff. A practical strategy for certifying GKS implementations. Computers and Graphics, 8 (2):125–133, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Biasotti:2022:FSIc

[BMP22] Silvia Biasotti, Ramanathan Muthuganapathy, and Jörg Peters. Foreword to the special issue on Shape Modeling International 2022 (SMI2022). Computers and Graphics, 107(??):A6– A8, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001613.

Barbosa:2023:SGH

[BMR23] Filipe Barbosa, Daniel Mendes, and Rui Rodrigues.

                                        Shape-A-Getti: a hap-

tic device for getting multiple shapes using a simple actuator. Computers and Graphics, 117(??):42– 50, December 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002510.

Brazil:2011:STD

[BMS^{+}11] Emilio Vital Brazil, Ives Macêdo, Mario Costa Sousa, Luiz Velho, and Luiz Henrique de Figueiredo. Shape and tone depiction for implicit surfaces. Computers and Graphics, 35(1):43– 53, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000169X.

Boulic:1996:RAC

[BMT96] Ronan Boulic, Ramon Mas, and Daniel Thalmann. A robust approach for the control of the center of mass with inverse kinetics. Computers and Graphics, 20(5):693–701, September– October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600043.

Barbieri:2016:IEC

[BMU^{+}16] Simone Barbieri, Pietro Meloni, Francesco Usai, L. Davide Spano, and Riccardo Scateni. An interactive editor for curve-skeletons: SkeletonLab. Computers and Graphics, 60(??):23– 33, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300905.

Baerentzen:2012:CSS

[BMW12] J. A. Bærentzen, M. K. Misztal, and K. Welnicka. Converting skeletal structures to quad dominant meshes. Computers and Graphics, 36 (5):555–561, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000556.

Boada:2003:TBH

[BN03] Imma Boada and Isabel Navazo. 3D texture-based hybrid visualizations. Computers and Graphics, 27 (1):41–49, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Baranoski:2006:NPS

[BN06] Gladimir V. G. Baranoski and Marcio Lobo Netto. Natural phenomena simulation. Computers and Graphics, 30(4):483–484, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000811.

Ball:2007:REI

[BN07] Robert Ball and Chris North. Realizing embodied interaction for visual analytics through large displays. Computers and Graphics, 31(3):380–400, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930700060X.

Bauer:2017:AAR

[BND^{+}17] Armelle Bauer, Debanga Raj Neog, Ali-Hamadi Dicko, Dinesh K. Pai, François Faure, Olivier Palombi, and Jocelyne Troccaz. Anatomical augmented reality with 3D commodity tracking and image-space alignment.

Computers and Graphics, 69(??):140–153, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317301747.

Berger:2010:FTM

[BNPS10] Matt Berger, Luis Gustavo Nonato, Valerio Pascucci, and Cláudio T. Silva. Fiedler trees for multiscale surface analysis. Computers and Graphics, 34(3): 272–281, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000488.

Black:2024:MRH

[BNS24] David Black, Mika Nogami, and Septimiu Salcudean. Mixed reality human teleoperation with device-agnostic remote ultrasound: Communication and user interaction. Computers and Graphics, 118(??):184–193, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000037.

Boardman:1978:ETC

[Boa78] D. Boardman. An experience in teaching CAD to undergraduate electronic engineering degree students.

Computers and Graphics, 3 (4):145–147, ???? 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849378900067.

Boender:1991:SIA

[Boe91] Edwin Boender. A survey of intersection algorithms for curved surfaces. Computers and Graphics, 15(1): 109–115, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Buttolo:1997:ASH

[BOH97] Pietro Buttolo, Roberto Oboe, and Blake Hannaford. Architectures for shared haptic virtual environments. Computers and Graphics, 21(4):421– 429, July–August 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=4&aid=9700019.

Bono:1995:IAE

[Bon95] Peter R. Bono. Introduction by the Associate Editor. Computers and Graphics, 19 (3):343–??, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bono:1997:IAE

[Bon97] Peter R. Bono. Introduction by the Associate Editor. Computers and Graphics, 21(5):535–??, September– October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bono:2001:E

[Bon01] Peter R. Bono. Editorial.

Computers and Graphics, 25(3):363–370, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/27/article.
pdf.

Bono:2003:AEI

[Bon03] Peter R. Bono. Associate editor’s introduction. Computers and Graphics, 27(4): 467, August 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bono:2005:AEI

[Bon05] Peter R. Bono. Associate editor’s introduction. Computers and Graphics, 29(3): 309–310, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Borges:1991:LAR

[Bor91] Rudolf Borges. Line algorithms for raster displays rescued from round-off errors. Computers and Graphics, 15(2):155–160, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bouvier:2002:PSG

[Bou02] Dennis J. Bouvier. From pixels to scene graphs in introductory computer graphics courses. Computers and Graphics, 26(4):603–608, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/40/abstract.
html.

Bourke:2006:CDL

[Bou06a] Paul Bourke. Constrained diffusion-limited aggregation in 3 dimensions. Computers and Graphics, 30 (4):646–649, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000896.

Bourke:2006:IAF

[Bou06b]  Paul Bourke.  An in-

troduction to the Apollonian fractal. Computers and Graphics, 30(1):134– 136, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002189.

Bourke:2009:ESL

[Bou09] Paul Bourke. Evaluating Second Life for the collaborative exploration of 3D fractals. Computers and Graphics, 33(1):113– 117, February 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000964.

Boulch:2020:PCC

[Bou20] Alexandre Boulch. ConvPoint:

Continuous convolutions for point cloud processing. Computers and Graphics, 88(??):24–34, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300224.

Bowman:1995:EPR

[Bow95a] Richard L. Bowman. Evaluating pseudo-random number generators. Computers and Graphics, 19(2): 315–324, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.

                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400158.

Bowman:1995:FMB

[Bow95b] Richard L. Bowman. Fractal metamorphosis: a brief student tutorial. Computers and Graphics, 19(1): 157–164, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        cag/cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1&aid=9400131.

Bao:1993:SML

[BP93] Hujun Bao and Qunsheng Peng. Shading models for linear and area light sources. Computers and Graphics, 17(2):137–145, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bao:1994:EFF

[BP94] Hujun Bao and Qunsheng Peng. An efficient formfactor evaluation algorithm for environments with curved surfaces. Computers and Graphics, 18(4):481–486, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Botsch:2008:GES

[BP08] Mario Botsch and Renato Pajarola. Guest editorial: Special section on the Symposium on Point-Based Graphics 2007. Computers and Graphics, 32(2): 187–188, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000290.

Bosch:2010:RTP

[BP10] Carles Bosch and Gustavo Patow. Real-time pathbased surface detail. Computers and Graphics, 34 (4):430–440, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000555.

Baravalle:2015:PBM

[BPD15] Rodrigo Baravalle, Gustavo Ariel Patow, and Claudio Delrieux. Procedural bread making. Computers and Graphics, 50(??): 13–24, August 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000503.

Blagojevic:2011:UDM

[BPGW11] Rachel Blagojevic, Beryl Plimmer, John Grundy, and Yong Wang. Using data mining for digital ink recognition: Dividing text and shapes in sketched diagrams. Computers and Graphics, 35 (5):976–991, October 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001312.

Bommes:2020:FSS

[BPH20] David Bommes, Nico Pietroni, and Ruizhen Hu. Foreword to the special section on Shape Modeling International 2020. Computers and Graphics, 90 (??):A4–A6, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301011.

Beacco:2015:FPM

[BPKB15] Alejandro Beacco, Nuria Pelechano, Mubbasir Kapadia, and Norman I. Badler. Footstep parameterized motion blending using barycentric coordinates. Computers and Graphics, 47(??): 105–112, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001551.

Blanchard:2007:VSH

[BPKG07] Julien Blanchard, Bruno Pinaud, Pascale Kuntz, and Fabrice Guillet. A 2D– 3D visualization support for human-centered rule mining. Computers and Graphics, 31(3):350–360, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700057X.

Barth:1990:PFH

[BPR90] W. Barth, W. Purgathofer, and Th. Rainer. Picture files for hierarchically structured pictures. Computers and Graphics, 14(2): 289–295, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bittner:2003:ERV

[BPS03] Jiří Bittner, Jan Přikryl, and Pavel Slavík. Exact regional visibility using line space partitioning. Computers and Graphics, 27 (4):569–580, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bilotta:2006:CGM

[BPS06] Eleonora Bilotta, Pietro Pantano, and Fausto Stranges. Computer graphics meets chaos and hyperchaos. Some key problems. Computers and Graphics, 30(3): 359–367, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000586.

Biasotti:2010:SAD

[BPS^{+}10] Silvia Biasotti, Giuseppe Patanè, Michela Spagnuolo, Bianca Falcidieno, and Gill Barequet. Shape approximation by differential properties of scalar functions. Computers and Graphics, 34(3):252–262, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000052X.

Bao:1989:QSL

[BR89] Paul G. Bao and Jon G. Rokne. Quadruple-step line generation. Computers and Graphics, 13(4): 461–469, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Beigl:1996:SSM

[BR96] Michael Beigl and Rimbert Rudisch. System support for mobile computing. Computers and Graphics, 20(5):619–625, September– October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=5&aid=9600033.

Barequet:2007:MAA

[BR07] Gill Barequet and Vadim Rogol. Maximizing the area of an axially symmetric polygon inscribed in a simple polygon. Computers and Graphics, 31(1):127– 136, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002238.

Bradley:1995:CEC

[Bra95] Elizabeth Bradley. Causes and effects of chaos. Computers and Graphics, 19 (5):755–778, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500057.

Berriel:2017:ALS

[BRdSOS17] Rodrigo F. Berriel, Franco Schmidt Rossi, Alberto F. de Souza, and Thiago Oliveira-Santos. Automatic large-scale data acquisition via crowdsourcing for crosswalk classification: a deep learning approach. Computers and Graphics, 68(??):32– 42, November 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301334.

Bresenham:2001:TGP

[Bre01] Jack Bresenham. Teaching the graphics processing pipeline: cosmetic and geometric attribute implications. Computers and Graphics, 25(2):343– 349, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/40/article.
                                        pdf.

Bedoucha:2020:VEL

[BRHB20] Pierre Bedoucha, Nathalie Reuter, Helwig Hauser, and Jan Byska. Visual exploration of large normal mode spaces to study protein flexibility. Computers and Graphics, 90(??): 73–83, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300790.

Briggs:1995:TMB

[Bri95] Keith Briggs. A torus map based on Jacobi’s sn. Computers and Graphics, 19(3): 451–453, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500015.

Bribiesca:2003:SCR

[Bri03] Ernesto Bribiesca. Scanningcurves representation for the coverage of surfaces using chain coding. Computers and Graphics, 27 (1):123–132, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brotz:1976:IPS

[Bro76]  D. K. Brotz.  Intersec-

tion polyhedra with successive planes. Computers and Graphics, 2(1):1–6, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brown:1979:FDW

[Bro79] C. M. Brown. Fast display of well-tessellated surfaces. Computers and Graphics, 4(2):77–85, 1979. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brodlie:1984:GCO

[Bro84] K. W. Brodlie. GKS certification — an overview. Computers and Graphics, 8 (1):13–17, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brown:1992:VLC

[Bro92] Judith R. Brown. The visualization lab: Crossroads between research and education. Computers and Graphics, 16(3):345–348, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brown:2000:EEC

[Bro00]  Judith R. Brown.  En-

abling educational collaboration a new shared reality. Computers and Graphics, 24(2):289–292, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/37/article.
                                        pdf.

Browne:2006:FBG

[Bro06a] Cameron Browne. Fractal board games. Computers and Graphics, 30 (1):126–133, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002165.

Browne:2006:WK

[Bro06b] Cameron Browne. Wild knots. Computers and Graphics, 30(6):1027–1032, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001555.

Browne:2007:EPT

[Bro07a]  Cameron Browne.  Ef-

ficient Pythagorean trees: Greed is good. Computers and Graphics, 31(4):610– 616, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000969.

Browne:2007:H

[Bro07b]  Cameron Browne.  Har-

monograms. Computers and Graphics, 31(2):292– 300, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600197X.

Browne:2007:IF

[Bro07c] Cameron Browne. Impossible fractals. Computers and Graphics, 31(4):659– 667, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000799.

Browne:2007:TVY

[Bro07d] Cameron Browne. Taiji variations: Yin and Yang in multiple dimensions. Computers and Graphics, 31 (1):142–146, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002263.

Browne:2008:GOG

[Bro08a] Cameron Browne. Gaudí’s organic geometry. Computers and Graphics, 32 (1):105–115, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001707.

Browne:2008:TCS

[Bro08b] Cameron Browne. Truchet curves and surfaces. Computers and Graphics, 32(2): 268–281, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001835.

Burska:2022:DDI

[BRO22] Karolína Dockalová Burská, Vít Rusnák, and Radek Oslejsek. Data-driven insight into the puzzle-based cybersecurity training. Computers and Graphics, 102 (??):441–451, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100203X.

Bui:2018:SDS

[BRPC18] Tu Bui, Leonardo Ribeiro, Moacir Ponti, and John Collomosse. Sketching out the details: Sketch-based image retrieval using convolutional neural networks with multi-stage regression. Computers and Graphics, 71(??):77–87, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317302194.

Ballester-Ripoll:2015:ATA

[BRSP15] Rafael Ballester-Ripoll, Susanne K. Suter, and Renato Pajarola. Analysis of tensor approximation for compression-domain volume visualization. Computers and Graphics, 47(??): 34–47, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001289.

Bruch:1975:UIC

[Bru75]  J. C. Bruch, Jr.  The

use of interactive computer graphics in the conformal mapping area. Computers and Graphics, 1(4):361–374, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brunet:1992:GEI

[Bru92] Pere Brunet. Guest Editor’s introduction. Computers and Graphics, 16 (4):353–??, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bruce:2014:EIP

[Bru14] Neil D. B. Bruce. ExpoBlend: Information preserving exposure blending based on normalized logdomain entropy. Computers and Graphics, 39(??): 12–23, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001428.

Bruckner:2010:HVC

[BRV^{+}10] Stefan Bruckner, Peter Rautek, Ivan Viola, Mike Roberts, Mario Costa Sousa, and M. Eduard Gröller. Hybrid visibility compositing and masking for illustrative rendering. Computers and Graphics, 34 (4):361–369, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000579.

Bryson:1993:VRS

[Bry93] Steve Bryson. Virtual reality in scientific visualization. Computers and Graphics, 17(6):679–685, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Biffle:1977:TDS

[BS77] J. H. Biffle and H. A. Sumlin. Three-dimensional structural analysis using interactive graphics. Computers and Graphics, 2(2):67–73, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Boursier:1982:PAC

[BS82] Patrice Boursier and Michel Scholl. Performance analysis of compaction techniques for map representation in geographic databases. Computers and Graphics, 6 (2):73–81, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Barth:1993:ERT

[BS93] W. Barth and W. Sturzlinger. Efficient ray tracing for Bézier and B-spline surfaces. Computers and Graphics, 17(4):423–430, July– August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bajaj:1998:TPD

[BS98] Chandrajit L. Bajaj and Daniel R. Schikore. Topology preserving data simplification with error bounds. Computers and Graphics, 22 (1):3–12, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/1/510.pdf.

Balsys:2001:VIS

[BS01a] Ronald J. Balsys and Kevin G. Suffern. Visualisation of implicit surfaces. Computers and Graphics, 25 (1):89–107, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/34/article.
pdf.

Bruyns:2001:ICS

[BS01b] Cynthia D. Bruyns and Steven Senger. Interactive cutting of 3D surface meshes. Computers and Graphics, 25(4):635– 642, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/37/abstract.
html.

Balfanz:2004:EVT

[BS04] D. Balfanz and D. Shrimpton. Editorial video technology and interactive broadcasting. Computers and Graphics, 28(2):137–138, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bleser:2009:ATT

[BS09] Gabriele Bleser and Didier Stricker. Advanced tracking through efficient image processing and visualinertial sensor fusion. Computers and Graphics, 33 (1):59–72, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001465.

Bartels:2011:MNS

[BS11] Richard Bartels and Faramarz Samavati. Multiresolutions numerically from subdivisions. Computers and Graphics, 35(2):185– 197, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001871.

Barth:2023:HAA

[BSAH^{+}23] Clara-Maria Barth, Jenny Schmid, Ibrahim Al-Hazwani, Madhav Sachdeva, Lena Cibulski, and Jürgen Bernard. How applicable are attributebased approaches for humancentered ranking creation? Computers and Graphics, 114(??):45–58, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000596.

Brusi:2002:TSO

[BSB^{+}02] Alex Brusi, Mateu Sbert, Philippe Bekaert, Robert Tobler, and Werner Purgathofer. Technical section: Optimal ray shooting in Monte Carlo radiosity. Computers and Graphics, 26(2):351– 354, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/40/abstract.
html.

Brossier:2023:MCE

[BSB^{+}23] Mathis Brossier, Robin Skånberg, Lonni Besançon, Mathieu Linares, Tobias Isenberg, Anders Ynnerman, and Alexander Bock. Moliverse: Contextually embedding the microcosm into the universe. Computers and Graphics, 112 (??):22–30, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000298.

Buchmuller:2021:PSC

[BSC^{+}21] Juri F. Buchmüller, Udo Schlegel, Eren Cakmak, Daniel A. Keim, and Evanthia Dimara. SpatialRugs: a compact visualization of space and time for analyzing collective movement data. Computers and Graphics, 101(??):23– 34, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001679.

Biasotti:2013:GRF

[BSF13] Silvia Biasotti, Michela Spagnuolo, and Bianca Falcidieno. Grouping real functions defined on 3D surfaces. Computers and Graphics, 37 (6):608–619, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000812.

Balfanz:2003:MSA

[BSGT03] Dirk Balfanz, Jürgen Schirmer, Matthias Grimm, and Mohammad-Reza Tazari. Mobile situation-awareness within the project map. Computers and Graphics, 27(6): 893–898, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bae:2002:PSA

[BSJC02] Seok-Hyung Bae, Hayong Shin, Won-Hyung Jung, and Byoung K. Choi. Parametricsurface adaptive tessellation based on degree reduction. Computers and Graphics, 26(5):709–719, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/34/abstract.
html.

Bordignon:2013:PBR

[BSL^{+}13] Alex Laier Bordignon, Luana Sá, Hélio Lopes, Sinésio Pesco, and Luiz Henrique de Figueiredo. Point-based rendering of implicit surfaces in R^{4}. Computers and Graphics, 37(7): 873–884, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001039.

Buck:2022:VGT

[BSM^{+}22] Valentin Buck, Flemming Stäbler, Jochen Mohrmann, Everardo González, and Jens Greinert. Visualising geospatial time series datasets in realtime with the Digital Earth Viewer. Computers and Graphics, 103 (??):121–128, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000103.

Bektas:2024:GEA

[BSMG24] Kenan Bektas, Jannis Strecker, Simon Mayer, and Kimberly Garcia. Gaze-enabled activity recognition for augmented reality feedback. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932400044X.

Bao:2010:BIV

[BSPD10] Fangxun Bao, Qinghua Sun, Jianxun Pan, and Qi Duan. A blending interpolator with value control and minimal strain energy. Computers and Graphics, 34(2): 119–124, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000178.

Bendavid:1977:CGS

[BSPR77] D. Bendavid, E. Somekh, A. Pipano, and A. Raibstein. Computer graphics in sizing and analysis of aircraft structures. Computers and Graphics, 2(2):81–89, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Brown:1978:PDN

[BSW78] M. D. Brown, S. W. Smoliar, and L. Weber. Preparing dance notation scores with a computer. Computers and Graphics, 3(1):1–7, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bonfatti:1978:DMT

[BT78] Flavio Bonfatti and Paolo Tiberio. Data management for thematic map generation. Computers and Graphics, 3(2–3):71–78, 1978. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bose:1994:GCA

[BT94] Prosenjit Bose and Godfried Toussaint. Geometric and computational aspects of manufacturing processes. Computers and Graphics, 18(4):487–497, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bruderlin:1994:PMA

[BTC94] Armin Bruderlin, Chor Guan Teo, and Tom Calvert. Procedural movement for articulated figure animation. Computers and Graphics, 18(4): 453–461, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Berretti:2022:FSS

[BTD^{+}22] Stefano Berretti, Theoharis Theoharis, Mohamed Daoudi, Claudio Ferrari, and Remco C. Veltkamp. Foreword to the special section on 3D Object Retrieval 2022 Symposium (3DOR2022). Computers and Graphics, 107(??):A3– A4, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001662.

Biasotti:2019:CAN

[BTS19] Silvia Biasotti, Elia Moscoso Thompson, and Michela Spagnuolo. Context-adaptive navigation of 3D model collections. Computers and Graphics, 79(??):1– 13, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301961.

Bonfatti:1983:NCP

[BTV83] F. Bonfatti, P. Tiberio, and A. Volta. A note on the conversion of polygonal to cellular maps. Computers and Graphics, 7(3–4): 355–360, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Busch:1997:WBT

[Bus97] Christoph Busch. Wavelet based texture segmentation of multi-modal tomographic images. Computers and Graphics, 21(3): 347–358, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=3&aid=9700012.

Busch:1998:GEI

[Bus98] Christoph Busch. Guest Editor’s introduction. Computers and Graphics, 22(4): 397–??, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bernard:2022:FSS

[BV22] Jürgen Bernard and Katerina Vrotsou. Foreword to special section on EuroVA 2021. Computers and Graphics, 107(??): A9, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001601.

Browne:2006:SP

[BvW06] Cameron Browne and Paul van Wamelen. Spiral packing. Computers and Graphics, 30(5):834–842, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001312.

Bronsvoort:1992:MCS

[BW92] Willem F. Bronsvoort and Jacob J. Waarts. A method for converting the surface of a generalized cylinder into a B-spline surface. Computers and Graphics, 16 (2):175–178, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bloor:1994:LCS

[BW94] M. I. G. Bloor and M. J. Wilson. Local control of surfaces generated using partial differential equations. Computers and Graphics, 18(2): 161–169, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bao:1998:CNL

[BW98] Paul Bao and Xiaolin Wu. L_{∞}-Constrained nearlossless image compression using weighted finite automata encoding. Computers and Graphics, 22(2–3):217– 223, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/cag/sub/1998/22/
2-3/544.pdf.

Basset:2020:CPS

[BWBM20] Jean Basset, Stefanie Wuhrer, Edmond Boyer, and Franck Multon. Contact preserving shape transfer: Retargeting motion from one shape to another. Computers and Graphics, 89(??): 11–23, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300406.

Bender:2013:FSC

[BWD13] Jan Bender, Daniel Weber, and Raphael Diziol. Fast and stable cloth simulation based on multi-resolution shape matching. Computers and Graphics, 37(8): 945–954, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001283.

Berretti:2013:MFS

[BWdBP13] Stefano Berretti, Naoufel Werghi, Alberto del Bimbo, and Pietro Pala. Matching 3D face scans using interest points and local histogram descriptors. Computers and Graphics, 37 (5):509–525, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000447.

Balreira:2018:SCI

[BWF18] Dennis G. Balreira, Marcelo Walter, and Dieter W. Fellner. A survey of the contents in introductory computer graphics courses. Computers and Graphics, 77 (??):88–96, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301560.

Bu:2024:LTM

[BWYZ24] Penghui Bu, Hang Wang, Tao Yang, and Hong Zhao. Linear time manageable edge-aware filtering on complementary tree structures. Computers and Graphics, 118(??):133–145, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784932300300X.

Bai:2018:SSC

[BWZ^{+}18] Dongdong Bai, Chaoqun Wang, Bo Zhang, Xiaodong Yi, and Xuejun Yang. Sequence searching with CNN features for robust and fast visual place recognition. Computers and Graphics, 70 (??):270–280, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301127.

Bao:1999:TSC

[BX99] Paul Bao and Dan Xu. Technical section — complex wavelet-based image mosaics using edge-preserving visual perception modeling. Computers and Graphics, 23(3):309–321, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        3/683.pdf.

Berezovsky:1988:SHI

[BY88] M. A. Berezovsky and A. K. Yablonsky. Synthesis of halftone 3D images on an array processor. Computers and Graphics, 12(3–4): 433–440, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bruder:2019:FSS

[BYC19] Gerd Bruder, Shunsuke Yoshimoto, and Sue Cobb. Foreword to the special section on the International Conference on Artificial Reality and Telexistence and Eurographics Symposium on Virtual Environments (ICAT-EGVE 2018). Computers and Graphics, 83 (??):A3–A4, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301268.

Bibissi:2022:DSI

[BYQZ22] Daryl L. Bibissi, Jiaqi Yang, Siwen Quan, and Yanning Zhang. Dual spinimage: a bi-directional spinimage variant using multiscale radii for 3D local shape description. Computers and Graphics, 103 (??):180–191, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000310.

Bai:2021:BFF

[BZYB21] Huidong Bai, Li Zhang, Jing Yang, and Mark Billinghurst. Bringing fullfeatured mobile phone interaction into virtual reality. Computers and Graphics, 97(??):42–53, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000467.

Chen:2001:DDS

[C^{+}01] Han Chen et al. Data distribution strategies for highresolution displays. Computers and Graphics, 25(5):811– 818, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/35/abstract.
html.

Cabiddu:2015:LMS

[CA15] Daniela Cabiddu and Marco Attene. Large mesh simplification for distributed environments. Computers and Graphics, 51(??):81– 89, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500062X.

Cabiddu:2017:MCD

[CA17] Daniela Cabiddu and Marco Attene. ϵ-maps: Characterizing, detecting and thickening thin features in geometric models. Computers and Graphics, 66(??): 143–153, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300638.

Celikcan:2020:DVS

[CAAC20] Ufuk Celikcan, Mehmet Bahadir Askin, Dilara Albayrak, and Tolga K. Capin. Deep into visual saliency for immersive VR environments rendered in real-time. Computers and Graphics, 88 (??):70–82, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300388.

Cebeci:2024:GDS

[CACC24] Berk Cebeci, Mehmet Bahadir Askin, Tolga K. Capin, and Ufuk Celikcan. Gazedirected and saliency-guided approaches of stereo camera control in interactive virtual reality. Computers and Graphics, 118(??):23– 32, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002492.

Cadik:2008:EEH

[Cad08] Martin Cadík. Erratum to “Evaluation of HDR tone mapping methods using essential perceptual attributes” [Comput. Graph. 32(3) (2008) 330–349]. Computers and Graphics, 32(6): 716–719, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000897.

Chuang:2004:PBG

[CAL^{+}04] Jen-Hui Chuang, Narendra Ahuja, Chien-Chou Lin, Chi-Hao Tsai, and Cheng-Hui Chen. A potentialbased generalized cylinder representation. Computers and Graphics, 28(6): 907–918, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Canright:1994:ESE

[Can94] D. Canright. Estimating the spatial extent of attractors of iterated function systems. Computers and Graphics, 18(2): 231–238, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Carson:1984:AFS

[Car84] G. S. Carson. An approach to the formal specification of computer graphics systems. Computers and Graphics, 8(1):51–57, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Caruso:1992:DPA

[Car92] Horacio A. Caruso. Dichotomous paths: An algorithm originally devised to represent botanical trees, herein used to move the arm of a robot and to draw figures with a small amount of information. Computers and Graphics, 16(2):203– 211, Summer 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Caruso:1993:AEM

[Car93] Horacio A. Caruso. The ability and efficiency of the method of dichotomous paths to simulate functions. Computers and Graphics, 17(4):439–445, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Carlson:1996:PDR

[Car96] Paul W. Carlson. Pseudo-3-D rendering methods for fractals in the complex plane. Computers and Graphics, 20(5):751–758, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600048.

Carlson:1999:TAO

[Car99a] Paul W. Carlson. Two artistic orbit trap rendering methods for Newton M-set fractals. Computers and Graphics, 23(6): 925–931, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
47/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/47/article.
pdf.

Cartwright:1999:NMF

[Car99b] Julyan H. E. Cartwright. Newton maps: fractals from Newton’s method for the circle map. Computers and Graphics, 23(4):607– 612, August 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/40/article.
                                        pdf.

Casciola:1987:IGS

[Cas87] Giulio Casciola. An interactive, graphics system for the study of functions. Computers and Graphics, 11(3): 275–280, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Casciola:1988:BCA

[Cas88] Giulio Casciola. Basic concepts to accelerate line algorithms. Computers and Graphics, 12(3–4):489–502, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Casey:1996:UDT

[Cas96] Stephen D. Casey. Using dimension theory to analyze and classify the generation of fractal sets. Computers and Graphics, 20(5):731–749, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600047.

Chapiro:2015:ADC

[CAS^{+}15] Alexandre Chapiro, Tunç Ozan Aydin, Nikolce Stefanoski, Simone Croci, Aljoscha Smolic, and Markus Gross. Art-directable Continuous Dynamic Range video. Computers and Graphics, 53 (part A)(??):54–62, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001405.

Cyrus:1978:GTT

[CB78] Mike Cyrus and Jay Beck.

Generalized two- and threedimensional clipping. Computers and Graphics, 3(1): 23–28, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:1997:DIF

[CB97] Yan Qiu Chen and Guoan Bi. 3-D IFS fractals as realtime graphics model. Computers and Graphics, 21(3): 367–370, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=3&aid=9700014.

Cunningham:2001:ELS

[CB01] Steve Cunningham and Michael J. Bailey. Education: Lessons from scene graphs: using scene graphs to teach hierarchical modeling. Computers and Graphics, 25(4):703– 711, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/33/44/abstract.
                                        html.

Cmolik:2010:LAO

[CB10] Ladislav Cmolík and Jirí Bittner. Layout-aware optimization for interactive labeling of 3D models. Computers and Graphics, 34 (4):378–387, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000695.

Cavus:2022:PPG

[ÇB22] Sencer Çavus and Mehmet Baran. Product path guiding with semi-adaptive spatiodirectional tree. Computers and Graphics, 103(??):

212–222, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002442.

Cohen:2019:GVF

[CBC19] David Cohen and Mirela Ben-Chen. Generalized volumetric foliation from inverted viscous flow. Computers and Graphics, 82 (??):152–162, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300767.

Castelo:2022:CMH

[CBG22] Antonio Castelo, Lucas Moutinho Bueno, and Marcio Gameiro. A combinatorial marching hypercubes algorithm. Computers and Graphics, 102 (??):67–77, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002405.

Cogalan:2022:LHV

[CBM^{+}22] Ugur Cogalan, Mojtaba Bemana, Karol Myszkowski, Hans-Peter Seidel, and Tobias Ritschel. Learning HDR video reconstruction for dual-exposure sensors with temporally-alternating exposures. Computers and Graphics, 105(??):57– 72, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000607.

Colombo:2023:EFR

[CBM23] Michele Colombo, Giacomo Boracchi, and Simone Melzi. Extracting a functional representation from a dictionary for non-rigid shape matching. Computers and Graphics, 113(??): 43–56, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300047X.

Christiansen:2015:CST

[CBNJ^{+}15] Asger Nyman Christiansen, J. Andreas Bærentzen, Morten Nobel-Jørgensen, Niels Aage, and Ole Sigmund. Combined shape and topology optimization of 3D structures. Computers and Graphics, 46 (??):25–35, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001095.

Crespin:2014:GME

[CBS^{+}14] Benoit Crespin, Richard Bézin, Xavier Skapin, Olivier Terraz, and Philippe Meseure. Generalized maps for erosion and sedimentation simulation. Computers and Graphics, 45(??):1–16, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000624.

Chaudhry:2015:DSD

[CBU^{+}15] E. Chaudhry, S. J. Bian, H. Ugail, Xiaogang Jin, L. H. You, and Jian J. Zhang. Dynamic skin deformation using finite difference solutions for character animation. Computers and Graphics, 46 (??):294–305, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001174.

Chaos:2001:ACD

[CC01a] Chaos and Graphics Gordon R. J. Cooper. Aspects of chaotic dynamics in the least-squares inversion of gravity data. Computers and Graphics, 25(4):691– 697, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/42/abstract.
html.

Chen:2001:NAL

[CC01b] Bei Chuan Chen and Yu-Tai Ching. A new antialiased line drawing algorithm.

Computers and Graphics, 25(2):187–193, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/27/article.
                                        pdf.

Chu:2003:BBI

[CC03] Hsueh-Ting Chu and Chaur-Chin Chen. On bounding boxes of iterated function system attractors. Computers and Graphics, 27(3): 407–414, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Castaneda:2004:DMP

[CC04] Miguel A. Padilla Castañeda and Fernando Arámbula Cosío. Deformable model of the prostate for TURP surgery simulation. Computers and Graphics, 28 (5):767–777, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2005:IOC

[CC05] Bo Chen and Harry H. Cheng. Interpretive OpenGL for computer graphics. Computers and Graphics, 29(3): 331–339, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2008:DOS

[CC08] Chun-Yen Chen and Kuo-Young Cheng. A directionoriented sharpness dependent filter for 3D polygon meshes. Computers and Graphics, 32(2):129– 140, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000307.

Campagnolo:2019:IDA

[CC19] Leonardo Q. Campagnolo and Waldemar Celes. Interactive directional ambient occlusion and shadow computations for volume ray casting. Computers and Graphics, 84(??):66– 76, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301372.

Campagnolo:2020:ESV

[CC20] Leonardo Quatrin Campagnolo and Waldemar Celes. An experimental study on volumetric visualization of black oil reservoir models. Computers and Graphics, 93 (??):84–94, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301424.

Chung:1997:EPC

[CCC97] Chung-Wen Chung, Jung-Hong Chuang, and Pei-Huan Chou. Efficient polygonization of CSG solids using boundary tracking. Computers and Graphics, 21(6):737–748, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700053.

Chung:2000:GMS

[CCC00] K. W. Chung, H. S. Y. Chan, and N. Chen. General Mandelbrot sets and Julia sets with color symmetry from equivariant mappings of the modular group. Computers and Graphics, 24 (6):911–918, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/34/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/34/34/article.
                                        pdf.

Chen:2016:ESC

[CCC^{+}16] Weifeng Chen, Wei Chen, Haidong Chen, Zhengfang Zhang, and Huamin Qu. An energy-saving color scheme for direct volume rendering. Computers and Graphics, 54 (??):57–64, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001181.

Company:2004:OBR

[CCCP04] Pedro Company, Manuel Contero, Julian Conesa, and Ana Piquer. An optimisation-based reconstruction engine for 3D modelling by sketching. Computers and Graphics, 28(6): 955–979, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Callieri:2008:MPB

[CCCS08] M. Callieri, P. Cignoni, M. Corsini, and R. Scopigno. Masked photo blending: Mapping dense photographic data set on high-resolution sampled 3D models. Computers and Graphics, 32 (4):464–473, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930800054X.

Chu:1994:ICC

[CCH94] Chyi-Hwa Chu, Ming-Guey Chern, and Yuang-Cheh Hsueh. Image compression by cardinality distribution. Computers and Graphics, 18(5):715–722, September– October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Choi:2012:CET

[CCI12] Byeongjun Choi, Byungjoon Chang, and Insung Ihm. Construction of efficient kdtrees for static scenes using voxel-visibility heuristic. Computers and Graphics, 36 (1):38–48, February 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001671.

Chaos:2001:FPT

[CCKN01] Chaos, Graphics J. M. Clausse, G. H. Kirby, and S. S. Nikiel. Fractal palettes for texturing. Computers and Graphics, 25(5):875– 881, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/40/abstract.
html.

Chang:2011:CMD

[CCKW11] Jung-Woo Chang, Yi-King Choi, Myung-Soo Kim, and Wenping Wang. Computation of the minimum distance between two Bézier curves/surfaces. Computers and Graphics, 35(3):

677–684, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000641.

Chen:2003:RCS

[CCL^{+}03] Chih-Chun Chen, Jung-Hong Chuang, Bo-Yin Lee, Wei-Wen Feng, and Ting Chiou. Rendering complex scenes using spatial subdivision and textured LOD meshes. Computers and Graphics, 27(2):189– 204, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chambel:2007:CVA

[CCM^{+}07] Teresa Chambel, Luís Correia, Jônatas Manzolli, Gonçalo Dias Miguel, Nuno A. C. Henriques, and Nuno Correia. Creating video art with evolutionary algorithms. Computers and Graphics, 31(6):837–847, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001549.

Chiang:2011:GAM

[CCM^{+}11] Patricia Chiang, Yiyu Cai, Koon Hou Mak, Ei Mon Soe, Chee Kong Chui, and Jianmin Zheng. A geometric approach to the modeling of the catheterheart interaction for VR simulation of intra-cardiac intervention. Computers and Graphics, 35(5):1013– 1022, October 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001361.

Cheng:2018:PMH

[CCM^{+}18] Zhi-Quan Cheng, Yin Chen, Ralph R. Martin, Tong Wu, and Zhan Song. Parametric modeling of 3D human body shape — a survey. Computers and Graphics, 71(??):88–100, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301929.

Chung:1997:TMG

[CCW97] K. W. Chung, H. S. Y. Chan, and B. N. Wang. Tessellations with the modular group from dynamics. Computers and Graphics, 21(4): 523–534, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700028.

Chung:1998:SSD

[CCW98] K. W. Chung, H. S. Y.

Chan, and B. N. Wang. ‘Smaller and Smaller’ from Dynamics. Computers and Graphics, 22(4):527–536, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
4/584.pdf.

Chung:1999:CGS

[CCW99] K. W. Chung, H. S. Y.

Chan, and B. N. Wang. Chaos and graphics — spiral tilings with colour symmetry from dynamics. Computers and Graphics, 23(3):439– 448, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/693.pdf.

Chung:2001:TSW

[CCW01] K. W. Chung, H. S. Y. Chan, and B. N. Wang. Tessellations with symmetries of the wallpaper groups and the modular group in the hyperbolic 3-space from dynamics. Computers and Graphics, 25(2):333–341, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
39/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/13/20/57/29/39/article.
                                        pdf.

Chen:2013:HAS

[CCW13] Kai-Chun Chen, Pei-Shan Chen, and Sai-Keung Wong. A heuristic approach to the simulation of water drops and flows on glass panes. Computers and Graphics, 37 (8):963–973, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001295.

Cai:2003:VST

[CCY^{+}03] Yiyu Cai, Cheekong Chui, Xiuzi Ye, Yaoping Wang, and James H. Anderson. VR simulated training for less invasive vascular intervention. Computers and Graphics, 27(2):215–221, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Culik:1993:BOC

[CD93] Karel Culik II and Simant Dube. Balancing order and chaos in image generation. Computers and Graphics, 17(4):465–486, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chladek:2015:PBS

[CD15] Michal Chládek and Roman Durikovic. Particlebased shallow water simulation for irregular and sparse simulation domains. Computers and Graphics, 53 (part B)(??):170–176, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000412.

Clemot:2023:NSI

[CD23] Mattéo Clémot and Julie Digne. Neural skeleton: Implicit neural representation away from the surface. Computers and Graphics, 114 (??):368–378, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001085.

Costagliola:2014:RAP

[CDF14] Gennaro Costagliola, Mattia De Rosa, and Vittorio Fuccella. Recognition and autocompletion of partially drawn symbols by using polar histograms as spatial relation descriptors. Computers and Graphics, 39 (??):101–116, April 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313002069.

Cunha:1984:DIC

[CDGA84] Geilberto J. Cunha, A. L.

Domingues, J. C. S. Gushi, and R. M. Araki. Development and implementation of CAD/CAM technology in the ITAU group. Computers and Graphics, 8(3): 255–263, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cong:1994:LBF

[CDGC94] Jason Cong, Yuzheng Ding, Tong Gao, and Kuang-Chien Chen. LUT-based FPGA technology mapping under arbitrary net-delay models. Computers and Graphics, 18(4):507–516, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Comic:2012:DIM

[CDI12] Lidija Comić, Leila De Floriani, and Federico Iuricich. Dimension-independent multiresolution Morse complexes. Computers and Graphics, 36 (5):541–547, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000490.

Comic:2016:CDM

[CDIM16] Lidija Comić, Leila De Floriani, Federico Iuricich, and Paola Magillo. Computing a discrete Morse gradient from a watershed decomposition. Computers and Graphics, 58(??):43–52, August 2016.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300632.

Callieri:2006:RIA

[CDPS06] M. Callieri, P. Debevec, J. Pair, and R. Scopigno. A realtime immersive application with realistic lighting: The Parthenon. Computers and Graphics, 30(3): 368–376, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000598.

Chaos:2001:VGF

[CDR01] Chaos, Graphics Jeffrey P.

Dumont, and Clifford A. Reiter. Visualizing generalized 3x + 1 function dynamics. Computers and Graphics, 25(5):883–898, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/41/abstract.
html.

Canino:2011:IAB

[CDW11] David Canino, Leila De Floriani, and Kenneth Weiss. IA*: an adjacency-based representation for nonmanifold simplicial shapes in arbitrary dimensions. Computers and Graphics, 35(3):

747–753, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000483.

Chin:1980:PLD

[CE80] S.-M. Chin and A. Eiger. Parking lot design by interactive computer graphics. Computers and Graphics, 5(2–4):115–121, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Caputo:2018:SPE

[CEG18] Fabio M. Caputo, Marco Emporio, and Andrea Giachetti. The Smart Pin: an effective tool for object manipulation in immersive virtual reality environments. Computers and Graphics, 74 (??):225–233, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300840.

Cossu:1989:ECF

[CEM89] R. Cossu, M. Ercoli, and L. Moltedo. An extension of CGI functions for generation and manipulation of raster images. Computers and Graphics, 13(1):39–48, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Conforti:2023:CSC

[CEN^{+}23] Pietro Manganelli Conforti, Matteo Emanuele, Pietro Nardelli, Giuseppe Santucci, and Marco An-

gelini. CryptoComparator:

Supporting cryptocurrencies analysis through visual analytics. Computers and Graphics, 114(??):276– 285, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000638.

Custodio:2013:PCM

[CEPS13] Lis Custodio, Tiago Etiene, Sinesio Pesco, and Claudio Silva. Practical considerations on Marching Cubes 33 topological correctness. Computers and Graphics, 37 (7):840–850, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000563.

Cetinaslan:2023:EES

[Cet23] Ozan Cetinaslan. ESBD: Exponential strain-based dynamics using XPBD algorithm. Computers and Graphics, 116(??):500–512, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002376.

Cheylan:1977:AEU

[CF77] J. P. Cheylan and L. Farin̆as. Analyse de l’évolution des utilisations des sols (construction de modèles explicatifs). Computers and Graphics, 2(4):255, ???? 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0097849377900280.

Costa:1996:AEB

[CF96] Mônica Costa and Bruno Feijó. Agents with emotions in behavioral animation. Computers and Graphics, 20(3):377–384, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=3&aid=9600006.

Cho:1999:IRT

[CF99] Franklin S. Cho and David Forsyth. Interactive ray tracing with the visibility complex. Computers and Graphics, 23(5):703–717, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/32/
                                        34/article.pdf;   http:
                                        //www.elsevier.nl/gej-

ng/10/13/20/24/34/34/abstract.
html.

Carmona:2011:ECR

[CF11] Rhadamés Carmona and Bernd Froehlich. Errorcontrolled real-time cut updates for multi-resolution volume rendering. Computers and Graphics, 35 (4):931–944, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000161.

Cui:2013:RTB

[CF13] Yuanmin Cui and Jieqing Feng. Real-time B-spline free-form deformation via GPU acceleration. Computers and Graphics, 37(1–2): 1–11, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001768.

Cheng:2015:PML

[CFB15] Irene Cheng, Amirhossein Firouzmanesh, and Anup Basu. Perceptually motivated LSPIHT for motion capture data compression. Computers and Graphics, 51(??):1–7, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000497.

Chai:2018:SBC

[CFH^{+}18] Shuangming Chai, Xiao-Ming Fu, Xin Hu, Yang Yang, and Ligang Liu. Sphere-based cut construction for planar parameterizations. Computers and Graphics, 74(??):66– 75, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300669.

Cugini:1984:PDR

[CFMP84] U. Cugini, G. Ferri, P. Mussio, and M. Protti. Patterndirected restoration and vectorization of digitized engineering drawings. Computers and Graphics, 8(4): 337–350, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cliburn:2002:DED

[CFMS02] Daniel C. Cliburn, Johannes J. Feddema, James R. Miller, and Terry A. Slocum. Design and evaluation of a decision support system in a water balance application. Computers and Graphics, 26 (6):931–949, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cheng:2019:PEB

[CFZC19] Xiao-Xiang Cheng, Xiao-Ming Fu, Chi Zhang, and Shuangming Chai. Practical error-bounded remeshing by adaptive refinement. Computers and Graphics, 82 (??):163–173, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300809.

Cheng:2016:VSG

[CFZL16] Xuan Cheng, Yuanli Feng, Ming Zeng, and Xinguo Liu. Video segmentation with L_{0} gradient minimization. Computers and Graphics, 54 (??):38–46, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001156.

Carey:1985:TRI

[CG85] R. J. Carey and D. P. Greenberg. Textures for realistic image synthesis. Computers and Graphics, 9(2): 125–138, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chamberlin:1987:GAS

[CG87] Donald D. Chamberlin and Charles F. Goldfarb. Graphics applications of the standard generalized markup language (SGML). Computers and Graphics, 11(4): 343–358, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Corbit:1993:CSM

[CG93] John D. Corbit and David J. Garbary. Computer simulation of the morphology and development of several species of seaweed using Lindenmayer systems. Computers and Graphics, 17(1):85– 88, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Citrin:1996:PBG

[CG96] Wayne V. Citrin and Mark D. Gross. PDA-based graphical interchange for field service and repair workers. Computers and Graphics, 20(5):641–649, September– October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=5&aid=9600038.

Canezin:2013:AIB

[CGB13] Florian Canezin, Gaël Guennebaud, and Loïc Barthe. Adequate inner bound for geometric modeling with compact field functions. Computers and Graphics, 37 (6):565–573, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300099X.

Cavalier:2019:LSN

[CGG19] Arthur Cavalier, Guillaume Gilet, and Djamchid Ghazanfarpour. Local spot noise for procedural surface details synthesis. Computers and Graphics, 85(??):92– 99, December 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301608.

Caputo:2020:PSN

[CGG^{+}20] Ariel Caputo, Andrea Giachetti, Franca Giannini, Katia Lupinetti, Marina Monti, Marco Pegoraro, and Andrea Ranieri. SFINGE 3D: a novel benchmark for online detection and recognition of heterogeneous hand gestures from 3D fingers’ trajectories. Computers and Graphics, 91 (??):232–242, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301163.

Chua:1994:MGM

[CGH94] Chee Kai Chua, Robert Gay, and Wolfgang Hoheisel. Method of generating motifs aligned along a circular arc. Computers and Graphics, 18(3):

353–362, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chua:1997:CADa

[CGH97a] Chee Kai Chua, Robert Gay, and Wolfgang Hoheisel. Computer aided decoration of ceramic tableware. Part I: 3-D decoration. Computers and Graphics, 21(5):641–653, September– October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700041.

Chua:1997:CADb

[CGH97b] Chee Kai Chua, Robert Gay, and Wolfgang Hoheisel. Computer aided decoration of ceramic tableware. Part II: rapid tooling. Computers and Graphics, 21(5):655–672, September– October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700042.

Cerimele:1991:VNS

[CGM91] M. M. Cerimele, F. R. Guarguaglini, and L. Moltedo.

Visualizations for a numerical simulation of a flame diffusion model. Computers and Graphics, 15(2): 231–235, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cignoni:2000:RTC

[CGMS00] P. Cignoni, F. Ganovelli, C. Montani, and R. Scopigno. Reconstruction of topologically correct and adaptive trilinear isosurfaces. Computers and Graphics, 24(3):399–418, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/36/article.
pdf.

Campbell-Grant:1987:IID

[CGR87] Ian R. Campbell-Grant and Peter J. Robinson. An introduction to ISO DIS, “office document architecture,” and its application to computer graphics. Computers and Graphics, 11(4): 325–341, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Carter:1998:CGF

[CGR98] Nathan C. Carter, Stephen M.

Grimes, and Clifford A. Reiter. Chaos and graphics — frieze and wallpaper chaotic attractors with a polar spin. Computers and Graphics, 22(6):765–779, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/631.pdf.

Coleman:1996:TTC

[CGS^{+}96] John Coleman, Ammo Goettsch, Andrei Savchenko, Hendrik Kollmann, Kui Wang, Edwin Klement, and Peter Bono. TeleInVivo^{T M}: towards collaborative volume visualization environments. Computers and Graphics, 20(6):801–811, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600053.

Caputo:2021:PSS

[CGS^{+}21] Ariel Caputo, Andrea Giachetti, Simone Soso, Deborah Pintani, Andrea D’Eusanio, Stefano Pini, Guido Borghi, Alessandro Simoni, Roberto Vezzani, Rita Cucchiara, Andrea Ranieri, Franca Giannini, Katia Lupinetti, Marina Monti, Mehran Maghoumi, Joseph J. LaViola, Jr., Minh-Quan Le, Hai-Dang Nguyen, and Minh-Triet Tran. SHREC 2021: Skeleton-based hand gesture recognition in the wild. Computers and Graphics, 99 (??):201–211, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001382.

Coto:2007:BBI

[CGW07] Ernesto Coto, Sören Grimm, and David Williams. Obuffer based IFT watershed from markers for large medical datasets. Computers and Graphics, 31(6): 848–863, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001641.

Chen:2016:AAO

[CGWW16] Jinming Chen, Shuming Gao, Rui Wang, and Haiyan Wu. An approach to achieving optimized complex sheet inflation under constraints. Computers and Graphics, 59 (??):39–56, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300449.

Chen:2022:PTN

[CGWZ22] Jiahao Chen, Yilin Ge, Quan Wang, and Yizhuo Zhang. TG-Net: Reconstruct visual wood texture with semantic attention. Computers and Graphics, 102 (??):546–553, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001928.

Chi:2015:RBE

[CGZZ15] Jing Chi, Shanshan Gao, Yunfeng Zhang, and Caiming Zhang. A region-based expression tracking algorithm for spacetime faces. Computers and Graphics, 51 (??):98–105, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000667.

Carlsen:1991:ICI

[CH91] Ingwer C. Carlsen and Detlef Haaks. IKSPFH — concept and implementation of an object-oriented framework for image processing. Computers and Graphics, 15(4):473–481, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Carlsson:1993:DPM

[CH93] Christer Carlsson and Olof Hagsand. DIVE — a platform for multi-user virtual environments. Computers and Graphics, 17(6):663– 669, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chu:1994:AEA

[CH94] Chyi-Hwa Chu and Yuang-Cheh Hsueh. Adaptive enhancement algorithm for quantized images. Computers and Graphics, 18(3): 335–342, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Colet:1996:WWI

[CH96] Marc Colet and Robert Herzog. WWW2GCG, a Web interface to the GCG biological sequences analysis software. Computers and Graphics, 20(3): 445–450, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600014.

Cohe:2012:BMU

[CH12] Aurélie Cohé and Martin Hachet. Beyond the mouse: Understanding user gestures for manipulating 3D objects from touchscreen inputs. Computers and Graphics, 36(8): 1119–1131, December 2012.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001562.

Collomosse:2015:FSS

[CH15] John Collomosse and Peter Hall. Foreword: Special section on visual media production. Computers and Graphics, 52(??): A2, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001090.

Chandler:1990:RTR

[Cha90] Richard E. Chandler. A recursive technique for rendering parametric curves. Computers and Graphics, 14(3– 4):477–479, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chandra:1997:TDT

[Cha97]  Satish Chandra.  A tu-

torial and diagnostic tool for chaotic oscillators and time series. Computers and Graphics, 21(2):253– 262, March–April 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.

cgi?year=1997&volume=21&
issue=2&aid=9600088.

Chen:2024:PRR

[CHC^{+}24] Yizhou Chen, Yushan Han, Jingyu Chen, Shiqian Ma, Ronald Fedkiw, and Joseph Teran. Primal residual reduction with extended position based dynamics and hyperelasticity. Computers and Graphics, 119(??): ??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000359.

Chen:2006:LAC

[Che06] Jim X. Chen. Learning abstract concepts through interactive playing. Computers and Graphics, 30 (1):10–19, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002037.

Cao:2011:ORP

[CHL^{+}11] Junjie Cao, Ying He, Zhiyang Li, Xiuping Liu, and Zhixun Su. Orienting raw point sets by global contraction and visibility voting. Computers and Graphics, 35(3):733–740, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000653.

Coeckelenbergh:1978:CDM

[CHMR78] Y. Coeckelenbergh, J. Hart, R. D. MacElroy, and Robert Rein. Computer display and manipulation of biological molecules. Computers and Graphics, 3(1):9–16, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chouraqui:1977:CE

[Cho77] Eugène Chouraqui. Le colloque EURIM 2. Computers and Graphics, 2(4): 253–254, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849377900267.

Choi:2006:ICD

[Cho06]  Kup-Sze Choi.  Interac-

tive cutting of deformable objects using force propagation approach and digital design analogy. Computers and Graphics, 30(2): 233–243, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000318.

Cani:2020:FSS

[CHPS20] Marie-Paule Cani, Edmond S. L. Ho, Tiberiu Popa, and Hubert P. H. Shum. Foreword to special section on motion, interactions and games. Computers and Graphics, 89(??): A4–A5, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300777.

Christensen:1978:NGP

[Chr78] A. H. J. Christensen. A note on geodesic polyhedra: Triangulation and contouring of spheres. Computers and Graphics, 3(4): 163–165, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2010:CBD

[CHSB10] Lu Chen, Jin Huang, Hanqiu Sun, and Hujun Bao. Cagebased deformation transfer. Computers and Graphics, 34(2):107–118, April 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000208.

Coppen:1995:DFB

[CHSD95] Derek Coppen, David Hawes, Mel Slater, and Allan Davison. A distributed frame buffer for rapid dynamic changes to 3D scenes. Computers and Graphics, 19(2): 247–250, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400150.

Chu:1990:BSI

[Chu90] Kai-Ching Chu. B3-splines for interactive curve and surface fitting. Computers and Graphics, 14(2): 281–288, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2023:PDP

[CHZ^{+}23] Jiazhou Chen, Keyu Huang, Xinding Zhu, Xianlong Qiu, Haidan Wang, and Xujia

                                        Qin.  Poetry4painting:

Diversified poetry generation for large-size ancient paintings based on data augmentation. Computers and Graphics, 116(??): 206–215, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001577.

Collomosse:2011:SSN

[CI11] John Collomosse and Tobias Isenberg. Special section on Non-Photorealistic Animation and Rendering (NPAR) 2010. Computers and Graphics, 35(1):iv– v, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001858.

Cibois:1977:ISH

[Cib77] Philippe Cibois. Informatique et sciences humaines: les enseignements d’un colloque — II. Computers and Graphics, 2(4):252– 253, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849377900255.

Chavez:1999:IAP

[CIK99] Esteban Chávez, Rüdiger Ide, and Thomas Kirste. Interactive applications of personal situation-aware assistants. Computers and Graphics, 23(6):903–915, December 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
45/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/45/article.
pdf.

Cornelis:2015:LBI

[CIPT15] Jens Cornelis, Markus Ihmsen, Andreas Peer, and Matthias Teschner. Liquid boundaries for implicit incompressible SPH. Computers and Graphics, 52 (??):72–78, November 2015.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001259.

CarrazzoneCavalcanti:2019:UET

[CIT^{+}19] Virgínia Carrazzone Cavalcanti, Maria Iziane de Santana Ferreira, Veronica Teichrieb, Ricardo Rossiter Barioni, Walter Franklin Marques Correia, and Alana Elza Fontes Da Gama. Usability and effects of text, image and audio feedback on exercise correction during augmented reality based motor rehabilitation. Computers and Graphics, 85(??): 100–110, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931930158X.

Cunningham:2009:CA

[CIW09] Douglas Cunningham, Victoria Interrante, and Brian Wyvill. Computational Aesthetics 08. Computers and Graphics, 33(4): 474, August 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000934.

Caldeira:2021:PIS

[CJAR21] Thais Caldeira, Paulo Rogério Julio, Simone Appenzeller, and Leticia Rittner. inCCsight: a software for exploration and visualization of DT-MRI data of the corpus callosum. Computers and Graphics, 99 (??):259–271, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001436.

Coors:1999:UVT

[CJJ99] Volker Coors, Uwe Jasnoch, and Volker Jung. Using the Virtual Table as an interaction platform for collaborative urban planning. Computers and Graphics, 23 (4):487–496, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/30/article.
pdf.

Casas:2019:FSS

[CJO19] Dan Casas, Adrián Jarabo, and Miguel Á. Otaduy. Foreword to the special section on CEIG 2019. Computers and Graphics, 83 (??):A7–A8, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301463.

Chen:1996:VDM

[CJT96] Min Chen, Mark W. Jones, and Peter Townsend. Volume distortion and morphing using disk fields. Computers and Graphics, 20(4): 567–575, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600027.

Chen:2023:RTS

[CJXZ23] Xiyu Chen, Tao Jia, Jiangjian Xiao, and Jiayan Zhuang. Real-time self-supervised tone curve estimation for HDR image. Computers and Graphics, 115(??):461– 471, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001632.

Collins:1975:ICG

[CK75] M. G. Collins and G. R. Kane. An interactive computer graphics approach to the design of marching band routines. Computers and Graphics, 1(4): 319–324, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849375900461.

Culik:1993:ICU

[CK93] Karel Culik II and Jarkko Kari. Image compression using weighted finite automata. Computers and Graphics, 17(3):305–313, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Culik:1996:FST

[CK96] Karel Culik II and Jarkko Kari. Finite state transformation of images. Computers and Graphics, 20(1): 125–135, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500072.

Case:2002:GEI

[CK02a] Colleen Case and Lars Kjelldahl. Guest editors’ introduction: Education. Computers and Graphics, 26(4):589–590, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/35/abstract.
html.

Choi:2002:SRP

[CK02b] Soo-Mi Choi and Myoung-Hee Kim. Shape reconstruction from partially missing data in modal space. Computers and Graphics, 26(5):701–708, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/33/abstract.
html.

Chang:2009:ETT

[CK09] Jung-Woo Chang and Myung-Soo Kim. Efficient triangletriangle intersection test for OBB-based collision detection. Computers and Graphics, 33(3):235– 240, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000430.

Choi:2009:MAL

[CKCK09] Young Choi, Ki-Youn Kwon, Soo-Won Chae, and Dong-Min Kim. A modified advancing layers mesh generation for thin threedimensional objects with variable thickness. Computers and Graphics, 33(3): 195–203, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000296.

Cremer:1996:SSS

[CKK96] James Cremer, Joseph Kearney, and Hyeongseok Ko. Simulation and scenario support for virtual environments. Computers and Graphics, 20(2):199–206, March–April 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500126.

Chambe:2023:HLI

[CKM^{+}23] Mathieu Chambe, Ewa Kijak, Zoltan Miklos, Olivier Le Meur, Rémi Cozot, and Kadi Bouatouch. HDR-LFNet: Inverse tone mapping using fusion network. Computers and Graphics, 114(??): 1–12, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000626.

Campagna:1998:GID

[CKS98] Swen Campagna, Leif Kobbelt, and Hans-Peter Seidel. Graphics in/for digital libraries

— enhancing digital documents by including 3Dmodels. Computers and Graphics, 22(6):655–666, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/cag/sub/1998/22/
                                        6/620.pdf.

Council:1992:VVR

[CL92] Edward Council and Robert Lee. Videotape from the 1990 Virtual Reality Conference. Computers and Graphics, 16(2):241–??, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chuang:1995:EGI

[CL95] Jung-Hong Chuang and Woan-Chiaun Lee. Efficient generation of isosurfaces in volume rendering. Computers and Graphics, 19(6):805–813, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=6&aid=9500051.

Chew:1996:SRA

[CL96] Hong-Gian Chew and Moung Liang. Software report: Alexis — giant jigsaw made easy. Computers and Graphics, 20(1):79–81, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.

cgi?year=1996&volume=20&
issue=1&aid=9688776.

Carrozzo:1997:GTD

[CL97] Mauro Carrozzo and Francesco Lacquaniti. Geometric transformations for displaying virtual objects on stereoscopic devices. Computers and Graphics, 21(3): 329–334, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700010.

Cant:2006:EAA

[CL06] R. J. Cant and C. S. Langensiepen. Efficient anti-aliased bump mapping. Computers and Graphics, 30(4):561– 580, August 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000744.

Choi:2018:RPS

[CL18] Min Gyu Choi and Jehee Lee. As-rigid-as-possible solid simulation with oriented particles. Computers and Graphics, 70(??): 1–7, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301206.

Claussen:1990:RPS

[Cla90]  Ute Claussen.  On re-

ducing the Phong shading method. Computers and Graphics, 14(1):73–81, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cohen:2013:VRA

[CLE13] Fernand Cohen, Zexi Liu, and Taslidere Ezgi. Virtual reconstruction of archeological vessels using expert priors and intrinsic differential geometry information. Computers and Graphics, 37(1–2):41–53, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001707.

Cai:2006:BEL

[CLF^{+}06] Yiyu Cai, Baifang Lu, Zhaowei Fan, Chandrasekaran Indhumathi, Kian Teck Lim, Ching Wern Chan, Yuan Jiang, and Lin Li. Bioedutainment: Learning life science through X gaming. Computers and Graphics, 30(1):3–9, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002013.

Chen:2012:VSS

[CLH12] Tao Chen, Aidong Lu, and Shi-Min Hu. Visual storylines: Semantic visualization of movie sequence. Computers and Graphics, 36(4):241–249, June 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000337.

Carlier:2016:SSD

[CLH^{+}16] Axel Carlier, Kathryn Leonard, Stefanie Hahmann, Geraldine Morin, and Misha Collins. The 2D shape structure dataset: a user annotated open access database. Computers and Graphics, 58(??):23–30, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300528.

Cui:2023:APF

[CLH^{+}23] Jiahao Cui, Shuai Li, Fei Hou, Aimin Hao, and Hong Qin. Analyzing part functionality via multi-modal latent space embedding and interweaving. Computers and Graphics, 115(??):1– 12, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001267.

Casti:2019:SBC

[CLM^{+}19] Sara Casti, Marco Livesu, Nicolas Mellado, Nadine Abu Rumman, Riccardo Scateni, Loïc Barthe, and Enrico Puppo. Skeleton based cage generation guided by harmonic fields. Computers and Graphics, 81(??): 140–151, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300457.

Castillo:2019:RAF

[CLMA19] Carlos Castillo, Jorge López-Moreno, and Carlos Aliaga. Recent advances in fabric appearance reproduction. Computers and Graphics, 84(??):103–121, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301256.

Chen:2016:SIC

[CLN^{+}16] Guangming Chen, Guiqing Li, Yongwei Nie, Chuhua Xian, and Aihua Mao. Stylistic indoor colour design via Bayesian network. Computers and Graphics, 60 (??):34–45, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301078.

Cheok:2007:HOA

[CLT07] Adrian David Cheok, Zheng Shawn Lim, and Roger Thomas KC Tan. Humanistic Oriental art created using automated computer processing and non-photorealistic rendering. Computers and Graphics, 31(2):280–291, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000076.

Cao:2009:SRU

[CLWQ09] Juan Cao, Xin Li, Guozhao Wang, and Hong Qin. Surface reconstruction using bivariate simplex splines on Delaunay configurations. Computers and Graphics, 33(3):341–350, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000247.

Cui:2019:LMV

[CLX^{+}19] Jiahao Cui, Shuai Li, Qing Xia, Aimin Hao, and Hong Qin. Learning multi-view manifold for single image based modeling. Computers and Graphics, 82(??): 275–285, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300925.

Cai:2022:AFR

[CLXJ22] Ruifan Cai, Honglin Li, Jun Xie, and Xiaogang Jin. Accurate floorplan reconstruction using geometric priors. Computers and Graphics, 102(??):360– 369, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002223.

Cuccu:1993:TMS

[CM93] Fabrizio Cuccu and Laura Moltedo. Texture mapping for scientific visualization environments. Computers and Graphics, 17(2): 131–136, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chatelier:2006:LCD

[CM06] Pierre Y. Chatelier and Rémy Malgouyres. A lowcomplexity discrete radiosity method. Computers and Graphics, 30(1):37– 45, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002074.

Caron:2014:TSU

[CM14] Jack Caron and David Mould. Texture synthesis using label assignment over a graph. Computers and Graphics, 39(??): 24–36, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001453.

Chen:2015:PQC

[CM15a] Xiao-Diao Chen and Weiyin Ma. A planar quadratic clipping method for computing a root of a polynomial in an interval. Computers and Graphics, 46 (??):89–98, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001022.

Chen:2015:RCC

[CM15b] Xiao-Diao Chen and Weiyin Ma. Rebuttal to “Comment on the ‘Coincidence condition of two Bézier curves of an arbitrary degree’ ”. Computers and Graphics, 53 (part B)(??): 167–169, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001685.

See comment [SR15].

Coros:2018:FSS

[CM18] Stelian Coros and Stefanie Mueller. Foreword to the special section on computational fabrication. Computers and Graphics, 75(??): A3, October 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301079.

Comic:2020:HCF

[CM20] Lidija Comić and Paola Magillo. On Hamiltonian cycles in the FCC grid. Computers and Graphics, 89(??):88–93, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300662.

Comic:2023:CFP

[CM23] Lidija Comić and Paola Magillo. Crossing-free paths in the square grid. Computers and Graphics, 114 (??):296–305, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001115.

Comninos:2010:ETA

[CMA10] P. Comninos, L. McLoughlin, and E. F. Anderson. Educating technophile artists and artophile technologists: a successful experiment in higher education. Computers and Graphics, 34(6): 780–790, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001330.

Carvalho:2017:DDL

[CMB17] Leonardo Carvalho, Ricardo Marroquim, and Emilio Vital Brazil. DiLight: Digital light table — inbetweening for 2D animations using guidelines. Computers and Graphics, 65(??): 31–44, June 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300390.

Cheverst:1999:DOM

[CMD99] Keith Cheverst, Keith Mitchell, and Nigel Davies. Design of an object model for a context sensitive tourist GUIDE. Computers and Graphics, 23 (6):883–891, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
43/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/43/article.
pdf.

Charamba:2021:CRA

[CMdL21] Luiz G. Charamba, Silvio Melo, and Ullayne de Lima. Cross ratio arrays: a descriptor invariant to severe projective deformation and robust to occlusion for planar shape recognition. Computers and Graphics, 100 (??):54–65, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001655.

Coutinho:2017:ACA

[CMDS17] Daniel Coutinho, Ricardo Marroquim, Matteo Dellepiane, and Roberto Scopigno. Assisted color acquisition for 3D models. Computers and Graphics, 68(??): 119–128, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301462.

Cai:2021:PLL

[CMLH21] Jun-Xiong Cai, Tai-Jiang Mu, Yu-Kun Lai, and Shi-Min Hu. LinkNet: 2D– 3D linked multi-modal network for online semantic segmentation of RGB-D videos. Computers and Graphics, 98(??):37–47, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000558.

Crippa:2011:GAM

[CMLR11] Alessandro Crippa, Natasha M. Maurits, Monicque M. Lorist, and Jos B. T. M. Roerdink.

Graph averaging as a means to compare multichannel EEG coherence networks and its application to the study of mental fatigue and neurodegenerative disease. Computers and Graphics, 35(2):265–274, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310002062.

Chentanez:2016:GAG

[CMM16] Nuttapong Chentanez, Matthias Müller, and Miles Macklin. GPU accelerated grid-free surface tracking. Computers and Graphics, 57(??): 1–11, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300152.

Cignoni:1998:CMS

[CMS98] P. Cignoni, C. Montani, and R. Scopigno. A comparison of mesh simplification algorithms. Computers and Graphics, 22(1):37–54, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
1/513.pdf.

Chalmers:2011:SSC

[CMS11] Alan Chalmers, Mark Mudge, and Luis Paulo Santos. Special section on cultural heritage. Computers and Graphics, 35(4): v–vi, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931100080X.

Chalmers:2012:CGJ

[CMS12] Alan Chalmers, Mark Mudge, and Luis Paulo Santos. Computers & Graphics journal special section on Cultural Heritage. Computers and Graphics, 36(1): iv, February 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931100166X.

Colom:2022:IVB

[CMS22] Arnau Colom, Ricardo Marques, and Luís Paulo Santos. Interactive VPLbased global illumination on the GPU using fuzzy clustering. Computers and Graphics, 108(??):74– 85, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001753.

Catalano:2011:SMC

[CMSF11] C. E. Catalano, M. Mortara, M. Spagnuolo, and B. Falcidieno. Semantics and 3D media: Current issues and perspectives. Computers and Graphics, 35 (4):869–877, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000793.

Codrea:2005:NAC

[CN05] Marius C. Codrea and Olli S.

Nevalainen. Note: An algorithm for contour-based region filling. Computers and Graphics, 29(3):441– 450, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2021:ERL

[CNC^{+}21] Xuechao Chen, Yu Nie, Jianwei Chen, Yanli Liu, and Yanci Zhang. Efficiently reconstruct light field probes from light-G-buffers. Computers and Graphics, 94 (??):144–151, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301837.

Castelo:2006:TAT

[CNS^{+}06] A. Castelo, L. G. Nonato, M. F. Siqueira, R. Minghim, and G. Tavares. The J_{1}^{a} triangulation: an adaptive triangulation in any dimension. Computers and Graphics, 30 (5):737–753, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001245.

Cosar:1988:SAP

[CO88] Ahmet Cosar and Bulent Ozguc. A simple animation package. Computers and Graphics, 12(3–4): 591–594, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cosar:1991:TIG

[CÖ91] Ahmet Cosar and Bülent Özgüc. A text, image, and graphics editor. Computers and Graphics, 15 (1):57–66, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cockerham:1979:FCD

[Coc79] G. Cockerham. Fitting computers into design. Computers and Graphics, 4(1):5–11, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Comninos:1985:CAI

[Com85] Peter P. Comninos. Computer animation in interior and industrial design. Computers and Graphics, 9(4): 449–453, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chiba:1994:GMH

[COM^{+}94] Norishige Chiba, Ken Ohshida, Kazunobu Muraoka, Mamoru Miura, and Nobuji Saito. A growth model having the abilities of growthregulations for simulating visual nature of botanical trees. Computers and Graphics, 18(4):469–479, July– August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Comninos:2001:IPB

[Com01] Peter Comninos. An interpolating piecewise bicubic surface with shape parameters. Computers and Graphics, 25(3):463–481, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/36/article.
pdf.

Connat:1977:PDG

[Con77] C. Connat. Presentation du GLERIAM. Computers and Graphics, 2(4):254– 255, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849377900279.

Conway:1991:CSG

[Con91] Damian Conway. Constructive solid geometry using the isoluminance contour model. Computers and Graphics, 15 (3):341–347, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Coons:1978:CLS

[Coo78] S. A. Coons. Constrained least-squares. Computers and Graphics, 3(1):43–47, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cooley:1979:CAD

[Coo79] P. Cooley. Computer-aided drafting with refresh graphics. Computers and Graphics, 4(3–4):139–147, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cooper:2000:FCP

[Coo00a] G. R. J. Cooper. Fractal convergence properties of geophysical inversion. Computers and Graphics, 24 (4):603–609, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-

ng/10/13/20/47/32/36/article.
pdf.

Cooper:2000:CBC

[Coo00b] Gordon R. J. Cooper.

Chaotic behaviour in the Carotid-Kundalini map function. Computers and Graphics, 24(3):465–470, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/41/article.
pdf.

Cooper:2001:CBC

[Coo01a] Crystal Cooper. Chaotic behavior in coupled Gierer– Meinhardt equations. Computers and Graphics, 25 (1):159–170, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/39/article.
pdf.

Cooper:2001:JSC

[Coo01b] Gordon R. J. Cooper. Julia sets of the complex Carotid– Kundalini function. Computers and Graphics, 25 (1):153–158, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/26/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/38/article.
                                        pdf.

Cooper:2002:CGD

[Coo02] Gordon R. J. Cooper. Chaos and graphics: Dynamics on the complex sphere and torus. Computers and Graphics, 26(1):151– 162, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/41/abstract.
                                        html.

Cooper:2007:NIC

[Coo07] Crystal Cooper. A note on on-off intermittency in a chaotic coin flip simulation. Computers and Graphics, 31 (1):137–141, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002251.

Conti:2017:ASP

[COPR17] Jessica Conti, Benoît Ozell, Eric Paquette, and Patrice Renaud. Adjusting stereoscopic parameters by evaluating the point of regard in a virtual environment. Computers and Graphics, 69 (??):24–35, December 2017.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301528.

Corliss:1976:CGA

[Cor76] G. Corliss. Computer graphics assisted numerical analysis instruction. Computers and Graphics, 2(1):11–13, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cortey:1982:IAS

[Cor82]  N. E. Cortey.  Interpo-

lation of arbitrary spaced points by closed surfaces. Computers and Graphics, 6 (1):19–21, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cordeiro:1984:ECA

[Cor84] Alvaro Augusto Cordeiro.

An experience of computeraided ship design systems development. Computers and Graphics, 8(3):237–245, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cortie:1993:DS

[Cor93]  M. B. Cortie.  Digi-

tal seashells. Computers and Graphics, 17(1):79–84, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cortez-Ordonez:2022:VTA

[COSEV22] Alexandra Cortez-Ordoñez, José Antonio Sanchez-Espigares, and Pere-Pau Vázquez. A visual tool for the analysis of usage trends of small and medium bicycle sharing systems. Computers and Graphics, 109(??):30– 41, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001777.

Cotton:1975:MCB

[Cot75a] I. W. Cotton. Methodologies for the cost-benefit analysis of computer graphics systems. Computers and Graphics, 1(1):33–44, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cotton:1975:SNG

[Cot75b] I. W. Cotton. Standards for network graphics communications. Computers and Graphics, 1(1):45–48, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Council:1992:ASV

[Cou92a] Edward Council. ACM/ SIGGRAPH video review issues 71 and 72. Computers and Graphics, 16(4):449– ??, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Council:1992:ACS

[Cou92b]  Edward Council.  Apple

Computer’s Siggraph Animation Reel. Computers and Graphics, 16(2):243–??, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chaturvedi:1996:PMT

[CP96] Alok K. Chaturvedi and Les A. Piegl. Procedural method for terrain surface interpolation. Computers and Graphics, 20(4): 541–566, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600026.

Clifton:1997:CPB

[CP97] Michael Clifton and Alex Pang. Cutting planes and beyond. Computers and Graphics, 21(5):563–575, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700036.

Chan:1998:NTT

[CP98] S. L. Chan and E. O. Purisima. A new tetrahedral tesselation scheme for isosurface generation. Computers and Graphics, 22(1):83–90, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        1/516.pdf.

Clement:2010:AAF

[CP10] Olivier Clément and Eric Paquette. Adaptable aging factory for multiple objects and colorations. Computers and Graphics, 34 (4):460–467, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000762.

Creus:2013:RRR

[CP13] Carles Creus and Gustavo A. Patow. R^{4}: Realistic rain rendering in realtime. Computers and Graphics, 37(1–2):33–40, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001781.

Cardoso:2019:SRL

[CP19a] Jorge C. S. Cardoso and André Perrotta. A survey of real locomotion techniques for immersive virtual reality applications on headmounted displays. Computers and Graphics, 85 (??):55–73, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301566.

Chaine:2019:FSS

[CP19b] Raphaelle Chaine and Giuseppe Patanè. Foreword to the special section on Shape Modelling International 2019. Computers and Graphics, 82 (??):A4–A6, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300998.

Cammarasana:2021:LSA

[CP21] Simone Cammarasana and Giuseppe Patané. Localised and shape-aware functions for spectral geometry processing and shape analysis: a survey and perspectives. Computers and Graphics, 97(??):1–18, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100042X.

Caputo:2018:CTS

[CPC^{+}18] Fabio M. Caputo, Pietro Prebianca, Alessandro Carcangiu, Lucio D. Spano, and Andrea Giachetti. Comparing 3D trajectories for simple mid-air gesture recognition. Computers and Graphics, 73(??):17–25, June 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300335.

Chung:1990:FBM

[CPCS90] J. C. H. Chung, D. R. Patel, R. L. Cook, and M. K. Simmons. Feature-based modeling for mechanical design. Computers and Graphics, 14(2):189–199, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chia:1994:TVC

[CPG94] L. T. Chia, D. J. Parish, and J. W. R. Griffiths. On the treatment of video cell loss in the transmission of motion-JPEG and JPEG images. Computers and Graphics, 18(1):11– 19, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Collin:2014:DOM

[CPLB14] Charly Collin, Sumanta Pattanaik, Patrick LiKamWa, and Kadi Bouatouch. Discrete ordinate method for polarized light transport solution and subsurface BRDF computation. Computers and Graphics, 45(??):17– 27, December 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400082X.

Calla:2019:MAF

[CPM19] Luciano A. Romero Calla, Lizeth J. Fuentes Perez, and Anselmo A. Montenegro. A minimalistic approach for fast computation of geodesic distances on triangular meshes. Computers and Graphics, 84 (??):77–92, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301426.

Caluya:2022:DOF

[CPS^{+}22] Nicko R. Caluya, Alexander Plopski, Christian Sandor, Yuichiro Fujimoto, Masayuki Kanbara, and Hirokazu Kato. Does overlay field of view in headmounted displays affect spatial memorization? Computers and Graphics, 102 (??):554–565, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001904.

Crosnier:1999:TST

[CR99] A. Crosnier and J. R. Rossignac. Technical section — tribox bounds for three-dimensional objects. Computers and Graphics, 23(3):429–437, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        3/704.pdf.

Coxe:2003:FHA

[CR03] Angela M. Coxe and Clifford A. Reiter. Fuzzy hexagonal automata and snowflakes. Computers and Graphics, 27(3):447– 454, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2007:BAR

[CR07] Ning Chen and Clifford A. Reiter. Best analogs for replacing missing image data. Computers and Graphics, 31 (4):617–624, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000957.

Cicirello:2013:FEA

[CR13] Vincent A. Cicirello and William C. Regli. A flexible and extensible approach to automated CAD/CAM format classification. Computers and Graphics, 37 (5):484–495, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000472.

Crawford:2002:AEC

[Cra02] Chris Crawford. Artists and engineers as cats and dogs: implications for interactive storytelling. Computers and Graphics, 26(1):13– 20, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/29/abstract.
html.

Corcoran:2010:PET

[CRD10] Andrew Corcoran, Niall Redmond, and John Dingliana. Perceptual enhancement of two-level volume rendering. Computers and Graphics, 34 (4):388–397, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000531.

Crow:1980:TMC

[Cro80]  F. C. Crow.  Toward

more complicated computer imagery. Computers and Graphics, 5(2–4):61–69, 1980. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Clarenz:2004:SPM

[CRT04] Ulrich Clarenz, Martin Rumpf, and Alexandru Telea. Surface processing methods for point sets using finite elements. Computers and Graphics, 28(6): 851–868, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cheek:1980:CMN

[CS80a] S. C. Cheek and D. Simpson. Common musical notation, computers and graphics. Computers and Graphics, 5(2–4):87–91, 1980. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Creutz:1980:ILC

[CS80b] G. Creutz and C. Schubert. An interactive line creation method using B-splines. Computers and Graphics, 5 (2–4):71–78, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Curtis:1985:DAT

[CS85] J. N. Curtis and D. H. Schwieder. Data analysis through a generalized computer animation method (DATICAM). Computers and Graphics, 9(2):153–157, 1985. CODEN COGRD2.

ISSN 0097-8493 (print), 1873-7684 (electronic).

Caudillo:1988:IGE

[CS88] R. J. Caudillo and R. Sanchez.

Interactive graphic editor for circuit design. Computers and Graphics, 12(3– 4):365–369, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chang:1998:AEA

[CS98] Chin-Chen Chang and Zen-Chung Shih. An accuracy enhancement algorithm for hierarchical radiosity. Computers and Graphics, 22(2–3):225–232, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/545.pdf.

Chaos:2001:MMD

[CS01] Chaos and Graphics Harry Seldom. From mundane to mandala: digital transformations of photographic art. Computers and Graphics, 25(4):699– 701, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/43/abstract.
html.

Cerezo:2003:ASS

[CS03] E. Cerezo and F. J. Seron. An approach to the simulation of the sea as participating medium. Computers and Graphics, 27 (4):487–501, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chorianopoulos:2004:UID

[CS04] Konstantinos Chorianopoulos and Diomidis Spinellis. User interface development for interactive television: extending a commercial DTV platform to the virtual channel API. Computers and Graphics, 28(2):157–166, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Coming:2006:KSP

[CS06] Daniel S. Coming and Oliver G. Staadt. Kinetic sweep and prune for multibody continuous motion. Computers and Graphics, 30(3):439–449, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000677.

Chen:2016:FPM

[CS16] Lujie Chen and Lawrence Sass. Fresh Press Modeler: a generative system for physically based low fidelity prototyping. Computers and Graphics, 54 (??):157–165, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001065.

Cui:2018:MAI

[CS18] Jian Cui and Alexei Sourin.

Mid-air interaction with optical tracking for 3D modeling. Computers and Graphics, 74(??):1–11, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300505.

Correa:2010:CIV

[CSC10] Carlos D. Correa, Deborah Silver, and Min Chen. Constrained illustrative volume deformation. Computers and Graphics, 34 (4):370–377, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000543.

Castello:2008:VDS

[CSCF08] P. Castelló, M. Sbert, M. Chover, and M. Feixas. Viewpoint-driven simplification using mutual information. Computers and Graphics, 32(4):451– 463, August 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000551.

Coeurjolly:2009:I

[CSD09] David Coeurjolly, Isabelle Sivignon, and Florent Dupont. Introduction. Computers and Graphics, 33(1): 1, February 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001520.

Coburn:2020:ETS

[CSF20] Joshua Coburn, John Salmon, and Ian Freeman. The effects of transition style for collaborative view sharing in immersive virtual reality. Computers and Graphics, 92 (??):44–54, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301217.

Camara:1996:SIR

[CSFG96] Gilberto Câmara, Ricardo Cartaxo Modesto Souza, Ubirajara Moura Freitas, and Juan Garrido. Spring: integrating remote sensing and GIS by object-oriented data modelling. Computers and Graphics, 20(3): 395–403, May–June 1996.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600008.

Chen:2017:LDP

[CSG^{+}17] Zhonggui Chen, Zifu Shen, Jianzhi Guo, Juan Cao, and Xiaoming Zeng. Line drawing for 3D printing. Computers and Graphics, 66 (??):85–92, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300687.

Castro:2008:ERP

[CSH08] Francesc Castro, Mateu Sbert, and John H. Halton. Efficient reuse of paths for random walk radiosity. Computers and Graphics, 32 (1):65–81, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001677.

Choi:2004:DSU

[CSHZ04] Kup-Sze Choi, Hanqiu Sun, Pheng-Ann Heng, and Jun Zou. Deformable simulation using force propagation model with finite element optimization. Computers and Graphics, 28 (4):559–568, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chheang:2021:CVR

[CSJ^{+}21] Vuthea Chheang, Patrick Saalfeld, Fabian Joeres, Christian Boedecker, Tobias Huber, Florentine Huettl, Hauke Lang, Bernhard Preim, and Christian Hansen. A collaborative virtual reality environment for liver surgery planning. Computers and Graphics, 99 (??):234–246, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001400.

Cho:1997:SNA

[CSK97] Sang-Young Cho, Phillip C-Y Sheu, and K. H. Kane Kim. A state network approach to parallel 3-D graphics. Computers and Graphics, 21(1):31– 49, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600068.

Cadik:2018:AOD

[CSL18] Martin Cadík, Daniel Sýkora, and Sungkil Lee. Automated outdoor depth-map generation and alignment. Computers and Graphics, 74 (??):109–118, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300608.

Chen:2023:TCT

[CSL23] Lifang Chen, Jianghu Su, and Shiyong Luo. TransPIFu: Combining transformer and pixel-aligned implicit function for single-view clothed human reconstruction. Computers and Graphics, 111 (??):1–13, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002321.

Chuang:2001:SMR

[CSLY01] Jen-Hui Chuang, Jin-Fa Sheu, Chien-Chou Lin, and Hui-Kuo Yang. Shape matching and recognition using a physically based object model. Computers and Graphics, 25(2): 211–222, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/29/article.
pdf.

Cheverst:2001:RSC

[CSM^{+}01] Keith Cheverst, Gareth Smith, Keith Mitchell, Adrian Friday, and Nigel Davies. The role of shared context in supporting cooperation between city visitors. Computers and Graphics, 25(4):555– 562, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/33/28/abstract.
                                        html.

Carra:2019:GBP

[CSP19] Edoardo Carra, Christian Santoni, and Fabio Pellacini. Grammar-based procedural animations for motion graphics. Computers and Graphics, 78(??):97– 107, February 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301924.

Chheang:2024:ALS

[CSS^{+}24] Vuthea Chheang, Danny Schott, Patrick Saalfeld, Lukas Vradelis, Tobias Huber, Florentine Huettl, Hauke Lang, Bernhard Preim, and Christian Hansen. Advanced liver surgery training in collaborative VR environments. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000050.

Choi:2000:EVR

[CSSC00] Jae Jeong Choi, Byeong-Seok Shin, Yeong Gil Shin, and Kevin Cleary. Efficient volumetric ray casting for isosurface rendering. Computers and Graphics, 24 (5):661–670, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/27/article.
pdf.

Cignoni:2005:SNE

[CST05] Paolo Cignoni, Roberto Scopigno, and Marco Tarini. A simple normal enhancement technique for interactive non-photorealistic renderings. Computers and Graphics, 29(1):125–133, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2019:PMS

[CSX^{+}19] Yin Chen, Zhan Song, Weiwei Xu, Ralph R. Martin, and Zhi-Quan Cheng. Parametric 3D modeling of a symmetric human body.

Computers and Graphics, 81(??):52–60, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300366.

Chao:1992:SMB

[CSZ92] Ming Chao, Yinghua Shen, and Wei Zhao. Solid modelling based on polyhedron approach. Computers and Graphics, 16(1): 101–105, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cohen:1975:ING

[CT75] D. Cohen and E. Taft. An interactive network graphics system. Computers and Graphics, 1(1):27–32, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2013:QDS

[CT13] Ming Chen and Kai Tang. Quasi-developable surface modeling of contours with curved triangular patches. Computers and Graphics, 37 (7):851–861, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000769.

Chae:2014:PBR

[CTJ^{+}14] Junghoon Chae, Dennis Thom, Yun Jang, SungYe Kim, Thomas Ertl, and David S. Ebert. Public behavior response analysis in disaster events utilizing visual analytics of microblog data. Computers and Graphics, 38(??):51– 60, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001490.

Chong:1994:GRS

[CTLG94] Michael M. S. Chong, Han Ngee Tan, Jun Liu, and Robert K. L. Gay. Geometric representation for structural analysis in image postprocessing. Computers and Graphics, 18(2): 209–218, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cerqueira:2017:NGB

[CTN^{+}17] Rômulo Cerqueira, Tiago Trocoli, Gustavo Neves, Sylvain Joyeux, Jan Albiez, and Luciano Oliveira. A novel GPU-based sonar simulator for real-time applications. Computers and Graphics, 68 (??):66–76, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301371.

Cao:2021:CSD

[CTP^{+}21] Lingxin Cao, Lihao Tian, Hao Peng, Yu Zhou, and Lin Lu. Constrained stacking in DLP 3D printing. Computers and Graphics, 95(??):60–68, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000030.

Canary:2014:VLD

[CTQ^{+}14] Hal Canary, Russell M. Taylor II, Cory Quammen, Scott Pratt, Facundo A. Gómez, Brian O’Shea, and Christopher G. Healey. Visualizing likelihood density functions via optimal region projection. Computers and Graphics, 41(??):62–71, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000296.

Cervenansky:2010:PGB

[CTS^{+}10] Michal Cervenanský, Zsolt Tóth, Juraj Starinský, Andrej Ferko, and Milos Srámek. Parallel GPUbased data-dependent triangulations. Computers and Graphics, 34(2):125– 135, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000166.

Conti:2006:VVS

[CUD06] Giuseppe Conti, Giuliana Ucelli, and Raffaele De Amicis. “Verba Volant Scripta Manent” a false axiom within virtual environments. A semi-automatic tool for retrieval of semantics understanding for speechenabled VR applications. Computers and Graphics, 30 (4):619–628, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000793.

Cullmann:1984:GPI

[Cul84] N. Cullmann. GKSGRAL — a portable implementation of the GKS standard. Computers and Graphics, 8 (3):303–307, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cummings:1989:CCE

[Cum89] J. B. Cummings. Computers, the cutting edge of learning. Computers and Graphics, 13(1):111–113, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cunningham:2000:RII

[Cun00] Steve Cunningham. Reinventing the introductory computer graphics course: providing tools for a wider audience. Computers and Graphics, 24(2): 293–296, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/38/article.
pdf.

Chouraqui:1977:CDG

[CV77] E. Chouraqui and J. Virbel. Cahiers du groupe de travail: Analyse et experimentation dans les sciences de l’homme par les méthodes informatiques. Computers and Graphics, 2(4): 249, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849377900231.

Culik:1997:FAB

[CV97] Karel Culik II and Vladimir Valenta. Finite automata based compression of bilevel and simple color images. Computers and Graphics, 21(1):61–68, January– February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600070.

Cervero:2016:MDA

[CVB16] M. Àngels Cerveró, Àlvar Vinacua, and Pere Brunet. 3D model deformations with arbitrary control points. Computers and Graphics, 57(??):92–101, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300231.

Chen:2003:ICA

[CVHM03] H. Chen, B. Vettermann, J. Hesser, and R. Männer. Innovative computer architecture for real-time volume rendering. Computers and Graphics, 27(5):715– 724, October 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cavet:2004:RCV

[CVL^{+}04] René Cavet, Stephan Volmer, Edda Leopold, Jörg Kindermann, and Gerhard Paaß. Revealing the connoted visual code: a new approach to video classification. Computers and Graphics, 28(3): 361–369, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cedrim:2016:DFQ

[CVP^{+}16] Douglas Cedrim, Viktor Vad, Afonso Paiva, M. Eduard Gröller, Luis Gustavo Nonato, and Antonio Castelo. Depth functions as a quality measure and for steering multidimensional projections. Computers and Graphics, 60 (??):93–106, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301066.

Chen:2003:IRP

[CW03] Hui Chen and Wenping Wang. On intrinsic representations of 3D polygons for shape blending. Computers and Graphics, 27 (1):133–141, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chen:2014:ISE

[CWC^{+}14] Haidong Chen, Ji Wang, Weifeng Chen, Huamin Qu, and Wei Chen. An image-space energy-saving visualization scheme for OLED displays. Computers and Graphics, 38(??): 61–68, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001611.

Coelho:2001:EMS

[CWGR01] André L. V. Coelho, Daniel Weingaertner, Ricardo R. Gudwin, and Ivan L. M. Ricarte. Emergence of multiagent spatial coordination strategies through artificial coevolution. Computers and Graphics, 25(6):1013–1023, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/34/abstract.
html.

Chen:2020:CDR

[CWL20] Mingjia Chen, Changbo Wang, and Ligang Liu. Cross-domain retrieving sketch and shape using cycle CNNs. Computers and Graphics, 89(??):50–58, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300698.

Cadik:2008:EHT

[CWNA08] Martin Cadík, Michael Wimmer, Laszlo Neumann, and Alessandro Artusi. Evaluation of HDR tone mapping methods using essential perceptual attributes. Computers and Graphics, 32(3):330–349, June 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000460.

Chang:2023:FAW

[CWT^{+}23] Jian Chang, Xiaokun Wang, Alexandru C. Telea, Jiří Kosinka, Feng Tian, Xiaojuan Ban, and Jian Jun Zhang. Foreword to AniNex workshop 2022. Computers and Graphics, 111(??): A4–A5, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000420.

Cheng:2008:IBT

[CWTL08] Chia-Ming Cheng, Shu-Fan Wang, Chin-Hung Teng, and Shang-Hong Lai. Image-based threedimensional model reconstruction for Chinese treasure — Jadeite Cabbage with Insects. Computers and Graphics, 32(6):682– 694, December 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000952.

Cao:2023:SSS

[CXCH23] Zhonghao Cao, Pengfei Xu, Zhuoyue Chen, and Hui Huang. Screen space shape manipulation by global structural optimization. Computers and Graphics, 115 (??):246–253, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001462.

Cao:2023:WNM

[CXGL23] Li Cao, Yike Xu, Jianwei Guo, and Xiaoping

Liu. WireframeNet: a novel

method for wireframe generation from point cloud. Computers and Graphics, 115 (??):226–235, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001449.

Chen:2018:OSP

[CXT18] Li Chen, Ke Xu, and Kai Tang. Optimized sequence planning for multi-axis hybrid machining of complex geometries. Computers and Graphics, 70(??): 176–187, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301115.

Chai:2020:TLM

[CXXW20] Yufei Chai, Yanning Xu, Maopu Xu, and Lu Wang. Two-layer microfacet model with diffraction. Computers and Graphics, 86(??): 71–80, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301451.

Chung:1994:FAC

[CY94] Kuo-Liang Chung and Wen-Ming Yan. Fast algorithm for cubic B-spline curve fitting. Computers and Graphics, 18(3): 327–334, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cychosz:1993:IRT

[Cyc93] Joseph M. Cychosz. An Introduction to Ray Tracing edited by Andrew S. Glassner. Computers and Graphics, 17(1):107–??, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chang:2009:SSP

[CYCL09] Hsing-Ching Chang, Chuan-Kai Yang, Jia-Wei Chiou, and Shih-Hsien Liu. Synthesizing solid particle textures via a visual hull algorithm. Computers and Graphics, 33 (5):648–658, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930900065X.

Chaudhry:2013:SMA

[CYJ^{+}13] E. Chaudhry, L. H. You, X. Jin, X. S. Yang, and J. J. Zhang. Shape modeling for animated characters using ordinary differential equations. Computers and Graphics, 37 (6):638–644, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000976.

Chung:2009:SBT

[CYKK09] Kyusik Chung, Chang-Hyo Yu, Donghyun Kim, and Lee-Sup Kim. Shader-based tessellation to save memory bandwidth in a mobile multimedia processor. Computers and Graphics, 33 (5):625–637, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000818.

Chen:2016:CCT

[CYM16] Xiao-Diao Chen, Chao Yang, and Weiyin Ma. Coincidence condition of two Bézier curves of an arbitrary degree. Computers and Graphics, 54 (??):121–126, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001168.

Cao:2015:PLG

[CYW15] Yuanhao Cao, Dong-Ming Yan, and Peter Wonka. Patch layout generation by detecting feature networks. Computers and Graphics, 46 (??):275–282, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001101.

Chen:2023:NFR

[CYWM23] Ruizhao Chen, Ran Yi, Tuanfeng Yang Wang, and Lizhuang Ma. Neural 3D face rendering conditioned on 2D appearance via GAN disentanglement method. Computers and Graphics, 116(??): 336–344, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001772.

Chen:1998:BSC

[CZ98] Ning Chen and Weiyong Zhu. Bud-sequence conjecture on M fractal image and M-J conjecture between C and Z planes from z ⟵ zw + c (w = a + ib). Computers and Graphics, 22(4):537–546, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/cag/sub/1998/22/
4/585.pdf.

Coeurjolly:2006:SMD

[CZ06] David Coeurjolly and Loutfi Zerarga. Supercover model, digital straight line recognition and curve reconstruction on the irregular isothetic grids. Computers and Graphics, 30(1):46– 53, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002086.

Chen:2022:BTS

[CZ22] Deqiang Chen and Lifeng Zhu. Branching tubular surfaces based on spherical Voronoi diagrams. Computers and Graphics, 105 (??):1–11, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000590.

Cai:2022:GBI

[CZB^{+}22] Shaoyu Cai, Lu Zhao, Yuki Ban, Takuji Narumi, Yue Liu, and Kening Zhu. GANbased image-to-friction generation for tactile simulation of fabric material. Computers and Graphics, 102 (??):460–473, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100193X.

Chen:2002:CGS

[CZC02] Ning Chen, X. L. Zhu, and K. W. Chung. Chaos and graphics: M and J sets from Newton’s transformation of the transcendental mapping F (z) = e^{zw} + c with vcps. Computers and Graphics, 26(2):371– 383, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/41/43/abstract.
                                        html.

Collins:2004:BSD

[CZCG04] Anne Collins, Afra Zomorodian, Gunnar Carlsson, and Leonidas J. Guibas. A barcode shape descriptor for curve point cloud data. Computers and Graphics, 28 (6):881–894, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Czech:1990:GOO

[Cze90] M. Czech. GKS in an objectoriented environment. Computers and Graphics, 14(3– 4):373–375, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cheng:2014:FPF

[CZL14] Xuan Cheng, Ming Zeng, and Xinguo Liu. Featurepreserving filtering with L_{0} gradient minimization. Computers and Graphics, 38 (??):150–157, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001660.

Chalmers:2022:IBI

[CZR22] Andrew Chalmers, Todd Zickler, and Taehyun Rhee. Illumination browser: an intuitive representation for radiance map databases. Computers and Graphics, 103 (??):101–108, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000061.

Chen:2022:ENB

[CZZ22] Qiaorui Chen, Shuai Zhang, and Yao Zheng. Enhanced narrow band surface reconstruction with anisotropic kernel. Computers and Graphics, 102 (??):280–288, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002326.

Dechau:2001:TCT

[D^{+}01]  J. Dechau et al.  The

Telebuddy^{T M}: collective tele-presence and tele-conversation through physical avatars.

Computers and Graphics, 25 (4):601–608, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/57/33/33/abstract.
                                        html.

deAmicis:2006:I

[dA06] Raffaele de Amicis. Introduction. Computers and Graphics, 30(3):321– 322, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000537.

Demir:2018:GPO

[DA18] Ilke Demir and Daniel G. Aliaga. Guided proceduralization: Optimizing geometry processing and grammar extraction for architectural models. Computers and Graphics, 74(??):257– 267, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300785.

Delgado:2022:HNL

[DAG22] Tiago Delgado, Tomás Alves, and Sandra Gama. How neuroticism and locus of control affect user performance in high-dimensional data visualization. Computers and Graphics, 109 (??):88–99, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001844.

Day:2004:CPS

[DAHF04] A. M. Day, D. B. Arnold, S. Havemann, and D. W. Fellner. Combining polygonal and subdivision surface approaches to modelling and rendering of urban environments. Computers and Graphics, 28 (4):497–507, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dai:1993:CAO

[Dai93]  Fan Dai.  Centralized,

application-oriented graphical interaction control using an example of planning robotic tasks. Computers and Graphics, 17(2): 155–163, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dallakyan:2000:NVM

[Dal00] Sargis Dallakyan. A note on the visualization of multiparametric bifurcations. Computers and Graphics, 24(2):269–270, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/13/20/47/27/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/34/article.
                                        pdf.

Danks:1978:ECG

[Dan78]  J. Danks.  Example in

computer graphics for engineering students. Computers and Graphics, 3(2– 3):79–84, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

dosAnjos:2018:NPD

[dAPG18] Rafael K. dos Anjos, João Pereira, and José Gaspar. A navigation paradigm driven classification for video-based rendering techniques. Computers and Graphics, 77(??): 205–216, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830178X.

Neves:2022:FRI

[dATNMC^{+}22] Tácito Trindade de Araújo Tiburtino Neves, Rafael Messias Martins, Danilo Barbosa Coimbra, Kostiantyn Kucher, Andreas Kerren, and Fernando V. Paulovich. Fast and reliable incremental dimensionality reduction for streaming data. Computers and Graphics, 102 (??):233–244, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001734.

deAguiar:2014:RMB

[dAU14] Edilson de Aguiar and Norimichi Ukita. Representing mesh-based character animations. Computers and Graphics, 38 (??):10–17, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001258.

Davidoff:1990:DF

[Dav90] Frank Davidoff. Dynamic fractals. Computers and Graphics, 14(1):135–136, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

David:1995:TFB

[Dav95] Hollister David. Two fractals based on Keplerian solids. Computers and Graphics, 19(6):885–888, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500075.

Day:1992:ACL

[Day92] J. D. Day. An algorithm for clipping lines in object and image space. Computers and Graphics, 16 (4):421–426, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dutton:1983:EIF

[DB83] R. D. Dutton and R. C. Brigham. Efficiently identifying the faces of a solid (computer graphics). Computers and Graphics, 7(2): 143–147, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Daniels:1992:LMP

[DBG92] Karen M. Daniels, R. Daniel Bergeron, and Georges G. Grinstein. Line monotonic partitioning of planar cubic B-splines. Computers and Graphics, 16(1):55–68, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dhurandhar:1993:APF

[DBG93] Sanjeev V. Dhurandhar, Virendra C. Bhavsar, and Uday G. Gujar. Analysis of Z-plane fractal images from z ← z^{α} + c for alpha < 0. Computers and Graphics, 17(1):89–94, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dafilis:2002:VCM

[DBLC02] Mathew P. Dafilis, Paul D.

Bourke, David T. J. Liley, and Peter J. Cadusch. Visualising chaos in a model of brain electrical activity. Computers and Graphics, 26 (6):971–976, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

DeLuca:2011:SBP

[DBS^{+}11] Livio De Luca, Chawee Busayarat, Chiara Stefani, Philippe Véron, and Michel Florenzano. A semanticbased platform for the digital analysis of architectural heritage. Computers and Graphics, 35(2): 227–241, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001780.

Debarba:2018:SAD

[DBS^{+}18] Henrique Galvan Debarba, Ronan Boulic, Roy Salomon, Olaf Blanke, and Bruno Herbelin. Self-attribution of distorted reaching movements in immersive virtual reality. Computers and Graphics, 76(??):142– 152, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301353.

Dunser:2012:EUH

[DBW^{+}12] Andreas Dünser, Mark Billinghurst, James Wen, Ville Lehtinen, and Antti Nurminen. Exploring the use of handheld AR for outdoor navigation. Computers and Graphics, 36(8): 1084–1095, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001586.

denBrok:2018:RMC

[dBWK18] Dennis den Brok, Michael Weinmann, and Reinhard Klein. Rapid material capture through sparse and multiplexed measurements. Computers and Graphics, 73(??):26–36, June 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300360.

deCastro:2014:ISC

[dCdLL14] Pedro Machado Manhães de Castro, Lucas Almeida Pereira de Lima, and Franklin Leandro Acioly Lucena. Invariances of single curved manifolds applied to mesh segmentation. Computers and Graphics, 38(??):399– 409, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313002057.

DeAraujo:2013:MBM

[DCJH13] Bruno R. De Araújo, Géry Casiez, Joaquim A. Jorge, and Martin Hachet. Mockup Builder: 3D modeling on and above the surface. Computers and Graphics, 37(3): 165–178, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001811.

Dong:2007:PST

[DCL07] Feng Dong, Gordon Clapworthy, and Hai Lin. Pseudo surface-texture synthesis. Computers and Graphics, 31(2):252–261, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000052.

Delanoy:2019:CVN

[DCLB19] Johanna Delanoy, David Coeurjolly, Jacques-Olivier Lachaud, and Adrien Bousseau. Combining voxel and normal predictions for multiview 3D sketching. Computers and Graphics, 82 (??):65–72, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300858.

Nascimento:2014:AIC

[dCNPdFS14] Filipe de Carvalho Nascimento, Afonso Paiva, Luiz Henrique de Figueiredo, and Jorge Stolfi. Approximating implicit curves on plane and surface triangulations with affine arithmetic. Computers and Graphics, 40(??): 36–48, May 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000144.

Darsa:1998:WCE

[DCV98] L. Darsa, B. Costa, and A. Varshney. Walkthroughs of complex environments using image-based simplification. Computers and Graphics, 22(1):55–69, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        1/514.pdf.

deJongh:1987:UIG

[dDH87] Robert de Jongh, John Duke, and Patrick Hillsdon. The use of interactive graphics in exploration and mine planning. Computers and Graphics, 11(4):475, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Danovaro:2006:LDD

[DDM^{+}06] E. Danovaro, L. De Floriani, P. Magillo, E. Puppo, and D. Sobrero. Level-of-detail for data analysis and exploration: a historical overview and some new perspectives. Computers and Graphics, 30(3):334–344, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000550.

Duan:1998:RCS

[DDPT98] Qi Duan, K. Djidjeli, W. G.

Price, and E. H. Twizell. A rational cubic spline based on function values. Computers and Graphics, 22(4):479–486, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
4/580.pdf.

Darmstaedter:1998:LCS

[DDQM98] V. Darmstaedter, J.-F. Delaigle, J. J. Quisquater, and B. Macq. Low cost spatial watermarking. Computers and Graphics, 22(4):417–424, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/cag/sub/1998/22/
                                        4/565.pdf.

Dey:2011:LCS

[DDW11] Tamal K. Dey, Ramsay Dyer, and Lei Wang. Localized Cocone surface reconstruction. Computers and Graphics, 35(3):483– 491, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000537.

DeGroot:1984:SCG

[De 84] H. De Groot. Selection criteria for graphics hardware. Computers and Graphics, 8 (3):295–301, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dzik:1992:RSV

[DE92] Steven Dzik and Jay Ezrielev. Representing surfaces with voxels. Computers and Graphics, 16(3):295–301, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Delgrande:1980:GTL

[Del80] J. P. Delgrande. A graphtheoretic language extension for an interactive computer graphics environment. Computers and Graphics, 5 (1):13–22, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dennis:1990:ORT

[Den90] Alan R. Dennis. An overview of rendering techniques. Computers and Graphics, 14 (1):101–115, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Desmontils:2000:ECS

[Des00] Emmanuel Desmontils. Expressing constraint satisfaction problems in declarative modeling using natural language and fuzzy sets. Computers and Graphics, 24 (4):555–568, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/33/article.
pdf.

DeAraujoBuck:1995:DSM

[DEST95] Thomas De Araujo Buck, Hans-Heino Ehricke, Wolfgang Strasser, and Lennart Thurfjell. 3-D segmentation of medical structures by integration of raycasting with anatomic knowledge. Computers and Graphics, 19(3): 441–449, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500014.

Denert:1975:GGL

[DEW75] E. Denert, G. Ernst, and H. Wetzel. GRAPHEX68: Graphical language features in ALGOL 68. Computers and Graphics, 1(2–3): 195–202, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

deFigueiredo:2024:VCR

[dF24] Luiz Henrique de Figueiredo. A vertex-centric representation for adaptive diamondkite meshes. Computers and Graphics, 119(??): ??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000451.

Diniz:2022:PCQ

[DFF22] Rafael Diniz, Pedro Garcia Freitas, and Mylene C. Q. Farias. Point cloud quality assessment based on geometry-aware texture descriptors. Computers and Graphics, 103(??):31– 44, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000036.

DeFloriani:1984:HSS

[DFNP84] L. De Floriani, B. Falcidieno, G. Nagy, and C. Pienovi. A hierarchical structure for surface approximation. Computers and Graphics, 8(2):183–193, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

deFigueiredo:2022:RRS

[dFP22] Luiz Henrique de Figueiredo and Afonso Paiva. Region reconstruction with the sphere-of-influence diagram. Computers and Graphics, 107(??):252–263, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001509.

Dey:2015:APM

[DFWW15] Tamal K. Dey, Bo Fu, Huamin Wang, and Lei Wang. Automatic posing of a meshed human model using point clouds. Computers and Graphics, 46 (??):14–24, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400106X.

Damodaran:1996:GEI

[DG96a] Murali Damodaran and Robert Kheng Leng Gay.

Guest Editors’ introduction. Computers and Graphics, 20 (1):7–??, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

DeFigueiredo:1996:SIO

[DG96b] Luiz Henrique De Figueiredo and Jonas Gomes. Sampling implicit objects with physically-based particle systems. Computers and Graphics, 20(3):365–375, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600005.

Demirer:1996:ATH

[DG96c] Mehmet Demirer and Richard L. Grimsdale. Approximation techniques for high performance texture mapping. Computers and Graphics, 20(4):483–490, July– August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600020.

Dischler:2001:ST

[DG01] J. M. Dischler and D. Ghazanfarpour. A survey of 3D texturing. Computers and Graphics, 25(1):135– 151, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/37/article.
pdf.

Dogrusoz:2006:MGA

[DG06] U. Dogrusoz and B. Genc.

A multi-graph approach to complexity management in interactive graph visualization. Computers and Graphics, 30(1):86– 97, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002128.

Durupinar:2007:PVK

[DG07] Funda Durupinar and Ugur Güdükbay. Procedural visualization of knitwear and woven cloth. Computers and Graphics, 31(5):778– 783, October 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001355.

Duarte:2017:RTS

[DG17] Rui P. M. Duarte and Abel J. P. Gomes. Realtime simulation of cumulus clouds through SkewT/ LogP diagrams. Computers and Graphics, 67(??): 103–114, October 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300936.

Dorner:2002:SBI

[DGA02] Ralf Dörner, Paul Grimm, and Daniel F. Abawi. Synergies between interactive training simulations and digital storytelling: a component-based framework. Computers and Graphics, 26(1):45–55, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/13/20/68/27/
                                        32/abstract.html.

Diaz-Garcia:2018:PRC

[DGBNV18] Jesús Díaz-García, Pere Brunet, Isabel Navazo, and Pere-Pau Vázquez. Progressive ray casting for volumetric models on mobile devices. Computers and Graphics, 73(??):1–16, June 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830030X.

Dubosc:2021:IAF

[DGC^{+}21] Charlotte Dubosc, Geoffrey Gorisse, Olivier Christmann, Sylvain Fleury, Killian Poinsot, and Simon Richir. Impact of avatar facial anthropomorphism on body ownership, attractiveness and social presence in collaborative tasks in immersive virtual environments. Computers and Graphics, 101(??):82– 92, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001758.

deGoes:2011:ECN

[dGGDV11] Fernando de Goes, Siome Goldenstein, Mathieu Desbrun, and Luiz Velho. Exoskeleton: Curve network abstraction for 3D shapes. Computers and Graphics, 35 (1):112–121, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001810.

deGoes:2008:SFF

[dGGV08] Fernando de Goes, Siome Goldenstein, and Luiz Velho. A simple and flexible framework to adapt dynamic meshes. Computers and Graphics, 32(2):141–148, April 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000174.

deBoer:1997:LHF

[dGHM97] M. de Boer, A. Gröpl, J. Hesser, and R. Männer. Latency- and hazard-free volume memory architecture for direct volume rendering. Computers and Graphics, 21(2):179–187, March–April 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=2&aid=9600081.

Dinev:2020:SMB

[DGKK20] Dimitar Dinev, Wenxian Guo, Petr Kadlecek, and Ladislav Kavan. Solving for muscle blending using data. Computers and Graphics, 92 (??):67–75, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301394.

Damiand:2018:HRR

[DGLRD18] Guillaume Damiand, Aldo Gonzalez-Lorenzo, Jarek Rossignac, and Florent Dupont. Hierarchical representation for rasterized planar face complexes. Computers and Graphics, 74 (??):161–170, August 2018.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300827.

Malheiros:2016:SSE

[dGMW16] Marcelo de Gomensoro Malheiros and Marcelo Walter. Spatial sorting: an efficient strategy for approximate nearest neighbor searching. Computers and Graphics, 57(??):112–126, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630019X.

Duvanenko:1993:SES

[DGR93] Victor J. Duvanenko, Ronald S.

Gyurcsik, and W. E. Robbins. Simple and efficient 2D and 3D span clipping algorithms. Computers and Graphics, 17(1):39–54, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

David:2024:STH

[DGV^{+}24] Erwan David, Jesús Gutiérrez, Melissa Lè-Hoa Võ, Antoine Coutrot, Matthieu Perreira Da Silva, and Patrick Le

Callet. The Salient360!

toolbox: Handling gaze data in 3D made easy. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000177.

deGroot:2009:LRB

[dGWvdW09] Erwin de Groot, Brian Wyvill, and Huub van de Wetering. Locally restricted blending of Blobtrees. Computers and Graphics, 33(6): 690–697, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000673.

Daehlen:1995:MDD

[DH95a] Morten Dæhlen and Per Gunnar Holm. Matrix decomposition and data reduction. Computers and Graphics, 19(2):203–214, March–April 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400144.

Doggett:1995:HAV

[DH95b] Michael C. Doggett and Graham R. Hellestrand. A hardware architecture for video rate smooth shading of volume data. Computers and Graphics, 19(5):695–704, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500048.

Dachselt:2007:TDM

[DH07] Raimund Dachselt and Anett Hübner. Threedimensional menus: a survey and taxonomy. Computers and Graphics, 31 (1):53–65, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001853.

deBoer:1997:ERT

[dHG^{+}97] M. de Boer, J. Hesser, A. Gröpl, T. Günther, C. Poliwoda, C. Reinhart, and R. Männer. Evaluation of a real-time direct volume rendering system. Computers and Graphics, 21(2): 189–198, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600082.

Dionisio:1997:VTH

[DHJ^{+}97] Jose Dionisio, Volker Henrich, Udo Jakob, Alexander Rettig, and Rolf Ziegler.

The virtual touch: Haptic interfaces in virtual environments. Computers and Graphics, 21(4):459– 468, July–August 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700029.

Dumont:1999:CAE

[DHJ^{+}99] Jeffrey P. Dumont, Flynn J. Heiss, Kevin C. Jones, Clifford A. Reiter, and Lisa M. Vislocky. Chaotic attractors and evolving planar symmetry. Computers and Graphics, 23(4):613–619, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/41/article.
pdf.

deHaan:2001:AWU

[dHT01] Geert de Haan and Jacques M. B. Terken. Agents and wearables — usability in the COMRIS system. Computers and Graphics, 25(4):571– 579, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

13/20/57/33/30/abstract.
html.

Dong:2020:VPB

[DHZL20] Zhetong Dong, Chuanfeng Hu, Chi Zhou, and Hongwei Lin. Vectorization of persistence barcode with applications in pattern classification of porous structures. Computers and Graphics, 90 (??):182–192, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300893.

Dantchev:2012:ECC

[DI12] Stefan Dantchev and Ioannis Ivrissimtzis. Efficient construction of the Cech complex. Computers and Graphics, 36(6):708– 713, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000398.

Diaz:1994:PNG

[Dia94] Bernard M. Diaz. Place for novelty in graphics and visualisation education. Computers and Graphics, 18(3): 281–286, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dror:1978:CADa

[DIE78] B. Dror, S. Isaac, and W. Eric. Computer aided design of aircraft electrical systems. Computers and Graphics, 3(1):35–42, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dixit:1991:QCI

[Dix91] Sudhir S. Dixit. Quantization of color images for display/printing on limited color output devices. Computers and Graphics, 15(4): 561–567, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Deng:2023:TDD

[DJC^{+}23] Ziyue Deng, Junfeng Jiang, Zhengming Chen, Wenxi Zhang, Qingqiang Yao, Chen Song, Yifan Sun, Zhenpei Yang, Siming Yan, Qixing Huang, and Chandrajit Bajaj. TAssembly: Datadriven fractured object assembly using a linear template model. Computers and Graphics, 113(??): 102–112, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000560.

DiGiacomo:2004:AVH

[DJG^{+}04] Thomas Di Giacomo, Chris Joslin, Stéphane Garchery, HyungSeok Kim, and Nadia Magnenat-Thalmann. Adaptation of virtual human animation and representation for MPEG. Computers and Graphics, 28 (4):485–495, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Deng:2023:SSR

[DJH^{+}23] Ziyue Deng, Junfeng Jiang, Rui Huang, Wenxi Zhang, Zhengming Chen, Kunjin He, and Qingqiang Yao. Synergistically segmenting and reducing fracture bones via whole-to-whole deep dense matching. Computers and Graphics, 116(??): 404–417, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002285.

Oliveira:2018:AAI

[dJONM18] Victor Adriel de Jesus Oliveira, Luciana Nedel, and Anderson Maciel. Assessment of an articulatory interface for tactile intercommunication in immersive virtual environments. Computers and Graphics, 76 (??):18–28, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301067.

DeMartino:1992:PRL

[DK92] Jose Mario De Martino and Rolf Köhling. Production rendering on a local area network. Computers and Graphics, 16(3): 317–324, Fall 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dyson:1997:LSF

[DK97] Mary C. Dyson and Gary J. Kipping. The legibility of screen formats: Are three columns better than one? Computers and Graphics, 21(6):703–712, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700048.

Duchowski:2024:FSS

[DK24] Andrew Duchowski and Krzysztof Krejtz. Foreword to the special section. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000475.

DiVerdi:2020:FSS

[DKFC20] Stephen DiVerdi, Craig S. Kaplan, Angus G. Forbes, and Chiara Eva Catalano. Foreword to the special section on the 8th ACM/EG Expressive symposium (Expressive 2019). Computers and Graphics, 87(??): A3–A4, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300261.

Djurcilov:2002:VSV

[DKLP02] Suzana Djurcilov, Kwansik Kim, Pierre Lermusiaux, and Alex Pang. Visualizing scalar volumetric data with uncertainty. Computers and Graphics, 26(2):239– 248, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/31/abstract.
html.

Dubey:2020:PAM

[DKM^{+}20] Rohit K. Dubey, Wei Ping Khoo, Michal Gath Morad, Christoph Hölscher, and Mubbasir Kapadia. AUTOSIGN: a multi-criteria optimization approach to computer aided design of signage layouts in complex buildings. Computers and Graphics, 88(??): 13–23, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300248.

Dvorak:2022:RPV

[DKV^{+}22] Jan Dvorák, Zuzana Kácereková, Petr Vanecek, Lukás Hruda, and Libor Vása. As-rigid-as-possible volume tracking for time-varying surfaces. Computers and Graphics, 102 (??):329–338, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002284.

Durikovic:1997:RDS

[DKY97] Roman Durikovic, Kazufumi Kaneda, and Hideo Yamashita. Reconstructing a 3-D structure with multiple deformable solid primitives. Computers and Graphics, 21(5):611–624, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700039.

Danihelka:2014:SGD

[DKZ14] Jiři Danihelka, Lukas Kencl, and Jiři Zara. Stateless generation of distributed virtual worlds. Computers and Graphics, 44(??):33– 44, November 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000636.

Dingeldein:1993:THL

[DL93] Dennis Dingeldein and Gregor Lux. THESEUS++: a high level user interface toolkit for graphical applications. Computers and Graphics, 17(2):147– 154, March–April 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dey:2009:PBH

[DL09] Tamal K. Dey and Kuiyu Li. Persistence-based handle and tunnel loops computation revisited for speed up. Computers and Graphics, 33(3):351–358, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000429.

Bicho:2012:SCB

[dLBRM^{+}12] Alessandro de Lima Bicho, Rafael Araújo Rodrigues, Soraia Raupp Musse, Cláudio Rosito Jung, Marcelo Paravisi, and Léo Pini Magalhães. Simulating crowds based on a space colonization algorithm. Computers and Graphics, 36(2): 70–79, April 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001713.

delaLosa:1999:TMV

[dlLC99] Arnaud de la Losa and Bernard Cervelle. 3D Topological modeling and visualisation for 3D GIS. Computers and Graphics, 23 (4):469–478, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/28/article.
pdf.

Duan:2018:EGS

[DLN^{+}18] Weiwei Duan, Jianxin Luo, Guiqiang Ni, Bin Tang, Qi Hu, and Yi Gao. Exclusive grouped spatial hashing. Computers and Graphics, 70 (??):71–79, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301474.

Dong:2010:CMS

[DLR^{+}10] Chu Yue (Stella) Dong, James T. Long, Clifford A. Reiter, Corey Staten, and Rytis Umbrasas. A cellular model for spatial population dynamics. Computers and Graphics, 34(2): 176–181, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000221.

Dyke:2020:PSS

[DLR^{+}20] Roberto M. Dyke, Yu-Kun Lai, Paul L. Rosin, Stefano Zappalà, Seana Dykes, Daoliang Guo, Kun Li, Riccardo Marin, Simone Melzi, and Jingyu Yang. SHREC'20: Shape correspondence with non-isometric deformations. Computers and Graphics, 92 (??):28–43, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301266.

Daniels:2011:TBQ

[DLS^{+}11] J. Daniels II, M. Lizier, M. Siqueira, C. T. Silva, and L. G. Nonato. Templatebased quadrilateral meshing. Computers and Graphics, 35(3):471–482, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931100063X.

DiGironimo:2006:CDQ

[DLV06] G. Di Gironimo, Antonio Lanzotti, and Amalia Vanacore. Concept design for quality in virtual environment. Computers and Graphics, 30(6):1011– 1019, December 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001531.

deLeeuw:2000:MLT

[dLvL00] Wim de Leeuw and Robert van Liere. Multi-level topology for flow visualization. Computers and Graphics, 24(3):325–331, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/30/28/article.
                                        pdf.

Dellinger:2023:DOS

[DLW23] Felix Dellinger, Xinye Li, and Hui Wang. Discrete orthogonal structures. Computers and Graphics, 114 (??):126–137, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000791.

Dong:2014:DSS

[DLZY14] Qi Dong, Yanli Liu, Qijun Zhao, and Hongyu Yang. Detecting soft shadows in a single outdoor image: From local edge-based models to global constraints. Computers and Graphics, 38 (??):310–319, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001842.

Delgrande:1979:ISC

[DM79] J. P. Delgrande and L. Mezei.

An interactive system for the construction and animation of systems dynamics models. Computers and Graphics, 4(3–4):155–160, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Debelov:2000:ISO

[DM00] Victor A. Debelov and Aleksandr M. Matsokin. Implementation of set operations and intersection of Bézier curves. Computers and Graphics, 24(1):53– 65, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/32/article.
pdf.

Drago:2001:VPG

[DM01] Frédéric Drago and Karol Myszkowski. Validation proposal for global illumination and rendering techniques. Computers and Graphics, 25(3):511–518, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/39/article.
                                        pdf.

deMiras:1999:TSI

[dMF99] J. Ruiz de Miras and F. R. Feito. Technical section

— inclusion test for freeform solids. Computers and Graphics, 23(2):255– 268, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        2/678.pdf.

Dischler:1999:RIC

[DMG99] J.-M. Dischler, L. Mostefaoui, and D. Ghazanfarpour. Radiosity including complex surfaces and geometric textures using solid irradiance and virtual surfaces. Computers and Graphics, 23(4):507–524, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/32/article.
                                        pdf.

Diaz:2020:IST

[DMG20] Jose Díaz, Fabio Marton, and Enrico Gobbetti. Interactive spatio-temporal exploration of massive timevarying rectilinear scalar volumes based on a variable bit-rate sparse representation over learned dictionaries. Computers and Graphics, 88(??):45–56, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300285.

delMolino:2019:PM

[dMM19] Fernando del Molino and Adolfo Muñoz. Polarization mapping. Computers and Graphics, 83(??): 42–50, October 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301050.

Dalmoro:2023:IIB

[DMM23] Bruna M. Dalmoro, Charles Monteiro, and Soraia R. Musse. Identifying influences between artists based on artwork faces and geographic proximity. Computers and Graphics, 114 (??):116–125, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000912.

Oliveira:2018:MC

[dMOHKO18] Marcelo da Mata Oliveira, Luiza A. Hagemann, Airton L. Kronbauer, and Manuel M. Oliveira. Mobile campimetry. Computers and Graphics, 76(??): 153–166, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301419.

Pinto:2008:VVEa

[dMPF08a] Francisco de M. Pinto and Carla M. D. S. Freitas. Volume visualization and exploration through flexible transfer function design. Computers and Graphics, 32 (4):420–429, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000496.

Pinto:2008:VVEb

[dMPF08b] Francisco de M. Pinto and Carla M. D. S. Freitas. Volume visualization and exploration through flexible transfer function design. Computers and Graphics, 32 (5):540–549, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001003.

Dias:2008:TMV

[DMS08] Paulo Dias, Joaquim Madeira, and Beatriz Sousa Santos.

Teaching 3D modelling and visualization using VTK. Computers and Graphics, 32(3):363–370, June 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000083.

Drakopoulos:2003:OPV

[DMT03] V. Drakopoulos, N. Mimikou, and T. Theoharis. An overview of parallel visualisation methods for Mandelbrot and Julia sets. Computers and Graphics, 27 (4):635–646, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

deMello:2021:DTL

[dMTB^{+}21] Jean Pablo Vieira de Mello, Lucas Tabelini, Rodrigo F. Berriel, Thiago M. Paixão, Alberto F. de Souza, Claudine Badue, Nicu Sebe, and Thiago Oliveira-Santos. Deep traffic light detection by overlaying synthetic context on arbitrary natural images. Computers and Graphics, 94(??):76– 86, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301461.

DeMartino:2006:FAB

[DMV06] José Mario De Martino, Léo Pini Magalhães, and Fábio Violaro. Facial animation based on contextdependent visemes. Computers and Graphics, 30(6): 971–980, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001518.

Damiand:2022:QRO

[DN22] Guillaume Damiand and Vincent Nivoliers. Queryreplace operations for topologically controlled 3D mesh editing. Computers and Graphics, 106(??):187–199, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001108.

Dodgson:2009:BES

[Dod09] N. A. Dodgson. Balancing the expected and the surprising in geometric patterns. Computers and Graphics, 33 (4):475–483, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000600.

Dohmen:1995:SCS

[Doh95] Maurice Dohmen. A survey of constraint satisfaction techniques for geometric modeling. Computers and Graphics, 19(6):831–845, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500055.

Dolenc:1995:RRP

[Dol95] A. Dolenc. Rapid recipes for parametric surface models. Computers and Graphics, 19(2):225–236, March– April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400146.

Domenech:1977:PLS

[Dom77] B. Domenech. Problèmes logico-sémantiques liés à l’expression des relations dynamiques entre la description d’un site et des exigences d’aménagement. Computers and Graphics, 2 (4):256–257, ???? 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849377900309.

Domik:1993:GCS

[Dom93] Gitta O. Domik. Guidelines for a curriculum in scientific visualization. Computers and Graphics, 17(2): 185–191, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Domik:1994:VE

[Dom94] Gitta O. Domik. Visualization education. Computers and Graphics, 18(3): 277–280, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dorr:1990:NAP

[Dor90] Michael Dorr. A new approach to parametric line clipping. Computers and Graphics, 14(3–4):449–464, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dorfmuller:1999:RTA

[Dor99]  Klaus Dorfmüller.  Ro-

bust tracking for augmented reality using retroreflective markers. Computers and Graphics, 23(6): 795–800, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/35/30/article.
                                        pdf.

dosSantos:2001:CGS

[dos01] Manuel Próspero dos Santos. Computer graphics in the scope of informatics engineering education. Computers and Graphics, 25(5):909– 915, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/43/abstract.
html.

Carretero:2005:VCF

[dPCOO^{+}05] Maria del Puy Carretero, David Oyarzun, Amalia Ortiz, Iker Aizpurua, and Jorge Posada. Virtual characters facial and body animation through the edition and interpretation of mark-up languages. Computers and Graphics, 29(2): 189–194, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Debattista:2010:PBN

[DPS10] Kurt Debattista, Alberto José Proença, and Luís Paulo Santos. Preface and biographic notes for the special issue on graphics for serious games. Computers and Graphics, 34(6): 641–642, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001494.

Damasio:2004:ITU

[DQF04] Manuel José Damásio, Célia Quico, and André Ferreira. Interactive television usage and applications: the Portuguese case-study. Computers and Graphics, 28(2): 139–148, April 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dodds:2009:UMG

[DR09] Trevor J. Dodds and Roy A. Ruddle. Using mobile group dynamics and virtual time to improve teamwork in large-scale collaborative virtual environments. Computers and Graphics, 33(2): 130–138, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000144.

Dimitrijevic:2015:ECP

[DR15] Aleksandar M. Dimitrijević and Dejan D. Rancić. Ellipsoidal clipmaps — a planet-sized terrain rendering algorithm. Computers and Graphics, 52(??):43– 61, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000916.

Drakopoulos:1998:CGA

[Dra98] V. Drakopoulos. Chaos and graphics — on the additional fixed points of Schröder iteration functions associated with a one-parameter family of cubic polynomials. Computers and Graphics, 22(5):629–634, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/605.pdf.

Debled-Rennesson:2006:OBS

[DRFRD06] I. Debled-Rennesson, F. Feschet, and J. Rouyer-Degli. Optimal blurred segments decomposition of noisy shapes in linear time. Computers and Graphics, 30(1):30– 36, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002062.

Dror:1978:CADb

[Dro78] B. Dror. Computer-aided design at Israel Aircraft Industries. Computers and Graphics, 3(2–3):93–105, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dharap:1993:MBG

[DS93] M. A. Dharap and G. R.

Shevare. Multi-block grid density control using nonrectangular topology and two stage mapping. Computers and Graphics, 17(4): 397–402, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Durikovic:2015:FSS

[DS15] Roman Durikovic and Luís Paulo Santos. Foreword to the special section on the Spring Conference on Computer Graphics 2015 (SCCG’2015). Computers and Graphics, 53 (part B)(??):A1, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000850.

Dey:2018:ECP

[DS18] Tamal K. Dey and Ryan Slechta. Edge contraction in persistence-generated discrete Morse vector fields. Computers and Graphics, 74(??):33–43, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931830061X.

Andrade:2015:PBF

[dSASP^{+}15] Luiz Fernando de Souza Andrade, Marcos Sandim, Fabiano Petronetto, Paulo Pagliosa, and Afonso Paiva. Particle-based fluids for viscous jet buckling. Computers and Graphics, 52(??): 106–115, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001247.

Dixon:1996:GGA

[DSB96] Stephen L. Dixon, Kevin L.

Steele, and Robert P. Burton. Generation and graphical analysis of Mandelbrot and Julia sets in more than four dimensions. Computers and Graphics, 20(3): 451–456, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600015.

dosSantos:2004:GUM

[dSB04] Selan dos Santos and Ken Brodlie. Gaining understanding of multivariate and multidimensional data through visualization. Computers and Graphics, 28(3): 311–325, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Caires:2007:TIF

[dSC07] Carlos Duarte de Sena Caires.

Towards the interactive filmic narrative — “Transparency”: An experimental approach. Computers and Graphics, 31(6): 800–808, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001550.

deSouza:2022:VSE

[dSdCLBC^{+}22] Carolina Veiga Ferreira de Souza, Priscila da Cunha Luz Barcellos, Lhaylla Crissaff, Marcio Cataldi, Fabio Miranda, and Marcos Lage. Visualizing simulation ensembles of extreme weather events. Computers and Graphics, 104(??):162–172, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322000073.

daSilva:2019:OOM

[dSEM19] Vinícius da Silva, Claudio Esperança, and Ricardo Marroquim. OMiCroN — oblique multipass hierarchy creation while navigating. Computers and Graphics, 84 (??):42–54, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930144X.

Dogan:2021:ACS

[DSG21] Yalim Dogan, Sinan Sonlu, and Ugur Güdükbay. An augmented crowd simulation system using automatic determination of navigable areas. Computers and Graphics, 95(??):141– 155, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000121.

Davison:2019:IEP

[DSJ19a] Timothy Davison, Faramarz Samavati, and Christian Jacob. Interactive examplepalettes for discrete element texture synthesis. Computers and Graphics, 78 (??):23–36, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301778.

Davison:2019:LPS

[DSJ19b] Timothy Davison, Faramarz Samavati, and Christian Jacob. LifeBrush: Painting, simulating, and visualizing dense biomolecular environments. Computers and Graphics, 82(??):232– 242, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300615.

daSilveira:2023:OVC

[dSJ23] Thiago L. T. da Silveira and Cláudio R. Jung. Omnidirectional visual computing: Foundations, challenges, and applications. Computers and Graphics, 113(??):89–101, June 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000663.

Junior:2018:PSO

[dSJdML18] José Guedes dos Santos Júnior and João Paulo Silva do Monte Lima. Particle swarm optimization for 3D object tracking in RGB-D images. Computers and Graphics, 76(??): 167–180, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301456.

Dixit:1999:CGI

[DSM^{+}99] Deepraj S. Dixit, Shirish H. Shanbhag, S. P. Mudur, Kurien Isaac, and Shirish Chinchalkar. Computer graphics in India — object oriented design of an interactive mechanism simulation system — Clodion. Computers and Graphics, 23(1):85–94, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/cag/sub/1999/23/
1/652.pdf.

deSouza:2023:RGA

[dSMBG23] Vinicius Luis Trevisan de Souza, Bruno Augusto Dorta Marques, Harlen Costa Batagelo, and João Paulo Gois. A review on Generative Adversarial Networks for image generation. Computers and Graphics, 114(??): 13–25, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300064X.

DeFanti:1975:CGW

[DSN75] T. A. DeFanti, D. J. Sandin, and T. H. Nelson. Computer graphics as a way of life. Computers and Graphics, 1(1):9–16, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nunes:2022:ADS

[dSNJA22] Mislene da Silva Nunes, Gastão Florêncio Miranda Junior, and Beatriz Trinchão Andrade. An appearancedriven space to create new BRDFs. Computers and Graphics, 102(??):245–256, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001722.

Doraiswamy:2013:TS

[DSNW13] Harish Doraiswamy, Nithin Shivashankar, Vijay Natarajan, and Yusu Wang. Topological saliency. Computers and Graphics, 37(7): 787–799, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000617.

Ducke:2011:MRA

[DSR11] Benjamin Ducke, David Score, and Joseph Reeves. Multiview 3D reconstruction of the archaeological site at Weymouth from image series. Computers and Graphics, 35(2):375– 382, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000070.

Demers:2015:SEC

[DTG15] Éric Demers, Christophe Tribes, and François Guibault. A selective eraser of curvature extrema for Bspline curves. Computers and Graphics, 51(??): 35–42, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000618.

Dani:1994:PRC

[DTWT94] Sandeep Dani, Joel E.

Tohline, Warren N. Waggenspack, and David E. Thompson. Parallel rendering of curvilinear volume data. Computers and Graphics, 18(3): 363–372, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Daurat:2009:AFS

[DTZ09] Alain Daurat, Mohamed Tajine, and Mahdi Zouaoui. About the frequencies of some patterns in digital planes. application to area estimators. Computers and Graphics, 33(1):11– 20, February 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001416.

Duchting:1990:TGS

[Düc90] Werner Düchting. Tumor growth simulation. Computers and Graphics, 14(3– 4):505–508, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Duchowski:2018:GBI

[Duc18] Andrew T. Duchowski.

Gaze-based interaction: A 30 year retrospective. Computers and Graphics, 73(??): 59–69, June 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300487.

Durand:1989:BMT

[Dur89] Charles X. Durand. Bit map transformations in computerized 2D animation. Computers and Graphics, 13(4): 433–440, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Durand:1991:TPR

[Dur91] Charles X. Durand. ‘TOON’ project. requirements for a computerized 2D animation system. Computers and Graphics, 15(2): 285–293, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

DeLuca:2006:REA

[DVF06] Livio De Luca, Philippe Veron, and Michel Florenzano. Reverse engineering of architectural buildings based on a hybrid modeling approach. Computers and Graphics, 30(2): 160–176, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000240.

Dagenais:2018:EVP

[DVG^{+}18] François Dagenais, Valentin Vervondel, Julián E. Guzmán, Alexander Hay, Sébastien Delorme, David Mould, and Eric Paquette. Extended virtual pipes for the stable and real-time simulation of small-scale shallow water. Computers and Graphics, 76 (??):84–95, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301341.

Diaz:2010:RTA

[DVND10] José Díaz, Pere-Pau Vázquez, Isabel Navazo, and Florent Duguet. Real-time ambient occlusion and halos with Summed Area Tables. Computers and Graphics, 34 (4):337–350, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000440.

dosSantosBrito:2018:RTB

[dVTT18] Caio José dos Santos Brito, André Luiz B. Vieira e Silva, João Marcelo Teixeira, and Veronica Teichrieb. Ray tracer based rendering solution for large scale fluid rendering. Computers and Graphics, 77(??):65– 79, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301547.

Danks:1982:ACG

[DW82] J. Danks and A. West. An application of computer graphics to the solution of an engineering design problem. Computers and Graphics, 6(1):1–5, 1982. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Devine:1989:SDS

[DW89] Michel P. Devine and Derick Wood. SEPARATION^{T M} in d dimensions or strip mining in asteroid fields. Computers and Graphics, 13(3): 329–336, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Day:2005:CTD

[DW05] A. M. Day and J. Willmott. Compound textures for dynamic impostor rendering. Computers and Graphics, 29 (1):109–124, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dey:2013:VBF

[DW13] Tamal K. Dey and Lei Wang. Voronoi-based feature curves extraction for sampled singular surfaces. Computers and Graphics, 37 (6):659–668, October 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000885.

DiVerdi:2009:AAM

[DWH09] Stephen DiVerdi, Jason Wither, and Tobias Höllerer. All around the map: Online spherical panorama construction. Computers and Graphics, 33(1):73– 84, February 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001453.

Dong:2015:APP

[DWH^{+}15] Wenqiang Dong, Fulai Wang, Yu Huang, Guangluan Xu, Zhi Guo, Xingyu Fu, and Kun Fu. An advanced prepositioning method for the force-directed graph visualization based on PageRank algorithm. Computers and Graphics, 47(??): 24–33, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001277.

Deng:2003:CRL

[DWL^{+}03] Ke Deng, Lifeng Wang, Zhouchen Lin, Tao Feng, and Zhidong Deng. Correction and rectification of light fields. Computers and Graphics, 27(2):169– 177, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ding:2023:FGV

[DWX^{+}23] Huiming Ding, Sen Wang, Zhifeng Xie, Mengtian Li, and Lizhuang Ma. A finegrained vision and language representation framework with graph-based fashion semantic knowledge. Computers and Graphics, 115 (??):216–225, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300153X.

Deng:2022:WSF

[DWZ^{+}22] An Deng, Yunchao Wu, Peng Zhang, Zhuheng Lu, Weiqing Li, and Zhiyong Su. A weakly supervised framework for real-world point cloud classification. Computers and Graphics, 102 (??):78–88, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002648.

Ding:2022:RFF

[DYW^{+}22] Hongkai Ding, Jianjun Yi, Zhuoran Wang, Jinzhen Mou, and Fei Han. Robust feature-free pose tracking and uncertainty-aware geometry reconstruction for spinning non-cooperative targets. Computers and Graphics, 102(??):30–44, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002430.

DeHaemer:1991:SOR

[DZ91] Michael J. DeHaemer, Jr.

and Michael J. Zyda. Simplification of objects rendered by polygonal approximations. Computers and Graphics, 15(2): 175–184, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

DeHaemer:1993:CGB

[DZ93] Michael J. DeHaemer and Michael J. Zyda. Computers and Graphics Best Paper Award (1991). Computers and Graphics, 17(1):1–??, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dai:2023:CNR

[DZD^{+}23] Zhenyu Dai, Junli Zhao, Xiaodan Deng, Fuqing Duan, Dantong Li, Zhenkuan Pan, and Mingquan Zhou. CR-Net: a robust craniofacial registration network by introducing Wasserstein distance constraint and geometric attention mechanism. Computers and Graphics, 116(??):194–205, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001851.

David:1979:TCA

[DZZ79] B. David, F. Zanelli, and F.-J. Z’Graggen. Teaching CAD in an architectural school with the SIGMA-ARCHI system. Computers and Graphics, 4(3–4): 193–199, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ebert:2000:PSG

[E^{+}00] David S. Ebert et al. Procedural shape generation for multi-dimensional data visualization. Computers and Graphics, 24(3):375– 384, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/30/33/article.
                                        pdf.

Erleben:2019:SIK

[EA19] Kenny Erleben and Sheldon Andrews. Solving inverse kinematics using exact Hessian matrices. Computers and Graphics, 78 (??):1–11, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301730.

Eskandar:2023:UUS

[EAAY23] George Eskandar, Mohamed Abdelsamad, Karim Armanious, and Bin Yang. USIS: Unsupervised semantic image synthesis. Computers and Graphics, 111(??): 14–23, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000018.

Eastman:1975:ESC

[Eas75] J. F. Eastman. An efficient scan conversion and hidden surface removal algorithm. Computers and Graphics, 1(2–3):215–220, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Eastman:1985:ACA

[Eas85] C. M. Eastman. Abstractions: a conceptual approach for structuring interaction with integrated CAD systems. Computers and Graphics, 9(2):97–105, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Eyiyurekli:2010:IFF

[EB10] Manolya Eyiyurekli and David Breen. Interactive free-form level-set surfaceediting operators. Computers and Graphics, 34 (5):621–638, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000907.

Entem:2015:MAS

[EBC^{+}15] Even Entem, Loic Barthe, Marie-Paule Cani, Frederic Cordier, and Michiel van de Panne. Modeling 3D animals from a side-view sketch. Computers and Graphics, 46 (??):221–230, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001253.

Eber:2000:CGC

[Ebe00] Dena Elisabeth Eber. Computer graphics curricula in the visual arts. Computers and Graphics, 24(6): 919–923, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
35/abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/13/20/47/34/35/article.
pdf.

Eichner:2014:ASB

[EBST14] Christian Eichner, Arne Bittig, Heidrun Schumann, and Christian Tominski. Analyzing simulations of biochemical systems with featurebased visual analytics. Computers and Graphics, 38 (??):18–26, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001301.

Emporio:2022:STO

[ECG^{+}22] Marco Emporio, Ariel Caputo, Andrea Giachetti, Marco Cristani, Guido Borghi, Andrea D’Eusanio, Minh-Quan Le, Hai-Dang Nguyen, Minh-Triet Tran, Felix Ambellan, Martin Hanik, Esfandiar Nava-Yazdani, and Christoph von Tycowicz. SHREC 2022 track on online detection of heterogeneous gestures. Computers and Graphics, 107(??):241–251, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001388.

Eckersley:1990:TCB

[Eck90] Michael Eckersley. Thoughts on a computer-based design apprentice. Computers and Graphics, 14(3–4): 511–517, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Edgar:2000:FFP

[Edg00] G. A. Edgar. The forest fractal puzzle. Computers and Graphics, 24 (1):133–141, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/27/38/article.
                                        pdf.

Ehricke:1994:VVV

[EDKS94] Hans-Heino Ehricke, Klaus Donner, Walter Koller, and Wolfgang Strasser. Visualization of vasculature from volume data. Computers and Graphics, 18(3): 395–406, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ehricke:1996:CGB

[EDKS96] Hans-Heino Ehricke, Klaus Donner, Walter Koller, and Wolfgang Strasser. Computers and Graphics Best Paper Award (1994). Computers and Graphics, 20(1):1– ??, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Encarnacao:2015:CGM

[EF15] José L. Encarnação and Dieter W. Fellner. Computer graphics “Made in Germany”: Darmstadt, the leading “Computer Graphics and Visual Computing Hub” in Europe: the way from 1975 to 2014. Computers and Graphics, 53 (part A) (??):13–27, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001491.

Ebert:2002:GEI

[EFP02] David S. Ebert, Jean M.

Favre, and Ronald Peikert. Guest editors’ introduction: Data visualization. Computers and Graphics, 26(2):207– 208, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/27/abstract.
html.

Eghbali:1980:IEU

[Egh80] Hassan J. Eghbali. Image enhancement using a high sequency ordered Hadamard Transform Filtering (HSHTF). Computers and Graphics, 5(1): 23–29, ???? 1980. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849380900047.

Eghbali:1983:CUA

[Egh83] H. J. Eghbali. Classification using axis symmetry feature vector developed from Hadamard transform. Computers and Graphics, 7(2): 155–160, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Egloff:1986:GCC

[Egl86a] Peter Egloff. Graphics and communication: Computer graphics in an open systems environment. Computers and Graphics, 10 (2):87–96, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Egloff:1986:GEI

[Egl86b] Peter Egloff. Guest Editor’s introduction: Graphics in networks. Computers and Graphics, 10(2): 85–86, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900312.

Erbacher:1995:EVR

[EGL^{+}95] Robert F. Erbacher, Georges Grinstein, John Peter Lee, Haim Levkowitz, Lisa Masterman, Ron Pickett, and Stuart Smith. Exploratory visualization research at the University of Massachusetts at Lowell. Computers and Graphics, 19(1): 131–139, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400128.

Elsasser:1996:ADP

[EH96] Bernhard Elsässer and Josef Hoschek. Approximation of digitized points by surfaces of revolution. Computers and Graphics, 20(1):85– 94, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500066.

Eitz:2010:EDL

[EHBA10] Mathias Eitz, Kristian Hildebrand, Tamy Boubekeur, and Marc Alexa. An evaluation of descriptors for large-scale image retrieval from sketched feature lines. Computers and Graphics, 34 (5):482–498, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001068.

Egan:1982:DMM

[EHBM82] J. T. Egan, J. Hart, S. K. Burt, and R. D. MacElroy. The display of molecular models with the Ames interactive modeling system (Aims). Computers and Graphics, 6(4):177–199, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ehlers:1985:UCH

[Ehl85] Hans Jürgen Ehlers. The use of color to help visualize information. Computers and Graphics, 9(2): 171–176, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Egan:1984:ACP

[EHM84] J. T. Egan, J. Hart, and R. D. MacElroy. Analyzing and comparing the performance of two real-time playback systems. Computers and Graphics, 8(1):67–79, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Edelsbrunner:2017:PMA

[EHSF17] Johannes Edelsbrunner, Sven Havemann, Alexei Sourin, and Dieter W. Fellner. Procedural modeling of architecture with round geometry. Computers and Graphics, 64(??):14–25, May 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300134.

Encarnacao:2015:PDI

[EJ15] José Luís Encarnação and J. Jorge. Professor Dr.- Ing. Dr.-Ing. E.h. Wolfgang Straßer (10.08.1941– 24.01.2015). Computers and Graphics, 47(??):165– 166, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000163.

Ernst:1996:HSB

[EJRW96] I. Ernst, D. Jackèl, H. Rüsseler, and O. Wittig. Hardwaresupported bump mapping. Computers and Graphics, 20(4):515–521, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600023.

Eklundh:1985:CGC

[EK85] J. O. Eklundh and L. Kjelldahl. Computer graphics and computer vision

— some unifying and discriminating features. Computers and Graphics, 9(4):

339–349, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Eicholtz:2015:CPI

[EK15] Matthew Eicholtz and Levent Burak Kara. Characterizing the performance of an image-based recognizer for planar mechanical linkages in textbook graphics and hand-drawn sketches. Computers and Graphics, 52 (??):1–17, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000874.

Eisemann:2018:FSS

[EK18] Elmar Eisemann and Paul G. Kry. Foreword to the special section on Graphics Interface 2017. Computers and Graphics, 71(??): A4–A5, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300074.

Elber:2022:SJP

[EK22] Gershon Elber and Myung-Soo Kim. Synthesis of 3D jigsaw puzzles over freeform 2-manifolds. Computers and Graphics, 102(??):339– 348, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002259.

Ehricke:2006:VMD

[EKG06] Hans-H. Ehricke, Uwe Klose, and Wolfgang Grodd. Visualizing MR diffusion tensor fields by dynamic fiber tracking and uncertainty mapping. Computers and Graphics, 30(2):255–264, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000331.

Ellzey:1993:FRI

[EKP93] Marion L. Ellzey, Jr., Vladik Kreinovich, and Julie Peña. Fast rotation of a 3D image about an arbitrary line. Computers and Graphics, 17(2):121–126, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ekman:2004:IGS

[EL04] Inger Ekman and Petri Lankoski. Integrating a game with a story—lessons from interactive television concept design. Computers and Graphics, 28(2):167– 177, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Evers:2022:MDP

[EL22] Marina Evers and Lars Linsen. Multi-dimensional parameter-space partitioning of spatio-temporal simulation ensembles. Computers and Graphics, 104 (??):140–151, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000577.

Elareef:1986:FSM

[Ela86] Taha I. Elareef. Flavor system and message passing as representation of knowledge for solid modeling in CAD expert system. Computers and Graphics, 10(4): 351–357, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Elber:2011:MSI

[Elb11]  Gershon Elber.  Model-

ing (seemingly) impossible models. Computers and Graphics, 35(3):632– 638, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000549.

Elber:2022:SDM

[Elb22] Gershon Elber. Simultaneous 3D dithering of multiple images by curves. Computers and Graphics, 105 (??):146–152, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000784.

Edmunds:2012:SBF

[ELC^{+}12] Matt Edmunds, Robert S.

Laramee, Guoning Chen, Nelson Max, Eugene Zhang, and Colin Ware. Surfacebased flow visualization. Computers and Graphics, 36 (8):974–990, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001355.

Elvins:1996:VVC

[Elv96]  T. Todd Elvins.  Vol-

ume visualization in a collaborative computing environment. Computers and Graphics, 20(2):219–222, March–April 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500128.

Encarnacao:1998:CSA

[EMB^{+}98] J. Encarnação, M. Mengel, P. Bono, K. Böhm, E. Borgmeier, J. Brisson-Lopes, C. Hornung, A. Knierriem-Jasnoch, E. Koch, D. Krömer, R. Lindner, C. Paris, A. Sandberg, M. Schnaider, D. Storck, J. Teixeira, B. Urban, and T. Wang. A concept and system architecture for IT-based Lifelong Learning. Computers and Graphics, 22(2–3):319–393, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/578.pdf.

Ezair:2015:OAO

[EME15] Ben Ezair, Fady Massarwi, and Gershon Elber. Orientation analysis of 3D objects toward minimal support volume in 3D-printing. Computers and Graphics, 51 (??):117–124, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000564.

Encarnacao:1983:ECE

[Enc83a]  J. Encarnação.  Editor-

in-Chief’s editorial. Computers and Graphics, 7(1): 1–2, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849383900389.

Encarnacao:1983:GG

[Enc83b]  J. Encarnação.  Graph-

ics glossary. Computers and Graphics, 7(1): 81, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784938390047X.

Encarnacao:1983:I

[Enc83c] J. Encarnação. Introduction. Computers and Graphics, 7(1):93, ???? 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900511.

Encarnacao:1983:CP

[Enc83d] JoséL. Encarnação. Call for papers. Computers and Graphics, 7(3–4):385– 386, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900353.

Encarnacao:1984:I

[Enc84a] J. Encarnação. Introduction. Computers and Graphics, 8(2):113–114, ???? 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784938490044X.

Encarnacao:1984:CP

[Enc84b] José L. Encarnação. Call for paper. Computers and Graphics, 8(2):221– 222, ???? 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/009784938490061X.

Encarnacao:1985:CCNa

[Enc85a] J. L. Encarnação. Call for contributions to the new software survey section to Computers & Graphics. Computers and Graphics, 9(3):i–iii, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900639.

Encarnacao:1985:CCNb

[Enc85b] J. L. Encarnação. Call for contributions to the new software survey section to Computers & Graphics. Computers and Graphics, 9(4):ix–xi, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900445.

Encarnacao:1986:CCNa

[Enc86a] J. L. Encarnação. Call for contributions to the new software survey section to Computers & Graphics. Computers and Graphics, 10(1):i–iii, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900737.

Encarnacao:1986:CCNb

[Enc86b] J. L. Encarnação. Call for contributions to the new software survey section to Computers & Graphics. Computers and Graphics, 10(4):i–iii, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900300.

Encarnacao:1986:EI

[Enc86c] J. L. Encarnação. Editor’s introduction. Computers and Graphics, 10 (1):iii, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900634.

Encarnacao:1993:EEC

[Enc93] Jose L. Encarnação. Editorial by the Editor-in-Chief. Computers and Graphics, 17(5):505–??, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Encarnacao:1995:MWK

[Enc95a] J. Encarnação. In memoriam: Wolfgang Kruger 1942–1995. Computers and Graphics, 19(6):III–??, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Encarnacao:1995:AGU

[Enc95b] L. Miguel Encarnação. Adaptivity in graphical user interfaces: An experimental framework. Computers and Graphics, 19(6):873–884, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500074.

Encarnacao:2007:E

[Enc07] José L. Encarnação. Editorial. Computers and Graphics, 31(3):313, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000805.

Encarnacao:2015:GEF

[Enc15a] José Luis Encarnação. Guest Editor foreword. Computers and Graphics, 53 (part A) (??):1–11, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001508.

Encarnacao:2015:MPD

[Enc15b] José Luis Encarnação. In memoriam Professor Dr.- Ing. Dr.-Ing. E.h. Wolfgang Straßer. Computers and Graphics, 53 (part A)(??): 12, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001478.

Enderle:1983:CS

[End83] G. Enderle. Core systems.

Computers and Graphics, 7 (1):87–90, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Enderle:1985:GEI

[End85] Günter Enderle. Guest Editor’s introduction: Computer graphics standards. Computers and Graphics, 9(1):1–8, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Engell-Norregaard:2011:PBT

[ENE11] Morten Engell-Nørregård and Kenny Erleben. A projected back-tracking linesearch for constrained interactive inverse kinematics. Computers and Graphics, 35(2):288–298, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310002098.

Engel:1993:MJS

[Eng93] Alejandro B. Engel. Morphosis of the Julia set of the real parameter family of complex quadratic maps. Computers and Graphics, 17(3): 315–319, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Entwisle:1977:IPA

[Ent77] Jeffrey Entwisle. An image processing approach to computer graphics. Computers and Graphics, 2(2): 111–117, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Entwistle:1989:JSA

[Ent89a] Ian D. Entwistle. Julia set art and fractals in the complex plane. Computers and Graphics, 13(3): 389–392, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Entwistle:1989:MDB

[Ent89b] Ian D. Entwistle. Methods of displaying the behaviour of the mapping z → z^{2} + μ. Computers and Graphics, 13(4):549–551, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Entem:2019:ASO

[EPB^{+}19] Even Entem, Amal Dev Parakkat, Loïc Barthe, Ramanathan Muthuganapathy, and Marie-Paule Cani. Automatic structuring of organic shapes from a single drawing. Computers and Graphics, 81(??):125– 139, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300536.

Egli:1996:MCF

[EPS96] R. Egli, C. Petit, and N. F.

Stewart. Moving coordinate frames for representation and visualization in four dimensions. Computers and Graphics, 20(6):905–919, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600060.

Einarsson:1987:DCD

[ER87] Goran Einarsson and Goran Roth. Data compression of digital color pictures. Computers and Graphics, 11(4): 409–426, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Evans:2014:GWS

[ERB^{+}14] Alun Evans, Marco Romeo, Arash Bahrehmand, Javi Agenjo, and Josep Blat. 3D graphics on the web: a survey. Computers and Graphics, 41(??):43–61, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000260.

Elek:2014:POP

[ERDS14] Oskar Elek, Tobias Ritschel, Carsten Dachsbacher, and Hans-Peter Seidel. Principal-Ordinates Propagation for real-time rendering of participating media. Computers and Graphics, 45 (??):28–39, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000776.

Elek:2012:ICR

[ERWS12] Oskar Elek, Tobias Ritschel, Alexander Wilkie, and Hans-Peter Seidel. Interactive cloud rendering using temporally coherent photon mapping. Computers and Graphics, 36(8): 1109–1118, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001598.

Etxezarreta:2022:RTC

[ES22] A. Etxezarreta and M. Sagardia. Real time contact surface prediction for grasping with complex geometries. Computers and Graphics, 107(??):66–72, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001248.

El-Sana:2002:IMP

[ESAH02] Jihad El-Sana, Nir Asis, and Ofer Hadar. Integrating motion perception with view-dependent rendering for dynamic environments. Computers and Graphics, 26(5):721–731, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/35/abstract.
html.

El-Said:1997:ATA

[ESFGDZ97] M. G. El-Said, G. Fischer, S. A. Gamalel-Din, and M. Zaki. ADDI: a tool for automating the design of visual inter-faces. Computers and Graphics, 21(1):79– 87, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=1&aid=9600072.

Elmqvist:2007:VPA

[ET07] Niklas Elmqvist and Philippas Tsigas. View-projection animation for 3D occlusion management. Computers and Graphics, 31(6): 864–876, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001719.

Egodagamage:2018:DMV

[ET18] Ruwan Egodagamage and Mihran Tuceryan. Distributed monocular visual SLAM as a basis for a collaborative augmented reality framework. Computers and Graphics, 71(??): 113–123, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300025.

Ehlers:2023:IRS

[EVRW23] Henry Ehlers, Anaïs Villedieu, Renata G. Raidou, and Hsiang-Yun Wu. Improving readability of static, straight-line graph drawings: a first look at edge crossing resolution through iterative vertex splitting. Computers and Graphics, 116(??): 448–463, November 2023.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002339.

Eastman:1975:GPE

[EW75] J. F. Eastman and D. R.

Wooten. A general purpose, expandable processor for real-time computer graphics. Computers and Graphics, 1(1):73–78, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Eberhardt:1999:PSA

[EW99] B. Eberhardt and A. Weber. A particle system approach to knitted textiles. Computers and Graphics, 23 (4):599–606, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/39/article.
pdf.

Ewins:2000:IAT

[EWWL00] Jon P. Ewins, Marcus D.

Waller, Martin White, and Paul F. Lister. Implementing an anisotropic texture filter. Computers and Graphics, 24(2):253–267, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/33/article.
                                        pdf.

Eisemann:2022:FSS

[EZ22] Elmar Eisemann and Fang-Lue Zhang. Foreword to the special section on Pacific Graphics 2020. Computers and Graphics, 107 (??):158–160, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001376.

Frame:1994:SNI

[FA94] Michael Frame and Maureen Angers. Some nonlinear iterated function systems. Computers and Graphics, 18(1):119–125, January– February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Falk:1989:TDM

[Fal89] Steve Falk. Three-dimensional modelling and Piero Della Francesca’s use of linear perspective. Computers and Graphics, 13(3):345–347, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fiser:2016:ADB

[FASS16] Jakub Fiser, Paul Asente, Stephen Schiller, and Daniel Sýkora. Advanced drawing beautification with Ship-Shape. Computers and Graphics, 56(??):46–58, May 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849316300127.

Fathauer:2001:FTB

[Fat01] Robert W. Fathauer. Fractal tilings based on kiteand dart-shaped prototiles. Computers and Graphics, 25(2):323–331, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/38/article.
pdf.

Fathauer:2002:CGF

[Fat02] Robert W. Fathauer. Chaos and graphics: Fractal tilings based on v-shaped prototiles. Computers and Graphics, 26(4):635–643, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/45/abstract.
html.

Fay:1985:CGI

[Fay85] D. Fay. Curve generation of implicit functions using DDA techniques. Computers and Graphics, 9(1):67, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fahim:2021:SVR

[FAZ21] George Fahim, Khalid Amin, and Sameh Zarif. Singleview 3D reconstruction: a survey of deep learning methods. Computers and Graphics, 94(??):164– 190, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301849.

Finke:2004:RAS

[FB04] Matthias Finke and Dirk Balfanz. A reference architecture supporting hypervideo content for ITV and the Internet domain. Computers and Graphics, 28(2):179–191, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Favreau:2011:TSC

[FB11] Jean-Marie Favreau and Vincent Barra. Tiling surfaces with cylinders using n-loops. Computers and Graphics, 35(1):35– 42, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001706.

Fernandez:2012:ILD

[FB12] Eduardo Fernández and Gonzalo Besuievsky. Inverse lighting design for interior buildings integrating natural and artificial sources. Computers and Graphics, 36(8): 1096–1108, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001550.

Fernandez:2014:SBM

[FB14] Eduardo Fernández and Gonzalo Besuievsky. A sample-based method for computing the radiosity inverse matrix. Computers and Graphics, 41(??): 1–12, June 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000259.

See erratum [FB15a].

Fernandez:2015:ESB

[FB15a] Eduardo Fernández and Gonzalo Besuievsky. Erratum to ‘A sample-based method for computing the radiosity inverse matrix’ [Computers & Graphics 41 (2014) 1–12]. Computers and Graphics, 49 (??):36, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000448.

See [FB14].

Fernandez:2015:IOD

[FB15b] Eduardo Fernández and Gonzalo Besuievsky. Inverse opening design with anisotropic lighting incidence. Computers and Graphics, 47(??):113–122, April 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000047.

Fang:2001:CFB

[FBH^{+}01] Xiang Fang, Hujun Bao, Pheng Ann Heng, TienTsin Wong, and Qunsheng Peng. Continuous field based freeform surface modeling and morphing. Computers and Graphics, 25(2):235– 243, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/31/article.
                                        pdf.

Fribourg:2021:DVT

[FBH^{+}21] Rebecca Fribourg, Evan Blanpied, Ludovic Hoyet, Anatole Lécuyer, and Ferran Argelaguet. Does virtual threat harm VR experience?: Impact of threat occurrence and repeatability on virtual embodiment and threat response. Computers and Graphics, 100(??): 125–136, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001485.

Fu:1996:ARA

[FBP96] Sheng Fu, Hujun Bao, and Qunsheng Peng. An accelerated rendering algorithm for stereoscopic display. Computers and Graphics, 20(2): 223–229, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500129.

Fondevilla:2017:PPR

[FBR^{+}17] Amélie Fondevilla, Adrien Bousseau, Damien Rohmer, Stefanie Hahmann, and Marie-Paule Cani. Patterns from photograph: Reverse-engineering developable products. Computers and Graphics, 66(??): 4–13, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300663.

Figueiredo:1993:AIT

[FBT93] Mauro Figueiredo, Klaus Bohm, and Jose Teixeira. Advanced interaction techniques in virtual environments. Computers and Graphics, 17(6):655–661, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fairen:2004:MPV

[FBT04] Marta Fairén, Pere Brunet, and Torsten Techmann. MiniVR: a portable virtual reality system. Computers and Graphics, 28(2):289– 296, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fulop-Balogh:2022:DSN

[FBTT^{+}22] Beatrix-Emőke Fülöp-Balogh, Eleanor Tursman, James Tompkin, Julie Digne, and Nicolas Bonneel. Dynamic scene novel view synthesis via deferred spatio-temporal consistency. Computers and Graphics, 107(??):220– 230, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200142X.

Fang:2000:HAV

[FC00a] Shiaofen Fang and Hongsheng Chen. Hardware accelerated voxelization. Computers and Graphics, 24(3): 433–442, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/38/article.
pdf.

Frame:2000:ICI

[FC00b] Michael Frame and Tatiana Cogevina. An infinite circle inversion limit set fractal. Computers and Graphics, 24 (5):797–804, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/38/article.
pdf.

Fuchs:2007:DED

[FCG^{+}07] Christian Fuchs, Tongbo Chen, Michael Goesele, Holger Theisel, and Hans-Peter Seidel. Density estimation for dynamic volumes. Computers and Graphics, 31(2):205–211, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002408.

Ferreira:2021:LDG

[FCG^{+}21] João P. Ferreira, Thiago M. Coutinho, Thiago L. Gomes, José F. Neto, Rafael Azevedo, Renato Martins, and Erickson R. Nascimento. Learning to dance: a graph convolutional adversarial network to generate realistic dance motions from audio. Computers and Graphics, 94 (??):11–21, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301436.

Fanni:2018:FOS

[FCM^{+}18] Filippo A. Fanni, Gianmarco Cherchi, Alessandro Muntoni, Alessandro Tola, and Riccardo Scateni. Fabrication oriented shape decomposition using polycube mapping. Computers and Graphics, 77(??): 183–193, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301717.

Ferreira:1990:GDR

[FCSB90] F. Nunes Ferreira, A. Cardoso Costa, A. Augusto Sousa, and V. Afonso Branco. 3D graphics developments and research at IN-ESC Norte. Computers and Graphics, 14(1):47–53, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fuchs:2010:RTT

[FCW^{+}10] Martin Fuchs, Tongbo Chen, Oliver Wang, Ramesh Raskar, Hans-Peter Seidel, and Hendrik P. A. Lensch. Realtime temporal shaping of high-speed video streams. Computers and Graphics, 34 (5):575–584, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000889.

Fulton:1975:LPP

[FD75] D. L. Fulton and R. T. Duquet. List processing primitives for BASIC. Computers and Graphics, 1(2– 3):203–210, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Franzetti:2003:FGA

[FDA03] T. Franzetti, A. M. Day, and D. B. Arnold. A fan growing algorithm for efficient vertex caching. Computers and Graphics, 27 (5):773–789, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fayer:2018:TIC

[FDGM18] Julien Fayer, Bastien Durix, Simone Gasparini, and Géraldine Morin. Texturing and inpainting a complete tubular 3D object reconstructed from partial views. Computers and Graphics, 74 (??):126–136, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300773.

Fischer:2007:HTM

[FEBS07] Jan Fischer, Michael Eichler, Dirk Bartz, and Wolfgang Straßer. A hybrid tracking method for surgical augmented reality. Computers and Graphics, 31 (1):39–52, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001841.

Figueiredo:2004:RDP

[FEJ04] Frederico C. Figueiredo, Dena E. Eber, and Joaquim A. Jorge. Refereed digital publication of computer graphics educational materials. Computers and Graphics, 28 (1):119–124, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fritsch:1975:MGS

[FEJM75] J. M. Fritsch, R. A. Ellis, T. H. Jacobi, and G. R. Marshall. A macromodular graphics system for protein structure research. Computers and Graphics, 1(2– 3):271–278, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Feng:1988:RBA

[Fen88] Lin Feng. Rule-based art pattern CAD. Computers and Graphics, 12(3–4): 323–327, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ferri:2001:VID

[Fer01] Lucilla Croce Ferri. Visualization of 3D information with digital holography using laser printers. Computers and Graphics, 25(2):309–321, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/37/article.
pdf.

Finkbeiner:2010:EVR

[FEVM10] Bernhard Finkbeiner, Alireza Entezari, Dimitri Van De Ville, and Torsten Möller.

Efficient volume rendering on the body centered cubic lattice using box splines. Computers and Graphics, 34 (4):409–423, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000245.

Fellner:1996:CGI

[FF96] Dieter W. Fellner and Martin Fischer. Computer graphics interface (CGI): a good concept and a valuable tool for research and teaching in computer graphics. Computers and Graphics, 20(2):341–346, March–April 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500135.

Fuvattanasilp:2021:PSG

[FFP^{+}21] Varunyu Fuvattanasilp, Yuichiro Fujimoto, Alexander Plopski, Takafumi Taketomi, Christian Sandor, Masayuki Kanbara, and Hirokazu Kato. SlidAR+: Gravityaware 3D object manipulation for handheld augmented reality. Computers and Graphics, 95(??): 23–35, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000054.

Fryazinov:2011:FBV

[FFV^{+}11] Oleg Fryazinov, Pierre-Alain Fayolle, Turlif Vilbrandt, Galina Pasko, and Alexander Pasko. Feature based volumes for implicit intersections. Computers and Graphics, 35(3): 524–531, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000689.

Fechter:1996:UCD

[FGES96] Jürgen Fechter, Thomas Grunert, L. Miguel Encarnação, and Wolfgang Straßer. User-centered development of medical visualization applications: Flexible interaction through communicating application objects. Computers and Graphics, 20(6):763–774, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600054.

Fu:2023:HPE

[FGL23] Huichen Fu, Junwei Gao, and Huabo Liu. Human pose estimation and action recognition for fitness movements. Computers and Graphics, 116(??): 418–426, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002315.

Fuhrmann:2003:IFD

[FGLW03] Arnulph Fuhrmann, Clemens Groß, Volker Luckas, and Andreas Weber. Interactionfree dressing of virtual humans. Computers and Graphics, 27(1):71–82, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Furmanova:2020:PVV

[FGM^{+}20] Katarína Furmanová, Nicolas Grossmann, Ludvig P. Muren, Oscar Casares-Magaz, Vitali Moiseenko, John P. Einck, M. Eduard Gröller, and Renata G. Raidou. VAPOR: Visual analytics for the exploration of pelvic organ variability in radiotherapy. Computers and Graphics, 91(??): 25–38, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300960.

Frosini:2022:FSS

[FGMR22] Patrizio Frosini, Daniela Giorgi, Simone Melzi, and Emanuele Rodolà. Foreword to the special section on Smart Tools and Applications in Graphics (STAG 2021). Computers and Graphics, 105(??): A12–A13, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000863.

Field:2010:ETC

[FGP^{+}10] M. Field, S. Gordon, E. Peterson, R. Robinson, T. Stahovich, and C. Alvarado. The effect of task on classification accuracy: Using gesture recognition techniques in free-sketch recognition. Computers and Graphics, 34 (5):499–512, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001056.

Feng:2022:STO

[FGZ^{+}22] Yifan Feng, Yue Gao, Xibin Zhao, Yandong Guo, Nihar Bagewadi, Nhat-Tan Bui, Hieu Dao, Shankar Gangisetty, Ripeng Guan, Xie Han, Cong Hua, Chidambar Hunakunti, Yu Jiang, Shichao Jiao, Yuqi Ke, Liqun Kuang, Anan Liu, Dinh-Huan Nguyen, Hai-Dang Nguyen, Weizhi Nie, Bang-Dang Pham, Karthik Raikar, Qingmei Tang, Minh-Triet Tran, Jialong Wan, Chenggang Yan, Haoxuan You, and Difei Zhu. SHREC’22 track: Open-set 3D object retrieval. Computers and Graphics, 107(??):231– 240, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001443.

Fellner:1994:BAG

[FH94] Dieter W. Fellner and Christoph Helmberg. Best approximate general ellipses on integer grids. Computers and Graphics, 18(2): 143–151, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Falcidieno:2011:SSSa

[FH11a] Bianca Falcidieno and Ivan Herman. Special section on semantic 3D media and content. Computers and Graphics, 35(2):vi–vii, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001883.

Falcidieno:2011:SSSb

[FH11b] Bianca Falcidieno and Ivan Herman. Special section on semantic 3D media and content. Computers and Graphics, 35(4): iii–iv, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001191.

Fuhrmann:1999:OCA

[FHFG99] Anton Fuhrmann, Gerd Hesina, François Faure, and Michael Gervautz. Occlusion in collaborative augmented environments. Computers and Graphics, 23(6): 809–819, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/33/article.
pdf.

Fellner:1998:GID

[FHM98] D. W. Fellner, S. Havemann, and G. Müller. Graphics in/for digital libraries

— modeling of and navigation in complex 3D documents. Computers and Graphics, 22(6):647–653, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/644.pdf.

Fowler:1989:MSP

[FHP89] Deborah R. Fowler, James Hanan, and Przemysław Prusinkiewicz. Modelling spiral phyllotaxis. Computers and Graphics, 13(3): 291–296, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Figueiredo:2021:DLB

[FIC21] Lucas Figueiredo, Paulo Ivson, and Waldemar Celes. Deep learning-based framework for shape instance registration on 3D CAD models. Computers and Graphics, 101(??):72– 81, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100176X.

Figueiredo:2023:UMI

[FIC23] Lucas Figueiredo, Paulo Ivson, and Waldemar Celes. Unsupervised method for identifying shape instances on 3D CAD models. Computers and Graphics, 116(??):228–238, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323001875.

Fiume:1987:BMG

[Fiu87] Eugene Fiume. Bit-mapped graphics: a semantics and theory. Computers and Graphics, 11(2):121–140, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fiume:1989:AOC

[Fiu89] Eugene Fiume. Active objects in the construction of graphical user interfaces. Computers and Graphics, 13 (3):321–327, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fonseca:2003:TCB

[FJ03] Manuel J. Fonseca and Joaquim A. Jorge. Towards content-based retrieval of technical drawings through high-dimensional indexing. Computers and Graphics, 27 (1):61–69, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fickel:2017:DME

[FJ17] Guilherme P. Fickel and Cláudio R. Jung. Disparity map estimation and view synthesis using temporally adaptive triangular meshes. Computers and Graphics, 68 (??):43–52, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730136X.

Feito:2008:ISI

[FJA08] Francisco R. Feito and Robert Joan-Arinyo. Introduction to this special issue. Computers and Graphics, 32 (4):383–384, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000770.

Florinabel:2011:CFS

[FJS11] D. Jemi Florinabel, S. Ebenezer Juliet, and V. Sadasivam. Combined frequency and spatial domain-based patch propagation for image completion. Computers and Graphics, 35(6): 1051–1062, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001488.

Feng:2011:VIC

[FJW11] Powei Feng, Tao Ju, and Joe Warren. View-independent contour culling of 3D density maps for far-field viewing of iso-surfaces. Computers and Graphics, 35(3): 561–568, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000744.

Frost:1982:AAP

[FK82] Charles E. Frost and Gary L. Kinzel. An automatic adjustment procedure for rational splines. Computers and Graphics, 6(4):171–176, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Finnegan:2004:SIE

[FK04] John Finnegan and Lars Kjelldahl. Special issue on education. Computers and Graphics, 28(3):381– 382, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fu:2011:NNB

[FK11] Luoting Fu and Levent Burak Kara. Neural networkbased symbol recognition using a few labeled samples. Computers and Graphics, 35 (5):955–966, October 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001300.

Fletcher:1991:SPC

[Fle91] Peter Fletcher. A SIMD parallel colour quantization algorithm. Computers and Graphics, 15(3): 365–373, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Li:2011:FHM

[fLhLfT11] Meng fei Li, Sheng hui Liao, and Ruo feng Tong. Facial hexahedral mesh transferring by volumetric mapping based on harmonic fields. Computers and Graphics, 35 (1):92–98, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001822.

Fuhrmann:2015:MIB

[FLM^{+}15] Simon Fuhrmann, Fabian Langguth, Nils Moehrle, Michael Waechter, and Michael Goesele. MVE — an image-based reconstruction environment. Computers and Graphics, 53 (part A) (??):44–53, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500148X.

Fu:2016:P

[FLMY16] Hongbo Fu, Xin Li, Lizhuang Ma, and Jun-Hai Yong. Preface. Computers and Graphics, 54(??):154–156, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001375.

FloresCampana:2023:VHV

[FLR^{+}23] Jose Luis Flores Campana, Luís Gustavo Lorgus Decker, Marcos Roberto e Souza, Helena de Almeida Maia, and Helio Pedrini. Variablehyperparameter visual transformer for efficient image inpainting. Computers and Graphics, 113(??): 57–68, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000614.

Fugacci:2020:CSP

[FLV20] Ulderico Fugacci, Claudia Landi, and Hanife Varli. Critical sets of PL and discrete Morse theory: a correspondence. Computers and Graphics, 90(??): 43–50, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300716.

Fels:1999:CGI

[FM99] Sidney Fels and Kenji Mase.

Chaos and graphics — Iamascope: a graphical musical instrument. Computers and Graphics, 23(2):277– 286, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/680.pdf.

Frame:2000:RBQ

[FM00] Michael Frame and Shontel Meachem. Reverse bifurcations in a quartic family. Computers and Graphics, 24 (1):143–149, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/27/39/article.
                                        pdf.

Fourey:2009:NED

[FM09] Sébastien Fourey and Rémy Malgouyres. Normals estimation for digital surfaces based on convolutions. Computers and Graphics, 33(1):2–10, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001489.

Francu:2017:PBS

[FM17] Mihai Frâncu and Florica Moldoveanu. Position based simulation of solids with accurate contact handling. Computers and Graphics, 69 (??):12–23, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301565.

Franke:2022:PGK

[FM22] Kai Franke and Heinrich Müller. Procedural generation of 3D karst caves with speleothems. Computers and Graphics, 102 (??):533–545, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002132.

Furmanova:2021:PPP

[FMCM^{+}21] Katarína Furmanová, Ludvig P. Muren, Oscar Casares-Magaz, Vitali Moiseenko, John P. Einck, Sara Pilskog, and Renata G. Raidou. PREVIS: Predictive visual analytics of anatomical variability for radiotherapy decision support. Computers and Graphics, 97(??): 126–138, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000522.

Feng:1996:NFF

[FMP96] Jieqing Feng, Lizhuang Ma, and Qunsheng Peng. A new free-form deformation through the control of parametric surfaces. Computers and Graphics, 20(4): 531–539, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600025.

Friedrich:1998:GID

[FMS98] M. Friedrich, M. Melle, and D. Saupe. Graphics in/for digital libraries

— ATLAS2000 — atlases of the future on the Internet. Computers and Graphics, 22(6):697–701, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/624.pdf.

Freytag:1999:RAW

[FN99] C. Freytag and L. Neumann. Resource adaptive WWW access for mobile applications. Computers and Graphics, 23(6): 841–848, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/38/article.
pdf.

Frame:2008:FTW

[FN08] Michael Frame and Nial Neger. Fractal tetrahedra: What’s left in, what’s left out, and how to build one in four dimensions. Computers and Graphics, 32(3): 371–381, June 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307002026.

Frame:2010:DPF

[FN10] Michael Frame and Nial Neger. Dimensions and the probability of finding odd numbers in Pascal’s triangle and its relatives. Computers and Graphics, 34(2): 158–166, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001162.

Ferstl:2020:AGG

[FNM20] Ylva Ferstl, Michael Neff, and Rachel McDonnell. Adversarial gesture generation with realistic gesture phasing. Computers and Graphics, 89(??): 117–130, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300455.

Fontanari:2021:SMS

[FO21] Thomas V. Fontanari and Manuel M. Oliveira. Simultaneous magnification of subtle motions and color variations in videos using Riesz pyramids. Computers and Graphics, 101(??): 35–45, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001795.

Fornari:1984:IAD

[For84] Monique Fornari. An integrated approach from design to manufacturing: CAD/ CAM at Avions Marcel Dassault-Breguet aviation. Computers and Graphics, 8 (3):275–280, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fossmeier:1987:AMA

[Fos87] Reinhard Fossmeier. Animating matrix algorithms. Computers and Graphics, 11 (3):309–311, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fournier:2011:MFA

[Fou11] Marc Fournier. Mesh filtering algorithm using an adaptive 3D convolution kernel applied to a volumebased vector distance field. Computers and Graphics, 35(3):668–676, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000586.

Futrelle:1975:SDG

[FP75] R. P. Futrelle and M. J. Potel. The system design for GALATEA, an interactive real-time computer graphics system for movie and video analysis. Computers and Graphics, 1(1): 115–122, May 1975. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fellner:1987:BOV

[FP87] W. D. Fellner and R. Posch.

Bildschirmtext — an open videotex network for text and graphic applications. Computers and Graphics, 11 (4):359–367, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fortescue:1989:NPU

[FP89] Peter D. Fortescue and Geoffrey S. Puterbaugh. Note on the practical use of the VGA monitor for the display of fractal images. Computers and Graphics, 13(4): 559–560, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fryazinov:2010:FRI

[FPC10] Oleg Fryazinov, Alexander Pasko, and Peter Comninos. Fast reliable interrogation of procedurally defined implicit surfaces using extended revised affine arithmetic. Computers and Graphics, 34(6): 708–718, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000107X.

Frame:1992:NSA

[FPR92] Michael Frame, A. G. Davis Philip, and Adam Robucci. A new scaling along the spike of the Mandelbrot set. Computers and Graphics, 16 (2):223–234, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Flaquer:1988:BOB

[FR88] J. Flaquer and J. L. Rodil. Boolean operations based on the planar polyhedral representation. Computers and Graphics, 12(1):59–64, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Flesia:1992:DRG

[FR92a] S. Flesia and M. Roche. 3-D reconstruction of generalized cylinders from biplane projections. Computers and Graphics, 16 (2):167–173, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Frame:1992:GMS

[FR92b] Michael Frame and James Robertson. A generalized Mandelbrot set and the role of critical points. Computers and Graphics, 16 (1):35–40, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Feito:1998:TSG

[FR98] F. R. Feito and M. Rivero.

Technical section — geometric modelling based on simplicial chains. Computers and Graphics, 22(5): 611–619, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/5/601.pdf.

Ferreira:2004:ISI

[FR04] P. Ferreira and H. Raffler. Introduction to the special issue. Computers and Graphics, 28(5):623– 624, October 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Franklin:1983:SAB

[Fra83]  W. R. Franklin.  Soft-

ware aspects of business graphics. Computers and Graphics, 7(1):5–12, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Franz:1986:CCB

[Fra86] X. Franz. CAD/CAM on the basis of programmed know-how in a geometrical data compound. Computers and Graphics, 10(4): 307–315, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Frame:1994:SCA

[Fra94] Michael Frame. Sensitivity in cellular automata: Some examples. Computers and Graphics, 18(5):733–737, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fontana:2006:COC

[FRC06] Marzia Fontana, Caterina Rizzi, and Umberto Cugini. A CAD-oriented cloth simulation system with stable and efficient ODE solution. Computers and Graphics, 30(3):391–406, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000616.

Fribert:2003:SMC

[Fri03] Miroslav Fribert. Separation model of colour regions in a halftone print. Computers and Graphics, 27 (5):801–806, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Figueiredo:2018:CED

[FRTT18] Lucas Figueiredo, Eduardo Rodrigues, João Teixeira, and Veronica Teichrieb. A comparative evaluation of direct hand and wand interactions on consumer devices. Computers and Graphics, 77(??):108–121, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301675.

Fruhauf:1991:VVW

[Frü91] Martin Frühauf. Volume visualization on workstations: Image quality and efficiency of different techniques. Computers and Graphics, 15(1): 101–107, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fruhauf:1994:IVV

[Fru94]  Thomas Fruhauf.  In-

teractive visualization of vector data in unstructured volumes. Computers and Graphics, 18(1):73– 80, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fei:2014:PBB

[FRWW14] Yun Fei, Guodong Rong, Bin Wang, and Wenping Wang. Parallel L-BFGS-B algorithm on GPU. Computers and Graphics, 40 (??):1–9, May 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000119.

Findler:1978:MPL

[FS78] Nicholas V. Findler and Jerry N. Shaw. Multi-Pierre

— a learning robot system. Computers and Graphics, 3(2–3):107–111, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fishman:1980:CDH

[FS80] B. Fishman and B. Schachter. Computer display of height fields. Computers and Graphics, 5(2–4):53–60, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fabrikant:1986:AGM

[FS86a] V. I. Fabrikant and T. S. Sankar. An algorithm for geometrical modelling of surfaces of revolution. Computers and Graphics, 10(3): 245–255, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Flasche:1986:DPD

[FS86b] Ute Flasche and Angela Scheller. Decentralized processing of documents. Computers and Graphics, 10(2): 119–131, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fuhs:1988:CPC

[FS88] Wilhelm Fuhs and Hellmuth Stachel. Circular pipeconnections. Computers and Graphics, 12(1):53–57, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Feng:1998:ACE

[FS98] Lin Feng and Seah Hock Soon. Algorithms corner — an effective 3D seed fill algorithm. Computers and Graphics, 22(5):641–644, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/607.pdf.

Fujishiro:2017:FSI

[FS17] Issei Fujishiro and Alexei Sourin. Foreword to the special issue on 2016 International Conference on Cyberworlds (CYBER-WORLDS 2016). Computers and Graphics, 64 (??):1–2, May 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300213.

Ferreira:2020:BBB

[FS20] Leonardo Souto Ferreira and Leonardo Sacht. Bounded biharmonic blending of Möbius transformations for flexible omnidirectional image rectification. Computers and Graphics, 93 (??):51–60, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030145X.

Fukumoto:1994:FPP

[FSM94] Masaaki Fukumoto, Yasuhito Suenaga, and Kenji Mase. “finger-pointer”: Pointing interface by image processing. Computers and Graphics, 18(5):633–642, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Fryazinov:2015:SCV

[FSP15] Oleg Fryazinov, Mathieu Sanchez, and Alexander Pasko. Shape conforming volumetric interpolation with interior distances. Computers and Graphics, 46 (??):149–155, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001162.

Fenster:2002:UIA

[FSS^{+}02] Aaron Fenster, Kathleen Surry, Wendy Smith, Jeremy Gill, and Dónal B. Downey. 3D ultrasound imaging: applications in image-guided therapy and biopsy. Computers and Graphics, 26(4): 557–568, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/55/32/abstract.
html.

Feng:1997:SMC

[FST97] Lin Feng, Seah Hock Soon, and Lee Yong Tsui. Structure modeling and contextfree grammar: Exploring a new approach for surface boundary construction. Computers and Graphics, 21(6):777–785, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700057.

Ferreira:2017:LMT

[FSV17] Leonardo Souto Ferreira, Leonardo Sacht, and Luiz Velho. Local Moebius transformations applied to omnidirectional images. Computers and Graphics, 68 (??):77–83, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301346.

Feito:1997:ITG

[FT97] F. R. Feito and J. C. Torres. Inclusion test for general polyhedra. Computers and Graphics, 21(1):23– 30, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=1&aid=9600067.

Fujishiro:2002:BPC

[FT02] Issei Fujishiro and Yuriko Takeshima. Best papers of CAD & CG 2001: Coherence-sensitive solid fitting. Computers and Graphics, 26(3):417–427, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/54/31/abstract.
                                        html.

Fernandes:2024:FSS

[FT24] Leandro A. F. Fernandes and Rafael Piccin Torchelsen. Foreword to the special section on SIB-GRAPI 2022 tutorials. Computers and Graphics, 118 (??):33, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002832.

Faraj:2012:VSF

[FTB12] Noura Faraj, Jean-Marc Thiery, and Tamy Boubekeur. VoxMorph: 3-scale freeform deformation of large voxel grids. Computers and Graphics, 36(5):562–568, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000593.

Faraj:2016:MMA

[FTB16] Noura Faraj, Jean-Marc Thiery, and Tamy Boubekeur. Multi-material adaptive volume remesher. Computers and Graphics, 58(??):150– 160, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300620.

Feito:1995:OSI

[FTU95] F. Feito, J. C. Torres, and A. Ureña. Orientation, simplicity, and inclusion test for planar polygons. Computers and Graphics, 19(4): 595–600, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500037.

Funkhouser:1999:VAH

[Fun99] Thomas A. Funkhouser. A visibility algorithm for hybrid geometry- and imagebased modeling and rendering. Computers and Graphics, 23(5):719–728, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/32/
                                        35/article.pdf;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/35/abstract.
                                        html.

Falcidieno:2006:SRU

[FV06] Bianca Falcidieno and Remco Veltkamp. Shape reasoning and understanding. Computers and Graphics, 30(2): 158–159, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000021.

Filip:2013:FMS

[FV13] Jirí Filip and Radomír Vávra. Fast method of sparse acquisition and reconstruction of view and illumination dependent datasets. Computers and Graphics, 37 (5):376–388, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000332.

Faloutsos:2001:VSD

[FvdPT01] Petros Faloutsos, Michiel van de Panne, and Demetri Terzopoulos. The virtual stuntman: dynamic characters with a repertoire of autonomous motor skills. Computers and Graphics, 25(6):933–953, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/29/abstract.
html.

Frerichs:2015:SOD

[FVG15] Dhana Frerichs, Andrew Vidler, and Christos Gatzidis. A survey on object deformation and decomposition in computer graphics. Computers and Graphics, 52 (??):18–32, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000898.

Feng:1997:MCF

[FWCS97] Dong Feng, Xiaomei Wang, Wenli Cai, and Jiaoying Shi. A mass conservative flow field visualization method. Computers and Graphics, 21(6):749–756, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700054.

Fellegara:2021:SDC

[FWD21] Riccardo Fellegara, Kenneth Weiss, and Leila De Floriani. The Stellar decomposition: a compact representation for simplicial complexes and beyond. Computers and Graphics, 98 (??):322–343, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000819.

Fink:2013:TMG

[FWW13] H. Fink, T. Weber, and M. Wimmer. Teaching a modern graphics pipeline using a shader-based software renderer. Computers and Graphics, 37(1–2): 12–20, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001628.

Feng:2018:NSP

[FWX^{+}18] Xiang Feng, Wanggen Wan, Richard Yi Da Xu, Stuart Perry, Pengfei Li, and Song Zhu. A novel spatial pooling method for 3D mesh quality assessment based on percentile weighting strategy. Computers and Graphics, 74(??):12–22, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300530.

Falby:1993:NHD

[FZPM93] John S. Falby, Michael J.

Zyda, David R. Pratt, and Randy L. Mackey. NPSNET: Hierarchical data structures for real-time threedimensional visual simulation. Computers and Graphics, 17(1):65–69, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Goldenstein:2001:SND

[G^{+}01] Siome Goldenstein et al.

Scalable nonlinear dynamical systems for agent steering and crowd simulation. Computers and Graphics, 25(6):983–998, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/32/abstract.
html.

Grabowski:1983:CTS

[GA83a] H. Grabowski and R. Anderl. CAD-turnkey-systems. Computers and Graphics, 7(1):91, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grabowski:1983:IDM

[GA83b] H. Grabowski and R. Anderl. Integration of the design and manufacture planning process based on a CAD system with a technology oriented volume model. Computers and Graphics, 7(2): 125–141, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grabowski:1988:ICS

[GA88] H. Grabowski and R. Anderl. Institute for Computer Science and Application in Planning and Design. Computers and Graphics, 12(3–4):583–585, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garcia:2007:CRP

[GA07] Raquel Paricio García and Juan Manuel Moreno Aróstegui. A cooperative robotic platform for adaptive and immersive artistic installations. Computers and Graphics, 31 (6):809–817, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001562.

Gonen:2012:SBM

[GA12] Ozgur Gonen and Ergun Akleman. Sketch based 3D modeling with curvature classification. Computers and Graphics, 36 (5):521–525, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000581.

Gabrielson:1977:MGT

[Gab77] Verlan K. Gabrielson. Mesh generation for two-dimensional regions using a DVST (Direct View Storage Tube) graphics terminal. Computers and Graphics, 2(2): 59–66, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gagalowicz:1995:TAT

[Gag95] André Gagalowicz. Tools for advanced telepresence systems. Computers and Graphics, 19(1):73–88, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400123.

Galligani:1988:LED

[Gal88] Ilio Galligani. A laboratory for experimental data processing. Computers and Graphics, 12(3–4):587–588, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guarese:2021:ASV

[GANM21] Renan Guarese, Pererik Andreasson, Emil Nilsson, and Anderson Maciel. Augmented situated visualization methods towards electromagnetic compatibility testing. Computers and Graphics, 94 (??):1–10, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301527.

Gudukbay:2002:TSV

[GAÖ02] Uǧur Güdükbay, Okan Arikan, and Bülent Özgüç. Technical section: Visualizer: a mesh visualization system using viewdependent refinement. Computers and Graphics, 26(3): 491–503, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/54/37/abstract.
                                        html.

Gur:1975:DMI

[GB75] Y. Gur and C. Biddle. Data management for an interactive graphics system oriented to planning application. Computers and Graphics, 1 (4):325–328, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gujar:1991:FCP

[GB91] Uday G. Gujar and Virendra C. Bhavsar. Fractals from z ← z^{α} + c in the complex c-plane. Computers and Graphics, 15(3): 441–449, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Goncalves:2018:FSSa

[GB18] Daniel Gonçalves and Maximino Bessa. Foreword to the special section on EPCGI 2016. Computers and Graphics, 71(??):A3, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300037.

Gao:2015:APT

[GBA15] Xihe Gao, Stephen Brooks, and Dirk V. Arnold. Automated parameter tuning for tone mapping using visual saliency. Computers and Graphics, 52(??): 171–180, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000837.

Gujar:1988:ITO

[GBD88] Uday G. Gujar, Virendra C.

Bhavsar, and Narendra N. Datar. Interpolation techniques for 3D object generation. Computers and Graphics, 12(3–4):541–555, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garnier:2014:DCB

[GBF14] Lionel Garnier, Hichem Barki, and Sebti Foufou. Dupin cyclide blends between non-natural quadrics of revolution and concrete shape modeling applications. Computers and Graphics, 42(??):31–41, August 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000417.

Goren-Bar:2004:FRT

[GBG04] Dina Goren-Bar and Oded Glinansky. FIT-recommending TV programs to family members. Computers and Graphics, 28(2):149– 156, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grimm:2004:RDA

[GBKG04] Sören Grimm, Stefan Bruckner, Armin Kanitsar, and Eduard Gröller. A refined data addressing and processing scheme to accelerate volume raycasting. Computers and Graphics, 28 (5):719–729, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guennebaud:2004:DSS

[GBP04] Gaël Guennebaud, Loïc Barthe, and Mathias Paulin. Dynamic surfel set refinement for high-quality rendering. Computers and Graphics, 28(6):827–838, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gribble:2008:PGI

[GBP08] Christiaan P. Gribble, Carson Brownlee, and Steven G. Parker. Practical global illumination for interactive particle visualization. Computers and Graphics, 32 (1):14–24, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001987.

Gissler:2017:GDF

[GBP^{+}17] Christoph Gissler, Stefan Band, Andreas Peer, Markus Ihmsen, and Matthias Teschner. Generalized drag force for particle-based simulations. Computers and Graphics, 69 (??):1–11, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301541.

Gujar:1992:FIC

[GBV92] Uday G. Gujar, Virendra C. Bhavsar, and Nagarjuna Vangala. Fractal images from z ← z^{α} + c in the complex z-plane. Computers and Graphics, 16(1):45–49, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gujar:1986:CES

[GC86] Uday G. Gujar and Claudia A. Crawford. Computer encoding, storage and transposition of musical scores. Computers and Graphics, 10 (1):37–49, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gisbert:2023:IHF

[GCC23] Guillaume Gisbert, Raphaëlle Chaine, and David Coeurjolly. Inpainting holes in folded fabric meshes. Computers and Graphics, 114 (??):201–209, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000808.

Guo:2014:NLS

[GCCZ14] Shihui Guo, Jian Chang, Yang Cao, and Jianjun Zhang. A novel locomotion synthesis and optimisation framework for insects. Computers and Graphics, 38 (??):78–85, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001623.

Gao:2022:CGP

[GCDL22] Depeng Gao, Jinhao Chen, Zhetong Dong, and Hongwei Lin. Connectivityguaranteed porous synthesis in free form model by persistent homology. Computers and Graphics, 106 (??):33–44, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000917.

Gao:2016:MIF

[GCLZ16] Bo Gao, Xiaowu Chen, Jianwei Li, and Dongqing Zou. Modeling interactive furniture from a single image. Computers and Graphics, 58 (??):102–108, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300498.

Guan:2023:LNI

[GCRN23] Yanran Guan, Andrei Chubarau, Ruby Rao, and Derek Nowrouzezahrai. Learning neural implicit representations with surface signal parameterizations. Computers and Graphics, 114 (??):257–264, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001103.

Gumbau:2011:VDP

[GCRR11] Jesús Gumbau, Miguel Chover, Inmaculada Remolar, and Cristina Rebollo. View-dependent pruning for real-time rendering of trees. Computers and Graphics, 35(2):364–374, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001834.

Gubins:2020:PSC

[GCvdS^{+}20] Ilja Gubins, Marten L. Chaillet, Gijs van der Schot, Remco C. Veltkamp, Friedrich Förster, Yu Hao, Xiaohua Wan, Xuefeng Cui, Fa Zhang, Emmanuel Moebel, Xiao Wang, Daisuke Kihara, Xiangrui Zeng, Min Xu, Nguyen P. Nguyen, Tommi White, and Filiz Bunyak. SHREC 2020: Classification in cryo-electron tomograms. Computers and Graphics, 91(??):279– 289, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301126.

Gu:2023:NCN

[GCW23] Pengfei Gu, Danny Z.

Chen, and Chaoli Wang. NeRVI: Compressive neural representation of visualization images for communicating volume visualization results. Computers and Graphics, 116(??): 216–227, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001930.

Gong:2023:MRE

[GCYX23] Wenyong Gong, Wenzhu Chen, Zhongwei Yu, and Xiaohua Xie. Multi-resolution edge-aware lighting enhancement network. Computers and Graphics, 116(??):55– 63, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001747.

Gurunathan:1987:ADC

[GD87] B. Gurunathan and S. G.

Dhande. Algorithms for development of certain classes of ruled surfaces. Computers and Graphics, 11(2): 105–112, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ghazanfarpour:1995:SAA

[GD95] D. Ghazanfarpour and J. M. Dischler. Spectral analysis for automatic 3-D texture generation. Computers and Graphics, 19(3): 413–422, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500011.

Gross:2000:DBE

[GD00] Mark D. Gross and Ellen Yi-Luen Do. Drawing on the back of an envelope: a framework for interacting with application programs by freehand drawing. Computers and Graphics, 24(6): 835–849, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/28/article.
pdf.

Gillies:2004:BRA

[GD04] Marco Gillies and Neil A. Dodgson. Behaviourally rich actions for user-controlled characters. Computers and Graphics, 28(6):945–954, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gawronski:2011:MSW

[GD11] Alexander Gawronski and Michel Dumontier. Mo-SuMo: a Semantic Web service to generate electrostatic potentials across solvent excluded protein surfaces and binding pockets. Computers and Graphics, 35 (4):823–830, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000781.

Gerstner:2013:PIA

[GDA^{+}13] Timothy Gerstner, Doug De-Carlo, Marc Alexa, Adam Finkelstein, Yotam Gingold, and Andrew Nealen. Pixelated image abstraction with integrated user constraints. Computers and Graphics, 37 (5):333–347, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000046.

Garcia-Dorado:2013:AUM

[GDDA13] Ignacio Garcia-Dorado, Ilke Demir, and Daniel G. Aliaga. Automatic urban modeling using volumetric reconstruction with surface graph cuts. Computers and Graphics, 37(7): 896–910, November 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001131.

Garcia:2003:FFS

[GdMF03] Á. L. García, J. Ruiz de Miras, and F. R. Fieto. Free-form solid modelling based on extended simplicial chains using triangular Bézier patches. Computers and Graphics, 27 (1):27–39, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gervautz:1986:TIR

[Ger86] Michael Gervautz. Three improvements of the ray tracing algorithm for CSG trees. Computers and Graphics, 10(4):333–339, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gerdes:1997:MCL

[Ger97] Paulus Gerdes. On mirror curves and Lundadesigns. Computers and Graphics, 21(3):371–378, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=3&aid=9700015.

Gerstner:2002:VVL

[Ger02] Thomas Gerstner. Visualization of very large datasets: Multiresolution extraction and rendering of transparent isosurfaces. Computers and Graphics, 26(2):219– 228, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/29/abstract.
html.

Golovinskiy:2009:CSM

[GF09] Aleksey Golovinskiy and Thomas Funkhouser. Consistent segmentation of 3D models. Computers and Graphics, 33(3):262– 269, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000454.

Gervautz:1996:CGP

[GG96] Michael Gervautz and Markus Gross. Computers and Graphics is planning a new section. Computers and Graphics, 20(1):5– ??, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gultepe:2014:RTV

[GG14] Umut Gültepe and Ugur Güdükbay. Real-time virtual fitting with body measurement and motion smoothing. Computers and Graphics, 43 (??):31–43, October 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000600.

Grossmann:2022:CSE

[GGW22] Nicolas Grossmann, Eduard Gröller, and Manuela Waldner. Concept splatters: Exploration of latent spaces based on human interpretable concepts. Computers and Graphics, 105(??):73–84, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000656.

Gemmar:1991:DIK

[GH91] P. Gemmar and G. Hofele. Design of an Iconic Kernel System. Computers and Graphics, 15(4): 483–493, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ghazanfarpour:1998:BTM

[GH98] D. Ghazanfarpour and J-M. Hasenfratz. A beam tracing method with precise antialiasing for polyhedral scenes. Computers and Graphics, 22(1):103–115, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
1/517.pdf.

Gibson:2003:IRV

[GHCH03] S. Gibson, R. J. Hubbold, J. Cook, and T. L. J. Howard. Interactive reconstruction of virtual environments from video sequences. Computers and Graphics, 27(2):293–301, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Groenewegen:2008:VWI

[GHFH08] Saskia Groenewegen, Stefanie Heinz, Bernd Fröhlich, and Anke Huckauf. Virtual world interfaces for special needs education based on props on a board. Computers and Graphics, 32 (5):589–596, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000769.

Gomes:2018:FSS

[GHK18] Abel Gomes, Stefanie Hahmann, and Evangelos Kalogerakis. Foreword to the Special Section on Shape Modeling International 2018. Computers and Graphics, 74 (??):A3–A5, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300876.

Grosskopf:1996:CAS

[GHM^{+}96] S. Großkopf, A. Hildebrand, R. Malkewitz, W. Müller, R. Ziegler, and G. Graschew. Computer aided surgery — vision and feasibility of an advanced operation theatre. Computers and Graphics, 20(6):825–838, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=6&aid=9600052.

Ghosh:1993:UCF

[Gho93] Pijush K. Ghosh. A unified computational framework for Minkowski operations. Computers and Graphics, 17(4):357–378, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gorke:2006:ILG

[GHS06] Johannes Görke, Frank Hanisch, and Wolfgang Straßer. Introducing live graphics gems to educational material. Computers and Graphics, 30(6): 1033–1041, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001567.

Gerl:2013:IEB

[GI13] Moritz Gerl and Tobias Isenberg. Interactive examplebased hatching. Computers and Graphics, 37(1–2): 65–80, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001744.

Giachetta:2015:FPL

[Gia15] Roberto Giachetta. A framework for processing large scale geospatial and remote sensing data in MapReduce environment. Computers and Graphics, 49(??): 37–46, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000278.

Gilbert:1994:NMM

[Gil94] William J. Gilbert. Newton’s method for multiple roots. Computers and Graphics, 18(2):227–229, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ginsberg:1993:CUS

[Gin93] Myron Ginsberg. Challenges to the use of supercomputers and scientific visualization for automotive applications. Computers and Graphics, 17(5):507–515, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gintz:2002:CGA

[Gin02] Terry W. Gintz. Chaos and graphics: Artist’s statement CQUATS—a nondistributive quad algebra for 3D renderings of Mandelbrot and Julia sets. Computers and Graphics, 26(2):367– 370, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/42/abstract.
html.

Girod:1993:SVM

[Gir93] B. Girod. Scalable video for multimedia workstations. Computers and Graphics, 17(3):269–276, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Graham:1995:IRB

[GIZ95] Phil Graham, S. Sitharama Iyengar, and Si-Qing Zheng. Improved recursive bisection line drawing algorithms. Computers and Graphics, 19(6):847–860, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500067.

Gedicke:2021:PFL

[GJN^{+}21] Sven Gedicke, Adalat Jabrayilov, Benjamin Niedermann, Petra Mutzel, and Jan-Henrik Haunert. Point feature label placement for multi-page maps on smallscreen devices. Computers and Graphics, 100(??):66– 80, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001564.

Gourret:1996:MCC

[GK96] Jean Paul Gourret and Jamal Khamlichi. A model for compression and classification of face data structures. Computers and Graphics, 20(6):863–879, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600057.

Guthe:2004:SHC

[GK04] M. Guthe and R. Klein.

Streaming HLODs: an outof-core viewer for network visualization of huge polygon models. Computers and Graphics, 28(1):43–50, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Govindaraju:2007:FCC

[GKLM07] Naga K. Govindaraju, Ilknur Kabul, Ming C. Lin, and Dinesh Manocha. Fast continuous collision detection among deformable models using graphics processors. Computers and Graphics, 31(1):5–14, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930600183X.

Goncalves:2002:DBA

[GKT02] Luiz Gonçalves, Marcelo Kallmann, and Daniel Thalmann. Defining behaviors for autonomous agents based on local perception and smart objects. Computers and Graphics, 26(6): 887–897, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guo:2024:HIA

[GKW^{+}24] Congwei Guo, Yongken Ke, Zhenkai Wan, Minrui Jia, Kai Wang, and Shuai Yang. Human image animation via semantic guidance. Computers and Graphics, 118 (??):102–110, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003035.

Gemballa:1983:MWB

[GL83] Renate Gemballa and Rolf Lindner. The multiple write BUS technique. A most promising technique for speeding up raster scan computation. Computers and Graphics, 7(2):189–191, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grimm:2010:F

[GL10] Cindy Grimm and Joseph J.

LaViola, Jr. Foreword. Computers and Graphics, 34(5): 481, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001111.

Gunpinar:2023:EMC

[GLA23] Erkan Gunpinar, Marco Livesu, and Marco Attene. Exploration of 3D motorcycle complexes from hexahedral meshes. Computers and Graphics, 114(??): 105–115, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000948.

Goude:2020:TMH

[GLC20] Ific Goudé, Jérémy Lacoche, and Rémi Cozot. Tone mapping high dynamic 3D scenes with global lightness coherency. Computers and Graphics, 91(??):243– 251, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301229.

Gross:1997:TMW

[GLDH97] M. H. Gross, L. Lippert, R. Dittrich, and S. Häring. Two methods for wavelet-based volume rendering. Computers and Graphics, 21(2): 237–252, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600087.

Gross:1995:NMA

[GLDK95] M. H. Gross, L. Lippert, A. Dreger, and R. Koch. A new method to approximate the volume-rendering equation using wavelet bases and piecewise polynomials. Computers and Graphics, 19(1):47–62, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400121.

Gross:1997:CGB

[GLDK97] M. H. Gross, L. Lippert, A. Dreger, and R. Koch. Computers and Graphics Best Paper Award (1995). Computers and Graphics, 21 (1):1–??, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gourlay:2000:VRT

[GLL00] D. Gourlay, K. C. Lun, and Guan Liya. Virtual reality and telemedicine for home health care. Computers and Graphics, 24 (5):695–699, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/30/article.
pdf.

Guingo:2020:CAT

[GLS^{+}20] Geoffrey Guingo, Frédéric Larue, Basile Sauvage, Nicolas Lutz, Jean-Michel Dischler, and Marie-Paule Cani. Content-aware texture deformation with dynamic control. Computers and Graphics, 91(??):95– 107, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301023.

Guan:1997:SAP

[GLT^{+}97] Zhidong Guan, Jin Ling, Ning Tao, Xi Ping, and Rongxi Tang. Study and application of physicsbased deformable curves and surfaces. Computers and Graphics, 21(3): 305–313, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=3&aid=9700007.

Glynn:1991:EGM

[Gly91] Earl F. Glynn. The evolution of the Gingerbread Man. Computers and Graphics, 15(4):579–582, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guan:2021:PLL

[GLZ^{+}21] Boliang Guan, Hanhui Li, Fan Zhou, Shujin Lin, and Ruomei Wang. LGCPNet: Local-global combined pointbased network for shape segmentation. Computers and Graphics, 97(??):

208–216, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000704.

Galligani:1978:ISE

[GM78] I. Galligani and L. Moltedo.

An interactive system for experimental data processing. Computers and Graphics, 3(4):135–140, 1978. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gagalowicz:1986:MDS

[GM86] Andre Gagalowicz and Song De Ma. Model driven synthesis of natural textures for 3-D scenes. Computers and Graphics, 10(2): 161–170, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gamito:2002:TSA

[GM02] Manuel N. Gamito and F. Kenton Musgrave. Technical section: An accurate model of wave refraction over shallow water. Computers and Graphics, 26(2):291– 307, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/36/abstract.
html.

Gobbetti:2004:LPC

[GM04] Enrico Gobbetti and Fabio Marton. Layered point clouds: a simple and efficient multiresolution structure for distributing and rendering gigantic pointsampled models. Computers and Graphics, 28(6): 815–826, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garcia:2005:VDD

[GM05] Inmaculada García and Ramón Mollá. Videogames decoupled discrete event simulation. Computers and Graphics, 29(2):195– 202, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ganovelli:2006:I

[GM06] Fabio Ganovelli and Cesar Mendoza. Introduction. Computers and Graphics, 30(3):407, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000628.

Gamito:2007:PRR

[GM07] Manuel N. Gamito and Steve C. Maddock. Progressive refinement rendering of implicit surfaces. Computers and Graphics, 31 (5):698–709, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001197.

Gomes:2013:EOR

[GMd^{+}13] Rafael Beserra Gomes, Bruno Marques Ferreira da Silva, Lourena Karin de Medeiros Rocha, Rafael Vidal Aroca, Luiz Carlos Pacheco Rodrigues Velho, and Luiz Marcos Garcia Gonçalves. Efficient 3D object recognition using foveated point clouds. Computers and Graphics, 37 (5):496–508, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000459.

Goncalves:2018:FSSb

[GMM18] Alexandrino Gonçalves, Luís Gonzaga Magalhães, and Pedro Miguel Moreira. Foreword to the special section on the 24th Portuguese meeting of computer graphics and interaction (EPCGI’17). Computers and Graphics, 77(??): A9–A10, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301729.

Goncalves:2023:RDL

[GMM^{+}23] Guilherme Gonçalves, Miguel Melo, Pedro Monteiro, Hugo Coelho, and Maximino Bessa. The role of different light settings on the perception of realism in virtual replicas in immersive virtual reality. Computers and Graphics, 117(??): 172–182, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002583.

Glencross:2021:FSS

[GMMP21] Mashhuda Glencross, Kenny Mitchell, Billinghurst Mark, and Ye Pan. Foreword to the special section on the Reality-Virtuality Continuum and its Applications (RVCA). Computers and Graphics, 97(??): A3–A4, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000947.

Grim:1993:SRP

[GMNS93] Patrick Grim, Gary Mar, Matthew Neiger, and Paul St.Denis. Self-reference and paradox in two and three dimensions. Computers and Graphics, 17(5):609–612, September–October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garcia:2022:FBC

[GMP22] Francisco González García, Ignacio Martin, and Gustavo Patow. Feature-based clustered geometry for interpolated ray-casting. Computers and Graphics, 102 (??):175–186, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001837.

Gujar:1980:AIL

[GN80] U. G. Gujar and A. R.

Nagesh. An APL implementation of level 2 CORE graphics system. Computers and Graphics, 5(1):1–12, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gujar:1989:CSO

[GN89] Uday G. Gujar and I. V.

Nagendra. Construction of 3D solid objects from orthographic views. Computers and Graphics, 13(4): 505–521, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gerfelder:1994:ICI

[GN94] Norbert Gerfelder and Christian Neuss. The image communication information board: a hypermedia information system for standards and activities in the area of image communication. Computers and Graphics, 18(1):35–46, January– February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gumhold:2004:IST

[GN04] S. Gumhold and Stefan Noll. Introduction to situation and task awareness computing. Computers and Graphics, 28(1):1– 2, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gnatz:1983:GE

[Gna83] R. Gnatz. Graphics and education. Computers and Graphics, 7(1):101–102, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gnatz:1984:AFF

[Gna84] R. Gnatz. Approaching a formal framework for graphics software standards. Computers and Graphics, 8(1): 39–50, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guo:2015:VDL

[GNL^{+}15] Dongliang Guo, Junlan Nie, Meng Liang, Yu Wang, Yanfen Wang, and Zhengping Hu. View-dependent levelof-detail abstraction for interactive atomistic visualization of biological structures. Computers and Graphics, 52 (??):62–71, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500093X.

Gallop:1985:EIG

[GO85] J. R. Gallop and C. D.

Osland. Experiences with implementing GKS on a PERQ and other computers. Computers and Graphics, 9(1):9–17, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gudukbay:1990:FFS

[GO90] Ugur Gudukbay and Bulent Ozguc. Free-form solid modeling using deformations. Computers and Graphics, 14 (3–4):491–500, 1990. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Giang:2006:ACR

[GO06] Thanh Giang and Carol O’Sullivan. Approximate collision response using closest feature maps. Computers and Graphics, 30(3): 423–431, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000653.

Garre:2010:HRO

[GO10] Carlos Garre and Miguel A. Otaduy. Haptic rendering of objects with rigid and deformable parts. Computers and Graphics, 34(6): 689–697, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001317.

Gobel:1993:VRDa

[Gob93a] Martin Gobel. Virtual reality demonstration centre. Computers and Graphics, 17 (3):333–??, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gobel:1993:VRG

[Gob93b] Martin Gobel. Virtual reality: Guest Editor’s introduction. Computers and Graphics, 17(6):625–??, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gama:2023:OML

[GOdSC23] Pedro Henrique Targino Gama, Hugo Oliveira, Jefersson A. dos Santos, and Roberto M. Cesar. An overview on metalearning approaches for fewshot weakly-supervised segmentation. Computers and Graphics, 113(??):77– 88, June 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000651.

Goertzel:1995:RGS

[Goe95] Ben Goertzel. Rapid generation of strange attractors with the eugenic genetic algorithm. Computers and Graphics, 19(1): 151–156, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400130.

Gohring:1984:TPI

[Goh84] Rolf Gohring. A training program for the introduction of CAD in a company. Computers and Graphics, 8 (3):325–328, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gomez:1985:TAS

[Gom85] J. E. Gomez. TWIXT: a 3D animation system. Computers and Graphics, 9(3): 291–298, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gomes:1990:CGP

[Gom90] Mario Rui Gomes. Computer graphics in Portugal. Computers and Graphics, 14 (1):5–??, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gomes:2014:CAP

[Gom14] Abel J. P. Gomes. A continuation algorithm for planar implicit curves with singularities. Computers and Graphics, 38(??):365– 373, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001854.

GonzalezRodriguez:1998:TRM

[Gon98] Jose Angel Gonzalez Rodriguez. A tutorial and recipe for moving fractal trees. Computers and Graphics, 22(2–3):301–305, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/548.pdf.

Gudukbay:1993:ASR

[GÖT93] Ug̃ur Güdükbay, Bülent Özgüç, and Yilmaz Tokad. An animation system for rigid and deformable models. Computers and Graphics, 17(1):71–77, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gudukbay:1997:SFF

[GÖT97] Ugur Güdükbay, Bülent Özgüç, and Yilmaz Tokad. A spring force formulation for elastically deformable models. Computers and Graphics, 21(3): 335–346, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700011.

Guid:1995:SIM

[GOZ95] Nikola Guid, Crtomir Oblonsek, and Borut Zalik. Surface interrogation methods. Computers and Graphics, 19(4):557–574, July– August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500034.

Ghosal:1986:PPS

[GP86] Dipak Ghosal and L. M.

Patnaik. Parallel polygon scan conversion algorithms: Performance evaluation on a shared bus architecture. Computers and Graphics, 10(1):7–25, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ghazanfarpour:1991:HQF

[GP91] D. Ghazanfarpour and B. Peroche. A high-quality filtering using forward texture mapping. Computers and Graphics, 15(4): 569–577, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gangnet:1984:PMP

[GPC84] M. Gangnet, D. Perny, and P. Coueignoux. Perspective mapping of planar textures. Computers and Graphics, 8 (2):115–123, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gimeno:2017:CTI

[GPC^{+}17] Jesús Gimeno, Cristina Portalés, Inmaculada Coma, Marcos Fernández, and Bibiana Martínez. Combining traditional and indirect augmented reality for indoor crowded environments. A case study on the Casa Batlló museum. Computers and Graphics, 69(??):92– 103, December 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730153X.

Gunther:1995:VMP

[GPR^{+}95] T. Günther, C. Poliwoda, C. Reinhart, J. Hesser, R. Männer, H.-P. Meinzer, and H.-J. Baur. VIRIM: a massively parallel processor for real-time volume visualization in medicine. Computers and Graphics, 19(5):705–710, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500049.

Giachetti:2018:FSS

[GPS18] Andrea Giachetti, Paolo Pingi, and Filippo Stanco. Foreword to the Special Section on Smart Tools and Applications in Computer Graphics 2017. Computers and Graphics, 74(??): A6–A7, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300682.

Gee:2002:TSN

[GPTB02] Andrew Gee, Richard Prager, Graham Treece, and Laurence Berman. Technical section: Narrow-band volume rendering for freehand 3D ultrasound. Computers and Graphics, 26(3):463– 476, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/54/35/abstract.
                                        html.

Glasser:2010:VAA

[GPTP10] Sylvia Glaßer, Uta Preim, Klaus Tönnies, and Bernhard Preim. A visual analytics approach to diagnosis of breast DCE-MRI data. Computers and Graphics, 34 (5):602–611, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000841.

Gopalsamy:1993:EEF

[GR93] S. Gopalsamy and T. S. Reddy. Energy and error function minimisation for computation of optimal shape parameters. Computers and Graphics, 17(4): 403–405, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

German:2009:ISB

[GR09] Daniel M. German and Jaume Rigau. Improving scans of black and white photographs by recovering the print maker’s artistic intent. Computers and Graphics, 33 (4):509–520, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000624.

Grabowski:1983:GEI

[Gra83] H. Grabowski. Guest Editor’s introduction: Computers and graphics in mechanical engineering. Computers and Graphics, 7(2): 109–110, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849383900018.

Grabowski:1985:UCG

[Gra85] H. Grabowski. Use of computer graphics instruments in application-oriented training. Computers and Graphics, 9(1):71–77, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grave:1993:DVF

[Gra93] Michel Grave. Distributed visualization in flow simulations. Computers and Graphics, 17(1):9–14, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Graf:2002:TSP

[Gra02] Frank Graf. Technical section: Providing security for eLearning. Computers and Graphics, 26(2):355– 365, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/41/abstract.
html.

Green:1996:SVE

[Gre96] Mark Green. Shared virtual environments: The implications for tool builders. Computers and Graphics, 20(2): 185–189, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500124.

Gagliardi:2022:SPL

[GRF^{+}22] Luca Gagliardi, Andrea Raffo, Ulderico Fugacci, Silvia Biasotti, Walter Rocchia, Hao Huang, Boulbaba Ben Amor, Yi Fang, Yuanyuan Zhang, Xiao Wang, Charles Christoffer, Daisuke Kihara, Apostolos Axenopoulos, Stelios Mylonas, and Petros Daras. SHREC 2022: Protein-ligand binding site recognition. Computers and Graphics, 107 (??):20–31, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001236.

Grinstein:1988:ULC

[Gri88] Georges Grinstein. University of Lowell computer graphics. Computers and Graphics, 12(1):125–126, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grinstein:1993:GEI

[Gri93] Georges Grinstein. Guest Editor’s introduction. Computers and Graphics, 17(1): 3–??, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gerl:2012:SAI

[GRIG12] Moritz Gerl, Peter Rautek, Tobias Isenberg, and Eduard Gröller. Semantics by analogy for illustrative volume visualization. Computers and Graphics, 36(3): 201–213, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931100152X.

Gross:1991:AVE

[Gro91] Markus Groß. The analysis of visibility — environmental interactions between computer graphics, physics, and physiology. Computers and Graphics, 15(3): 407–415, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gross:1992:IAA

[Gro92] Markus Groß. Image analysis for advertisement purposes: a computational model of visual perception. Computers and Graphics, 16 (2):213–221, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Groeller:1994:MRN

[Gro94] Eduard Groeller. Modeling and rendering of nonlinear iterated function systems. Computers and Graphics, 18(5):739–748, September– October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garcia:2008:MFL

[GRPR08] Marcos García, Oscar D. Robles, Luis Pastor, and Angel Rodríguez. MSRS: a fast linear solver for the real-time simulation of deformable objects. Computers and Graphics, 32(3): 293–306, June 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000162.

Gruber:1987:SID

[Gru87] Wolfgang Gruber. Structures of an integrated document workstation. Computers and Graphics, 11(4): 377–387, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gerstner:2000:EIM

[GRW00] Thomas Gerstner, Martin Rumpf, and Ulrich Weikard. Error indicators for multilevel visualization and computing on nested grids. Computers and Graphics, 24(3): 363–373, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/32/article.
pdf.

Grabowski:1983:PDR

[GS83] H. Grabowski and W. Seiler.

Preliminary design requirements to communication processing in CAD. Computers and Graphics, 7(2): 111–123, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Goldberg:1984:CAC

[GS84] T. Goldberg and G. F.

Schrack. Computer-aided correlation of music and graphic image. Computers and Graphics, 8(1):109–111, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gujar:1987:SOC

[GS87] Uday G. Gujar and Joseph W. P. So. A student oriented CAD system in APL. Computers and Graphics, 11 (4):483–495, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guerrini:1989:IRA

[GS89] C. Guerrini and G. Spaletta. Image reconstruction algorithm in tomography. A version for the Cray X-MP vector computer. Computers and Graphics, 13(3): 367–372, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gobel:1993:VRDb

[GS93] Martin Göbel and Jiaoying Shi. The virtual reality demonstration centre. Computers and Graphics, 17(6):627–631, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gobel:1996:GEI

[GS96] Martin Göbel and Jiaoying Shi. Guest Editors’ introduction. Computers and Graphics, 20(2):181–??, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Goodman-Strauss:1999:DST

[GS99] Chaim Goodman-Strauss.

Dodecafoam and substitution tilings. Computers and Graphics, 23(6):917– 924, December 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
46/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/46/article.
pdf.

Goncalves:2001:CML

[GS01a] Luiz M. G. Gonçalves and Fernando W. V. Silva. Control mechanisms and local perception to support autonomous behavior in virtual animated agents. Computers and Graphics, 25(6):965–982, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/31/abstract.
html.

Gotsman:2001:GIF

[GS01b] Craig Gotsman and Vitaly Surazhsky. Guaranteed intersection-free polygon morphing. Computers and Graphics, 25(1):67– 75, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/26/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/32/article.
                                        pdf.

Guthe:2004:ATH

[GS04] S. Guthe and W. Strasser. Advanced techniques for high-quality multi-resolution volume rendering. Computers and Graphics, 28 (1):51–58, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Geigel:2005:UPM

[GS05] Joe Geigel and Nan C. Schaller. Using photography as a metaphor for teaching image synthesis. Computers and Graphics, 29(2):257–265, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grimm:2011:SCN

[GS11] Cindy Grimm and William D. Smart. Shape classification and normal estimation for non-uniformly sampled, noisy point data. Computers and Graphics, 35 (4):904–915, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000756.

Gargantini:1989:ADL

[GSA89] I. Gargantini, G. Schrack, and H. H. Atkinson. Adaptive display of linear octrees. Computers and Graphics, 13(3):337–343, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gumilar:2021:CSI

[GSB^{+}21] Ihshan Gumilar, Ekansh Sareen, Reed Bell, Augustus Stone, Ashkan Hayati, Jingwen Mao, Amit Barde, Anubha Gupta, Arindam Dey, Gun Lee, and Mark Billinghurst. A comparative study on inter-brain synchrony in real and virtual environments using hyperscanning. Computers and Graphics, 94(??):62– 75, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301540.

Gotsman:1999:OOC

[GSF99] Craig Gotsman, Oded Sudarsky, and Jeffrey A. Fayman. Optimized occlusion culling using fivedimensional subdivision.

Computers and Graphics, 23 (5):645–654, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
29/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/29/abstract.
html.

Ganglberger:2019:BVA

[GSF^{+}19] Florian Ganglberger, Nicolas Swoboda, Lisa Frauenstein, Joanna Kaczanowska, Wulf Haubensak, and Katja Bühler. BrainTrawler: a visual analytics framework for iterative exploration of heterogeneous big brain data. Computers and Graphics, 82 (??):304–320, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300949.

Guy:2022:FSS

[GSKZ22] Stephen J. Guy, Shinjiro Sueda, Ioannis Karamouzas, and Victor B. Zordan. Foreword to the special section on motion, interaction, and games 2020. Computers and Graphics, 102(??): A3, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002570.

Gutierrez:2006:SAP

[GSMA06] Diego Gutierrez, Francisco J.

Seron, Adolfo Munoz, and Oscar Anson. Simulation of atmospheric phenomena. Computers and Graphics, 30(6):994–1010, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001026.

Goulev:2004:CAE

[GSME04] Petar Goulev, Lisa Stead, Ebrahim Mamdani, and Caroline Evans. Computer aided emotional fashion. Computers and Graphics, 28 (5):657–666, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gumbau:2013:SSB

[GSSK^{+}13] Jesús Gumbau, Mateu Sbert, László Szirmay-Kalos, Miguel Chover, and Carlos González. Smooth shadow boundaries with exponentially warped Gaussian filtering. Computers and Graphics, 37(3): 214–224, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200180X.

Gomes:2018:RTD

[GSV^{+}18] George A. M. Gomes, Emanuele Santos, Creto A.

Vidal, Ticiana L. Coelho da Silva, and Jose Antonio F. Macedo. Real-time discovery of hot routes on trajectory data streams using interactive visualization based on GPU. Computers and Graphics, 76(??): 129–141, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301420.

Gregory:1994:ICD

[GSY94] John A. Gregory, Muhammad Sarfraz, and P. K. Yuen. Interactive curve design using C^{2} rational splines. Computers and Graphics, 18(2): 153–159, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gomes:1991:FFM

[GT91] Abel J. P. Gomes and Jose C. G. Teixeira. Form feature modelling in a hybrid CSG/BRep scheme. Computers and Graphics, 15(2): 217–229, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Garcia:2018:TTC

[GTdS^{+}18] Rafael Garcia, Alexandru C. Telea, Bruno Castro da Silva, Jim Tørresen, and João Luiz Dihl Comba. A task-and-technique centered survey on visual analytics for deep learning model engineering. Computers and Graphics, 77 (??):30–49, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301535.

Guiducci:2023:LGB

[GTFB23] Niccolò Guiducci, Claudio Tortorici, Claudio Ferrari, and Stefano Berretti. Learning graph-based features for relief patterns classification on mesh manifolds. Computers and Graphics, 115 (??):69–80, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001334.

Glotin:2006:SRM

[GTG06] Hervé Glotin, Sabrina Tollari, and Pascale Giraudet. Shape reasoning on missegmented and mis-labeled objects using approximated Fisher criterion. Computers and Graphics, 30(2): 177–184, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000252.

Gledhill:2003:PIR

[GTTC03] Duke Gledhill, Gui Yun Tian, Dave Taylor, and David Clarke. Panoramic imaging—-a review. Computers and Graphics, 27(3): 435–445, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gudukbay:1997:MJM

[Güd97] Ugur Güdükbay. A movable jaw model for the human face. Computers and Graphics, 21(5):549–554, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=5&aid=9700045.

Geist:1989:PBD

[GV89] Daniel Geist and Michael W. Vannier. PC-based 3-D reconstruction of medical images. Computers and Graphics, 13(2):135–143, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gross:2007:MFE

[GV07] Markus Gross and Christian Voegeli. A multimedia framework for effective language training. Computers and Graphics, 31 (5):761–777, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001665.

Garrison:2020:IVE

[GVC^{+}20] Laura Garrison, Jakub Vašíček, Alexander R. Craven, Renate Grüner, Noeska N. Smit, and Stefan Bruckner. Interactive visual exploration of metabolite ratios in MR spectroscopy studies. Computers and Graphics, 92 (??):1–12, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301199.

Gonzalez:2018:SMM

[GvK18] Diego Gonzalez and Oliver van Kaick. 3D synthesis of man-made objects based on fine-grained parts. Computers and Graphics, 74 (??):150–160, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300815.

Gelas:2009:VIS

[GVPN09] Arnaud Gelas, Sébastien Valette, Rémy Prost, and Wiesław L. Nowinski. Variational implicit surface meshing. Computers and Graphics, 33(3):312–320, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000326.

Grau:2010:EPS

[GVTA10] Sergi Grau, Eduard Vergés, Dani Tost, and Dolors Ayala. Exploration of porous structures with illustrative visualizations. Computers and Graphics, 34(4):398– 408, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000683.

Germs:1999:MVV

[GVVJ99] Rick Germs, Gert Van Maren, Edward Verbree, and Frederik W. Jansen. A multiview VR interface for 3D GIS. Computers and Graphics, 23(4):497–506, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/31/article.
                                        pdf.

Georgii:2006:MFR

[GW06] Joachim Georgii and Rüdiger Westermann. A multigrid framework for real-time simulation of deformable bodies. Computers and Graphics, 30(3):408–415, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600063X.

Gingold:2017:FSS

[GW17] Yotam Gingold and Holger Winnemöller. Foreword to the special section on Expressive 2016. Computers and Graphics, 65 (??):A1, June 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730047X.

Gu:2017:EGB

[GWBD17] Yi Gu, Chaoli Wang, Robert Bixler, and Sidney D’Mello. ETGraph: a graph-based approach for visual analytics of eye-tracking data. Computers and Graphics, 62(??):1–14, ???? 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301340.

Gee:2011:TSW

[GWEA^{+}11] Andrew P. Gee, Matthew Webb, Jorge Escamilla-Ambrosio, Walterio Mayol-Cuevas, and Andrew Calway. A topometric system for wide area augmented reality. Computers and Graphics, 35(4):854– 868, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001099.

Guo:2020:BSS

[GWL^{+}20] Lingchao Guo, Changjian Wang, Fangzhao Li, Hongjun He, and Fen Li. Binary surface smoothing for abnormal lung segmentation. Computers and Graphics, 89(??):68–76, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300625.

Gao:2000:ASM

[GWP00] Shuming Gao, Huagen Wan, and Qunsheng Peng. An approach to solid modeling in a semi-immersive virtual environment. Computers and Graphics, 24(2): 191–202, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/28/article.
                                        pdf.

Grieger:1984:ICM

[GWW84] I. Grieger, G. Wagemann, and G. Wu. Interactive creation and modification of Chinese characters. Computers and Graphics, 8(1):

81–92, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ganglberger:2022:SDD

[GWW^{+}22] Florian Ganglberger, Monika Wißmann, Hsiang-Yun Wu, Nicolas Swoboda, Andreas Thum, Wulf Haubensak, and Katja Bühler. Spatial-data-driven layouting for brain network visualization. Computers and Graphics, 105 (??):12–24, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000668.

Gao:2018:SEB

[GWX^{+}18] Fengyi Gao, Guangshun Wei, Shiqing Xin, Shanshan Gao, and Yuanfeng Zhou. 2D skeleton extraction based on heat equation. Computers and Graphics, 74 (??):99–108, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300645.

Gao:2019:ECS

[GY19] Qiqi Gao and Yasushi Yamaguchi. Extraction of coherent and smooth feature lines from meshes with fine details. Computers and Graphics, 82(??):222– 231, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300810.

Gevins:1975:IAD

[GYD75] A. S. Gevins, C. L. Yeager, and S. L. Diamond. Interactive analysis and display of the electro-encephalogram (EEG) in real time. Computers and Graphics, 1(4): 329–336, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Guo:2015:EMP

[GYJZ15] Jianwei Guo, Dong-Ming Yan, Xiaohong Jia, and Xiaopeng Zhang. Efficient maximal Poisson-disk sampling and remeshing on surfaces. Computers and Graphics, 46(??):72– 79, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001034.

Gao:2023:SPC

[GYK^{+}23] Yang Gao, Honglin Yuan, Tao Ku, Remco C. Veltkamp, Georgios Zamanakos, Lazaros Tsochatzidis, Angelos Amanatiadis, Ioannis Pratikakis, Aliki Panou, Ioannis Romanelis, Vlassis Fotis, Gerasimos Arvanitis, and Konstantinos Moustakas. SHREC 2023: Point cloud change detection for city scenes. Computers and Graphics, 115 (??):35–42, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001206.

Guo:2013:IDM

[GYL^{+}13] Jianwei Guo, Dong-Ming Yan, Er Li, Weiming Dong, Peter Wonka, and Xiaopeng Zhang. Illustrating the disassembly of 3D models. Computers and Graphics, 37 (6):574–581, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000940.

Ge:2023:PLM

[GYY^{+}23] Mengyuan Ge, Junfeng Yao, Baorong Yang, Ningna Wang, Zhonggui Chen, and Xiaohu Guo. Point2MM: Learning medial mesh from point clouds. Computers and Graphics, 115(??):511– 521, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001498.

GomesdeSa:1999:TSV

[GZ99] Antonino Gomes de Sá and Gabriel Zachmann. Technical section — virtual reality as a tool for verification of assembly and maintenance processes. Computers and Graphics, 23(3):389– 403, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        3/690.pdf.

Guo:2021:NRA

[GZL21] Guanhui Guo, Yanni Zou, and Peter X. Liu. A new rendering algorithm based on multi-space for living soft tissue. Computers and Graphics, 98(??):242– 254, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001187.

Guo:2014:EVM

[GZLW14] Yanwen Guo, Guiping Zhang, Zili Lan, and Wenping Wang. Efficient view manipulation for cuboid-structured images. Computers and Graphics, 38(??):174–182, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001799.

Gan:2020:QPC

[GZSZ20] Yuan Gan, Yan Zhang, Zhengxing Sun, and Hao Zhang. Qualitative photo collage by quartet analysis and active learning. Computers and Graphics, 88(??): 35–44, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300236.

Guo:2012:IED

[GZW12] Tiantian Guo, Hui Zhang, and Yamei Wen. An improved example-driven symbol recognition approach in engineering drawings. Computers and Graphics, 36(7): 835–845, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001173.

Gong:2006:RCW

[GZZS06] Jie-Hui Gong, Gui-Fang Zhang, Hui Zhang, and Jia-Guang Sun. Reconstruction of 3D curvilinear wire-frame from three orthographic views. Computers and Graphics, 30(2): 213–224, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600029X.

Haase:2000:MMC

[H^{+}00a] H. Haase et al. Meteorology meets computer graphics a look at a wide range of weather visualisations for diverse audiences. Computers and Graphics, 24(3): 391–397, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/35/article.
pdf.

Hastreiter:2000:RTA

[H^{+}00b] P. Hastreiter et al. Registration techniques for the analysis of the brain shift in neurosurgery. Computers and Graphics, 24(3): 385–389, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/34/article.
pdf.

Hollerer:2001:UIM

[H^{+}01] Tobias Höllerer et al. User interface management techniques for collaborative mobile augmented reality. Computers and Graphics, 25(5):799–810, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
13/20/57/34/34/abstract.
html.

Haas:1994:OV

[Haa94]  Stefan Haas.  Objects

and views. Computers and Graphics, 18(2):193– 199, March–April 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Holmes:1975:MSA

[HAB75] H. H. Holmes, D. M. Austin, and W. H. Benson. The MA-PEDIT system for automatic map digitization. Computers and Graphics, 1(2–3): 251–256, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hu:2020:CGS

[HAL20] Chuanfeng Hu, Jiaming Ai, and Hongwei Lin. Curve guided T-spline skinning for surface and solid generation. Computers and Graphics, 90(??):84–94, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300741.

Heinrich:2021:EDI

[HAL^{+}21] Florian Heinrich, Vikram Apilla, Kai Lawonn, Christian Hansen, Bernhard Preim, and Monique Meuschke.

Estimating depth information of vascular models: a comparative user study between a virtual reality and a desktop application. Computers and Graphics, 98 (??):210–217, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001138.

Hanusa:1984:DMA

[Han84] H. Hanusa. Dialogue monitoring and analysis. Computers and Graphics, 8(1): 103–104, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hansmann:1995:CGC

[Han95] Werner Hansmann. Computer graphics curricula for a wide range of disciplines. Computers and Graphics, 19(2):327–330, March–April 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400160.

Hansmann:1997:SCGa

[Han97a] Werner Hansmann. A survey of computer graphics education at German universities. Computers and Graphics, 21(1):113–116, January– February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600090.

Hansmann:1997:SCGb

[Han97b] Werner Hansmann. A survey of computer graphics in art education in Germany. Computers and Graphics, 21(1):117–119, January– February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600091.

Hardie:1983:VSC

[Har83]  P. F. Hardie.  Volume,

space, and colour. Computers and Graphics, 7(1): 105–107, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hart:2000:CGR

[Har00] George W. Hart. Chaos and graphics — reticulated geodesic constructions. Computers and Graphics, 24 (6):907–910, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/34/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/34/33/article.
                                        pdf.

Hazony:1977:ITC

[Haz77] Yehonathan Hazony. Interactive tools for conceptual programming using APL-Graphics: the integral and derivative operators. Computers and Graphics, 2(4): 209–218, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hazony:1979:APP

[Haz79a] Y. Hazony. Algorithms for parallel processing: Curve and surface definition with Q-splines. Computers and Graphics, 4(3–4):165–176, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hazony:1979:ICA

[Haz79b] Yehonathan Hazony. Interactive cartography — an analytical tool for management information systems. Computers and Graphics, 4 (1):63–75, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hofmann:1991:BCI

[HB91] Georg Rainer Hofmann and Christof Blum. The basic concepts for ISO/IEC’s Image Interchange Format (IIF). Computers and Graphics, 15(4):507–514, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hildebrand:2013:OSA

[HBA13] Kristian Hildebrand, Bernd Bickel, and Marc Alexa. Orthogonal slicing for additive manufacturing. Computers and Graphics, 37 (6):669–675, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300085X.

Heinzl:2014:GEU

[HBG14] Christoph Heinzl, Stefan Bruckner, and Eduard Gröller. Guest editorial — uncertainty and parameter space analysis in visualization. Computers and Graphics, 41(??):A1–A2, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000375.

Hamilton:1982:ICG

[HBJ82] P. S. Hamilton, R. E. Barr, and D. Juricic. Implementation of computer graphics exercises in freshman engineering graphics education. Computers and Graphics, 6(1):29–34, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hernandez-Bautista:2023:DLC

[HBM23] Marina Hernández-Bautista and Francisco Javier Melero. Deep learning of curvature features for shape completion. Computers and Graphics, 115(??):204–215, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300136X.

Haciomeroglu:2013:GAH

[HBOS13] Murat Haciomeroglu, Oner Barut, Cumhur Y. Ozcan, and Hayri Sever. A GPUassisted hybrid model for real-time crowd simulations. Computers and Graphics, 37 (7):862–872, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000800.

Huang:1995:TBB

[HC95] Chi-Yen Huang and Kuo-Liang Chung. Transformations between bincodes and the DF-expression. Computers and Graphics, 19(4): 601–610, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500038.

Hsueh:1991:IRC

[HCC91] Yuang-Chen Hsueh, Ming-Guey Chern, and Chyi-Hwa Chu. Image requantization by cardinality distribution. Computers and Graphics, 15(3):397–405, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hsiao:2013:EAH

[HCC13] Chih-Chieh Hsiao, Slo-Li Chu, and Chen-Yu Chen. Energy-aware hybrid precision selection framework for mobile GPUs. Computers and Graphics, 37 (5):431–444, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000423.

Huang:2016:MAT

[HCLC16] Yi-Jheng Huang, Shu-Yuan Chan, Wen-Chieh Lin, and Shan-Yu Chuang. Making and animating transformable 3D models. Computers and Graphics, 54 (??):127–134, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931500117X.

Hachette:2022:ASA

[HCV^{+}22] Olivier Hachette, Florian Canezin, Rodolphe Vaillant, Nicolas Mellado, and Loïc Barthe. Automatic shape adjustment at joints for the implicit skinning. Computers and Graphics, 102 (??):300–308, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002296.

Han:2023:MPT

[HCX^{+}23] Xianjun Han, Qianqian Chen, Zhaoyang Xie, Xuejun Li, and Hongyu Yang. Multiscale progressive text prompt network for medical image segmentation. Computers and Graphics, 116(??):262–274, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323002170.

Han:2023:USS

[HDZR23] Bing Han, Lixiang Deng, Yi Zheng, and Shuang Ren. UPU-SNet: Siamese network for unsupervised point cloud upsampling based on spatialaware transformers. Computers and Graphics, 115 (??):484–499, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001668.

Herbison-Evans:1980:HMH

[HE80] D. Herbison-Evans. How to merge hidden arcs and then not draw them. Computers and Graphics, 5(2– 4):79–81, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Havran:2015:EBT

[HE15] Vlastimil Havran and Petr Egert. Extensions to bidirectional texture function compression with multi-level vector quantization in OpenCL. Computers and Graphics, 48 (??):1–10, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000060.

Hartung:1998:DWM

[HEG98] Frank Hartung, Peter Eisert, and Bernd Girod. Digital watermarking of MPEG-4 facial animation parameters. Computers and Graphics, 22(4):425–435, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/cag/sub/1998/22/
                                        4/566.pdf.

Hettinga:2022:AIV

[HEK22] Gerben J. Hettinga, Jose Echevarria, and Jirí Kosinka. Adaptive image vectorisation and brushing using mesh colours. Computers and Graphics, 105(??): 119–130, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200067X.

Helmick:1995:VVD

[Hel95] Richard Helmick. Virtues of verisimilitude in design and art. Computers and Graphics, 19(4): 505–507, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=4&aid=9500027.

Herbert:1983:ACG

[Her83a] F. Herbert. The art of computer graphics — computer graphics in art. Computers and Graphics, 7(3– 4):381–383, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Herbert:1983:SMU

[Her83b] F. Herbert. Solid modelling using a volume encoding data structure. Computers and Graphics, 7(1): 97–99, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Herbert:1985:SMA

[Her85] F. Herbert. Solid modeling for architectural design using octpaths. Computers and Graphics, 9(2): 107–116, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Herbert:2018:DCC

[HEW^{+}18] Bradley Herbert, Barrett Ens, Amali Weerasinghe, Mark Billinghurst, and Grant Wigley. Design considerations for combining augmented reality with intelligent tutors. Computers and Graphics, 77(??): 166–182, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301523.

Havemann:2013:CCC

[HEWF13] Sven Havemann, Johannes Edelsbrunner, Philipp Wagner, and Dieter Fellner. Curvature-controlled curve editing using piecewise clothoid curves. Computers and Graphics, 37(6):764–773, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000915.

Hu:1985:PPA

[HF85] Mei-Cheng Hu and James D. Foley. Parallel processing approaches to hidden-surface removal in image space. Computers and Graphics, 9 (3):303–317, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hetherington:2006:IRT

[HFP06] Robina Hetherington, Brian Farrimond, and Steve Presland. Information rich temporal virtual models using X3D. Computers and Graphics, 30(2):287– 298, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000355.

Hollerer:1999:EMD

[HFT^{+}99] Tobias Höllerer, Steven Feiner, Tachio Terauchi, Gus Rashid, and Drexel Hallaway. Exploring MARS: developing indoor and outdoor user interfaces to a mobile augmented reality system. Computers and Graphics, 23 (6):779–785, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/28/article.
pdf.

Hladuvka:2002:VVL

[HG02] Jiří Hladůvka and Eduard Gröller. Visualization of very large datasets: Smallest second-order derivatives for efficient volume-data representation. Computers and Graphics, 26(2):229– 238, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/30/abstract.
html.

Hegde:2021:PPN

[HG21] Sindhu Hegde and Shankar Gangisetty. PIG-Net: Inception based deep learning architecture for 3D point cloud segmentation. Computers and Graphics, 95(??): 13–22, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000042.

Hu:2022:LLA

[HG22] Xuegang Hu and Juelin Gong. LARFNet: Lightweight asymmetric refining fusion network for real-time semantic segmentation. Computers and Graphics, 109 (??):55–64, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001807.

Hulsmann:2018:CME

[HGH^{+}18] Felix Hülsmann, Jan Philip Göpfert, Barbara Hammer, Stefan Kopp, and Mario Botsch. Classification of motor errors to provide realtime feedback for sports coaching in virtual reality

— a case study in squats and Tai Chi pushes. Computers and Graphics, 76 (??):47–59, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301304.

Holloway:2016:TBS

[HGJ16] Michelle Holloway, Cindy Grimm, and Tao Ju. Templatebased surface reconstruction from cross-sections. Computers and Graphics, 58 (??):84–91, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300553.

Ho:2021:NAS

[HGJC21] Kary Ho, Andrew Gilbert, Hailin Jin, and John Collomosse. Neural architecture search for deep image prior. Computers and Graphics, 98 (??):188–196, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001126.

Heinemann:2023:TBC

[HGS23] Birte Heinemann, Sergej Görzen, and Ulrik Schroeder. Teaching the basics of computer graphics in virtual reality. Computers and Graphics, 112(??):1–12, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000304.

Hermosilla:2016:HQI

[HGVV16] P. Hermosilla, V. Guallar, A. Vinacua, and P. P. Vázquez. High quality illustrative effects for molecular rendering. Computers and Graphics, 54(??): 113–120, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500120X.

Heim:2024:AVE

[HGW^{+}24] Anja Heim, Alexander Gall, Manuela Waldner, Eduard Gröller, and Christoph

Heinzl. AccuStripes: Vi-

sual exploration and comparison of univariate data distributions using color and binning. Computers and Graphics, 119(??): ??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000414.

Huang:1988:OOT

[HH88] Samuel D. Huang and Denise M. Hammonds. “to be of one’s own time”: Hightech meets tradition in computer graphics art. Computers and Graphics, 12(1): 133–139, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hsu:1991:AGS

[HH91] Wei-I Hsu and J. L. Hock. An algorithm for the general solution of hidden line removal for intersecting solids. Computers and Graphics, 15 (1):67–86, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hubner:1993:GEI

[HH93] Wolfgang Hubner and Christoph Hornung. Guest editors’ introduction. Computers and Graphics, 17(3):201–??, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hernandez:2017:AFR

[HHCM17] Matthias Hernandez, Tal Hassner, Jongmoo Choi, and Gerard Medioni. Accurate 3D face reconstruction via prior constrained structure from motion. Computers and Graphics, 66 (??):14–22, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300572.

Herz:1997:CPC

[HHG97] Jacky Herz, Roger D. Hersch, and Jakob Gonczarowski. Coherent processing of character skeletal forms. Computers and Graphics, 21 (6):727–736, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700050.

Hernandez:2013:TBS

[HHK^{+}13] Edwin Alexander Peraza Hernandez, Shiyu Hu, Han Wei Kung, Darren Hartl, and Ergun Akleman. Towards building smart selffolding structures. Computers and Graphics, 37 (6):730–742, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000964.

Hohmann:2010:GSG

[HHKF10] Bernhard Hohmann, Sven Havemann, Ulrich Krispel, and Dieter Fellner. A GML shape grammar for semantically enriched 3D building models. Computers and Graphics, 34(4):322– 334, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000749.

He:1999:FAD

[HHL99] Yu-Xin He, YaLing He, and Hua Li. Fast and accurate determination of the spatial boundary of IFS attractors. Computers and Graphics, 23 (4):547–553, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/35/article.
                                        pdf.

Herr:2017:HBP

[HHLE17] Dominik Herr, Qi Han, Steffen Lohmann, and Thomas Ertl. Hierarchy-based projection of high-dimensional labeled data to reduce visual clutter. Computers and Graphics, 62(??):28– 40, ???? 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301388.

Huang:2023:EBF

[HHN^{+}23] Xue Huang, Bin Han, Yaqian Ning, Jie Cao, and Ying Bi. Edge-based feature extraction module for 3D point cloud shape classification. Computers and Graphics, 112(??): 31–39, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000328.

Holzapfel:2022:FDV

[HHZ^{+}22] Constantin Holzapfel, Miriam Hoene, Xinjie Zhao, Chunxiu Hu, Cora Weigert, Andreas Niess, Guowang Xu, Rainer Lehmann, Andreas Dräger, and Michael Krone.

FluxomicsExplorer: Dif-

ferential visual analysis of Flux Sampling based on Metabolomics. Computers and Graphics, 108(??):11– 21, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001583.

Han:2005:GEA

[HIK05] Chang-Young Han, Yeon-Ho Im, and Lee-Sup Kim. Geometry engine architecture with early backface culling hardware. Computers and Graphics, 29(3):415– 425, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hayashi:1983:MMI

[HIS83] I. Hayashi, K. Itoh, and S. Suzuki. Man-machine interactive guidance for urban railway networks passenger information system. Computers and Graphics, 7 (1):59–72, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hourizi:2003:TEP

[HJ03] Rachid Hourizi and Peter Johnson. Towards an explanatory, predictive account of awareness. Computers and Graphics, 27(6): 859–872, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huang:2016:ESD

[HJ16] Zhiyang Huang and Tao Ju. Extrinsically smooth direction fields. Computers and Graphics, 58(??):

109–117, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300589.

Hill:1995:IRV

[HJDR95] Doug L. Hill, Scott Juskiw, Nelson G. Durdle, and V. James Raso. Interactive rendering of volumetric data sets. Computers and Graphics, 19(5):685–693, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500047.

Huynh:1993:POO

[HJL^{+}93] Duong Le Huynh, Morten Jensen, Rolf Larsen, Jonathan Southard, Y. F. Wang, Yulun Wang, and Amante Mangaser. PIX: An objectoriented network graphics environment. Computers and Graphics, 17(3):295– 304, May–June 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hubbold:2007:FSI

[HJL07] Roger Hubbold, Joaquim A.

Jorge, and Ming Lin. Foreword to: “special issue on virtual environments”. Computers and Graphics, 31(1):

2–4, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001798.

Huang:2013:PGB

[HJT^{+}13] Chun-Yen Huang, Wan-Ting Jheng, Wen-Kai Tai, Chin-Chen Chang, and Der-Lor Way. Procedural grape bunch modeling. Computers and Graphics, 37(4): 225–237, June 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000034.

Heo:1997:ETC

[HJW97] Junhyeok Heo, Soonki Jung, and Kwangyun Wohn. Exploiting temporally coherent visibility for accelerated walkthroughs. Computers and Graphics, 21(4): 507–517, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=4&aid=9700026.

Han:2008:MFF

[HJW^{+}08] Wei Han, Zhidong Jin, Li Wang, Min Hu, Zhangye Wang, and Qunsheng Peng. Molecular field feature extraction and analysis with level set method. Computers and Graphics, 32(6): 632–638, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001192.

Herzner:1993:MSM

[HK93] Wolfgang Herzner and Matthias Kummer. MMV — synchronizing multimedia documents: An extension of CDA for synchronization and presentation of multimedia documents. Computers and Graphics, 17(3): 229–241, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hubbold:1999:LNL

[HK99] Roger Hubbold and Martin Keates. Landmarking for navigation of large models. Computers and Graphics, 23 (5):729–738, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
36/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/36/abstract.
html.

Hellenschmidt:2004:SSS

[HK04] Michael Hellenschmidt and Thomas Kirste. Software solutions for self-organizing multimedia-appliances. Computers and Graphics, 28 (5):643–655, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hermann:2015:VAP

[HK15] Max Hermann and Reinhard Klein. A visual analytics perspective on shape analysis: State of the art and future prospects. Computers and Graphics, 53 (part A) (??):63–71, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001429.

Herholz:2017:UMG

[HKBA17] Philipp Herholz, Sebastian Koch, Tamy Boubekeur, and Marc Alexa. Unsharp masking geometry improves 3D prints. Computers and Graphics, 66 (??):135–142, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300675.

Harding:2002:VVL

[HKCL02] Chris Harding, Ioannis A. Kakadiaris, John F. Casey, and R. Bowen Loftin. Visualization of very large datasets: a multi-sensory system for the investigation of geoscientific data. Computers and Graphics, 26(2):259– 269, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/33/abstract.
html.

Heckel:2011:IMI

[HKHP11] Frank Heckel, Olaf Konrad, Horst Karl Hahn, and Heinz-Otto Peitgen. Interactive 3D medical image segmentation with energy-minimizing implicit functions. Computers and Graphics, 35(2): 275–287, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001925.

Hao:2023:HBH

[HKL^{+}23] Chenhui Hao, Dehui Kong, Jinghua Li, Caixia Liu, and

Baocai Yin. HyperGraph

based human mesh hierarchical representation and reconstruction from a single image. Computers and Graphics, 115(??):339– 347, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001395.

Hartl:1998:CAC

[HKPL98] Martin Hartl, Ivan Křupka, Radek Poliščuk, and Miroslav Liška. Computer-aided chromatic interferometry. Computers and Graphics, 22(2– 3):203–208, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/2-3/542.pdf.

Hildebrand:2000:CSI

[HKS00] A. Hildebrand, M. H. Kim, and G. Sakas. C&G special issue advanced and dynamic medical visualisation. Computers and Graphics, 24 (5):657–659, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/26/article.
pdf.

Herfet:2001:EMA

[HKS01] Thorsten Herfet, Thomas Kirste, and Michael Schnaider. EMBASSI multimodal assistance for infotainment and service infrastructures. Computers and Graphics, 25 (4):581–592, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/33/31/abstract.
html.

Hirayama:2001:AIM

[HKYM01] H. Hirayama, K. Kaneda, H. Yamashita, and Y. Monden. An accurate illumination model for objects coated with multilayer films. Computers and Graphics, 25(3):391–400, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/30/article.
pdf.

Herter:1993:ADD

[HL93] Thomas Herter and Klaus Lott. Algorithms for decomposing 3-D orthogonal matrices into primitive rotations. Computers and Graphics, 17(5):517–527, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hu:1996:SVG

[HL96] Zeng-Jia Hu and Zhi-Kui Ling. Swept volumes generated by the natural quadric surfaces. Computers and Graphics, 20(2): 263–274, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=20&
                                        issue=2&aid=9500100.

Han:1997:EDE

[HL97] Wei-Yu Han and Ja-Chen Lin. Edge detection and edge-preserved compression for error- diffused images. Computers and Graphics, 21(6):757–767, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700055.

Hunkins:2002:ARR

[HL02] Dalton Hunkins and David B. Levine. Additional rich resources for computer graphics educators. Computers and Graphics, 26(4): 609–614, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/55/41/abstract.
                                        html.

Hui:2006:EMB

[HL06] K. C. Hui and H. C. Leung. Expression modeling — a boundary element approach. Computers and Graphics, 30 (6):981–993, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600152X.

Huang:2006:ELD

[HLB^{+}06] Jin Huang, Xinguo Liu, Hujun Bao, Baining Guo, and Heung-Yeung Shum. An efficient large deformation method using domain decomposition. Computers and Graphics, 30(6): 927–935, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001488.

Hernandez-Luna:1988:ODC

[HLCF88] Alberto A. Hernandez-Luna and Luis Cardenas-Franco. Optimal design of class molds using CAD/CAE and response surface methodology techniques. Computers and Graphics, 12(3–4): 391–399, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liao:2009:GFB

[hLfTxDdZ09] Sheng hui Liao, Ruo feng Tong, Jin xiang Dong, and Fu dong Zhu. Gradient field based inhomogeneous volumetric mesh deformation for maxillofacial surgery simulation. Computers and Graphics, 33(3): 424–432, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000351.

Hu:2018:WBP

[HLL^{+}18] Ling Hu, Qinsong Li, Shengjun Liu, Xinru Liu, and Zheng Wang. Waveletbased polygon soup consolidation. Computers and Graphics, 70(??):39– 50, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301218.

Henz:2016:ICC

[HLO16] Bernardo Henz, Frederico A. Limberger, and Manuel M. Oliveira. Independent color-channel adjustment for seamless cloning based on Laplacian-membrane modulation. Computers and Graphics, 57(??):46–54, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849316300176.

Huang:1989:DHF

[HLS89] Samuel D. Huang, Raul Lopez, and Jan Siemieniewski. Discovery of hidden facial images in the paintings of Vincent Van Gogh and Paul Gauguin by using computer graphics enhancement techniques. Computers and Graphics, 13(4):

571–578, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huang:2023:TPT

[HLXL23] Huiwen Huang, Xiaonan Luo, Songhua Xu, and Youxing Li. Twin pseudo-training for semi-supervised semantic segmentation. Computers and Graphics, 115 (??):348–358, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001383.

He:2019:DDH

[HLY^{+}19] Huayun He, Guiqing Li, Zehao Ye, Aihua Mao, Chuhua Xian, and Yongwei Nie. Data-driven 3D human head reconstruction. Computers and Graphics, 80(??): 85–96, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300317.

Hu:2017:ETM

[HLZ^{+}17] Shaojun Hu, Zhengrong Li, Zhiyi Zhang, Dongjian He, and Michael Wimmer. Efficient tree modeling from airborne LiDAR point clouds. Computers and Graphics, 67(??):1–13, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300420.

Haedecke:2023:SVA

[HMA23] Elena Haedecke, Michael Mock, and Maram Akila. ScrutinAI: a visual analytics tool supporting semantic assessments of object detection models. Computers and Graphics, 114 (??):265–275, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300105X.

Hildebrand:1995:TVC

[HMdM^{+}95] A. Hildebrand, L. P. Magalhães, J. M. de Martino, F. Seibert, R. Strack, C. L. Tozzi, and S. T. Wu. Towards a visual computing and communication reference model. Computers and Graphics, 19(1):141–149, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1&aid=9400129.

Herzog:1998:GID

[HMHA98] O. Herzog, A. Miene, Th. Hermes, and P. Alshuth. Graphics in/for digital libraries — integrated information mining for texts, images, and videos. Computers and Graphics, 22(6):675–685, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/622.pdf.

Hubo:2008:SSB

[HMHB08] Erik Hubo, Tom Mertens, Tom Haber, and Philippe Bekaert. Self-similarity based compression of point set surfaces with application to ray tracing. Computers and Graphics, 32(2): 221–234, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000241.

Hatfield:1991:MEL

[HMW91] Donald Hatfield, Richard A.

Miner, and C. Thomas Wilkes. A mathematical expression language for imaging applications. Computers and Graphics, 15(4): 495–506, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hourcade:1985:AAC

[HN85] J. C. Hourcade and A. Nicolas. Algorithms for antialiased cast shadows. Computers and Graphics, 9(3): 259–265, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hauenstein:2020:DEM

[HN20] Jacob D. Hauenstein and Timothy S. Newman. Descriptions and evaluations of methods for determining surface curvature in volumetric data. Computers and Graphics, 86(??):52– 70, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301724.

Higa:1988:CGD

[HO88] A. Afuso Higa and M. Vega Ortiz. Computer graphics in distribution system planning. Computers and Graphics, 12(3–4):361–363, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hodgkinson:1994:CIM

[HO94] Ian J. Hodgkinson and R. P. O’Shea. Constraints imposed by Mach bands on shape from shading. Computers and Graphics, 18(4): 531–536, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hachaj:2012:VPA

[HO12] Tomasz Hachaj and Marek R.

Ogiela. Visualization of perfusion abnormalities with GPU-based volume rendering. Computers and Graphics, 36(3):163–169, May 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849312000039.

Hu:2007:RMT

[HOCN07] G. H. Hu, S. K. Ong, Y. P.

Chen, and A. Y. C. Nee. Reflectance modeling for a textured object under uncontrolled illumination from high dynamic range maps. Computers and Graphics, 31(2):262–270, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000064.

Hodgkinson:1991:MBO

[Hod91] Ian J. Hodgkinson. Mach band optical ray illusion

— drawing with the second differential. Computers and Graphics, 15(4): 545–547, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hollasch:1994:AAR

[Hol94] Steven Hollasch. Advanced animation and rendering techniques by Alan Watt and Mark Watt. Computers and Graphics, 18(2): 249–??, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Holmes:2003:WDC

[Hol03] Tiffany Holmes. What do computers eat? Teaching beginners to think critically about technology and art. Computers and Graphics, 27(3):361–368, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hooper:1991:NSI

[Hoo91] Kenneth J. Hooper. Note on some internal structures of the Mandelbrot set. Computers and Graphics, 15(2): 295–297, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hopgood:1986:ETA

[Hop86] Bob Hopgood. Eurographics 86 tutorials: August 25–26 1986. Computers and Graphics, 10(1):81–83, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900725.

Hoppe:1998:EIP

[Hop98] Hugues Hoppe. Efficient implementation of progressive meshes. Computers and Graphics, 22(1):27–36, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
1/512.pdf.

Hornung:1982:ACM

[Hor82] C. Hornung. An approach to a calculation-minimized hidden line algorithm. Computers and Graphics, 6(3): 121–126, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Horak:1983:IMT

[Hor83]  W. Horak.  Interchang-

ing mixed text image documents in the office environment. Computers and Graphics, 7(1):13–29, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Howden:1975:SPI

[How75]  W. E. Howden.  Solu-

tion plans and interactive problem solving. Computers and Graphics, 1(1):21– 26, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Howie:1979:GST

[How79] J. M. Howie. GD3 — a system for two-dimensional computer graphics. Computers and Graphics, 4(3– 4):161–164, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Howard:1988:SCD

[How88] Toby Howard. A shareable centralised database for KRT^{3}: a hierarchical graphics system based on PHIGS. Computers and Graphics, 12 (2):201–211, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Holden:1991:GIS

[HP91] Arun V. Holden and A. V. Panfilov. Graphical identification of spatio-temporal chaos. Computers and Graphics, 15(2):301–302, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hornby:2001:ESG

[HP01] Gregory S. Hornby and Jordan B. Pollack. Evolving Lsystems to generate virtual creatures. Computers and Graphics, 25(6):1041–1048, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/36/abstract.
html.

Havran:2003:CRS

[HP03] Vlastimil Havran and Werner Purgathofer. On comparing ray shooting algorithms. Computers and Graphics, 27 (4):593–604, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hollemeersch:2010:IIE

[HPD^{+}10] Charles-Frederik Hollemeersch, Bart Pieters, Aljosha Demeulemeester, Frederik Cornillie, Bert Van Semmertier, Erik Mannens, Peter Lambert, Piet Desmet, and Rik Van de Walle. Infinitex: an interactive editing system for the production of large texture data sets. Computers and Graphics, 34 (6):643–654, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001457.

Hong:2019:TOS

[HPKE19] Q Youn Hong, Youngjin Park, Myung-Soo Kim, and Gershon Elber. Trimming offset surface selfintersections around nearsingular regions. Computers and Graphics, 82(??): 84–94, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300779.

Hu:2014:NPM

[HQW14] Yong Hu, Yue Qi, and Shan Wang. Non-parametric modeling and editing for captured SVBRDF. Computers and Graphics, 39(??): 74–88, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931300143X.

Herman:1988:MIG

[HR88] Ivan Herman and Janos Reviczky. A means to improve the GKS-3D/PHIGS output pipeline implementation. Computers and Graphics, 12(2):191–199, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Haanpaa:1997:AHS

[HR97] Douglas P. Haanpaa and Gerald P. Roston. An advanced haptic system for improving man-machine interfaces. Computers and Graphics, 21(4):443–449, July–August 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=4&aid=9700017.

Hewgill:2004:PTS

[HR04] Adam Hewgill and Brian J.

Ross. Procedural 3D texture synthesis using genetic programming. Computers and Graphics, 28(4):569– 584, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hou:2007:CCS

[HR07] Suyu Hou and Karthik Ramani. Classifier combination for sketch-based 3D part retrieval. Computers and Graphics, 31(4):598– 609, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000970.

Hopgood:1988:CGR

[HRGD88] F. R. A. Hopgood, K. Robinson, J. R. Gallop, and D. A. Duce. Computer graphics at Rutherford Appleton laboratory. Computers and Graphics, 12(2):285–287, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Herman:1986:CGM

[HRTK86] Iván Herman, János Reviczky, and Tibor Tolnay-Knefély. A concept for a GKS machine. Computers and Graphics, 10(2): 177–182, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hersch:1997:GEI

[HS97] Roger D. Hersch and Jurgen Schonhut. Guest Editors’ introduction. Computers and Graphics, 21(6):689–??, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huttner:1999:FTV

[HS99] Tobias Hüttner and Wolfgang Strasser. FlyAway: a 3D terrain visualization system using multiresolution principles. Computers and Graphics, 23(4):479– 485, August 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/29/article.
pdf.

Hitchner:2000:ACG

[HS00] Lewis E. Hitchner and Henry A. Sowizral. Adapting computer graphics curricula to changes in graphics. Computers and Graphics, 24(2):283–288, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/13/20/47/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/36/article.
pdf.

Hanisch:2003:AIF

[HS03] Frank Hanisch and Wolfgang Straßer. Adaptability and interoperability in the field of highly interactive Web-based courseware. Computers and Graphics, 27 (4):647–655, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hanisch:2005:HIV

[HS05] Frank Hanisch and Wolfgang Straßer. How to include visuals and interactivities in an educational computer graphics repository. Computers and Graphics, 29(2):237–243, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hong:2008:PRS

[HS08] Yuan Hong and Han-Wei Shen. Parallel reflective symmetry transformation for volume data. Computers and Graphics, 32 (1):41–54, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700204X.

Herold:2011:STS

[HS11] James Herold and Thomas F. Stahovich. SpeedSeg: a technique for segmenting pen strokes using pen speed. Computers and Graphics, 35(2):250–264, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001895.

Herold:2014:MLA

[HS14] James Herold and Thomas F. Stahovich. A machine learning approach to automatic stroke segmentation. Computers and Graphics, 38 (??):357–364, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001465.

Hoang:2010:VIW

[HSB^{+}10] Roger V. Hoang, Matthew R. Sgambati, Timothy J. Brown, Daniel S. Coming, and Frederick C. Harris, Jr. VFire: Immersive wildfire simulation and visualization. Computers and Graphics, 34(6): 655–664, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001470.

Haase:1996:VTI

[HSD96] Helmut Haase, Johannes Strassner, and Fan Dai. VR techniques for the investigation of molecule data. Computers and Graphics, 20(2): 207–217, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500127.

Hilsmann:2010:RCA

[HSE10] Anna Hilsmann, David C.

Schneider, and Peter Eisert. Realistic cloth augmentation in single view video under occlusions. Computers and Graphics, 34 (5):567–574, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931000083X.

Hasler:2009:EBS

[HSR^{+}09] Nils Hasler, Carsten Stoll, Bodo Rosenhahn, Thorsten Thormählen, and Hans-Peter Seidel. Estimating body shape of dressed humans. Computers and Graphics, 33(3):211–216, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000399.

Hinkenjann:2020:FSS

[HSTR20] André Hinkenjann, Christian Sandor, João Marcelo Teixeira, and Rafael Rieder. Foreword to the special section on the Symposium on Virtual and Augmented Reality 2019 (SVR 2019). Computers and Graphics, 86 (??):A3–A4, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300066.

Hasegawa:1996:SSP

[HT96] Júlio Kiyoshi Hasegawa and Clésio Luis Tozzi. Shape from shading with perspective projection and camera calibration. Computers and Graphics, 20(3): 351–364, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=3&aid=9600004.

Herman:1988:TDG

[HTKRW88] I. Herman, T. Tolnay-Knefely, J. Reviczky, and F. L. Westhoff. Threedimensional graphic standards and CGI. Computers and Graphics, 12(2):

149–154, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Herman:1984:XMI

[HTKV84] I. Herman, T. Tolnay-Knefely, and A. Vincze. XGKS — a multitask implementation of GKS. Computers and Graphics, 8(2): 141–144, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huang:2019:NHS

[HTW^{+}19] Dongjin Huang, Pengbin Tang, Xianglong Wang, Tao Ruan Wan, and Wen Tang. New haptic syringe device for virtual angiography training. Computers and Graphics, 80(??): 97–103, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300354.

Hunter:1977:CAS

[Hun77] Gregory M. Hunter. Computer animation survey. Computers and Graphics, 2 (4):225–229, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hvistendahl:1986:GEI

[Hvi86a] H. Hvistendahl. Guest Editor’s introduction: Computer graphics in Australia. Computers and Graphics, 10 (3):205, ???? 1986. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849386900014.

Hvistendahl:1986:BMC

[Hvi86b] Harry Hvistendahl. The Brisbane media centre. Computers and Graphics, 10(3): 219–220, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hvistendahl:1986:DGG

[Hvi86c] Harry Hvistendahl. Designing good graphs. Computers and Graphics, 10(3): 221–224, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huijsmans:1987:WMG

[HvK87] D. P. Huijsmans, A. van Delft, and C. A. C. Kuip. WAALSURF: Molecular graphics on a personal computer. Computers and Graphics, 11(4):449–458, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hanusa:1985:GEI

[HW85] H. Hanusa and H. R. Weber. Guest Editors’ introduction. Computers and Graphics, 9(3):183–185, ???? 1985.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900457.

Hunter:1989:BFQ

[HW89] Andrew Hunter and Philip J.

Willis. Breadth-first quad encoding for networked picture browsing. Computers and Graphics, 13(4): 419–432, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hsieh:2016:TSV

[HW16] Cheng-You Hsieh and Yu-Shuen Wang. Traffic situation visualization based on video composition. Computers and Graphics, 54(??): 1–7, February 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001107.

Han:2022:TVG

[HW22] Jun Han and Chaoli Wang.

TSR-VFD: Generating temporal super-resolution for unsteady vector field data. Computers and Graphics, 103(??):168–179, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000115.

Herbert:2022:CLC

[HWEB22] Bradley Herbert, Grant Wigley, Barrett Ens, and Mark Billinghurst. Cognitive load considerations for augmented reality in network security training. Computers and Graphics, 102 (??):566–591, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001874.

He:2009:DCA

[HWFQ09] Ying He, Hongyu Wang, Chi-Wing Fu, and Hong Qin. A divide-and-conquer approach for automatic polycube map construction. Computers and Graphics, 33(3):369–380, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000338.

Hou:2023:MCN

[HWR^{+}23] Ruijie Hou, Zhihao Wang, Ruimin Ren, Yang Cao, and Zhao Wang. Multichannel network: Constructing efficient GCN baselines for skeleton-based action recognition. Computers and Graphics, 110 (??):111–117, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200231X.

Hepperle:2019:SCS

[HWSW19] Daniel Hepperle, Yannick Weiß, Andreas Siess, and Matthias Wölfel. 2D, 3D or speech? A case study on which user interface is preferable for what kind of object interaction in immersive virtual reality. Computers and Graphics, 82 (??):321–331, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300974.

Han:2021:SDA

[HWYL21] Xianjun Han, Huabin Wang, Hongyu Yang, and Xuejun Li. Small data assisting face image illumination normalization. Computers and Graphics, 98 (??):82–92, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000674.

Hu:2012:PMQ

[HXA^{+}12] Shiyu Hu, Qing Xing, Ergun Akleman, Jianer Chen, and Jonathan Gross. Pattern mapping with quad-pattern-coverable quad-meshes. Computers and Graphics, 36 (5):455–465, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000647.

Han:2023:LDD

[HXC^{+}23] Xianjun Han, Zhaoyang Xie, Qianqian Chen, Xuejun Li, and Hongyu Yang. Learning the degradation distribution for medical image superresolution via sparse swin transformer. Computers and Graphics, 114 (??):168–178, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000900.

Hu:2024:DRT

[HXH24] Kaidi Hu, Zongxia Xie, and Qinghua Hu. Dualresolution transformer combined with multi-layer separable convolution fusion network for real-time semantic segmentation. Computers and Graphics, 118 (??):220–232, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323003096.

Hwang:1993:EVS

[HY93a] Sam C. Hwang and Hyun S. Yang. Efficient view sphere tessellation method based on halfedge data structure and quadtree. Computers and Graphics, 17(5):575–581, September–October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Hwang:1993:DAK

[HY93b] Sam Chung Hwang and Hyun Seung Yang. Discrete approximation of the Koch curve. Computers and Graphics, 17(1):95–102, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huang:2003:NCC

[HY03] Ding Huang and Hong Yan.

NURBS curve controlled modelling for facial animation. Computers and Graphics, 27(3):373–385, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Huang:2023:SND

[HY23] You-Liang Huang and Xue-

Feng Yuan. StyleTerrain:

a novel disentangled generative model for controllable high-quality procedural terrain generation. Computers and Graphics, 116(??): 373–382, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300225X.

Hu:2020:PCE

[HYP^{+}20] Kaidong Hu, Sejong Yoon, Vladimir Pavlovic, Petros Faloutsos, and Mubbasir Kapadia. Predicting crowd egress and environment relationships to support building design optimization. Computers and Graphics, 88(??): 83–96, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300376.

Huang:2024:SSW

[HYP^{+}24] Hao Huang, Shuaihang Yuan, Zheng Peng, Yu Hao, Congcong Wen, and Yi Fang. A single 3D shape waveletbased generative model. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000189.

Han:2015:CBO

[HZ15] Pengfei Han and Gang Zhao. CAD-based 3D objects recognition in monocular images for mobile augmented reality. Computers and Graphics, 50(??): 36–46, August 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000680.

Huang:2018:GBM

[HZ18] Weijie Huang and Guoshan Zhang. GPU-based multilayer invariant EKF for camera localization. Computers and Graphics, 76 (??):29–36, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301298.

Hu:2023:LLB

[HZ23] Xuegang Hu and Baoman Zhou. LBARNet: Lightweight bilateral asymmetric residual network for real-time semantic segmentation. Computers and Graphics, 116 (??):1–12, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001681.

Higgins:2022:SDI

[HZC^{+}22] Darragh Higgins, Katja Zibrek, Joao Cabral, Donal Egan, and Rachel McDonnell. Sympathy for the digital: Influence of synthetic voice on affinity, social presence and empathy for photorealistic virtual humans. Computers and Graphics, 104(??):116–128, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000474.

Huang:2019:ACR

[HZD^{+}19] Ruikun Huang, Junli Zhao, Fuqing Duan, Xin Li, Celong Liu, Xiaodan Deng, Zhenkuan Pan, Zhongke Wu, and Mingquan Zhou. Automatic craniofacial registration based on radial curves. Computers and Graphics, 82 (??):264–274, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300883.

Hua:2022:VSD

[HZLC22] Chunjian Hua, Xintong Zou, Yan Ling, and Ying Chen. Visual saliency detection via a recurrent residual convolutional neural network based on densely aggregated features. Computers and Graphics, 104(??): 72–85, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000516.

Hu:2020:ABR

[HZLQ20] Zhiyu Hu, Dongbo Zhang, Shuai Li, and Hong Qin. Attention-based relation and context modeling for point cloud semantic segmentation. Computers and Graphics, 90(??):126–134, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030090X.

Iwata:1983:ISP

[IA83] K. Iwata and E. Arai. Input system for part model in CAD from coloured handwritten illustration with use of ITV. Computers and Graphics, 7(3–4):259–268, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ibaroudene:1991:CRB

[IA91] Djaffer Ibaroudene and Raj Acharya. Coordinate relationships between vertices of linear octree nodes and corners of the universe. Computers and Graphics, 15(3): 375–381, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Icart:1999:TSA

[IA99] Isabelle Icart and Didier Arquès. Technical section

— an approach to geometrical and optical simulation of soap froth. Computers and Graphics, 23(3):405– 418, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/691.pdf.

Isenberg:2006:GSC

[IB06] Tobias Isenberg and Angela Brennecke. G-strokes: a concept for simplifying line stylization. Computers and Graphics, 30(5):754– 766, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001257.

Isenberg:2022:SCS

[IB22] Tobias Isenberg and Stefan Bruckner. Special CG session at EuroVis 2022. Computers and Graphics, 107(??):A5, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001674.

Ishimaru:1987:TIF

[IC87] John M. Ishimaru and Dale E. Calkins. Transportation interface facilities design using interactive high-level computer graphics. Computers and Graphics, 11(3):255–267, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ip:1996:SBF

[IC96] Horace H. S. Ip and C. S. Chan. Script-based facial gesture and speech animation using a NURBS based face model. Computers and Graphics, 20(6):881–891, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600058.

Ivo:2021:MFR

[ICNV21] Douglas Ivo, Joaquim Cavalcante-Neto, and Creto Vidal. A model for flexible representation of social groups in crowd simulation. Computers and Graphics, 101 (??):7–22, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001692.

Iuricich:2017:HFT

[ID17] Federico Iuricich and Leila De Floriani. Hierarchical Forman Triangulation: a multiscale model for scalar field analysis. Computers and Graphics, 66(??):113– 123, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730064X.

Ihlenfeldt:1998:GID

[IE98] Wolf-D. Ihlenfeldt and Klaus Engel. Graphics in/ for digital libraries — visualizing chemical data in the Internet — datadriven and interactive graphics. Computers and Graphics, 22(6):703–714, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/625.pdf.

Iuricich:2015:TCS

[IFD15] Federico Iuricich, Ulderico Fugacci, and Leila De Floriani. Topologically-consistent simplification of discrete Morse complex. Computers and Graphics, 51(??): 157–166, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000540.

Ishikawa:2022:IBV

[II22] Shudai Ishikawa and Takumi Ikenaga. Image-based virtual try-on system with clothing extraction module that adapts to any posture. Computers and Graphics, 106 (??):161–173, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001091.

Inglis:2012:OAR

[IIK12] Tiffany C. Inglis, Stephen Inglis, and Craig S. Kaplan. Op Art rendering with lines and curves. Computers and Graphics, 36 (6):607–621, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000428.

Itoh:2021:FSS

[IK21] Takayuki Itoh and Kiyoshi Kiyokawa. Foreword to the special section on 2019 international conference on cyberworlds (Cyberworlds 2019). Computers and Graphics, 94 (??):A7–A8, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000091.

Ivanov:2000:EIB

[IKB00] D. Ivanov, E. Kuzmin, and S. Burtsev. An efficient integer-based skeletonization algorithm. Computers and Graphics, 24(1):41– 51, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/13/20/47/27/31/article.
                                        pdf.

Iwainsky:1990:CGL

[IKM90] A. Iwainsky, D. Kaiser, and M. May. Computer graphics and layout design in documentation processes. Computers and Graphics, 14(3– 4):377–388, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ikeda:2020:SIO

[IKM^{+}20] Sei Ikeda, Yuto Kimura, Shinnosuke Manabe, Asako Kimura, and Fumihisa Shibata. Shadow induction on optical see-through headmounted displays. Computers and Graphics, 91 (??):141–152, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300996.

Ito:2022:ISB

[IKTS22] Tomohiko Ito, Teruyoshi Kaneko, Yoshiki Tanaka, and Sato Saga. An interactive sketch-based CAD interface realizing geometrical and topological editing across multiple objects based on fuzzy logic. Computers and Graphics, 103 (??):153–167, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000243.

Ihm:1997:IDS

[IL97] Insung Ihm and Rae Kyoung Lee. Indexing data structures for faster volume rendering. Computers and Graphics, 21(4): 497–506, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700025.

Ip:2001:AHM

[ILLC01] Horace H. S. Ip, Maria S. W.

Lam, Ken C. K. Law, and Sam C. S. Chan. Animation of hand motion from target posture images using an anatomy-based hierarchical model. Computers and Graphics, 25 (1):121–133, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/36/article.
pdf.

Iurgel:2007:EPR

[IM07] Ido A. Iurgel and Adérito F.

Marcos. Employing personalityrich virtual persons-new tools required. Computers and Graphics, 31(6): 827–836, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001616.

Ibrahim:2022:DPM

[IMG22] Muhammad Twaha Ibrahim, Aditi Majumder, and M. Gopi. Dynamic projection mapping on deformable stretchable materials using boundary tracking. Computers and Graphics, 103(??): 61–74, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000048.

Itoh:1982:GEA

[IMMS82] K. Itoh, K. Muramatsu, M. Matsui, and S. Suzuki. Graphical editing and analysis system for network system (GEANS). Computers and Graphics, 6(2):47–61, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Isler:1991:FRT

[IO91] Veysi Isler and Bulent Ozguc. Fast ray tracing 3D models. Computers and Graphics, 15(2): 205–216, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). includes survey of efficiency schemes.

Interian:2017:CSF

[IORM17] Ruben Interian, Juan M.

Otero, Celso C. Ribeiro, and Anselmo A. Montenegro. Curve and surface fitting by implicit polynomials: Optimum degree finding and heuristic refinement. Computers and Graphics, 67 (??):14–23, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300444.

Iehl:2003:TPC

[IP03] J. C. Iehl and B. Péroche.

Towards perceptual control of physically based spectral rendering. Computers and Graphics, 27(5):747– 762, October 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Incahuanaco:2023:SRM

[IP23] Filomen Incahuanaco and Afonso Paiva. Surface reconstruction method for particle-based fluids using discrete indicator functions. Computers and Graphics, 114(??):26–35, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000705.

Izmailov:1996:VP

[IPV96] Rauf Izmailov, Alexei Pokrovskii, and Alexander Vladimirov. Visualization of polynomials. Computers and Graphics, 20(1):95–105, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1&aid=9500068.

Ip:2006:OCD

[IR06] Cheuk Yiu Ip and William C. Regli. A 3D object classifier for discriminating manufacturing processes. Computers and Graphics, 30(6): 903–916, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001464.

Isern:2005:SIC

[IS05] Jordi Regincós Isern and Francisco Serón. Special issue: Computer graphics in Spain. Computers and Graphics, 29(2):177– 178, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Isenberg:2021:SCG

[Ise21]  Tobias Isenberg.  Spe-

cial C&G and G&VC session at EuroVis. Computers and Graphics, 96(??): A3–A4, May 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000832.

Islam:2017:RED

[ISPS17] A. B. M. Tariqul Islam, Christian Scheel, Renato Pajarola, and Oliver Staadt. Robust enhancement of depth images from depth sensors. Computers and Graphics, 68(??):53– 65, November 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301322.

Itskovich:2011:SPM

[IT11] Arik Itskovich and Ayellet Tal. Surface partial matching and application to archaeology. Computers and Graphics, 35(2): 334–341, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001792.

Imre:2018:INE

[ITW18] Martin Imre, Jun Tao, and Chaoli Wang. Identifying nearly equally spaced isosurfaces for volumetric data sets. Computers and Graphics, 72(??):82–97, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300220.

Imre:2020:SPS

[ITW^{+}20] Martin Imre, Jun Tao, Yongyu Wang, Zhiqiang Zhao, Zhuo Feng, and Chaoli Wang. Spectrum-preserving sparsification for visualization of big graphs. Computers and Graphics, 87(??): 89–102, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300212.

Inoue:2009:HMS

[IU09] Kohei Inoue and Kiichi Urahama. Halftoning with minimum spanning trees and its application to mazelike images. Computers and Graphics, 33(5):638– 647, October 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000661.

Izmailov:1993:DAS

[IV93] Rauf Izmailov and Alexander Vladimirov. Dimension of aliasing structures. Computers and Graphics, 17(5):539–547, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ivo:2020:ISR

[IVCN20] Rafael Fernandes Ivo, Creto Augusto Vidal, and Joaquim Bento Cavalcante-Neto. Improved silhouette rendering and detection of splatbased models. Computers and Graphics, 93(??):39– 50, December 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301448.

Ip:1994:FCC

[IWM94] Horace H. S. Ip, Helena T. F. Wong, and Florence Y. Mong. Fractal coding of Chinese scalable calligraphic fonts. Computers and Graphics, 18(3): 343–351, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Israel:2009:ITD

[IWM^{+}09] J. H. Israel, E. Wiese, M. Mateescu, C. Zöllner, and R. Stark. Investigating three-dimensional sketching for early conceptual designresults from expert discussions and user studies. Computers and Graphics, 33 (4):462–473, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000855.

Ihmsen:2013:LFS

[IWT13] Markus Ihmsen, Arthur Wahl, and Matthias Teschner. A Lagrangian framework for simulating granular material with high detail. Computers and Graphics, 37(7): 800–808, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000629.

Ibrahim:2018:FFS

[IY18] Mohamed Ibrahim and Dong-Ming Yan. Fold and fit: Space conserving shape editing. Computers and Graphics, 70 (??):316–326, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301310.

Iwata:1997:LIS

[IYH97] Hiroo Iwata, Hiroaki Yano, and Wataru Hashimoto. LHX: An integrated software tool for haptic interface. Computers and Graphics, 21(4):413–420, July–August 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-

bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700018.

Jenness:1984:GSP

[JA84a] J. D. Jenness and G. C. Andrews. Graphic simulation of planar dynamic mechanisms. Computers and Graphics, 8 (4):411–421, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Joshi:1984:GGS

[JA84b] R. R. Joshi and H. Arunachalam. GRASP — a 3D graphics system for PAS-CAL users. Computers and Graphics, 8(1):93–102, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Joshi:1984:MPS

[JA84c] R. R. Joshi and H. Arunachalam. Multilevel picture segmentation in the graphics system — GRASP. Computers and Graphics, 8(2): 167–176, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jacobson:1993:AVR

[Jac93]  Robert Jacobson.  After

the “virtual reality” gold rush: The virtual worlds paradigm. Computers and Graphics, 17(6):695–698, November–December 1993.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jacob:1995:TDI

[Jac95] Marie-Andrée Jacob. Transformation of digital images by discrete affine applications. Computers and Graphics, 19(3):373– 389, May–June 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500008.

Joan-Arinyo:2003:GAR

[JALS03] R. Joan-Arinyo, M. V. Luzón, and A. Soto. Genetic algorithms for root multiselection in constructive geometric constraint solving. Computers and Graphics, 27 (1):51–60, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jarvis:1975:GDS

[Jar75] J. F. Jarvis. A graphical display system utilizing plasma panels. Computers and Graphics, 1(2– 3):175–180, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jassim:1988:NAG

[Jas88] S. A. Jassim. A noniterative algorithm for generating connected quadrilaterals of specified area. Computers and Graphics, 12(2): 229–233, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Joan-Arinyo:1997:CRB

[JAS97] R. Joan-Arinyo and A. Soto.

A correct rule-based geometric constraint solver. Computers and Graphics, 21(5):599–609, September– October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700038.

Jensen:1995:PMB

[JC95] Henrik Wann Jensen and Niels Jørgen Christensen. Photon maps in bidirectional Monte Carlo ray tracing of complex objects. Computers and Graphics, 19(2): 215–224, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400145.

Junior:2018:MMB

[JCFN18] Sergio N. Silva Junior, Felipe C. Chamone, Renato C. Ferreira, and Erickson R. Nascimento. A 3D modeling methodology based on a concavity-aware geometric test to create 3D textured coarse models from concept art and orthographic projections. Computers and Graphics, 76(??):73– 83, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301365.

Jain:2015:MLS

[JCT^{+}15] Arjun Jain, Chao Chen, Thorsten Thormählen, Dimitris Metaxas, and Hans-Peter Seidel. Multi-layer stencil creation from images. Computers and Graphics, 48(??):11–22, May 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000114.

Jervert:1975:MCI

[JD75] J. C. Jervert and R. M. Dunn. Multi-console intelligent satellite graphics. Computers and Graphics, 1(2– 3):181–190, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

James:1999:TSH

[JD99] A. James and A. M.

Day. Technical section — the hidden face determination tree. Computers and Graphics, 23(3):377– 387, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/689.pdf.

Joshi:1988:HHT

[JDGS88] R. C. Joshi, H. Darbari, Goel, and S. Sasikumaran. A hierarchical hex-tree representational technique for solid modelling. Computers and Graphics, 12(2): 235–238, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jiang:2023:CCS

[JDT^{+}23] Zhaoyi Jiang, Luyun Ding, Gary K. L. Tam, Chao Song, Frederick W. B. Li, and Bailin Yang. C2SPoint: a classification-to-saliency network for point cloud saliency detection. Computers and Graphics, 115 (??):274–284, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001322.

Jeffrey:1992:CGV

[Jef92]  H. Joel Jeffrey.  Chaos

game visualization of sequences. Computers and Graphics, 16(1):25–33, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jern:1985:RGA

[Jer85] M. Jern. The raster graphics approach in mapping. Computers and Graphics, 9 (4):373–381, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jern:1995:CWI

[Jer95] Mikael Jern. Custom widgets for interactive visualization using X and Motif. Computers and Graphics, 19(2): 189–197, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400142.

Jorge:2000:GEI

[JG00] Joaquim Jorge and Ephraim P. Glinert. Guest Editors’ introduction. Computers and Graphics, 24(6):817– 818, December 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/26/article.
pdf.

Jamin:2009:CVC

[JGA09] Clément Jamin, Pierre-Marie Gandoin, and Samir Akkouche. CHuMI viewer: Compressive huge mesh interactive viewer. Computers and Graphics, 33 (4):542–553, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000582.

Jense:1989:IVB

[JH89] G. J. Jense and D. P.

Huijsmans. Interactive voxel-based graphics for 3D reconstruction of biological structures. Computers and Graphics, 13(2): 145–150, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jenny:2011:SCH

[JH11] Bernhard Jenny and Lorenz Hurni. Studying cartographic heritage: Analysis and visualization of geometric distortions. Computers and Graphics, 35(2): 402–411, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000069.

Jung:2012:EMA

[JHL^{+}12] Jinki Jung, Jaewon Ha, Sang-Wook Lee, Francisco A. Rojas, and Hyun S. Yang. Efficient mobile AR technology using scalable recognition and tracking based on server-client model. Computers and Graphics, 36(3): 131–139, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000052.

Jo:2011:AFI

[JHPhR11] Hyungeun Jo, Sungjae Hwang, Hyunwoo Park, and Jung hee Ryu. Aroundplot: Focus+ context interface for off-screen objects in 3D environments. Computers and Graphics, 35 (4):841–853, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001087.

Jordan:1997:AHP

[JJL97] Stephen D. Jordan, Philip E. Jensen, and Barthold B. A. Lichtenbelt. An architecture for high-performance 2-D image display. Computers and Graphics, 21(2): 151–157, March–April 1997.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600078.

Jurado:2022:EMA

[JJPP^{+}22] Juan M. Jurado, J. Roberto Jiménez-Pérez, Luís Pádua, Francisco R. Feito, and Joaquim J. Sousa. An efficient method for acquisition of spectral BRDFs in real-world scenarios. Computers and Graphics, 102 (??):154–163, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001850.

Jansen:1984:IFD

[JK84] H. Jansen and F. L. Krause.

Interpretation of freehand drawings for mechanical design processes. Computers and Graphics, 8(4):351–369, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jetschny:1990:GPA

[JK90] W. Jetschny and W. Krug.

Graphic presentation and animation in the MSOKS kernel system for modelling, simulation, and optimization in the computer-aided production. Computers and Graphics, 14(3–4):419–427, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ji:2015:IGD

[JK15] Joonghyun Ji and Kwang-Hee Ko. Improved gloss depiction using the empirical highlight un-distortion method for 3D imagewarping-based stereo rendering. Computers and Graphics, 50(??):1–12, August 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849315000485.

Jiang:2021:DRC

[JK21] Giulio Jiang and Bernhard Kainz. Deep radiance caching: Convolutional autoencoders deeper in ray tracing. Computers and Graphics, 94(??): 22–31, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301412.

Jeong:2002:TSD

[JKK02] Moonja Jeong, Gi Ok Kim, and Seong-A Kim. Technical section: Dynamics of Newton’s method for solving some equations. Computers and Graphics, 26(2):271– 279, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/34/abstract.
html.

Jin:2023:RHM

[JL23] Pengle Jin and Xinguo Liu.

Robust human motion estimation using bidirectional motion prior model and spatiotemporal progressive motion optimization. Computers and Graphics, 115 (??):331–338, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001450.

Jin:2000:GCD

[JLP00] Xiaogang Jin, Youfu Li, and Qunsheng Peng. General constrained deformations based on generalized metaballs. Computers and Graphics, 24(2):219–231, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/30/article.
pdf.

Jackson:1988:EDC

[JM88] Alden W. Jackson and Joel M. Morris. Enhancement of diglib: Computer graphics software for animated computer-generated video movies. Computers and Graphics, 12(2): 271–283, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Johnson:2004:LSI

[JMC^{+}04] Andrew Johnson, Thomas Moher, Yong-Joo Cho, Danny Edelson, and Eric Russell. Learning science inquiry skills in a virtual field. Computers and Graphics, 28(3):409–416, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Joensson:1990:CGD

[JMV90] D. Joensson, J. Müglitz, and M. Vogel. Computer graphics for design processes in mechanical engineering. Computers and Graphics, 14 (3–4):413–417, 1990. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jansen:1985:HHF

[JNR85] H. Jansen, E. Nullimeir, and K. H. Roediger. Handsketching as a human factor aspect in graphical interaction. Computers and Graphics, 9(3):195–210, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Joel:1992:SA

[Joe92] William J. Joel. A simple attraction. Computers and Graphics, 16(1):41–43, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

John:2018:FSS

[Joh18] Nigel W. John. Foreword to the special section on the 2017 International Conference on Cyberworlds. Computers and Graphics, 72(??): A2–A3, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300293.

Jo:2007:SVI

[JOK^{+}07] Yoshitomo Jo, Masafumi Oka, Akinori Kimura, Kyoko Hasegawa, Ayumu Saitoh, Susumu Nakata, Akihiro Shibata, and Satoshi Tanaka. Stochastic visualization of intersection curves of implicit surfaces. Computers and Graphics, 31(2): 230–242, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000040.

Jones:1990:TUR

[Jon90] John Dewey Jones. Three unconventional representations of the Mandelbrot set. Computers and Graphics, 14(1):127–129, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jones:1991:MS

[Jon91] John Dewey Jones. The method of secants. Computers and Graphics, 15(3): 451–454, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jorge:2007:E

[Jor07]  Joaquim Jorge.  Edito-

rial. Computers and Graphics, 31(4):537, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001495.

Jorge:2010:Ea

[Jor10a]  Joaquim Jorge.  Edito-

rial. Computers and Graphics, 34(2):93, April 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000257.

Jorge:2010:Eb

[Jor10b] Joaquim Jorge. Editorial. Computers and Graphics, 34(6):640, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001639.

Jorge:2011:E

[Jor11] Joaquim Jorge. Editorial.

Computers and Graphics, 35 (6):1070, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001622.

Jorge:2018:ENa

[Jor18a]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 71(??):A1– A2, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300347.

Jorge:2018:ENb

[Jor18b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 74(??):A1– A2, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300992.

Jorge:2018:LEC

[Jor18c] Joaquim Jorge. A letter from the Editor-in-Chief. Computers and Graphics, 72 (??):A1, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931830044X.

Jorge:2018:NECa

[Jor18d] Joaquim Jorge. A note from the Editor-in-Chief. Computers and Graphics, 70 (??):A1–A2, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730208X.

Jorge:2018:NECb

[Jor18e] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 75 (??):ii, October 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931830133X.

Jorge:2018:NECc

[Jor18f] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 76 (??):A1–A2, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301511.

Jorge:2018:NECd

[Jor18g] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 77 (??):A1–A2, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301912.

Jorge:2019:ENa

[Jor19a]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 83(??):A1– A2, October 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301335.

Jorge:2019:ENb

[Jor19b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 84(??):A1– A2, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930161X.

Jorge:2019:ENc

[Jor19c]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 85(??):A1–A2, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930175X.

Jorge:2019:NECa

[Jor19d] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 78 (??):A1–A2, February 2019.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300111.

Jorge:2019:NECb

[Jor19e] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 79(??):A1–A2, April 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931930024X.

Jorge:2019:NECc

[Jor19f] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 80(??):A1–A2, May 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300548.

Jorge:2019:NECd

[Jor19g] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 81(??):A1–A2, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300871.

Jorge:2019:NECe

[Jor19h] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 82 (??):A1–A3, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301086.

Jorge:2020:ENa

[Jor20a]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 86(??):A1– A2, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300108.

Jorge:2020:ENb

[Jor20b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 87(??):A1– A2, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300297.

Jorge:2020:ENc

[Jor20c]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 89(??):A1– A3, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300807.

Jorge:2020:ENd

[Jor20d]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 90(??):A1– A3, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932030100X.

Jorge:2020:NECa

[Jor20e] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 88(??):A1–A3, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300479.

Jorge:2020:NECb

[Jor20f] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 91 (??):A1–A2, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301291.

Jorge:2020:NECc

[Jor20g] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 92 (??):A1–A3, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301576.

Jorge:2020:NECd

[Jor20h] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 93 (??):A1–A2, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301680.

Jorge:2021:E

[Jor21a] Joaquim Jorge. Editorial.

Computers and Graphics, 101(??):1–2, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002508.

Jorge:2021:ENa

[Jor21b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 98(??):A1– A3, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001461.

Jorge:2021:ENb

[Jor21c]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 99(??):A1– A4, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001941.

Jorge:2021:ENc

[Jor21d]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 100(??):A1–A4, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002302.

Jorge:2021:NECa

[Jor21e] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 94 (??):A1–A2, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000157.

Jorge:2021:NECb

[Jor21f] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 95(??):A1–A2, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000327.

Jorge:2021:NECc

[Jor21g] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 96(??):A1–A2, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000820.

Jorge:2021:NECd

[Jor21h] Joaquim Jorge. A note from the Editor in Chief. Computers and Graphics, 97(??):A1–A2, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000960.

Jorge:2022:ENa

[Jor22a]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 102(??):A1–A2, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200019X.

Jorge:2022:ENb

[Jor22b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 103(??):A1– A2, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000383.

Jorge:2022:ENc

[Jor22c]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 105(??):A1– A3, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000942.

Jorge:2022:ENd

[Jor22d]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 106(??):A1–A3, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001297.

Jorge:2022:ENe

[Jor22e]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 107(??):A1–A2, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001686.

Jorge:2022:ENf

[Jor22f]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 108(??):A1–A3, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200190X.

Jorge:2022:ENg

[Jor22g]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 109(??):A1–A2, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322002084.

Jorge:2023:ENa

[Jor23a]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 110(??):A1–A3, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000158.

Jorge:2023:ENb

[Jor23b]  Joaquim Jorge.  Edito-

rial note. Computers and Graphics, 111(??):A1– A3, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000493.

Jorge:2023:INEa

[Jor23c]  Joaquim Jorge.  Issue

112: a note from the Editor-in-Chief. Computers and Graphics, 112(??): A1–A4, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000833.

Jorge:2023:INEb

[Jor23d]  Joaquim Jorge.  Issue

113C: a note from the Editor in Chief. Computers and Graphics, 113(??): A1–A2, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000973.

Jorge:2023:NCGb

[Jor23e] Joaquim Jorge. Note Computers and Graphics issue

117. Computers and Graphics, 117(??):A1–A2, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002923.

Jorge:2023:NCGa

[Jor23f] Joaquim Jorge. Note computers and graphics issue 115. Computers and Graphics, 115(??):A1–A3, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300256X.

Juan:2010:UAV

[JP10] M. Carmen Juan and David Pérez. Using augmented and virtual reality for the development of acrophobic scenarios. Comparison of the levels of presence and anxiety. Computers and Graphics, 34(6):756– 766, December 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001263.

Jesus:2018:GSD

[JPCS18] Diego Jesus, Gustavo Patow, António Coelho, and António Augusto Sousa. Generalized selections for direct control in procedural buildings. Computers and Graphics, 72(??): 106–121, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300232.

Jones:2001:TDF

[JPP01] Gareth T. Jones, David J.

Parish, and Iain W. Phillips. A transform domain feature detection and concealment algorithm for errors in DCT encoded images. Computers and Graphics, 25(4):671– 680, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/40/abstract.
html.

Jones:2000:CAC

[JR00] Kevin C. Jones and Clifford A. Reiter. Chaotic attractors with cyclic symmetry revisited. Computers and Graphics, 24(2): 271–282, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/

                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/35/article.
                                        pdf.

Jurado-Rodriguez:2022:SSC

[JRJP^{+}22] David Jurado-Rodríguez, Juan M. Jurado, Luís Pádua, Alexandre Neto, Rafael Muñoz-Salinas, and Joaquim J. Sousa. Semantic segmentation of 3D car parts using UAV-based images. Computers and Graphics, 107(??):93–103, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200125X.

Joshi:2021:TFT

[JRS21] Piyush Joshi, Alireza Rastegarpanah, and Rustam Stolkin. A training free technique for 3D object recognition using the concept of vibration, energy and frequency. Computers and Graphics, 95(??): 92–105, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000145.

Joshi:2017:VLC

[JRSM17] Sarang Joshi, Yoshida Rao, Bharath Ram Sundar, and Ramanathan Muthuganapathy. On the visibility locations for continuous curves. Computers and Graphics, 66(??):34–44, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300729.

Jovane:2023:WCA

[JRZ^{+}23] Alberto Jovane, Pierre Raimbaud, Katja Zibrek, Claudio Pacchierotti, Marc Christie, Ludovic Hoyet, Anne-Hélène Olivier, and Julien Pettré. Warping character animations using visual motion features. Computers and Graphics, 110 (??):38–48, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002023.

Johnstone:1992:CIP

[JS92] John K. Johnstone and Ching-Kuang Shene. Computing the intersection of a plane and a natural quadric. Computers and Graphics, 16 (2):179–186, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jimenez:2008:CDB

[JS08] Juan J. Jiménez and Rafael J.

Segura. Collision detection between complex polyhedra. Computers and Graphics, 32 (4):402–411, August 2008.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000526.

Janicke:2009:SVD

[JS09] H. Jänicke and G. Scheuermann. Steady visualization of the dynamics in fluids using ϵ-machines. Computers and Graphics, 33 (5):597–606, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000892.

Ju:2004:RTP

[JSG04] Tao Ju, Scott Schaefer, and Ron Goldman. Recursive turtle programs and iterated affine transformations. Computers and Graphics, 28(6): 991–1004, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Janetzko:2014:ADV

[JSMK14] Halldór Janetzko, Florian Stoffel, Sebastian Mittelstädt, and Daniel A. Keim. Anomaly detection for visual analytics of power consumption data. Computers and Graphics, 38 (??):27–37, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001477.

Jin:2003:SII

[JSP03] Xiaogang Jin, Hanqiu Sun, and Qunsheng Peng. Subdivision interpolating implicit surfaces. Computers and Graphics, 27(5):763– 772, October 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jaillet:1998:DOR

[JSV98] F. Jaillet, B. Shariat, and D. Vandorpe. Deformable object reconstruction with particle systems. Computers and Graphics, 22(2–3):189– 194, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/540.pdf.

Joormann:1986:ACC

[JT86] Otto Joormann and Geert Teunis. The application of CAD/CAM systems at Volkswagen. Computers and Graphics, 10(4):317–325, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jin:2002:TSA

[JT02] Xiaogang Jin and Chiew-Lan Tai. Technical section: Analytical methods for polynomial weighted convolution surfaces with various kernels. Computers and Graphics, 26(3):437– 447, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/33/abstract.
html.

Jimenez:2001:CDS

[JTT01] P. Jiménez, F. Thomas, and C. Torras. 3D collision detection: a survey. Computers and Graphics, 25(2):269–285, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/34/article.
pdf.

Jung:1994:HGI

[Jun94]  Volker Jung.  HIST: a

geographic information system for the support of historic research. Computers and Graphics, 18(2): 219–225, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Juttler:1994:VMO

[Jut94] Bert Juttler. Visualization of moving objects using dual quaternion curves. Computers and Graphics, 18(3): 315–326, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jansen:1984:PTR

[JV84] F. W. Jansen and J. J. Van Wijk. Previewing techniques in raster graphics. Computers and Graphics, 8(2): 149–161, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jansen:1991:SAC

[JV91] Frederik W. Jansen and Arno N. T. Van Der Zalm. A shadow algorithm for CSG. Computers and Graphics, 15(2):237–247, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jansen:1995:CGE

[Jv95] Frederik W. Jansen and Peter R. van Nieuwenhuizen. Computer graphics education at Delft University of Technology. Computers and Graphics, 19(3): 461–465, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500017.

Jovanovic:2024:CTP

[JVS^{+}24] Marko Jovanović, Marko Vucić, Vesna Stojaković, Bojan Tepavcević, and Mirko Raković. Creating terracotta panels through grayscale image processing and robotic hotwire cut molds. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000402.

Jin:2012:APC

[JWL12] Yong Jin, Qingbiao Wu, and Ligang Liu. Aesthetic photo composition by optimal crop-and-warp. Computers and Graphics, 36(8): 955–965, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001380.

Jin:2023:MGR

[JWZ23] Chuan Jin, Tieru Wu, and Junsheng Zhou. Multigrid representation with field regularization for selfsupervised surface reconstruction from point clouds. Computers and Graphics, 114(??):379–386, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323001097.

Joy:1996:RFC

[JX96] Gregory Joy and Zhigang Xiang. Reducing false contours in quantized color images. Computers and Graphics, 20(2):231–242, March–April 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500098.

Jia:2022:RTF

[JXJ22] Huijie Jia, Zhongjun Xiao, and Peng Ji. Real-time fatigue driving detection system based on multimodule fusion. Computers and Graphics, 108(??):22– 33, November 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001625.

Jiang:2010:CCS

[JXM^{+}10] Hao Jiang, Wenbin Xu, Tianlu Mao, Chunpeng Li, Shihong Xia, and Zhaoqi Wang. Continuum crowd simulation in complex environments. Computers and Graphics, 34(5):537– 544, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000804.

Jiang:2022:PCC

[JXW^{+}22] Xiaotong Jiang, Benlian Xu, Mingqiang Wei, Ke Wu, Siyuan Yang, Longgen Qian, Ningzhong Liu, and Qingjin

                                        Peng.  C2F-3DToothSeg:

Coarse-to-fine 3D tooth segmentation via intuitive single clicks. Computers and Graphics, 102(??):601– 609, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001825.

Jie:1987:CCG

[JXY87] Pang Yun Jie, Yang Shu Xun, and Chi Yiu. Combining computer graphics with Chinese traditional painting. Computers and Graphics, 11(1):63–68, 1987. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jung:1998:WCB

[JY98] Jin H. Jung and Hyun S. Yang. Window capturingbased application sharing under heterogeneous window systems. Computers and Graphics, 22(2–3):243–254, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/547.pdf.

Jiang:2023:AEM

[JYC^{+}23] Pengfei Jiang, Xiaoyan Yang, Yuanjie Chen, Wenjie Song, and Yang Li.

AdaptMVSNet:  Efficient

multi-view stereo with adaptive convolution and attention fusion. Computers and Graphics, 116(??): 128–138, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001838.

Jiang:2017:HMZ

[JYL17] Xiaolei Jiang, Hongxun Yao, and Shaohui Liu. How many zero crossings? a method for structure-texture image decomposition. Computers and Graphics, 68(??): 129–141, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301279.

Jia:2023:WGI

[JYL23] Hongbin Jia, Qingbo Yin, and Mingyu Lu. Weighted guided image filtering with entropy evaluation weighting. Computers and Graphics, 117(??):114–123, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002595.

Jia:2024:SKW

[JYL24] Hongbin Jia, Qingbo Yin, and Mingyu Lu. Steering kernel weighted guided image filtering with gradient constraint. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000438.

Jin:2014:GID

[JYLW14] Jianqiu Jin, Bailing Yang, Kewei Liang, and Xun Wang. General image denoising framework based on compressive sensing theory. Computers and Graphics, 38 (??):382–391, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001908.

Jaillet:2017:FSS

[JZ17] Fabrice Jaillet and Florence Zara. Foreword to the special section on VRIPHYS 2017. Computers and Graphics, 69 (??):A1–A2, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301759.

Ji:2023:FFC

[JZLP23] Baoning Ji, Jie Zhang, Yuan Li, and Jiazhen Pang. Freeform CAD model retrieval approach for engineering reuse based on local feature segmentation. Computers and Graphics, 111 (??):111–121, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000080.

Jia:1988:CAD

[JZR88] Minghua Jia, Xianzhang Zeng, and Xueyu Ruan. Computer-aided design of extrusion dies. Computers and Graphics, 12(3–4): 335–340, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Jia:2021:CRB

[JZX^{+}21] Bin Jia, Junli Zhao, Shiqing Xin, Fuqing Duan, Zhenkuan Pan, Zhongke Wu, Jinhua Li, and Mingquan Zhou. Craniofacial reconstruction based on heat flow geodesic grid regression (HF-GGR) model. Computers and Graphics, 97(??): 258–267, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000716.

Jiang:2023:IBM

[JZY^{+}23] Nanhe Jiang, Yucun Zhang, Fang Yan, Xianbin Fu, and Tao Kong. Image blind motion deblurring method with longitudinal channel and wavelet dynamic convolution. Computers and Graphics, 116(??):275–286, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001917.

Jaillet:2016:FSS

[JZZ16] Fabrice Jaillet, Florence Zara, and Gabriel Zachmann. Foreword to special section on VRIPHYS 2015. Computers and Graphics, 57(??):A3–A4, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300334.

Kehl:2000:HMB

[K^{+}00a] Hans Gerd Kehl et al. 3D heart modelling from biplane, rotational angiocardiographic X-ray sequences. Computers and Graphics, 24 (5):731–739, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/33/article.
pdf.

Klimenko:2000:VST

[K^{+}00b] Stanislav Klimenko et al. Visualization in string theory. Computers and Graphics, 24 (1):23–30, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/29/article.
pdf.

Kaufman:1985:TST

[KA85] A. Kaufman and S. Azaria.

Texture synthesis techniques for computer graphics. Computers and Graphics, 9(2): 139–145, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kuhlmann:1986:CGS

[KA86] H. W. Kuhlmann and B. W.

Alheit. Communicating via graphics standard interfaces. Computers and Graphics, 10 (2):97–102, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kamal:2022:DTA

[KA22] Ahmed Kamal and Carlos Andujar. Designing, testing and adapting navigation techniques for the immersive web. Computers and Graphics, 106(??):66– 76, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000875.

Kunii:1975:IFD

[KAAO75] T. L. Kunii, T. Amano, H. Arisawa, and S. Okada. An interactive fashio design system ‘INFADS’. Computers and Graphics, 1(4): 297–302, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kaya:2018:LFP

[KAFB18] Erdem Kaya, Sema Alacam, Yasin Findik, and Selim Balcisoy. Low-fidelity prototyping with simple collaborative tabletop computeraided design systems. Computers and Graphics, 70 (??):307–315, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730119X.

Kalantari:2004:PAA

[Kal04] Bahman Kalantari. Polynomiography and applications in art, education, and science. Computers and Graphics, 28(3):417– 430, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kamat:1993:STS

[Kam93] V. V. Kamat. A survey of techniques for simulation of dynamic collision detection and response. Computers and Graphics, 17(4): 379–385, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kansy:1985:EG

[Kan85]  K. Kansy.  3D exten-

sion to GKS. Computers and Graphics, 9(3):267–273, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kurc:1998:OSP

[KAÖ98] Tahsin M. Kurç, Cevdet Aykanat, and Bülent Özgüç. Object-space parallel polygon rendering on hypercubes. Computers and Graphics, 22(4):487–503, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
4/581.pdf.

Karakas:1992:AVC

[Kar92] Chris Karakas. Applications of visibility on the computation of offset curve patterns. Computers and Graphics, 16 (2):159–165, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kasprzak:1987:LAO

[Kas87] Wlodzimierz Kasprzak. A linguistic approach to 3D object recognition. Computers and Graphics, 11(4): 427–443, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kaufman:1988:EAS

[Kau88] Arie Kaufman. Efficient algorithms for scan-converting 3D polygons. Computers and Graphics, 12(2): 213–219, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Khachaturov:1988:CAO

[KAV^{+}88] V. R. Khachaturov, N. D. Astakhov, V. E. Veselovskij, A. V. Zlotov, I. A. Krylov, and I. H. Sigal. Computeraided oilfield assimilation design system: Purpose and major capabilities. Computers and Graphics, 12(3– 4):467–475, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kolling:2023:MBF

[KAVM23] Camila Kolling, Victor Araujo, Adriano Veloso, and Soraia Raupp Musse. Mitigating bias in facial analysis systems by incorporating label diversity. Computers and Graphics, 116(??): 173–184, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001899.

Kolingerova:2002:TSI

[KB02] Ivana Kolingerová and BorutŽalik.

Technical section: Improvements to randomized incremental Delaunay insertion. Computers and Graphics, 26(3):477–490, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/36/abstract.
html.

Kobbelt:2004:SPB

[KB04] Leif Kobbelt and Mario Botsch. A survey of pointbased techniques in computer graphics. Computers and Graphics, 28(6): 801–814, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:2005:FEP

[KB05] Ku-Jin Kim and Nakhoon Baek. Fast extraction of polyhedral model silhouettes from moving viewpoint on curved trajectory. Computers and Graphics, 29(3): 393–402, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kerautret:2006:RSM

[KB06] Bertrand Kerautret and Achille Braquelaire. A reversible and statistical method for discrete surfaces smoothing. Computers and Graphics, 30 (1):54–61, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002098.

Kimmel:2010:SAS

[KB10] Bradley W. Kimmel and Gladimir V. G. Baranoski. Simulating the appearance of sandy landscapes. Computers and Graphics, 34 (4):441–448, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000567.

Kurz:2012:HAR

[KB12] Daniel Kurz and Selim Benhimane. Handheld Augmented Reality involving gravity measurements. Computers and Graphics, 36(7): 866–883, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000866.

Kuric:2015:AEI

[KB15] Eduard Kuric and Maria Bielikova. ANNOR: Efficient image annotation based on combining local and global features. Computers and Graphics, 47(??): 1–15, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400123X.

Kristiansen:2020:VIV

[KB20] Yngve Sekse Kristiansen and Stefan Bruckner. Visception: an interactive visual framework for nested visualization design. Computers and Graphics, 92 (??):13–27, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301254.

Kappe:2022:TBF

[KBL22] Christopher Kappe, Michael Böttinger, and Heike Leitte. Topology-based feature analysis of scalar field ensembles: an application to climate (change) analysis. Computers and Graphics, 104(??):59–71, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200036X.

Kaplan:2007:RFG

[KC07] Matthew Kaplan and Elaine Cohen. Reconstructing the frontal geometry of drawings of arbitrary surfaces. Computers and Graphics, 31 (4):568–579, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000994.

Kremer:2022:AEP

[KCH^{+}22] Melissa Kremer, Peter Caruana, Brandon Haworth, Mubbasir Kapadia, and Petros Faloutsos. Automatic estimation of parametric saliency maps (PSMs) for autonomous pedestrians. Computers and Graphics, 104(??):86–94, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000486.

Kim:2017:CTU

[KCK17] Kyungyoon Kim, John V. Carlis, and Daniel F. Keefe. Comparison techniques utilized in spatial 3D and 4D data visualizations: a survey and future directions. Computers and Graphics, 67 (??):138–147, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300481.

Kim:2018:EMS

[KCL18] Yong Hwi Kim, Junho Choi, and Kwan H. Lee. An efficient method for specularenhanced BTF compression. Computers and Graphics, 75(??):A4, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300918.

Kuhnapfel:2000:EST

[KÇM00] U. Kühnapfel, H. K. Çakmak, and H. Maaß. Endoscopic surgery training using virtual reality and deformable tissue simulation. Computers and Graphics, 24 (5):671–682, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/28/article.
pdf.

Klopfer:2002:NPS

[KCR02] Eric Klopfer, Vanessa Colella, and Mitchel Resnick. New paths on a StarLogo adventure. Computers and Graphics, 26(4):615–622, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        13/20/68/55/42/abstract.
                                        html.

Kushwaha:2022:EHP

[KCS22] Mohit Kushwaha, Jaytrilok Choudhary, and Dhirendra Pratap Singh. Enhancement of human 3D pose estimation using a novel concept of depth prediction with pose alignment from a single 2D image. Computers and Graphics, 107 (??):172–185, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001455.

Kushwaha:2024:PEN

[KCS24] Mohit Kushwaha, Jaytrilok Choudhary, and Dhirendra Pratap Singh. 3DPMesh: an enhanced and novel approach for the reconstruction of 3D human meshes from a single 2D image. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000219.

Khan:2022:RAV

[KCU^{+}22] Dawar Khan, Zhanglin Cheng, Hideaki Uchiyama, Sikandar Ali, Muhammad Asshad, and Kiyoshi Kiyokawa. Recent advances in visionbased indoor navigation: a systematic literature review. Computers and Graphics, 104(??):24–45, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000371.

Konstantinovic:1986:GSP

[KD86] Zoran Konstantinovic and Ljiljana B. Damnjanovic. A graphics software package for interactive graphics terminal. Computers and Graphics, 10(3):257–261, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kanev:1994:FGO

[KD94a] Kamen Kanev and Kris Dockx. A framework for graphically-oriented human computer interactions in intelligent operator support systems. Computers and Graphics, 18(4): 563–570, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kromker:1994:GEI

[KD94b] Detlef Kromker and David Duce. Guest editors’ introduction. Computers and Graphics, 18(1):3– ??, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kopylov:2000:LPV

[KD00] Edward A. Kopylov and Kirill A. Dmitriev. Light propagation visualization as a tool for 3D scene analysis in lighting design. Computers and Graphics, 24 (1):31–39, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/27/30/article.
                                        pdf.

Kirste:2003:IST

[KD03] T. Kirste and D. Duke. Introduction: Situation- and task-awareness computing. Computers and Graphics, 27 (6):835–836, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kels:2011:ROC

[KD11] Shay Kels and Nira Dyn. Reconstruction of 3D objects from 2D cross-sections with the 4-point subdivision scheme adapted to sets. Computers and Graphics, 35(3):741–746, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931100046X.

Kurtek:2015:CSF

[KD15] Sebastian Kurtek and Hassen Drira. A comprehensive statistical framework for elastic shape analysis of 3D faces. Computers and Graphics, 51(??):52– 59, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000746.

Krishnan:1996:OEB

[KDG96] R. Krishnan, A. Das, and B. Gurumoorthy. Octree encoding of B-Rep based objects. Computers and Graphics, 20(1):107–114, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500069.

Kautz:2004:AEM

[KDS04] Jan Kautz, Katja Daubert, and Hans-Peter Seidel. Advanced environment mapping in VR applications. Computers and Graphics, 28 (1):99–104, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Keith:1989:ACD

[Kei89]  Michael Keith.  Artistic

choice in the display of chaotic dynamics. Computers and Graphics, 13(3): 393–395, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kelly:1986:FTC

[Kel86] Mary B. Kelly. First-time computer graphics. Computers and Graphics, 10(2): 157–159, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kelley:2000:LTF

[Kel00] Alice Kelley. Layering techniques in fractal art. Computers and Graphics, 24 (4):611–616, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/37/article.
pdf.

Khosravi:2022:MSV

[KES22] Hooman Khosravi, Katayoon Etemad, and Faramarz F. Samavati. Mass simulation in VR using vibrotactile feedback and a co-located physically-based virtual hand. Computers and Graphics, 102(??):120– 132, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002582.

Kuipers:2018:HPL

[KEVD18] Tim Kuipers, Willemijn Elkhuizen, Jouke Verlinden, and Eugeni Doubrovski. Hatching for 3D prints: Line-based halftoning for dual extrusion fused deposition modeling. Computers and Graphics, 74(??): 23–32, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300578.

Klinger:1988:PDA

[KF88] Allen Klinger and Warren K.

Fox. Point data analysis. Computers and Graphics, 12(3–4):557–564, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:2002:ESI

[KF02] M. H. Kim and A. Fischer. Editorial: Special issue: Geometric modeling and computer graphics. Computers and Graphics, 26(5):653–654, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

13/20/68/56/26/abstract.
html.

Klein:2009:VCM

[KFH^{+}09] Jan Klein, Ola Friman, Markus Hadwiger, Bernhard Preim, Felix Ritter, Anna Vilanova, Gabriel Zachmann, and Dirk Bartz. Visual computing for medical diagnosis and treatment. Computers and Graphics, 33 (4):554–565, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000685.

Khojasteh:2023:GGM

[KFN23] Mohammad Hossein Khojasteh, Nastaran Moradzadeh Farid, and Ahmad Nickabadi. GMFIM: a generative mask-guided facial image manipulation model for privacy preservation. Computers and Graphics, 112 (??):81–91, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000365.

Kanzler:2016:LDC

[KFW16] Mathias Kanzler, Florian Ferstl, and Rüdiger Westermann. Line density control in screen-space via balanced line hierarchies. Computers and Graphics, 61 (??):29–39, December 2016.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300899.

Karni:2004:CSB

[KG04] Zachi Karni and Craig Gotsman. Compression of softbody animation sequences. Computers and Graphics, 28 (1):25–34, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kersting:2020:DTS

[KG20] Felipe Einsfeld Kersting and Eduardo S. L. Gastal. Domain transform for spherical geometry images. Computers and Graphics, 93 (??):71–83, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301503.

Ku:2021:PSP

[KGB^{+}21] Tao Ku, Sam Galanakis, Bas Boom, Remco C. Veltkamp, Darshan Bangera, Shankar Gangisetty, Nikolaos Stagakis, Gerasimos Arvanitis, and Konstantinos Mous-

takas. SHREC 2021: 3D

point cloud change detection for street scenes. Computers and Graphics, 99 (??):192–200, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001369.

Kohler:2019:BSG

[KGGP19] Benjamin Köhler, Matthias Grothoff, Matthias Gutberlet, and Bernhard Preim. Bloodline: a system for the guided analysis of cardiac 4D PC-MRI data. Computers and Graphics, 82 (??):32–43, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300573.

Kostopoulos:2007:HPB

[KGK^{+}07] S. Kostopoulos, D. Glotsos, G. C. Kagadis, A. Daskalakis, P. Spyridonos, I. Kalatzis, M. Karamessini, T. Petsas, D. Cavouras, and G. Nikiforidis. A hybrid pixelbased classification method for blood vessel segmentation and aneurysm detection on CTA. Computers and Graphics, 31(3): 493–500, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000532.

King:1975:TAD

[KGM75] R. G. King, F. S. Gregorski, and P. C. Marth. Traffic analysis and display for the San Francisco Vessel Traffic System. Computers and Graphics, 1(4): 347–360, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kromker:1986:RMR

[KH86] Detlef Krömker and Georg Rainer Hofmann. Reconstructing and modelling Raphael’s School of Athens. Computers and Graphics, 10(4): 381–385, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kreylos:2000:DSO

[KH00] Oliver Kreylos and Bernd Hamann. Data structures for optimizing linear spline approximations. Computers and Graphics, 24(3): 353–361, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/31/article.
pdf.

Kim:2003:EBV

[KH03] Young J. Kim and Christoph M.

Hoffmann. Enhanced battlefield visualization for situation awareness. Computers and Graphics, 27(6): 873–885, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:2018:EEG

[KHK18] SeongKi Kim, Takahiro Harada, and Young J. Kim. Energy-efficient global illumination algorithms for mobile devices using dynamic voltage and frequency scaling. Computers and Graphics, 70(??): 198–205, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301103.

Kojekine:2003:STU

[KHS03] Nikita Kojekine, Ichiro Hagiwara, and Vladimir Savchenko. Software tools using CSRBFs for processing scattered data. Computers and Graphics, 27(2): 311–319, April 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Koepnick:2010:RRI

[KHS^{+}10] Steven Koepnick, Roger V. Hoang, Matthew R. Sgambati, Daniel S. Coming, Evan A. Suma, and William R. Sherman. RIST: Radiological Immersive Survey Training for two simultaneous users. Computers and Graphics, 34(6):665– 676, December 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931000141X.

Karaduzovic-Hadziabdic:2017:AME

[KHTM17] Kanita Karaduzović-Hadziabdić, Jasminka Hasić Telalović, and Rafal K. Mantiuk. Assessment of multi-exposure HDR image deghosting methods. Computers and Graphics, 63(??):1–17, April 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849317300110.

Kirste:1993:SIH

[Kir93] Thomas Kirste. SpacePicture — an interactive hypermedia satellite image archival system. Computers and Graphics, 17(3): 251–260, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kirste:1996:GEI

[Kir96] Thomas Kirste. Guest Editor’s introduction. Computers and Graphics, 20(5): 611–??, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kitching:1977:AGA

[Kit77]  Alan Kitching.  Antics

— graphic animation by computer. Computers and Graphics, 2(4):219–223, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kiyokawa:2019:FSS

[Kiy19] Kiyoshi Kiyokawa. Foreword to the special section on Cyberworlds 2018. Computers and Graphics, 82 (??):A7–A8, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301062.

Knierriem-Jasnoch:2001:EAC

[KJ01] Anette Knierriem-Jasnoch. Education: An approach to classify IT-based teaching and learning environments. Computers and Graphics, 25(5):899–907, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/42/abstract.
html.

Kye:2008:AMB

[KJ08] Heewon Kye and DongKyun Jeong. Accelerated MIP based on GPU using block clipping and occlusion query. Computers and Graphics, 32(3):283–292, June 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307002038.

Kjelldahl:1995:GEI

[Kje95] Lars Kjelldahl. Guest Editor’s introduction. Computers and Graphics, 19(2): 177–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Koa:2017:ISF

[KJS17] Ming Di Koa, Henry Johan, and Alexei Sourin. Interactive screenspace fragment rendering for direct illumination from area lights using gradient aware subdivision and radial basis function interpolation. Computers and Graphics, 64(??): 37–50, May 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300122.

Koa:2018:IRT

[KJS18] Ming Di Koa, Henry Johan, and Alexei Sourin. Interactive rendering of translucent materials under area lights using voxels and Poisson disk samples. Computers and Graphics, 71 (??):101–112, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300013.

Knierriem-Jasnoch:1996:RWF

[KJTS96] A. Knierriem-Jasnoch, B. Tritsch, and U. Schroeder. Reflection on WWW functionalities for educational purposes. Computers and Graphics, 20(3): 435–443, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=3&aid=9600013.

Kim:2008:AEP

[KK08] Donghyun Kim and Lee-Sup Kim. Area-efficient pixel rasterization and texture coordinate interpolation. Computers and Graphics, 32(6): 669–681, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000988.

Krecklau:2012:IMP

[KK12] Lars Krecklau and Leif Kobbelt. Interactive modeling by procedural highlevel primitives. Computers and Graphics, 36(5):376– 386, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000672.

Knott:2016:AAC

[KK16] Thomas C. Knott and Torsten W. Kuhlen. Accurate and adaptive contact modeling for multi-rate multi-point haptic rendering of static and deformable environments. Computers and Graphics, 57(??): 68–80, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300206.

Kiuchi:2021:SAM

[KK21] Shunji Kiuchi and Naoya Koizumi. Simulating the appearance of mid-air imaging with micro-mirror array plates. Computers and Graphics, 96(??):14– 23, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000297.

Kneiphof:2022:RTI

[KK22] Tom Kneiphof and Reinhard Klein. Real-time imagebased lighting of metallic and pearlescent car paints. Computers and Graphics, 105(??):36–45, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000619.

Kim:1994:VRI

[KKC94] Yong-Yil Kim, Heedong Ko, and Byung-Hong Choe. Virtual reality infrastructure and its application to telerobotics. Computers and Graphics, 18(5):667–673, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kohlhammer:2007:I

[KKE07] Jörn Kohlhammer, Daniel Keim, and David Ebert. Introduction. Computers and Graphics, 31(3):314– 315, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000659.

Kruppa:2023:STM

[KKH23] Kinga Kruppa, Roland Kunkli, and Miklós Hoffmann. A skinning technique for modeling artistic disk Bspline shapes. Computers and Graphics, 115(??):96– 106, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001255.

Kirchner:2003:SVU

[KKHS03] Peter D. Kirchner, James T. Klosowski, Peter Hochschild, and Richard Swetz. Scalable visualization using a network-attached video framebuffer. Computers and Graphics, 27(5):669– 680, October 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kendre:2023:SGA

[KKJ^{+}23] Prasad Pralhad Kendre, Kamalesh Kumar Kosalaraman, Sanjay Santhosh Kumar Jayasree, Sreehari Rajan, Akash Jayan, and Ramanathan Muthuganapathy.

SketchCADGAN: a generative

approach for completing partially drawn query sketches of engineering shapes to enhance retrieval system performance. Computers and Graphics, 115(??):55– 68, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001243.

Kim:2002:SSU

[KKL02] Sun-Jeong Kim, Chang-Hun Kim, and David Levin. Surface simplification using a discrete curvature norm. Computers and Graphics, 26(5):657–663, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/29/abstract.
html.

Keim:2006:ESW

[KKMT06] Daniel Keim, Jörn Kohlhammer, Thorsten May, and Jim Thomas. Event summary of the Workshop on Visual Analytics: June 4, 2005, Darmstadt Germany, jointly organized by University of Konstanz and Fraunhofer IGD. Computers and Graphics, 30(2):284–286, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000045.

Kozlikova:2021:FSS

[KKN^{+}21] Barbora Kozlíková, Michael Krone, Kay Nieselt, Renata G. Raidou, and Noeska N. Smit. Foreword: Special section on the Eurographics Workshop on Visual Computing for Biology and Medicine (EG VCBM) 2020. Computers and Graphics, 101(??):5–6, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001680.

Klimov:1988:SSM

[KKNT88] V. E. Klimov, V. V. Klishin, A. V. Neder, and A. S. Terentiev. A system of solid modeling for low-cost CAD systems. Computers and Graphics, 12(3–4): 407–413, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kiyokawa:2001:OST

[KKO01] Kiyoshi Kiyokawa, Yoshinori Kurata, and Hiroyuki Ohno. An optical see-through display for mutual occlusion with a real-time stereovision system. Computers and Graphics, 25(5):765– 779, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/31/abstract.
html.

Kwiatkowski:2023:URL

[KKPC23] Ariel Kwiatkowski, Vicky Kalogeiton, Julien Pettré, and Marie-Paule Cani. Understanding reinforcement learned crowds. Computers and Graphics, 110(??): 28–37, February 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002035.

Klimenko:1986:CGI

[KKS86] S. V. Klimenko, V. N.

Kochin, and A. V. Samarin. Computer graphics at IHEP. Computers and Graphics, 10 (1):63–69, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:1993:IGB

[KKS93] Hoi Sub Kim, Hong Oh Kim, and Sung Yong Shin. Image generation by Blaschke products in the unit disk. Computers and Graphics, 17(4): 489–494, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kohout:2004:POP

[KKŽ04] Josef Kohout, Ivana Kolingerová, and Jiří Žára. Practically oriented parallel Delaunay triangulation in E^{2} for computers with shared memory. Computers and Graphics, 28 (5):703–718, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Koo:2002:WAO

[KL02] Seungbum Koo and Kunwoo Lee. Wrap-around operation to make multi-resolution model of part and assembly. Computers and Graphics, 26(5):687–700, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/56/32/abstract.
                                        html.

Kang:2007:RTC

[KL07] Moon Koo Kang and Jeongjin Lee. A real-time cloth draping simulation algorithm using conjugate harmonic functions. Computers and Graphics, 31(2): 271–279, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001865.

Klassen:1991:SPT

[Kla91] R. Victor Klassen. A SIMD parallel trapezoid rasterization algorithm. Computers and Graphics, 15(4): 553–559, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Klein:1986:MSS

[Kle86] A. Klein. Modelling 3D shaded solids of arbitrary shape using an edge-oriented algorithm. Computers and Graphics, 10(4):327–331, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Klimenko:2000:GEI

[Kli00] S. V. Klimenko. Guest Editors’ Introduction. Computers and Graphics, 24(1): 5–7, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/27/article.
pdf.

Kim:2011:ECH

[KLKE11] Yong-Joon Kim, Jieun Lee, Myung-Soo Kim, and Gershon Elber. Efficient convex hull computation for planar freeform curves. Computers and Graphics, 35(3): 698–705, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000677.

Kuang:2015:MFT

[KLL^{+}15] Zhenzhong Kuang, Zongmin Li, Qian Lv, Tian Weiwei, and Yujie Liu. Modal function transformation for isometric 3D shape representation. Computers and Graphics, 46(??):209– 220, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001216.

Kim:2001:FEB

[KLP01] Jun Sung Kim, Jong Hyun Lee, and Kyu Ho Park. A fast and efficient bump mapping algorithm by angular perturbation. Computers and Graphics, 25(3): 401–407, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
31/abstract.html;   http:

//www.elsevier.nl/gej-
ng/10/13/20/57/32/31/article.
pdf.

Kim:2012:RTI

[KLW12] Kiyoung Kim, Vincent Lepetit, and Woontack Woo. Real-time interactive modeling and scalable multiple object tracking for AR. Computers and Graphics, 36(8): 945–954, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200132X.

Kim:2014:FBS

[KM14] Byung Chul Kim and Duhwan Mun. Feature-based simplification of boundary representation models using sequential iterative volume decomposition. Computers and Graphics, 38 (??):97–107, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001726.

Kougianos:2021:LSR

[KM21] Gerasimos Kougianos and Konstantinos Moustakas. Large-scale ray traced water caustics in real-time using cascaded caustic maps. Computers and Graphics, 98 (??):255–267, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001230.

Kumar:1999:HBF

[KMGL99] Subodh Kumar, Dinesh Manocha, William Garrett, and Ming Lin. Hierarchical back-face computation. Computers and Graphics, 23 (5):681–692, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
32/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/32/abstract.
html.

Kanus:1997:ICT

[KMS^{+}97] U. Kanus, M. Meißner, W. Straßer, H. Pfister, A. Kaufman, R. Amerson, R. J. Carter, B. Culbertson, P. Kuekes, and G. Snider. Implementations of Cube-4 on the Teramac custom computing machine. Computers and Graphics, 21(2): 199–208, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600083.

Konev:2018:FCV

[KMV^{+}18] Artem Konev, Manuel Matusich, Ivan Viola, Hendrik Schulze, Daniel Cornel, and Jürgen Waser. Fast cutaway visualization of subterrain tubular networks. Computers and Graphics, 75 (??):11–24, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318301031.

Kuchera-Morin:2014:IFS

[KMWW^{+}14] JoAnn Kuchera-Morin, Matthew Wright, Graham Wakefield, Charles Roberts, Dennis Adderton, Behzad Sajadi, Tobias Höllerer, and Aditi Majumder. Immersive full-surround multi-user system design. Computers and Graphics, 40(??): 10–21, May 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000090.

Klement:1988:EMP

[KN88] Kornel Klement and Horst Nowacki. Exchange of model presentation information between CAD systems. Computers and Graphics, 12(2): 173–180, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:2011:BAC

[KNC11] Junho Kim, Changwoo Nam, and Sungyul Choe. Bayesian AD coder: Mesh-aware valence coding for multiresolution meshes. Computers and Graphics, 35(3): 713–718, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000628.

Kuhlen:2020:FSS

[KNDT20] Torsten W. Kuhlen, Huyen Nguyen, Thierry Duval, and Tomas Trescak. Foreword to the special section on the 25th Symposium on Virtual Reality Software and Technology. Computers and Graphics, 89(??): A6–A7, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300467.

Knittel:1995:SAV

[Kni95a] Günter Knittel. A scalable architecture for volume rendering. Computers and Graphics, 19(5):653–665, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500044.

Knittel:1995:VDF

[Kni95b] Günter Knittel. A VLSIdesign for fast vector normalization. Computers and Graphics, 19(2):261– 271, March–April 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400152.

Kwon:2014:ETA

[KNMP14] Hyuck-Joo Kwon, Jae-Ho Nah, Dinesh Manocha, and Woo-Chan Park. Effective traversal algorithms and hardware architecture for pyramidal inverse displacement mapping. Computers and Graphics, 38(??): 140–149, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001659.

Knott:1987:CPF

[Kno87] Gary D. Knott. Computing polygon fill-lines. Computers and Graphics, 11 (1):21–25, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Koh:1996:WDV

[Koh96a] Eng-Kiat Koh. WAVEVI-SIONS: a desktop virtual reality software. Computers and Graphics, 20(1):69– 75, January–February 1996.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1&aid=9500094.

Kohl:1996:HCA

[Koh96b] Hans Werner Kohl. Hiddencurve algorithm for correct grid surface representation of functions of two variables. Computers and Graphics, 20(2): 243–261, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=2&aid=9500099.

Kohl:1997:FHS

[Koh97] Hans Werner Kohl. Firsthit speed up in ray tracing surfaces of functions of two variables. Computers and Graphics, 21(6):825–839, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700061.

Kong:1989:DFG

[Kon89] T. Y. Kong. A digital fundamental group. Computers and Graphics, 13(2): 159–166, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kim:2017:NIS

[KONS17] Seungki Kim, Yutaka Ohtake, Yukie Nagai, and Hiromasa Suzuki. A novel interpolation scheme for dual marching cubes on octree volume fraction data. Computers and Graphics, 66 (??):169–178, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300705.

Kornhauser:1987:ETP

[Kor87] Alain L. Kornhauser. Exploring transportation problems using interactive computer graphics. Computers and Graphics, 11(3): 231–239, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kornel:1990:PT

[Kor90] Klement Kornel. 2D and 3D perspective transformations. Computers and Graphics, 14 (1):117–124, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kotzauer:1990:GEI

[Kot90] A. Kotzauer. Guest Editor’s introduction. Computers and Graphics, 14(3– 4):365–??, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kjelldahl:1995:SHD

[KP95] Lars Kjelldahl and Martin Prime. A study on how depth perception is affected by different presentation methods of 3D objects on a 2D display. Computers and Graphics, 19(2): 199–202, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400143.

Korfiatis:1998:TDO

[KP98] Ioannis Korfiatis and Yakup Paker. Three-dimensional object metamorphosis through energy minimization. Computers and Graphics, 22(2– 3):195–202, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/2-3/541.pdf.

Kjelldahl:2005:SIE

[KP05] Lars Kjelldahl and Zhigeng Pan. Special issue: Education. Computers and Graphics, 29(2):235–236, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Karciauskas:2009:ACC

[KP09] Keştutis Karciauskas and Jörg Peters. Assembling curvature continuous surfaces from triangular patches. Computers and Graphics, 33(3):204–210, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000302.

Karciauskas:2015:SMS

[KP15] Kestutis Karciauskas and Jörg Peters. Smooth multisided blending of biquadratic splines. Computers and Graphics, 46(??):172–185, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000922.

Karciauskas:2018:RCS

[KP18] Kestutis Karciauskas and Jörg Peters. Rapidly contracting subdivision yields finite, effectively C^{2} surfaces. Computers and Graphics, 74 (??):182–190, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300839.

Karciauskas:2019:RSS

[KP19] Kestutis Karciauskas and Jörg Peters. Refinable smooth surfaces for locally quad-dominant meshes with T -gons. Computers and Graphics, 82(??):193– 202, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300743.

Karciauskas:2022:IRR

[KP22a] Kestutis Karciauskas and Jörg Peters. An improved refinement rule for multisided faces. Computers and Graphics, 102(??):370– 379, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002193.

Karciauskas:2022:LRP

[KP22b] Kestutis Karciauskas and Jörg Peters. Localized remeshing for polyhedral splines. Computers and Graphics, 106(??):58–65, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000905.

Kammoun:2012:SBO

[KPA12] Aymen Kammoun, Frédéric Payan, and Marc Antonini. Sparsity-based optimization of two lifting-based wavelet transforms for semi-regular mesh compression. Computers and Graphics, 36(4): 272–282, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000209.

Kylanpaa:1996:NAI

[KPB96] Markku Kylänpää, Olli Pihlajamaa, and Martin Bergenwall. Nomadic access to information services by a GSM phone. Computers and Graphics, 20(5):651–658, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600039.

Kozlikova:2020:FSS

[KPBR20] Barbora Kozlíková, Bernhard Preim, Katja Bühler, and Renata Raidou. Foreword: Special section on the Eurographics Workshop on Visual Computing for Biology and Medicine (EG VCBM) 2019. Computers and Graphics, 92(??):A4– A5, November 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301278.

Karabassi:2003:EMM

[KPFT03] E. A. Karabassi, G. Papaioannou, C. Fretzagias, and T. Theoharis. Exploiting multiresolution models to accelerate ray tracing. Computers and Graphics, 27 (1):91–98, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kuester:2005:ICF

[KPH^{+}05] Falko Kuester, Mark E. Phair, Tara C. Hutchinson, Visualization, and Interactive Systems Group. Image centric finite element simulation. Computers and Graphics, 29(3):379– 392, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kolhoff:2008:CBI

[KPL08] Philipp Kolhoff, Jacqueline Preuß, and Jörn Loviscach. Content-based icons for music files. Computers and Graphics, 32(5):550– 560, October 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930800006X.

Kodama:2018:CEP

[KPMT18] Shuhei Kodama, Pierre Poulin, Tomoaki Moriya, and Tokiichiro Takahashi. Creativity enhancement of painterly rendering using a suggestive interface. Computers and Graphics, 71(??): 42–54, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301760.

Keim:2004:PBV

[KPSN04] Daniel A. Keim, Christian Panse, Mike Sips, and Stephen C. North. Pixel based visual data mining of geo-spatial data. Computers and Graphics, 28(3): 327–344, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kaul:1992:SID

[KR92] Anil Kaul and Jarek Rossignac.

Solid-interpolating deformations. construction and animation of PIPS. Computers and Graphics, 16(1): 107–115, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kirste:1996:PMM

[KR96] Thomas Kirste and Uwe Rauschenbach. A presentation model for mobile information visualization. Computers and Graphics, 20(5):669–681, September– October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600041.

Krause:2010:FFP

[Kra10] Philipp Klaus Krause. ftc — floating precision texture compression. Computers and Graphics, 34 (5):594–601, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000890.

KuffnerdosAnjos:2023:ESD

[KRA^{+}23a] Rafael Kuffner dos Anjos, Richard Andrew Roberts, Benjamin Allen, Joaquim Jorge, and Ken Anjyo. Erratum to “Saliency Detection for Large-Scale Mesh Decimation” [Comput. Graph. 111 (2023) 63–76]. Computers and Graphics, 114 (??):414, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002066.

See [KRA^{+}23b].

KuffnerdosAnjos:2023:SDL

[KRA^{+}23b] Rafael Kuffner dos Anjos, Richard Andrew Roberts, Benjamin Allen, Joaquim Jorge, and Ken Anjyo. Saliency detection for largescale mesh decimation. Computers and Graphics, 111 (??):63–76, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000134.

See erratum [KRA^{+}23a].

Kress:1993:EVT

[Kre93] James G. Kress. Experiences in Visual Thinking by Robert H. McKim. Computers and Graphics, 17(1):105– ??, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kil:2006:WBM

[KRK^{+}06] Yong Joo Kil, Pietro Renzulli, Oliver Kreylos, Bernd Hamann, Guiseppe Monno, and Oliver G. Staadt. 3D warp brush modeling. Computers and Graphics, 30 (4):610–618, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000781.

Kromker:1991:GEI

[Kro91] Detlef Kromker. Guest Editor’s introduction. Computers and Graphics, 15(4): 463–??, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kromker:1996:GEI

[Kro96] Detlef Kromker. Guest Editor’s introduction. Computers and Graphics, 20 (3):413–??, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Krober:1998:SGP

[Krö98]  Günter Kröber.  Struc-

ture generation by palindromization. Computers and Graphics, 22(2–3):307–317, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/573.pdf.

Klein:1991:TBU

[KRP^{+}91] Michael Klein, Otto E. Rössler, Jürgen Parisi, Joachim Peinke, Gerold Baier, Claus Kahlert, and John L. Hudson. Toward a better understanding of fractality in nature. Computers and Graphics, 15(4):583–596, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kry:2012:IKE

[KRRS12] Paul G. Kry, Cyrus Rahgoshay, Amir Rabbani, and Karan Singh. Inverse kinodynamics: Editing and constraining kinematic approximations of dynamic motion. Computers and Graphics, 36 (8):904–915, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200146X.

Kruckeberg:1984:PCT

[Kru84]  F. Kruckeberg.  Princi-

ples of conformity testing. Computers and Graphics, 8 (1):63–65, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kruszewski:1999:AST

[Kru99a] Paul Kruszewski. An algorithm for sculpting trees. Computers and Graphics, 23 (5):739–749, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
37/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/37/abstract.
html.

Kruszewski:1999:CGP

[Kru99b] Paul Kruszewski. Chaos and graphics — a probabilistic technique for the synthetic imagery of lightning. Computers and Graphics, 23(2):287–293, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        2/681.pdf.

Kry:2014:FSS

[Kry14] Paul G. Kry. Foreword to Special Section on Graphics Interaction. Computers and Graphics, 45(??): A2–A3, December 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001241.

Karg:1984:HRC

[KS84] R. Karg and J. Schroter. High-resolution colour graphic workstation. Computers and Graphics, 8(2):195–197, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kumar:1986:GSP

[KS86] M. P. Subodh Kumar and Y. N. Srikant. Graphical simulation of Petri nets. Computers and Graphics, 10 (3):225–228, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kerr:1991:SSR

[KS91] Randall L. Kerr and Phillip C.-Y. Sheu. SPARQ. spatial relationship query based on spatial decomposition. Computers and Graphics, 15(2): 267–284, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kjelldahl:1996:EYS

[KS96] Lars Kjelldahl and Yngve Sundblad. Experience from 10 years of student projects oriented towards graphic interaction. Computers and Graphics, 20(3): 463–471, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600017.

Kanev:1998:DSI

[KS98] Kamen Kanev and Tomoyuki Sugiyama. Design and simulation of interactive 3D computer games. Computers and Graphics, 22(2–3):281–300, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/572.pdf.

Kim:2001:FVR

[KS01a] Tae-Young Kim and Yeong Gil Shin. Fast volume rendering with interactive classification. Computers and Graphics, 25(5):819–831, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/34/36/abstract.
                                        html.

Kirste:2001:EII

[KS01b] Thomas Kirste and Heidrun Schumann. Editorial: Intelligent interactive assistance and mobile multimedia computing. Computers and Graphics, 25(4):551– 553, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/33/27/abstract.
                                        html.

Kaufmann:2003:MGE

[KS03a] Hannes Kaufmann and Dieter Schmalstieg. Mathematics and geometry education with collaborative augmented reality. Computers and Graphics, 27(3): 339–345, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Knittel:2003:I

[KS03b] Günter Knittel and Bengt-Olaf Schneider. Introduction. Computers and Graphics, 27(5):667–668, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kjelldahl:2004:VPC

[KS04] Lars Kjelldahl and Beatriz Sousa Santos. Visual perception in computer graphics courses. Computers and Graphics, 28(3):451– 456, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kirste:2006:I

[KS06] Thomas Kirste and Heidi Schumann. Introduction. Computers and Graphics, 30 (5):679–680, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600118X.

Kyratzi:2009:EPS

[KS09] Sofia Kyratzi and Nickolas Sapidis. Extracting a polyhedron from a singleview sketch: Topological construction of a wireframe sketch with minimal hidden elements. Computers and Graphics, 33(3): 270–279, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900020X.

Kramer:2015:MSA

[KS15] Michel Krämer and Ivo Senner. A modular software architecture for processing of big geospatial data in the cloud. Computers and Graphics, 49(??): 69–81, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000138.

Kita:2020:CDG

[KS20] Naoki Kita and Takafumi Saito. Computational design of generalized centrifugal puzzles. Computers and Graphics, 90(??):21– 28, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030056X.

Knibbe:2015:VST

[KSF15] Jarrod Knibbe, Sue Ann Seah, and Mike Fraser. VideoHandles: Searching through action camera videos by replicating hand gestures. Computers and Graphics, 48(??):99–106, May 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000059.

Klemm:2017:HAP

[KSH17] Martin Klemm, Fabian Seebacher, and Harald Hoppe. High accuracy pixel-wise spatial calibration of optical see-through glasses. Computers and Graphics, 64(??): 51–61, May 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730016X.

Kim:2019:BWI

[KSH^{+}19] Kangsoo Kim, Ryan Schubert, Jason Hochreiter, Gerd Bruder, and Gregory Welch. Blowing in the wind: Increasing social presence with a virtual human via environmental airflow interaction in mixed reality. Computers and Graphics, 83 (??):23–32, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301001.

Knittel:1996:HST

[KSKS96] G. Knittel, A. Schilling, A. Kugler, and W. Straßer. Hardware for superior texture performance. Computers and Graphics, 20(4): 475–481, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600019.

Kara:2007:EUE

[KSM07] Levent Burak Kara, Kenji Shimada, and Sarah D. Marmalefsky. An evaluation of user experience with a sketch-based 3D modeling system. Computers and Graphics, 31(4):580– 597, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000982.

Klinker:1999:OBD

[KSR99] G. Klinker, D. Stricker, and D. Reiners. Optically based direct manipulation for augmented reality. Computers and Graphics, 23(6): 827–830, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/35/35/article.
                                        pdf.

Kjelldahl:1995:E

[KT95] Lars Kjelldahl and Jose Teixeira. Editorial. Computers and Graphics, 19(2): 325–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Katz:2017:VER

[KT17] Sagi Katz and Ayellet Tal.

On visibility and emptyregion graphs. Computers and Graphics, 66(??): 45–52, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300560.

Knecht:2012:RSM

[KTMW12] Martin Knecht, Christoph Traxler, Oliver Mattausch, and Michael Wimmer. Reciprocal shading for mixed reality. Computers and Graphics, 36(7):846–856, November 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001148.

Kotzauer:1990:GSG

[KUMW90] A. Kotzauer, B. Urban, M. Mikut, and J. Winkler. Graphics standards in the GDR. Computers and Graphics, 14(3–4): 367–371, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kuo:2001:AEQ

[Kuo01] M. H. Kuo. Automatic extraction of quadric surfaces from wire-frame models. Computers and Graphics, 25(1):109–119, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/35/article.
pdf.

Kurzhals:2024:AET

[Kur24] Kuno Kurzhals. Anonymizing eye-tracking stimuli with stable diffusion. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000256.

Ku:2020:PSP

[KVB^{+}20] Tao Ku, Remco C. Veltkamp, Bas Boom, David Duque-Arias, Santiago Velasco-Forero, Jean-Emmanuel Deschaud, Francois Goulette, Beatriz Marcotegui, Sebastián Ortega, Agustín Trujillo, José Pablo Suárez, José Miguel Santana, Cristian Ramírez, Kiran Akadas, and Shankar Gangisetty. SHREC 2020: 3D point cloud semantic segmentation for street scenes. Computers and Graphics, 93(??):13– 24, December 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301400.

Koh:1996:SRD

[KW96] Aik-Siong Koh and Alan Wegienka. Software report: Dynamic designer

— a multibody dynamics software. Computers and Graphics, 20(1):77– 78, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9688775.

Kwok:2014:DCV

[KW14] Tsz-Ho Kwok and Charlie C. L. Wang. Domain construction for volumetric cross-parameterization. Computers and Graphics, 38 (??):86–96, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001672.

Krumpen:2017:OCS

[KWK17] Stefan Krumpen, Michael Weinmann, and Reinhard Klein. OctreeBTFs — a compact, seamless and distortion-free reflectance representation. Computers and Graphics, 68(??):21– 31, November 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301309.

Kluge:2023:PRI

[KWK23] Markus Kluge, Tim Weyrich, and Andreas Kolb. Progressive refinement imaging with depth-assisted disparity correction. Computers and Graphics, 115 (??):446–460, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001656.

Ko:1997:FMG

[KY97] Byong K. Ko and Hyun S. Yang. Finger mouse and gesture recognition system as a new human computer interface. Computers and Graphics, 21(5):555–561, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=5&aid=9700034.

Kang:2019:FRC

[KYKK19] Yunku Kang, Seung-Hyun Yoon, Min-Ho Kyung, and Myung-Soo Kim. Fast and robust computation of the Hausdorff distance between triangle mesh and quad mesh for near-zero cases. Computers and Graphics, 81(??): 61–72, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300378.

Kim:2015:PTP

[KYL15] Young J. Kim, Mincheol Yoon, and Taekhee Lee. Probabilistic triangles for point set surfaces. Computers and Graphics, 51 (??):26–34, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000552.

Kwon:2012:PRM

[KYM12] Yunmi Kwon, Heekyung Yang, and Kyungha Min. Pencil rendering on 3D meshes using convolution. Computers and Graphics, 36 (8):930–944, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001379.

Krichenbauer:2017:EEP

[KYT^{+}17] Max Krichenbauer, Goshiro Yamamoto, Takafumi Taketomi, Christian Sandor, Hirokazu Kato, and Steven Feiner. Evaluating the effect of positional head-tracking on task performance in 3D modeling user interfaces. Computers and Graphics, 65(??):22–30, June 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300407.

Klein:2004:PCS

[KZ04] Jan Klein and Gabriel Zachmann. Point cloud surfaces using geometric proximity graphs. Computers and Graphics, 28(6): 839–850, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lammarsch:2014:MTU

[LAB^{+}14] T. Lammarsch, W. Aigner, A. Bertone, S. Miksch, and A. Rind. Mind the time: Unleashing temporal aspects in pattern discovery. Computers and Graphics, 38 (??):38–50, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001489.

Luu:2019:ESC

[LAE^{+}19] Thu Huong Luu, Christian Altenhofen, Tobias Ewald, André Stork, and Dieter Fellner. Efficient slicing of Catmull–Clark solids for 3D printed objects with functionally graded material. Computers and Graphics, 82 (??):295–303, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300846.

Lafon:1994:GPP

[Laf94] Jean-Claude Lafon. 2D and 3D graphics with PostScript and PHIGS. Computers and Graphics, 18(3):295– 298, May–June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lakhtakia:1989:SGD

[Lak89] Akhlesh Lakhtakia. A simple gasket derived from prime numbers. Computers and Graphics, 13(1):57–58, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lakhtakia:1991:JSS

[Lak91] Akhlesh Lakhtakia. Julia sets of switched processes. Computers and Graphics, 15 (4):597–599, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lee:2011:DSS

[LAL11] Hyunjun Lee, Minsu Ahn, and Seungyong Lee. Displaced subdivision surfaces of animated meshes. Computers and Graphics, 35(3): 532–541, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000719.

Lamiche:1983:TG

[Lam83] D. Lamiche. Teletel graphics. Computers and Graphics, 7(1):45–50, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lam:1987:PFC

[Lam87] K. Lam. Potential flow calculation by surface vorticity method and computer graphics. Computers and Graphics, 11(1):35–47, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lam:1999:MRI

[Lam99] Stephen Wang-Cheung Lam. Multiresolution representation of interval surfaces using subdivision wavelet transform and linear programming. Computers and Graphics, 23(4):555– 572, August 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/36/article.
pdf.

Lam:2000:EGP

[Lam00] Stephen Wang-Cheung Lam.

Extensions of the general polar value based control point specification method in constructing tensor product B-spline surfaces. Computers and Graphics, 24 (4):493–507, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/27/article.
pdf.

Larsson:2006:DBV

[LAM06] Thomas Larsson and Tomas Akenine-Möller. A dynamic bounding volume hierarchy for generalized collision detection. Computers and Graphics, 30(3): 450–459, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000689.

Lang:1988:ISV

[Lan88] Johann Lang. On illuminations of C^{2}-surfaces in vector graphic description. Computers and Graphics, 12 (1):33–38, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Landini:1997:FAL

[Lan97]  Gabriel Landini.  Frac-

tal anamorphosis: Look up table transform using self-affine series. Computers and Graphics, 21(1): 105–111, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        cag/cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=1&aid=9600074.

Laramee:2003:FFI

[Lar03] Robert S. Laramee. FIRST: a flexible and interactive resampling tool for CFD simulation data. Computers and Graphics, 27(6): 905–916, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lastra:1991:TSS

[Las91] Gerardo Leon Lastra. Topological sort of SSI with minimal searching. Computers and Graphics, 15(3): 389–396, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Laurillard:1977:DDC

[Lau77] D. M. Laurillard. Design and development of CAL materials in undergraduate science. Computers and Graphics, 2(4):241–247, 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

LaViola:2007:IEM

[LaV07] Joseph J. LaViola, Jr.

An initial evaluation of MathPad^{2}: a tool for creating dynamic mathematical illustrations. Computers and Graphics, 31 (4):540–553, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700101X.

LeCureux:1975:GMU

[LB75] F. LeCureux and J. Burnett. Graphical methods used in the numerical solution of Jeffery–Hamel flow at fixed flow rates. Computers and Graphics, 1(2– 3):233–239, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lockyer:2012:AMT

[LB12] Matt Lockyer and Lyn Bartram. Affective motion textures. Computers and Graphics, 36(6):776– 790, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001100.

Luciano:2019:GGF

[LB19] Lorenzo Luciano and A. Ben Hamza. A global geometric framework for 3D shape retrieval using deep learning. Computers and Graphics, 79(??):14–23, April 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830195X.

Litman:2011:DGM

[LBB11] Roee Litman, Alexander M. Bronstein, and Michael M. Bronstein. Diffusion-geometric maximally stable component detection in deformable shapes. Computers and Graphics, 35(3):549–560, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000501.

Litman:2012:SVF

[LBB12] R. Litman, A. M. Bronstein, and M. M. Bronstein. Stable volumetric features in deformable shapes. Computers and Graphics, 36 (5):569–576, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000738.

Le:2017:MVR

[LBD17] Truc Le, Giang Bui, and Ye Duan. A multi-view recurrent neural network for 3D mesh segmentation. Computers and Graphics, 66 (??):103–112, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300602.

Langrana:1997:HPU

[LBLD97] Noshir Langrana, Grigore Burdea, Jumoke Ladeji, and Michael Dinsmore. Human performance using virtual reality tumor palpation simulation. Computers and Graphics, 21(4): 451–458, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700021.

Laycock:2011:AAM

[LBLD11] S. D. Laycock, P. G. Brown, R. G. Laycock, and A. M. Day. Aligning archive maps and extracting footprints for analysis of historic urban environments. Computers and Graphics, 35(2): 242–249, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000033.

Leon:2016:CSD

[LBLV16] Vincent Léon, Nicolas Bonneel, Guillaume Lavoué, and Jean-Philippe Vandeborre. Continuous semantic description of 3D meshes. Computers and Graphics, 54 (??):47–56, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001211.

Ludwig:2015:STM

[LBTM15] Michael Ludwig, Seth Berrier, Michael Tetzlaff, and Gary Meyer. 3D shape and texture morphing using 2D projection and reconstruction. Computers and Graphics, 51 (??):146–156, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000527.

Liu:2014:ETB

[LBV14] Tianyang Liu, Fatma Bouali, and Gilles Venturini. EXOD: a tool for building and exploring a large graph of open datasets. Computers and Graphics, 39(??): 117–130, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001933.

Liarokapis:2018:FSS

[LC18] Fotis Liarokapis and Yiorgos L. Chrysanthou. Foreword to the special section on serious games and virtual environments. Computers and Graphics, 75 (??):A1–A2, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318301055.

Liarokapis:2023:FSS

[LC23] Fotis Liarokapis and Yiorgos Chrysanthou. Foreword to the special section on emerging computer graphics. Computers and Graphics, 116 (??):A4–A5, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002613.

Livesu:2019:SMT

[LCA19] Marco Livesu, Daniela Cabiddu, and Marco Attene. slice2mesh: a meshing tool for the simulation of additive manufacturing processes. Computers and Graphics, 80(??): 73–84, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300275.

Lira:2002:MMF

[LCCM02] William M. Lira, Paulo Roma Cavalcanti, Luiz C. G. Coelho, and Luiz F. Martha. A modeling methodology for finite element mesh generation of multi-region models with parametric surfaces. Computers and Graphics, 26 (6):907–918, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Leon:2004:RTD

[LCCS04] Enrique Leon, Graham Clarke, Victor Callaghan, and Francisco Sepulveda. Real-time detection of emotional changes for inhabited environments. Computers and Graphics, 28 (5):635–642, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2015:IBA

[LCD15] Baoquan Liu, Gordon J. Clapworthy, and Feng Dong. IsoBAS: a binary accelerating structure for fast isosurface rendering on GPUs. Computers and Graphics, 48 (??):60–70, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000102.

Lenoir:2006:IPB

[LCDN06] Julien Lenoir, Stephane Cotin, Christian Duriez, and Paul Neumann. Interactive physically-based simulation of catheter and guidewire. Computers and Graphics, 30(3):416–422, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000641.

LeNaour:2019:KMS

[LCG19] Thibaut Le Naour, Nicolas Courty, and Sylvie Gibet. Kinematics in the metric space. Computers and Graphics, 84(??):13– 23, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301232.

Liu:2024:LGP

[LCGD24] Shanghuan Liu, Xunhao Chen, Shaoyan Gai, and Feipeng Da. Local geometryperceptive mesh convolution with multi-ring receptive field. Computers and Graphics, 118(??):123– 132, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003047.

LuqueRuiz:1992:GIM

[LCGN92] I. Luque Ruiz, J. L. Cruz Soto, and M. A. Gomez-Nieto. GESIM. image management system for TIFF structures. Computers and Graphics, 16(3):325– 329, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Larsson:2016:PCO

[LCK16] Thomas Larsson, Gabriele Capannini, and Linus Källberg. Parallel computation of optimal enclosing balls by iterative orthant scan. Computers and Graphics, 56 (??):1–10, May 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300024.

Levallois:2015:SSF

[LCL15] Jérémy Levallois, David Coeurjolly, and Jacques-Olivier Lachaud. Scale-space feature extraction on digital surfaces. Computers and Graphics, 51(??):177– 189, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000709.

Li:2021:AHA

[LCL^{+}21] Zhong Li, Lele Chen, Celong Liu, Fuyao Zhang, Zekun Li, Yu Gao, Yuanzhou Ha, Chenliang Xu, Shuxue Quan, and Yi Xu. Animated 3D human avatars from a single image with GAN-based texture inference. Computers and Graphics, 95(??):81–91, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000029.

Liu:2023:SAE

[LCL^{+}23] Yifan Liu, Jincai Chen, Ping Lu, Chuanbo Zhu, Yugen Jian, Chao Sun, and Han Liang. Snowed autoencoders are efficient snow removers. Computers and Graphics, 114(??):73–85, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000729.

Liu:2018:PMN

[LCW^{+}18] Bin Liu, Junjie Cao, Weiming Wang, Ning Ma, Bo Li, Ligang Liu, and Xiuping Liu. Propagated mesh normal filtering. Computers and Graphics, 74(??):119– 125, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300621.

Li:2014:CFI

[LCWZ14] Faming Li, Xiaowu Chen, Lin Wang, and Qinping Zhao. Canopy-frame interactions for umbrella simulation. Computers and Graphics, 38(??):320–327, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001635.

Liu:2023:PNP

[LCX^{+}23] Yilin Liu, Ruiqi Cui, Ke Xie, Minglun Gong, and Hui Huang. PA-Net: Plane attention network for realtime urban scene reconstruction. Computers and Graphics, 115(??):254–262, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001516.

Lu:2020:PPD

[LCXL20] Qiang Lu, Chao Chen, Wenjun Xie, and Yuetong Luo. PointNGCNN: Deep convolutional networks on 3D point clouds with neighborhood graph filters. Computers and Graphics, 86 (??):42–51, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301748.

Li:2011:DQD

[LCZ^{+}11] Er Li, Wujun Che, Xiaopeng Zhang, Yi-Kuan Zhang, and Bo Xu. Direct quad-dominant meshing of point cloud via global parameterization. Computers and Graphics, 35(3): 452–460, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000604.

Lee:1978:CGH

[LD78] Mansuk Lee and Robert V.

DiMarco. Computer graphics with hidden surfaces

— graphs. Computers and Graphics, 3(2–3):85–91, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Laycock:2003:RGU

[LD03] R. G. Laycock and A. M.

Day. Rapid generation of urban models. Computers and Graphics, 27(3):423– 433, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Laycock:2005:IHF

[LD05] S. D. Laycock and A. M.

Day. Incorporating haptic feedback for the simulation of a deformable tool in a rigid scene. Computers and Graphics, 29(3):341– 351, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lavoue:2009:SSS

[LD09] Guillaume Lavoué and Florent Dupont. Semi-sharp subdivision surface fitting based on feature lines approximation. Computers and Graphics, 33(2):151– 161, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000168.

Laura:2011:CAM

[LD11] Cristina Oyarzun Laura and Klaus Drechsler. Computer assisted matching of anatomical vessel trees. Computers and Graphics, 35(2): 299–311, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310002074.

Lister:2012:SBA

[LD12] Wayne Daniel Lister and Andy Day. Stream-based animation of real-time crowd scenes. Computers and Graphics, 36(6):651–657, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000374.

Lavoue:2007:SSW

[LDD07] Guillaume Lavoué, Florence Denis, and Florent Dupont. Subdivision surface watermarking. Computers and Graphics, 31(3): 480–492, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000507.

Ling:1996:MAM

[LDG96] Li Ling, M. Damodaran, and Robert K. L. Gay. A model for animating the motion of cloth. Computers and Graphics, 20(1): 137–156, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500073.

Lemeunier:2022:RLM

[LDLD22] Clément Lemeunier, Florence Denis, Guillaume Lavoué, and Florent Dupont. Representation learning of 3D meshes using an Autoencoder in the spectral domain. Computers and Graphics, 107(??):131–143, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001285.

Lemeunier:2023:SST

[LDLD23] Clément Lemeunier, Florence Denis, Guillaume Lavoué, and Florent Dupont. SpecTrHuMS: Spectral transformer for human mesh sequence learning. Computers and Graphics, 115 (??):191–203, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001309.

Langlotz:2011:RDT

[LDM^{+}11] Tobias Langlotz, Claus Degendorfer, Alessandro Mulloni, Gerhard Schall, Gerhard Reitmayr, and Dieter Schmalstieg. Robust detection and tracking of annotations for outdoor augmented reality browsing. Computers and Graphics, 35 (4):831–840, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001075.

Lan:2021:PPW

[LDS^{+}21] Yuqing Lan, Yao Duan, Yifei Shi, Hui Huang, and Kai Xu. 3DRM: Pair-wise relation module for 3D object detection. Computers and Graphics, 98(??): 58–70, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000753.

Lee:2008:NPP

[LdSP^{+}08] WeeSan Lee, Ruwanee de Silva, Eric J. Peterson, Robert C. Calfee, and Thomas F. Stahovich. Newton’s Pen: a pen-based tutoring system for statics. Computers and Graphics, 32(5):511– 524, October 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000587.

Lorenzetto:2002:FTT

[LDT02] Gian Paolo Lorenzetto, Amitava Datta, and Richard C. Thomas. A fast trapezoidation technique for planar polygons. Computers and Graphics, 26(2):281– 289, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/35/abstract.
html.

LeMaitre:1977:PRE

[Le 77] J. Le Maitre. A propos d’une rationalisation éventuelle des systèmes et réseaux de données relatifs aux collections d’objets, matériels en sciences humaines. Computers and Graphics, 2(4):

256, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849377900292.

Leach:1985:GCS

[Lea85] R. J. Leach. Graphical control systems and multiple displays. Computers and Graphics, 9(4):415–422, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Leach:1987:EPU

[Lea87] Ronald J. Leach. Evaluating the performance of a user interface. Computers and Graphics, 11(2): 141–146, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lee:2001:DRB

[Lee01] Jintae Lee. Diffusion rendering of black ink paintings using new paper and ink models. Computers and Graphics, 25(2):295–308, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/29/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/29/36/article.
                                        pdf.

Les:2001:PMS

[Les01] Z. Les. The processing method as a set of the image transformations in shape understanding. Computers and Graphics, 25(2): 223–233, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/30/article.
pdf.

Les:2002:SUS

[Les02] Z. Les. Shape understanding system: Understanding the thin object. Computers and Graphics, 26(6): 951–970, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849302001826.

Lewallen:1975:RGP

[Lew75a] R. A. Lewallen. Reducing graphics programming effort through structured support. Computers and Graphics, 1(2–3):191–194, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lewis:1975:TIS

[Lew75b] J. W. Lewis. TREE: An interactive system for editing tree structures. Computers and Graphics, 1(1): 65–68, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Leys:2002:BAA

[Ley02] Jos Leys. Biomorphic art: an artist’s statement. Computers and Graphics, 26(6): 977–979, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Leys:2005:SIF

[Ley05] Jos Leys. Sphere inversion fractals. Computers and Graphics, 29(3):463– 466, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Leys:2007:DEI

[Ley07] Jos Leys. The Droste effect image transformation. Computers and Graphics, 31(3):516–523, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306002433.

Li:2022:GNN

[LF22] Zijie Li and Amir Barati Farimani. Graph neural network-accelerated Lagrangian fluid simulation. Computers and Graphics, 103(??):201–211, April 2022.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000206.

Liu:2002:VVL

[LFL02] Zhiyan Liu, Adam Finkelstein, and Kai Li. Visualization of very large datasets: Improving progressive view-dependent isosurface propagation. Computers and Graphics, 26(2):209– 218, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/28/abstract.
html.

Ladeveze:2010:IPP

[LFP10] N. Ladeveze, Jean-Yves Fourquet, and Bernard Puel. Interactive path planning for haptic assistance in assembly tasks. Computers and Graphics, 34(1):17– 25, February 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001356.

Liao:2021:ADE

[LFY^{+}21] Jie Liao, Yanping Fu, Qingan Yan, Fei Luo, and Chunxia Xiao. Adaptive depth estimation for pyramid multi-view stereo. Computers and Graphics, 97 (??):268–278, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000583.

Lopes:1989:CAP

[LG89] Pedro Faria Lopes and Mario Rui Gomes. Computer animation in Portugal. Computers and Graphics, 13(3):381–387, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liang:1994:JHI

[LG94] Jiandong Liang and Mark Green. JDCAD: a highly interactive 3D modeling system. Computers and Graphics, 18(4):499–506, July– August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2003:IRV

[LG03] Jianghong Liu and Haiming Gu. Image retrieval in various domains. Computers and Graphics, 27 (5):807–812, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2023:LSR

[LG23] Fuyan Liu and Junwen Gan. Light subpath reservoir for interactive raytraced global illumination.

Computers and Graphics, 111(??):37–46, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000055.

Liu:2016:SES

[LGLK16] Dongwei Liu, Haokun Geng, Tieying Liu, and Reinhard Klette. Star-effect simulation for photography. Computers and Graphics, 61 (??):19–28, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301091.

Lawonn:2014:ARC

[LGRP14] Kai Lawonn, Rocco Gasteiger, Christian Rössl, and Bernhard Preim. Adaptive and robust curve smoothing on surface meshes. Computers and Graphics, 40(??): 22–35, May 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000132.

Leimer:2017:RBP

[LGWM17] Kurt Leimer, Lukas Gersthofer, Michael Wimmer, and Przemysław Musialski. Relation-based parametrization and exploration of shape collections. Computers and Graphics, 67 (??):127–137, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300948.

Li:2021:BMD

[LGZ^{+}21] Menghang Li, Shanshan Gao, Chenhao Zhang, Minfeng Xu, and Caiming Zhang. Blind motion deblurring via L0 sparse representation. Computers and Graphics, 97(??):248– 257, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000662.

Loveluck:1983:ITR

[LH83] J. M. Loveluck and F. R. A. Hopgood. The ICL/Three Rivers PERQ and distributed interactive computing. Computers and Graphics, 7(2):193–198, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liao:1991:FGB

[LH91] Chia-Wei Liao and Jun S. Huang. Font generation by beta-spline curve. Computers and Graphics, 15(4): 527–534, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Li:2000:FRT

[LH00] C. L. Li and K. C. Hui.

Feature recognition by template matching. Computers and Graphics, 24(4):569– 582, August 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/34/article.
pdf.

Li:2014:ESJ

[LH14] Yong Li and Zhangjin Huang. Efficient schemes for joint isotropic and anisotropic total variation minimization for deblurring images corrupted by impulsive noise. Computers and Graphics, 38(??):108– 116, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300160X.

Lynch:2012:CEC

[LHC12] Sean Lynch, Jonathan Haber, and Sheelagh Carpendale. ColourVis: Exploring colour in digital images. Computers and Graphics, 36 (6):696–707, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200026X.

Luecke:1997:FIS

[LhCE97] Greg R. Luecke, Young ho Chai, and James C. Edwards. Force interactions in the synthetic environment using the ISU force reflecting exoskeleton. Computers and Graphics, 21(4): 431–442, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700020.

Liao:2005:FRD

[LHCL05] Horng-Shyang Liao, Tan-Chi Ho, Jung-Hong Chuang, and Cheng-Chung Lin. Fast rendering of dynamic clouds. Computers and Graphics, 29 (1):29–40, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Li:2021:RPD

[LHG21] Teng Li, Shijie Hao, and Yanrong Guo. Robust pencil drawing generation via fast Retinex decomposition. Computers and Graphics, 97(??):67–77, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000509.

Lin:2021:SID

[LHH^{+}21] Xiao Lin, Qi Huang, Wei Huang, Xin Tan, Meie Fang, and Lizhuang Ma. Single image deraining via detailguided efficient channel attention network. Computers and Graphics, 97(??): 117–125, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100056X.

Liu:2023:SGT

[LHL23] Chun Liu, Jingsong Hu, and Hong Lin. SWF-GAN: a textto-image model based on sentence-word fusion perception. Computers and Graphics, 115(??):500–510, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300167X.

Lastra:1987:URD

[LHS87] Gerardo Leon Lastra, Gabriel Ruiz Huerta, and Julio S. Santana Sepulveda. The universal ruler for 2D drafting systems. Computers and Graphics, 11(1):27–31, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2023:LML

[LHY23] Houxuan Liu, Xiao Han, and Lu Yang. LS-MVSNet:

Lightweight self-supervised multi-view stereo. Computers and Graphics, 117(??): 183–191, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002704.

Li:1989:CPG

[Li89] Wentian Li. Complex patterns generated by next nearest neighbors cellular automata. Computers and Graphics, 13(4):531–537, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Libes:1991:MDS

[Lib91] Don Libes. Modeling dynamic surfaces with octrees. Computers and Graphics, 15 (3):383–387, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liewald:1985:IGE

[Lie85]  M. H. Liewald.  Initial

Graphics Exchange Specification: Successes and evolution. Computers and Graphics, 9(1):47–50, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lindner:1979:PRT

[Lin79] Rolf Lindner. A processor for real time TV raster scan conversion. Computers and Graphics, 4(1):23–28, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Linhart:1990:QPP

[Lin90] Johann Linhart. A quick point-in-polyhedron test. Computers and Graphics, 14 (3–4):445–447, 1990. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ling:1997:ASM

[Lin97]  Li Ling.  Animation of

stochastic motion of 3-D cloth objects. Computers and Graphics, 21 (6):769–775, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700056.

Lindeman:2010:FSS

[Lin10] Robert Lindeman. Foreword to special section on IEEE VR 2009. Computers and Graphics, 34(1):1– 2, February 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900154X.

Lin:2015:BHP

[Lin15] Wei-Chin Lin. Boundary handling and porous flow for fluid-hair interactions. Computers and Graphics, 52 (??):33–42, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000904.

Lister:1995:GEI

[Lis95] Paul Lister. Guest Editor’s introduction. Computers and Graphics, 19(2): 237–??, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2021:FSS

[Liu21] Ligang Liu. Foreword to the special section on Chinagraph 2020. Computers and Graphics, 96(??): A5–A6, May 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000303.

Livesu:2018:HFB

[Liv18]  Marco Livesu.  A heat

flow based relaxation scheme for n dimensional discrete hyper surfaces. Computers and Graphics, 71(??): 124–131, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300049.

Livesu:2023:TRP

[Liv23] Marco Livesu. Towards a robust and portable pipeline for quad meshing: Topological initialization of injective integer grid maps. Computers and Graphics, 112(??):50–59, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000341.

Li:2004:ESS

[LJCW04] Xueyi Li, Hong Jiang, Song Chen, and Xiaochun Wang. An efficient surface–surface intersection algorithm based on geometry characteristics. Computers and Graphics, 28 (4):527–537, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lu:2018:PCQ

[LJH18] Lizheng Lu, Chengkai Jiang, and Qianqian Hu. Planar cubic G^{1} and quintic G^{2} Hermite interpolations via curvature variation minimization. Computers and Graphics, 70 (??):92–98, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301000.

Lopez:2022:GHP

[LJJPF22] Alfonso López, Juan M. Jurado, J. Roberto Jiménez-Pérez, and Francisco R. Feito. Generation of hyperspectral point clouds: Mapping, compression and rendering. Computers and Graphics, 106(??):267–276, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001145.

Liu:2007:EBA

[LJWcH07] Shengjun Liu, Xiaogang Jin, Charlie C. L. Wang, and Kin chuen Hui. Ellipsoidalblob approximation of 3D models and its applications. Computers and Graphics, 31(2):243–251, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000325.

Lee:2000:SSP

[LK00] Jin-Aeon Lee and Lee-Sup Kim. SPARP: a single pass antialiased rasterization processor. Computers and Graphics, 24(2): 233–243, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/31/article.
pdf.

Lee:2018:HMR

[LK18] Wonseop Lee and Hyeong-Seok Ko. Heuristic misfit reduction: A programmable approach for 3D garment fit customization. Computers and Graphics, 71(??): 1–13, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931730170X.

Liu:1994:IAP

[LKC94] Wen-Bang Liu, Ming-Tat Ko, and Ruei-Chuan Chang. An interactive approach to planning snake motion. Computers and Graphics, 18(4):537–542, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:1998:SSC

[LKC98] Jen-Duo Liu, Ming-Tat Ko, and Ruei-Chuan Chang. A simple self-collision avoidance for cloth animation. Computers and Graphics, 22 (1):117–128, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/

                                        cas/tree/store/cag/sub/
                                        1998/22/1/518.pdf.

Li:2019:ETC

[LKHM19] Xiang Li, Adarsh Krishnamurthy, Iddo Hanniel, and Sara McMains. Edge topology construction of Voronoi diagrams of spheres in nongeneral position. Computers and Graphics, 82 (??):332–342, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301013.

Lee:2002:MPV

[LKL02a] Yunjin Lee, Hyoung Seok Kim, and Seungyong Lee. Mesh parameterization with a virtual boundary. Computers and Graphics, 26(5): 677–686, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/56/31/abstract.
                                        html.

Li:2002:TCS

[LKL^{+}02b] Yong-Qing Li, Ying-Lin Ke, Wei-Shi Li, Qun-Sheng Peng, and Jian-Rong Tan. Termination criterion for subdivision of triangular Bézier patch. Computers and Graphics, 26(1):67–74, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/34/abstract.
html.

Lengauer:2020:SAR

[LKL^{+}20] Stefan Lengauer, Alexander Komar, Arniel Labrada, Stephan Karl, Elisabeth Trinkl, Reinhold Preiner, Benjamin Bustos, and Tobias Schreck. A sketch-aided retrieval approach for incomplete 3D objects. Computers and Graphics, 87 (??):111–122, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300182.

Li:2016:MSC

[LKLW16] Zongmin Li, Zhenzhong Kuang, Yujie Liu, and Jiayan Wang. Multiscale shape context and re-ranking for deformable shape retrieval. Computers and Graphics, 54 (??):8–17, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001053.

Lee:2007:EGB

[LKS07] WeeSan Lee, Levent Burak Kara, and Thomas F. Stahovich. An efficient graphbased recognizer for handdrawn symbols. Computers and Graphics, 31(4):554– 567, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001008.

Li:1991:SCR

[LL91] Hua Li and Shen-Quan Liu. Shape controls in rational Beta-splines. Computers and Graphics, 15(1):25–28, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lei:1992:PAF

[LL92] Chin-Laung Lei and Horng-Twu Liaw. A parallel algorithm for finding congruent regions. Computers and Graphics, 16(3):289– 294, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Li:2000:SHW

[LL00] Ling Li and Xiaoyan Liu. Simulating human walking on special terrain: up and down slopes. Computers and Graphics, 24(3): 453–463, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/30/40/article.
                                        pdf.

Les:2004:UCP

[LL04] Z. Les and M. Les. Understanding the curve-polygon object. Computers and Graphics, 28(6):919–936, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Les:2005:UCP

[LL05] Zbigniew Les and Magdalena Les. Understanding of the concave polygon object in the shape understanding system. Computers and Graphics, 29(3):365– 378, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Les:2006:COE

[LL06] Zbigniew Les and Magdalena Les. The cyclic object: The example of visual reasoning in the shape understanding system. Computers and Graphics, 30 (5):787–799, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001282.

Li:2011:SDA

[LL11] Jituo Li and Guodong Lu.

Skeleton driven animation based on implicit skinning. Computers and Graphics, 35 (5):945–954, October 2011.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001348.

Lou:2012:CIU

[LL12] Qi Lou and Ligang Liu. Curve intersection using hybrid clipping. Computers and Graphics, 36(5):309– 320, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200060X.

Liao:2004:TCR

[LLC04] Shih-Kuan Liao, Jim Z. C. La, and Yeh-Ching Chung. Time-critical rendering for time-varying volume data. Computers and Graphics, 28(2):279–288, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ling:2009:TBS

[LLC09] Ruotian Ling, Xiaonan Luo, and Zhongxian Chen. Ternary butterfly subdivision. Computers and Graphics, 33(4):566–575, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000648.

Lugo:2022:LHR

[LLC^{+}22] Gabriel Lugo, Ryan Li, Rutvik Chauhan, Zihao Wang, Palak Tiwary, Utkarsh Pandey, Archi Patel, Steve Rombough, Rod Schatz, and Irene Cheng. LiSurveying: a high-resolution TLS-LiDAR benchmark. Computers and Graphics, 107(??):116– 130, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001273.

Lu:2012:DP

[LLGA12] Yanyan Lu, Jyh-Ming Lien, Mukulika Ghosh, and Nancy M. Amato. α-decomposition of polygons. Computers and Graphics, 36(5):466– 476, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200057X.

Lawonn:2017:ISP

[LLH17] Kai Lawonn, Maria Luz, and Christian Hansen. Improving spatial perception of vascular models using supporting anchors and illustrative visualization. Computers and Graphics, 63(??): 37–49, April 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300171.

Lilley:1994:DDD

[LLHH94] C. C. Lilley, F. Lin, W. T. Hewitt, and T. L. J. Howard. Design and development of distance learning materials for graphics and visualisation. Computers and Graphics, 18(3):269– 275, May–June 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2015:PCO

[LLL^{+}15] Xiuping Liu, Shuhua Li, Risheng Liu, Jun Wang, Hui Wang, and Junjie Cao. Properly constrained orthonormal functional maps for intrinsic symmetries. Computers and Graphics, 46 (??):198–208, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001046.

Liu:2023:RAC

[LLL^{+}23] Xinqi Liu, Jituo Li, Guodong Lu, Dongliang Zhang, and Shihai Xing. Robust and automatic clothing reconstruction based on a single RGB image. Computers and Graphics, 110 (??):98–110, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002059.

Li:2011:RWG

[LLLC11] Li Li, Fei Liu, Congbo Li, and Guoan Chen. Realistic wrinkle generation for 3D face modeling based on automatically extracted curves and improved shape control functions. Computers and Graphics, 35(1):175– 184, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001287.

Li:2016:HMR

[LLLZ16] Meng Li, Howard Leung, Zhiguang Liu, and Liuyang Zhou. 3D human motion retrieval using graph kernels based on adaptive graph construction. Computers and Graphics, 54 (??):104–112, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001089.

Lempert:1990:PCH

[LLM^{+}90] H. Lempert, M. Lutz, L. A.

Messina, W. S. Ting, and C. Sanger. A prototype for configuring hardware and software. Computers and Graphics, 14(2): 303–310, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lin:2016:CAS

[LLM^{+}16] Shuai Lin, Yu-Kun Lai, Ralph R. Martin, Shiyao Jin, and Zhi-Quan Cheng. Coloraware surface registration. Computers and Graphics, 58(??):31–42, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300504.

Liu:2021:MPP

[LLP^{+}21] Peiqing Liu, An Liu, Hao Peng, Lihao Tian, Jikai Liu, and Lin Lu. Mechanical property profiles of microstructures via asymptotic homogenization. Computers and Graphics, 100(??): 106–115, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100162X.

Li:2022:PFG

[LLP^{+}22] Guangyao Li, Liangfu Li, Yingdan Pu, Nan Wang, and Xi Zhang. Progressive feature generation with mask awareness for image inpainting. Computers and Graphics, 107(??):197– 207, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001467.

Lang:1993:SVS

[LLR93] Ulrich Lang, Ruth Lang, and Roland Rühle. Scientific visualization in a supercomputer network at RUS. Computers and Graphics, 17(1):15–22, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2016:SCB

[LLS^{+}16] Xianyong Liu, Ligang Liu, Weijie Song, Yanping Liu, and Lizhuang Ma. Shape context based mesh saliency detection and its applications: a survey. Computers and Graphics, 57(??): 12–30, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300140.

Liu:2019:DSD

[LLW^{+}19] Yuxin Liu, Guiqing Li, Yupan Wang, Yongwei Nie, and Aihua Mao. Discrete shell deformation driven by adaptive sparse localized components. Computers and Graphics, 78 (??):76–86, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301833.

Liang:2023:MUU

[LLW^{+}23] Hui Liang, Shiqing Liu, Yi Wang, Junjun Pan, Yazhou Zhang, and Xiaohang Dong. Multi-user upper limb rehabilitation training system integrating social interaction. Computers and Graphics, 111(??): 103–110, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000079.

Lu:2015:ECE

[LLX^{+}15] Huina Lu, Guiqing Li, Chuhua Xian, Zhibang Zhang, and Mengxiao Yin. EC–CageR: Error controllable cage reverse for animated meshes. Computers and Graphics, 46(??): 138–148, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000946.

Li:2011:MQG

[LLZ^{+}11] Er Li, Bruno Lévy, Xiaopeng Zhang, Wujun Che, Weiming Dong, and Jean-Claude Paul. Meshless quadrangulation by global parameterization. Computers and Graphics, 35(5): 992–1000, October 2011.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931100121X.

Lu:2023:MDM

[LLZ^{+}23] Shitao Lu, Shice Liu, Keyue Zhang, Mingang Chen, Xin Tan, and Lizhuang Ma. Multi-domain mixup for scenario-universal face antispoofing. Computers and Graphics, 116(??):327–335, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001760.

Lakhtakia:1989:SSS

[LM89] Akhlesh Lakhtakia and Russell Messier. Self-similar sequences and chaos from Gauss sums. Computers and Graphics, 13(1):59–62, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2016:TAR

[LM16] Xianyong Liu and Lizhuang Ma. Topological analysis for 3D real, symmetric secondorder tensor fields using Deviatoric Eigenvalue Wheel. Computers and Graphics, 54 (??):28–37, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001120.

Leimer:2022:ARO

[LM22] Kurt Leimer and Przemyslaw Musialski. Analysis of a reduced-order model for the simulation of elastic geometric zigzag-spring meta-materials. Computers and Graphics, 102(??): 187–198, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002181.

Lee:2013:API

[LMC13] Yongjoon Lee, Jaehwan Ma, and Sunghee Choi. Automatic pose-independent 3D garment fitting. Computers and Graphics, 37(7): 911–922, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001234.

Leonhardt:1996:LSM

[LMD96] Ulf Leonhardt, Jeff Magee, and Paul Dias. Location service in mobile computing environments. Computers and Graphics, 20(5):627–632, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600036.

Lopez-Moreno:2010:CIT

[LMHRG10] Jorge Lopez-Moreno, Sunil Hadap, Erik Reinhard, and Diego Gutierrez. Compositing images through light source detection. Computers and Graphics, 34(6): 698–707, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001299.

Lopez-Moreno:2011:NPD

[LMJH^{+}11] Jorge Lopez-Moreno, Jorge Jimenez, Sunil Hadap, Ken Anjyo, Erik Reinhard, and Diego Gutierrez. Nonphotorealistic, depth-based image editing. Computers and Graphics, 35(1):99– 111, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001767.

Lechlek:2019:IHI

[LMR^{+}19] Loubna Lechlek, Daniel Meneveaux, Mickaël Ribardière, Romuald Perrot, and Mohamed Chaouki Babahenini. Interactive HDR image-based rendering from unstructured LDR photographs. Computers and Graphics, 84(??):1– 12, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301281.

Li:2006:SOA

[LMW06] Z. Li, D. S. Meek, and D. J. Walton. A smooth, obstacleavoiding curve. Computers and Graphics, 30(4):581– 587, August 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000756.

Li:2021:CJE

[LMY^{+}21] Yangke Li, Dongyang Ma, Yuhang Yu, Guangshun Wei, and Yuanfeng Zhou. Compact joints encoding for skeleton-based dynamic hand gesture recognition. Computers and Graphics, 97(??):191–199, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000595.

Landgraf:1990:SRD

[LMZ90] G. Landgraf, K. H. Modler, and M. Ziegenhorn. System to represent doublycurved surfaces. Computers and Graphics, 14(3–4): 435–443, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Laporte:1995:GSB

[LNFC95] H. Laporte, E. Nyiri, M. Froumentin, and C. Chaillou. A graphics system based on quadrics. Computers and Graphics, 19(2): 251–260, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400151.

Le:2023:TTB

[LNL^{+}23a] Trung-Nghia Le, Tam V.

Nguyen, Minh-Quan Le, Trong-Thuan Nguyen, Viet-Tham Huynh, Trong-Le Do, Khanh-Duy Le, Mai-Khiem Tran, Nhat Hoang-Xuan, Thang-Long Nguyen-Ho, Vinh-Tiep Nguyen, Tuong-Nghiem Diep, Khanh-Duy Ho, Xuan-Hieu Nguyen, Thien-Phuc Tran, Tuan-Anh Yang, Kim-Phat Tran, Nhu-Vinh Hoang, Minh-Quang Nguyen, E-Ro Nguyen, Minh-Khoi Nguyen-Nhat, Tuan-An To, Trung-Truc Huynh-Le, Nham-Tan Nguyen, Hoang-Chau Luong, Truong Hoai Phong, Nhat-Quynh Le-Pham, Huu-Phuc Pham, Trong-Vu Hoang, Quang-Binh Nguyen, Hai-Dang Nguyen, Akihiro Sugimoto, and Minh-Triet Tran. TextANIMAR:

Text-based 3D animal finegrained retrieval. Computers and Graphics, 116(??): 162–172, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001553.

Le:2023:SSB

[LNL^{+}23b] Trung-Nghia Le, Tam V. Nguyen, Minh-Quan Le, Trong-Thuan Nguyen, Viet-Tham Huynh, Trong-Le Do, Khanh-Duy Le, Mai-Khiem Tran, Nhat Hoang-Xuan, Thang-Long Nguyen-Ho, Vinh-Tiep Nguyen, Nhat-Quynh Le-Pham, Huu-Phuc Pham, Trong-Vu Hoang, Quang-Binh Nguyen, Trong-Hieu Nguyen-Mau, Tuan-Luc Huynh, Thanh-Danh Le, Ngoc-Linh Nguyen-Ha, Tuong-Vy Truong-Thuy, Truong Hoai Phong, Tuong-Nghiem Diep, Khanh-Duy Ho, Xuan-Hieu Nguyen, Thien-Phuc Tran, Tuan-Anh Yang, Kim-Phat Tran, Nhu-Vinh Hoang, Minh-Quang Nguyen, Hoai-Danh Vo, Minh-Hoa Doan, Hai-Dang Nguyen, Akihiro Sugimoto, and Minh-Triet Tran. SketchANIMAR: Sketch-based 3D animal fine-grained retrieval. Computers and Graphics, 116(??):150–161, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001644.

Lidal:2013:GS

[LNP^{+}13] Endre M. Lidal, Mattia Natali, Daniel Patel, Helwig Hauser, and Ivan Viola. Geological storytelling. Computers and Graphics, 37 (5):445–459, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000125.

Li:2016:RBM

[LNSW16] Minglei Li, Liangliang Nan, Neil Smith, and Peter Wonka. Reconstructing building mass models from UAV images. Computers and Graphics, 54(??):84– 93, February 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001077.

Lodha:2021:SNR

[Lod21]  Ishaan Lodha.  Subjec-

tive and no-reference quality metric of domain independent images and videos. Computers and Graphics, 95(??):123–129, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000133.

Lopes:2002:RAP

[LOdF02] Hélio Lopes, João Batista Oliveira, and Luiz Henrique de Figueiredo. Robust adaptive polygonal approximation of implicit curves. Computers and Graphics, 26(6): 841–852, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lohmann:1995:AST

[Loh95] Gabriele Lohmann. Analysis and synthesis of textures: a co-occurrence-based approach. Computers and Graphics, 19(1):29– 36, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1&aid=9400119.

Lopes:1992:DRP

[Lop92] Artur O. Lopes. On the dynamics of real polynomials on the plane. Computers and Graphics, 16(1):15–23, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Loyless:1991:ARA

[Loy91] James E. Loyless. Autumn — a recipe for artistic fractal images. Computers and Graphics, 15(1):87–88, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lozover:1983:ACC

[LP83] O. Lozover and K. Preiss.

Automatic construction of a cubic B-spline representation for a general curve. Computers and Graphics, 7 (2):149–153, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lin:1992:SBA

[LP92] Feng Lin and Yunhe Pan.

A stack-based approach for shading of regions. Computers and Graphics, 16 (1):79–84, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lakhtakia:1993:CBP

[LP93] Akhlesh Lakhtakia and Dann E. Passoja. On the congruence of binary patterns generated by modular arithmetic on a parent array. Computers and Graphics, 17(5):613–617, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lindemeier:2013:ISP

[LPD13] Thomas Lindemeier, Sören Pirk, and Oliver Deussen. Image stylization with a painting machine using semantic hints. Computers and Graphics, 37(5):293– 301, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000071.

Lavoue:2018:FSS

[LPD^{+}18] Guillaume Lavoué, Ioannis Pratikakis, Florent Dupont, Maks Ovsjanikov, and Michela Spagnuolo. Foreword to the special section on Eurographics Workshop on 3D Object Retrieval 2017. Computers and Graphics, 71(??): A6–A7, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931730198X.

Lu:2005:VLE

[LPL^{+}05] Jianfeng Lu, Zhigeng Pan, Hai Lin, Mingmin Zhang, and Jiaoying Shi. Virtual learning environment for medical education based on VRML and VTK. Computers and Graphics, 29(2): 283–288, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Langenfeld:2020:PSM

[LPL^{+}20] Florent Langenfeld, Yuxu Peng, Yu-Kun Lai, Paul L. Rosin, Tunde Aderinwale, Genki Terashi, Charles Christoffer, Daisuke Kihara, Halim Benhabiles, Karim Hammoudi, Adnane Cabani, Feryal Windal, Mahmoud Melkemi, Andrea Giachetti, Stelios Mylonas, Apostolos Axenopoulos, Petros Daras, Ekpo Otu, Reyer Zwiggelaar, David Hunter, Yonghuai Liu, and Matthieu Montès. SHREC 2020: Multidomain protein shape retrieval challenge. Computers and Graphics, 91 (??):189–198, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301151.

Liu:2020:POG

[LPO20] Chang Liu, Alexander Plopski, and Jason Orlosky. OrthoGaze: Gaze-based three-dimensional object manipulation using orthogonal planes. Computers and Graphics, 89(??): 1–10, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300431.

Linhares:2019:SNO

[LPP^{+}19] Cláudio D. G. Linhares, Jean R. Ponciano, Fabíola S. F. Pereira, Luis E. C. Rocha, Jose Gustavo S. Paiva, and Bruno A. N. Travençolo. A scalable node ordering strategy based on community structure for enhanced temporal network visualization. Computers and Graphics, 84(??):185– 198, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301347.

Lopes:2007:VTM

[LPPM07] A. A. Lopes, R. Pinho, F. V. Paulovich, and R. Minghim. Visual text mining using association rules. Computers and Graphics, 31(3): 316–326, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000544.

Lieng:2014:CSS

[LPR^{+}14] Henrik Lieng, Tania Pouli, Erik Reinhard, Jirí Kosinka, and Neil A. Dodgson. Cornsweet surfaces for selective contrast enhancement. Computers and Graphics, 42(??):1–13, August 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000405.

Livesu:2019:FSS

[LPS19] Marco Livesu, Gianni Pintore, and Alberto Signoroni. Foreword to the special section on Smart Tools and Applications in Computer Graphics (STAG 2018). Computers and Graphics, 82 (??):A10–A11, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300950.

Luo:1992:ARS

[LPV92] Yuhua Luo, Francisco J.

Perales Lopez, and Juan J. Villanueva Pipaon. An automatic rotoscopy system for human motion based on a biomechanic graphical model. Computers and Graphics, 16(4):355– 362, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lennings:1995:EIA

[LPV95] A. F. Lennings, J. C. Peters, and J. S. M. Vergeest. An efficient integration of algorithms to evaluate the quality of freeform surfaces. Computers and Graphics, 19(6):861–872, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500070.

Liu:2021:SIT

[LPZ^{+}21] Rui Liu, Chao Peng, Yunbo Zhang, Hannah Husarek, and Qi Yu. A survey of immersive technologies and applications for industrial product development. Computers and Graphics, 100(??):137–151, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001643.

Li:2012:CAT

[LQ12] Bo Li and Hong Qin. Component-aware tensorproduct trivariate splines of arbitrary topology. Computers and Graphics, 36 (5):329–340, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000465.

Li:2015:CMS

[LQ15] Kang Li and Xiaoping Qian. Covariance matrix of a shape population: a tale on spline setting. Computers and Graphics, 47(??): 89–104, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000023.

Liu:2008:DCD

[LQOW08] Wanli Liu, Xinghua Qu, Jianfei Ouyang, and Zhankui Wang. Design and CADdirected inspection planning of laser-guided measuring robot. Computers and Graphics, 32(6): 617–623, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001209.

Ludwig:1990:GMS

[LR90] M. Ludwig and Ch. Richter.

The 3D geometry modelling system (GEMO). Computers and Graphics, 14(3–4): 389–394, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2016:DFS

[LR16] Xin Liu and Jon G. Rokne.

Depth of field synthesis from sparse views. Computers and Graphics, 55(??): 21–32, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001831.

Laycock:2007:AGT

[LRD07] R. G. Laycock, G. D. G.

Ryder, and A. M. Day. Automatic generation, texturing and population of a reflective real-time urban environment. Computers and Graphics, 31 (4):625–635, August 2007.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000945.

Loaiza:2011:MCC

[LRG11] Manuel E. Loaiza, Alberto B. Raposo, and Marcelo Gattass. Multicamera calibration based on an invariant pattern. Computers and Graphics, 35(2): 198–207, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310002050.

Luboschik:2014:SIV

[LRHS14] Martin Luboschik, Stefan Rybacki, Fiete Haack, and Hans-Jörg Schulz. Supporting the integrated visual analysis of input parameters and simulation trajectories. Computers and Graphics, 39(??):37–47, April 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001416.

Lashkari:1992:PPI

[LRMS92] Y. Z. Lashkari, S. S. S. P. Rao, V. Y. Mhaskar, and A. C. Shelat. PLX: a proposal to implement a general broadcasting facility in a distributed environment running X windows.

Computers and Graphics, 16 (2):143–149, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Loganathan:1987:IGG

[LRR87] K. Loganathan, N. V. Raman, and A. Rajaraman. Interactive graphics for guyed tower analysis. Computers and Graphics, 11(3): 297–303, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lopes:2003:ESI

[LRS^{+}03] Hélio Lopes, Jarek Rossignac, Alla Safonova, Andrzej Szymczak, and Geovan Tavares. Edgebreaker: a simple implementation for surfaces with handles. Computers and Graphics, 27 (4):553–567, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2024:FRS

[LRY^{+}24] Yang Liu, Teng Ran, Liang Yuan, Kai Lv, and Guoquan Zheng. 3D face reconstruction from a single image based on hybrid-level contextual information with weak supervision. Computers and Graphics, 118 (??):80–89, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002881.

Lawrence:1979:CAD

[LS79] D. J. Lawrence and D. J. Sheppard. Computer aided design education — the more the education, the less the use. Computers and Graphics, 4(3–4):189–192, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lester:1984:CGH

[LS84] Lewis N. Lester and John Sandor. Computer graphics on a hexagonal grid. Computers and Graphics, 8 (4):401–409, ???? 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849384900384.

Lee:1990:APB

[LS90] Sukhan Lee and Yeong Gil Shin. Assembly planning based on geometric reasoning. Computers and Graphics, 14(2):237–250, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lin:2005:FNI

[LS05] Shengyou Lin and Jiaoying Shi. Fast natural image matting in perceptual color space. Computers and Graphics, 29(3):403– 414, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:2006:FBS

[LS06] Qi Liu and Alexei Sourin.

Function-based shape modelling extension of the Virtual Reality Modelling Language. Computers and Graphics, 30(4):629–645, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600080X.

Levinski:2007:IFB

[LS07] Konstantin Levinski and Alexei Sourin. Interactive function-based shape modelling. Computers and Graphics, 31(1):66– 76, January 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001804.

Lu:2008:SHR

[LS08] Xiong Lu and Aiguo Song.

Stable haptic rendering with detailed energy-compensating control. Computers and Graphics, 32(5):561–567, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000757.

Levy:2018:NOT

[LS18] Bruno Lévy and Erica L. Schwindt. Notions of optimal transport theory and how to implement them on a computer. Computers and Graphics, 72(??): 135–148, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300098.

Lage:2019:FSS

[LS19] Marcos Lage and Filip Sadlo. Foreword to special section on SIBGRAPI 2019. Computers and Graphics, 85 (??):A3–A6, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301712.

Liu:2023:PBU

[LSCJ23] Hao Liu, Yiran Sun, Xiangyun Cheng, and Dong Jiang. Prior-based 3D U-Net: a model for kneecartilage segmentation in MRI images. Computers and Graphics, 115(??):167– 180, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001413.

Liao:2018:DAD

[LSE18] Jingtang Liao, Shuheng Shen, and Elmar Eisemann. Depth annotations: Designing depth of a single image for depth-based effects. Computers and Graphics, 71(??):180–188, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301899.

Lopez-Sastre:2013:ESP

[LSGFRC^{+}13] R. J. López-Sastre, A. García-Fuertes, C. Redondo-Cabrera, F. J. Acevedo-Rodríguez, and S. Maldonado-Bascón. Evaluating 3D spatial pyramids for classifying 3D shapes. Computers and Graphics, 37(5):473–483, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000551.

Li:2012:IIV

[LSH^{+}12] Ping Li, Hanqiu Sun, Chen Huang, Jianbing Shen, and Yongwei Nie. Interactive image/video retexturing using GPU parallelism. Computers and Graphics, 36(8): 1048–1059, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001331.

Lichtenberg:2018:RTF

[LSHL18] Nils Lichtenberg, Noeska Smit, Christian Hansen, and Kai Lawonn. Real-time field aligned stripe patterns. Computers and Graphics, 74 (??):137–149, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300591.

Li:2010:RNE

[LSK^{+}10] Bao Li, Ruwen Schnabel, Reinhard Klein, Zhiquan Cheng, Gang Dang, and Shiyao Jin. Robust normal estimation for point clouds with sharp features. Computers and Graphics, 34(2):94–106, April 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000021X.

Lumley:2022:FCA

[LSR22] Sam Lumley, Renee Sieber, and Robert Roth. A framework and comparative analysis of web-based climate change visualization tools. Computers and Graphics, 103(??):19–30, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002636.

Li:1997:TRM

[LSS97] Minglu Li, Yongqiang Sun, and Huanye Sheng. Temporal relations in multimedia systems. Computers and Graphics, 21(3): 315–320, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700008.

Lombardi:2021:CDA

[LSS21] Marco Lombardi, Mattia Savardi, and Alberto Signoroni. Cross-domain assessment of deep learningbased alignment solutions for real-time 3D reconstruction. Computers and Graphics, 99 (??):54–69, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001266.

Lin:1996:DVM

[LST96] Feng Lin, Seah Hock Soon, and Lee Yong Tsui. Deformable volumetric model and isosurface: Exploring a new approach for surface boundary construction. Computers and Graphics, 20(1):33–40, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1&aid=9500090.

Luo:2012:FAS

[LSW12] Chuanjiang Luo, Issam Safa, and Yusu Wang. Featureaware streamline generation of planar vector fields via topological methods. Computers and Graphics, 36 (6):754–766, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931200115X.

Liang:2015:GSC

[LSW15] Luming Liang, Andrzej Szymczak, and Mingqiang Wei. Geodesic spin contour for partial near-isometric matching. Computers and Graphics, 46(??):156–171, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001125.

Lang:2023:MSB

[LSW^{+}23] Matěj Lang, Clemens Strobel, Felix Weckesser, Danielle Langlois, Enkelejda Kasneci, Barbora Kozlíková, and Michael Krone. A multimodal smartwatch-based interaction concept for immersive environments. Computers and Graphics, 117 (??):85–95, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002479.

Liu:2013:MAC

[LSWL13] Li Liu, Zhuo Su, Ruomei Wang, and Xiaonan Luo. Material-aware cloth simulation via constrained geometric deformation. Computers and Graphics, 37(1–2): 21–32, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200163X.

Liu:2023:HFG

[LSWZ23] Fukang Liu, Mingwen Shao, Fan Wang, and Lixu Zhang. High-fidelity GAN inversion by frequency domain guidance. Computers and Graphics, 114(??):286–295, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001176.

Lee:2011:DTT

[LSY11] Hochang Lee, Sanghyun Seo, and Kyunghyun Yoon. Directional texture transfer with edge enhancement. Computers and Graphics, 35 (1):81–91, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001779.

Li:2021:FTL

[LSZQ21] Jiachen Li, Xiuqiang Song, Fan Zhong, and Xueying Qin. Fast 3D texture-less object tracking with geometric contour and local region. Computers and Graphics, 97(??):225–235, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000546.

Limaiem:1995:GAI

[LT95] Anis Limaiem and François Trochu. Geometric algorithms for the intersection of curves and surfaces. Computers and Graphics, 19(3): 391–403, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500009.

Liu:2013:NEM

[LTBZ13] Zhenbao Liu, Sicong Tang, Shuhui Bu, and Hao Zhang. New evaluation metrics for mesh segmentation. Computers and Graphics, 37 (6):553–564, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000952.

Li:2020:PAN

[LTC^{+}20] Chen Li, Yusong Tan, Wei Chen, Xin Luo, Yulin He, Yuanming Gao, and Fei Li. ANU-Net: Attention-based nested U-Net to exploit full resolution features for medical image segmentation. Computers and Graphics, 90(??):11–20, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300546.

Lin:2019:EBF

[LTH^{+}19] Tsung-Ying Lin, Yu-Ting Tsai, Tsung-Shian Huang, Wen-Chieh Lin, and Jung-Hong Chuang. Exemplarbased freckle retouching and skin tone adjustment. Computers and Graphics, 78 (??):54–63, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301808.

Luvizon:2019:HPR

[LTP19] Diogo C. Luvizon, Hedi Tabia, and David Picard. Human pose regression by combining indirect part detection and contextual information. Computers and Graphics, 85(??):15–22, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301475.

Lopez:1996:VOS

[LTPN96] J. López, D. Tost, A. Puig, and I. Navazo. VolDmi: An open system for volume modeling and visualization. Computers and Graphics, 20(5):703–712, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600044.

Liu:2014:PCE

[LTR^{+}14] Baoquan Liu, Alexandru C. Telea, Jos B. T. M. Roerdink, Gordon J. Clapworthy, David Williams, Po Yang, Feng Dong, Valeriu Codreanu, and Alessandro Chiarini. Parallel centerline extraction on the GPU. Computers and Graphics, 41(??):72–83, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000272.

Li:1988:ISO

[LTS88] Xinyou Li, Zesheng Tang, and Jiaguang Sun. The implementation of set operation for regularized geometric object. Computers and Graphics, 12(3–4): 309–318, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lee:1996:MMS

[LTS96] Y. T. Lee, S. B. Tor, and E. L. Soo. Mathematical modelling and simulation of pop-up books. Computers and Graphics, 20(1):21– 31, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500089.

Leon:2008:VOI

[LTV08] Alejandro León, Juan Carlos Torres, and Francisco Velasco. Volume octree with an implicitly defined dual grid. Computers and Graphics, 32 (4):393–401, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000538.

Lovset:2013:RBM

[LUB^{+}13] Tyge Løvset, Dag Magne Ulvang, Tor Christian Bekkvik, Kåre Villanger, and Ivan Viola. Rule-based method for automatic scaffold assembly from 3D building models. Computers and Graphics, 37(4):256–268, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000095.

Luciani:2004:PSB

[LUMC04] Annie Luciani, Daniela Urma, Sylvain Marlière, and Joël Chevrier. PRESENCE: the sense of believability of inaccessible worlds. Computers and Graphics, 28 (4):509–517, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lesage:2002:TST

[LV02] P.-L. Lesage and M. Visvalingam. Technical section: Towards sketch-based exploration of terrain. Computers and Graphics, 26(2):309– 328, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/37/49/show/
index.htt;   http://www.
elsevier.com/gej-ng/10/
13/20/68/41/37/abstract.
html.

Lagae:2010:PIS

[LVLD10] Ares Lagae, Peter Vangorp, Toon Lenaerts, and Philip Dutré. Procedural isotropic stochastic textures by example. Computers and Graphics, 34 (4):312–321, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000713.

Lewiner:2011:ICH

[LVM^{+}11] Thomas Lewiner, Thales Vieira, Dimas Martínez, Adelailson Peixoto, Vinícius Mello, and Luiz Velho. Interactive 3D caricature from harmonic exaggeration. Computers and Graphics, 35(3):586–595, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000446.

Luyten:2006:DDU

[LVVC06] Kris Luyten, Jan Van den Bergh, Chris Vandervelpen, and Karin Coninx. Designing distributed user interfaces for ambient intelligent environments using models and simulations. Computers and Graphics, 30 (5):702–713, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600121X.

Liu:1988:AGP

[LW88] A. C. Liu and K. S. Wong. Automatic generation of process flow diagrams. Computers and Graphics, 12(3– 4):525–539, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lethbridge:1989:SHB

[LW89] Timothy C. Lethbridge and Colin Ware. A simple heuristically-based method for expressive stimulusresponse animation. Computers and Graphics, 13(3): 297–303, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Layman:1990:LMI

[LW90] John W. Layman and Tad E. Womack. Linear Markov iterated function systems. Computers and Graphics, 14 (2):343–353, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Liu:1999:TDS

[LW99] Li-Gang Liu and Guo-Jin Wang. Three-dimensional shape blending: intrinsic solutions to spatial interpolation problems. Computers and Graphics, 23 (4):535–545, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/34/article.
pdf.

Liu:2010:OUP

[LW10] Shengjun Liu and Charlie C. L. Wang. Orienting unorganized points for surface reconstruction. Computers and Graphics, 34(3): 209–218, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000403.

Leonard:2024:IBR

[LW24] Ludwic Leonard and Rüdiger Westermann. Image-based reconstruction of heterogeneous media in the presence of multiple light-scattering. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000049.

Liu:2018:RIO

[LWD^{+}18] Fuchang Liu, Shuangjian Wang, Dandan Ding, Qingshu Yuan, Zhengwei Yao, Zhigeng Pan, and Haisheng Li. Retrieving indoor objects: 2D–3D alignment using single image and interactive ROI-based refinement.

Computers and Graphics, 70 (??):108–117, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931730122X.

Liu:2023:GFN

[LWFZ23] Dezhi Liu, Weibing Wan, Zhijun Fang, and Xiuyuan Zheng. GsNeRF: Fast novel view synthesis of dynamic radiance fields. Computers and Graphics, 116(??): 491–499, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300239X.

Liu:2023:FAS

[LWG^{+}23] Yong-Jin Liu, Baodong Wang, Lin Gao, Junli Zhao, Ran Yi, Minjing Yu, Zhenkuan Pan, and Xianfeng Gu. 4D facial analysis: a survey of datasets, algorithms and applications. Computers and Graphics, 115(??):423– 445, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001437.

Li:2022:ISI

[LWJ^{+}22] Shujie Li, Lei Wang, Wei Jia, Yang Zhao, and Liping Zheng. An iterative solution for improving the generalization ability of unsupervised skeleton motion retargeting. Computers and Graphics, 104(??):129–139, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200053X.

Li:2011:SOQ

[LWLT11] Yufei Li, Wenping Wang, Ruotian Ling, and Changhe Tu. Shape optimization of quad mesh elements. Computers and Graphics, 35(3): 444–451, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000768.

Lu:2002:BPC

[LWP02] Guodong Lu, Xuanhui Wu, and Qunsheng Peng. Best papers of CAD & CG 2001: An efficient line clipping algorithm based on adaptive line rejection. Computers and Graphics, 26(3):409– 415, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/30/abstract.
html.

Li:2015:EFF

[LWS15] Yifei Li, Chaoli Wang, and Ching-Kuang Shene. Extracting flow features via supervised streamline segmentation. Computers and Graphics, 52(??):79– 92, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000886.

Li:2007:PPT

[LWW07] Jing Li, Wencheng Wang, and Enhua Wu. Pointin-polygon tests by convex decomposition. Computers and Graphics, 31 (4):636–648, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000921.

Li:2008:NAC

[LWW08] Sheng Li, Guoping Wang, and Enhua Wu. A new approach for construction and rendering of dynamic light shaft. Computers and Graphics, 32(6):660– 668, December 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001210.

Lipp:2010:PGM

[LWW10] Markus Lipp, Peter Wonka, and Michael Wimmer. Parallel generation of multiple L-systems. Computers and Graphics, 34(5):585– 593, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000828.

Li:2020:FIS

[LWW^{+}20] Yuerong Li, Xingce Wang, Zhongke Wu, Guoshuai Li, Shaolong Liu, and Mingquan Zhou. Flexible indoor scene synthesis based on multiobject particle swarm intelligence optimization and user intentions with 3D gesture. Computers and Graphics, 93 (??):1–12, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301205.

Li:2022:PLF

[LWWY22] Qingyu Li, Xin Wang, Tian Wu, and Huijun Yang. Point-line feature fusion based field real-time RGB-D SLAM. Computers and Graphics, 107(??):10– 19, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001169.

Li:2023:DFS

[LWZ^{+}23a] Tan Li, Hong Wang, Bin Zhou, Ziyang Li, Zhouping Chen, Qin Lan, and Dongchuan Fan. Design and feasibility study of a HCPS framework-based VR alpine skiing decision-making training system. Computers and Graphics, 114(??):138– 149, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000961.

Liu:2023:PAV

[LWZ^{+}23b] Hongbo Liu, Ziliang Wu, Erqing Zhang, Zhaosong Huang, Mingliang Xu, Lechao Cheng, Minfeng Zhu, and Wei Chen. A privacy-aware visual query approach for location-based data. Computers and Graphics, 115 (??):263–273, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001590.

Luo:2008:LDS

[LX08] Zhongxuan Luo and Junxiao Xue. Layered deformation of solid model using conformal mapping. Computers and Graphics, 32(6): 695–703, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930800126X.

Liu:2015:ISA

[LXB^{+}15] Zhenbao Liu, Caili Xie, Shuhui Bu, Xiao Wang, Junwei Han, Hongwei Lin, and Hao Zhang. Indirect shape analysis for 3D shape retrieval. Computers and Graphics, 46(??): 110–116, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001265.

Liu:2018:DHI

[LXCW18] Hui Liu, Jiazhi Xia, Jianer Chen, and Jianxin Wang. Detection of hierarchical intrinsic symmetry structure in 3D models. Computers and Graphics, 70(??): 8–16, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301280.

Liu:2021:SAD

[LXJL21] Xin Liu, Chuhua Xian, Shuo Jin, and Guiqing Li. Surface attributes driven volume segmentation for 3D-printing. Computers and Graphics, 100(??):43– 53, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001618.

Liu:2006:SMA

[LXPP06] Yong Liu, Congfu Xu, Zhigeng Pan, and Yunhe Pan. Semantic modeling for ancient architecture of digital heritage. Computers and Graphics, 30(5):800– 814, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001294.

Lin:2023:IDB

[LXT^{+}23] Xiao Lin, Duojiu Xu, Peiwen Tan, Lizhuang Ma, and Zhi-Jie Wang. Image deraining based on dual-channel component decomposition. Computers and Graphics, 116 (??):93–101, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001802.

Li:2010:FAH

[LXW^{+}10] Xin Li, Huanhuan Xu, Shenghua Wan, Zhao Yin, and Wuyi Yu. Featurealigned harmonic volumetric mapping using MFS. Computers and Graphics, 34(3):242–251, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000415.

Li:2022:SWD

[LXY22] Jie Li, Dan Xu, and Shaowen Yao. Sliced Wasserstein distance for neural style transfer. Computers and Graphics, 102(??):89– 98, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002600.

Liu:2008:GOV

[LY08] Yong-Jin Liu and Matthew Ming-Fai Yuen. Geometryoptimized virtual human head and its applications. Computers and Graphics, 32 (6):624–631, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001179.

Lai:2015:BMO

[LY15] Lip M. Lai and Matthew M. F. Yuen. Blending of mesh objects to parametric surface. Computers and Graphics, 46(??): 283–293, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001186.

Liu:2015:DCS

[LYC^{+}15] Yin Liu, Xiaosong Yang, Yang Cao, Zhao Wang, Biaosong Chen, Jianjun Zhang, and Hongwu Zhang. Dehydration of core/shell fruits. Computers and Graphics, 47(??):68–77, April 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849314001411.

Li:2017:HCD

[LYL^{+}17] Shuai Li, Dehui Yan, Xiangyang Li, Aimin Hao, and Hong Qin. Hessianconstrained detail-preserving 3D implicit reconstruction from raw volumetric dataset. Computers and Graphics, 64 (??):3–13, May 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300109.

Lyman:1989:EM

[Lym89] Kenvin Lyman. An elegant merging. Computers and Graphics, 13(1): 127–128, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lu:2016:BSC

[LYS^{+}16] Yang Lu, Jun-Hai Yong, Kan-Le Shi, Hai-Chuan Song, and Tian-Yu Ye. 3D B-spline curve construction from orthogonal views with self-overlapping projection segments. Computers and Graphics, 54 (??):18–27, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001132.

Lin:2019:AGP

[LYS^{+}19] Shih-Syun Lin, Juo-Yu Yang, Huang-Sin Syu, Chao-Hung Lin, and Tun-Wen Pai. Automatic generation of puzzle tile maps for spatialtemporal data visualization. Computers and Graphics, 82(??):1–12, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930055X.

Li:2010:FGM

[LYW^{+}10] Jituo Li, Juntao Ye, Yangsheng Wang, Li Bai, and Guodong Lu. Fitting 3D garment models onto individual human models. Computers and Graphics, 34(6): 742–755, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001147.

Li:2011:STG

[LYW^{+}11] Xin Li, Zhao Yin, Li Wei, Shenghua Wan, Wei Yu, and Maoqing Li. Symmetry and template guided completion of damaged skulls. Computers and Graphics, 35 (4):885–893, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000240.

Li:2013:PSI

[LYW^{+}13] Guiqing Li, Liang Yang, Shihao Wu, Wenshuang Tan, Xinyu Chen, and Chuhua Xian. Planar shape interpolation using relative velocity fields. Computers and Graphics, 37(5):364– 375, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000344.

Liao:2018:EID

[LYX18] Bin Liao, Peng Yin, and Chunxia Xiao. Efficient image dehazing using boundary conditions and local contrast. Computers and Graphics, 70(??):242–250, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301097.

Liu:2019:IGB

[LYXY19] Congying Liu, Zexi Yang, Feng Xu, and Jun-Hai Yong. Image generation from bounding box-represented semantic labels. Computers and Graphics, 81(??): 32–40, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300329.

Liang:2015:PPN

[LYZ15] Yunfeng Liang, Hao Yang, and Hui Zhang. A perpixel noise detection approach for example-based photometric stereo. Computers and Graphics, 46 (??):327–335, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000995.

Liu:2022:NHH

[LYZ^{+}22] Chunxiao Liu, Shuangshuang Ye, Lideng Zhang, Haiyong Bao, Xun Wang, and Fanding Wu. Nonhomogeneous haze data synthesis based real-world image dehazing with enhancementand-restoration fused CNNs. Computers and Graphics, 106(??):45–57, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000802.

Lin:2024:IMF

[LYZ^{+}24] Yunhui Lin, Guoying Yang, Yuefeng Ze, Lekai Zhang, Baixi Xing, Xinya Liu, and Ruimin Lyu. The impact of motion features of hand-drawn lines on emotional expression: an experimental study. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000244.

Liashtchenko:1988:IGS

[LZ88] A. Liashtchenko and V. Zaitsev. Interactive graphic systems design on the GKSstandard and unified language processor basis. Computers and Graphics, 12(3– 4):425–432, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lamot:2003:FPT

[LŽ03] Marko Lamot and Borut Žalik. A fast polygon triangulation algorithm based on uniform plane subdivision. Computers and Graphics, 27(2):239–253, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lin:2011:EIF

[LZ11] Hongwei Lin and Zhiyu Zhang. An extended iterative format for the progressive-iteration approximation. Computers and Graphics, 35(5):967–975, October 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001324.

Liu:2014:AIP

[LZ14] Xiao-Dan Liu and Chang-Wen Zheng. Adaptive importance photon shooting technique. Computers and Graphics, 38(??):158– 166, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001684.

Long:2024:ISC

[LZ24] Jianwu Long and Chen Zhang. Image smoothing combining edge-consistency with region-piecewise flatting. Computers and Graphics, 118(??):90–101, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300290X.

Liu:2015:QPC

[LZC^{+}15] Xiuping Liu, Jie Zhang, Junjie Cao, Bo Li, and Ligang Liu. Quality point cloud normal estimation by guided least squares representation. Computers and Graphics, 51 (??):106–116, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000710.

Li:2023:EEB

[LZG^{+}23] Chunhui Li, Mingquan Zhou, Guohua Geng, Yifei Xie, Yuhe Zhang, and Yangyang Liu. EPCS: Endpoint-based part-aware curve skeleton extraction for low-quality point clouds. Computers and Graphics, 117(??):209–221, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323002601.

Li:2023:MFS

[LZKJ23] Lei Li, Tianfang Zhang, Zhongfeng Kang, and Xikun Jiang. Mask-FPAN: Semisupervised face parsing in the wild with de-occlusion and UV GAN. Computers and Graphics, 116(??): 185–193, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001735.

Liu:2015:LRM

[LZL^{+}15] Xiuping Liu, Jie Zhang, Risheng Liu, Bo Li, Jun Wang, and Junjie Cao. Lowrank 3D mesh segmentation and labeling with structure guiding. Computers and Graphics, 46(??):99– 109, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001071.

Liao:2019:IAE

[LZL^{+}19] Bin Liao, Yao Zhu, Chao Liang, Fei Luo, and Chunxia Xiao. Illumination animating and editing in a single picture using scene structure estimation. Computers and Graphics, 82(??): 53–64, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300627.

Luo:2024:RDS

[LZL^{+}24] Xiaozhong Luo, Han Zhong, Junjie Lu, Chen Meng, and

Xu Han. RepDehazeNet:

Dual subnets image dehazing network based on structural re-parameterization. Computers and Graphics, 118 (??):71–79, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002893.

Liu:2018:HQC

[LZLS18] Zhiguang Liu, Liuyang Zhou, Howard Leung, and Hubert P. H. Shum. Highquality compatible triangulations and their application in interactive animation. Computers and Graphics, 76 (??):60–72, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301018.

Li:2004:WMS

[LZP^{+}04] Li Li, David Zhang, Zhigeng Pan, Jiaoying Shi, Kun Zhou, and Kai Ye. Watermarking 3D mesh by spherical parameterization. Computers and Graphics, 28(6): 981–989, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lo:2022:SSS

[LZR22] Wei Hong Lo, Stefanie Zollmann, and Holger Regenbrecht. Stats on-site

— sports spectator experience through situated visualizations. Computers and Graphics, 102(??):99– 111, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100265X.

Lamboray:2003:IMS

[LZSG03] Edouard Lamboray, Aaron Zollinger, Oliver G. Staadt, and Markus Gross. Interactive multimedia streams in distributed applications. Computers and Graphics, 27 (5):735–745, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lin:2024:GIG

[LZT^{+}24] Dayong Lin, Chunhui Zhao, Qihang Tian, Yunfei Xu, Ruilin Wang, and Zonghua Qu. GMM-ICQ: a GMM vertex-optimization-based implicitly-connected quadrilateral format for 3D mesh storage. Computers and Graphics, 118(??):34–47, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002820.

Li:2021:MRT

[LZW^{+}21a] Shi Li, Chuankun Zheng, Rui Wang, Yuchi Huo, Wenting Zheng, Hai Lin, and Hujun Bao. Multiresolution terrain rendering using summed-area tables. Computers and Graphics, 95(??):130–140, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000236.

Li:2021:HAG

[LZW^{+}21b] Zhaoxin Li, Xiaoge Zhang, Kangkan Wang, Hao Jiang, and Zhaoqi Wang. High accuracy and geometryconsistent confidence prediction network for multiview stereo. Computers and Graphics, 97(??):148– 159, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000625.

Li:2023:CDP

[LZY^{+}23] Mingyu Li, Lifeng Zhu, Yibing Yan, Ziyi Zhao, and Aiguo Song. Computational design of planet regolith sampler based on Bayesian optimization. Computers and Graphics, 116(??):464– 473, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002352.

Lin:2022:SAC

[LZYQ22] Yanna Lin, Wei Zeng, Yu Ye, and Huamin Qu. Saliencyaware color harmony models for outdoor signboard. Computers and Graphics, 105(??):25–35, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000644.

Li:2019:BRA

[LZZ^{+}19] Shujie Li, Yang Zhou, Haisheng Zhu, Wenjun Xie, Yang Zhao, and Xiaoping Liu. Bidirectional recurrent autoencoder for 3D skeleton motion data refinement. Computers and Graphics, 81(??):92–103, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300330.

Marhl:1994:CDR

[MA94] Marko Marhl and Varol Akman. On the correct determination of rotational angles for twisted-profiled sweep objects. Computers and Graphics, 18(5):691–694, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miksch:2014:MTA

[MA14] Silvia Miksch and Wolfgang Aigner. A matter of time: Applying a data-users-tasks design triangle to visual analytics of time-oriented data. Computers and Graphics, 38 (??):286–290, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001817.

Macedo:2015:FPC

[MA15] Márcio C. F. Macedo and Antônio L. Apolinário, Jr. Focus plus context visualization based on volume clipping for markerless onpatient medical data visualization. Computers and Graphics, 53 (part B)(??): 196–209, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001521.

Mousas:2017:RTP

[MA17] Christos Mousas and Christos-Nikolaos Anagnostopoulos. Real-time performancedriven finger motion synthesis. Computers and Graphics, 65(??):1–11, June 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300304.

Macedo:2018:IAA

[MA18] Márcio Macedo and Antônio Apolinário. Improved antialiasing for Euclidean distance transform shadow mapping. Computers and Graphics, 71(??):166–179, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301905.

Mahdavi-Amiri:2015:SDE

[MAAS15] Ali Mahdavi-Amiri, Troy Alderson, and Faramarz Samavati. A survey of Digital Earth. Computers and Graphics, 53 (part B) (??):95–117, December 2015.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001399.

Machover:1977:BSI

[Mac77]  Carl Machover.  Back-

ground and source information about computer graphics. Computers and Graphics, 2(2):119–122, 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Machgeels:1978:INI

[Mac78] C. Machgeels. Iota: a noninteractive graphics system using the pseudo-machine concept. Computers and Graphics, 3(1):29–34, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

MacanAichinnigh:1985:CHL

[Mac85] M. Mac an Aichinnigh. The context: a high level structuring concept for GKS input. Computers and Graphics, 9(3):211–220, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Machover:1995:GEI

[Mac95] Carl Machover. Guest Editor’s introduction. Computers and Graphics, 19(4): 489–??, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marinov:2019:BCM

[MAC19] Martin Marinov, Marco Amagliani, and Peter Charrot. Boundary conforming mesh to T-NURCC surface conversion. Computers and Graphics, 82(??): 95–105, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300731.

Maddock:1994:PCW

[Mad94] S. C. Maddock. Personal CAL workbook. Computers and Graphics, 18(3): 299–304, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Maddock:1995:PCW

[Mad95]  S. C. Maddock.  Per-

sonal CAL workbooks: The next chapters. Computers and Graphics, 19(4): 631–635, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500041.

Madhu:2000:MSM

[Mad00] K. P. Madhu. Meditations on the sutras of modern physics.

Computers and Graphics, 24 (4):629–631, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/40/article.
pdf.

Martin:2019:ADC

[MAdS^{+}19] Domingo Martín, Germán Arroyo, Vicente del Sol, Celia Romo, and Tobias Isenberg. Analysis of drawing characteristics for reproducing traditional handmade stippling. Computers and Graphics, 80(??): 1–16, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300226.

Melero:2019:FCD

[MAF19] Francisco Javier Melero, Ángel Aguilera, and Francisco Ramón Feito. Fast collision detection between high resolution polygonal models. Computers and Graphics, 83 (??):97–106, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301244.

Magalhaes:2016:PFE

[MAFL16] Salles V. G. Magalhães, Marcus V. A. Andrade, W. Randolph Franklin, and Wenli Li. PinMesh — fast and exact 3D point location queries using a uniform grid. Computers and Graphics, 58(??): 1–11, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300607.

Maguire:1984:VTG

[Mag84] M. C. Maguire. Visual testing of GKS at the human interface. Computers and Graphics, 8(1):19–27, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Maguire:1985:RHF

[Mag85] M. C. Maguire. A review of human factors guidelines and techniques for the design of graphical humancomputer interfaces. Computers and Graphics, 9(3): 221–235, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marton:2012:NEM

[MAG^{+}12] Fabio Marton, Marco Agus, Enrico Gobbetti, Giovanni Pintore, and Marcos Balsa Rodriguez. Natural exploration of 3D massive models on large-scale light field displays using the FOX proximal navigation technique. Computers and Graphics, 36 (8):893–903, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001215.

Mankoff:2000:OTS

[MAH00] Jennifer Mankoff, Gregory D. Abowd, and Scott E. Hudson. OOPS: a toolkit supporting mediation techniques for resolving ambiguity in recognition-based interfaces. Computers and Graphics, 24(6):819– 834, December 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/27/article.
pdf.

Majewski:1998:TRV

[Maj98] Mirek Majewski. A tutorial on the realistic visualization of 3D Sierpinski fractals. Computers and Graphics, 22(1):129–142, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/cag/sub/1998/22/
                                        1/519.pdf.

Mallya:1987:CPC

[Mal87] Prabhakar Mallya. Computational problems in the construction of plane curves. Computers and Graphics, 11 (2):79–86, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Martin:2011:SDE

[MALI11] Domingo Martín, Germán Arroyo, M. Victoria Luzón, and Tobias Isenberg. Scaledependent and examplebased grayscale stippling. Computers and Graphics, 35 (1):160–174, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001755.

Maio:2024:PAR

[MAM^{+}24] Rafael Maio, Tiago Araújo, Bernardo Marques, André Santos, Pedro Ramalho, Duarte Almeida, Paulo Dias, and Beatriz Sousa Santos. Pervasive augmented reality to support real-time data monitoring in industrial scenarios: Shop floor visualization evaluation and user study. Computers and Graphics, 118(??):11– 22, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002686.

Manohar:1999:CGI

[Man99]  Swami Manohar.  Com-

puter graphics in India — advances in volume graphics. Computers and Graphics, 23(1):73–84, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/651.pdf.

Miranda:2012:SES

[MAO^{+}12] José Carlos Miranda, Xenxo Alvarez, João Orvalho, Diego Gutierrez, A. Augusto Sousa, and Verónica Orvalho. Sketch express: a sketching interface for facial animation. Computers and Graphics, 36(6):585– 595, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000416.

Marovac:1976:SND

[Mar76] N. Marovac. The structures of the network definition language NEDLAN. its use in defining networks in CAD using interactive computer graphics. Computers and Graphics, 2(1):23–29, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marcus:1978:CGC

[Mar78] Aaron Marcus. A computer graphics course for the two cultures. Computers and Graphics, 3(1):17–22, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1979:ICA

[Mar79] N. Marovac. Interactive computer aided 3D engineering and art design. Computers and Graphics, 4(2): 87–93, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1980:NOI

[Mar80] Nenad Marovac. A network oriented information structure: Networks, semantics and structures. Computers and Graphics, 5(2–4): 41–52, ???? 1980. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849380900060.

Marovac:1982:NOI

[Mar82] Nenad Marovac. A network oriented information structure: networks, semantics and structures. Computers and Graphics, 5(2– 4):41–52, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1986:AAO

[Mar86] Nenad Marovac. Architecture of application-oriented systems using interactive computer graphics. Computers and Graphics, 10(4): 371–377, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1987:HFE

[Mar87]  Nenad Marovac.  Han-

dling fonts in electronic publishing systems. Computers and Graphics, 11(3): 289–295, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1991:PDL

[Mar91] Nenad Marovac. Page description language INTER-PRESS in electronic publishing environment. Computers and Graphics, 15(3): 423–434, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Martyn:2001:ERT

[Mar01] Tomek Martyn. Efficient ray tracing affine IFS attractors. Computers and Graphics, 25(4):665–670, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/39/abstract.
html.

Martyn:2002:TSE

[Mar02a] Tomek Martyn. Technical section: An elementary proof of correctness of the Chaos Game for IFS and its hierarchical and recurrent generalizations. Computers and Graphics, 26(3):505– 510, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/38/abstract.
html.

Martyn:2002:TSA

[Mar02b] Tomek Martyn. Technical section: On approximation accuracy of the Chaos Game’s finite-time activity. Computers and Graphics, 26(5):753–764, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/38/abstract.
html.

Martyn:2003:TBB

[Mar03]  Tomek Martyn.  Tight

bounding ball for affine IFS attractor. Computers and Graphics, 27(4):535– 552, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Martyn:2004:NAM

[Mar04] Tomek Martyn. A new approach to morphing 2D affine IFS fractals. Computers and Graphics, 28(2): 249–272, April 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marcos:2007:ESI

[Mar07] Adérito Fernandes Marcos.

Editorial of the special issue “technology and digital art”. Computers and Graphics, 31 (6):785–787, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001525.

Martyn:2009:AWA

[Mar09]  Tomasz Martyn.  The

attractor-wrapping approach to approximating convex hulls of 2D affine IFS attractors. Computers and Graphics, 33(1):104–112, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000976.

Martyn:2010:RRI

[Mar10] Tomasz Martyn. Realistic rendering 3D IFS fractals in real-time with graphics accelerators. Computers and Graphics, 34(2): 167–175, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001150.

Martin:2017:SDS

[MARI17] Domingo Martín, Germán Arroyo, Alejandro Rodríguez, and Tobias Isenberg. A survey of digital stippling. Computers and Graphics, 67 (??):24–44, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300432.

Mason:1992:ACA

[Mas92] William K. Mason. Art from cellular automata and symmetrized dot-patterns. Computers and Graphics, 16 (4):439–441, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mahdavi-Amiri:2014:ACM

[MAS14] Ali Mahdavi-Amiri and Faramarz Samavati. Atlas of connectivity maps. Computers and Graphics, 39 (??):1–11, April 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001325.

McManus:1997:OSD

[MB97a] Donald McManus and Carl Beckmann. Optimal static 2-dimensional screen subdivision for parallel rasterization architectures. Computers and Graphics, 21(2): 159–169, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600079.

Melville:1997:PHG

[MB97b] John D. Melville and Robert P.

Burton. Piles for hyperdimensional graphics. Computers and Graphics, 21(1): 51–60, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600069.

Miyai:2010:FSS

[MB10] Ayumi Miyai and Gladimir V. G. Baranoski. Foreword to the special section on education. Computers and Graphics, 34(6): 779, December 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001500.

Merillou:2014:FSS

[MB14] Stéphane Merillou and Carles Bosch. Foreword to the special section on aging and weathering. Computers and Graphics, 45 (??):A1, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000831.

Mathew:2020:ODA

[MBA20] C. D. Tharindu Mathew, Bedrich Benes, and Daniel G. Aliaga. An output-driven approach to design a swarming model for architectural indoor environments. Computers and Graphics, 87 (??):103–110, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300200.

Morais:2023:EIE

[MBC^{+}23] Lucas Zanusso Morais, Victor Kunde Bergmann, Eduarda Abreu Carvalho, Raquel Zimmer, Marcelo Gomes Martins, Luciana Porcher Nedel, Anderson Maciel, and Rafael Piccin Torchelsen. An enhanced interactive endoscope model based on position-based dynamics and Cosserat rods for colonoscopy simulation. Computers and Graphics, 116 (??):345–353, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323001905.

Messina:1989:TDE

[MBGK89] Luiz Ary Messina, Annelore Buhmann, Marion Günther, and Georg Koberle. Teachware development for education in CAD. Computers and Graphics, 13(2): 237–241, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mirzendehdel:2021:OBO

[MBN21] Amir M. Mirzendehdel, Morad Behandish, and Saigopal Nelaturi. Optimizing build orientation for support removal using multiaxis machining. Computers and Graphics, 99(??):247– 258, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001424.

Muguercia:2014:FMC

[MBP14] Lien Muguercia, Carles Bosch, and Gustavo Patow. Fracture modeling in computer graphics. Computers and Graphics, 45 (??):86–100, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000806.

Maimone:2012:EPA

[MBPF12] Andrew Maimone, Jonathan Bidwell, Kun Peng, and Henry Fuchs. Enhanced personal autostereoscopic telepresence system using commodity depth cameras. Computers and Graphics, 36 (7):791–807, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001124.

Mesika:2022:CRW

[MBST22] Adi Mesika, Yizhak Ben-Shabat, and Ayellet Tal.

                                        CloudWalker:  Random

walks for 3D point cloud shape analysis. Computers and Graphics, 106(??): 110–118, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001030.

Morel:2018:SRI

[MBV18] Jules Morel, Alexandra Bac, and Cédric Véga. Surface reconstruction of incomplete datasets: a novel Poisson surface approach based on CSRBF. Computers and Graphics, 74 (??):44–55, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300633.

Martin:2010:VPC

[MC10] Tobias Martin and Elaine Cohen. Volumetric parameterization of complex objects by respecting multiple materials. Computers and Graphics, 34(3): 187–197, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000506.

Makri:2023:CBR

[MC23] Stela Makri and Panayiotis Charalambous. Curriculum based reinforcement learning for traffic simulations. Computers and Graphics, 113(??):32–42, June 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000468.

Maglo:2012:PCM

[MCAH12] Adrien Maglo, Clément Courbet, Pierre Alliez, and Céline Hudelot. Progressive compression of manifold polygon meshes. Computers and Graphics, 36 (5):349–359, August 2012.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000623.

McCormick:1984:PFC

[McC84] J. J. McCormick. Present and future color display technologies for graphics. Computers and Graphics, 8(3): 281–293, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McConnell:1995:CGE

[McC95] Jeffrey J. McConnell. Computer graphics education: Issues from multiple perspectives. Computers and Graphics, 19(2):331–334, March–April 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400161.

McConnell:1996:AGL

[McC96] Jeffrey J. McConnell. Active and group learning techniques and their use in graphics education. Computers and Graphics, 20(1): 177–180, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.com/

cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500103.

McClure:2002:CGS

[McC02] Mark McClure. Chaos and graphics: a stochastic cellular automaton for threecoloring Penrose tiles. Computers and Graphics, 26(3): 519–524, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/54/40/abstract.
html.

McClure:2008:IVK

[McC08] Mark McClure. Images of a vibrating Koch drum. Computers and Graphics, 32(6): 711–715, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000691.

McDonald:2006:FCT

[McD06] John McDonald. Fractal classifications of trajectories in a nonlinear massspring system. Computers and Graphics, 30(5):815– 833, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001300.

McGraw:2008:GRD

[McG08] Tim McGraw. Generalized reaction-diffusion textures. Computers and Graphics, 32 (1):82–92, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001689.

Monteiro:2023:EUE

[MCG^{+}23] Pedro Monteiro, Hugo Coelho, Guilherme Gonçalves, Miguel Melo, and Maximino Bessa. Exploring the user experience of hands-free VR interaction methods during a Fitts’ task. Computers and Graphics, 117(??): 1–12, December 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932300242X.

Martin:2012:MEV

[MCK12] Tobias Martin, Elaine Cohen, and Robert M. Kirby. Mixed-element volume completion from NURBS surfaces. Computers and Graphics, 36(5):548–554, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000477.

Mikovec:2006:BTI

[MCKS06] Zdenek Mikovec, Ladislav Cmolik, Jiri Kopsa, and Pavel Slavik. Beyond traditional interaction in a mobile environment: New approach to 3D scene rendering. Computers and Graphics, 30 (5):714–726, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001221.

Monteiro:2018:ASL

[MCM^{+}18] Pedro Monteiro, Diana Carvalho, Miguel Melo, Frederico Branco, and Maximino Bessa. Application of the steering law to virtual reality walking navigation interfaces. Computers and Graphics, 77 (??):80–87, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301584.

Martins:2014:VAD

[MCMT14] Rafael Messias Martins, Danilo Barbosa Coimbra, Rosane Minghim, and A. C. Telea. Visual analysis of dimensionality reduction quality for parameterized projections. Computers and Graphics, 41(??): 26–42, June 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000235.

Marques:2022:SCC

[MCMV22] Bruno Augusto Dorta Marques, Esteban Walter Gonzalez Clua, Anselmo Antunes Montenegro, and Cristina Nader Vasconcelos. Spatially and color consistent environment lighting estimation using deep neural networks for mixed reality. Computers and Graphics, 102(??):257–268, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001710.

Miola:2022:CAM

[MCP^{+}22] Marianna Miola, Daniela Cabiddu, Simone Pittaluga, Michela Mortara, Marino Vetuschi Zuccolini, and Gianmario Imitazione. A computational approach for 3D modeling and integration of heterogeneous geo-data. Computers and Graphics, 105(??):105–118, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000681.

Musse:2021:FSS

[MCPW21] Soraia Raupp Musse, Roberto Cesar, Jr., Nuria Pelechano, and Zhangyang (Atlas) Wang. Foreword to the special section on SIBGRAPI-Conference on Graphics, Patterns and Images is an international conference 2020. Computers and Graphics, 94 (??):A5–A6, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030176X.

Mykhalchuk:2013:LTM

[MCS13] Vasyl Mykhalchuk, Frederic Cordier, and Hyewon Seo. Landmark transfer with minimal graph. Computers and Graphics, 37 (5):539–552, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000575.

Meng:2018:ECF

[MCS^{+}18] Wenlong Meng, Shuangmin Chen, Zhenyu Shu, Shi-Qing Xin, Hongbo Fu, and Changhe Tu. Efficiently computing feature-aligned and high-quality polygonal offset surfaces. Computers and Graphics, 70(??): 62–70, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300961.

Melero:2005:CSO

[MCT05] F. J. Melero, P. Cano, and J. C. Torres. Combining SP-Octrees and impostors for the visualisation of multiresolution models. Computers and Graphics, 29(2):225–233, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Melero:2008:BPO

[MCT08] F. J. Melero, P. Cano, and J. C. Torres. Boundingplanes Octree: a new volume-based LOD scheme. Computers and Graphics, 32 (4):385–392, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000733.

Maule:2011:SRB

[MCTB11] Marilena Maule, João L. D. Comba, Rafael P. Torchelsen, and Rui Bastos. A survey of raster-based transparency techniques. Computers and Graphics, 35(6): 1023–1034, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931100135X.

Maule:2014:MOO

[MCTB14] Marilena Maule, João L. D. Comba, Rafael Torchelsen, and Rui Bastos. Memoryoptimized order-independent transparency with Dynamic Fragment Buffer. Computers and Graphics, 38(??): 1–9, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001246.

Marques:2018:DSH

[MCV18] Bruno Augusto Dorta Marques, Esteban Walter Gonzalez Clua, and Cristina Nader Vasconcelos. Deep spherical harmonics light probe estimator for mixed reality games. Computers and Graphics, 76(??):96– 106, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301377.

McWhinnie:1987:PCI

[McW87] Harold J. McWhinnie. The personal computer as an interactive technology for the artist (simulation vs. abstraction). Computers and Graphics, 11(3): 315–319, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1988:EM

[McW88a] Harold J. McWhinnie. The electronic museum. Computers and Graphics, 12 (2):269, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1988:TUC

[McW88b] Harold J. McWhinnie. Technology update: Computer graphics software. Computers and Graphics, 12(1): 129–131, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1989:DVL

[McW89] Harold J. McWhinnie. Development of visual languages with interactive video disks. Computers and Graphics, 13(1):115–116, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1990:TIC

[McW90] Harold James McWhinnie. Two individual computer workstations. Computers and Graphics, 14(3–4): 527–530, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1991:CRS

[McW91a] Harold J. McWhinnie. The computer and the right side of the brain II. Computers and Graphics, 15(2): 313–318, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

McWhinnie:1991:SAQ

[McW91b] Harold J. McWhinnie. Some aesthetic questions on computerbased art and design. Computers and Graphics, 15(1): 139–142, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Merrett:1985:RSP

[MD85] T. H. Merrett and B. Düchting.

Relational storage and processing of two-dimensional diagrams. Computers and Graphics, 9(3):247– 258, ???? 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849385900512.

Mak:1999:TSC

[MD99a] Brenda L. Mak and Al Degennaro. Technical section

— computer graphics for art creation: cultural biases against its acceptance in education. Computers and Graphics, 23(3):419– 427, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/692.pdf.

Marsan:1999:CGI

[MD99b] Anne L. Marsan and Debasish Dutta. Computer graphics in India — computational techniques for automatically tiling and skinning branched objects. Computers and Graphics, 23(1):111–126, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/654.pdf.

Margala:1995:MBG

[MDJ^{+}95] Martin Margala, Nelson G. Durdle, Scott Juskiw, V. James Raso, and Doug L. Hill. A 33 MHz 16-bit gradient calculator for real-time volume imaging. Computers and Graphics, 19(5):679–684, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500046.

Manda:2021:PCA

[MDM^{+}21] Bharadwaj Manda, Shubham Dhayarkar, Sai Mitheran, V. K. Viekash, and Ramanathan Muthuganapathy.

`CADSketchNet' — an an-

notated sketch dataset for 3D CAD model retrieval with deep neural networks.

Computers and Graphics, 99 (??):100–113, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001333.

Ma:2021:SRC

[MDS^{+}21] Zhao Ma, Simon Duenser, Christian Schumacher, Romana Rust, Moritz Bächer, Fabio Gramazio, Matthias Kohler, and Stelian Coros. Stylized robotic clay sculpting. Computers and Graphics, 98(??):150–164, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321000959. See

[MDS^{+}22].

Ma:2022:ESR

[MDS^{+}22] Zhao Ma, Simon Duenser, Christian Schumacher, Romana Rust, Moritz Bächer, Fabio Gramazio, Matthias Kohler, and Stelian Coros. Erratum to “Stylized robotic clay sculpting” [Comput. Graph. 98 (2021) 150–164]. Computers and Graphics, 102(??):646, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001813.

See [MDS^{+}21].

Matsokin:1988:MPC

[MDSU88] A. M. Matsokin, V. A. Debelov, V. G. Sirotin, and S. A. Upol’nikov. Multipurpose computer graphics system SMOG-85. Computers and Graphics, 12(3– 4):441–456, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1977:NOL

[ME77] N. Marovac and W. S. Elliot. A network-oriented language — a new approach to network design, using interactive graphics. Computers and Graphics, 2(4): 235–239, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1983:ICA

[ME83] N. Marovac and W. S. Elliott. Interactive computer aided design using computer graphics. Computers and Graphics, 7(2): 177–187, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mascarenhas:1992:IRS

[ME92] Nelson D. A. Mascarenhas and Guaraci J. Erthal. Image registration by sequential tests of hypotheses. the relationship between Gaussian and binomial models.

Computers and Graphics, 16(3):259–264, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Morrical:2017:PQC

[ME17] Nathan Morrical and John Edwards. Parallel quadtree construction on collections of objects. Computers and Graphics, 66(??):162– 168, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300730.

Medland:1986:AAS

[Med86] A. J. Medland. The alignment of associated space in an automated CADkinematics procedure. Computers and Graphics, 10(3): 239–244, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Meintzen:1983:SYL

[Mei83] H. Meintzen. Some years of lectures and practicals in computer graphics. Computers and Graphics, 7(3– 4):369–379, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Melzi:2019:SRS

[Mel19] Simone Melzi. Sparse representation of step functions on manifolds. Computers and Graphics, 82 (??):117–128, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300718.

Menard:1985:ESD

[Men85] D. Menard. An electrical schematic design station. Computers and Graphics, 9 (2):159–162, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Malik:2019:SED

[MENS19] Jameel Malik, Ahmed Elhayek, Fabrizio Nunnari, and Didier Stricker. Simple and effective deep hand shape and pose regression from a single depth image. Computers and Graphics, 85 (??):85–91, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301591.

Messina:1984:JCS

[Mes84] Luiz Ary Messina. Justification of CAD systems. Computers and Graphics, 8(1): 105–106, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mestetskii:2000:FCR

[Mes00]  L. M. Mestetskii.  Fat

curves and representation of planar figures. Computers and Graphics, 24(1):9– 21, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/28/article.
pdf.

Metzner:1985:IV

[Met85]  R. Metzner.  Interac-

tive video. Computers and Graphics, 9(2):163–164, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Meyer:1979:LCS

[Mey79] B. Meyer. A low cost satellite for fast interactive graphics in a timesharing environment. Computers and Graphics, 4(3– 4):177–183, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Molina:2002:TSM

[MF02] Andrés Molina and Francisco R. Feito. Technical section: a method for testing anisotropy and quantifying its direction in digital images. Computers and Graphics, 26(5):771–784, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/40/abstract.
html.

Meng:2011:CEF

[MFL11] Min Meng, Lubin Fan, and Ligang Liu. A comparative evaluation of foreground/background sketchbased mesh segmentation algorithms. Computers and Graphics, 35(3):650– 660, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931100077X.

Mukai:1994:ADS

[MFOK94] Shinji Mukai, Susumu Furukawa, Makoto Obi, and Fumihiko Kimura. An algorithm for deciding similarities of convex polyhedra. Computers and Graphics, 18(2):171–176, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miao:2011:VSG

[MFP11] Yongwei Miao, Jieqing Feng, and Renato Pajarola. Visual saliency guided normal enhancement technique for 3D shape depiction. Computers and Graphics, 35(3):

706–712, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000562.

Ma:1986:DLC

[MG86] Song De Ma and A. Gagalowicz. Determination of local coordinate systems for texture synthesis on 3-D surfaces. Computers and Graphics, 10(2):171–176, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marcheix:1998:NMI

[MG98] David Marcheix and Stefka Gueorguieva. Nibble meshing: Incremental triangulation of non-manifold solid boundary. Computers and Graphics, 22(2–3):181–188, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/539.pdf.

Merillou:2008:SAW

[MG08] S. Mérillou and D. Ghazanfarpour. A survey of aging and weathering phenomena in computer graphics. Computers and Graphics, 32(2):159–174, April 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000058.

Martinek:2009:ORA

[MG09] Michael Martinek and Roberto Grosso. Optimal rotation alignment of 3D objects using a GPU-based similarity function. Computers and Graphics, 33(3): 291–298, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000314.

Maglo:2013:PCP

[MGH13] Adrien Maglo, Ian Grimstead, and Céline Hudelot. POMAR: Compression of progressive oriented meshes accessible randomly. Computers and Graphics, 37 (6):743–752, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000794.

Mohammadi:2022:GAL

[MGH22] Iman Soltani Mohammadi, Mohammad Ghanbari, and Mahmoud Reza Hashemi. GAMORRA: an API-level workload model for rasterizationbased graphics pipeline architecture. Computers and Graphics, 106(??):9– 19, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000796.

Moench:2011:CAM

[MGJ^{+}11] Tobias Moench, Rocco Gasteiger, Gabor Janiga, Holger Theisel, and Bernhard Preim. Context-aware mesh smoothing for biomedical applications. Computers and Graphics, 35 (4):755–767, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001142.

Mittenentzwei:2023:IUB

[MGM^{+}23] Sarah Mittenentzwei, Laura A.

Garrison, Eric Mörth, Kai Lawonn, Stefan Bruckner, Bernhard Preim, and Monique Meuschke. Investigating user behavior in slideshows and scrollytelling as narrative genres in medical visualization. Computers and Graphics, 114 (??):229–238, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001061.

Marto:2022:SMV

[MGMB22] Anabela Marto, Alexandrino Gonçalves, Miguel Melo, and Maximino Bessa. A survey of multisensory VR and AR applications for cultural heritage. Computers and Graphics, 102 (??):426–440, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002120.

Marhl:1996:ESS

[MGOH96] Marko Marhl, Nikola Guid, Crtomir Oblonsek, and Matjaz Horvat. Extensions of sweep surface constructions. Computers and Graphics, 20(6):893–903, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600059.

Meuschke:2022:NMV

[MGS^{+}22] Monique Meuschke, Laura A. Garrison, Noeska N. Smit, Benjamin Bach, Sarah Mittenentzwei, Veronika Weiß, Stefan Bruckner, Kai Lawonn, and Bernhard Preim. Narrative medical visualization to communicate disease data. Computers and Graphics, 107(??):144– 157, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200139X.

Martin-Gutierrez:2010:DVA

[MGSC^{+}10] Jorge Martín-Gutiérrez, José Luís Saorín, Manuel Contero, Mariano Alcañiz, David C. Pérez-López, and Mario Ortega. Design and validation of an augmented book for spatial abilities development in engineering students. Computers and Graphics, 34 (1):77–91, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001514.

Markus:1989:LEL

[MH89] Mario Markus and Benno Hess. Lyapunov exponents of the logistic map with periodic forcing. Computers and Graphics, 13(4): 553–558, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Matulis:2021:RAD

[MH21] Marius Matulis and Carlo Harvey. A robot arm digital twin utilising reinforcement learning. Computers and Graphics, 95(??): 106–114, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100011X.

Martinez:2015:CSO

[MHCL15] Jonas Martinez, Samuel Hornus, Frédéric Claux, and Sylvain Lefebvre. Chained segment offsetting for raybased solid representations. Computers and Graphics, 46 (??):36–47, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001058.

Millar:1999:CGI

[MHK99] R. J. Millar, J. R. P. Hanna, and S. M. Kealy. Computer graphics in India — a review of behavioural animation. Computers and Graphics, 23(1):127–143, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/655.pdf.

Mahmudi:2016:AOC

[MHLB16] Mentar Mahmudi, Pawan Harish, Benoît Le Callennec, and Ronan Boulic. Artistoriented 3D character posing from 2D strokes. Computers and Graphics, 57(??): 81–91, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300218.

Mones-Hattal:1995:EVT

[MHM95] Barbara Mones-Hattal and Evans Mandes. Enhancing visual thinking and learning with computer graphics and virtual environment design. Computers and Graphics, 19(6):889–894, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500086.

Ma:2010:ADC

[MHW10] Lujie Ma, Zhengdong Huang, and Yanwei Wang. Automatic discovery of common design structures in CAD models. Computers and Graphics, 34(5):545– 555, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000853.

Meng:2023:VFF

[MHYN23] Qingkuo Meng, Yongjian Huai, Jiawei You, and Xiaoying Nie. Visualization of 3D forest fire spread based on the coupling of multiple weather factors. Computers and Graphics, 110 (??):58–68, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322002254.

McCartney:1994:LTS

[MHZ94] J. McCartney, B. K. Hinds, and J. J. Zhang. Leather texture synthesis and rendering. Computers and Graphics, 18(1):87–92, January– February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Michelitsch:1990:CMG

[Mic90] Michael Michelitsch. Color maps generated by “trigonometric iteration loops”. Computers and Graphics, 14 (1):125–126, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mohanto:2022:IVF

[MIGS22] Bipul Mohanto, ABM Tariqul Islam, Enrico Gobbetti, and Oliver Staadt. An integrative view of foveated rendering. Computers and Graphics, 102(??):474–501, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002211.

Miller:1975:GBD

[Mil75] I. M. Miller. Graphics in business decision making. Computers and Graphics, 1(2–3):293–296, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miller:1992:CGF

[Mil92a] James C. Miller. Computer Graphics by Francis S. Hill. Computers and Graphics, 16(4):451–??, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miller:1992:CGP

[Mil92b] James C. Miller. Computer Graphics Principles and Practice, Second Edition, by James Foley, Andries van Dam, Steven Feiner and John Hughes. Computers and Graphics, 16 (2):239–??, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miller:1993:TVF

[Mil93] G. Todd Miller. A tutorial on the visualization of forward orbits associated with Siegel disks in the quadratic Julia sets. Computers and Graphics, 17(3): 321–324, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miller:2003:RAC

[Mil03] James R. Miller. The remote application controller. Computers and Graphics, 27 (4):605–615, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

MirandaICanals:1987:CIC

[Mir87] Jaume Miranda I.Canals. Cartographic institute of Catalonia. Computers and Graphics, 11(4):477, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mitacek:1977:UCS

[Mit77] Paul Mitacek, Jr. Use of computer simulated experimental data in an advanced laboratory course. Computers and Graphics, 2(3): 175–177, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mittelman:1987:AWN

[Mit87] Jeffrey P. Mittelman. The ARF workstation: a new application of human-computer interface principles. Computers and Graphics, 11(2): 147–156, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marcilio-Jr:2021:CAS

[MJEG21] Wilson E. Marcílio-Jr, Danilo M. Eler, and Rogério E. Garcia. Contrastive analysis for scatterplot-based representations of dimensionality reduction. Computers and Graphics, 101(??):46– 58, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001783.

Managaki:1983:PMM

[MK83] M. Managaki and K. Kawagoe. A parametric man/ machine interaction with semantic data. Computers and Graphics, 7(3–4): 233–242, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Maciejewski:1985:SAS

[MK85] A. A. Maciejewski and C. A.

Klein. SAM — animation software for simulating articulated motion. Computers and Graphics, 9(4): 383–391, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Margalit:1989:ACU

[MK89] Avraham Margalit and Gary D. Knott. Algorithm for computing the union, intersection or difference of two polygons. Computers and Graphics, 13(2): 167–183, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mohler:2003:SIE

[MK03] James L. Mohler and Lars Kjelldahl. Special issue on education. Computers and Graphics, 27(3):327– 328, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marroquim:2008:EIR

[MKC08] Ricardo Marroquim, Martin Kraus, and Paulo Roma Cavalcanti. Efficient image reconstruction for pointbased and line-based rendering. Computers and Graphics, 32(2):189–203, April 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000265.

Manda:2022:SDL

[MKDM22] Bharadwaj Manda, Prasad Pralhad Kendre, Subhrajit Dey, and Ramanathan Muthuganapathy. SketchCleanNet

— a deep learning approach to the enhancement and correction of query sketches for a 3D CAD model retrieval system. Computers and Graphics, 107(??): 73–83, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001212.

Mroz:2000:MIP

[MKG00] Lukas Mroz, Andreas König, and Eduard Gröller. Maximum intensity projection at warp speed. Computers and Graphics, 24(3): 343–352, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/30/article.
pdf.

Miranda:2001:ALA

[MKHN01] Fábio Roberto Miranda, João Eduardo Kögler, Jr., Emílio Del Moral Hernandez, and Márcio Lobo Netto. An artificial life approach for the animation of cognitive characters. Computers and Graphics, 25(6):955–964, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/30/abstract.
html.

Muraleedharan:2018:RCP

[MKKM18] Lakshmi Priya Muraleedharan, Shyam Sundar Kannan, Ameya Karve, and Ramanathan Muthuganapathy. Random cutting plane approach for identifying volumetric features in a CAD mesh model. Computers and Graphics, 70(??):51– 61, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301188.

Muraleedharan:2019:ABP

[MKM19] Lakshmi Priya Muraleedharan, Shyam Sundar Kannan, and Ramanathan Muthuganapathy. Autoencoder-based part clustering for part-in-whole retrieval of CAD models. Computers and Graphics, 81(??):41–51, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300391.

Methirumangalath:2017:HDP

[MKPM17] Subhasree Methirumangalath, Shyam Sundar Kannan, Amal Dev Parakkat, and Ramanathan Muthuganapathy. Hole detection in a planar point set: an empty disk approach. Computers and Graphics, 66(??):124– 134, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300559.

Moss:1979:PSC

[ML79] R. Moss and A. Lindgard. Parametric spline curves in integer arithmetic designed for use in microcomputer controlled plotters. Computers and Graphics, 4(1): 51–61, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Montero:2012:SPC

[ML12] Andrés Solís Montero and Jochen Lang. Skeleton pruning by contour approximation and the integer medial axis transform. Computers and Graphics, 36 (5):477–487, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000684.

Manolas:2022:AGF

[MLC^{+}22] Iason Manolas, Francesco Laccone, Gianmarco Cherchi, Luigi Malomo, and Paolo Cignoni. Automated generation of flat tileable patterns and 3D reduced model simulation. Computers and Graphics, 106 (??):141–151, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000929.

Marin-Lora:2023:CSI

[MLCMGR23] Carlos Marín-Lora, Miguel Chover, Micaela Y. Martín, and Linda García-Rytman. Comparative study of interaction methods for mobile gaming while running on a treadmill. Computers and Graphics, 117(??): 164–171, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002558.

Ma:2003:RAH

[MLM03] Kwan-Liu Ma, Eric B. Lum, and Shigeru Muraki. Recent advances in hardwareaccelerated volume rendering. Computers and Graphics, 27(5):725–734, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Machado:2017:VSM

[MLM^{+}17] Vinicius Machado, Roger Leite, Felipe Moura, Sergio Cunha, Filip Sadlo, and João L. D. Comba. Visual soccer match analysis using spatiotemporal positions of players. Computers and Graphics, 68 (??):84–95, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301358.

Min:2001:CBP

[MLP01] Kyung Ha Min, In-Kwon Lee, and Chan-Mo Park. Component-based polygonal approximation of soft objects. Computers and Graphics, 25(2):245–257, April 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/29/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/29/32/article.
pdf.

Mancinelli:2019:CMG

[MLP19] Claudio Mancinelli, Marco Livesu, and Enrico Puppo. A comparison of methods for gradient field estimation on simplicial meshes. Computers and Graphics, 80 (??):37–50, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300287.

Miao:2002:BPC

[MLPB02] Lanfang Miao, Xinguo Liu, Qunsheng Peng, and Hujun Bao. Best papers of CAD & CG 2001: BRDC: binary representation of displacement code for line. Computers and Graphics, 26(3):401– 408, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/29/abstract.
html.

Manocha:2018:FSI

[MLX18] Dinesh Manocha, Ligang Liu, and Kai Xu. Foreword to the special issue on CAD/Graphics 2017. Computers and Graphics, 70 (??):A3–A6, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301930.

Merrell:2010:EBC

[MM10] Paul Merrell and Dinesh Manocha. Example-based curve synthesis. Computers and Graphics, 34(4):304– 311, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000737.

Mahmood:2018:EVP

[MM18] Salman Mahmood and Klaus Mueller. An exploded view paradigm to disambiguate scatterplots. Computers and Graphics, 73(??): 37–46, June 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300323.

Martin:2001:TBF

[MMALRA01] Marcos Martín, Miguel Martín, Carlos Alberola-López, and Juan Ruiz-Alzola. A topology-based filling algorithm. Computers and Graphics, 25(3):

493–509, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/38/article.
pdf.

Males:2020:IFV

[MMD^{+}20] Jan Males, Eva Monclús, Jose Díaz, Isabel Navazo, and Pere-Pau Vázquez. Interactive framework for the visual exploration of colonic data. Computers and Graphics, 91(??):39– 51, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300947.

Mello:2022:UEB

[MMdOE^{+}22] Claudio Dornelles Mello, Bryan Umpierre Moreira, Paulo Jefferson Dias de Oliveira Evald, Paulo Jorge Lilles Drews, and Silvia Silva da Costa Botelho. Underwater enhancement based on a self-learning strategy and attention mechanism for high-intensity regions. Computers and Graphics, 107(??):264–276, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001492.

Marcelino:2003:CMS

[MMF03] Luis Marcelino, Norman Murray, and Terrence Fernando. A constraint manager to support virtual maintainability. Computers and Graphics, 27(1):19– 26, February 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mindek:2017:DSV

[MMGB17] Peter Mindek, Gabriel Mistelbauer, Eduard Gröller, and Stefan Bruckner. Datasensitive visual navigation. Computers and Graphics, 67 (??):77–85, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300614.

Murphy:2021:AGG

[MMH^{+}21] Christian Murphy, Sudhir Mudur, Daniel Holden, Marc-André Carbonneau, Donya Ghafourzadeh, and Andre Beauchamp. Artist guided generation of video game production quality face textures. Computers and Graphics, 98(??):268– 279, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001199.

Meseth:2004:RFB

[MMK04] Jan Meseth, Gero Müller, and Reinhard Klein. Reflectance field based realtime, high-quality rendering of bidirectional texture functions. Computers and Graphics, 28(1):105– 112, February 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Morrison:2011:CUM

[MML^{+}11] Ann Morrison, Alessandro Mulloni, Saija Lemmelä, Antti Oulasvirta, Giulio Jacucci, Peter Peltonen, Dieter Schmalstieg, and Holger Regenbrecht. Collaborative use of mobile augmented reality with paper maps. Computers and Graphics, 35 (4):789–799, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001129.

Mould:2012:ERV

[MML12] David Mould, Regan L.

Mandryk, and Hua Li. Emotional response and visual attention to nonphotorealistic images. Computers and Graphics, 36 (6):658–672, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931200088X.

Melzi:2020:IEE

[MMM^{+}20] Simone Melzi, Riccardo Marin, Pietro Musoni, Filippo Bardon, Marco Tarini, and Umberto Castellani. Intrinsic/extrinsic embedding for functional remeshing of 3D shapes. Computers and Graphics, 88 (??):1–12, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300194.

Mas:2018:HDI

[MMP18] Albert Mas, Ignacio Martín, and Gustavo Patow. Heuristic driven inverse reflector design. Computers and Graphics, 77(??):1– 15, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301444.

Major:1989:DBO

[MMS89] F. Major, J. Malenfant, and N. F. Stewart. Distance between objects represented by octtrees defined in different coordinate systems. Computers and Graphics, 13(4):497–503, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marcos:2004:VKD

[MMS04] Adèrito Fernandes Marcos, Wolfgang Müller, and Heidi Schumann. Visual knowledge discovery. Computers and Graphics, 28(3):309– 310, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Michels:2015:PBA

[MMS15] Dominik L. Michels, J. Paul T.

Mueller, and Gerrit A. Sobottka. A physically based approach to the accurate simulation of stiff fibers and stiff fiber meshes. Computers and Graphics, 53 (part B)(??):136–146, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001624.

Mendes:2017:DEN

[MMS^{+}17] Daniel Mendes, Daniel Medeiros, Maurício Sousa, Eduardo Cordeiro, Alfredo Ferreira, and Joaquim A. Jorge. Design and evaluation of a novel out-of-reach selection technique for VR using iterative refinement. Computers and Graphics, 67 (??):95–102, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300778.

Michahelles:2003:AWS

[MMSS03] Florian Michahelles, Peter Matter, Albrecht Schmidt, and Bernt Schiele. Applying wearable sensors to avalanche rescue. Computers and Graphics, 27(6): 839–847, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Musleh:2023:UGP

[MMT^{+}23] Maath Musleh, Ludvig Paul Muren, Laura Toussaint, Anne Vestergaard, Eduard Gröller, and Renata G. Raidou. Uncertainty guidance in proton therapy planning visualization. Computers and Graphics, 111 (??):166–179, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000250.

Mitani:1988:DCA

[MMV88] Yoshito N. Mitani, H. R. Raya Merchand, and E. Ortiz Velazquez. Development of computer-aided design systems for metal-forming processes. Computers and Graphics, 12(3–4):371–379, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mura:2014:ARD

[MMV^{+}14] Claudio Mura, Oliver Mattausch, Alberto Jaspe Villanueva, Enrico Gobbetti, and Renato Pajarola. Automatic room detection and reconstruction in cluttered indoor environments with complex room layouts. Computers and Graphics, 44 (??):20–32, November 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000636.

Murakami:1990:UFM

[MN90] Tamotsu Murakami and Naomasa Nakajima. Using features for machine design problems. Computers and Graphics, 14(2): 201–210, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mikami:2016:EGJ

[MNI^{+}16] Koji Mikami, Yosuke Nakamura, Akinori Ito, Motonobu Kawashima, Taichi Watanabe, Yoshihiro Kishimoto, and Kunio Kondo. Effectiveness of Game Jambased iterative program for game production in Japan. Computers and Graphics, 61 (??):1–10, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300863.

Muntoni:2019:MFS

[MNS^{+}19] Alessandro Muntoni, Stefano Nuvoli, Andreas Scalas, Alessandro Tola, Luigi Malomo, and Riccardo Scateni. Mill and fold: Shape simplification for fabrication. Computers and Graphics, 80 (??):17–28, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300263.

Mudur:1999:CGIb

[MNSJ99] S. P. Mudur, Niranjan Nayak, Shrinath Shanbhag, and R. K. Joshi. Computer graphics in India — an architecture for the shaping of Indic texts. Computers and Graphics, 23(1):7–24, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        1/647.pdf.

Martins:1990:AOU

[MO90] F. Mario Martins and J. Nuno Oliveira. Archetypeoriented user interfaces. Computers and Graphics, 14 (1):17–28, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Marovac:1992:HTC

[MO92] Nenad Marovac and Larry Osburn. HyperNet: a tool to choreograph worldwide distributed hypermedia documents. Computers and Graphics, 16(2):197– 202, Summer 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mendoza:2006:ICD

[MO06] C. Mendoza and C. O’Sullivan.

Interruptible collision detection for deformable objects. Computers and Graphics, 30(3):432–438, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000665.

Mohilner:1977:RPM

[Moh77] Patricia R. Mohilner. Retrace prevention in multicurve graphs. Computers and Graphics, 2(3):163–167, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mokrzycki:1987:PMG

[Mok87] Wojciech Mokrzycki. Peak and medium-global errors of discretization of algebraic curves on square grids. Computers and Graphics, 11(3): 269–273, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mokrzycki:1988:ADA

[Mok88] Wojciech Mokrzycki. Algorithms of discretization of algebraic spatial curves on homogeneous cubical grids. Computers and Graphics, 12 (3–4):477–487, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Molnar:1996:PTI

[Mol96] Steven Molnar. The PixelFlow texture and image subsystem. Computers and Graphics, 20(4): 491–502, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600021.

Montefusco:1987:NPS

[Mon87] Laura Bacchelli Montefusco. An nteractive procedure for shape preserving cubic sspline interpolation. Computers and Graphics, 11(4): 389–392, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Morris:1975:RTA

[Mor75] G. C. Morris. Real time animation of dynamic processes. Computers and Graphics, 1(2–3):221–226, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Morris:1976:FTT

[Mor76] L. R. Morris. Fast transformation of three-dimensional graphics structures via mixedpoint arithmetic. Computers and Graphics, 2(1):7–10, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Manni:2021:PSI

[MOS^{+}21] Alessandro Manni, Damiano Oriti, Andrea Sanna, Francesco De Pace, and Federico Manuri. Snap2cad: 3D indoor environment reconstruction for AR/VR applications using a smartphone device. Computers and Graphics, 100(??): 116–124, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100145X.

Mould:2013:IVA

[Mou13] David Mould. Image and video abstraction using cumulative range geodesic filtering. Computers and Graphics, 37(5):413–430, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000356.

Mould:2015:FSS

[Mou15] David Mould. Foreword to the Special Section on Expressive 2014. Computers and Graphics, 52(??): A1, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000862.

Miller:1989:GDC

[MP89] Gavin Miller and Andrew Pearce. Globular dynamics: a connected particle system for animating viscous fluids. Computers and Graphics, 13(3):305–309, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ma:1993:PVV

[MP93] Kwan-Liu Ma and James S. Painter. Parallel volume visualization on workstations. Computers and Graphics, 17 (1):31–37, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Martin:2019:RRP

[MP19] Ignacio Martin and Gustavo Patow. Ruleset-rewriting for procedural modeling of buildings. Computers and Graphics, 84(??):93– 102, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301311.

Mancinelli:2022:VGS

[MP22] Claudio Mancinelli and Enrico Puppo. Vector graphics on surfaces using straightedge and compass constructions. Computers and Graphics, 105 (??):46–56, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000589.

Meunier:2010:CLI

[MPA^{+}10] Sylvain Meunier, Romuald Perrot, Lilian Aveneau, Daniel Meneveaux, and Djamchid Ghazanfarpour. Cosine lobes for interactive direct lighting in dynamic scenes. Computers and Graphics, 34(6): 767–778, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001275.

Munoz-Pandiella:2023:ADC

[MPAC^{+}23] Imanol Munoz-Pandiella, Carlos Andujar, Begonya Cayuela, Xavier Pueyo, and Carles Bosch. Automated digital color restitution of mural paintings using minimal art historian input. Computers and Graphics, 114(??):316– 325, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001164.

Marion-Poty:2002:WDS

[MPL02] V. Marion-Poty and Wilfrid Lefer. A wavelet decomposition scheme and compression method for streamline-based vector field visualizations. Computers and Graphics, 26 (6):899–906, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849302001784.

Meuschke:2021:ASV

[MPL21] Monique Meuschke, Bernhard Preim, and Kai Lawonn. Aneulysis — a system for the visual analysis of aneurysm data. Computers and Graphics, 98 (??):197–209, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001151.

Methirumangalath:2015:UAT

[MPM15] Subhasree Methirumangalath, Amal Dev Parakkat, and Ramanathan Muthuganapathy. A unified approach towards reconstruction of a planar point set. Computers and Graphics, 51 (??):90–97, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000722.

Menezes:1996:CST

[MPOL96] V. M. Menezes, W. P. Paula Filho, A. A. Oliveira, and N. C. Lima Filho. A CAD system for telecommunications engineering in a GIS environment. Computers and Graphics, 20(3): 405–411, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600009.

Martin:1998:TSS

[MPP98] Ignacio Martin, Frederic Pérez, and Xavier Pueyo. Technical section — the SIR rendering architecture. Computers and Graphics, 22(5):601–609, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/600.pdf.

Meng:2018:RTF

[MPQG18] Xiangfei Meng, Junjun Pan, Hong Qin, and Pu Ge. Realtime fish animation generation by monocular camera. Computers and Graphics, 71(??):55–65, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317302170.

Mathias:1989:SAC

[MPR89] P. C. Mathias, L. M. Patnaik, and Sudha Ramesh. Systolic architectures in curve generation. Computers and Graphics, 13(4): 561–567, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Maver:1985:VIA

[MPS85] T. W. Maver, C. Purdie, and D. Stearn. Visual impact analysis — modelling and viewing the natural and built environment. Computers and Graphics, 9(2): 117–124, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mortara:2006:GSH

[MPS06] M. Mortara, G. Patané, and M. Spagnuolo. From geometric to semantic human body models. Computers and Graphics, 30(2): 185–196, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000264.

Meister:2021:STB

[MPSB21] Daniel Meister, Adam Pospísil, Imari Sato, and Jirí Bittner. Spatio-temporal BRDF: Modeling and synthesis. Computers and Graphics, 97(??):279–291, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000431.

Munoz-Pandiella:2022:GCA

[MPTA^{+}22] Imanol Munoz-Pandiella, Marc Comino Trinidad, Carlos Andújar, Oscar Argudo, Carles Bosch, Antonio Chica, and Beatriz Martínez. Gain compensation across LIDAR scans. Computers and Graphics, 106(??):174– 186, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001054.

Miles:2012:RVE

[MPW^{+}12] Helen C. Miles, Serban R.

Pop, Simon J. Watt, Gavin P. Lawrence, and Nigel W. John. A review of virtual environments for training in ball sports. Computers and Graphics, 36 (6):714–726, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000957.

Mao:2009:SBA

[MQW09] Chen Mao, Sheng Feng Qin, and David Wright. A sketchbased approach to human body modelling. Computers and Graphics, 33(4):521– 541, August 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000594.

Michelitsch:1989:NFH

[MR89] M. Michelitsch and O. E. Rossler. New feature in Henon’s map. Computers and Graphics, 13(2): 263–265, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mick:1990:IHP

[MR90a] S. Mick and O. Röschel. Interpolation of helical patches by kinematic rational Bézier patches. Computers and Graphics, 14(2):275–280, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Motyka:1990:CNM

[MR90b] Mark A. Motyka and Clifford A. Reiter. Chaos and Newton’s method on systems. Computers and Graphics, 14(1):131–134, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Michelitsch:1992:BSQ

[MR92] Michael Michelitsch and Otto E. Rössler. The “burning ship” and its quasi-Julia sets. Computers and Graphics, 16(4):435–438, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mascarenhas:1995:AVU

[MR95] Edward Mascarenhas and Vernon Rego. An architecture for visualization and user interaction in parallel environments. Computers and Graphics, 19 (5):739–753, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500056.

Meier:1996:FRC

[MR96] John Meier and Clifford A.

Reiter. Fractal representations of Cayley Graphs. Computers and Graphics, 20(1):163–170, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1&aid=9500101.

Mania:2005:EER

[MR05] Katerina Mania and Andrew Robinson. An experimental exploration of the relationship between subjective impressions of illumination and physical fidelity. Computers and Graphics, 29 (1):49–56, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mould:2017:DAB

[MR17] David Mould and Paul L. Rosin. Developing and applying a benchmark for evaluating image stylization. Computers and Graphics, 67 (??):58–76, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300742.

Mecca:2015:RPS

[MRC15] R. Mecca, E. Rodolà, and D. Cremers. Realistic photometric stereo using partial differential irradiance equation ratios. Computers and Graphics, 51(??): 8–16, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000679.

Martinez:2006:MRD

[MRF06] Francisco Martínez, Antonio J. Rueda, and Francisco R. Feito. The multi-L-REP decomposition and its application to a point-in-polygon inclusion test. Computers and Graphics, 30(6): 947–958, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001506.

Malqui:2019:HDS

[MRG^{+}19] Jose Luis Sotomayor Malqui, Noemí Maritza Lapa Romero, Rafael Garcia, Hande Alemdar, and João L. D. Comba. How do soccer teams coordinate consecutive passes? A visual analytics system for analysing the complexity of passing sequences using soccer flow motifs. Computers and Graphics, 84(??): 122–133, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301384.

Magalhaes:1998:TSA

[MRR98] Léo P. Magalhães, Alberto B. Raposo, and Ivan L. M. Ricarte. Technical section — animation modeling with Petri nets. Computers and Graphics, 22(6):735–743, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        6/628.pdf.

Machado:2007:DEA

[MRS^{+}07] Penousal Machado, Juan Romero, Antonino Santos, Amílcar Cardoso, and Alejandro Pazos. On the development of evolutionary artificial artists. Computers and Graphics, 31(6): 818–826, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001574.

Moltaji:2017:SMP

[MRS17] Amirhessam Moltaji, Adam Runions, and Faramarz F. Samavati. Subdivision and multiresolution for PUPs. Computers and Graphics, 62(??):53–66, ???? 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301352.

Mejia:2018:HGT

[MRSS^{+}18] Daniel Mejia, Oscar Ruiz-Salguero, Jairo R. Sánchez, Jorge Posada, Aitor Moreno, and Carlos A. Cadavid.

Hybrid geometry/topology based mesh segmentation for reverse engineering. Computers and Graphics, 73(??): 47–58, June 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300372.

Magnor:2010:F

[MRT10] Marcus Magnor, Bodo Rosenhahn, and Holger Theisel. Foreword. Computers and Graphics, 34(5):565–566, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000816.

Maack:2021:FUA

[MRW^{+}21] Robin G. C. Maack, Michael L.

Raymer, Thomas Wischgoll, Hans Hagen, and Christina Gillmann. A framework for uncertainty-aware visual analytics of proteins. Computers and Graphics, 98 (??):293–305, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001102.

Miao:2023:WSS

[MRWL23] Yongwei Miao, Guoxiang Ren, Jinrong Wang, and Fuchang Liu. Weakly supervised semantic segmentation for point cloud based on view-based adversarial training and selfattention fusion. Computers and Graphics, 116 (??):46–54, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001759.

Matsumoto:1982:GDA

[MS82] N. Matsumoto and S. Suzuki. Graphics for dynamic analysis/synthesis of vibration systems with arbitrary degrees-of-freedom (Davis). Computers and Graphics, 6 (4):159–169, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Montani:1990:RTC

[MS90] C. Montani and R. Scopigno. Ray tracing CSG trees using the sticks representation scheme. Computers and Graphics, 14(3–4):481–490, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mortara:2001:SMB

[MS01] Michela Mortara and Michela Spagnuolo. Similarity measures for blending polygonal shapes. Computers and Graphics, 25(1):13– 27, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/28/article.
pdf.

Masood:2007:CDS

[MS07] Asif Masood and M. Sarfraz. Corner detection by sliding rectangles along planar curves. Computers and Graphics, 31(3): 440–448, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000520.

Masood:2008:ETC

[MS08] A. Masood and M. Sarfraz. An efficient technique for capturing 2D objects. Computers and Graphics, 32 (1):93–104, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001847.

McCrae:2009:SPC

[MS09a] James McCrae and Karan Singh. Sketching piecewise clothoid curves. Computers and Graphics, 33 (4):452–461, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000843.

Mortara:2009:SDB

[MS09b] Michela Mortara and Michela Spagnuolo. Semantics-driven best view of 3D shapes. Computers and Graphics, 33(3):280–290, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000259.

Marroquim:2016:FSS

[MS16] Ricardo Marroquim and Pedro Sander. Foreword to special section on SIB-GRAPI 2015. Computers and Graphics, 57(??): A5–A6, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300309.

Muller:2001:FFY

[MSAR01] W. Müller, U. Spierling, M. Alexa, and Th. Rieger. Face-to-face with your assistant. Realization issues of animated user interface agents for home appliances. Computers and Graphics, 25(4):593–600, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

13/20/57/33/32/abstract.
html.

McCracken:1975:ETD

[MSD75] T. E. McCracken, B. W.

Sherman, and S. J. Dwyer, III. An economical tonal display for interactive graphics and image analysis data. Computers and Graphics, 1 (1):79–94, May 1975. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ma:2017:QAV

[MSE17] Bo Ma, Susanne K. Suter, and Alireza Entezari. Quality assessment of volume compression approaches using isovalue clustering. Computers and Graphics, 63(??): 18–27, April 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300158.

Machchhar:2020:CSS

[MSE20] Jinesh Machchhar, Henry Segerman, and Gershon Elber. Conjugate shape simplification via precise algebraic planar sweeps toward gear design. Computers and Graphics, 90(??): 1–10, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300686.

Meyers:1995:PRP

[MSF95] Stephan Meyers, Ellen Sandor, and Janine Fron. PHSColograms^{REG} and rotated PHSColograms. Computers and Graphics, 19(4): 513–522, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500029.

Men:2022:GBR

[MSHL22] Qianhui Men, Hubert P. H. Shum, Edmond S. L. Ho, and Howard Leung. GAN-based reactive motion synthesis with class-aware discriminators for human-human interaction. Computers and Graphics, 102(??):634– 645, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002089.

Meuschke:2019:ESV

[MSL^{+}19] Monique Meuschke, Noeska N. Smit, Nils Lichtenberg, Bernhard Preim, and Kai Lawonn. EvalViz — surface visualization evaluation wizard for depth and shape perception tasks. Computers and Graphics, 82 (??):250–263, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300834.

Matsumoto:2019:QCC

[MSMK19] Taishi Matsumoto, Koichiro Sato, Yoshiyuki Matsuoka, and Takeo Kato. Quantification of “complexity” in curved surface shape using total absolute curvature. Computers and Graphics, 78 (??):108–115, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301705.

Murray:2012:TAF

[MSMP12] Naila Murray, Sandra Skaff, Luca Marchesotti, and Florent Perronnin. Toward automatic and flexible concept transfer. Computers and Graphics, 36(6):622– 634, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000210.

Muller:2020:VAE

[MSO^{+}20] Juliane Müller, Matthaeus Stoehr, Alexander Oeser, Jan Gaebel, Marc Streit, Andreas Dietz, and Steffen Oeltze-Jafra. A visual approach to explainable computerized clinical decision support. Computers and Graphics, 91(??): 1–11, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300935.

Mikamo:2013:PIA

[MSR^{+}13] Michihiro Mikamo, Marcos Slomp, Bisser Raytchev, Toru Tamaki, and Kazufumi Kaneda. Perceptually inspired afterimage synthesis. Computers and Graphics, 37(4):247–255, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000319.

Montesdeoca:2017:ADW

[MSRB17] S. E. Montesdeoca, H. S. Seah, H.-M. Rall, and D. Benvenuti. Art-directed watercolor stylization of 3D animations in real-time. Computers and Graphics, 65(??):60–72, June 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300316.

Marques:2022:VCE

[MST^{+}22] Bernardo Marques, Samuel Silva, António Teixeira, Paulo Dias, and Beatriz Sousa Santos. A vision for contextualized evaluation of remote collaboration supported by AR. Computers and Graphics, 102 (??):413–425, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100220X.

Ma:1988:RBO

[MT88] Dechang Ma and Rongxi Tang. Realizing the Boolean operations in solid modeling technique via directed loops. Computers and Graphics, 12 (3–4):319–322, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Maurel:2000:HSM

[MT00] Walter Maurel and Daniel Thalmann. Human shoulder modeling including scapulothoracic constraint and joint sinus cones. Computers and Graphics, 24(2):203– 218, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/29/article.
pdf.

Musse:2018:FSS

[MTB18] Soraia Raupp Musse, Daniel Thalmann, and Rafael Bidarra. Foreword to the special section on XVII Brazilian symposium on computer games and digital entertainment (SBGames 2018). Computers and Graphics, 76 (??):A5–A6, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301481.

Meinzer:2002:CPL

[MTC02] Hans-Peter Meinzer, Matthias Thorn, and Carlos E. Cárdenas. Computerized planning of liver surgery— an overview. Computers and Graphics, 26(4):569–576, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/33/abstract.
html.

Mukundan:2022:PAC

[MTM22] Manoj Kumar Mukundan, Safeer Babu Thayyil, and Ramanathan Muthuganapathy. A parallel algorithm for computing Voronoi diagram of a set of spheres using restricted lower envelope approach and topology matching. Computers and Graphics, 106(??):210– 221, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000899.

Menin:2022:EVT

[MTN22] Aline Menin, Rafael Torchelsen, and Luciana Nedel. The effects of VR in training simulators: Exploring perception and knowledge gain. Computers and Graphics, 102 (??):402–412, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002119.

Martinek:2021:STF

[MTS21] Magdalena Martinek, Philip Thiemann, and Marc Stamminger. Spatio-temporal filtered motion DAGs for path-tracing. Computers and Graphics, 99(??):224– 233, October 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001394.

Marques:2022:CAR

[MTS^{+}22] Bernardo Marques, António Teixeira, Samuel Silva, João Alves, Paulo Dias, and Beatriz Sousa Santos. A critical analysis on remote collaboration mediated by augmented reality: Making a case for improved characterization and evaluation of the collaborative process. Computers and Graphics, 102 (??):619–633, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001709.

Mehra:2010:VNP

[MTSM10] Ravish Mehra, Pushkar Tripathi, Alla Sheffer, and Niloy J. Mitra. Visibility of noisy point cloud data. Computers and Graphics, 34(3):219–230, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000397.

Magnenat-Thalmann:1982:SUP

[MTT82] N. Magnenat-Thalmann and D. Thalmann. Some unusual primitives in the Mira graphical extensions of Pascal. Computers and Graphics, 6(3):127–139, 1982. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Magnenat-Thalmann:1984:CAS

[MTT84] N. Magnenat-Thalmann and D. Thalmann. Computer animated scenes. Computers and Graphics, 8(3): 331–333, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Magnenat-Thalmann:1982:MGS

[MTTL82] N. Magnenat-Thalmann, D. Thalmann, and A. Larouche. A multilevel graphics system based on top-down methodology. Computers and Graphics, 6(3):97–100, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Magnenat-Thalmann:1982:GIG

[MTTLL82] N. Magnenat-Thalmann, D. Thalmann, A. Larouche, and L. Lorrain. GRAFEDIT: an interactive general-purpose graphics editor. Computers and Graphics, 6(1):41–46, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mudur:1993:GEI

[Mud93] S. P. Mudur. Guest editor’s introduction. Computers and Graphics, 17(4): 345–??, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mudur:1999:CGIa

[Mud99] S. P. Mudur. Computer graphics in India — Guest Editors’ introduction. Computers and Graphics, 23(1): 3–??, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Miyata:2010:EFC

[MUH10] Kazunori Miyata, Katsuhiro Umemoto, and Takeo Higuchi. An educational framework for creating VR application through groupwork. Computers and Graphics, 34(6):811–819, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001329.

Mukundan:1986:EOA

[Muk86] R. Mukundan. An edge oriented algorithm for the evaluation of projected surface parameters. Computers and Graphics, 10(1):1–5, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mullineux:2001:CRU

[Mul01] Glen Mullineux. Constraint resolution using optimisation techniques. Computers and Graphics, 25(3): 483–492, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/37/article.
pdf.

Matsuka:1983:CTI

[MUS83] H. Matsuka, S. Uno, and K. Sugimoto. Concept and tools for interactive graphics. Computers and Graphics, 7(3–4):215–224, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mesquita:2021:NOG

[MVCNI21] Erison Miller Santos Mesquita, Creto Augusto Vidal, Joaquim Bento Cavalcante-Neto, and Rafael Fernandes Ivo. Non-overlapping geometric shadow map. Computers and Graphics, 101(??):59–71, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849321001771.

Meuschke:2021:SVM

[MVG^{+}21] Monique Meuschke, Samuel Voß, Franziska Gaidzik, Bernhard Preim, and Kai Lawonn. Skyscraper visualization of multiple timedependent scalar fields on surfaces. Computers and Graphics, 99(??):22–42, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100090X.

Munz:2022:VBI

[MVK^{+}22] Tanja Munz, Dirk Väth, Paul Kuznecov, Ngoc Thang Vu, and Daniel Weiskopf. Visualization-based improvement of neural machine translation. Computers and Graphics, 103(??):45– 60, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002594.

Meuschke:2018:EBF

[MVPL18] Monique Meuschke, Samuel Voß, Bernhard Preim, and Kai Lawonn. Exploration of blood flow patterns in cerebral aneurysms during the cardiac cycle. Computers and Graphics, 72(??): 12–25, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300128.

Melo:2018:PCI

[MVRB18] Miguel Melo, José Vasconcelos-Raposo, and Maximino Bessa. Presence and cybersickness in immersive content: Effects of content type, exposure time and gender. Computers and Graphics, 71(??):159–165, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301917.

Maximo:2014:AMC

[MVS14] Andre Maximo, Luiz Velho, and Marcelo Siqueira. Adaptive multi-chart and multiresolution mesh representation. Computers and Graphics, 38(??):332–340, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001921.

Massarwi:2018:UPC

[MvSE18] Fady Massarwi, Boris van Sosin, and Gershon Elber. Untrimming: Precise conversion of trimmed-surfaces to tensor-product surfaces. Computers and Graphics, 70 (??):80–91, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301383.

Murray:1983:CXT

[MW83] J. L. Murray and P. M.

Wilson. CAM-X turnkey CAE system. Computers and Graphics, 7(1):95–96, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Meek:1992:NPM

[MW92] D. S. Meek and D. J.

Walton. A note on planar minimax arc splines. Computers and Graphics, 16 (4):431–433, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Madi:1999:TSM

[MW99] M. M. Madi and D. J. Walton. Technical section — modeling and visualization of layered objects. Computers and Graphics, 23(3):331– 342, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/685.pdf.

Muller-Wittig:2005:E

[MW05] Wolfgang Müller-Wittig. Editorial. Computers and Graphics, 29(1):1–2, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ma:2012:IAN

[MW12] Weiyin Ma and Huawei Wang. Interpolating an arbitrary number of joint B-spline curves by loop surfaces. Computers and Graphics, 36(5):321–328, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000489.

Mihai:2014:VSC

[MW14] Mihaela Mihai and Rüdiger Westermann. Visualizing the stability of critical points in uncertain scalar fields. Computers and Graphics, 41(??):13–25, June 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000247.

Masia:2013:DAC

[MWA^{+}13] Belen Masia, Gordon Wetzstein, Carlos Aliaga, Ramesh Raskar, and Diego Gutierrez. Display adaptive 3D content remapping. Computers and Graphics, 37(8): 983–996, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001027.

Masia:2013:SCD

[MWDG13] Belen Masia, Gordon Wetzstein, Piotr Didyk, and Diego Gutierrez. A survey on computational displays: Pushing the boundaries of optics, computation, and perception. Computers and Graphics, 37(8): 1012–1038, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001441.

Ma:2022:RTS

[MWLZ22] Jing Ma, Jin Wang, Jituo Li, and Dongliang Zhang. Real-time skeletonization for sketch-based modeling. Computers and Graphics, 102(??):56–66, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002478.

Mei:2004:IDT

[MWS04] Chunhui Mei, Fuli Wu, and Jiaoying Shi. Illuminationdependent texture. Computers and Graphics, 28 (5):747–756, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Mikami:2010:CTP

[MWY^{+}10] Koji Mikami, Taichi Watanabe, Katsunori Yamaji, Kenji Ozawa, Akinori Ito, Motonobu Kawashima, Ryota Takeuchi, Kunio Kondo, and Mitsuru Kaneko. Construction trial of a practical education curriculum for game development by industry-university collaboration in Japan. Computers and Graphics, 34(6): 791–799, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001482.

Ma:2012:EAE

[MX12] Li-Qian Ma and Kun Xu. Efficient antialiased edit propagation for images and videos. Computers and Graphics, 36(8):1005–1012, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001367.

Ma:2014:EMP

[MX14] Li-Qian Ma and Kun Xu.

Efficient manifold preserving edit propagation with adaptive neighborhood size. Computers and Graphics, 38 (??):167–173, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001544.

Martin:2019:UPB

[MXK^{+}19] Rodrigo Martín, Min Xue, Reinhard Klein, Matthias B. Hullin, and Michael Weinmann. Using patchbased image synthesis to measure perceptual texture similarity. Computers and Graphics, 81(??): 104–116, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300421.

Mould:1997:MWC

[MY97] David Mould and Yee-Hong Yang. Modeling water for computer graphics. Computers and Graphics, 21(6):801–814, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700059.

Miyai:2016:NAC

[MY16] Ayumi Miyai and Yasushi Yamaguchi. New approach to camerawork skills education for 3D/S3D computer graphics animation. Computers and Graphics, 59 (??):119–129, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300838.

Mine:2015:PID

[MYC15] Mark Mine, Arun Yoganandan, and Dane Coffey. Principles, interactions and devices for real-world immersive modeling. Computers and Graphics, 48(??): 84–98, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000126.

Masuda:2006:IMD

[MYF06] H. Masuda, Y. Yoshioka, and Y. Furukawa. Interactive mesh deformation using equality-constrained least squares. Computers and Graphics, 30(6): 936–946, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600149X.

Ma:2023:ISM

[MYL^{+}23] Yuanyuan Ma, Zenghao Yang, Tao Li, Lige Xu, and Yaqiong Qiao. Image steganalysis method based on cover selection and adaptive filtered residual network. Computers and Graphics, 115(??):43– 54, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001292.

Millham:1989:LPH

[MZ89] C. B. Millham and J. L.

Zheng. A linear pivoting heuristic procedure for computing the curve of intersection of two bicubic surface patches. Computers and Graphics, 13(1):25–38, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ma:2023:TAT

[MZ23] Jing Ma and Dongliang Zhang. TARig: Adaptive template-aware neural rigging for humanoid characters. Computers and Graphics, 114(??):158– 167, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000730.

McDonnell:2021:MPP

[MZCD21] Rachel McDonnell, Katja Zibrek, Emma Carrigan, and Rozenn Dahyot. Model for predicting perception of facial action unit activation using virtual humans. Computers and Graphics, 100 (??):81–92, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001631.

Ma:2016:FMH

[MZPZ16] Long Ma, Yuanfeng Zhou, Hao Pan, and Caiming Zhang. A framework for modeling high quality tension-determined surfaces. Computers and Graphics, 61 (??):50–59, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301200.

Nasri:2002:BPC

[NA02] Ahmad H. Nasri and A. Abbas. Best papers of CAD & CG 2001: Designing Catmull–Clark subdivision surfaces with curve interpolation constraints. Computers and Graphics, 26(3):393– 400, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/28/abstract.
html.

Nah:2023:QHI

[Nah23] Jae-Ho Nah. QuickETC2-HQ:

Improved ETC2 encoding techniques for real-time, high-quality texture compression. Computers and Graphics, 116(??):308–316, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002236.

Northam:2013:SIS

[NAK13] Lesley Northam, Paul Asente, and Craig S. Kaplan. Stereoscopic 3D image stylization. Computers and Graphics, 37 (5):389–402, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200177X.

Nanci:1977:ASH

[Nan77] Dominique Nanci. Apports des sciences de l’homme à la perception et modélisation de la réalité dans les bases de données. Computers and Graphics, 2(4):249– 251, ???? 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849377900243.

Natali:2013:SDP

[NAO13] Mattia Natali, Marco Attene, and Giulio Ottonello. Steepest descent paths on simplicial meshes of arbitrary dimensions. Computers and Graphics, 37 (6):687–696, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000770.

Nappi:1995:AIT

[Nap95] Maureen Nappi. Aesthetic intention, technology and the art praxis. Computers and Graphics, 19(4): 509–512, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500028.

Narahara:2015:DET

[Nar15] Taro Narahara. Design exploration through interactive prototypes using sensors and microcontrollers. Computers and Graphics, 50(??):25–35, August 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000473.

Neukom:2018:RTG

[NAS18] Benjamin Neukom, Stefan Müller Arisona, and Simon Schubiger. Real-time GIS-based snow cover approximation and rendering for large terrains. Computers and Graphics, 71(??): 14–22, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301693.

Navazo:1989:EOR

[Nav89] I. Navazo. Extended octtree representation of general solids with plane faces: Model structure and algorithms. Computers and Graphics, 13(1):5–16, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Newman:2004:HPS

[NBE^{+}04] Timothy S. Newman, J. Brad Byrd, Pavan Emani, Amit Narayanan, and Abouzar Dastmalchi. High performance SIMD marching cubes isosurface extraction on commodity computers. Computers and Graphics, 28(2):213–233, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nadalutti:2007:VAU

[NC07] Daniele Nadalutti and Luca Chittaro. Visual analysis of users’ performance data in fitness activities. Computers and Graphics, 31(3):429–439, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000635.

Nam:2012:SSM

[NC12] Sanghun Nam and Youngho Chai. SPACESKETCH: Shape modeling with 3D meshes and control curves in stereoscopic environments. Computers and Graphics, 36 (5):526–533, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000519.

Normoyle:2023:FSS

[ND23] Aline Normoyle and Zhigang Deng. Foreword to the special section on motion, interaction, and games, 2022. Computers and Graphics, 110(??):A4– A5, February 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000031.

Novello:2020:GIN

[NdSV20a] Tiago Novello, Vincius da Silva, and Luiz Velho. Global illumination of non-Euclidean spaces. Computers and Graphics, 93 (??):61–70, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301485.

Novello:2020:VNS

[NdSV20b] Tiago Novello, Vinícius da Silva, and Luiz Velho. Visualization of Nil, Sol, and SL2(R) geometries. Computers and Graphics, 91 (??):219–231, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301187.

Negroponte:1977:RSA

[Neg77] Nicholas Negroponte. Raster scan approaches to computer graphics. Computers and Graphics, 2(3):179–193, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Neroth:1975:GPS

[Ner75] C. C. Neroth. A graphical programming system with speech input. Computers and Graphics, 1(2–3): 227–232, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nilsson:2006:MBU

[NFHS06] Erik G. Nilsson, Jacqueline Floch, Svein Hallsteinsen, and Erlend Stav. Model-based user interface adaptation. Computers and Graphics, 30(5):692– 701, October 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001208.

Nadler:1999:QQV

[NFLYCO99] Boaz Nadler, Gadi Fibich, Shuly Lev-Yehudi, and Daniel Cohen-Or. A qualitative and quantitative visibility analysis in urban scenes. Computers and Graphics, 23 (5):655–666, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
30/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/30/abstract.
html.

Nguyen:2024:DIV

[NFW^{+}24] Anh Nguyen, Michael Francis, Emma Windfeld, Guillaume Lhermie, and Kangsoo Kim. Developing an immersive virtual farm simulation for engaging and effective public education about the dairy industry. Computers and Graphics, 118 (??):173–183, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003059.

Nagendra:1988:OOV

[NG88] I. V. Nagendra and U. G.

Gujar. 3D objects from 2D orthographic views — a survey. Computers and Graphics, 12(1):111–114, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nettleton:1995:EF

[NG95a] David John Nettleton and Roberto Garigliano. Evolving fractals. Computers and Graphics, 19(5):779–782, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500058.

Ng:1995:AFR

[Ng95b] Adelene Ng. Assessment of five radiosity acceleration techniques. Computers and Graphics, 19(5):727–738, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500054.

Nikiel:2003:GVE

[NG03] Slawek Nikiel and Adam Goinski. Generation of volumetric escape time fractals. Computers and Graphics, 27 (6):977–982, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nikiel:2005:RSS

[NG05] Slawomir Nikiel and Adam Goinski. A recursive subdivision scheme for isosurface construction. Computers and Graphics, 29 (1):155–164, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ng:1995:SMV

[NGA95] H. N. Ng, R. L. Grimsdale, and W. G. Allen. A system for modelling and visualization of cloth material. Computers and Graphics, 19(3): 423–430, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500012.

Nunes:2023:SRU

[NGAS23] Francisco Nunes, Petra Rato Grego, Ricardo Araújo, and Paula Alexandra Silva. Selfreport user interfaces for patients with rheumatic and musculoskeletal diseases: App review and usability experiments with mobile user interface components. Computers and Graphics, 117 (??):61–72, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002480.

Nahas:1983:SOI

[NH83] M. Nahas and H. Huitric.

Scenes of our imagination. Computers and Graphics, 7 (2):205–207, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nonnemann:2022:DDP

[NHR^{+}22] Lars Nonnemann, Marius Hogräfer, Martin Röhlig, Heidrun Schumann, Bodo Urban, and Hans-Jörg Schulz. A data-driven platform for the coordination of independent visual analytics tools. Computers and Graphics, 106(??):152–160, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001017.

Nickel:1984:SGI

[Nic84]  Randy Nickel.  Silicon

Graphics IRIS system. Computers and Graphics, 8(4): 433–435, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nagayasu:2008:DPA

[NIH08] Daisuke Nagayasu, Fumihiko Ino, and Kenichi Hagihara. A decompression pipeline for accelerating outof-core volume rendering of time-varying data. Computers and Graphics, 32(3): 350–362, June 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000502.

Nijholt:2004:WCD

[Nij04] Anton Nijholt. Where computers disappear, virtual humans appear. Computers and Graphics, 28(4):467– 476, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nikiel:1998:CGT

[Nik98] Slawomir S. Nikiel. Chaos and graphics — true-colour images and iterated function systems. Computers and Graphics, 22(5):635–640, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        5/606.pdf.

Nikiel:2006:IIF

[Nik06] Slawomir Nikiel. Integration of iterated function systems and vector graphics for aesthetics. Computers and Graphics, 30(2): 277–283, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000033.

Netto:2001:EAL

[NK01] Márcio Lobo Netto and João Eduardo Kögler, Jr. Editorial: Artificial life: towards new generation of computer animation. Computers and Graphics, 25(6): 929–931, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/35/28/abstract.
html.

Nakazawa:1983:CSM

[NKA83] M. Nakazawa, K. Kiruchi, and K. Arai. CAE system in mechanical engineering. Computers and Graphics, 7(3–4):315–325, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

NGuyen:2023:EBS

[NKMI23] Franck N’Guyen, Toufik Kanit, F. Maisonneuve, and Abdellatif Imad. Efficient boundary surface reconstruction from multilabel volumetric data with mathematical morphology. Computers and Graphics, 117(??):192–208, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323002406.

Nakartsuyama:1983:CGI

[NKNN83] M. Nakartsuyama, K. Kanno, H. Nagahashi, and N. Nishizuka. Curve generation of implicit functions by incremental computers. Computers and Graphics, 7(2): 161–167, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nascimento:2023:SRL

[NLdAL^{+}23] Valfride Nascimento, Rayson Laroca, Jorge de A. Lambert, William Robson Schwartz, and David Menotti. Superresolution of license plate images using attention modules and sub-pixel convolution layers. Computers and Graphics, 113(??): 69–76, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000602.

Naert:2020:SAS

[NLG20] Lucie Naert, Caroline Larboulette, and Sylvie Gibet.

A survey on the animation of signing avatars: From sign representation to utterance synthesis. Computers and Graphics, 92 (??):76–98, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301370.

Nunes:2023:SRO

[NLOdS23] Ian Nunes, Camila Laranjeira, Hugo Oliveira, and Jefersson A. dos Santos. A systematic review on openset segmentation. Computers and Graphics, 115 (??):296–308, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001218.

Neperud:2007:VAW

[NLS07] Bryan Neperud, John Lowther, and Ching-Kuang Shene. Visualizing and animating the winged-edge data structure. Computers and Graphics, 31 (6):877–886, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001653.

Ning:2011:CDS

[NLSN11] Xie Ning, Hamid Laga, Suguru Saito, and Masayuki Nakajima. Contour-driven Sumi-e rendering of real photos. Computers and Graphics, 35(1):122–134, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000186X.

Newman:1985:CAV

[NM85] H. Newman and Z. Muscati. Coding aspects of the virtual device metafile and other picture standards. Computers and Graphics, 9 (1):39–45, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nam:2009:HDF

[NMM09] Julia EunJu Nam, Mauricio Maurer, and Klaus Mueller. A high-dimensional feature clustering approach to support knowledge-assisted visualization. Computers and Graphics, 33(5):607– 615, October 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000910.

Nowrouzezahrai:2015:FSS

[NN15] Derek Nowrouzezahrai and Diego Nehab. Forward to the special section on SIB-GRAPI 2014. Computers and Graphics, 53 (part B)(??):A3, December 2015.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001697.

Nieselt:2022:F

[NOJS^{+}22] Kai Nieselt, Steffen Oeltze-Jafra, Thomas Schultz, Noeska Smit, and Bjorn Sommer. Foreword. Computers and Graphics, 108 (??):A4–A5, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001893.

Noll:1995:BCA

[Nol95] A. Michael Noll. The beginnings of computer art in the United States: a memoir. Computers and Graphics, 19(4):495–503, July–August 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500026.

Norton:1989:JSQ

[Nor89] Alan Norton. Julia sets in the quaternions. Computers and Graphics, 13(2): 267–278, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nagai:2015:TSR

[NOS15] Yukie Nagai, Yutaka Ohtake, and Hiromasa Suzuki. Tomographic surface reconstruction from point cloud. Computers and Graphics, 46 (??):55–63, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001228.

Novins:2003:FGR

[Nov03] Kevin L. Novins. Filling the gaps: a research-driven approach to learning computer graphics. Computers and Graphics, 27(3):347– 351, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nowacki:1986:MN

[Now86]  H. Nowacki.  Modeling

in networks. Computers and Graphics, 10(2): 113–117, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Noll:1988:PGI

[NP88] S. Noll and J. Poller. PHI-GKS: An implementation to support GKS-3D and PHIGS. Computers and Graphics, 12(2):163–172, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nahed:1996:DHM

[NP96] A. Nahed and B. Peroche. A 21/2-D hierarchical model based on planar maps. Computers and Graphics, 20(1):115–123, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500071.

Nakatsu:2012:P

[NP12] Ryhohei Nakatsu and Zhigeng Pan. Preface. Computers and Graphics, 36(3): iii–iv, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000179.

Ning:2007:CMT

[NR07a] Chen Ning and Clifford A.

Reiter. A cellular model for three-dimensional snow crystallization. Computers and Graphics, 31(4):668– 677, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000787.

Ning:2007:GBD

[NR07b] Chen Ning and Clifford A.

Reiter. Generalized Binet dynamics. Computers and Graphics, 31(2): 301–307, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001968.

Noser:1995:NDA

[NRTT95] Hansrudi Noser, Olivier Renault, Daniel Thalmann, and Nadia Magnenat Thalmann. Navigation for digital actors based on synthetic vision, memory, and learning. Computers and Graphics, 19(1):7–19, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400117.

Nair:1987:AGM

[NS87] K. N. Ramachandran Nair and R. Sankar. An approach to geometric modeling of solids bounded by sculptured surfaces. Computers and Graphics, 11(2): 113–120, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Naef:2005:MIB

[NSG05] Martin Naef, Oliver Staadt, and Markus Gross. Multimedia integration into the blue-c API. Computers and Graphics, 29(1):3–15, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nah:2016:BAB

[NSL16] Jae-Ho Nah, Youngsun Suh, and Yeongkyu Lim. L-Bench: an Android benchmark set for low-power mobile GPUs. Computers and Graphics, 61(??):40– 49, December 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630111X.

Novello:2022:EDG

[NSS^{+}22] Tiago Novello, Guilherme Schardong, Luiz Schirmer, Vinícius da Silva, Hélio Lopes, and Luiz Velho. Exploring differential geometry in neural implicits. Computers and Graphics, 108 (??):49–60, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001649.

Nakamae:1995:PCG

[NT95] Eihachiro Nakamae and Katsumi Tadamura. Photorealism in computer graphics — past and present. Computers and Graphics, 19(1):119–130, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400127.

Newman:2000:AEV

[NT00] Timothy S. Newman and Ning Tang. Approaches that exploit vector-parallelism for three rendering and volume visualization techniques. Computers and Graphics, 24 (5):755–774, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/35/article.
pdf.

Noh:2020:SRR

[NTAI20] Seung-Tak Noh, Kenichi Takahashi, Masahiko Adachi, and Takeo Igarashi. Shape refinement and rigging of raw-scanned 3D volume by a user-specified skeleton. Computers and Graphics, 87(??):80–88, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030011X.

Nedel:2018:FSS

[NTB18] Luciana Nedel, Verônica Teichrieb, and Mark Billinghurst. Foreword to the special section on SVR 2018. Computers and Graphics, 77 (??):A7–A8, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301754.

Nunes:2021:FSS

[NTTF21] Fatima Nunes, Indira Thouvenin, João Marcelo Teixeira, and Pablo Figueroa. Foreword to the special section on the Symposium on Virtual and Augmented Reality 2020 (SVR 2020). Computers and Graphics, 94 (??):A3–A4, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301771.

Nugent:1991:NPP

[Nug91] Jim Nugent. Notes on Pascal’s pyramid for personal computer users. Computers and Graphics, 15(2): 303–311, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nicolas:2024:SEH

[NUM24] Wagner Nicolas, Schwanecke Ulrich, and Botsch

Mario.  SparseSoftDECA

— efficient high-resolution physics-based facial animation from sparse landmarks. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000360.

Newman:2006:SMC

[NY06] Timothy S. Newman and Hong Yi. A survey of the marching cubes algorithm. Computers and Graphics, 30 (5):854–879, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001336.

Nakamae:1983:CCG

[NYKN83] Eihachiro Nakamae, Hideo Yamashita, Nobuyuki Kawano, and Shinji Nakano. Color computer graphics in magnetic field analysis by means of the finite element method. Computers and Graphics, 7 (3–4):295–306, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Nie:2021:ERV

[NZL^{+}21] Jianhui Nie, Zhaochen Zhang, Ye Liu, Hao Gao, Feng Xu, and Wenkai Shi. Enhancement of ridge-valley features in point cloud based on position and normal guidance. Computers and Graphics, 99 (??):212–223, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001345.

Ni:2021:BFD

[NZZ^{+}21] Jiancheng Ni, Susu Zhang, Zili Zhou, Lijun Hou, Jie Hou, and Feng Gao. Background and foreground disentangled generative adversarial network for scene image synthesis. Computers and Graphics, 97(??): 54–66, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000455.

Oguz:2010:ECS

[OAYG10] Oguzcan Oguz, Ates Akaydin, Türker Yilmaz, and Ugur Güdükbay. Emergency crowd simulation for outdoor environments. Computers and Graphics, 34(2): 136–144, April 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001563.

Olivier:2023:FFA

[OBD^{+}23] Nicolas Olivier, Kelian Baert, Fabien Danieau, Franck Multon, and Quentin

Avril. FaceTuneGAN: Face

autoencoder for convolutional expression transfer using neural generative adversarial networks. Computers and Graphics, 110 (??):69–85, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002278.

Ollila:2000:BAC

[OC00] Mark Ollila and Eva Carling. Bringing art into computer graphics education. Computers and Graphics, 24 (4):617–622, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/38/article.
pdf.

Oti:2021:IST

[OC21] Alfred Oti and Nathan Crilly. Immersive 3D sketching tools: Implications for visual thinking and communication. Computers and Graphics, 94(??):111– 123, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301667.

Ouyang:2012:VHP

[OCCZ12] Peichang Ouyang, Dongsheng Cheng, Yanhua Cao, and Xiaogen Zhan. The visualization of hyperbolic patterns from invariant mapping method. Computers and Graphics, 36(2): 92–100, April 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001725.

Olivier:2019:NEM

[OCR^{+}19] Pauline Olivier, Renaud Chabrier, Damien Rohmer, Eric de Thoisy, and Marie-Paule Cani. Nested explorative maps: a new 3D canvas for conceptual design in architecture. Computers and Graphics, 82 (??):203–213, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300895.

Ortega:2002:CGP

[OdlCA02] Alfonso Ortega, Marina de la Cruz, and Manuel Alfonseca. Chaos and graphics: Parametric 2-dimensional L systems and recursive fractal images: Mandelbrot set, Julia sets and biomorphs. Computers and Graphics, 26(1):143– 149, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/40/abstract.
                                        html.

Olson:2011:PSS

[ODZS11] Matt Olson, Ramsay Dyer, Hao Zhang, and Alla Sheffer. Point set silhouettes via local reconstruction. Computers and Graphics, 35(3): 500–509, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000732.

Osorio:2011:NAD

[OFP^{+}11] Maria Fernanda Osorio, Pablo Figueroa, Flavio Prieto, Pierre Boulanger, and Eduardo Londoño. A novel approach to documenting artifacts at the Gold Museum in Bogota. Computers and Graphics, 35 (4):894–903, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000239.

Ortiz-Garcia:2007:AGV

[OGSSLM^{+}07] Emilio G. Ortiz-García, Sancho Salcedo-Sanz, Jose M. Leiva-Murillo, Angel M. Pérez-Bellido, and José A. Portilla-Figueras. Automated generation and visualization of picture-logic puzzles. Computers and Graphics, 31(5):750–760, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700163X.

Ohara:1983:ICC

[OH83] Masashi Ohara and Masatake Higashi. Integration of CAD/CAM systems in automotive body engineering. Computers and Graphics, 7(3–4):307–314, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Oikarinen:1998:TSU

[Oik98] Jarkko Oikarinen. Technical section — using 2- and 2(1/2)-dimensional seed filling in view lattice to accelerate volumetric rendering. Computers and Graphics, 22(6):745–757, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/629.pdf.

Oh:1999:TSS

[OK99] Beom-Soo Oh and Chang-Hun Kim. Technical section — systematic reconstruction of 3D curvilinear objects from two-view drawings. Computers and Graphics, 23(3):343–352, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        3/686.pdf.

Osinga:2002:VSC

[OK02] Hinke M. Osinga and Bernd Krauskopf. Visualizing the structure of chaos in the Lorenz system. Computers and Graphics, 26(5): 815–823, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/56/44/abstract.
                                        html.

Oh:2003:PRO

[OK03] Beom-Soo Oh and Chang-Hun Kim. Progressive reconstruction of 3D objects from a single freehand line drawing. Computers and Graphics, 27 (4):581–592, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Orbay:2012:SBS

[OK12] Günay Orbay and Levent Burak Kara. Sketchbased surface design using malleable curve networks. Computers and Graphics, 36 (8):916–929, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001446.

Osato:2020:COS

[OK20] Yui Osato and Naoya Koizumi. Compact optical system displaying midair images movable in depth by rotating light source and mirror. Computers and Graphics, 91(??):290– 300, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301242.

Ozturk:2008:LAB

[OKBG08] Aydin Ozturk, Murat Kurt, Ahmet Bilgili, and Cengiz Gungor. Linear approximation of Bidirectional Reflectance Distribution Functions. Computers and Graphics, 32(2): 149–158, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000071.

Okino:1983:TGP

[OKK83] N. Okino, Y. Kakazu, and H. Kubo. Theories for graphics processors in TIPS-

1. Computers and Graphics, 7(3–4):243–258, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ohbuchi:2001:BSU

[OKT01] Ryutarou Ohbuchi, Yoshiyuki Kokojima, and Shigeo Takahashi. Blending shapes by using subdivision surfaces. Computers and Graphics, 25 (1):41–58, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/30/article.
pdf.

Ong:1996:SRW

[OL96] Hwee-Leng Ong and Hing-Yan Lee. Software report: WINVIZ — a visual data analysis tool. Computers and Graphics, 20(1):83– 84, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9688777.

Oliveira:2008:CGB

[Oli08] Manuel M. Oliveira. Computer graphics in Brazil: a selection of papers from SIBGRAPI 2006. Computers and Graphics, 32(5): 525, October 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930800099X.

Ouyang:1996:PSB

[OM96] S. Ouyang and D. E.

Maynard. Phong shading by binary interpolation. Computers and Graphics, 20(6):839–848, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600055.

Owen:1993:DHT

[OMF93] G. Scott Owen, J. Morgan Morris, and Martin D. Fraser. The development of a hypermedia training system for a water treatment plant. Computers and Graphics, 17(3):243– 249, May–June 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Oprea:2019:VRG

[OMGGG^{+}19] Sergiu Oprea, Pablo Martinez-Gonzalez, Alberto Garcia-Garcia, John A. Castro-Vargas, Sergio Orts-Escolano, and Jose Garcia-Rodriguez. A visually realistic grasping system for object manipulation and interaction in virtual reality environments. Computers and Graphics, 83 (??):77–86, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301098.

Ou:2018:KDA

[OMP^{+}18] Jifei Ou, Zhao Ma, Jannik Peters, Sen Dai, Nikolaos Vlavianos, and Hiroshi Ishii. KinetiX — designing auxetic-inspired deformable material structures. Computers and Graphics, 75 (??):59–71, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301006.

Ohrhallinger:2013:MEL

[OMW13] Stefan Ohrhallinger, Sudhir Mudur, and Michael Wimmer. Minimizing edge length to connect sparsely sampled unstructured point sets. Computers and Graphics, 37 (6):645–658, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000903.

OGrady:2004:GGA

[OO04] M. J. O’Grady and G. M. P. O’Hare. Gulliver’s Genie: agency, mobility, adaptivity. Computers and Graphics, 28 (5):677–689, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Otaduy:2022:FSS

[OOC22] Miguel Otaduy, Lidia Ortega, and Antoni Chica. Foreword to the special section on CEIG 2021. Computers and Graphics, 102 (??):A4–A5, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001898.

Ozacar:2023:VRB

[ÖOK23] Kasim Özacar, Yasin Ortakci, and Muhammed Yusuf

Küçükkara. VRArchEducation:

Redesigning building survey process in architectural education using collaborative virtual reality. Computers and Graphics, 113 (??):1–9, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000456.

Oliva:2013:NNO

[OP13] R. Oliva and N. Pelechano.

NEOGEN: Near optimal generator of navigation meshes for 3D multi-layered environments. Computers and Graphics, 37(5):403– 412, August 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000435.

Oliva:2015:CDA

[OP15] Ramon Oliva and Nuria Pelechano. Clearance for diversity of agents’ sizes in navigation meshes. Computers and Graphics, 47(??): 48–58, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001423.

OSullivan:2019:FSS

[OP19] Carol O’Sullivan and Julien Pettré. Foreword to the special section on motion in games. Computers and Graphics, 78(??): A5, February 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301985.

Ovtcharova:1992:PFC

[OPR92] Jivka Ovtcharova, Gerhard Pahl, and Joachim Rix. A proposal for feature classification in feature-based design. Computers and Graphics, 16(2):187–195, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Olsen:2007:MCS

[OSB07] L. Olsen, F. F. Samavati, and R. H. Bartels. Multiresolution for curves and surfaces based on constraining wavelets. Computers and Graphics, 31(3): 449–462, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000519.

Osinga:2005:TDI

[Osi05] Hinke M. Osinga. Twodimensional invariant manifolds in four-dimensional dynamical systems. Computers and Graphics, 29(2):289– 297, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Olsen:2009:SBM

[OSSJ09] Luke Olsen, Faramarz F.

Samavati, Mario Costa Sousa, and Joaquim A. Jorge. Sketch-based modeling: a survey. Computers and Graphics, 33 (1):85–103, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001295.

Oliveira:2016:VAB

[OST^{+}16] Guilherme N. Oliveira, Jose L.

Sotomayor, Rafael P. Torchelsen, Cláudio T. Silva, and João L. D. Comba. Visual analysis of bike-sharing systems. Computers and Graphics, 60(??):119–129, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300991.

Owen:2000:DDL

[OSZ00] G. Scott Owen, Raj Sunderraman, and Yanqing Zhang. The development of a digital library to support the teaching of computer graphics and visualization. Computers and Graphics, 24 (4):623–627, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/39/article.
pdf.

Occena:1988:GCB

[OT88] Luis G. Occena and Jose M. A. Tanchoco. GSS — a CAD-based graphic sawing simulator for hardwood logs. Computers and Graphics, 12(3–4):565–578, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ongun:2021:PLL

[ÖT21] Cihan Öngün and Alptekin Temizel. LPMNet: Latent part modification and generation for 3D point clouds. Computers and Graphics, 96 (??):1–13, May 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000285.

Oku:2018:EFM

[OUZS18] Hiromasa Oku, Takahiro Uji, Yiting Zhang, and Kumi Shibahara. Edible fiducial marker made of edible retroreflector. Computers and Graphics, 77(??): 156–165, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301572.

Ochmann:2016:ARP

[OVWK16] Sebastian Ochmann, Richard Vock, Raoul Wessel, and Reinhard Klein. Automatic reconstruction of parametric building models from indoor point clouds. Computers and Graphics, 54 (??):94–103, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001119.

Owen:1994:TIS

[Owe94] G. Scott Owen. Teaching image synthesis as a physical science. Computers and Graphics, 18(3): 305–308, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Poulin:1991:SSL

[PA91] Pierre Poulin and John Amanatides. Shading and shadowing with linear light sources. Computers and Graphics, 15(2):259–265, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Payan:2007:TWB

[PA07] Frédéric Payan and Marc Antonini. Temporal waveletbased compression for 3D animated models. Computers and Graphics, 31 (1):77–88, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001774.

Perez-Aixendri:2023:MYL

[PACSG^{+}23] M. Pérez-Aixendri, P. Casanova-Salas, J. Gimeno, C. Portalés, and M. Fernández. Model of yarn-level visualization of historical fabrics based on the use of levels of detail in Sphere Tracing. Computers and Graphics, 115(??):321–330, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001565.

Peterson:2009:OSA

[PAE09] Stephen D. Peterson, Magnus Axholt, and Stephen R. Ellis. Objective and subjective assessment of stereoscopically separated labels in augmented reality. Computers and Graphics, 33 (1):23–33, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001507.

Pfeuffer:2021:PAD

[PAE^{+}21] Ken Pfeuffer, Yasmeen Abdrabou, Augusto Esteves, Radiah Rivu, Yomna Abdelrahman, Stefanie Meitner, Amr Saadi, and Florian Alt. ARtention: a design space for gaze-adaptive user interfaces in augmented reality. Computers and Graphics, 95(??):1–12, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000017.

Peltier:2006:CHG

[PAFL06] Samuel Peltier, Sylvie Alayrangues, Laurent Fuchs, and Jacques-Olivier Lachaud. Computation of homology groups and generators. Computers and Graphics, 30(1):62– 69, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002104.

Pan:2019:TIR

[PAJ19] Xian Pan, Victor Arellano, and Adrian Jarabo. Transient instant radiosity for efficient time-resolved global illumination. Computers and Graphics, 83(??):107– 113, October 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931930127X.

Pan:2006:SIE

[Pan06] Zhigeng Pan. Special issue on edutainment (Elearning and game). Computers and Graphics, 30(1): 1–2, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849305002001.

Paquette:2005:CGE

[Paq05] Eric Paquette. Computer graphics education in different curricula: analysis and proposal for courses. Computers and Graphics, 29(2):245–255, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parke:1975:MTD

[Par75a] F. I. Parke. Measuring threedimensional surfaces with a two-dimensional data tablet. Computers and Graphics, 1 (1):5–8, May 1975. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parke:1975:MHF

[Par75b]  F. I. Parke.  A model

for human faces that allows speech synchronized animation. Computers and Graphics, 1(1):3–4, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parker:1986:OOO

[Par86] H. Houston Parker. Overview of the Omni 1000 and Omni 1200 graphics display controllers. Computers and Graphics, 10(4): 369–370, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parker:1988:EVR

[Par88] J. R. Parker. Extracting vectors from raster images. Computers and Graphics, 12 (1):75–79, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parida:1993:CTG

[Par93] Laxmi Parida. A computational technique for general shape deformations for use in font design. Computers and Graphics, 17(4): 349–356, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Patane:2015:VHK

[Pat15] Giuseppe Patané. Volumetric heat kernel: Padé– Chebyshev approximation, convergence, and computation. Computers and Graphics, 46(??):64–71, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S009784931400096X.

Paukowitsch:1988:FIC

[Pau88] Peter Paukowitsch. Fundamental ideas for computersupported descriptive geometry. Computers and Graphics, 12(1):3–14, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pratt:1996:SPQ

[PB96] Paul Pratt and Martin Berzins. Shock preserving quadratic interpolation for visualization on triangular meshes. Computers and Graphics, 20(5):723–730, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600046.

Peng:2010:F

[PB10] Qunsheng Peng and Hujun Bao. Foreword. Computers and Graphics, 34(5): 528, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001135.

Preim:2011:SSV

[PB11] Bernhard Preim and Charl P.

Botha. Special section on visual computing in biology and medicine. Computers and Graphics, 35(2): iv–v, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000185.

Paiva:2022:FSS

[PB22] Afonso Paiva and Gladimir Baranoski. Foreword to the special section on SIBGRAPI 2021. Computers and Graphics, 103 (??):A10–A11, April 2022.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000255.

Paquette:2023:FSS

[PB23] Eric Paquette and Jean-Jacques Bourdin. Foreword to the special section on best papers of the Eurographics 2022 Education Papers Program. Computers and Graphics, 112(??): A6–A7, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000407.

Pintus:2021:FSS

[PBB21] Ruggero Pintus, Silvia Biasotti, and Stefano Berretti. Foreword to the special section on Smart Tool and Applications for Graphics (STAG 2020). Computers and Graphics, 101(??): 3–4, December 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001667.

Patterson:2014:HCF

[PBG^{+}14] Robert E. Patterson, Leslie M. Blaha, Georges G. Grinstein, Kristen K. Liggett, David E. Kaveney, Kathleen C. Sheldon, Paul R. Havig, and Jason A. Moore. A human cognition framework for information visualization. Computers and Graphics, 42 (??):42–58, August 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000387.

Padia:2019:SGS

[PBH19] Kalpesh Padia, Kaveen Herath Bandara, and Christopher G. Healey. A system for generating storyline visualizations using hierarchical task network planning. Computers and Graphics, 78 (??):64–75, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301821.

Park:2013:MEI

[PBK13] Hanwook Park, Haewon Byun, and Changhun Kim. Multi-exemplar inhomogeneous texture synthesis. Computers and Graphics, 37(1–2):54–64, February/ April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001616.

Payten:1997:OSF

[PBN97] W. M. Payten and B. Ben-Nissan. Optimal structure formation using a chaotic self- organisational algorithm. Computers and Graphics, 21(5):685–688, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700044.

Pujol:2023:AAS

[PC23] Eduard Pujol and Antonio Chica. Adaptive approximation of signed distance fields through piecewise continuous interpolation. Computers and Graphics, 114 (??):337–346, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001139.

Potenziani:2015:HHO

[PCD^{+}15] Marco Potenziani, Marco Callieri, Matteo Dellepiane, Massimiliano Corsini, Federico Ponchio, and Roberto Scopigno. 3DHOP: 3D Heritage Online Presenter. Computers and Graphics, 52(??):129–141, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849315001041.

Pingi:2015:FSA

[PCGS15] Paolo Pingi, Massimiliano Corsini, Fabio Ganovelli, and Roberto Scopigno. Fast and simple automatic alignment of large sets of range maps. Computers and Graphics, 47(??): 78–88, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001447.

Polasek:2023:VUL

[PCKB23] Tomas Polasek, Martin Cadík, Yosi Keller, and Bedrich Benes. Vision UFormer: Long-range monocular absolute depth estimation. Computers and Graphics, 111(??):180–189, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000262.

Prakash:2015:NAB

[PCPW15] Mahesh Prakash, Paul W.

Cleary, Soon Hyoung Pyo, and Fletcher Woolard. A new approach to boiling simulation using a discrete particle based method. Computers and Graphics, 53 (part B)(??):118–126, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001442.

Pina:2000:CAA

[PCS00] Alfredo Pina, Eva Cerezo, and Francisco J. Serón. Computer animation: from avatars to unrestricted autonomous actors (A survey on replication and modelling mechanisms). Computers and Graphics, 24(2): 297–311, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/39/article.
                                        pdf.

Plumed:2016:DMS

[PCV16] Raquel Plumed, Pedro Company, and Peter A. C. Varley. Detecting mirror symmetry in single-view wireframe sketches of polyhedral shapes. Computers and Graphics, 59(??): 1–12, October 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300565.

Pradelle:2023:LIF

[PCWD23] Olivier Pradelle, Raphaëlle Chaine, David Wendland, and Julie Digne. Lightweight integration of 3D features to improve 2D image segmentation. Computers and Graphics, 114(??):326– 336, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000936.

Pan:2006:VRM

[PCY^{+}06] Zhigeng Pan, Adrian David Cheok, Hongwei Yang, Jiejie Zhu, and Jiaoying Shi. Virtual reality and mixed reality for virtual learning environments. Computers and Graphics, 30(1):20– 28, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002025.

Paiva:2006:RVS

[PdFS06] Afonso Paiva, Luiz Henrique de Figueiredo, and Jorge Stolfi. Robust visualization of strange attractors using affine arithmetic. Computers and Graphics, 30(6): 1020–1026, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001543.

Peters:2016:FSS

[PDK16] Christopher Peters, Michael Doggett, and Lars Kjelldahl. Foreword to special section on SIGGRAD 2015. Computers and Graphics, 57(??):A1–A2, June 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300322.

Piao:2021:PMA

[PDL^{+}21] Haiyin Piao, Pengyuan Du, Qi Liu, Letian Yu, Yang Sun, Yuxin Wang, and Xin Yang. MBKD: Acceleration structure designed for moving primitives. Computers and Graphics, 98(??):126– 137, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000807.

Perez-Diaz:2021:CBA

[PDS21] Sonia Pérez-Díaz and Li-Yong Shen. Computing the μ-bases of algebraic monoid curves and surfaces. Computers and Graphics, 97(??): 78–87, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000534.

Pessoa:2012:RSR

[PdSMdML^{+}12] Saulo A. Pessoa, Guilherme de S. Moura, Joao Paulo S. do M. Lima, Veronica Teichrieb, and Judith Kelner. RPR-SORS: Realtime photorealistic rendering of synthetic objects into real scenes. Computers and Graphics, 36(2): 50–69, April 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001701.

Pocco:2022:ESL

[PdSP^{+}22] Ximena Pocco, Tiago da Silva, Jorge Poco, Luis Gustavo Nonato, and Erick Gomez-Nieto. Exploring scientific literature by textual and image content using DRIFT. Computers and Graphics, 103(??):140–152, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000218.

Pe:2004:XF

[Pe04] Joseph L. Pe. The 3 x +1 fractal. Computers and Graphics, 28(3):431– 435, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pfluger:2016:STV

[PE16] H. Pflüger and T. Ertl.

Sifting through visual arts collections. Computers and Graphics, 57(??):127– 138, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931630022X.

Pearce:2002:EAN

[Pea02] Celia Pearce. Emergent authorship: the next interactive revolution. Computers and Graphics, 26(1):21– 29, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/30/abstract.
                                        html.

Ping:2022:VES

[PECW22] Guiju Ping, Mahdi Abolfazli Esfahani, Jiaying Chen, and Han Wang. Visual enhancement of single-view 3D point cloud reconstruction. Computers and Graphics, 102 (??):112–119, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000012.

Perrill:1977:PPC

[Per77] Walter A. Perrill. Packaging printed circuit boards with interactive graphics. Computers and Graphics, 2 (2):39–50, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Perez:1984:CGA

[Per84] R. M. V. Perez. Computer graphics as an aid to a robot dynamical simulation analysis. Computers and Graphics, 8(3):265–268, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Perlin:2002:BAC

[Per02] Ken Perlin. Better acting in computer games: the use of procedural methods. Computers and Graphics, 26(1):3–11, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/27/28/abstract.
html.

Peters:2018:GSB

[Pet18] Jörg Peters. On G^{1} stitched bi-cubic Bézier patches with arbitrary topology. Computers and Graphics, 71 (??):154–156, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301966.

See response [ASC18].

Pottmann:2015:AG

[PEVW15] Helmut Pottmann, Michael Eigensatz, Amir Vaxman, and Johannes Wallner. Architectural geometry. Computers and Graphics, 47 (??):145–164, April 2015.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400140X.

Potmesil:1980:ITH

[PF80] M. Potmesil and H. Freeman. Implementation of two hidden-line algorithms. Computers and Graphics, 5 (1):31–40, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pintado:1989:GFD

[PF89] Xavier Pintado and Eugene Fiume. GRAFIELDS: Fielddirected dynamic splines for interactive motion control. Computers and Graphics, 13 (1):77–82, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Park:1997:FDB

[PF97] Jihun Park and Donald S. Fussell. Forward dynamics based realistic animation of rigid bodies. Computers and Graphics, 21(4): 483–496, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700024.

Patane:2009:CSA

[PF09] Giuseppe Patanè and Bianca Falcidieno. Computing smooth approximations of scalar functions with constraints. Computers and Graphics, 33(3):399–413, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000284.

Pelechano:2016:HPF

[PF16] Nuria Pelechano and Carlos Fuentes. Hierarchical path-finding for navigation meshes (HNA^{∗}). Computers and Graphics, 59(??): 68–78, October 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300668.

Pfaff:1983:CVG

[Pfa83] Günther E. Pfaff. Certification/validation of graphics software. Computers and Graphics, 7(2):203–204, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pfaff:1984:FCT

[Pfa84] Günther E. Pfaff. Functional conformance testing of graphics software using a configurable reference system. Computers and Graphics, 8(1):29–37, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Philip:1994:WMM

[PFR94] A. G. Davis Philip, Michael Frame, and Adam Robucci. Warped midgets in the Mandelbrot set. Computers and Graphics, 18(2):239– 248, March–April 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pascual:1986:IPC

[PGB86] J. Pascual, J. Giralt, and P. Brunet. An interactive package for the computeraided design of woven fabrics. Computers and Graphics, 10(4):359–368, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Peng:2023:MMU

[PGL^{+}23] Hao-Yang Peng, Meng-Hao Guo, Zheng-Ning Liu, Yong-Liang Yang, and Tai-Jiang Mu. MWFormer: Mesh understanding with windowbased transformer. Computers and Graphics, 115 (??):382–391, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001589.

Patnaik:1983:IPR

[PGR83] L. M. Patnaik, B. A. Gadkari, and L. Ramakrishnan. Implementation of placement and routing algorithms for computer aided design of printed circuit boards. Computers and Graphics, 7 (3–4):333–347, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Portales:2023:MRA

[PGS^{+}23] Cristina Portalés, Jesús Gimeno, Antonio Salvador, Alfonso García-Fadrique, and Sergio Casas-Yrurzum. Mixed reality annotation of robotic-assisted surgery videos with real- time tracking and stereo matching. Computers and Graphics, 110(??):125–140, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322002291.

Pla-Garcia:2006:SPB

[PGVACN06] N. Pla-Garcia, M. Vigo-Anglada, and J. Cotrina-Navau. N-sided patches with B-spline boundaries. Computers and Graphics, 30(6): 959–970, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001002.

Popsel:1990:HSL

[PH90] J. Pöpsel and Ch. Hornung. Highlight shading: Lighting and shading in a PHIGS+/ PEX-environment. Computers and Graphics, 14 (1):55–64, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pham:1989:QBS

[Pha89] Binh Pham. Quadratic Bsplines for automatic curve and surface fitting. Computers and Graphics, 13(4): 471–475, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pharasi:1995:CCL

[Pha95]  B. Pharasi.  Connected

components labelling using Murray polygons. Computers and Graphics, 19(3): 405–411, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500010.

Philip:1992:FLM

[Phi92] Kenelm W. Philip. Field lines in the Mandelbrot set. Computers and Graphics, 16 (4):443–447, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Peng:2015:HBB

[PHLW15] Shu-Juan Peng, Gao-Feng He, Xin Liu, and Hua-Zhen Wang. Hierarchical block-based incomplete human mocap data recovery using adaptive nonnegative matrix factorization. Computers and Graphics, 49(??): 10–23, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000436.

Patotskaya:2023:AVH

[PHO^{+}23] Yuliya Patotskaya, Ludovic Hoyet, Anne-Hélène Olivier, Julien Pettré, and Katja Zibrek. Avoiding virtual humans in a constrained environment: Exploration of novel behavioural measures. Computers and Graphics, 110(??):162–172, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300002X.

Pickover:1984:UCD

[Pic84] C. A. Pickover. The use of computer-drawn faces as an educational aid for the presentation of statistical concepts. Computers and Graphics, 8(2):163–166, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1987:MBT

[Pic87a] Clifford A. Pickover. Mathematics and beauty: Timediscrete phase planes associated with the cyclic system, ẋ(t) = −f(y(t)), ẏ(t) = f(x(t)). Computers and Graphics, 11(2):217–226, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1987:MBV

[Pic87b] Clifford A. Pickover. Mathematics and beauty V: Turbulent complex curls. Computers and Graphics, 11(4): 499–508, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1988:NRS

[Pic88a] Clifford A. Pickover. A note on rendering 3D strangeattractors. Computers and Graphics, 12(2):263–267, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1988:RST

[Pic88b] Clifford A. Pickover. Rendering of the Shroud of Turin using sinusoidal pseudocolor and other image processing techniques. Computers and Graphics, 12(1):81–90, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1989:CKN

[Pic89] Clifford A. Pickover. Circles with kiss: a note on osculatory packing. Computers and Graphics, 13(1):63–67, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1992:IAE

[Pic92a] Clifford A. Pickover. Introduction by the Associate Editor. Computers and Graphics, 16(1):3–??, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1992:VMA

[Pic92b] Clifford A. Pickover. A vacation on Mars — an artist’s journey in a computer graphics world. Computers and Graphics, 16(1):9–13, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1994:APG

[Pic94] Clifford A. Pickover. Automatic parallel generation of aeolian fractals on the IBM Power Visualization System. Computers and Graphics, 18(3):407–416, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1995:CFB

[Pic95a] Clifford A. Pickover. The crying of fractal Batrachion 1,489. Computers and Graphics, 19(4):611– 615, July–August 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500039.

Pickover:1995:LC

[Pic95b] Clifford A. Pickover. The loom of creation. Computers and Graphics, 19(4): 523–527, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500030.

Piegl:1983:FAP

[Pie83] L. Piegl. A fast algorithm for parametric curve plotting. Computers and Graphics, 7(3–4):361–367, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piegl:1984:DCC

[Pie84]  Leslie Piegl.  Defining

C^{1} curves containing conic segments. Computers and Graphics, 8(2):177–182, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piegl:1988:CTP

[Pie88]  L. Piegl.  Coons-type

patches. Computers and Graphics, 12(2):221–228, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piegl:1989:NEU

[Pie89] L. Piegl. Negative experiment with univariate blending functions. Computers and Graphics, 13(2): 217–222, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piegl:1990:PDU

[Pie90] Les Piegl. Program development for a University course in geometric modeling. Computers and Graphics, 14(3–4):501–504, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Park:2000:SAV

[PiP00] Jinah Park and Sang il Park.

Strain analysis and visualization: left ventricle of a heart. Computers and Graphics, 24 (5):701–714, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/

31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/31/article.
pdf.

Panchal:2021:CPD

[PJ21] Dakshata Panchal and Deepak Jayaswal. Computational paradigms for direct triangular surface remeshing. Computers and Graphics, 94 (??):87–110, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301679.

Prevost:2016:LSP

[PJJSH16] Romain Prévost, Alec Jacobson, Wojciech Jarosz, and Olga Sorkine-Hornung. Large-scale painting of photographs by interactive optimization. Computers and Graphics, 55(??):108– 117, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500182X.

Pratt:2023:BLN

[PJP23] Louis Pratt, Andrew Johnston, and Nico Pietroni. Bending the light: Next generation anamorphic sculptures. Computers and Graphics, 114(??):210–218, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000778.

Pintore:2024:DSE

[PJVH^{+}24] Giovanni Pintore, Alberto Jaspe-Villanueva, Markus Hadwiger, Jens Schneider, Marco Agus, Fabio Marton, Fabio Bettio, and Enrico Gobbetti. Deep synthesis and exploration of omnidirectional stereoscopic environments from a single surround-view panoramic image. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849324000426.

Pickover:1985:CGG

[PK85] C. A. Pickover and E. Khorasani. Computer graphics generated from the iteration of algebraic transformations in the complex plane. Computers and Graphics, 9 (2):147–151, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1986:FCS

[PK86] Clifford A. Pickover and Al Khorasani. Fractal characterization of speech waveform graphs. Computers and Graphics, 10(1):51–61, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parida:1991:CFS

[PK91] Laxmi Parida and Pramod Koparkar. A closed form solution to the problem of tangential circles, lines, points with extension to 3D. Computers and Graphics, 15 (1):49–55, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Popa:2017:FSS

[PK17] Tiberiu Popa and Paul G. Kry. Foreword to the special section on Graphics Interface 2016. Computers and Graphics, 62(??): A1–A2, ???? 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301418.

Park:2003:NNA

[PKK03] ChangHoon Park, Heedong Ko, and Taiyun Kim. NAVER: Networked and Augmented Virtual Environment aRchitecture; design and implementation of VR framework for Gyeongju VR Theater. Computers and Graphics, 27(2):223– 230, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pillwein:2021:DFM

[PKRM21] Stefan Pillwein, Johanna Kübert, Florian Rist, and Przemyslaw Musialski. Design and fabrication of multi-patch elastic geodesic grid structures. Computers and Graphics, 98(??): 218–230, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001163.

Park:1997:HDT

[PL97] Sangkun Park and Kunwoo Lee. High-dimensional trivariate NURBS representation for analyzing and visualizing fluid flow data. Computers and Graphics, 21(4):473–482, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=4&aid=9700023.

Plath:2000:RMT

[Pla00] Jan Plath. Realistic modelling of textiles using interacting particle systems. Computers and Graphics, 24 (6):897–905, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/

                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/34/32/article.
                                        pdf.

Ponciano:2021:ONT

[PLFT21] Jean R. Ponciano, Claudio D. G. Linhares, Elaine R. Faria, and Bruno A. N. Travençolo. An online and nonuniform timeslicing method for network visualisation. Computers and Graphics, 97(??):170– 182, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000480.

Park:2013:FAF

[PLJ^{+}13] Min Ki Park, Seung Joo Lee, In Yeop Jang, Yong Yi Lee, and Kwan H. Lee. Feature-aware filtering for point-set surface denoising. Computers and Graphics, 37 (6):589–595, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000782.

Park:2015:ERD

[PLJL15] Min Ki Park, Yong Yi Lee, In Yeop Jang, and Kwan H. Lee. Enhancement of range data using a structure-aware filter. Computers and Graphics, 48(??):48–59, May 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500014X.

Pithioux:2005:PNP

[PLM^{+}05] M. Pithioux, O. López, U. Meier, C. Monserrat, M. C. Juan, and M. Alcañiz. ParSys: a new particle system for the introduction of on-line physical behaviour to three-dimensional synthetic objects. Computers and Graphics, 29(1):135– 144, February 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Perazzo:2023:DGB

[PLVT23] Daniel Perazzo, João Paulo Lima, Luiz Velho, and Veronica Teichrieb. DirectVoxGO++: Grid-based fast object reconstruction using radiance fields. Computers and Graphics, 114(??):96– 104, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000924.

Pasemann:1984:ERC

[PM84a] K. Pasemann and E. Metzner. Educational requirements for CAD in industry. Computers and Graphics, 8(2):211–213, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pereira:1984:DIR

[PM84b] J. A. G. Pereira and N. D. A. Mascarenhas. Digital image registration by sequential analysis. Computers and Graphics, 8(3):247–253, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pickover:1990:VCC

[PM90] Clifford A. Pickover and Kevin McCarty. Visualizing Cantor cheese construction. Computers and Graphics, 14(2):337–341, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pradhan:1991:ACD

[PM91] B. S. S. Pradhan and A. Mukhopadhyay. Adaptive cell division for ray tracing. Computers and Graphics, 15(4):549–552, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pattanaik:1993:EPE

[PM93] S. N. Pattanaik and S. P. Mudur. Efficient potential equation solutions for global illumination computation. Computers and Graphics, 17(4):387–396, July– August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prakash:1995:VRU

[PM95] C. E. Prakash and S. Manohar.

Volume rendering of unstructured grids — a voxelization approach. Computers and Graphics, 19 (5):711–726, September– October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500053.

Porter:2010:DMS

[PM10] Benjamin Porter and Jon McCormack. Developmental modelling with SDS. Computers and Graphics, 34 (4):294–303, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000750.

Pytel:2013:SOA

[PM13] Alexei Pytel and Stephen Mann. Self-organized approach to modeling hydraulic erosion features. Computers and Graphics, 37(4):280–292, June 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000083.

Preim:2020:SMA

[PM20] Bernhard Preim and Monique Meuschke. A survey of medical animations. Computers and Graphics, 90 (??):145–168, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300923.

See corrigendum [PM22a].

Preim:2022:CSM

[PM22a] Bernhard Preim and Monique Meuschke. Corrigendum to “A Survey of Medical Animations” [Comput. Graph. 90 (2020) 145–168]. Computers and Graphics, 107(??): 303, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001558.

See [PM20].

Preim:2022:SMA

[PM22b] Bernhard Preim and Monique Meuschke. A survey of medical animations. Computers and Graphics, 107 (??):304–328, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200156X.

Pham:2014:VCS

[PMBS14] Tuan Pham, Ronald Metoyer, Katerina Bezrukova, and Chester Spell. Visualization of cluster structure and separation in multivariate mixed data: a case study of diversity faultlines in work teams. Computers and Graphics, 38 (??):117–130, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001507.

Pfaff:1985:VCC

[PMK85] G. E. Pfaff, R. Muller, and B. Kirsch. A verifier for checking the conformance of programs with the GKS standard. Computers and Graphics, 9(1):19–25, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parakkat:2018:PLS

[PMM18] Amal Dev Parakkat, Subhasree Methirumangalath, and Ramanathan Muthuganapathy. Peeling the longest: a simple generalized curve reconstruction algorithm. Computers and Graphics, 74 (??):191–201, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300803.

Perez-Monte:2016:MFL

[PMPR^{+}16] Cristian F. Perez-Monte, Mauricio D. Perez, Silvio Rizzi, Fabiana Piccoli, and Cristian Luciano. Modelling frame losses in a parallel alternate frame rendering system with a computational best-effort scheme. Computers and Graphics, 60 (??):76–82, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630098X.

Patel:1987:SRM

[PMS87] Nitin R. Patel, L. Mohan, and S. Sudarshan. Statistical representation on maps

— an interactive graphics approach. Computers and Graphics, 11(2):95–100, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Park:2001:NSI

[PMTK01] Joo-Young Park, Tim McInerney, Demetri Terzopoulos, and Myoung-Hee Kim. A non-self-intersecting adaptive deformable surface for complex boundary extraction from volumetric images. Computers and Graphics, 25(3):421–440, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/33/article.
pdf.

Pernot:2006:FHM

[PMV06] Jean-Philippe Pernot, George Moraru, and Philippe Véron. Filling holes in meshes using a mechanical model to simulate the curvature variation minimization. Computers and Graphics, 30(6): 892–902, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001452.

Pan:1997:CFC

[PMZS97] Zhigeng Pan, Xiaohu Ma, Mingmin Zhang, and Jiaoying Shi. Chinese font composition method based on algebraic system of geometric shapes. Computers and Graphics, 21(3): 321–328, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700009.

Pirsch:1983:TGL

[PN83] P. Pirsch and A. N. Netravali. Transmission of gray level images by multilevel dither techniques. Computers and Graphics, 7 (1):31–44, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pretto:2009:ISF

[POBB09] P. Pretto, M. Ogier, H. H. Bülthoff, and J.-P. Bresciani. Influence of the size of the field of view on motion perception. Computers and Graphics, 33(2): 139–146, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000132.

Polhemus:1983:SGA

[Pol83] N. W. Polhemus. Statistical graphics in APL: structuring a graphical support system for interactive data analysis. Computers and Graphics, 7(1):73–80, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Popentiu:1993:IIR

[Pop93] Florin Popentiu. Iterative identification and restoration of images. Computers and Graphics, 17(4): 497–??, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Posdamer:1977:CPC

[Pos77a] Jeffrey L. Posdamer. Concurrent processing for computer graphics. Computers and Graphics, 2(4):259–263, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Posdamer:1977:SCE

[Pos77b] Jeffrey L. Posdamer. Some criteria for the evaluation of graphics implementation languages. Computers and Graphics, 2(2):91–95, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Potts:1975:CGW

[Pot75] J. Potts. Computer graphics — whence and hence. Computers and Graphics, 1(2– 3):137–156, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pottle:1977:CLI

[Pot77a] Christopher Pottle. Computer as a laboratory instrument: Experiments for a third-year laboratory in electrical engineering. Computers and Graphics, 2(3): 169–173, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Potts:1977:CGC

[Pot77b] J. Potts. Crystal gazing in computer graphics. Computers and Graphics, 2(4): 197–200, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Potts:1978:SIG

[Pot78] Jackie Potts. Share, Inc., graphics glossary. Computers and Graphics, 3(4): 123–128, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prospero:1990:PIG

[PP90] Manuel Joao Prospero and Paulo Jorge Pereira. On programming an interactive graphical application in logic. Computers and Graphics, 14(1):7–16, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prabhu:1999:CGI

[PP99] B. S. Prabhu and S. S. Pande. Computer graphics in India — intelligent interpretation of CADD drawings. Computers and Graphics, 23(1):25–44, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/648.pdf.

Pozzer:2002:PSS

[PP02] Cesar Tadeu Pozzer and Sérgio Roberto Matiello Pellegrino. Procedural solidspace techniques for modeling and animating waves. Computers and Graphics, 26 (6):877–885, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Paille:2012:CPD

[PP12] Gilles-Philippe Paillé and Pierre Poulin. As-conformal-as-possible discrete volumetric mapping. Computers and Graphics, 36(5):427– 433, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000532.

Pandey:2016:NAM

[PP16] Achala Pandey and Umesh Chandra Pati. A novel approach to multi-scale blending based on saliency mapping for multimedia image compositing applications. Computers and Graphics, 59 (??):93–106, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630070X.

Pereira:2020:PRV

[PP20] Mateus M. Pereira and Fernando V. Paulovich. RankViz: a visualization framework to assist interpretation of Learning to Rank algorithms. Computers and Graphics, 93 (??):25–38, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301515.

Pierson:2022:PBC

[PPD22] Emery Pierson, Juan-Carlos Álvarez Paiva, and Mohamed Daoudi. Projection-based classification of surfaces for 3D human mesh sequence retrieval. Computers and Graphics, 102(??):45–55, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002235.

Pintore:2018:REC

[PPG^{+}18] Giovanni Pintore, Ruggero Pintus, Fabio Ganovelli, Roberto Scopigno, and Enrico Gobbetti. Recovering 3D existing-conditions of indoor structures from spherical images. Computers and Graphics, 77 (??):16–29, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830147X.

Paultre:1991:CCA

[PPL91] Patrick Paultre, Jean Proulx, and Pierre Leger. CAL/CGI

— an application of graphics for matrix structural analysis education. Computers and Graphics, 15(1): 131–135, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Parakkat:2018:DTB

[PPM18] Amal Dev Parakkat, Uday Bondi Pundarikaksha, and Ramanathan Muthuganapathy. A Delaunay triangulation based approach for cleaning rough sketches. Computers and Graphics, 74 (??):171–181, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300761.

Paz:2023:MNM

[PPN^{+}23] Hallison Paz, Daniel Perazzo, Tiago Novello, Guilherme Schardong, Luiz Schirmer, Vinícius da Silva, Daniel Yukimura, Fabio Chagas, Hélio Lopes, and Luiz Velho. MR-Net: Multiresolution sinusoidal neural networks. Computers and Graphics, 114(??):387– 400, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000699.

Pasko:1988:GMA

[PPP88] A. A. Pasko, V. V. Pilyugin, and V. N. Pokrovskiy. Geometric modeling in the analysis of trivariate functions. Computers and Graphics, 12 (3–4):457–465, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Paccini:2020:ASA

[PPS20] Martina Paccini, Giuseppe Patané, and Michela Spagnuolo. Analysis of 3D segmented anatomical districts through grey-levels mapping. Computers and Graphics, 91(??):179–188, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301175.

Pueyo:1996:OOA

[PPSS96] Xavier Pueyo, Frederic Pérez, Toni Sellarès, and Josep Suy. An objectoriented approach for teaching visibility computation algorithms. Computers and Graphics, 20(4):605– 610, July–August 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600031.

ProsperodosSantos:2003:IAC

[PPV03] Manuel Próspero dos Santos, Xavier Pueyo, and Luiz Velho. Ibero–American co-operation in computer graphics. Computers and Graphics, 27(1):1–3, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Patow:2007:UGI

[PPV07] Gustavo Patow, Xavier Pueyo, and Alvar Vinacua. User-guided inverse reflector design. Computers and Graphics, 31(3):501– 515, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000039.

Puig:2003:STM

[PPVT03] Anna Puig, Lluís Perez-Vidal, and Dani Tost. 3D simulation of tool machining. Computers and Graphics, 27 (1):99–106, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pouli:2010:RDI

[PPZ^{+}10] Tania Pouli, Martin Prazák, Pavel Zemcík, Diego Gutierrez, and Erik Reinhard. Rendering fur directly into images. Computers and Graphics, 34(5):612–620, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000877.

Peters:2010:HMP

[PQ10] Christopher Peters and Adam Qureshi. A head movement propensity model for animating gaze shifts and blinks of virtual characters. Computers and Graphics, 34 (6):677–687, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001408.

Peng:2023:AST

[PQCT23] Haiyuan Peng, Wenhua Qian, Jinde Cao, and Shan Tang. Arbitrary style transfer based on attention and covariancematching. Computers and Graphics, 116(??):298–307, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002108.

Patnaik:1982:IIR

[PR82a] L. M. Patnaik and N. Ramesh. Implementation of an interactive relational graphics database. Computers and Graphics, 6(3):93–96, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Peirasamy:1982:AHL

[PR82b] M. Peirasamy and D. K. Ravindra. An algorithm for hidden line ruled surface display using microprocessor. Computers and Graphics, 6(4):141–157, 1982. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piegl:1993:ADS

[PR93] Les A. Piegl and Arnaud M.

Richard. Algorithm and data structure for triangulating multiply connected polygonal domains. Computers and Graphics, 17(5):563–574, September–October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pulsifer:1996:OTE

[PR96] John E. Pulsifer and Clifford A. Reiter. One tub, eight blocks, twelve blinkers and other views of life. Computers and Graphics, 20(3): 457–462, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600016.

Pai:1997:HIM

[PR97] Dinesh K. Pai and L.-M.

Reissell. Haptic interaction with multiresolution image curves. Computers and Graphics, 21(4): 405–411, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=4&aid=9700031.

Pouli:2011:PCT

[PR11] Tania Pouli and Erik Reinhard. Progressive color transfer for images of arbitrary dynamic range. Computers and Graphics, 35 (1):67–80, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931000172X.

Puchert:2023:AIA

[PR23] Patrik Puchert and Timo Ropinski. A3GC-IP: Attentionoriented adjacency adaptive recurrent graph convolutions for human pose estimation from sparse inertial measurements. Computers and Graphics, 117 (??):96–104, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002327.

Pratt:1992:NEA

[Pra92] Leonard Pratt. A note on earthworm algebra and computer graphics. Computers and Graphics, 16(3):339– 340, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prakash:1999:CGI

[Pra99] B. G. Prakash. Computer graphics in India — AUTO-LAY — a GUI-based design and development software for laminated composite components. Computers and Graphics, 23(1):95–110, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/653.pdf.

Pastor:2004:CBM

[PRÁM04] G. Pastor, M. Romera, G. Álvarez, and F. Montoya. Chaotic bands in the Mandelbrot set. Computers and Graphics, 28 (5):779–784, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Prakash:2022:DSS

[PRBD22] Siddhant Prakash, Gilles Rainer, Adrien Bousseau, and George Drettakis. Deep scene-scale material estimation from multi-view indoor captures. Computers and Graphics, 109(??):15– 29, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001789.

Preiss:1984:CSR

[Pre84a] K. Preiss. Constructing the solid representation from engineering projections. Computers and Graphics, 8(4): 381–389, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Preiss:1984:GEI

[Pre84b] Kenneth Preiss. Guest Editor’s introduction. Computers and Graphics, 8(4): 335–336, ???? 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849384900323.

Prasse:2024:ICS

[PRM^{+}24] Paul Prasse, David R. Reich, Silvia Makowski, Tobias Scheffer, and Lena A. Jäger. Improving cognitivestate analysis from eye gaze with synthetic eyemovement data. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000281.

Prospero:1985:EGI

[Pro85] M. J. Prospero. Evaluation of a graphical interaction technique using GKS. Computers and Graphics, 9 (3):319–321, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Paneels:2013:PHD

[PRRR13] Sabrina A. Panëels, Panagiotis D. Ritsos, Peter J. Rodgers, and Jonathan C. Roberts. Prototyping 3D haptic data visualizations. Computers and Graphics, 37(3):179–192, May 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000113.

Pester:2022:UMM

[PRW^{+}22] Britta Pester, Benjamin Russig, Oliver Winke, Carolin Ligges, Raimund Dachselt, and Stefan Gumhold. Understanding multi-modal brain network data: an immersive 3D visualization approach. Computers and Graphics, 106(??):88–97, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001029.

Puchtler:1986:EDC

[PS86] Wolfgang Puchtler and Juergen Schoenhut. ERLNET

— distributed components of the Erlangen Graphics System. Computers and Graphics, 10(2):143–149, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pham:1991:PAS

[PS91] Binh Pham and H. Schröder. Parallel algorithms and a systolic device for cubic B-spline curve and surface generation. Computers and Graphics, 15(3): 349–354, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Park:2002:TSP

[PS02] Sang C. Park and Hayong Shin. Technical section: Polygonal chain intersection. Computers and Graphics, 26(2):341–350, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/41/39/abstract.
                                        html.

Pan:2003:GEI

[PS03] Zhigeng Pan and Jiaoying Shi. Guest editor’s introduction. Computers and Graphics, 27(2):163– 167, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Petrik:2008:STE

[PS08] Slavomir Petrik and Vaclav Skala. Space and time efficient isosurface extraction.

Computers and Graphics, 32 (6):704–710, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001258.

Patane:2012:LAS

[PS12] Giuseppe Patanè and Michela Spagnuolo. Local approximation of scalar functions on 3D shapes and volumetric data. Computers and Graphics, 36(5):387– 397, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000507.

Patane:2013:HDK

[PS13a] Giuseppe Patané and Michela Spagnuolo. Heat diffusion kernel and distance on surface meshes and point sets. Computers and Graphics, 37 (6):676–686, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000939.

Patane:2013:IAH

[PS13b] Giuseppe Patané and Michela Spagnuolo. An interactive analysis of harmonic and diffusion equations on discrete 3D shapes. Computers and Graphics, 37 (5):526–538, August 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000460.

Peters:2015:PSS

[PS15a] Jörg Peters and Martin Sarov. Polynomial spline surfaces with rational linear transitions. Computers and Graphics, 51(??): 43–51, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000606.

Petersen:2015:CAR

[PS15b] Nils Petersen and Didier Stricker. Cognitive augmented reality. Computers and Graphics, 53 (part A) (??):82–91, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001430.

Preim:2018:SVH

[PS18] Bernhard Preim and Patrick Saalfeld. A survey of virtual human anatomy education systems. Computers and Graphics, 71(??): 132–153, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300050.

Pala:2019:ESF

[PSBD19] Pietro Pala, Lorenzo Seidenari, Stefano Berretti, and Alberto Del Bimbo. Enhanced skeleton and face 3D data for person re-identification from depth cameras. Computers and Graphics, 79(??): 69–80, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300135.

Peterlik:2010:RTV

[PSBM10] Igor Peterlík, Mert Sedef, Cagatay Basdogan, and Ludek Matyska. Real-time visio-haptic interaction with static soft tissue models having geometric and material nonlinearity. Computers and Graphics, 34 (1):43–54, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001332.

Patel:2009:KAV

[PSH^{+}09] Daniel Patel, Øyvind Sture, Helwig Hauser, Christopher Giertsen, and M. Eduard Gröller. Knowledge-assisted visualization of seismic data. Computers and Graphics, 33 (5):585–596, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000880.

Pernod:2011:MFE

[PSK^{+}11] E. Pernod, M. Sermesant, E. Konukoglu, J. Relan, H. Delingette, and N. Ayache. A multi-front eikonal model of cardiac electrophysiology for interactive simulation of radiofrequency ablation. Computers and Graphics, 35(2): 431–440, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000173.

Ponciano:2016:GBI

[PSM16] Daniel Ponciano, Marcos Seefelder, and Ricardo Marroquim. Graph-based interactive volume exploration. Computers and Graphics, 60 (??):55–65, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300802.

Padua:2014:IEP

[PSMD14] Luciana Padua, Hendrik Schulze, Kresimir Matković, and Claudio Delrieux. Interactive exploration of parameter space in data mining: Comprehending the predictive quality of large decision tree collections. Computers and Graphics, 41(??): 99–113, June 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000284.

Pueyo:2020:SCL

[PSP^{+}20] Oriol Pueyo, Albert Sabrià, Xavier Pueyo, Gustavo Patow, and Michael Wimmer. Shrinking city layouts. Computers and Graphics, 86 (??):15–26, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301736.

Prakash:1993:CGD

[PSR^{+}93] B. G. Prakash, T. G. A.

Simha, D. D. Ravindranath, K. G. Shastry, and K. Sundararaju. Computer graphics in the design and manufacture of composite laminate components. Computers and Graphics, 17(4): 407–413, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pyun:2004:EWC

[PSS04] Hyewon Pyun, Hyun Joon Shin, and Sung Yong Shin. On extracting the wire curves from multiple face models for facial animation. Computers and Graphics, 28 (5):757–765, October 2004.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Poston:1996:GDV

[PSSP96] Tim Poston, Luis Serra, Meiyappan Solaiyappan, and Ann Heng Pheng. The graphics demands of virtual medicine. Computers and Graphics, 20(1):61– 68, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500093.

Pan:1996:DGS

[PSZ96] Zhigeng Pan, Jiaoying Shi, and Mingmin Zhang. Distributed graphics support for virtual environments. Computers and Graphics, 20(2): 191–197, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500125.

Peng:2016:FMM

[PT16] Chao Peng and Sabin Timalsena. Fast mapping and morphing for genus-zero meshes with cross spherical parameterization. Computers and Graphics, 59 (??):107–118, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300681.

Pesco:2004:SAM

[PTL04] Sinésio Pesco, Geovan Tavares, and Hélio Lopes. A stratification approach for modeling two-dimensional cell complexes. Computers and Graphics, 28(2):235– 247, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pottinger:1992:PPP

[PTR^{+}92] David Pottinger, Stephen Todd, Iain Rodrigues, Tom Mullin, and Anne Skeldon. Phase portraits for parametrically excited pendula. an exercise in multidimensional data visualisation. Computers and Graphics, 16(3):331–337, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pianykh:1998:TSI

[PTW98] O. S. Pianykh, J. M. Tyler, and W. N. Waggenspack, Jr. Technical section — improved Monte Carlo form factor integration. Computers and Graphics, 22(6): 723–734, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/6/627.pdf.

Polvi:2016:SPM

[PTY^{+}16] Jarkko Polvi, Takafumi Taketomi, Goshiro Yamamoto, Arindam Dey, Christian Sandor, and Hirokazu Kato. SlidAR: a 3D positioning method for SLAM-based handheld augmented reality. Computers and Graphics, 55(??): 33–43, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001806.

Pumar:1996:ZTI

[Pum96]  M. A. Pumar.  Zoom-

ing of terrain imagery using fractal-based interpolation. Computers and Graphics, 20(1):171–176, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1&aid=9500102.

Purgathofer:1987:IHL

[Pur87a] Werner Purgathofer. Identification in a high-level graphics programming language. Computers and Graphics, 11 (1):3–10, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Purgathofer:1987:SMA

[Pur87b] Werner Purgathofer. A statistical method for adaptive stochastic sampling. Computers and Graphics, 11(2): 157–162, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Plumed:2022:EDR

[PVCM22] Raquel Plumed, Peter A. C.

Varley, Pedro Company, and Ralph Martin. Extracting datums to reconstruct CSG models from 2D engineering sketches of polyhedral shapes. Computers and Graphics, 102(??): 349–359, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002247.

Pasman:1999:AOM

[PvdSLJ99] W. Pasman, A. van der Schaaf, R. L. Lagendijk, and F. W. Jansen. Accurate overlaying for mobile augmented reality. Computers and Graphics, 23(6): 875–881, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/13/20/24/34/
                                        42/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/35/42/article.
                                        pdf.

Parreiras:2022:PCM

[PVM^{+}22] Emanuel Antônio Parreiras, Marcelo Bernardes Vieira, Arthur Gonze Machado, Marcelo Caniato Renhe, and Gilson Antônio Giraldi. A particle-in-cell method for anisotropic fluid simulation. Computers and Graphics, 102(??):220– 232, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001746.

Pardikar:1987:GIC

[PVR87] Shishir Pardikar, Mahesh Vaidya, and S. S. S. P. Rao. GKS implementation: a case study. Computers and Graphics, 11(2):75–78, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Piovarci:2018:DAD

[PWJ^{+}18] Michal Piovarči, Michael Wessely, Michał Jagielski, Marc Alexa, Wojciech Matusik, and Piotr Didyk. Design and analysis of directional front projection screens. Computers and Graphics, 74(??):213–224, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300694.

Pfister:1995:GES

[PWK95] Hanspeter Pfister, Frank Wessels, and Arie Kaufman. Gradient estimation and sheared interpolation for the cube architecture. Computers and Graphics, 19(5):667–677, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=5&aid=9500045.

Polatsek:2018:EVA

[PWV^{+}18] Patrik Polatsek, Manuela Waldner, Ivan Viola, Peter Kapec, and Wanda Benesova. Exploring visual attention and saliency modeling for task-based visual analysis. Computers and Graphics, 72(??): 26–38, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300104.

Pourazady:2006:DMN

[PX06] M. Pourazady and X. Xu.

Direct manipulations of NURBS surfaces subjected to geometric constraints. Computers and Graphics, 30 (4):598–609, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930600077X.

Pan:2003:ESB

[PXH^{+}03] Zhigeng Pan, Weiwei Xu, Jin Huang, Mingmin Zhang, and Jiaoying Shi. Easybowling: a small bowling machine based on virtual simulation. Computers and Graphics, 27(2):231–238, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Peng:2005:NAB

[PYD^{+}05] Yu Peng, Jun-Hai Yong, Wei-Ming Dong, Hui Zhang, and Jia-Guang Sun. A new algorithm for Boolean operations on general polygons. Computers and Graphics, 29 (1):57–70, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pan:2005:ILC

[PZH^{+}05] Zhigeng Pan, Jiejie Zhu, Weihua Hu, Hung Pak Lun, and Xin Zhou. Interactive learning of CG in networked virtual environments. Computers and Graphics, 29(2): 273–281, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

PiniMagalhaes:1996:GEI

[PZK96] Leo Pini Magalhaes, Joao Antonio Zuffo, and Marcelo Knorich Zuffo. Guest Editors’ introduction. Computers and Graphics, 20 (3):347–??, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Patel:2023:VDP

[PZM^{+}23] Dhyey Patel, Houssem Zouaghi, Sudhir Mudur, Eric Paquette, Serge Laforest, Martin Rouillard, and Tiberiu Popa. Visual dubbing pipeline with localized lip-sync and two-pass identity transfer. Computers and Graphics, 110 (??):19–27, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001984.

Qian:1992:DHC

[QB92] Kai Qian and Prabir Bhattacharya. Determining holes and connectivity in binary images. Computers and Graphics, 16(3):283– 288, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Qian:2003:HOM

[QD03] Xiaoping Qian and Deba Dutta. Heterogeneous object modeling through direct face neighborhood alteration. Computers and Graphics, 27(6):943–961, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Qin:1997:NMS

[QGGW97] Kaihuai Qin, Minglun Gong, Youjiang Guan, and Wenping Wang. A new method for speeding up ray tracing NURBS surfaces. Computers and Graphics, 21 (5):577–586, September– October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9787263.

Qin:2008:SIC

[QGW08] Hong Qin, Baining Guo, and Guoping Wang. Special issue of CAD/Graphics’ 2007. Computers and Graphics, 32 (6):615–616, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001222.

Qingsong:2001:COB

[QKS01] Zou Qingsong, Kwoh Chee Keong, and Ng Wan Sing. Convex object based volume visualization: a formal proof and example. Computers and Graphics, 25(5):857– 873, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/39/abstract.
html.

Que:2023:LFC

[QL23] Shicheng Que and Yue Li. Lightweight fully connected network-based fast CU size decision for videobased point cloud compression. Computers and Graphics, 117(??):20–30, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002443.

Quiros:1996:GSU

[QLCV96] R. Quirós, J. Lluch, M. Chover, and R. Vivó. Geometric substitution using random L-systems. Computers and Graphics, 20(5):713–721, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=20&
                                        issue=5&aid=9600045.

Quarles:2009:SLM

[QLF^{+}09] John Quarles, Samsun Lampotang, Ira Fischler, Paul Fishwick, and Benjamin Lok. Scaffolded learning with mixed reality. Computers and Graphics, 33 (1):34–46, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001490.

Qammar:1996:FBC

[QM96] Helen Qammar and F. Mossayebi. Fractal basins in the control of the logistic equation. Computers and Graphics, 20(4):589–596, July– August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600029.

Qamar:1991:DFN

[QMHH91] Ihtzaz Qamar, Nurul Mustafa, S. Wilayat Husain, and F. H. Hashmi. The design of a flat negative pattern to make grooves on a spherical spiral groove bearing. Computers and Graphics, 15(3): 417–421, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Qiao:1998:CME

[QN98] Lintian Qiao and Klara Nahrstedt. Comparison of MPEG encryption algorithms. Computers and Graphics, 22(4):437–448, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
4/567.pdf.

Qiao:2023:MCP

[QSLS23] Yuanjian Qiao, Mingwen Shao, Huan Liu, and Kai Shang. Mutual channel prior guided dual-domain interaction network for single image raindrop removal. Computers and Graphics, 112(??):132–142, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000432.

Qi:2022:SPG

[QSXT22] Yuhang Qi, Wanjuan Su, Qingshan Xu, and Wenbing Tao. Sparse prior guided deep multi-view stereo. Computers and Graphics, 107(??):1–9, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001157.

Queau:1993:TMT

[Que93] Philippe Queau. Televirtuality: The merging of telecommunications and virtual reality. Computers and Graphics, 17(6):691–693, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Quien:1991:SRI

[Qui91] Norbert Quien. Some results of the interplay between computer graphics and mathematics. Computers and Graphics, 15(4): 515–518, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Qammar:1995:MCC

[QV95] Helen Qammar and A. Venkatesan. Manifolds and control of chaotic systems. Computers and Graphics, 19(3): 455–460, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500016.

Qian:2014:AFF

[QWC14] Kanglai Qian, Bin Wang, and Huarong Chen. Automatic flexible face replacement with no auxiliary data. Computers and Graphics, 45 (??):64–74, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000752.

Qin:2022:STS

[QYC^{+}22] Jie Qin, Shuaihang Yuan, Jiaxin Chen, Boulbaba Ben Amor, Yi Fang, Nhat Hoang-Xuan, Chi-Bien Chu, Khoi-Nguyen Nguyen-Ngoc, Thien-Tri Cao, Nhat-Khang Ngo, Tuan-Luc Huynh, Hai-Dang Nguyen, Minh-Triet Tran, Haoyang Luo, Jianning Wang, Zheng Zhang, Zihao Xin, Yang Wang, Feng Wang, Ying Tang, Haiqin Chen, Yan Wang, Qunying Zhou, Ji Zhang, and Hongyuan Wang. SHREC’22 track: Sketch-based 3D shape retrieval in the wild. Computers and Graphics, 107(??):104–115, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322001261.

Rodriguez:2003:FNO

[RA03] J. Rodríguez and D. Ayala. Fast neighborhood operations for images and volume data sets. Computers and Graphics, 27(6): 931–942, December 2003.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Richter:2015:BM

[RA15a] Ronald Richter and Marc Alexa. Beam meshes. Computers and Graphics, 53 (part A)(??):28–36, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001417.

Richter:2015:MCV

[RA15b] Ronald Richter and Marc Alexa. Mahalanobis centroidal Voronoi tessellations. Computers and Graphics, 46 (??):48–54, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000971.

Romera:2008:DCE

[RAA^{+}08] M. Romera, G. Alvarez, D. Arroyo, A. B. Orue, V. Fernandez, and G. Pastor. Drawing and computing external rays in the multiple-spiral medallions of the Mandelbrot set. Computers and Graphics, 32 (5):597–610, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000484.

Radai:1996:CAM

[Rad96] Yisrael Radai. Computer art from the Mandelbrot set. Computers and Graphics, 20(6):925–926, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600062.

Ronnow:2021:FAM

[RAF21] Mads J. L. Rønnow, Ulf Assarsson, and Marco Fratarcangeli. Fast analytical motion blur with transparency. Computers and Graphics, 95(??):36–46, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000066.

Raggett:1979:ICG

[Rag79] G. F. Raggett. Interactive computer graphics for the generation of phase plane portraits. Computers and Graphics, 4(3–4): 149–154, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Raggett:1980:TDI

[Rag80] G. F. Raggett. Towards the development of an interactive graphics package for the solution of optimal control problems. Computers and Graphics, 5(2– 4):93–98, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rodriguez:2005:VNA

[RAG05] J. Rodríguez, D. Ayala, and S. Grau. VolumeEVM: a new approach for surface/ volume integration. Computers and Graphics, 29(2): 217–224, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ragulskis:2009:TAP

[Rag09] Minvydas Ragulskis. Timeaveraged patterns produced by stochastic moiré gratings. Computers and Graphics, 33(2):147–150, April 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001246.

Rasheed:1988:SAF

[RAHA88] Sarbast M. Rasheed, Bakir H. Al-Hashemy, and Ahmad I. Sheikh Ahmad. Synthesis and analysis of flow about airfoils with interactive graphics. Computers and Graphics, 12(1): 99–110, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reh:2015:FFT

[RAK^{+}15] Andreas Reh, Aleksandr Amirkhanov, Johann Kastner, Eduard Gröller, and Christoph Heinzl. Fuzzy feature tracking: Visual analysis of industrial 4D-XCT data. Computers and Graphics, 53 (part B)(??): 177–184, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000400.

Ramesh:1989:GTS

[Ram89] S. R. Ramesh. Graphics of truncated sinusoids. Computers and Graphics, 13(4): 545–547, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rankin:1987:GHL

[Ran87a] John R. Rankin. A geometric hidden-line processing algorithm. Computers and Graphics, 11(1):11–19, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rankin:1987:NMP

[Ran87b] John R. Rankin. A note on multi-polygon area filling. Computers and Graphics, 11(4):445–447, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rankin:1988:CPD

[Ran88] John R. Rankin. Classes of polyhedra defined by jet graphics. Computers and Graphics, 12(2): 239–254, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rankin:1991:RBL

[Ran91] John R. Rankin. Recursive bisection line algorithm. Computers and Graphics, 15(1):1–8, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rauschenbach:2006:ITN

[Rau06] Uwe Rauschenbach. Interactive TV: a new application for mobile computing. Computers and Graphics, 30 (5):727–736, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001233.

Reisert:2006:SOS

[RB06] Marco Reisert and Hans Burkhardt. Second order 3D shape features: an exhaustive study. Computers and Graphics, 30(2): 197–206, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000276.

Rodriguez:2008:EII

[RB08] I. Rodriguez and R. Boulic.

Evaluating the influence of induced passive torques in the simulation of timevarying human poses. Computers and Graphics, 32 (4):474–484, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000721.

Raffo:2020:DDQ

[RB20] Andrea Raffo and Silvia Biasotti. Data-driven quasi-interpolant spline surfaces for point cloud approximation. Computers and Graphics, 89(??):144– 155, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300558.

Raviv:2011:AIG

[RBB^{+}11] Dan Raviv, Alexander M.

Bronstein, Michael M. Bronstein, Ron Kimmel, and Nir Sochen. Affine-invariant geodesic geometry of deformable 3D shapes. Computers and Graphics, 35(3): 692–697, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000690.

Ravaglia:2017:ETS

[RBF17] Joris Ravaglia, Alexandra Bac, and Richard A. Fournier. Extraction of tubular shapes from dense point clouds and application to tree reconstruction from laser scanned data. Computers and Graphics, 66(??):23–33, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300651.

Romanengo:2020:HBI

[RBF20] Chiara Romanengo, Silvia Biasotti, and Bianca Falcidieno. HT-based identification of 3D feature curves and their insertion into 3D meshes. Computers and Graphics, 89(??): 105–116, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300637.

Ruiz:2010:VIC

[RBFS10] M. Ruiz, I. Boada, M. Feixas, and M. Sbert. Viewpoint information channel for illustrative volume rendering. Computers and Graphics, 34 (4):351–360, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000439.

Reuter:2009:DLB

[RBG^{+}09] Martin Reuter, Silvia Biasotti, Daniela Giorgi, Giuseppe Patanè, and Michela Spagnuolo. Discrete Laplace– Beltrami operators for shape analysis and segmentation. Computers and Graphics, 33(3):381–390, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000272.

Renaud:1995:MPH

[RBL95] C. Renaud, F. Bricout, and E. Leprêtre. Massively parallel hemispherical projection for progressive radiosity. Computers and Graphics, 19(2): 273–279, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400153.

Rodrigues:2021:WQR

[RBLB21] Ariane Moraes Bueno Rodrigues, Gabriel Diniz Junqueira Barbosa, Hélio Côrtes Vieira Lopes, and Simone Diniz Junqueira Barbosa. What questions reveal about novices’ attempts to make sense of data visualizations: Patterns and misconceptions. Computers and Graphics, 94(??):32– 42, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301497.

Rutledge:1996:MOH

[RBP96] Lloyd Rutledge, John Buford, and Roger Price. Mobile objects and the HyOctane distributed hyperdocument server. Computers and Graphics, 20(5):633–639, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600037.

Romera:2001:SLP

[RBP^{+}01] M. Romera, V. Bañuls, G. Pastor, G. Álvarez, and F. Montoya. Snail-like pattern generation with the Hénon family of maps. Computers and Graphics, 25(3): 529–537, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/41/article.
pdf.

Rix:1994:SMP

[RBUB94] Joachim Rix, Andreas Burkert, Max Ungerer, and Monika Bräckelmann. STEP meets PREMO: Product modelling and new presentation techniques. Computers and Graphics, 18(4): 553–562, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Regenbrecht:2001:TAD

[RBW01] Holger Regenbrecht, Gregory Baratoff, and Michael Wagner. A tangible AR desktop environment. Computers and Graphics, 25 (5):755–763, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/57/34/30/abstract.
html.

Rosenstein:1994:VEF

[RC94] Michael T. Rosenstein and James J. Collins. Visualizing the effects of filtering chaotic signals. Computers and Graphics, 18(4): 587–592, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ripolles:2008:OMC

[RC08] O. Ripollés and M. Chover.

Optimizing the management of continuous level of detail models on GPU. Computers and Graphics, 32(3): 307–319, June 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000319.

Ranon:2015:ACC

[RCB15] Roberto Ranon, Luca Chittaro, and Fabio Buttussi. Automatic camera control meets emergency simulations: an application to aviation safety. Computers and Graphics, 48(??): 23–34, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000291.

Rodriguez:2010:NAC

[RCBS10] Rafael Rodriguez, Eva Cerezo, Sandra Baldassarri, and Francisco J. Seron. New approaches to culling and LOD methods for scenes with multiple virtual actors. Computers and Graphics, 34(6): 729–741, December 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001123.

Romao:2004:AAE

[RCD^{+}04] Teresa Romão, Nuno Correia, Eduardo Dias, José Danado, Adelaide Trabuco, Carlos Santos, Rossana Santos, António Câmara, and Edmundo Nobre. ANTS— augmented environments. Computers and Graphics, 28 (5):625–633, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ruiz:2005:SSC

[RCG^{+}05] Oscar E. Ruiz, Carlos A.

Cadavid, Miguel Granados, Sebastián Peña, and Eliana Vásquez. 2D shape similarity as a complement for Voronoi–Delone methods in shape reconstruction. Computers and Graphics, 29 (1):81–94, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Roussel:2018:EDM

[RCLM18] Robin Roussel, Marie-Paule Cani, Jean-Claude Léon, and Niloy J. Mitra. Exploratory design of mechanical devices with motion constraints. Computers and Graphics, 74 (??):244–256, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300888.

Reina:2020:MTV

[RCM^{+}20] Guido Reina, Hank Childs, Krešimir Matković, Katja Bühler, Manuela Waldner, David Pugmire, Barbora Kozlíková, Timo Ropinski, Patric Ljung, Takayuki Itoh, Eduard Gröller, and Michael Krone. The moving target of visualization software for an increasingly complex world. Computers and Graphics, 87(??):12–29, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300078.

Rigiroli:2001:MRC

[RCPB01] P. Rigiroli, P. Campadelli, A. Pedotti, and N. Alberto Borghese. Mesh refinement with color attributes. Computers and Graphics, 25(3):449–461, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/35/article.
pdf.

Roberts:2021:DTS

[RdAMA21] Richard A. Roberts, Rafael Kuffner dos Anjos, Akinobu Maejima, and Ken Anjyo. Deformation transfer survey. Computers and Graphics, 94 (??):52–61, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301552.

Raposo:2001:CCC

[RdCAM01] Alberto B. Raposo, Adailton J. A. da Cruz, Christian M. Adriano, and Léo P. Magalhães. Coordination components for collaborative virtual environments. Computers and Graphics, 25(6):1025–1039, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/35/abstract.
html.

Ribeiro:2016:HHD

[RdCVL16] Paula Ceccon Ribeiro, Haroldo Fraga de Campos Velho, and Hélio Lopes. Helmholtz–Hodge decomposition and the analysis of 2D vector field ensembles. Computers and Graphics, 55(??):80–96, April 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316000029.

Rallis:2018:STS

[RDD^{+}18] Ioannis Rallis, Nikolaos Doulamis, Anastasios Doulamis, Athanasios Voulodimos, and Vassilios Vescoukis. Spatiotemporal summarization of dance choreographies. Computers and Graphics, 73(??): 88–101, June 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300499.

Rueda:2008:GBR

[RdMF08] A. J. Rueda, J. Ruiz de Miras, and F. R. Feito. GPU-based rendering of curved polygons using simplicial coverings. Computers and Graphics, 32 (5):581–588, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000915.

Retondaro:2022:OBT

[RE22] Luis C. S. C. Retondaro and Claudio Esperança. Optimized 2D ball trees for shape layout applications. Computers and Graphics, 103 (??):129–139, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000231.

Reed:1985:SVD

[Ree85] T. N. Reed. Standardization of the virtual device metafile and the virtual device interface. Computers and Graphics, 9(1):33–38, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reeves:1991:NHM

[Ree91] Ray Reeves. A note on Halley’s method. Computers and Graphics, 15 (1):89–90, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reeves:1992:FIH

[Ree92] Ray Reeves. Further insights into Halley’s method. Computers and Graphics, 16 (2):235–236, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ruder:1989:LSI

[REG^{+}89] H. Ruder, T. Ertl, F. Geyer, H. Herold, and U. Kraus. Line-of-sight integration: a powerful tool for visualization of three-dimensional scalar fields. Computers and Graphics, 13(2): 223–228, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rehwald:1985:VIT

[Reh85] P. Rehwald. VDAFS — an interface to transfer surface description data between CAD systems. Computers and Graphics, 9(1): 69–70, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reid:1975:SOT

[Rei75] A. C. Reid. Searching for oil through an interactive graphic terminal. Computers and Graphics, 1(1):129– 136, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reiter:1992:VDR

[Rei92]  Clifford A. Reiter.  Vi-

sualizing the dynamics of the Rayleigh quotient iteration. Computers and Graphics, 16(3):341–344, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reid:1995:VCS

[Rei95] John D. Reid. Visualizing cross section forces. Computers and Graphics, 19(3): 475–480, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500019.

Reiter:1997:ADS

[Rei97a]  Clifford A. Reiter.  At-

tractors with dueling symmetry. Computers and Graphics, 21(2):263–271, March–April 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.

cgi?year=1997&volume=21&
issue=2&aid=9600089.

Reiter:1997:CAS

[Rei97b] Clifford A. Reiter. Chaotic attractors with the symmetry of a tetrahedron. Computers and Graphics, 21(6):841–848, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700062.

Reiter:1997:GLH

[Rei97c] Clifford A. Reiter. The Game of Life on a hyperbolic domain. Computers and Graphics, 21(5):673–683, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700043.

Reinhard:1998:MRS

[Rei98]  Klein Reinhard.  Mul-

tiresolution representations for surfaces meshes based on the vertex decimation method. Computers and Graphics, 22(1):13–26, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        1/511.pdf.

Reiter:2002:CGC

[Rei02] Clifford A. Reiter. Chaos and graphics: Chaotic attractors exhibiting quasicrystalline structure. Computers and Graphics, 26(3): 511–517, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/13/20/68/54/39/abstract.
                                        html.

Reiners:2004:SIO

[Rei04a]  D. Reiners.  Special is-

sue on the OpenSG Symposium and OpenSG Plus. Computers and Graphics, 28 (1):59–61, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reiter:2004:VFD

[Rei04b] Clifford A. Reiter. Views of Fibonacci dynamics. Computers and Graphics, 28(2): 297–300, April 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reiter:2010:MSC

[Rei10] Clifford A. Reiter. Medley of spirals from cyclic cellular automata. Computers and Graphics, 34 (1):72–76, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001137.

RuizDeMiras:1997:ITC

[RF97] J. Ruiz De Miras and F. R. Feito. Inclusion test for curved-edge polygons. Computers and Graphics, 21(6):815–824, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700060.

Rivero:2000:BOG

[RF00] M. Rivero and F. R. Feito.

Boolean operations on general planar polygons. Computers and Graphics, 24(6): 881–896, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/34/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/31/article.
pdf.

Raffo:2021:PSR

[RFB^{+}21] Andrea Raffo, Ulderico Fugacci, Silvia Biasotti, Walter Rocchia, Yonghuai Liu, Ekpo Otu, Reyer Zwiggelaar, David Hunter, Evangelia I. Zacharaki, Eleftheria Psatha, Dimitrios Laskos, Gerasimos Arvanitis, Konstantinos Moustakas, Tunde Aderinwale, Charles Christoffer, Woong-Hee Shin, Daisuke Kihara, Andrea Giachetti, Huu-Nghia Nguyen, Tuan-Duy Nguyen, Vinh-Thuyen Nguyen-Truong, Danh Le-Thanh, Hai-Dang Nguyen, and Minh-Triet Tran. SHREC 2021: Retrieval and classification of protein surfaces equipped with physical and chemical properties. Computers and Graphics, 99(??):1–21, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001254.

Raffo:2023:GNG

[RFB23] Andrea Raffo, Ulderico Fugacci, and Silvia Biasotti. GEO-Nav: a geometric dataset of voltage-gated sodium channels. Computers and Graphics, 115 (??):285–295, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001188.

Rueda:2002:TBR

[RFR02] A. J. Rueda, F. R. Feito, and M. Rivero. A trianglebased representation for polygons and its applications. Computers and Graphics, 26(5):805–814, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/43/abstract.
html.

Rink:2022:SSV

[RFS22] Karsten Rink, Kathrin Feige, and Gerik Scheuermann. Special section on visualization in environmental sciences. Computers and Graphics, 104(??): A4–A5, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000632.

Ren:2017:RTD

[RFZ^{+}17] Pu Ren, Yachun Fan, Mingquan Zhou, Zhe Wang, Guoguang Du, and Lu Qian. Rapid three-dimensional scene modeling by sketch retrieval and auto-arrangement. Computers and Graphics, 64 (??):26–36, May 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300183.

Ramana:1993:SRM

[RG93] N. V. Ramana and Indira Ghosh. Software report: Molecular graphics software — MOGRA. Computers and Graphics, 17(4): 415–416, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ramachandran:2018:JPP

[RGdL^{+}18] Srinivasan Ramachandran, Donya Ghafourzadeh, Martin de Lasa, Tiberiu Popa, and Eric Paquette. Joint planar parameterization of segmented parts and cage deformation for dense correspondence. Computers and Graphics, 74(??):202– 212, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300864.

Rotard:2007:SLS

[RGE07] Martin Rotard, Mark Giereth, and Thomas Ertl. Semantic lenses: Seamless augmentation of Web pages with context information from implicit queries. Computers and Graphics, 31(3): 361–369, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000581.

Rougeron:2002:SIP

[RGGB02] Gilles Rougeron, François Gaudaire, Yannick Gabillet, and Kadi Bouatouch. Simulation of the indoor propagation of a 60GHz electromagnetic wave with a time-dependent radiosity algorithm. Computers and Graphics, 26(1):125– 141, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/39/abstract.
html.

Ristovski:2019:UAV

[RGH^{+}19] Gordan Ristovski, Nicole Garbers, Horst K. Hahn, Tobias Preusser, and Lars Linsen. Uncertainty-aware visual analysis of radiofrequency ablation simulations. Computers and Graphics, 79(??):24–35, April 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301973.

Ross:2018:FSS

[RGJdQ18] Arun Ross, Eduardo S. L.

Gastal, Joaquim Jorge, and Ricardo L. de Queiroz. Foreword to the special section on SIBGRAPI 2018. Computers and Graphics, 77 (??):A5–A6, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301663.

Royo:2022:NLS

[RGMJ22] Diego Royo, Jorge García, Adolfo Muñoz, and Adrian Jarabo. Non-line-of-sight transient rendering. Computers and Graphics, 107 (??):84–92, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001200.

Ridel:2015:PCM

[RGRG15] Brett Ridel, Gaël Guennebaud, Patrick Reuter, and Xavier Granier. Paraboliccylindrical moving least squares surfaces. Computers and Graphics, 51 (??):60–66, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000539.

Rehwald:1985:AVM

[RH85] Peter Rehwald and Christoph Hornung. An analytical visibility method for displaying parametrically defined surfaces. Computers and Graphics, 9(3):275–281, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ross:1995:VMB

[RHBS95] Th. Roß, H. Handels, U. Breuer, and K. Szabó. 3D visualization of microvascular blood vessel networks.

Computers and Graphics, 19 (1):89–96, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400124.

Rohmer:2015:RTC

[RHC15] Damien Rohmer, Stefanie Hahmann, and Marie-Paule Cani. Real-time continuous self-replicating details for shape deformation. Computers and Graphics, 51 (??):67–73, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000588.

Riveiro:2014:EVU

[RHFL14] Maria Riveiro, Tove Helldin, Göran Falkman, and Mikael Lebram. Effects of visualizing uncertainty on decisionmaking in a target identification scenario. Computers and Graphics, 41(??): 84–98, June 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000302.

Rajan:2020:DPF

[RHK^{+}20] Krishna Rajan, Soheil Hashemi, Ulya Karpuzcu, Michael Doggett, and Sherief Reda. Dual-precision fixed-point arithmetic for low-power ray-triangle intersections. Computers and Graphics, 87(??):72–79, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932030008X.

Regenbrecht:2012:VMM

[RHM^{+}12] Holger Regenbrecht, Simon Hoermann, Graham McGregor, Brian Dixon, Elizabeth Franz, Claudia Ott, Leigh Hale, Thomas Schubert, and Julia Hoermann. Visual manipulations for motor rehabilitation. Computers and Graphics, 36(7): 819–834, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001136.

Reithinger:2003:SMI

[RHN03] Norbert Reithinger, Gerd Herzog, and Alassane Ndiaye. Situated multimodal interaction in SmartKcom. Computers and Graphics, 27 (6):899–903, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Robertson:1994:MDC

[RHS^{+}94] P. K. Robertson, M. Hutchins, D. R. Stevenson, S. Barrass, C. Gunn, and D. Smith. Mapping data into colour gamuts: Using interaction to increase usability and reduce complexity. Computers and Graphics, 18(5):653–665, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Richards:1989:GRP

[Ric89] T. Richards. Graphical representation of pseudorandom sequences. Computers and Graphics, 13(2): 261–262, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Riesenfeld:1978:CTC

[Rie78] Richard F. Riesenfeld. Current trends in computer graphics. Computers and Graphics, 3(4):115–122, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rix:1983:GRD

[Rix83] J. Rix. GKS raster driver.

Computers and Graphics, 7 (3–4):349–350, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rix:1984:DGD

[Rix84] J. Rix. On developing a GKS driver architecture for raster workstations. Computers and Graphics, 8(2):

145–148, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rousseau:2006:RRT

[RJG06] Pierre Rousseau, Vincent Jolivet, and Djamchid Ghazanfarpour. Realistic realtime rain rendering. Computers and Graphics, 30 (4):507–518, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000859.

Rossgatterer:2012:MDB

[RJKV12] M. Rossgatterer, B. Jüttler, M. Kapl, and G. Della Vecchia. Medial design of blades for hydroelectric turbines and ship propellers. Computers and Graphics, 36 (5):434–444, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000611.

Roth:1998:ACK

[RJS98] Volker Roth and Mehrdad Jalali-Sohi. Access control and key management for mobile agents. Computers and Graphics, 22(4):457–461, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/cag/sub/1998/22/
4/569.pdf.

Rauschenbach:2001:GRF

[RJS01] Uwe Rauschenbach, Stefan Jeschke, and Heidrun Schumann. General rectangular fisheye views for 2D graphics. Computers and Graphics, 25(4):609– 617, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/34/abstract.
html.

Roussopoulos:1984:GGD

[RK84] N. Roussopoulos and S. Kelley. GDDT — a graphical design and documentation tool for software development. Computers and Graphics, 8 (3):309–323, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ke:1993:SBP

[rKC93] Hao ren Ke and Ruei-Chuan Chang. Sample buffer: a progressive refinement raycasting algorithm for volume rendering. Computers and Graphics, 17(3): 277–283, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Reischl:2022:PNO

[RKG22] Maximilian Reischl, Christian Knauer, and Michael Guthe. Parallel nearoptimal pathfinding based on landmarks. Computers and Graphics, 102(??): 1–8, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002466.

Reisner-Kollmann:2013:RSB

[RKMP13]  Irene  Reisner-Kollmann,

Stefan Maierhofer, and Werner Purgathofer. Reconstructing shape boundaries with multimodal constraints. Computers and Graphics, 37(3):137–147, May 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000022.

Ribelles:2002:TSM

[RLB^{+}02] J. Ribelles, A. López, O. Belmonte, I. Remolar, and M. Chover. Technical section: Multiresolution modeling of arbitrary polygonal surfaces: a characterization. Computers and Graphics, 26(3):449–462, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/34/abstract.
html.

Reiner:2012:RCI

[RLD^{+}12] Tim Reiner, Sylvain Lefebvre, Lorenz Diener, Ismael García, Bruno Jobard, and Carsten Dachsbacher. A runtime cache for interactive procedural modeling. Computers and Graphics, 36 (5):366–375, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000702.

Ritz:2012:HRA

[RLS^{+}12] M. Ritz, F. Langguth, M. Scholz, M. Goesele, and A. Stork. High resolution acquisition of detailed surfaces with lens-shifted structured light. Computers and Graphics, 36(1):16– 27, February 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001506.

Roberto:2016:TMD

[RLT16] Rafael Roberto, João Paulo Lima, and Veronica Teichrieb. Tracking for mobile devices: a systematic mapping study. Computers and Graphics, 56(??): 20–30, May 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300115.

Roberto:2019:GSI

[RLU^{+}19] Rafael Roberto, João Paulo Lima, Hideaki Uchiyama, Veronica Teichrieb, and Rin ichiro Taniguchi. Geometrical and statistical incremental semantic modeling on mobile devices. Computers and Graphics, 84(??): 199–211, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301542.

Ruhlmann:1991:LSN

[RM91] George M. Ruhlmann and John C. McKeeman. Local search. A new hidden line elimination algorithm to display spherical coordinate equations. Computers and Graphics, 15(4): 535–544, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rilling:2005:VTS

[RM05] Juergen Rilling and S. P. Mudur. 3D visualization techniques to support slicing-based program comprehension. Computers and Graphics, 29(3):311– 329, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rakhimol:2022:RAT

[RM22] V. Rakhimol and P. Uma Maheswari. Restoration of ancient temple murals using cGAN and PConv networks. Computers and Graphics, 109(??):100–110, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001881.

Reiner:2011:IMI

[RMD11] Tim Reiner, Gregor Mückl, and Carsten Dachsbacher. Interactive modeling of implicit surfaces using a direct visualization approach with signed distance functions. Computers and Graphics, 35(3):596–603, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000495.

Rodrigues:2015:CBS

[RMG15] Rui S. V. Rodrigues, José F. M. Morgado, and Abel J. P. Gomes. A contourbased segmentation algorithm for triangle meshes in 3D space. Computers and Graphics, 49(??): 24–35, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000424.

Raina:2019:SFP

[RMP19] Prashant Raina, Sudhir Mudur, and Tiberiu Popa. Sharpness fields in point clouds using deep learning. Computers and Graphics, 78 (??):37–53, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830181X.

Rodrigues:2022:FSS

[RMSB22] Nuno Rodrigues, Daniel Mendes, Luís Paulo Santos, and Kadi Bouatouch. Foreword to the special section on recent advances in graphics and interaction. Computers and Graphics, 102 (??):A8–A9, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002624.

Robson:2011:CAG

[RMSC11] Cody Robson, Ron Maharik, Alla Sheffer, and Nathan Carr. Contextaware garment modeling from sketches. Computers and Graphics, 35(3): 604–613, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000410.

Ristovski:2017:UVI

[RMW^{+}17] Gordan Ristovski, Jose Matute, Thomas Wehrum, Andreas Harloff, Horst K. Hahn, and Lars Linsen. Uncertainty visualization for interactive assessment of stenotic regions in vascular structures. Computers and Graphics, 69(??): 116–130, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301681.

Ruprecht:1995:SFF

[RNM95] Detlef Ruprecht, Ralf Nagel, and Heinrich Müller. Spatial free-form deformation with scattered data interpolation methods. Computers and Graphics, 19(1):63– 71, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400122.

Roberto:2019:CBC

[RNM^{+}19] Guilherme F. Roberto, Marcelo Z.

Nascimento, Alessandro S. Martins, Thaína A. A. Tosta, Paulo R. Faria, and Leandro A. Neves. Classification of breast and colorectal tumors based on percolation of color normalized images. Computers and Graphics, 84(??): 134–143, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301360.

Robertson:1987:CGT

[RO87] P. K. Robertson and J. F. O’Callaghan. Colour graphics and three-dimensional scene synthesis in image display. Computers and Graphics, 11(4):469–473, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ren:2013:SFG

[RO13] Gang Ren and Eamonn O’Neill. 3D selection with freehand gesture. Computers and Graphics, 37(3): 101–120, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001823.

Roberts:1978:UCG

[Rob78] Michael C. Roberts. Use of computer graphics in the evaluation of sanitary landfill sites. Computers and Graphics, 3(4):167–169, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Robertson:1991:GEI

[Rob91] Philip K. Robertson. Guest Editor’s introduction. Computers and Graphics, 15(3): 329–??, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Roche:1989:DFG

[Roc89]  Martin Roche.  Defin-

ing a faceted generalized cylinder by projections of cross sections. Computers and Graphics, 13(3): 349–354, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Roelandt:2000:TDE

[Roe00] Jos R. T. C. Roelandt.

Three-dimensional echocardiography: the future today! Computers and Graphics, 24 (5):715–729, October 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/32/article.
pdf.

Rogers:1983:V

[Rog83] G. Rogers. Viewing. Computers and Graphics, 7(1): 83–85, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rogers:1985:GEI

[Rog85] Gary Rogers. Guest Editor’s introduction: Computer graphics in architecture. Computers and Graphics, 9(2):81–83, ???? 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849385900159.

Rojas:1991:TEC

[Roj91] Raul Rojas. A tutorial on efficient computer graphic representations of the Mandelbrot set. Computers and Graphics, 15(1):91–100, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rojas:1994:OIP

[Roj94] Rauel Rojas. Oscillating iteration paths in neural networks learning. Computers and Graphics, 18(4): 593–597, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rokita:1993:FGD

[Rok93a] Przemysław Rokita. Fast generation of depth of field effects in computer graphics. Computers and Graphics, 17(5):593–595, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rokita:1993:MRH

[Rok93b]  Przemysław Rokita.  A

model for rendering high intensity lights. Computers and Graphics, 17(4): 431–437, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ropinski:2011:SGB

[ROP11] Timo Ropinski, Steffen Oeltze, and Bernhard Preim. Survey of glyph-based visualization techniques for spatial multivariate medical data. Computers and Graphics, 35(2):392– 401, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000203.

Rossignac:1990:IFB

[Ros90] Jarosław R. Rossignac. Issues on feature-based editing and interrogation of solid models. Computers and Graphics, 14(2): 149–172, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rossignac:1992:CGB

[Ros92] Jarosław R. Rossignac. Computers and Graphics Best Paper Award (1990). Computers and Graphics, 16 (1):1–2, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rowe:1982:ISE

[Row82] P. P. Rowe. Inhibition, selection and enhancement operators for picture processing. Computers and Graphics, 6(1):39–40, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rowe:1986:MCG

[Row86] J. H. Rowe. Metafiles and computer graphics. Computers and Graphics, 10(2): 103–106, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Roy:2018:HED

[RP18] Bruno Roy and Pierre Poulin. A hybrid Eulerian-DFSPH scheme for efficient surface band liquid simulation. Computers and Graphics, 77(??):194– 204, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301687.

Rahmani:2020:MAP

[RP20] Vahid Rahmani and Nuria Pelechano. Multi-agent parallel hierarchical path finding in navigation meshes (MA-HNA*). Computers and Graphics, 86(??): 1–14, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301694.

Rahmani:2022:THL

[RP22] Vahid Rahmani and Nuria Pelechano. Towards a human-like approach to path finding. Computers and Graphics, 102(??):164–174, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001849.

Romera:2000:GCE

[RPAM00] M. Romera, G. Pastor, G. Alvarez, and F. Montoya. Growth in complex exponential dynamics. Computers and Graphics, 24 (1):115–131, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/37/article.
pdf.

Romera:2004:EAD

[RPÁM04] M. Romera, G. Pastor, G. Álvarez, and F. Montoya. External arguments of Douady cauliflowers in the Mandelbrot set. Computers and Graphics, 28(3): 437–449, June 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Romera:2006:EAM

[RPAM06] M. Romera, G. Pastor, G. Alvarez, and F. Montoya. External arguments in the multiple-spiral medallions of the Mandelbrot set. Computers and Graphics, 30(3):460–469, June 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000690.

Razdan:2001:VVM

[RPFC01] Anshuman Razdan, Kamal Patel, Gerald E. Farin, and David G. Capco. Volume visualization of multicolor laser confocal microscope data. Computers and Graphics, 25(3):371–382, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/28/article.
                                        pdf.

Ristovski:2014:UMV

[RPHL14] Gordan Ristovski, Tobias Preusser, Horst K. Hahn, and Lars Linsen. Uncertainty in medical visualization: Towards a taxonomy. Computers and Graphics, 39(??):60–73, April 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001568.

Rodriguez-Pardo:2023:NNF

[RPKLMG23] Carlos Rodriguez-Pardo, Konstantinos Kazatzis, Jorge Lopez-Moreno, and Elena Garces. NeuBTF: Neural fields for BTF encoding and transfer. Computers and Graphics, 114 (??):239–246, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001140.

Romera:1996:GTA

[RPM96] M. Romera, G. Pastor, and F. Montoya. Graphic tools to analyse one-dimensional quadratic maps. Computers and Graphics, 20(2): 333–339, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500134.

Romera:1997:SCE

[RPM97] M. Romera, G. Pastor, and F. Montoya. A scaling constant equal to unity in 1-D quadratic maps. Computers and Graphics, 21(6):849–857, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700063.

Roy:2020:PUF

[RPP20] Bruno Roy, Eric Paquette, and Pierre Poulin. Particle upsampling as a flexible post-processing approach to increase details in animations of splashing liquids. Computers and Graphics, 88(??):57–69, May 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300273.

Rogla:2021:PCG

[RPP21] Otger Rogla, Gustavo A. Patow, and Nuria Pelechano. Procedural crowd generation for semantically augmented virtual cities. Computers and Graphics, 99 (??):83–99, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001291.

Rodriguez-Pardo:2019:AES

[RPSP^{+}19]  Carlos  Rodriguez-Pardo,

Sergio Suja, David Pascual, Jorge Lopez-Moreno, and Elena Garces. Automatic extraction and synthesis of regular repeatable patterns. Computers and Graphics, 83 (??):33–41, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301049.

Ramprasad:1992:CLA

[RR92] V. V. Ramprasad and E. G.

Rajan. Cellular logic array processing (CLAP). Computers and Graphics, 16 (1):1–??, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Robinson:2001:LSA

[RR01] John A. Robinson and Charles Robertson. The LivePaper system: augmenting paper on an enhanced tabletop. Computers and Graphics, 25(5):731– 743, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/28/abstract.
html.

Richter:2015:DAP

[RR15] Stephan R. Richter and Stefan Roth. A discriminative approach to perspective shape from shading in uncalibrated illumination. Computers and Graphics, 53 (part A)(??): 72–81, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001387.

Romanengo:2022:SFR

[RRB^{+}22] Chiara Romanengo, Andrea Raffo, Silvia Biasotti, Bianca Falcidieno, Vlassis Fotis, Ioannis Romanelis, Eleftheria Psatha, Konstantinos Moustakas, Ivan Sipiran, Quang-Thuc Nguyen, Chi-Bien Chu, Khoi-Nguyen Nguyen-Ngoc, Dinh-Khoi Vo, Tuan-An To, Nham-Tan Nguyen, Nhat-Quynh Le-Pham, Hai-Dang Nguyen, Minh-Triet Tran, Yifan Qie, and Nabil Anwer. SHREC 2022: Fitting and recognition of simple geometric primitives on point clouds. Computers and Graphics, 107(??):32– 49, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001224.

Roussellet:2018:DIM

[RRC^{+}18] Valentin Roussellet, Nadine Abu Rumman, Florian Canezin, Nicolas Mellado, Ladislav Kavan, and Loïc Barthe. Dynamic implicit muscles for character skinning. Computers and Graphics, 77(??):227– 239, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301742.

Romanengo:2022:PFP

[RRQ^{+}22] Chiara Romanengo, Andrea Raffo, Yifan Qie, Nabil Anwer, and Bianca Falcidieno. Fit4CAD: a point cloud benchmark for fitting simple geometric primitives in CAD objects. Computers and Graphics, 102 (??):133–143, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002053.

Rieber:1975:IPG

[RS75] J. E. Rieber and A. C.

Shaw. Interactive picture generation and manipulation through formal descriptions. Computers and Graphics, 1 (1):95–108, May 1975. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Richardson:1996:PP

[RS96] Ryan Richardson and Christine Shannon. Palindrome pictures. Computers and Graphics, 20(4):597–603, July–August 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600030.

Rauschenbach:1999:DDI

[RS99] Uwe Rauschenbach and Heidrun Schumann. Demanddriven image transmission with levels of detail and regions of interest. Computers and Graphics, 23(6): 857–866, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/40/article.
pdf.

Ragulskis:2008:CPU

[RS08] Minvydas Ragulskis and Miguel A. F. Sanjuan. Chaotic pattern of unsmoothed isochromatics around the regions of concentrated stresses. Computers and Graphics, 32(1):116– 119, February 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001690.

Rohmer:2022:FSS

[RS22] Damien Rohmer and Karan Singh. Foreword to the special section on MIG 2021. Computers and Graphics, 106(??):A4–A5, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000620.

Riffnaller-Schiefer:2018:PBD

[RSAF18] Andreas Riffnaller-Schiefer, Ursula H. Augsdörfer, and Dieter W. Fellner. Physicsbased deformation of subdivision surfaces for shared virtual worlds. Computers and Graphics, 71(??): 66–76, April 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317302182.

Ribeiro:2019:VEE

[RSB^{+}19] Paula Ceccon Ribeiro, Guilherme G. Schardong, Simone D. J. Barbosa, Clarisse Sieckenius de Souza, and Hélio Lopes. Visual exploration of an ensemble of classifiers. Computers and Graphics, 85 (??):23–41, December 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301402.

Rink:2022:EES

[RŞH^{+}22] Karsten Rink, Özgür Ozan Şen, Marco Hannemann, Uta Ködel, Erik Nixdorf, Ute Weber, Ulrike Werban, Martin Schrön, Thomas Kalbacher, and Olaf Kolditz. An environmental exploration system for visual scenario analysis of regional hydro-meteorological systems. Computers and Graphics, 103(??):192–200, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000309.

Raidou:2024:FSS

[RSK^{+}24] Renata G. Raidou, Bjorn Sommer, Torsten W. Kuhlen, Michael Krone, Thomas Schultz, and Hsiang-Yun Wu. Foreword special section on VSI: C&G VCBM 2022. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000517.

Rodrigues:2007:ISS

[RSN^{+}07] Maria Andréia F. Rodrigues, Wendel B. Silva, Milton E. Barbosa Neto, Duncan F. Gillies, and Isabel M. M. P. Ribeiro. An interactive simulation system for training and treatment planning in orthodontics. Computers and Graphics, 31 (5):688–697, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001161.

Rompapas:2019:TLS

[RSP^{+}19] Damien Constantine Rompapas, Christian Sandor, Alexander Plopski, Daniel Saakes, Joongi Shin, Takafumi Taketomi, and Hirokazu Kato. Towards large scale high fidelity collaborative augmented reality. Computers and Graphics, 84(??):24– 41, November 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301359.

Ramos:2018:CAR

[RTB^{+}18] Saulo Ramos, Diogo Fernando Trevisan, Harlen C. Batagelo, Mario Costa Sousa, and João Paulo Gois. Contour-aware 3D reconstruction of side-view sketches. Computers and Graphics, 77(??):97–107, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301432.

Rakkolainen:2001:CIM

[RV01] Ismo Rakkolainen and Teija Vainio. A 3D City Info for mobile users. Computers and Graphics, 25(4):619– 625, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/57/33/35/abstract.
                                        html.

Romeiro:2008:SGR

[RVdF08] Fabiano Romeiro, Luiz Velho, and Luiz Henrique de Figueiredo. Scalable GPU rendering of CSG models. Computers and Graphics, 32 (5):526–539, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000745.

Roth:2004:MTC

[RVR04] Marcus Roth, Gerrit Voss, and Dirk Reiners. Multithreading and clustering for scene graph systems. Computers and Graphics, 28 (1):63–66, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Roch:1987:CGC

[RW87] Michel Roch and Jacques Weber. Computer graphics in chemical education and research. Computers and Graphics, 11(1):55–60, 1987.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ribarsky:2014:SMA

[RWD14] William Ribarsky, Derek Xiaoyu Wang, and Wenwen Dou. Social media analytics for competitive advantage. Computers and Graphics, 38 (??):328–331, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001829.

Rotard:2005:CIC

[RWE05] Martin Rotard, Daniel Weiskopf, and Thomas Ertl. A combined introductory course on human– computer interaction and computer graphics. Computers and Graphics, 29(2): 267–272, April 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Rao:2023:RGD

[RYNJ23] Yunbo Rao, Yuling Yi, Obed Tettey Nartey, and Saeed Ullah Jan. Relevance gradient descent for parameter optimization of image enhancement. Computers and Graphics, 117(??): 124–133, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002534.

Ran:2019:ICC

[RZF19] Qing Ran, Wenjing Zhao, and Jieqing Feng. Intrinsic color correction for stereo matching. Computers and Graphics, 82(??): 22–31, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300639.

Ran:2020:HPH

[RZY^{+}20] Qing Ran, Kaimao Zhou, Yong-Liang Yang, Junpeng Kang, Linan Zhu, Yizhi Tang, and Jieqing Feng. High-precision human body acquisition via multi-view binocular stereopsis. Computers and Graphics, 87(??): 43–61, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300054.

Schroeder:2001:CNI

[S^{+}01] Ralph Schroeder et al. Collaborating in networked immersive spaces: as good as being there together? Computers and Graphics, 25 (5):781–788, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-

ng/10/13/20/57/34/32/abstract.
html.

Sobhanpanah:1986:PMQ

[SA86] Cathy Sobhanpanah and Ian O. Angell. Polygonal mesh and quad-tree display algorithms for nonconvex crystal structures. Computers and Graphics, 10(4): 341–349, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sobhanpanah:1987:CDP

[SA87] Farzin Sobhanpanah and Ian O. Angell. On the construction and display of a polytopal mesh for the ndimensional hypercube and (n + 1) simplex. Computers and Graphics, 11(3): 281–288, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Santos:2004:MGR

[SA04] Maribel Yasmina Santos and Luís Alfredo Amaral. Mining geo-referenced data with qualitative spatial reasoning strategies. Computers and Graphics, 28(3):371– 379, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sharma:2017:SDB

[SA17] Ojaswa Sharma and Nidhi Agarwal. Signed distance based 3D surface reconstruction from unorganized planar cross-sections. Computers and Graphics, 62(??): 67–76, ???? 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301364.

Sykora:2019:FSS

[SA19] Daniel Sýkora and Tunç O. Aydin. Foreword to the special section on Expressive 2018. Computers and Graphics, 82(??):A9, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300986.

Serin:2012:APG

[SAB12] Ekrem Serin, Serdar Hasan Adali, and Selim Balcisoy. Automatic path generation for terrain navigation. Computers and Graphics, 36(8): 1013–1024, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001422.

Sacht:2022:SAB

[Sac22] Leonardo Sacht. Structureaware bottle cap art. Computers and Graphics, 107 (??):277–288, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001510.

Sahillioglu:2015:SDA

[Sah15]  Yusuf Sahillioglu.  A

shape deformation algorithm for constrained multidimensional scaling. Computers and Graphics, 53 (part B)(??):156–165, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001648.

Suzuki:1990:GCR

[SAK90] Hiromasa Suzuki, Hidetoshi Ando, and Fumihiko Kimura. Geometric constraints and reasoning for geometrical CAD systems. Computers and Graphics, 14(2): 211–224, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sakas:2002:TMI

[Sak02] Georgios Sakas. Trends in medical imaging: from 2D to 3D. Computers and Graphics, 26(4):577–587, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/34/abstract.
html.

Sidell:1975:CGS

[SAKB75] P. M. Sidell, D. U. Anderson, T. J. Knopp, and J. B. Bassingthwaighte. Computer graphics in simulation of cardiovascular transport phenomena. Computers and Graphics, 1(2–3): 289–292, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Salzman:1985:ECC

[Sal85] Roy M. Salzman. The evolution from CAD/CAM to CIM: Possibilities, problems and strategies for the future. Computers and Graphics, 9(4):435–439, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Samet:1989:IRT

[Sam89] Hanan Samet. Implementing ray tracing with octrees and neighbor finding. Computers and Graphics, 13(4):445–460, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). includes code.

Sarfraz:1997:PMS

[SAMA97] Muhammad Sarfraz, Muhammed Al-Mulhem, and Farooq Ashraf. Preserving monotonic shape of the data using piece-wise rational cubic functions. Computers and Graphics, 21(1):5– 14, January–February 1997.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600065.

Sandor:1985:ODS

[San85] J. Sandor. Octree data structures and perspective imagery. Computers and Graphics, 9(4):393–405, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Santos:1993:CCR

[San93] Adelino Santos. CoMEdiA: Conceptualisation and realisation of a cooperative hypermedia editing architecture. Computers and Graphics, 17(3):261– 268, May–June 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Santos:2000:ICV

[San00] Beatriz Sousa Santos. An introductory course on Visualization. Computers and Graphics, 24(1):163– 169, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
42/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/13/20/47/27/42/article.
                                        pdf.

Santos:2006:ICH

[San06] Beatriz Sousa Santos. An introductory course on humancomputer interaction: Programme, bibliography, practical classes and assignments. Computers and Graphics, 30(4):658–668, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000872.

Sarfraz:1992:RCS

[Sar92a] Muhammad Sarfraz. A C^{2} rational cubic spline alternative to the nurbs. Computers and Graphics, 16 (1):69–77, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:1992:IRC

[Sar92b] Muhammad Sarfraz. Interpolatory rational cubic spline with biased, point and interval tension. Computers and Graphics, 16 (4):427–430, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:1993:DCS

[Sar93] Muhammad Sarfraz. Designing of curves and surfaces using rational cubics.

Computers and Graphics, 17(5):529–538, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:1994:CSC

[Sar94a] Muhammad Sarfraz. Cubic spline curves with shape control. Computers and Graphics, 18(5):707–713, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:1994:GGI

[Sar94b] Muhammad Sarfraz. Generalized geometric interpolation for rational cubic splines. Computers and Graphics, 18(1):61– 72, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:2000:RCS

[Sar00] M. Sarfraz. A rational cubic spline for the visualization of monotonic data. Computers and Graphics, 24 (4):509–516, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/28/article.
pdf.

Sarfraz:2002:LES

[Sar02] M. Sarfraz. Letter to the editor: Some remarks on a rational cubic spline for the visualization of monotonic data. Computers and Graphics, 26(1):193– 197, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/44/abstract.
                                        html.

Sarfraz:2003:RCS

[Sar03] M. Sarfraz. A rational cubic spline for the visualization of monotonic data: an alternate approach. Computers and Graphics, 27 (1):107–121, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarfraz:2004:WNS

[Sar04] Muhammad Sarfraz. Weighted Nu splines with local support basis functions. Computers and Graphics, 28 (4):539–549, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sasmor:2004:FFR

[Sas04] Joshua C. Sasmor. Fractals for functions with rational exponent. Computers and Graphics, 28(4):601– 615, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sawchuk:1977:IGA

[SB77] W. Sawchuk and H. G.

Bown. Interactive graphics applied to symbol communication for non-speaking children. Computers and Graphics, 2(4):201–204, 1977. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Spruth:1983:PT

[SB83] G. W. Spruth and H. G.

Bahr. Printing technologies. Computers and Graphics, 7(1):51–58, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Slater:1984:GIP

[SB84] M. Slater and R. J. Baker.

GRAPH — an interactive program based on the graphical kernel system. Computers and Graphics, 8(2): 135–140, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Segenreich:1986:ONG

[SB86] Solly Andy Segenreich and Leda Maria P. Faria Braga. Optimal nesting of general plane figures: a Monte Carlo heuristical approach. Computers and Graphics, 10(3): 229–237, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schwarz:1994:REI

[SB94] A. J. Schwarz and B. G. Blundell. Regions of extreme image distortion in rotatingscreen volumetric display systems. Computers and Graphics, 18(5):643–652, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sakas:1996:GEI

[SB96] Georgios Sakas and Peter Bono. Guest Editors’ introduction. Computers and Graphics, 20(6):759–??, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Srinivasan:1997:HVE

[SB97] Mandayam A. Srinivasan and Cagatay Basdogan. Haptics in virtual environments: taxonomy, research status, and challenges. Computers and Graphics, 21(4):393–404, July–August 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/free/1997/21/
                                        4/9700030.pdf.

Strack:1994:ICO

[SBD^{+}94] Rüdiger Strack, Christof Blum, David Duce, Dale Sutcliffe, Narciso García, María J. Pérez-Luque, Eckhard Moeller, and Hauke Peyn. Image communication open architecture. Computers and Graphics, 18(1):21– 34, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Skytt:2015:LRS

[SBD15a] Vibeke Skytt, Oliver Barrowclough, and Tor Dokken. Locally refined spline surfaces for representation of terrain data. Computers and Graphics, 49(??): 58–68, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000308.

Srikanth:2015:CRI

[SBD15b] Manohar Srikanth, Kavita Bala, and Frédo Durand. Computational rim illumination of dynamic subjects using aerial robots. Computers and Graphics, 52(??): 142–154, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500031X.

Sbert:1998:RWR

[Sbe98] Mateu Sbert. Random walk radiosity with infinite path length. Computers and Graphics, 22(2–3):161–166, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/536.pdf.

Shi:2020:PSS

[SBE20] Peiteng Shi, Markus Billeter, and Elmar Eisemann.

SalientGaze:  Saliency-

based gaze correction in virtual reality. Computers and Graphics, 91(??): 83–94, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300972.

Schmidt:1999:TMC

[SBG99] Albrecht Schmidt, Michael Beigl, and Hans-W. Gellersen. There is more to context than location. Computers and Graphics, 23(6): 893–901, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
44/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/44/article.
pdf.

Sarfraz:2001:VSD

[SBH01] M. Sarfraz, S. Butt, and M. Z. Hussain. Visualization of shaped data by a rational cubic spline interpolation. Computers and Graphics, 25(5):833– 845, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/37/abstract.
html.

Stoddard:2007:TDI

[SBH07] Jacob Stoddard, R. Daniel Bergeron, and Donald House. Tangent driven interpolative subdivision. Computers and Graphics, 31(5):737– 749, October 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001501.

Sharma:2022:DEA

[SBHC22] Vandit Sharma, Kaushal Kumar Bhagat, Huai-Hsuan Huang, and Nian-Shing Chen. The design and evaluation of an AR-based serious game to teach programming. Computers and Graphics, 103(??):1–18, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000024.

Steinicke:2010:GTT

[SBHS10] Frank Steinicke, Gerd Bruder, Klaus Hinrichs, and Anthony Steed. Gradual transitions and their effects on presence and distance estimation. Computers and Graphics, 34 (1):26–33, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001551.

Sachdeva:2023:LTS

[SBKB23] Madhav Sachdeva, Jan Burmeister, Jörn Kohlhammer, and Jürgen Bernard. LFPeers: Temporal similarity search and result exploration. Computers and Graphics, 115(??):81– 95, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001048.

Sawchuk:1978:IIC

[SBOT78] W. Sawchuk, H. G. Bown, C. D. O’Brien, and G. W. Thorgeirson. An interactive image communication system using narrowband lines. Computers and Graphics, 3 (4):129–134, 1978. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sugathan:2022:LVE

[SBR^{+}22] Sherin Sugathan, Hauke Bartsch, Frank Riemer, Renate Grüner, Kai Lawonn, and Noeska Smit. Longitudinal visualization for exploratory analysis of multiple sclerosis lesions. Computers and Graphics, 107 (??):208–219, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001479.

Santos:2023:SFP

[SBR23] Wallas H. S. Santos, Emilio Vital Brazil, and Alberto Raposo. ShapeGraMM: On the fly procedural generation of massive models for realtime visualization. Computers and Graphics, 116(??): 239–250, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001887.

Staneker:2004:OCO

[SBS04] Dirk Staneker, Dirk Bartz, and Wolfgang Straßer. Occlusion culling in OpenSG PLUS. Computers and Graphics, 28(1):87–92, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sipiran:2013:DAP

[SBS13] Ivan Sipiran, Benjamin Bustos, and Tobias Schreck. Data-aware 3D partitioning for generic shape retrieval. Computers and Graphics, 37 (5):460–472, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000484.

Schmidt:2019:EVA

[SBS19] Susanne Schmidt, Gerd Bruder, and Frank Steinicke. Effects of virtual agent and object representation on experiencing exhibited artifacts. Computers and Graphics, 83(??):1–10, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300962.

Sorgente:2022:PKC

[SBS22] Tommaso Sorgente, Silvia Biasotti, and Michela Spagnuolo. Polyhedron kernel computation using a geometric approach. Computers and Graphics, 105(??): 94–104, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000693.

Soetebier:2004:MLP

[SBSL04] Ingo Soetebier, Horst Birthelmer, Jörg Sahm, and Volker Luckas. Managing large progressive meshes. Computers and Graphics, 28 (5):691–701, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Saleem:2011:VCS

[SBWS11] Waqar Saleem, Alexander Belyaev, Danyi Wang, and Hans-Peter Seidel. On visual complexity of 3D shapes. Computers and Graphics, 35(3):580–585, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000458.

Seversky:2011:HPC

[SBY11] Lee M. Seversky, Matt S.

Berger, and Lijun Yin. Harmonic point cloud orientation. Computers and Graphics, 35(3):492–499, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000513.

Spitaleri:1997:CAR

[SC97] Rosa Maria Spitaleri and Rossella Cossu. A comparative analysis of reference models for visual and computational integrated environments. Computers and Graphics, 21(5):537–547, September–October 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=5&aid=9700033.

Sharf:2012:FSS

[SC12] Andrei Sharf and Baoquan Chen. Foreword to special section. Computers and Graphics, 36(4): 215, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000386.

Segear:2024:VFG

[SCB^{+}24] Sydney Segear, Vuthea Chheang, Lauren Baron, Jicheng Li, Kangsoo Kim, and Roghayeh Leila Barmaki. Visual feedback and guided balance training in an immersive virtual reality environment for lower extremity rehabilitation. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000062.

Stanculescu:2011:FSM

[SCC11] Lucian Stãnculescu, Raphaëlle Chaine, and Marie-Paule Cani. Freestyle: Sculpting meshes with self-adaptive topology. Computers and Graphics, 35(3):614–622, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000720.

Stanculescu:2013:SMD

[SCCS13] Lucian Stãnculescu, Raphaëlle Chaine, Marie-Paule Cani, and Karan Singh. Sculpting multi-dimensional nested structures. Computers and Graphics, 37(6):753– 763, October 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000848.

Su:2016:CCO

[SCFF16] Xiaoyu Su, Xiaowu Chen, Qiang Fu, and Hongbo Fu. Cross-class 3D object synthesis guided by reference examples. Computers and Graphics, 54(??): 145–153, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001028.

Serrano:2023:FSS

[SCG23] Ana Serrano, Marc Comino, and Jesús Gimeno. Foreword to the special section on CEIG 2023. Computers and Graphics, 115(??): A10–A11, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002571.

Schaefer:1975:ITI

[Sch75]  L. J. Schaefer.  Imple-

mentation of transform invoking elements. Computers and Graphics, 1(1):69– 72, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schaefer:1985:MIA

[Sch85a] Thomas J. Schaefer. A modular and integratable approach to CIM. Computers and Graphics, 9(4): 441–447, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schilcher:1985:IGD

[Sch85b] M. Schilcher. Interactive graphic data processing in cartography. Computers and Graphics, 9(1):57–66, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Scheller:1986:IPT

[Sch86a] A. Scheller. Integrated processing of text and graphics within the German Research Network. Computers and Graphics, 10(2): 133–142, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schoenhut:1986:CGM

[Sch86b] J. Schoenhut. Classification of graphics metafile encodings. Computers and Graphics, 10(2):107–111, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schuster:1986:GEI

[Sch86c] R. Schuster. Guest Editor’s introduction. Computers and Graphics, 10 (4):275–276, ???? 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849386900178.

Schaal:1988:CGT

[Sch88] Hermann Schaal. Computer graphical treatments of perspective pictures. Computers and Graphics, 12 (1):15–31, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schwartz:1995:MTF

[Sch95] Lillian F. Schwartz. Morphing the three faces of Mona: The decision-making steps Leonardo used to create his Mona Lisa. Computers and Graphics, 19(4): 529–539, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500031.

Schneider:1998:DSB

[Sch98] Uwe Schneider. DaType: a stroke-based typeface design system. Computers and Graphics, 22(4):515–526, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
4/583.pdf.

Schlei:2012:VES

[Sch12]  B. R. Schlei.  Volume-

enclosing surface extraction. Computers and Graphics, 36(2):111–130, April 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001816.

Su:2018:PSA

[SCH^{+}18] Zhilong Su, Lujie Chen, Xiaoyuan He, Fujun Yang, and Lawrence Sass. Planar structures with automatically generated bevel joints. Computers and Graphics, 72(??):98–105, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300256.

Sims-Coomber:1994:ILE

[SCM94] Helen Sims-Coomber and Ralph Martin. An implementation of LOGO for elliptic geometry. Computers and Graphics, 18(4): 543–552, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sims-Coomber:1991:NEI

[SCMT91] Helen Sims-Coomber, Ralph Martin, and Michael Thorne. A non-Euclidean implementation of LOGO. Computers and Graphics, 15(1): 117–130, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Spehar:2003:UAF

[SCNT03] Branka Spehar, Colin W. G.

Clifford, Ben R. Newell, and Richard P. Taylor. Universal aesthetic of fractals. Computers and Graphics, 27 (5):813–820, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Scowen:1984:CTS

[Sco84]  R. S. Scowen.  Confor-

mance testing of software. Computers and Graphics, 8 (1):5–12, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sujar:2018:RTA

[SCSG18] Aaron Sujar, Juan Jose Casafranca, Antoine Serrurier, and Marcos Garcia. Real-time animation of human characters’ anatomy. Computers and Graphics, 74 (??):268–277, August 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300906.

Sundar:2014:FDM

[SCT^{+}14] Bharath Ram Sundar, Abhijith Chunduru, Rajat Tiwari, Ashish Gupta, and Ramanathan Muthuganapathy. Footpoint distance as a measure of distance computation between curves and surfaces. Computers and Graphics, 38 (??):300–309, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001738.

Sampaio:2016:TCS

[SCVCN16] Artur P. Sampaio, Raphaëlle Chaine, Creto A. Vidal, and Joaquim B. Cavalcante-Neto. Temporally coherent sculpture of composite objects. Computers and Graphics, 58(??):118– 127, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300541.

St-Denis:1990:LDK

[SD90] Richard St-Denis. LGV:

a domain knowledge validation environment. Computers and Graphics, 14(2): 311–320, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sezgin:2008:SRI

[SD08] T. M. Sezgin and R. Davis.

Sketch recognition in interspersed drawings using timebased graphical models. Computers and Graphics, 32 (5):500–510, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000575.

Semmo:2015:IIF

[SD15] Amir Semmo and Jürgen Döllner. Interactive image filtering for level-of-abstraction texturing of virtual 3D scenes. Computers and Graphics, 52(??): 181–198, November 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000096.

Spagnuolo:2016:FSS

[SD16] Michela Spagnuolo and Roman Durikovic. Foreword to the special section on the Spring Conference on Computer Graphics 2016 (SCCG 2016). Computers and Graphics, 59(??): A1, October 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300875.

Scott:2021:EBT

[SD21] Joshua J. Scott and Neil A. Dodgson. Example-based terrain synthesis with pit removal. Computers and Graphics, 99(??):43–53, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001278.

Shen:1995:PIR

[SDD95] Li-Sheng Shen, Ed F. Deprettere, and P. Dewilde. A parallel image-rendering algorithm and architecture based on ray tracing and radiosity shading. Computers and Graphics, 19(2): 281–296, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400154.

Simari:2013:GED

[SDIM13] Patricio Simari, Leila De Floriani, Federico Iuricich, and Mohammed M. Mesmoudi. Generalized extrinsic distortion and applications. Computers and Graphics, 37 (6):582–588, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000927.

Slater:1989:LRT

[SDS89] Mel Slater, Allan Davison, and Mark Smith. Liberation from rectangles: a tiling method for dynamic modification of objects on raster displays. Computers and Graphics, 13(1):83–89, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Silva:2019:TRT

[SdSR^{+}19] Luiz José Schirmer Silva, Djalma Lúcio Soares da Silva, Alberto Barbosa Raposo, Luiz Velho, and Hélio Côrtes Vieira Lopes. Tensorpose: Real-time pose estimation for interactive applications. Computers and Graphics, 85(??):1– 14, December 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301414.

Saiti:2021:CTR

[SDT21] Evdokia Saiti, Antonios Danelakis, and Theoharis Theoharis. Cross-time registration of 3D point clouds. Computers and Graphics, 99 (??):139–152, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001357.

Sommer:1999:WIV

[SDWE99] Ove Sommer, Alexander Dietz, Rüdiger Westermann, and Thomas Ertl. WSCG ’98 — an interactive visualization and navigation tool for medical volume data. Computers and Graphics, 23(2):233– 244, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/676.pdf.

Surazhsky:2001:MFF

[SE01] Tatiana Surazhsky and Gershon Elber. Matching freeform surfaces. Computers and Graphics, 25(1):3– 12, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/27/article.
pdf.

Shi:1990:ELC

[SEC90] Kaijian Shi, J. A. Edwards, and D. C. Cooper. An efficient line clipping algorithm. Computers and Graphics, 14 (2):297–301, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sanchez-Elez:2003:AOM

[SEDT^{+}03] Marcos Sanchez-Elez, Haitao Du, Nozar Tabrizi, Yun Long, Nader Bagherzadeh, and Milagros Fernandez. Algorithm optimizations and mapping scheme for interactive ray tracing on a reconfigurable architecture. Computers and Graphics, 27 (5):701–713, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Segenreich:1988:HBL

[Seg88] Solly A. Segenreich. How to build a lattice nesting in any arbitrary direction. Computers and Graphics, 12 (2):255–259, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Seibert:1999:ARU

[Sei99] Frank Seibert. Augmenting reality by using uncalibrated optical tracking. Computers and Graphics, 23 (6):801–804, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/31/article.
pdf.

Subramanian:2019:DLB

[SEKA19] Sai Ganesh Subramanian, Mathew Eng, Vinayak R. Krishnamurthy, and Ergun Akleman. Delaunay Lofts: a biologically inspired approach for modeling space filling modular structures. Computers and Graphics, 82(??):73–83, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300822.

Seldom:1989:ICP

[Sel89] Harry Seldom. Iteration as a creative process in visual art. Computers and Graphics, 13(3):397–??, 1989. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Song:2005:UCC

[SEMWC05] Meehae Song, Thomas Elias, Wolfgang Mueller-Wittig, and Tony K. Y. Chan. Using the Chinese Calligraphy brush as a tangible user interface tool in virtual heritage scenarios. Computers and Graphics, 29 (1):41–48, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sen:1998:CGP

[Sen98] Asok K. Sen. Chaos and graphics — a Product-Delay algorithm for graphic design. Computers and Graphics, 22(6):759–764, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/630.pdf.

Sen:1999:CGP

[Sen99] Asok K. Sen. Chaos and graphics — the productdelay algorithm: Graphic design with amplitude- and frequency-modulated waveforms. Computers and Graphics, 23(1):169–174, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
1/658.pdf.

Sen:2000:MP

[Sen00] Asok K. Sen. Moiré patterns. Computers and Graphics, 24(3):471–475, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/42/article.
pdf.

Sen:2003:TDA

[Sen03] Asok K. Sen. A threedimensional approach to graphic design. Computers and Graphics, 27(2):303– 310, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sequin:2013:MGT

[Séq13] Carlo H. Séquin. Making graphics tangible. Computers and Graphics, 37(3): 148–164, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000137.

Siska:1988:SMS

[SET^{+}88] J. Carlos Siska, Ricardo H.

Espinosa, David A. Trevino, Junipero Rodriguez, Juan D. Sanchez, and Hector Ramirez. 3D solid modeling software development for industrial and academic purposes. Computers and Graphics, 12 (3–4):381–389, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sezgin:2016:FSS

[Sez16] Tevfik Metin Sezgin. Foreword to the special section on Expressive 2015. Computers and Graphics, 59(??):A2, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300784.

Suffern:1991:IMC

[SF91] Kevin G. Suffern and Edward D. Fackerell. Interval methods in computer graphics. Computers and Graphics, 15(3):331–340, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Stytz:1992:VPB

[SF92] Martin R. Stytz and Ophir Frieder. Volume-primitive based three-dimensional medical image rendering. customized architectural approaches. Computers and Graphics, 16(1):85–100, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Segura:1998:TSA

[SF98] Rafael J. Segura and Francisco R. Feito. Technical section — an algorithm for determining intersection segment-polygon in 3D. Computers and Graphics, 22(5):587–592, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/598.pdf.

Szwarcman:2001:GOD

[SFC01] D. Szwarcman, B. Feijó, and M. Costa. Goaloriented dead reckoning for autonomous characters. Computers and Graphics, 25(6):999–1011, December 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/35/33/abstract.
html.

Storz:2006:SCS

[SFD06] Oliver Storz, Adrian Friday, and Nigel Davies. Supporting content scheduling on situated public displays. Computers and Graphics, 30 (5):681–691, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001191.

Schatz:2021:ASP

[SFS^{+}21] Karsten Schatz, Florian Frieß, Marco Schäfer, Patrick C. F. Buchholz, Jürgen Pleiss, Thomas Ertl, and Michael Krone. Analyzing the similarity of protein domains by clustering Molecular Surface Maps. Computers and Graphics, 99 (??):114–127, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001229.

Sanchez:2013:MSG

[SFVP13] Mathieu Sanchez, Oleg Fryazinov, Turlif Vilbrandt, and Alexander Pasko. Morphological shape generation through user-controlled group metamorphosis. Computers and Graphics, 37 (6):620–627, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000836.

Sakas:1992:SAA

[SG92] Georgios Sakas and Matthias Gerth. Sampling and antialiasing of discrete 3-D volume density textures. Computers and Graphics, 16(1): 121–134, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shoja:2015:GBP

[SG15] Ehsan Shoja and Mohammad Ghodsi. GPU-based parallel algorithm for computing point visibility inside simple polygons. Computers and Graphics, 49 (??):1–9, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000242.

Spierling:2002:SSP

[SGBI02] Ulrike Spierling, Dieter Grasbon, Norbert Braun, and Ido Iurgel. Setting the scene: playing digital director in interactive storytelling and creation. Computers and Graphics, 26(1):31–44, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/31/abstract.
                                        html.

Saalfeld:2017:FFF

[SGBP17] Patrick Saalfeld, Sylvia Glaßer, Oliver Beuing, and Bernhard Preim. The FAUST framework: Freeform annotations on unfolding vascular structures for treatment planning. Computers and Graphics, 65(??): 12–21, June 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300389.

Schmidt:2000:CVV

[SGC00] Ana Elisa F. Schmidt, Marcelo Gattass, and Paulo Cezar P. Carvalho. Combined 3D visualization of volume data and polygonal models using a Shear-Warp algorithm. Computers and Graphics, 24 (4):583–601, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/35/article.
pdf.

Sarcinelli:2019:HPS

[SGC^{+}19] Renan Sarcinelli, Rânik Guidolini, Vinicius B. Cardoso, Thiago M. Paixão, Rodrigo F. Berriel, Pedro Azevedo, Alberto F. De Souza, Claudine Badue, and Thiago Oliveira-Santos. Handling pedestrians in self-driving cars using image tracking and alternative path generation with Frenét frames. Computers and Graphics, 84(??): 173–184, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301323.

Stein:2012:CFP

[SGES12] Ayal Stein, Eran Geva, and Jihad El-Sana. CudaHull: Fast parallel 3D convex hull on the GPU. Computers and Graphics, 36(4):265–271, June 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000350.

Seron:2005:IMC

[SGGC05] F. J. Seron, D. Gutierrez, G. Gutierrez, and E. Cerezo. Implementation of a method of curved ray tracing for inhomogeneous atmospheres. Computers and Graphics, 29 (1):95–108, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sandler:1997:HADa

[SGM97a] Eugene A. Sandler, Dmitri A. Gusev, and Gregory Y. Milman. Hybrid algorithms for digital halftoning and their application to medical imaging. Computers and Graphics, 21(1):69– 78, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600071.

Sandler:1997:HADb

[SGM97b] Eugene A. Sandler, Dmitri A.

Gusev, and Gregory Y. Milman. Hybrid algorithms for digital halftoning and their application to medical imaging. Computers and Graphics, 21(6):859, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700103.

Salinas:2020:PRV

[SGPC20] Mario Salinas, Daniela Giorgi, Federico Ponchio, and Paolo Cignoni. ReviewerNet: a visualization platform for the selection of academic reviewers. Computers and Graphics, 89(??): 77–87, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300443.

Shikhare:1999:CGI

[SGR^{+}99] Dinesh Shikhare, S. Gopalsamy, T. Sathi Reddy, Ashwini Patgawkar, Satyashree Mahapatra, S. P. Mudur, K. P. Singh, Indira Narayanswamy, and Laxmi Ravishankar. Computer graphics in India — Zeus: Surface modeling, surface grid generation, tetrahedral volume discretization. Computers and Graphics, 23(1):59–72, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        1/650.pdf.

Shirsat:1999:CGI

[SGS99] Amit Shirsat, Sandeep Gupta, and Gopal R. Shevare. Computer graphics in India — generation of multi-block topology for discretisation of threedimensional domains. Computers and Graphics, 23(1): 45–57, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1999/
                                        23/1/649.pdf.

Sun:2010:ABR

[SGZ10] Rui Sun, Shuming Gao, and Wei Zhao. An approach to B-rep model simplification based on region suppression.

Computers and Graphics, 34 (5):556–564, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001032.

Semwal:1994:BMI

[SH94] Sudhanshu K. Semwal and John J. Hallauer. Biomechanical modeling: Implementing line-of-action algorithm for human muscles and bones using generalized cylinders. Computers and Graphics, 18(1): 105–112, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sterzbach:1996:MVB

[SH96] Bernhard Sterzbach and Wolfgang A. Halang. A mobile vehicle on-board computing and communication system. Computers and Graphics, 20(5):659–667, September–October 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=5&aid=9600040.

Sherman:2003:DMR

[SH03] Paul Sherman and John C.

Hart. Direct manipulation of recurrent models.

Computers and Graphics, 27 (1):143–151, February 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schaefer:2012:FSM

[SH12] Scott Schaefer and John C. Hart. Foreword to Shape Modeling International 2012. Computers and Graphics, 36 (5):305, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000908.

Stanko:2016:SCN

[SHBSS16] Tibor Stanko, Stefanie Hahmann, Georges-Pierre Bonneau, and Nathalie Saguin-Sprynski. Surfacing curve networks with normal control. Computers and Graphics, 60(??): 1–8, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300814.

Stanko:2017:SSC

[SHBSS17] Tibor Stanko, Stefanie Hahmann, Georges-Pierre Bonneau, and Nathalie Saguin-Sprynski. Shape from sensors: Curve networks on surfaces from 3D orientations. Computers and Graphics, 66(??):74–84, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300626.

Stumpfegger:2022:GAS

[SHCW22] Josef Stumpfegger, Kevin Höhlein, George Craig, and Rüdiger Westermann. GPU accelerated scalable parallel coordinates plots. Computers and Graphics, 109(??): 111–120, December 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001868.

Steed:2017:FVA

[SHD^{+}17] Chad A. Steed, William Halsey, Ryan Dehoff, Sean L. Yoder, Vincent Paquit, and Sarah Powers. Falcon: Visual analysis of large, irregularly sampled, and multivariate time series data in additive manufacturing. Computers and Graphics, 63(??):50–64, April 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300201.

Sheu:1988:OOG

[She88] Phillip Chen-Yu Sheu. Objectoriented graphics knowledge bases. Computers and Graphics, 12(1):115–123, 1988. CODEN COGRD2.

ISSN 0097-8493 (print),

1873-7684 (electronic).

Sherard:1993:JSQ

[She93] Paul K. Sherard. Julia sets and quasi-stable orbits in the complex plane. Computers and Graphics, 17(2): 175–184, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shesh:2012:HLA

[She12] Amit Shesh. High-level application development for non-computer science majors using image processing. Computers and Graphics, 36(3):170–177, May 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000088.

Su:1998:GID

[SHG98] Jonathan K. Su, Frank Hartung, and Bernd Girod. Graphics in/for digital libraries — digital watermarking of text, image, and video documents. Computers and Graphics, 22(6):687–695, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        6/623.pdf.

Shirley:1992:TCM

[Shi92] Peter Shirley. Time complexity of Monte Carlo radiosity. Computers and Graphics, 16(1):117–120, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shirriff:1993:FSP

[Shi93a] Ken Shirriff. Fractals from simple polynomial composite functions. Computers and Graphics, 17(6):701–703, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shirriff:1993:GFV

[Shi93b] Ken Shirriff. Generating fractals from Voronoi diagrams. Computers and Graphics, 17(2):165–167, March–April 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shirriff:1993:IFG

[Shi93c] Ken W. Shirriff. An investigation of fractals generated by z → z^{−n} + c. Computers and Graphics, 17(5):603–607, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shirriff:1993:RCM

[Shi93d] Ken W. Shirriff. The RBW color model. Computers and Graphics, 17(5):597–602, September–October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shi:2002:GEI

[Shi02] Jiaoying Shi. Guest editor’s introduction: Best papers of CAD & CG 2001. Computers and Graphics, 26(3):391– 392, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/27/abstract.
html.

Shirehjini:2004:NIM

[Shi04] Ali. A. Nazari Shirehjini. A novel interaction metaphor for personal environment control: direct manipulation of physical environment based on 3D visualization. Computers and Graphics, 28 (5):667–675, October 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shih:2006:RAC

[Shi06] Naai-Jung Shih. RP-aided computer modeling for architectural education. Computers and Graphics, 30 (1):137–144, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002177.

Schier:2018:FTM

[SHK18] Alexander Schier, Stefan Hartmann, and Reinhard Klein. Fast texture mapping for triangle soups using electrostatic monopole field lines. Computers and Graphics, 77(??):140– 155, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301559.

Shlien:1983:RPC

[Shl83] S. Shlien. Raster to polygon conversion of images. Computers and Graphics, 7 (3–4):327–332, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Simon:2024:DEU

[SHL^{+}24] Cassandre Simon, Lucas Herfort, Flavien Lebrun, Elsa Brocas, Samir Otmane, and Amine Chellali. Design and evaluation of UltRASim: an immersive simulator for learning ultrasound-guided regional anesthesia basic skills. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000074.

Sinclair:1989:CCO

[SHLW89] B. Sinclair, A. G. Hannam, A. A. Lowe, and W. W. Wood. Complex contour organization for surface reconstruction. Computers and Graphics, 13(3): 311–319, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Simon:2023:SCM

[SHOC23] Cassandre Simon, Manel Boukli Hacene, Samir Otmane, and Amine Chellali. Study of communication modalities to support teaching tool manipulation skills in a shared immersive environment. Computers and Graphics, 117 (??):31–41, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002340.

Szabo:2023:FAI

[SHS^{+}23] Attila Szabo, Georg Haaser, Harald Steinlechner, Andreas Walch, Stefan Maierhofer, Thomas Ortner, and M. Eduard Gröller. Featureassisted interactive geometry reconstruction in 3D point clouds using incremental region growing. Computers and Graphics, 111 (??):213–224, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000274.

Suarez:2019:ACS

[SHZ19] Alvaro Javier Fuentes Suarez, Evelyne Hubert, and Cedric Zanni. Anisotropic convolution surfaces. Computers and Graphics, 82 (??):106–116, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300792.

Strauss:2014:GMC

[SIE14] Rouven Strauss, Florin Isvoranu, and Gershon Elber. Geometric multi-covering. Computers and Graphics, 38 (??):222–229, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001593.

Sifaqui:1999:TSS

[Sif99] Christian Sifaqui. Technical section — structuring user interfaces with a meta-model of mental models. Computers and Graphics, 23(3):323– 330, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/684.pdf.

SIGGRAPH:2002:FEW

[SIG02] SIGGRAPH. Fourth Eurographics Workshop on Parallel Graphics and Visualization (pending) — in cooperation with ACM SIGGRAPH September 9–10, 2002, Blaubeuren, Germany. Computers and Graphics, 26(3): 527, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/44/abstract.
html;   http://www.gris.
uni-tuebingen.de/conf/
egpgv02.

Singleton:1987:IGS

[Sin87] Karen Singleton. An implementation of the GKS-3D/PHIGS viewing pipeline. Computers and Graphics, 11 (2):163–183, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schetinger:2017:IFD

[SIPO17] Victor Schetinger, Massimo Iuliani, Alessandro Piva, and Manuel M. Oliveira. Image forgery detection confronts image composition. Computers and Graphics, 68(??): 152–163, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301498.

Shene:1994:CIP

[SJ94] Ching-Kuang Shene and John K. Johnstone. Computing the intersection of a plane and a revolute quadric. Computers and Graphics, 18 (1):47–59, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Song:2009:MOR

[SJ09] Xinghua Song and Bert Jüttler. Modeling and 3D object reconstruction by implicitly defined surfaces with sharp features. Computers and Graphics, 33(3): 321–330, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000405.

Strodthoff:2015:LRG

[SJ15] B. Strodthoff and B. Jüttler.

Layered Reeb graphs for three-dimensional manifolds in boundary representation. Computers and Graphics, 46 (??):186–197, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001149.

Sperrle:2021:CAV

[SJB^{+}21] Fabian Sperrle, Astrik Jeitler, Jürgen Bernard, Daniel Keim, and Mennatallah El-Assady. Co-adaptive visual data analysis and guidance processes. Computers and Graphics, 100(??): 93–105, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100131X.

Stein:2019:IDC

[SJG19] Oded Stein, Alec Jacobson, and Eitan Grinspun. Interactive design of castable shapes using two-piece rigid molds. Computers and Graphics, 80 (??):51–62, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300238.

Schubert:2020:FPN

[SJT20] Herman R. Schubert, Andrei C. Jalba, and Alexandru C. Telea. Feature preserving noise removal for binary voxel volumes using 3D surface skeletons. Computers and Graphics, 87(??):30–42, April 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300017.

Song:2023:DSM

[SJZ^{+}23] Dan Song, Xue-Jing Jiang, Yue Zhang, Fang-Lue Zhang, Yao Jin, and Yun Zhang.

Domain-specific modeling and semantic alignment for image-based 3D model retrieval. Computers and Graphics, 115(??):25–34, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001280.

Shen:2007:GBI

[SJZW07] Jianbing Shen, Xiaogang Jin, Chuan Zhou, and Charlie C. L. Wang. Gradient based image completion by solving the Poisson equation. Computers and Graphics, 31 (1):119–126, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600224X.

Schilling:1998:GID

[SK98] Andreas Schilling and Reinhard Klein. Graphics in/for digital libraries — rendering of multiresolution models with texture. Computers and Graphics, 22(6):667–674, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/621.pdf.

Shirman:1999:WNL

[SK99] Leon Shirman and Yakov Kamen. WSCG ’98 — a new look at mipmap level estimation techniques. Computers and Graphics, 23(2):223– 231, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        2/675.pdf.

Sangole:2003:VRO

[SK03] Archana Sangole and George K. Knopf. Visualization of randomly ordered numeric data sets using spherical self-organizing feature maps. Computers and Graphics, 27 (6):963–976, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sovakar:2004:ADA

[SK04] Abhijit Sovakar and Leif Kobbelt. API design for adaptive subdivision schemes. Computers and Graphics, 28(1):67–72, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Stelldinger:2006:CPD

[SK06] Peer Stelldinger and Ullrich Köthe. Connectivity preserving digitization of blurred binary images in 2D and 3D. Computers and Graphics, 30(1):70– 76, February 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002116.

Soares:2012:SSV

[SK12] Luciano Pereira Soares and Judith Kelner. Special section on virtual reality in Brazil 2011. Computers and Graphics, 36 (2):49, April 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001658.

Sung:2013:FBC

[SK13] Min-Hyuk Sung and Junho Kim. Finding the Mbest consistent correspondences between 3D symmetric objects. Computers and Graphics, 37(1–2): 81–92, February/April 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001732.

Skorkovska:2016:CMM

[SK16] Vera Skorkovská and Ivana Kolingerová. Complex multi-material approach for dynamic simulations. Computers and Graphics, 56(??): 11–19, May 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300103.

Savva:2022:FVS

[SK22] Manolis Savva and Paul G. Kry. Foreword to the virtual special section on Graphics Interface 2021. Computers and Graphics, 102(??): A10–A11, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002612.

Sharma:2023:CAD

[SK23] Ritesh Sharma and Marcelo Kallmann. Computing and analyzing decision boundaries from shortest path maps. Computers and Graphics, 117(??):73–84, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002431.

Skala:1993:EAL

[Ska93] Vaclav Skala. An efficient algorithm for line clipping by convex polygon. Computers and Graphics, 17(4): 417–421, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Skala:1994:LCA

[Ska94] Vaclav Skala. O(lg N) line clipping algorithm in E^{2}.

Computers and Graphics, 18(4):517–524, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Skala:1996:LCP

[Ska96] Václav Skala. Line clipping in E^{2} with O(1) processing complexity. Computers and Graphics, 20(4): 523–530, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=9600024.

Skala:1997:FAL

[Ska97]  Václav Skala.  A fast

algorithm for line clipping by convex polyhedron in E^{3}. Computers and Graphics, 21(2): 209–214, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600084.

Skala:1998:GEI

[Ska98] Vaclav Skala. Guest Editor’s introduction. Computers and Graphics, 22(2– 3):151–??, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Skala:1999:WGE

[Ska99] Vaclav Skala. WSCG ’98 — Guest Editors’ introduction. Computers and Graphics, 23 (2):191–??, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Skala:2008:BCC

[Ska08] Vaclav Skala. Barycentric coordinates computation in homogeneous coordinates. Computers and Graphics, 32 (1):120–127, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930700180X.

Szirmay-Kalos:1999:WID

[SKCP99] László Szirmay-Kalos, Balázs Csébfalvi, and Werner Purgathofer. WSCG ’98 — importance driven quasirandom walk solution of the rendering equation. Computers and Graphics, 23(2):203– 211, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        2/673.pdf.

Spur:1983:BG

[SKH83] G. Spur, F.-L. Krause, and H. Hoffmann. Baustein GEOMETRIE. Computers and Graphics, 7(2):199–201, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schirski:2005:VTE

[SKH^{+}05] Marc Schirski, Torsten Kuhlen, Martin Hopp, Philipp Adomeit, Stefan Pischinger, and Christian Bischof. Virtual Tubelets —efficiently visualizing large amounts of particle trajectories. Computers and Graphics, 29 (1):17–27, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Steinberger:2012:RPU

[SKH^{+}12] Markus Steinberger, Bernhard Kainz, Stefan Hauswiesner, Rostislav Khlebnikov, Denis Kalkofen, and Dieter Schmalstieg. Ray prioritization using stylization and visual saliency. Computers and Graphics, 36 (6):673–684, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000854.

Sakamoto:2010:IPB

[SKKN10] Noahisa Sakamoto, Takuma Kawamura, Koji Koyamada, and Kazunori Nozaki. Improvement of particle-based volume rendering for visualizing irregular volume data sets. Computers and Graphics, 34(1):34– 42, February 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001538.

Suma:2013:AUI

[SKL^{+}13] Evan A. Suma, David M. Krum, Belinda Lange, Sebastian Koenig, Albert Rizzo, and Mark Bolas. Adapting user interfaces for gestural interaction with the flexible action and articulated skeleton toolkit. Computers and Graphics, 37(3): 193–201, May 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001756.

Szirmay-Kalos:1998:ACW

[SKM98] László Szirmay-Kalos and Gábor Márton. Analysis and construction of worstcase optimal ray shooting algorithms. Computers and Graphics, 22(2–3):167–174, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/537.pdf.

Shirai:1983:ADM

[SKO83] Y. Shirai, K. Koshikawa, and M. Oshima. Application of 3-D models to computer vision. Computers and Graphics, 7(3–4):269–275, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szirmay-Kalos:1999:WGR

[SKP99] László Szirmay-Kalos and Werner Purgathofer. WSCG ’98 — global ray-bundle tracing with infinite number of rays. Computers and Graphics, 23(2):193– 202, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/672.pdf.

Sunden:2015:MVI

[SKR15] Erik Sundén, Sathish Kottravel, and Timo Ropinski. Multimodal volume illumination. Computers and Graphics, 50(??):47– 60, August 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000515.

Shen:2009:FML

[SKS09] Li Shen, Sungeun Kim, and Andrew J. Saykin. Fourier method for largescale surface modeling and registration. Computers and Graphics, 33(3):299– 311, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000223.

Szirmay-Kalos:2017:ISM

[SKS17] László Szirmay-Kalos and László Szécsi. Improved stratification for Metropolis light transport. Computers and Graphics, 68 (??):11–20, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301255.

Shinagawa:1995:MCT

[SKSI95] Yoshihisa Shinagawa, Tosiyasu L. Kunii, Hideyuki Sato, and Masumi Ibusuki. Modeling contact of two complex objects, with an application to characterizing dental articulations. Computers and Graphics, 19(1):21– 28, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1&aid=9400118.

Schill:1999:TAA

[SKSZ99] Alexander Schill, Sascha Kümmel, Thomas Springer, and Thomas Ziegert. Two approaches for an adaptive multimedia transfer service for mobile environments. Computers and Graphics, 23 (6):849–856, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/39/article.
pdf.

Stephenson:2001:RLL

[SL01] Peter Stephenson and Bruce Litow. Running the line: Line drawing using runs and runs of runs. Computers and Graphics, 25(4):681– 690, August 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/33/41/abstract.
html.

Spirkovska:2002:SAA

[SL02] Lilly Spirkovska and Suresh K.

Lodha. Systems: AWE: aviation weather data visualization environment. Computers and Graphics, 26(1):169– 191, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/43/abstract.
html.

Shim:2012:ACR

[SL12a] Hyunjung Shim and Seungkyu Lee. Automatic color realism enhancement for computer generated images. Computers and Graphics, 36 (8):966–973, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001471.

Silk:2012:HDR

[SL12b] Simon Silk and Jochen Lang. High dynamic range image deghosting by fast approximate background modelling. Computers and Graphics, 36(8):1060–1071, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001483.

Stahovich:2016:EDM

[SL16a] Thomas F. Stahovich and Hanlung Lin. Enabling data mining of handwritten coursework. Computers and Graphics, 57(??): 31–45, June 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300012.

Sterling:2016:IMI

[SL16b] Auston Sterling and Ming C. Lin. Integrated multimodal interaction using texture representations. Computers and Graphics, 55(??): 118–129, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001715.

Skola:2018:EVE

[SL18] Filip Skola and Fotis Liarokapis.

Embodied VR environment facilitates motor imagery brain-computer interface training. Computers and Graphics, 75(??): 44–58, October 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931830089X.

Slater:1992:ASI

[Sla92] Mel Slater. An algorithm to support 3D interaction on relatively low performance graphics systems. Computers and Graphics, 16(3):311– 315, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Steed:2009:EI

[SLCN09] Anthony Steed, Ming C.

Lin, and Carolina Cruz-Neira. Editor’s introduction. Computers and Graphics, 33 (1):21–22, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001519.

Sun:2024:ERS

[SLF^{+}24] Shilei Sun, Ming Liu, Zhongyi Fan, Qingliang Jiao, Yuxue Liu, Liquan Dong, and Lingqin Kong. Efficient ray sampling for radiance fields reconstruction. Computers and Graphics, 118 (??):48–59, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002868.

Sun:1997:LMG

[SLG97] Haijian Sun, Lin Liu, and Aike Guo. Logistic map graph set. Computers and Graphics, 21(1): 89–103, January–February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
cag/cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600073.

Shao:2023:HTM

[SLGQ23] Mingwen Shao, Fukang Liu, Zihao Guo, and Yuanjian Qiao. Hairstyle transfer via manipulating decoupled latent codes of StyleGAN2. Computers and Graphics, 116(??):363– 372, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002261.

Sterzik:2023:EMV

[SLK^{+}23] Anna Sterzik, Nils Lichtenberg, Michael Krone, Daniel Baum, Douglas W. Cunningham, and Kai Lawonn. Enhancing molecular visualization: Perceptual evaluation of line variables with application to uncertainty visualization. Computers and Graphics, 114(??):401– 413, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300095X.

Semmo:2016:ISI

[SLKD16] Amir Semmo, Daniel Limberger, Jan Eric Kyprianidis, and Jürgen Döllner. Image stylization by interactive oil paint filtering. Computers and Graphics, 55(??):157–171, April 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315002083.

She:2017:LFL

[SLL^{+}17] Jiangfeng She, Jianlong Liu, Chuang Li, Jiaqi Li, and Qiujun Wei. A line-feature label placement algorithm for interactive 3D map. Computers and Graphics, 67 (??):86–94, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300766.

Sipiran:2021:PSR

[SLL^{+}21] Ivan Sipiran, Patrick Lazo, Cristian Lopez, Milagritos Jimenez, Nihar Bagewadi, Benjamin Bustos, Hieu Dao, Shankar Gangisetty, Martin Hanik, Ngoc-Phuong Ho-Thi, Mike Holenderski, Dmitri Jarnikov, Arniel Labrada, Stefan Lengauer, Roxane Licandro, Dinh-Huan Nguyen, Thang-Long Nguyen-Ho, Luis A. Perez Rey, Bang-Dang Pham, Minh-Khoi Pham, Reinhold Preiner, Tobias Schreck, Quoc-Huy Trinh, Loek Tonnaer, Christoph von Tycowicz, and The-Anh Vu-

Le.  SHREC 2021: Re-

trieval of cultural heritage objects. Computers and Graphics, 100(??):1– 20, November 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001412.

Song:2023:ISR

[SLL^{+}23] Xiaogang Song, Wanbo Liu, Li Liang, Weiwei Shi, Guo Xie, Xiaofeng Lu, and Xinhong Hei. Image super-resolution with multiscale fractal residual attention network. Computers and Graphics, 113(??): 21–31, June 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000444.

Sohn:2022:AEE

[SLM^{+}22] Samuel S. Sohn, Mihee Lee, Seonghyeon Moon, Gang Qiao, Muhammad Usman, Sejong Yoon, Vladimir Pavlovic, and Mubbasir Kapadia. A2X: an endto-end framework for assessing agent and environment interactions in multimodal human trajectory prediction. Computers and Graphics, 106(??):130– 140, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000826.

Si:2019:VNF

[SLQ^{+}19] Weixin Si, Xiangyun Liao, Yinling Qian, Qiong Wang, and Pheng-Ann Heng. Versatile numerical fractures removal for SPH-based free surface liquids. Computers and Graphics, 81(??): 1–8, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931930038X.

Sutherland:2016:FDI

[SLRP16] Craig J. Sutherland, Andrew Luxton-Reilly, and Beryl Plimmer. Freeform digital ink annotations in electronic documents: a systematic mapping study. Computers and Graphics, 55(??): 1–20, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001818.

Schmittler:2003:VMA

[SLS03] Jörg Schmittler, Alexander Leidinger, and Philipp Slusallek. A virtual memory architecture for realtime ray tracing hardware. Computers and Graphics, 27 (5):693–699, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shi:2016:DDC

[SLX^{+}16] Yifei Shi, Pinxin Long, Kai Xu, Hui Huang, and Yueshan Xiong. Data-driven contextual modeling for 3D scene understanding. Computers and Graphics, 55(??): 55–67, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315002009.

Shum:1997:SRO

[SLYY97] Simon S. P. Shum, W. S.

Lau, Matthew M. F. Yuen, and K. M. Yu. Solid reconstruction from orthographic opaque views using incremental extrusion. Computers and Graphics, 21(6):787–800, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700058.

Stocker:1975:GGP

[SM75] F. R. Stocker and T. Minsker.

Graphics geometric perception and communications. Computers and Graphics, 1(2–3):161–174, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sharma:1992:LCR

[SM92] N. C. Sharma and S. Manohar.

Line clipping revisited. two efficient algorithms based on simple geometric observations. Computers and Graphics, 16(1):51–54, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sethia:1998:TSM

[SM98] Saurabh Sethia and S. Manohar.

Technical section — Minkowski operators for voxel based sculpting. Computers and Graphics, 22(5):593–600, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/599.pdf.

Schneider:1999:AFG

[SM99] Bengt-Olaf Schneider and Ioana M. Martin. An adaptive framework for 3D graphics over networks. Computers and Graphics, 23(6): 867–874, December 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/41/article.
pdf.

Sarfraz:2007:COP

[SM07] M. Sarfraz and A. Masood. Capturing outlines of planar images using Bézier cubics. Computers and Graphics, 31 (5):719–729, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001331.

Sasaki:2022:SIL

[SM22] Kosuke Sasaki and Jun Mitani. Simple implementation and low computational cost simulation of curved folds based on rulingaware triangulation. Computers and Graphics, 102 (??):213–219, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002041.

Settgast:2004:ATS

[SMFF04] Volker Settgast, Kerstin Müller, Christoph Fünfzig, and Dieter Fellner. Adaptive tesselation of subdivision surfaces. Computers and Graphics, 28(1):73–78, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Silberstein:1977:NIE

[SMG77] Peter T. Silberstein, David N.

Mastronarde, and Stephen A. George. Nerve impulse equations: Computer solutions and displays. Computers and Graphics, 2(3): 129–133, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smith:1975:EPA

[Smi75] L. B. Smith. An example of a pragmatic approach to portable interactive graphics. Computers and Graphics, 1(1):49–54, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smith:1977:UKP

[Smi77] P. R. Smith. U.K. Project in Computer Assisted Learning in Engineering Science. Computers and Graphics, 2 (3):151–154, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smith:1978:CCE

[Smi78] P. R. Smith. CAL is CAD ED. Computers and Graphics, 3(4):141–143, 1978. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smith:1993:PPC

[Smi93] Stuart Smith. Parallel processing for computer graphics. Computers and Graphics, 17(4):495–??, July– August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schnabel:2008:FVQ

[SMK08] Ruwen Schnabel, Sebastian Möser, and Reinhard Klein. Fast vector quantization for efficient rendering of compressed point-clouds. Computers and Graphics, 32(2): 246–259, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000186.

Sundar:2020:UAT

[SMM20] Bharath Ram Sundar, Manoj Kumar Mukundan, and Ramanathan Muthuganapathy. A unified approach towards computing Voronoi diagram, medial axis, Delaunay graph and α-hull of planar closed curves using touching discs. Computers and Graphics, 89(??):131–143, June 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300613.

Silva:2022:FR

[SMMG22] Paula Alexandra Silva, Luís Gonzaga Magalhães, Daniel Mendes, and Andrea Giachetti. Foreword RAGI. Computers and Graphics, 109(??):A5– A6, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002072.

Sanna:2001:VVF

[SMMS01] A. Sanna, B. Montrucchio, P. Montuschi, and A. Sparavigna. Visualizing vector fields: the thick oriented stream-line algorithm (TOSL). Computers and Graphics, 25(5):847– 855, October 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/57/34/38/abstract.
html.

Smolka:2003:NRA

[Smo03] Bogdan Smolka. On the new robust algorithm of noise reduction in color images. Computers and Graphics, 27 (4):503–513, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Silva:2009:EPI

[SMS09a] Samuel Silva, Joaquim Madeira, and Beatriz Sousa Santos. Erratum to “Poly-MeCo — An integrated environment for polygonal mesh analysis and comparison” [Comput. Graphics 33 (2009) 181–191]. Computers and Graphics, 33(5): 624, October 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001174.

Silva:2009:PIE

[SMS09b] Samuel Silva, Joaquim Madeira, and Beatriz Sousa Santos. PolyMeCo — an integrated environment for polygonal mesh analysis and comparison. Computers and Graphics, 33(2):181– 191, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001271.

Siegl:2013:SAS

[SMSS13] Christian Siegl, Quirin Meyer, Gerd Sußner, and Marc Stamminger. Solving aliasing from shading with selective shader supersampling. Computers and Graphics, 37(8):955– 962, December 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001271.

Sahin:2022:POU

[SMU22] Yusuf H. Sahin, Alican Mertan, and Gozde Unal. ODFNet: Using orientation distribution functions to characterize 3D point clouds. Computers and Graphics, 102(??):610–618, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001801.

Shum:2017:FSS

[SNB17] Hubert P. H. Shum, Michael Neff, and Ronan Boulic. Foreword to the special section on motion in games 2016. Computers and Graphics, 69(??):A3–A4, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301772.

Shin:2006:PBT

[SNS06a] Seunghyup Shin, Tomoyuki Nishita, and Sung Yong Shin. On pixel-based texture synthesis by nonparametric sampling. Computers and Graphics, 30 (5):767–778, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001269.

Szilvasi-Nagy:2006:GCC

[SNS06b] Márta Szilvási-Nagy and Ildikó Szabó. Generalization of Coons’ construction. Computers and Graphics, 30 (4):588–597, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000768.

Staudhammer:1975:CGH

[SO75] J. Staudhammer and Deborah J. Ogden. Computer graphics for half-tone three-dimensional object images. Computers and Graphics, 1(1):109–114, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sobhanpanah:1989:EBR

[Sob89] Cathy Sobhanpanah. Extension of a boundary representation technique for the description of m dimensional polytopes. Computers and Graphics, 13(1):17–23, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sandim:2019:BPR

[SOC^{+}19] Marcos Sandim, Nicolas Oe, Douglas Cedrim, Paulo Pagliosa, and Afonso Paiva. Boundary particle resampling for surface reconstruction in liquid animation. Computers and Graphics, 84 (??):55–65, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301396.

Schetinger:2017:HEF

[SOdSC17] Victor Schetinger, Manuel M.

Oliveira, Roberto da Silva, and Tiago J. Carvalho. Humans are easily fooled by digital images. Computers and Graphics, 68(??): 142–151, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301450.

See corrigendum [SOdSC18].

Schetinger:2018:CHE

[SOdSC18] Victor Schetinger, Manuel M. Oliveira, Roberto da Silva, and Tiago J. Carvalho. Corrigendum to “Humans are easily fooled by digital images” [Computers & Graphics 68 (2017) 142–151]. Computers and Graphics, 70(??):327, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301723.

See [SOdSC17].

Steinemann:2008:TES

[SOG08] Denis Steinemann, Miguel A. Otaduy, and Markus Gross. Tight and efficient surface bounds in meshless animation. Computers and Graphics, 32(2):235– 245, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000198.

Southard:1992:TSV

[Sou92] David A. Southard. Transformations for stereoscopic visual simulation. Computers and Graphics, 16 (4):401–410, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Southard:1993:GCG

[Sou93] David A. Southard. The Geometry of Computer Graphics by Walter F. Taylor. Computers and Graphics, 17(1):103–??, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Southard:1994:CGB

[Sou94] David A. Southard. Computers and Graphics Best Paper Award (1992). Computers and Graphics, 18 (1):1–??, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sprott:1995:AGG

[SP95] Julien C. Sprott and Clifford A. Pickover. Automatic generation of general quadratic map basins. Computers and Graphics, 19(2): 309–313, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400157.

Shi:1997:GEI

[SP97] Jiaoying Shi and Zhigeng Pan. Guest Editors’ introduction. Computers and Graphics, 21(3):273–??, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sadarjoen:2000:DQT

[SP00a] I. Ari Sadarjoen and Frits H. Post. Detection, quantification, and tracking of vortices using streamline geometry. Computers and Graphics, 24(3):333–341, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/29/article.
pdf.

Schlender:2000:MLD

[SP00b] D. Schlender and O. H. Peters. Managing levels of detail with fuzzy control. Computers and Graphics, 24(2):245–251, April 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/29/32/article.
pdf.

Sainz:2004:PBR

[SP04] Miguel Sainz and Renato Pajarola. Point-based rendering techniques. Computers and Graphics, 28(6): 869–879, December 2004.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sarov:2016:RPG

[SP16] Martin Sarov and Jörg Peters. Refinable polycube G-splines. Computers and Graphics, 58(??): 92–101, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300644.

Sforza:2023:NAE

[SP23] Davide Sforza and Fabio Pellacini. Numerical approximations for energy preserving microfacet models. Computers and Graphics, 114 (??):36–44, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000675.

Scopigno:1993:PDM

[SPGR93] R. Scopigno, A. Paoluzzi, S. Guerrini, and G. Rumolo. Parallel depth-merge: a paradigm for hidden surface removal. Computers and Graphics, 17(5):583–592, September–October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Spierling:2002:EDS

[Spi02] Ulrike Spierling. Editorial: Digital storytelling. Computers and Graphics, 26(1): 1–2, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/27/abstract.
html.

Schultz:2019:FSS

[SPK19] Thomas Schultz, Anna Puig, and Bernhard Kainz. Foreword to the special section on the Eurographics Workshop on Visual Computing for Biology and Medicine (VCBM) at Medical Image Computing and Computer Assisted Intervention (MICCAI) 2018. Computers and Graphics, 83 (??):A5–A6, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301293.

Stahovich:2014:ECB

[SPL14] Thomas F. Stahovich, Eric J. Peterson, and Hanlung Lin. An efficient, classificationbased approach for grouping pen strokes into objects. Computers and Graphics, 42(??):14–30, August 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000399.

Snelgrove:2013:PWL

[SPMA13] Xavier Snelgrove, Thiago Pereira, Wojciech Matusik, and Marc Alexa. Parallax Walls: Light fields from occlusion on height fields. Computers and Graphics, 37 (8):974–982, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300112X.

Serrano:2022:FSS

[SPO22] Ana Serrano, Jorge Posada, and Miguel Otaduy. Foreword to the special section on CEIG 2022. Computers and Graphics, 107(??): A10–A11, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001595.

Sprecher:1975:ICG

[Spr75] L. T. Sprecher. Interactive computer graphics for assisting human programmers. Computers and Graphics, 1(1):17–20, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sprott:1993:AGS

[Spr93] J. C. Sprott. Automatic generation of strange attractors.

Computers and Graphics, 17(3):325–332, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sprott:1994:AGI

[Spr94] J. C. Sprott. Automatic generation of iterated function systems. Computers and Graphics, 18(3): 417–425, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sprott:1996:SAS

[Spr96] J. C. Sprott. Strange attractor symmetric icons. Computers and Graphics, 20(2): 325–332, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500133.

Sprott:1998:ANN

[Spr98]  J. C. Sprott.  Artifi-

cial neural net attractors. Computers and Graphics, 22(1):143–149, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
1/520.pdf.

Sprott:2004:MAM

[Spr04] J. C. Sprott. A method for approximating missing data in spatial patterns. Computers and Graphics, 28 (1):113–117, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sourin:1996:URF

[SPS96] A. Sourin, A. Pasko, and V. Savchenko. Using real functions with application to hair modelling. Computers and Graphics, 20(1):11– 19, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500088.

Smith:2012:ENV

[SPS12a] J. Smith, G. Petrova, and S. Schaefer. Encoding normal vectors using optimized spherical coordinates. Computers and Graphics, 36 (5):360–365, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000568.

Smith:2012:PEC

[SPS12b] J. Smith, G. Petrova, and S. Schaefer. Progressive encoding and compression of surfaces generated from point cloud data. Computers and Graphics, 36 (5):341–348, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000714.

Sfikas:2018:EPB

[SPT18] Konstantinos Sfikas, Ioannis Pratikakis, and Theoharis Theoharis. Ensemble of PANORAMA-based convolutional neural networks for 3D model classification and retrieval. Computers and Graphics, 71 (??):208–218, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301978.

Sobkow:1987:FTD

[SPY87] Mark S. Sobkow, Paul Pospisil, and Yee-Hong Yang. A fast two-dimensional line clipping algorithm via line encoding. Computers and Graphics, 11(4): 459–467, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sanchez-Reyes:1989:LPR

[SR89] Javier Sanchez-Reyes. Laser printers for rendering surfaces. Computers and Graphics, 13(1):49–54, 1989. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Schneider:1995:MBF

[SR95] Bengt-Olaf Schneider and Jarek Rossignac. M-buffer: a flexible MISD architecture for advanced graphics. Computers and Graphics, 19(2): 239–246, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400149.

Shamir:1997:QED

[SR97] Ariel Shamir and Ari Rappoport. Quality enhancements of digital outline fonts. Computers and Graphics, 21(6):713–725, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=6&aid=9700049.

Sarfraz:2002:TSA

[SR02] M. Sarfraz and M. F. A.

Razzak. Technical section: An algorithm for automatic capturing of the font outlines. Computers and Graphics, 26(5):795–804, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/56/42/abstract.
                                        html.

Sanchez-Reyes:2015:CCC

[SR15] J. Sánchez-Reyes. Comment on the “Coincidence condition of two Bézier curves of an arbitrary degree”. Computers and Graphics, 53 (part B)(??): 166, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001673.

See rebuttal [CM15b].

Sheharyar:2019:VAR

[SRA^{+}19] Ali Sheharyar, Alexander Ruh, Maria Aristova, Michael Scott, Kelly Jarvis, Mohammed Elbaz, Ryan Dolan, Susanne Schnell, Kai Lin, James Carr, Michael Markl, Othmane Bouhali, and Lars Linsen. Visual analysis of regional myocardial motion anomalies in longitudinal studies. Computers and Graphics, 83 (??):62–76, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301219.

Santos:2008:PGV

[SRF08] Luis Paulo Santos, Dirk Reiners, and Jean Favre. Parallel graphics and visualization. Computers and Graphics, 32(1):1–2, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000022.

Sridharan:2002:ECM

[Sri02]  K. Sridharan.  Efficient

computation of a measure of depth between convex objects for graphics applications. Computers and Graphics, 26(5):785–793, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/41/abstract.
html.

Schmitz:2023:IPS

[SRZK23] Caro Schmitz, Constantin Rösch, Domenic Zingsheim, and Reinhard Klein. Interactive pose and shape editing with simple sketches from different viewing angles. Computers and Graphics, 114(??):347–356, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300119X.

Spahr:1975:ICG

[SS75a] H. R. Spahr and H. A. Sumlin. Interactive computer graphics applied to the theoretical aircraft/store separation problem. Computers and Graphics, 1(2–3): 263–270, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Steward:1975:IDS

[SS75b] D. V. Steward and M. J. Stedwell. Interactive design system to be taught 3D graphical-mathematical procedures. Computers and Graphics, 1(2–3):257–262, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Singh:1982:FSO

[SS82] S. Singh and W. R. Spillers. FEM shape optimization: a case for interactive graphics. Computers and Graphics, 6(2):63–72, 1982. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shallit:1989:TMG

[SS89a] Jeffrey Shallit and Jorge Stolfi. Two methods for generating fractals. Computers and Graphics, 13(2): 185–191, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shepard:1989:SH

[SS89b] Sherri Shepard and Andrew Simoson. Scouts in hyperspace. Computers and Graphics, 13(2): 253–260, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sheu:1993:EOO

[SS93] P. C. Y. Sheu and D. Silver.

Extending object-oriented databases with problem solving and visualization. Computers and Graphics, 17(4): 447–455, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shih:1996:GMA

[SS96] Naai-Jung Shih and Wei-Der Shih. Gesture modeling for architectural design. Computers and Graphics, 20(6):849–862, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600056.

Schneider:1997:GEI

[SS97] Bengt-Olaf Schneider and Andreas Schilling. Guest Editors’ introduction. Computers and Graphics, 21(2): 125–127, March–April 1997.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9786686.

Skov:2002:DIN

[SS02] Mikael B. Skov and Jan Stage. Designing interactive narrative systems: is objectorientation useful? Computers and Graphics, 26(1):57– 66, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/33/abstract.
html.

Soon:2004:GEI

[SS04a] Seah Hock Soon and Alexei Sourin. Guest editor’s introduction. Computers and Graphics, 28(4):465– 466, August 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sung:2004:TAT

[SS04b] Kelvin Sung and Peter Shirley. A top-down approach to teaching introductory computer graphics. Computers and Graphics, 28(3):383–391, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Stahovich:2007:I

[SS07] Thomas F. Stahovich and Mario Costa Sousa. Introduction. Computers and Graphics, 31(4):538– 539, August 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001379.

Sadeghi:2009:SRS

[SS09] Javad Sadeghi and Faramarz F. Samavati. Smooth reverse subdivision. Computers and Graphics, 33(3): 217–225, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000235.

Schmidt:2012:CM

[SS12a] Ryan Schmidt and Patricio Simari. Consensus meshing. Computers and Graphics, 36 (5):488–497, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000696.

Shamir:2012:IVQ

[SS12b] Ariel Shamir and Alla Stolpnik. Interactive visual queries for multivariate graphs exploration. Computers and Graphics, 36(4): 257–264, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000234.

Silva:2013:IHF

[SS13] Nuno Silva and Luís Paulo Santos. Interactive high fidelity visualization of complex materials on the GPU. Computers and Graphics, 37 (7):809–819, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001040.

Saporta:2022:URD

[SS22] Tsahi Saporta and Andrei Sharf. Unsupervised recursive deep fitting of 3D primitives to points. Computers and Graphics, 102 (??):289–299, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002314.

Sung:1996:RTD

[SSA96] Kelvin Sung, Jason Loh Jen Shiuan, and A. L. Ananda. Ray tracing in a distributed environment. Computers and Graphics, 20(1):41– 49, January–February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500091.

Sahm:2004:ERS

[SSB04] J. Sahm, I. Soetebier, and H. Birthelmer. Efficient representation and streaming of 3D scenes. Computers and Graphics, 28 (1):15–24, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Shirley:2008:FRT

[SSB^{+}08] Peter Shirley, Kelvin Sung, Erik Brunvand, Alan Davis, Steven Parker, and Solomon Boulos. Fast ray tracing and the potential effects on graphics and gaming courses. Computers and Graphics, 32(2):260–267, April 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000150.

Surazhsky:2001:BPS

[SSBT01] Tatiana Surazhsky, Vitaly Surazhsky, Gill Barequet, and Ayellet Tal. Blending polygonal shapes with different topologies. Computers and Graphics, 25 (1):29–39, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/

                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/26/29/article.
                                        pdf.

Schafer:2012:MEL

[SSDS12] Henry Schäfer, Jochen Süßmuth, Cornelia Denk, and Marc Stamminger. Memory efficient light baking. Computers and Graphics, 36(3): 193–200, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001683.

Scharfig:2016:CLA

[SSH16] Randolf Schärfig, Marc Stamminger, and Kai Hormann. Creating light atlases with multi-bounce indirect illumination. Computers and Graphics, 55(??): 97–107, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001661.

Sawkar:1987:CPS

[SSK87] D. G. Sawkar, G. R. Shevare, and S. P. Koruthu. Contour plotting for scattered data. Computers and Graphics, 11(2):101–104, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Saxena:1987:HLS

[SSM87] Sanjaya K. Saxena, Sanjiv Saxena, and Rao V. S. Malladi. The hierarchical logic schematic capture. Computers and Graphics, 11 (2):87–93, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Silva:2011:ACC

[SSM11a] Jeferson R. Silva, Thiago T.

Santos, and Carlos H. Morimoto. Automatic camera control in virtual environments augmented using multiple sparse videos. Computers and Graphics, 35(2):412–421, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000215.

Silva:2011:UCV

[SSM11b] Samuel Silva, Beatriz Sousa Santos, and Joaquim Madeira. Using color in visualization: a survey. Computers and Graphics, 35(2): 320–333, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001846.

Shang:2024:FAN

[SSQL24] Kai Shang, Mingwen Shao, Yuanjian Qiao, and Huan Liu. Frequency-aware network for low-light image enhancement. Computers and Graphics, 118(??):210– 219, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003084.

Sousa:1990:OOS

[SSS90] Joao Pedro Sousa, Cristina Sernadas, and Amilcar Sernadas. An object-oriented specification tool for graphical interfaces. Computers and Graphics, 14(1):29–40, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Steinicke:2015:FBG

[SSS15] Frank Steinicke, Wolfgang Stuerzlinger, and Evan Suma. Foreword to Computers & Graphics special section on spatial user interaction (SUI). Computers and Graphics, 48(??): A1–A2, May 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000266.

Sankaranarayanan:2007:FAN

[SSV07] Jagan Sankaranarayanan, Hanan Samet, and Amitabh Varshney. A fast all nearest neighbor algorithm for applications involving large point-clouds. Computers and Graphics, 31(2):157– 174, April 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002378.

Sun:2023:GVD

[SSW^{+}23] Yanan Sun, Zhiyao Sun, Yu-Hui Wen, Sheng Ye, Tian Lv, Minjing Yu, Ran Yi, Lin Gao, and Yong-Jin Liu. Generation of virtual digital human for customer service industry. Computers and Graphics, 115 (??):359–370, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001474.

Schmalstieg:1997:ECD

[ST97] Dieter Schmalstieg and Robert F. Tobler. Exploiting coherence in 2.5-D visibility computation. Computers and Graphics, 21(1):121–123, January– February 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=1&aid=9600092.

Shih:2002:PBV

[ST02] Naai-Jung Shih and Yu-Tun Tsai. A photogrammetrybased verification of assumptions applied in the interpretation of paper architecture. Computers and Graphics, 26(1):109–124, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/38/abstract.
                                        html.

Saiti:2020:AIR

[ST20] E. Saiti and T. Theoharis. An application independent review of multimodal 3D registration methods. Computers and Graphics, 91 (??):153–178, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932030114X.

Saiti:2022:MRA

[ST22] E. Saiti and T. Theoharis. Multimodal registration across 3D point clouds and CT-volumes. Computers and Graphics, 106 (??):259–266, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001121.

Sundt:2023:MMA

[ST23] Peder Bergebakken Sundt and Theoharis Theoharis.

MARF: the Medial Atom Ray Field object representation. Computers and Graphics, 115(??):122–136, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001279.

Stark:1987:IDC

[Sta87a] Jurgen Stark. The integrated display controller (IDC) for VLSI-design workstations. Computers and Graphics, 11 (2):185–192, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Stash:1987:MSA

[Sta87b] Susan Stash. A multiple standards approach. Computers and Graphics, 11(4): 479–481, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Starkov:2019:USA

[STBG19] Rastislav Starkov, Christine Tanner, Michael Bajka, and Orcun Goksel. Ultrasound simulation with animated anatomical models and on-the-fly fusion with real images via path-tracing. Computers and Graphics, 82(??):44–52, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300603.

Smelik:2011:DAP

[STdKB11] R. M. Smelik, T. Tutenel, K. J. de Kraker, and R. Bidarra. A declarative approach to procedural modeling of virtual worlds. Computers and Graphics, 35(2):352–363, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001809.

Stellman:1975:ATD

[Ste75] S. D. Stellman. Application of three-dimensional interactive graphics in X-ray crystallographic analysis. Computers and Graphics, 1(2– 3):279–288, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Stewart:1999:CVF

[Ste99] A. James Stewart. Computing visibility from folded surfaces. Computers and Graphics, 23(5):693–702, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/32/
                                        33/article.pdf;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/33/abstract.
                                        html.

Stepien:2009:IBM

[Ste09] C. Stepień. An IFS-based method for modelling horns, seashells and other natural forms. Computers and Graphics, 33(4):576– 581, August 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000193.

Seron:2004:GVM

[STM^{+}04] F. J. Seron, J. J. Torrens, J. A. Magallon, A. Turon, and S. Baldassarri. Geometric and visual modelling of complex stratigraphic structures. Computers and Graphics, 28 (4):585–599, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Somers:1995:OOI

[STN95] Fergal Somers, Ghee S. Teo, and Francis Neelamkavil. Object-oriented implementation of the OSF/Motif^{T M} widget set in Eiffel. Computers and Graphics, 19(4): 575–583, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500035.

Schreck:2020:FSS

[STP^{+}20] Tobias Schreck, Theoharis Theoharis, Ioannis Pratikakis, Michela Spagnuolo, and Remco C. Veltkamp. Foreword to the special section on 3D Object Retrieval 2020 workshop (3DOR2020). Computers and Graphics, 93 (??):A3–A4, December 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301564.

Strasser:1983:GEI

[Str83a] W. Strasser. Guest Editor’s introduction: Graphics in the office. Computers and Graphics, 7(1): 3–4, ???? 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849383900390.

Strasser:1983:V

[Str83b]  W. Strasser.  Videodisc.

Computers and Graphics, 7 (3–4):351–353, 1983. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Straayer:1985:STD

[Str85a]  D. H. Straayer.  The

standardization of threedimensional graphics systems. Computers and Graphics, 9(1):27–32, 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Straub:1985:CFT

[Str85b] D. Straub. A comparison of and future trends in personal computer-based computer-aided design and drafting software. Computers and Graphics, 9(3): 323–329, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Strickland:1986:AGC

[Str86] Robin N. Strickland. An algorithm for generating connected quadrilaterals of specified area. Computers and Graphics, 10(2): 151–155, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Strasser:1995:GEI

[Str95] Wolfgang Strasser. Guest Editor’s introduction. Computers and Graphics, 19(5): 651–??, September–October 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Strasser:1996:GEI

[Str96] Wolfgang Strasser. Guest Editor’s introduction. Computers and Graphics, 20(4): 473–??, July–August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Salazar:2018:HBB

[STT^{+}18] Steeven Villa Salazar, Jose Abel Ticona, Rafael Torchelsen, Luciana Nedel, and Anderson Maciel. Heat-based bidirectional phase shifting simulation using positionbased dynamics. Computers and Graphics, 76(??): 107–116, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301389.

Stuedell:1991:BM

[Stu91] David Stuedell. Biomorphic mitosis. Computers and Graphics, 15(3):455–??, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sturzlinger:1998:CGI

[Stü98] Wolfgang Stürzlinger. Calculating global illumination for glossy surfaces. Computers and Graphics, 22(2–3):175– 180, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/538.pdf.

Shephard:1982:ASF

[STW82] M. S. Shephard, C. N. Tonias, and T. J. Weidner. Attribute specification for finite element models. Computers and Graphics, 6(2):83–91, 1982. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Slater:1993:SPV

[SU93] Mel Slater and Martin Usoh.

Simulating peripheral vision in immersive virtual environments. Computers and Graphics, 17(6):643–653, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Suffern:1988:PVP

[Suf88] Kevin G. Suffern. Perspective views of polar coordinate functions. Computers and Graphics, 12(3–4): 515–524, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Staadt:2018:FSS

[SUF^{+}18] Oliver Staadt, Jonas Unger, Morten Fjeld, Marco Fratarcangeli, and Daniel Sjölie. Foreword to the special section on the 23rd ACM symposium on virtual reality software and technology 2017. Computers and Graphics, 77(??):A3– A4, December 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301791.

Sugihara:1983:RDT

[Sug83]  K. Sugihara.  A robust

description of time-varying scenes for computer animation. Computers and Graphics, 7(3–4):277–284, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sugihara:1984:IAP

[Sug84] K. Sugihara. Interpretation of an axonometric projection of a polyhedron. Computers and Graphics, 8(4): 391–400, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sulonen:1985:GEI

[Sul85] Reijo “Shosta” Sulonen. Guest Editor’s introduction: Arctic views on computer graphics. Computers and Graphics, 9(4):337– 338, ???? 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849385900287.

Sidiropoulos:2006:URS

[SV06] George Sidiropoulos and Athanasios Vasilakos. Ultrareal or symbolic visualization? The case of the city through time. Computers and Graphics, 30(2): 299–310, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000367.

Salvi:2018:MSB

[SV18] Péter Salvi and Tamás Várady. Multi-sided Bézier surfaces over concave polygonal domains. Computers and Graphics, 74(??): 56–65, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300657.

Sabino:2023:BOB

[SVCNM23] Rodolfo Sabino, Creto Augusto Vidal, Joaquim Bento Cavalcante-Neto, and José Gilvan Rodrigues Maia. Building oriented bounding boxes by the intermediate use of ODOPs. Computers and Graphics, 116(??): 251–261, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002091.

Spalter:2003:PUC

[SvD03] Anne Morgan Spalter and Andries van Dam. Problems with using components in educational software. Computers and Graphics, 27(3): 329–337, June 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smit:2009:SBA

[SvL09] F. A. Smit and R. van Liere. A simulator-based approach to evaluating optical trackers. Computers and Graphics, 33(2): 120–129, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900017X.

Smit:2010:SSG

[SvLBF10] Ferdi Smit, Robert van Liere, Stephan Beck, and Bernd Froehlich. A sharedscene-graph image-warping architecture for VR: Low latency versus image quality. Computers and Graphics, 34(1):3–16, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001344.

Saona-Vazquez:1999:VOD

[SVNB99] C. Saona-Vázquez, I. Navazo, and P. Brunet. The visibility octree: a data structure for 3D navigation. Computers and Graphics, 23 (5):635–643, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/

28/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/28/abstract.
html.

Srikant:1982:IGS

[SVP82] Y. N. Srikant, D. Vidyasagar, and L. M. Patnaik. An interactive graphics system for 2D drawing and design. Computers and Graphics, 6 (1):23–27, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Smit:2007:GTP

[SvRvL07] F. A. Smit, A. van Rhijn, and R. van Liere. Graphtracker: a topology projection invariant optical tracker. Computers and Graphics, 31(1):26–38, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001828.

Sipos:2020:MSI

[SVSV20] Ágoston Sipos, Tamás Várady, Péter Salvi, and Márton Vaitkus. Multi-sided implicit surfacing with I-patches. Computers and Graphics, 90(??):29–42, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300601.

Schuster:1986:UCD

[SVT86] R. Schuster, E. Voge, and D. Trippner. The use of computers in design and planning — integration via interface management. Computers and Graphics, 10(4): 277–295, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Soto:1992:TCC

[SVV92] Antoni Soto, Sebastia Vila, and Alvar Vinacua. A toolkit for constructing command driven graphics programs. Computers and Graphics, 16 (4):375–382, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Salvi:2023:CMM

[SVV23] Péter Salvi, Márton Vaitkus, and Tamás Várady. Constrained modeling of multisided patches. Computers and Graphics, 114(??): 86–95, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000754.

Soriano-Vargas:2017:VAT

[SVVS^{+}17] Aurea Soriano-Vargas, Bruno C. Vani, Milton H. Shimabukuro, João F. G. Monico, Maria Cristina F. Oliveira, and Bernd Hamann. Visual analytics of time-varying multivariate ionospheric scintillation data. Computers and Graphics, 68(??): 96–107, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301486.

Silva:2023:ITB

[SVW23] Daniel Silva, Raphael Voltoline, and Shin-Ting Wu. An interactive triangle-based ODF glyph rendering for high angular resolution diffusion imaging. Computers and Graphics, 116(??): 393–403, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300208X.

Spackman:1991:SNR

[SW91] John Spackman and Philip Willis. The SMART navigation of a ray through an octtree. Computers and Graphics, 15(2):185–194, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Scheiblauer:2011:CSE

[SW11] Claus Scheiblauer and Michael Wimmer. Out-of-core selection and editing of huge point clouds. Computers and Graphics, 35(2): 342–351, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000057.

Siess:2019:UCT

[SW19] Andreas Siess and Matthias Wölfel. User color temperature preferences in immersive virtual realities. Computers and Graphics, 81(??): 20–31, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931930041X.

Shang:2021:API

[SW21] Yunyi Shang and Hon-Cheng Wong. Automatic portrait image pixelization. Computers and Graphics, 95(??): 47–59, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100008X.

Sassatelli:2020:NIS

[SWF^{+}20] Lucile Sassatelli, Marco Winckler, Thomas Fisichella, Antoine Dezarnaud, Julien Lemaire, Ramon Aparicio-Pardo, and Daniela Trevisan. New interactive strategies for virtual reality streaming in degraded context of use. Computers and Graphics, 86(??): 27–41, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301621.

Shen:2017:PBA

[SWH^{+}17] Yijun Shen, He Wang, Edmond S. L. Ho, Longzhi Yang, and Hubert P. H. Shum. Posture-based and action-based graphs for boxing skill visualization. Computers and Graphics, 69(??): 104–115, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301590.

Su:2016:RDT

[SWL^{+}16] Tianyun Su, Wen Wang, Zhihan Lv, Wei Wu, and Xinfang Li. Rapid Delaunay triangulation for randomly distributed point cloud data using adaptive Hilbert curve. Computers and Graphics, 54 (??):65–74, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001223.

SotoSanchez:2021:IRP

[SWMdF21] José Ezequiel Soto Sánchez, Tim Weyrich, Asla Medeiros e Sá, and Luiz Henrique de Figueiredo. An integer representation for periodic tilings of the plane by regular polygons. Computers and Graphics, 95(??): 69–80, April 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000078.

Shore:1975:IMS

[SWS75] S. Shore, J. L. Wilson, and G. A. Semsarzadeh. Interactive modeling system for bridges. Computers and Graphics, 1(4):337–346, December 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sugihara:2010:WTE

[SWS10] Masamichi Sugihara, Brian Wyvill, and Ryan Schmidt. WarpCurves: a tool for explicit manipulation of implicit surfaces. Computers and Graphics, 34(3): 282–291, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000476.

Smit:1995:DRT

[SWvB95] J. Smit, H. J. Wessels, A. van der Horst, and M. J. Bentum. On the design of a real-time volume rendering engine. Computers and Graphics, 19(2): 297–300, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=2&aid=9400155.

Shi:2023:RDG

[SWZZ23] Jinhua Shi, Shuhong Wang, Naming Zhang, and Jianguo Zhu. A residualsdistribution-guided local optimization approach to Bspline fitting in capturing image outlines. Computers and Graphics, 112(??): 105–115, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000389.

Siqueira:2009:NCS

[SXG^{+}09] Marcelo Siqueira, Dianna Xu, Jean Gallier, Luis Gustavo Nonato, Dimas Martínez Morera, and Luiz Velho. A new construction of smooth surfaces from triangle meshes using parametric pseudo-manifolds. Computers and Graphics, 33(3): 331–340, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900034X.

Shen:2023:UUI

[SXL^{+}23] Zhen Shen, Haiyong Xu, Ting Luo, Yang Song, and Zhouyan He. UDAformer: Underwater image enhancement based on dual attention transformer. Computers and Graphics, 111(??): 77–88, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000109.

Song:2014:PPP

[SXSY14] Hai-Chuan Song, Xin Xu, Kan-Le Shi, and Jun-Hai Yong. Projecting points onto planar parametric curves by local biarc approximation. Computers and Graphics, 38 (??):183–190, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931300174X.

Sun:2022:MVS

[SXW^{+}22] Shang Sun, Dan Xu, Hao Wu, Haocong Ying, and Yurui Mou. Multi-view stereo for large-scale scene reconstruction with MRFbased depth inference. Computers and Graphics, 106 (??):248–258, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200111X.

Saito:2023:CME

[SY23] Takafumi Saito and Norimasa Yoshida. Curvature monotonicity evaluation functions on rational Bézier curves. Computers and Graphics, 114 (??):219–228, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000742.

Sun:2021:TMM

[SYMW21] Riming Sun, Yichen Yang, Yongfeng Ma, and Shengfa Wang. A transmission model for motion estimation of instability space targets. Computers and Graphics, 98 (??):29–36, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000510.

Shao:1995:NCF

[SZ95] Lejun Shao and Hao Zhou.

A new contour fill algorithm for outlined character image generation. Computers and Graphics, 19(4): 551–556, July–August 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=4&aid=9500033.

Sun:2009:GFC

[SZ09] Chuan Sun and Huanxi Zhao. Generating fair, C^{2} continuous splines by blending conics. Computers and Graphics, 33(2): 173–180, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001283.

Shi:1993:MRA

[SZEG93] Jiaoying Shi, Aidong Zhang, Jose Encarnação, and Martin Göbel. A modified radiosity algorithm for integrated visual and auditory rendering. Computers and Graphics, 17(6):633–642, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Su:2013:TAC

[SZL^{+}13] Zhiyong Su, Lang Zhou, Weiqing Li, Yuewei Dai, and Weiqing Tang. Topology authentication for CAPD models based on Laplacian coordinates. Computers and Graphics, 37(4): 269–279, June 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000320.

Song:2010:AVS

[SZW10] Ruixia Song, Zhaoxia Zhao, and Xiaochun Wang. An application of the V-system to the clustering of Chernoff faces. Computers and Graphics, 34(5):529– 536, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000865.

Song:2014:FCS

[SZW^{+}14] Chao Song, Hongxin Zhang, Xun Wang, Jianwei Han, and Huiyan Wang. Fast corotational simulation for example-driven deformation. Computers and Graphics, 40 (??):49–57, May 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000120.

Szyszkowicz:1989:CGG

[Szy89a]  Mieczysław  Szyszkowicz.

Computer graphics generated by numerical iteration. Computers and Graphics, 13(1):121–126, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szyszkowicz:1989:IN

[Szy89b] Mieczysław Szyszkowicz. Images of nonlinearity. Computers and Graphics, 13(1): 119–120, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szyszkowicz:1990:CAG

[Szy90a]  Mieczysław  Szyszkowicz.

Computer art generated by the method of secants in the complex plane. Computers and Graphics, 14 (3–4):509, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szyszkowicz:1990:SGD

[Szy90b] Mieczysław Szyszkowicz. A simple gasket derived from the logistic parabola. Computers and Graphics, 14(2): 335–336, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szyszkowicz:1991:PGL

[Szy91]  Mieczysław  Szyszkowicz.

Patterns generated by logical operators. Computers and Graphics, 15(2): 299–300, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Szyszkowicz:1992:GRP

[Szy92]  Mieczysław  Szyszkowicz.

Graphical representation of pseudorandom numbers. Computers and Graphics, 16 (2):237–??, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tomiyama:1990:GEI

[TA90] Tetsuo Tomiyama and Farhad Arbab. Guest Editors’ introduction: Features and geometric reasoning. Computers and Graphics, 14(2): 147–??, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2022:NGF

[TACS22] Jingwei Tang, Vinicius C.

Azevedo, Guillaume Cordonnier, and Barbara Solenthaler. Neural Green’s function for Laplacian systems. Computers and Graphics, 107(??):186–196, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001406.

Thaller:2016:PMF

[TAF16] Wolfgang Thaller, Ursula Augsdörfer, and Dieter W. Fellner. Procedural mesh features applied to subdivision surfaces using graph grammars. Computers and Graphics, 58(??):184– 192, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300577.

Takats:1977:TEC

[Tak77] Martha C. Takats. Teaching elementary concepts in quantum mechanics with computer graphics. Computers and Graphics, 2(3): 125–128, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tanimoto:1980:CMT

[Tan80] S. L. Tanimoto. Colormapping techniques for computer-aided design and verification of VLSI systems. Computers and Graphics, 5(2–4):103–113, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:1994:PCC

[Tan94] Zesheng Tang, editor. Proceedings of the Conference on Computer Aided Design and Computer Graphics, volume 18(4) of Computers and Graphics. Pergamon, New York, NY, USA, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Taponecco:2006:DTB

[Tap06] Francesca Taponecco. Dense texture-based visualization of unsteady and multivariate vector fields. Computers and Graphics, 30(3): 353–358, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000574.

Tarini:2022:CFQ

[Tar22]  Marco Tarini.  Closed-

form quadrangulation of nsided patches. Computers and Graphics, 107(??):60– 65, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001170.

Tominski:2009:CIG

[TAS09] Christian Tominski, James Abello, and Heidrun Schumann. CGV-an interactive graph visualization system. Computers and Graphics, 33 (6):660–678, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000879.

Taxen:2004:TCG

[Tax04]  Gustav Taxén.  Teach-

ing computer graphics constructively. Computers and Graphics, 28(3):393– 399, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Taylor:1987:AIC

[Tay87] Michael A. P. Taylor. Applying interactive colour graphics in traffic planning. Computers and Graphics, 11(3): 241–248, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Trescak:2018:SCS

[TB18] Tomas Trescak and Anton Bogdanovych. Simulating complex social behaviours of virtual agents through casebased planning. Computers and Graphics, 77(??): 122–139, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301596.

Thompson:2019:RCP

[TB19] Elia Moscoso Thompson and Silvia Biasotti. Retrieving color patterns on surface meshes using edgeLBP descriptors. Computers and Graphics, 79(??):46– 57, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930010X.

Thompson:2020:PMP

[TBDC20] Elia Moscoso Thompson, Silvia Biasotti, Julie Digne, and Raphaëlle Chaine. mpLBP: a point-based representation for surface pattern description. Computers and Graphics, 86(??):81– 92, February 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301840.

Thompson:2020:PSR

[TBG^{+}20] Elia Moscoso Thompson, Silvia Biasotti, Andrea Giachetti, Claudio Tortorici, Naoufel Werghi, Ahmad Shaker Obeid, Stefano Berretti, Hoang-Phuc Nguyen-Dinh, Minh-Quan Le, Hai-Dang Nguyen, Minh-Triet Tran, Leonardo Gigli, Santiago Velasco-Forero, Beatriz Marcotegui, Ivan Sipiran, Benjamin Bustos, Ioannis Romanelis, Vlassis Fotis, Gerasimos Arvanitis, Konstantinos Moustakas, Ekpo Otu, Reyer Zwiggelaar, David Hunter, Yonghuai Liu, Yoko Arteaga, and Ramamoorthy Luxman. SHREC 2020: Retrieval of digital surfaces with similar geometric reliefs. Computers and Graphics, 91 (??):199–218, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301138.

Thomas:2017:QTV

[TBLH17] Dean P. Thomas, Rita Borgo, Robert S. Laramee, and Simon J. Hands. QCD-Vis: a tool for the visualisation of Quantum Chromodynamics (QCD) data. Computers and Graphics, 67 (??):115–126, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300997.

Tomov:2004:AIP

[TBM^{+}04] Stanimire Tomov, Robert Bennett, Michael McGuigan, Arnold Peskin, Gordon Smith, and John Spiletic. Application of interactive parallel visualization for commodity-based clusters using visualization APIs. Computers and Graphics, 28(2):273–278, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tadeja:2023:ERP

[TBS^{+}23] Sławomir K. Tadeja, Luca O. Solari Bozzi, Kerr D. G. Samson, Sebastian W. Pattinson, and Thomas Bohné. Exploring the repair process of a 3D printer using augmented reality-based guidance. Computers and Graphics, 117(??):134–144, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002546.

Turcotte:1993:DDI

[TC93] Louis H. Turcotte and Bradley M. Comes. Delivering data interpretation: From GFLOPS to insight. Computers and Graphics, 17(1):23–30, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tsai:2000:RFA

[TC00] Yao-Hong Tsai and Kuo-Liang Chung. Region-filling algorithm on bincode-based contour and its implementation. Computers and Graphics, 24(4):529–537, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/31/article.
pdf.

Tasoren:2024:NAD

[TC24] Ali Egemen Tasoren and Ufuk Celikcan. NOVAction23: Addressing the data diversity gap by uniquely generated synthetic sequences for real-world human action recognition. Computers and Graphics, 118 (??):1–10, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002467.

Tao:2015:MSR

[TCL^{+}15a] Pingping Tao, Junjie Cao, Shuhua Li, Xiuping Liu, and Ligang Liu. Mesh saliency via ranking unsalient patches in a descriptor space. Computers and Graphics, 46 (??):264–274, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001113.

Tran:2015:ECE

[TCL15b] Trung-Thien Tran, Van-Toan Cao, and Denis Laurendeau. Extraction of cylinders and estimation of their parameters from point clouds. Computers and Graphics, 46(??):345– 357, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001150.

Turner:2000:SS

[TCP00] Alasdair Turner, David Chapman, and Alan Penn. Sketching space. Computers and Graphics, 24(6): 869–879, December 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/34/

30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/34/30/article.
pdf.

Tiddeman:2001:GMO

[TDR01] Bernard Tiddeman, Neil Duffy, and Graham Rabey. A general method for overlap control in image warping. Computers and Graphics, 25 (1):59–66, February 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/31/article.
pdf.

Takouachet:2017:PNG

[TDR^{+}17] Nawel Takouachet, Samuel Delepoulle, Christophe Renaud, Nesrine Zoghlami, and João Manuel R. S. Tavares. Perception of noise and global illumination: Toward an automatic stopping criterion based on SVM. Computers and Graphics, 69 (??):49–58, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301607.

Tasse:2014:FBT

[TEC^{+}14] Flora Ponjou Tasse, Arnaud Emilien, Marie-Paule Cani, Stefanie Hahmann, and Neil Dodgson. Featurebased terrain editing from complex sketches. Computers and Graphics, 45(??): 101–115, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000818.

Teixeira:1996:ETC

[Tei96] José Carlos Teixeira. Environments for teaching computer graphics: An experience. Computers and Graphics, 20(6):927–935, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=6&aid=9600063.

Teshima:1984:ACA

[Tes84] Aldo Norio Teshima. Application of CAD in aircraft wiring diagrams in the Brazilian aeronautical industry. Computers and Graphics, 8(3):231–236, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2018:MSS

[TF18] Yizhi Tang and Jieqing Feng. Multi-scale surface reconstruction based on a curvature-adaptive signed distance field. Computers and Graphics, 70(??): 28–38, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301085.

Texler:2020:AST

[TFF^{+}20] Ondřej Texler, David Futschik, Jakub Fišer, Michal Lukáč, Jingwan Lu, Eli Shechtman, and Daniel Sýkora. Arbitrary style transfer using neurally-guided patchbased synthesis. Computers and Graphics, 87(??): 62–71, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300042.

Tal:2006:E

[TFS06] Ayellet Tal, Thomas Funkhouser, and Ariel Shamir. Editorial. Computers and Graphics, 30(6):891, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001440.

Tanaka:2000:SAD

[TFY00] Sotoshi Tanaka, Yasushi Fukuda, and Hiroaki Yamamoto. Stochastic algorithm for detecting intersection of implicit surfaces. Computers and Graphics, 24 (4):523–528, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/32/30/article.
                                        pdf.

Thon:2002:OWS

[TG02] Sébastien Thon and Djamchid Ghazanfarpour. Ocean waves synthesis and animation using real world information. Computers and Graphics, 26(1):99–108, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/27/37/abstract.
                                        html.

Tewari:2006:MGP

[TGG06] Geetika Tewari, Craig Gotsman, and Steven J. Gortler. Meshing genus-1 point clouds using discrete one-forms. Computers and Graphics, 30 (6):917–926, December 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001476.

Tapadia:1990:UFB

[TH90] Rajendra K. Tapadia and Mark R. Henderson. Using a feature-based model for automatic determination of assembly handling codes. Computers and Graphics, 14 (2):251–262, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2009:STM

[TH09] Yuk-Ming Tang and Kin-Chuen Hui. Simulating tendon motion with axial massspring system. Computers and Graphics, 33(2): 162–172, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000156.

tenHagen:1990:DGW

[tHd90] P. J. W. ten Hagen, I. Herman, and J. R. G. de Vries. A dataflow graphics workstation. Computers and Graphics, 14(1):83–93, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Editors:2015:GEF

[The15] The Guest Editors SMI 2015. Guest Editor foreword. Computers and Graphics, 51 (??):A1–A2, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000849.

Thimbleby:1985:FTU

[Thi85] Harold Thimbleby. Failure in the technical user-interface design process. Computers and Graphics, 9(3):187– 193, ???? 1985. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0097849385900469.

Truong-Hong:2015:QES

[THL15] Linh Truong-Hong and Debra F. Laefer. Quantitative evaluation strategies for urban 3D model generation from remote sensing data. Computers and Graphics, 49(??):82–91, June 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000254.

Thompson:1984:GCE

[Tho84]  K. Thompson.  Graph-

ics certification at the European Community level. Computers and Graphics, 8 (1):59–61, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tao:2016:VWI

[THQ^{+}16] Jun Tao, Xiaoke Huang, Feng Qiu, Chaoli Wang, Jingfeng Jiang, Ching-Kuang Shene, Ye Zhao, and Daphne Yu. VesselMap: a web interface to explore multivariate vascular data. Computers and Graphics, 59 (??):79–92, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931630067X.

terHaar:2010:EME

[tHV10] Frank B. ter Haar and Remco C. Veltkamp. Expression modeling for expressioninvariant face recognition. Computers and Graphics, 34(3):231–241, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931000049X.

Tindle:1986:FSD

[Tin86] G. L. Tindle. Fermi surface display. Computers and Graphics, 10(1):77, 79, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tamminen:1985:IFR

[TJ85] M. Tamminen and F. W.

Jansen. An integrity filter for recursive subdivision meshes. Computers and Graphics, 9(4):351–363, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tagliasacchi:2020:FSS

[TK20] Andrea Tagliasacchi and Paul Kry. Foreword to the special section on graphics interface. Computers and Graphics, 88(??): A6–A7, May 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030025X.

Turdukulov:2007:DVE

[TKB07] Ulanbek D. Turdukulov, Menno-Jan Kraak, and Connie A. Blok. Designing a visual environment for exploration of time series of remote sensing data: In search for convective clouds. Computers and Graphics, 31(3):370–379, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000593.

Tasse:2016:HWD

[TKD16] Flora Ponjou Tasse, Jirí Kosinka, and Neil Dodgson. How well do saliencybased features perform for shape retrieval? Computers and Graphics, 59 (??):57–67, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300346.

Trevisan:2022:FSS

[TKdJO22] Daniela Gorski Trevisan, Regis Kopper, and Victor Adriel de Jesus Oliveira. Foreword to the special section on the 2021 Symposium on Virtual and Augmented Reality (SVR 2021). Computers and Graphics, 102(??):A12–A13, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002491.

Trunz:2024:NIP

[TKM^{+}24] Elena Trunz, Jonathan Klein, Jan Müller, Lukas Bode, Ralf Sarlette, Michael Weinmann, and Reinhard Klein. Neural inverse procedural modeling of knitting yarns from images. Computers and Graphics, 118 (??):161–172, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003072.

Titov:2024:CAO

[TKOD24] Andrey Titov, Marta Kersten-Oertel, and Simon Drouin. Contextual ambient occlusion: a volumetric rendering technique that supports real-time clipping. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000116.

Tatzgern:2011:MPC

[TKS11] Markus Tatzgern, Denis Kalkofen, and Dieter Schmalstieg. Multi-perspective compact explosion diagrams. Computers and Graphics, 35 (1):135–147, February 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001743.

Thaller:2013:SGC

[TKZ^{+}13] Wolfgang Thaller, Ulrich Krispel, René Zmugg, Sven Havemann, and Dieter W. Fellner. Shape grammars on convex polyhedra. Computers and Graphics, 37 (6):707–717, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000861.

Thien:2002:TSS

[TL02] Chih-Ching Thien and Ja-Chen Lin. Technical section: Secret image sharing. Computers and Graphics, 26(5):765–770, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

13/20/68/56/39/abstract.
html.

Tresset:2013:PDP

[TL13] Patrick Tresset and Frederic Fol Leymarie. Portrait drawing by Paul the Robot. Computers and Graphics, 37 (5):348–363, August 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000149.

Turrill:1975:XEI

[TM75] C. N. Turrill and W. R.

Mallgren. XPLG — experiences in implementing an experimental interactive graphics programming system. Computers and Graphics, 1(1):55–64, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tomov:2005:BIP

[TMB^{+}05] Stanimire Tomov, Michael McGuigan, Robert Bennett, Gordon Smith, and John Spiletic. Benchmarking and implementation of probability-based simulations on programmable graphics cards. Computers and Graphics, 29(1):71– 80, February 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tong:2020:FSS

[TMH20] Xin Tong, Karol Myszkowski, and Jin Huang. Foreword to the special section on the international conference on computer-aided design and computer graphics (CAD/Graphics) 2019. Computers and Graphics, 86 (??):A5–A6, February 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301852.

Tijerino:1994:IDC

[TMK94] Yuri A. Tijerino, Kenji Mochizuki, and Fumio Kishino. Interactive 3-D computer graphics driven through verbal instructions: Previous and current activities at ATR. Computers and Graphics, 18(5):621–631, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tanaka:2000:SIS

[TMN^{+}00] Satoshi Tanaka, Akio Morisaki, Satoru Nakata, Yasushi Fukuda, and Hiroaki Yamamoto. Sampling implicit surfaces based on stochastic differential equations with converging constraint. Computers and Graphics, 24(3): 419–431, June 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/37/article.
pdf.

Telles:2007:NCD

[TMP07] G. P. Telles, R. Minghim, and F. V. Paulovich. Normalized compression distance for visual analysis of document collections. Computers and Graphics, 31(3): 327–337, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000556.

Trutoiu:2009:CLC

[TMSPB09] Laura C. Trutoiu, Betty J.

Mohler, Jörg Schulte-Pelkum, and Heinrich H. Bülthoff. Circular, linear, and curvilinear vection in a largescreen virtual environment with floor projection. Computers and Graphics, 33 (1):47–58, February 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001477.

Taxen:2002:SEO

[TN02] Gustav Taxén and Ambjörn Naeve. A system for exploring open issues in VR-based education. Computers and Graphics, 26(4):593–598, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        13/20/68/55/38/abstract.
                                        html.

Takahashi:2014:FSV

[TNF14] Tetsuya Takahashi, Tomoyuki Nishita, and Issei Fujishiro. Fast simulation of viscous fluids with elasticity and thermal conductivity using position-based dynamics. Computers and Graphics, 43 (??):21–30, October 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000612.

Tournier:2014:VBA

[TNFG14] Maxime Tournier, Matthieu Nesme, François Faure, and Benjamin Gilles. Velocitybased adaptivity of deformable models. Computers and Graphics, 45 (??):75–85, December 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000788.

Tanaka:2001:ASS

[TNU^{+}01] Satoshi Tanaka, Tomoharu Nakamura, Miharu Ueda, Hiroaki Yamamoto, and Kisou Shino. Application of the stochastic sampling method to various implicit surfaces. Computers and Graphics, 25(3): 441–448, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/32/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/32/34/article.
pdf.

Taketomi:2014:CPE

[TOY^{+}14] Takafumi Taketomi, Kazuya Okada, Goshiro Yamamoto, Jun Miyazaki, and Hirokazu Kato. Camera pose estimation under dynamic intrinsic parameter change for augmented reality. Computers and Graphics, 44 (??):11–19, November 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000648.

Theoharis:1989:PRD

[TP89] Theoharis Theoharis and Ian Page. Polygon rendering on a dual-paradigm parallel processor. Computers and Graphics, 13(2): 207–216, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Torchelsen:2017:FSS

[TP17] Rafael Torchelsen and Daniele Panozzo. Foreword to the special section on SIB-GRAPI 2017. Computers and Graphics, 68(??):A1– A6, November 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301553.

Thomaszewski:2008:PTP

[TPB08] Bernhard Thomaszewski, Simon Pabst, and Wolfgang Blochinger. Parallel techniques for physically based simulation on multi-core processor architectures. Computers and Graphics, 32 (1):25–40, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307002002.

Treece:1999:RMT

[TPG99] G. M. Treece, R. W. Prager, and A. H. Gee. Regularised marching tetrahedra: improved iso-surface extraction. Computers and Graphics, 23(4):583–598, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/38/article.
pdf.

Tonnis:2013:RIC

[TPK13] Marcus Tönnis, David A.

Plecher, and Gudrun Klinker. Representing information

— classifying the Augmented Reality presentation space. Computers and Graphics, 37(8):997– 1011, December 2013. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001313.

Tonneau:2014:UTE

[TPM14] Steve Tonneau, Julien Pettré, and Franck Multon. Using task efficient contact configurations to animate creatures in arbitrary environments. Computers and Graphics, 45(??):40– 50, December 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931400079X.

Tost:1995:VVA

[TPN95] D. Tost, A. Puig, and I. Navazo. A volume visualization algorithm using a coherent extended weight matrix. Computers and Graphics, 19(1):37– 45, January–February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1&aid=9400120.

Toss:2018:PMQ

[TPRC18] Julio Toss, Cícero A. L. Pahins, Bruno Raffin, and João L. D. Comba. Packed-Memory Quadtree: a cacheoblivious data structure for visual exploration of streaming spatiotemporal big data. Computers and Graphics, 76(??):117–128, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301390.

Thomas:1995:VMV

[TR95] Andrew D. H. Thomas and Michael G. Rodd. Verifying machine vision systems by digital montage. Computers and Graphics, 19(3): 365–371, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500007.

Thompson:2022:SPC

[TRB^{+}22] Elia Moscoso Thompson, Andrea Ranieri, Silvia Biasotti, Miguel Chicchon, Ivan Sipiran, Minh-Khoi Pham, Thang-Long Nguyen-Ho, Hai-Dang Nguyen, and Minh-Triet Tran. SHREC 2022: Pothole and crack detection in the road pavement using images and RGB-D data. Computers and Graphics, 107(??):161– 171, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001418.

Tao:2022:LVM

[TRLX22] Rui Tao, Hongxiang Ren, Jun Liu, and Fangbing Xiao. A Lagrangian vortex method for smoke simulation with two-way fluidsolid coupling. Computers and Graphics, 107(??):289– 302, October 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001571.

Truckenbrod:1986:CCI

[Tru86] Joan Truckenbrod. Creative computer imaging. Computers and Graphics, 10(2): 191–197, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tripp:1975:ITD

[TS75] E. R. Tripp III and J. R.

Suttle. An interactive threedimensional color graphics system. Computers and Graphics, 1(2–3):211–214, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tung:1987:ILU

[TS87] Robert Shieng-I Tung and Jerry Schneider. The investigation of land-use/ transportation interrelationships: a CAD approach. Computers and Graphics, 11 (3):249–254, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Teixeira:1994:CGC

[TS94] Jose Carlos Teixeira and Joaquim Silvestre Madeira. Computer graphics curriculum at the University of Coimbra. Computers and Graphics, 18(3):309–314, May–June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tollmar:1995:DBG

[TS95] Konrad Tollmar and Yngve Sundblad. The design and building of the graphic user interface for the collaborative desktop. Computers and Graphics, 19(2): 179–188, March–April 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.

cgi?year=1995&volume=19&
issue=2&aid=9400141.

Tian:2016:FSS

[TSC16] Feng. Tian, Minghui Sun, and Maiga Chang. Foreword to the special section on the international conference on e-learning and games 2016 (edutainment ’16). Computers and Graphics, 61 (??):A2–A3, December 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301212.

Tamma:1987:ERR

[TSD87] Kumar K. Tamma, William B.

Squire, and Brian L. Dowler. An entire real range plotting algorithm via computer graphics. Computers and Graphics, 11(3): 305–308, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2007:MBM

[TSD^{+}07] H. Tang, H. Z. Shu, J. L.

Dillenseger, X. D. Bao, and L. M. Luo. Momentbased metrics for mesh simplification. Computers and Graphics, 31(5):710– 718, October 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930700132X.

Tai:1998:RGB

[TSK98] Chiew-Lan Tai, Yoshihisa Shinagawa, and Tosiyasu L. Kunii. A Reeb graphbased representation for non-sequential construction of topologically complex shapes. Computers and Graphics, 22(2–3):255–268, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/570.pdf.

Tsuchie:2017:RUS

[Tsu17] Shoichi Tsuchie. Reconstruction of underlying surfaces from scanned data using lines of curvature. Computers and Graphics, 68(??): 108–118, November 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301504.

Taketomi:2011:RTA

[TSY11] Takafumi Taketomi, Tomokazu Sato, and Naokazu Yokoya. Real-time and accurate extrinsic camera parameter estimation using feature landmark database for augmented reality. Computers and Graphics, 35(4):768– 777, August 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001105.

Tran-Thong:1982:SLA

[TT82] Tran-Thong. A symmetric linear algorithm for line segment generation. Computers and Graphics, 6(1):15–17, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tran-Thong:1983:EAE

[TT83] Tran-Thong. Ellipse, arc of ellipse and elliptic spline. Computers and Graphics, 7 (2):169–175, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tu:2012:HCD

[TT12] Shih-Chun Tu and Wen-Kai Tai. A high-capacity datahiding approach for polygonal meshes using maximum expected level tree. Computers and Graphics, 36 (6):767–775, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001185.

Telea:2019:FSS

[TT19] Alexandru Telea and Theoharis Theoharis. Foreword to the special section on 3D Object Retrieval (3DOR2018). Computers and Graphics, 79(??):A3–A4, April 2019.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300159.

Ton-That:2023:PBN

[TTKA23] Quoc-Minh Ton-That, Paul G. Kry, and Sheldon Andrews. Parallel block neo-Hookean XPBD using graph clustering. Computers and Graphics, 110 (??):1–10, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200187X.

Taranta:2016:SAG

[TVL16] Eugene M. Taranta II, Andrés N. Vargas, and Joseph J. LaViola, Jr. Streamlined and accurate gesture recognition with Penny Pincher. Computers and Graphics, 55(??): 130–142, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001788.

Triantafyllidis:2003:EAE

[TVS^{+}03] G. A. Triantafyllidis, M. Varnuska, D. Sampson, D. Tzovaras, and M. G. Strintzis. An efficient algorithm for the enhancement of JPEGcoded images. Computers and Graphics, 27(4):529– 534, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2024:SIS

[TW24] Kaiyuan Tang and Chaoli Wang. STSR-INR: Spatiotemporal super-resolution for multivariate time-varying volumetric data via implicit neural representation. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000013.

Tang:2003:RBC

[TWBP03] Ying Tang, Jin Wang, Hujun Bao, and Qunsheng Peng. RBF-based constrained texture mapping. Computers and Graphics, 27(3):415– 422, June 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tang:2022:PCD

[TWNL22] Jing Tang, Yinghui Wang, Xiaojuan Ning, and Ke Lv. Point cloud decomposition by internal and external critical points. Computers and Graphics, 102(??):18– 29, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100248X.

Tricoche:2002:VVL

[TWSH02] X. Tricoche, T. Wischgoll, G. Scheuermann, and H. Hagen. Visualization of very large datasets: Topology tracking for the visualization of time-dependent twodimensional flows. Computers and Graphics, 26(2):249– 257, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/32/abstract.
html.

Tsuchie:2024:CEL

[TY24] Shoichi Tsuchie and Norimasa Yoshida. α-curves: Extended log-aesthetic curves with variable shape parameter. Computers and Graphics, 118(??):60–70, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300287X.

Tang:2012:GAC

[TyZfTM12] Min Tang, Jie yi Zhao, Ruo feng Tong, and Dinesh Manocha. GPU accelerated convex hull computation. Computers and Graphics, 36 (5):498–506, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000544.

Tan:2022:CFP

[TZT^{+}22] Xuefeng Tan, Dejun Zhang, Long Tian, Yiqi Wu, and Yilin Chen. Coarse-to-fine pipeline for 3D wireframe reconstruction from point cloud. Computers and Graphics, 106(??):288– 298, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001194.

Tian:2021:UMA

[TZvD^{+}21] Zonglin Tian, Xiaorui Zhai, Daan van Driel, Gijs van Steenpaal, Mateus Espadoto, and Alexandru Telea. Using multiple attribute-based explanations of multidimensional projections to explore high-dimensional data. Computers and Graphics, 98 (??):93–104, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000765.

Ugail:1999:MPS

[UBW99] Hassan Ugail, Malcolm I. G.

Bloor, and Michael J. Wilson. Manipulation of PDE surfaces using an interactively defined parameterisation. Computers and Graphics, 23(4):525–534, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/24/34/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/24/34/33/article.
                                        pdf.

Ugail:2006:MTP

[Uga06]  Hassan Ugail.  Method

of trimming PDE surfaces. Computers and Graphics, 30(2):225–232, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000306.

Haque:2006:DLU

[uHRBK06] Asif ul Haque, Mohammad Saifur Rahman, Mehedi Bakht, and M. Kaykobad. Drawing lines by uniform packing. Computers and Graphics, 30(2): 207–212, April 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000288.

Unger:2013:SVI

[UKL^{+}13] Jonas Unger, Joel Kronander, Per Larsson, Stefan Gustavson, Joakim Löw, and Anders Ynnerman. Spatially varying image based lighting using HDR-video.

Computers and Graphics, 37 (7):923–934, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001052.

Unterguggenberger:2023:VAW

[UKW23] Johannes Unterguggenberger, Bernhard Kerbl, and Michael Wimmer. Vulkan all the way: Transitioning to a modern low-level graphics API in academia. Computers and Graphics, 111 (??):155–165, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000249.

Urribarri:2022:VTA

[UL22] Dana K. Urribarri and Martín L. Larrea. A visualization technique to assist in the comparison of large meteorological datasets. Computers and Graphics, 104 (??):1–10, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000322.

Uchiki:1983:CDM

[UOT83] T. Uchiki, T. Ohashi, and M. Tokoro. Collision detection in motion simulation. Computers and Graphics, 7(3–4):285–293, 1983.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Urena:1997:FCG

[UPT97] C. Ureña, X. Pueyo, and J. C. Torres. A formalization and classification of global illumination methods. Computers and Graphics, 21(2): 225–236, March–April 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=2&aid=9600086.

Urena:1992:OOA

[UPTd92] C. Urena, J. Parets, J. C. Torres, and V. del Sol. An object-oriented approach to ray tracing image synthesis implementations. Computers and Graphics, 16 (4):363–368, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Uchida:2020:SLD

[US20] Mitsuhiro Uchida and Suguru Saito. Stylized linedrawing of 3D models using CNN with line property encoding. Computers and Graphics, 91(??):252– 264, October 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301114.

Ucoluk:1999:ARB

[ÜT99] Göktürk Üçoluk and I. Hakk Toroslu. Automatic reconstruction of broken 3-D surface objects. Computers and Graphics, 23 (4):573–582, August 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/37/article.
pdf.

Ullman:1990:IDM

[UWC90] David G. Ullman, Stephen Wood, and David Craig. The importance of drawing in the mechanical design process. Computers and Graphics, 14(2):263–274, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vandewalle:1996:TML

[VA96] N. Vandewalle and M. Ausloos. A toy model for life at the “edge of chaos”. Computers and Graphics, 20(6):921–923, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/

                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=6&aid=9600061.

Vaananen:1993:IHC

[Vää93]  Kaisa Väänänen.  Inter-

faces to hypermedia: Communicating the structure and interaction possibilities to the users. Computers and Graphics, 17(3): 219–228, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Verges:2008:RQP

[VAGT08] Eduard Vergés, Dolors Ayala, Sergi Grau, and Dani Tost. 3D reconstruction and quantification of porous structures. Computers and Graphics, 32 (4):438–444, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000459.

VanWijk:1985:RTO

[Van85] J. J. VanWijk. Ray tracing objects defined by sweeping a sphere. Computers and Graphics, 9(3):283–290, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

vanEmmerik:1989:CMP

[van89a] Maarten J. G. M. van Emmerik. Creation and modification of parametrized solid models by graphical interaction. Computers and Graphics, 13(1):71–76, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vandoni:1989:DLG

[Van89b] Carlo E. Vandoni. Development of a large graphicsbased application package. Computers and Graphics, 13 (2):243–252, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

VanArsdale:1994:HTM

[Van94] Daniel VanArsdale. Homogeneous transformation matrices for computer graphics. Computers and Graphics, 18(2):177–191, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

VanLoocke:2009:NLI

[Van09] Philip Van Loocke. Nonlinear iterated function systems and the creation of fractal patterns over regular polygons. Computers and Graphics, 33(6): 698–704, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930900082X.

VanLoocke:2010:CBO

[Van10] Philip Van Loocke. Combination of basic origami with fractal iteration. Computers and Graphics, 34 (1):66–71, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309001149.

Varma:1992:CGP

[Var92] Jayanth Rama Varma. Computer graphics, peripheral vision and non-Euclidean geometry. Computers and Graphics, 16(3):253– 258, Fall 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Visvalingam:1999:CGD

[VB99] M. Visvalingam and C. I. Brown. Chaos and graphics — the deconstruction of teragons into decogons. Computers and Graphics, 23(1):155–167, February 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1999/23/
                                        1/657.pdf.

Viola:2017:FSS

[VB17] Ivan Viola and Jirí Bittner. Foreword to the special section on SCCG 2017.

Computers and Graphics, 67(??):A1, October 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301036.

Varona:2005:HFT

[VBP05] Javier Varona, José M.

Buades, and Francisco J. Perales. Hands and face tracking for VR applications. Computers and Graphics, 29(2):179–187, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vallet:2015:TIU

[VBS^{+}15] Bruno Vallet, Mathieu Brédif, Andres Serna, Beatriz Marcotegui, and Nicolas Paparoditis. TerraMobilita/iQmulus urban point cloud analysis benchmark. Computers and Graphics, 49(??):126–133, June 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500028X.

vanBlokland:2020:ISD

[vBT20a] Bart Iver van Blokland and Theoharis Theoharis. An indexing scheme and descriptor for 3D object retrieval based on local shape querying. Computers and Graphics, 92(??):55– 66, November 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932030128X.

vanBlokland:2020:RIC

[vBT20b] Bart Iver van Blokland and Theoharis Theoharis. Radial intersection count image: a clutter resistant 3D shape descriptor. Computers and Graphics, 91 (??):118–128, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301096.

vanBlokland:2021:POR

[vBT21] Bart Iver van Blokland and Theoharis Theoharis. Partial 3D object retrieval using local binary QUICCI descriptors and dissimilarity tree indexing. Computers and Graphics, 100 (??):32–42, November 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001497.

Vais:2013:LFL

[VBTW13] Alexander Vais, Daniel Brandes, Hannes Thielhelm, and Franz-Erich Wolter. Laplacians on flat line bundles over 3-manifolds. Computers and Graphics, 37 (6):718–729, October 2013.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000873.

Velikhov:1988:CAS

[VBVS88] E. P. Velikhov, V. B. Betelin, Ju. S. Vishnjakov, and A. I. Stavitskij. Computer-aided strength calculation system. Computers and Graphics, 12 (3–4):415–424, 1988. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vais:2012:SCN

[VBW12] Alexander Vais, Benjamin Berger, and Franz-Erich Wolter. Spectral computations on nontrivial line bundles. Computers and Graphics, 36(5):398–409, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000660.

Vera:2007:CGA

[VCHR07] Lucia Vera, Ruben Campos, Gerardo Herrera, and Cristina Romero. Computer graphics applications in the education process of people with learning difficulties. Computers and Graphics, 31 (4):649–658, August 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000933.

Vivo:1992:SPM

[VCQ92] Roberto Vivo, Emilio Camahort, and Ricardo Quiros. A secondary parametric model for CSG. Computers and Graphics, 16 (4):369–373, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Visvalingam:1998:ASS

[VD98] Mahes Visvalingam and Kurt Dowson. Algorithms for sketching surfaces. Computers and Graphics, 22(2–3):269–280, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/571.pdf.

vandenBerg:2007:PCB

[vdBB07] Eelco van den Berg and Willem F. Bronsvoort. Parameterised, constraint-based wrapping of freeform shapes. Computers and Graphics, 31(1):89–99, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306001786.

vanderLinden:2023:SVT

[vdLdFvdEV23] Sanne van der Linden, Evie de Fouw, Stef van den Elzen, and Anna Vilanova. A survey of visualization techniques for comparing event sequences. Computers and Graphics, 115 (??):522–542, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000717.

vanDam:2002:EIV

[vDLS02] Andries van Dam, David H.

Laidlaw, and Rosemary Michelle Simpson. Experiments in immersive virtual reality for scientific visualization. Computers and Graphics, 26(4):535–555, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/31/abstract.
html.

Vock:2019:FTM

[VDOK19] Richard Vock, Alexander Dieckmann, Sebastian Ochmann, and Reinhard Klein. Fast template matching and pose estimation in 3D point clouds. Computers and Graphics, 79(??): 36–45, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300081.

vandePanne:2008:CGS

[vdPS08] Michiel van de Panne and Eric Saund. Computers and Graphics special issue on EG SBIM 2007. Computers and Graphics, 32(5): 485, October 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308000940.

vanderHelm:1998:MMS

[vEB98] Aadjan van der Helm, Peter Ebell, and Willem F. Bronsvoort. Modelling mollusc shells with generalized cylinders. Computers and Graphics, 22(4):505–513, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        4/582.pdf.

Vazquez:2006:RAS

[VFSL06] Pere-Pau Vázquez, Miquel Feixas, Mateu Sbert, and Antoni Llobet. Realtime automatic selection of good molecular views. Computers and Graphics, 30 (1):98–110, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002141.

Valette:2004:CTM

[VGP04] Sébastien Valette, Alexandre Gouaillard, and Rémy Prost. Compression of 3D triangular meshes with progressive precision. Computers and Graphics, 28 (1):35–42, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vaya:2002:TSF

[VH02] Ramón Mollá Vayá and Roberto Vivó Hernando. Technical section: Fixedpoint digital differential analyser with antialiasing (FDDAA). Computers and Graphics, 26(2):329–339, April ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/41/38/abstract.
html.

Vierjahn:2015:SRG

[VH15] Tom Vierjahn and Klaus Hinrichs. Surface-reconstructing growing neural gas: a method for online construction of textured triangle meshes. Computers and Graphics, 51(??):190– 201, October 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000631.

Viola:2010:ENS

[VHE10] Ivan Viola, Helwig Hauser, and David Ebert. Editorial note for special section on illustrative visualization. Computers and Graphics, 34 (4):335–336, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000786.

VanAssche:2004:MCP

[VHON04] Steven Van Assche, Filip Hendrickx, Nico Oorts, and Lode Nachtergaele. Multichannel publishing of interactive multimedia presentations. Computers and Graphics, 28(2):193– 206, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vosough:2018:PHV

[VHR^{+}18] Zana Vosough, Marius Hogräfer, Loïc A. Royer, Rainer Groh, and Hans-Jörg Schulz. Parallel hierarchies: a visualization for cross-tabulating hierarchical categories. Computers and Graphics, 76 (??):1–17, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301080.

Vinkler:2012:VDB

[VHS12] Marek Vinkler, Vlastimil Havran, and Jirí Sochor. Visibility driven BVH build up algorithm for ray tracing. Computers and Graphics, 36(4):283–296, June 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000362.

vanWijk:1984:RRG

[vJ84] Jarke J. van Wijk and Frederik W. Jansen. Realism in raster graphics. Computers and Graphics, 8(2): 217–219, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

VanLoocke:2006:SBF

[VJ06] Philip Van Loocke and Yannick Joye. Symmetry breaking in fields as a methodology for three-dimensional fractal form generation. Computers and Graphics, 30 (5):843–853, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001324.

Vanecek:2007:CTS

[VK07] Petr Vanecek and Ivana Kolingerová. Comparison of triangle strips algorithms. Computers and Graphics, 31 (1):100–118, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306002226.

Vizcay:2023:DEC

[VKA^{+}23] Sebastian Vizcay, Panagiotis Kourtesis, Ferran Argelaguet, Claudio Pacchierotti, and Maud Marchal. Design, evaluation and calibration of wearable electrotactile interfaces for enhancing contact information in virtual reality. Computers and Graphics, 111 (??):199–212, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000146.

Veron:1998:TSS

[VL98] P. Veron and J. C. Leon. Technical section — shape preserving polyhedral simplification with bounded error. Computers and Graphics, 22(5):565–585, October 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
5/597.pdf.

Vidal:2015:LBH

[VLD15] Vincent Vidal, Guillaume Lavoué, and Florent Dupont.

Low budget and high fidelity relaxed 567-remeshing. Computers and Graphics, 47(??): 16–23, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001307.

vanLiere:2009:FSIb

[vLLSM09] Robert van Liere, Robert S.

Laramee, Gerik Scheuermann, and Kwan-Liu Ma. Foreword to special issue on knowledge assisted visualization. Computers and Graphics, 33(5):583– 584, October 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000922.

vanLiere:2009:FSIa

[vLM09] Robert van Liere and Betty Mohler. Foreword to: Special issue on virtual environments. Computers and Graphics, 33(2): 119, April 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000181.

vanLankveld:2011:IRL

[vLvKV11] Thijs van Lankveld, Marc van Kreveld, and Remco Veltkamp. Identifying rectangles in laser range data for urban scene reconstruction. Computers and Graphics, 35(3):719–725, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000434.

Venkataraman:2015:ACP

[VM15] Vishwanath A. Venkataraman and Ramanathan Muthuganapathy. Algorithm for computing positive α-hull for a set of planar closed curves. Computers and Graphics, 51 (??):125–135, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315000655.

Vieira:2016:EAI

[VMAL16] Thales Vieira, Dimas Martínez, Maria Andrade, and Thomas Lewiner. Estimating affineinvariant structures on triangle meshes. Computers and Graphics, 60(??):83– 92, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300887.

Verhulst:2023:DWE

[VNMP23] Adrien Verhulst, Jean-Marie Normand, Guillaume Moreau, and Gustavo Patow. Deep weathering effects. Computers and Graphics, 112(??):40–49, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849323000353.

vanOnzenoodt:2020:PIS

[vOHR20] Christian van Onzenoodt, Anke Huckauf, and Timo Ropinski. On the perceptual influence of shape overlap on data-comparison using scatterplots. Computers and Graphics, 90 (??):169–181, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300881.

vanOosterom:2015:MPC

[vOMRI^{+}15] Peter van Oosterom, Oscar Martinez-Rubi, Milena Ivanova, Mike Horhammer, Daniel Geringer, Siva Ravada, Theo Tijssen, Martin Kodde, and Romulo Gonçalves. Massive point cloud data management: Design, implementation and execution of a point cloud benchmark. Computers and Graphics, 49(??):92– 125, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000084.

Veen:1977:TCG

[VP77] Arthur Veen and Lee D. Peachey. Trots: a computer graphics system for threedimensional reconstruction from serial sections. Computers and Graphics, 2(3): 135–150, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Voyatzis:1998:DIW

[VP98] G. Voyatzis and I. Pitas. Digital image watermarking using mixing systems. Computers and Graphics, 22(4): 405–416, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/4/564.pdf.

Vigo:2000:CDC

[VP00] Marc Vigo and Núria Pla. Computing directional constrained Delaunay triangulations. Computers and Graphics, 24(2): 181–190, April 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/47/27/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/47/29/27/article.
                                        pdf.

Vrolijk:2006:ICI

[VP06] Benjamin Vrolijk and Frits H.

Post. Interactive out-of-core isosurface visualisation in time-varying data sets. Computers and Graphics, 30(2):265–276, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000343.

Velho:2002:ASM

[VPBY02] Luiz Velho, Ken Perlin, Henning Biermann, and Lexing Ying. Algorithmic shape modeling with subdivision surfaces. Computers and Graphics, 26(6): 865–875, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Valette:2006:SPS

[VPLL06] Gilles Valette, Stéphanie Prévost, Laurent Lucas, and Joël Léonard. SoDA project: a simulation of soil surface degradation by rainfall. Computers and Graphics, 30 (4):494–506, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000823.

Vaidyanathan:1984:Z

[VR84] C. V. Vaidyanathan and N. V. Raman. Zooming.

Computers and Graphics, 8 (2):203–205, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Varsa:2006:SBL

[VR06] Petri Varsa and Jon Rokne. Simulation of the ball lightning phenomenon. Computers and Graphics, 30 (4):485–493, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000860.

Vinayak:2016:EHG

[VR16] Vinayak and Karthik Ramani. Extracting hand grasp and motion for intent expression in mid-air shape deformation: a concrete and iterative exploration through a virtual pottery application. Computers and Graphics, 55(??):143–156, April 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500179X.

vonRadziewsky:2016:OSD

[vRESH16] Philipp von Radziewsky, Elmar Eisemann, Hans-Peter Seidel, and Klaus Hildebrandt. Optimized subspaces for deformation-based modeling and shape interpolation. Computers and Graphics, 58(??):128– 138, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300590.

Vicent:2005:PPB

[VRV05] María J. Vicent, Vicente Rosell, and Roberto Vivó. A polar-plane-based method for natural illumination of plants and trees. Computers and Graphics, 29(2):203– 208, April 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Villar:2003:IPP

[VSKG03] Nicolas Villar, Albrecht Schmidt, Gerd Kortuem, and Hans-Werner Gellersen. Interacting with proactive public displays. Computers and Graphics, 27(6): 849–857, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Veelaert:2006:CSA

[VT06] Peter Veelaert and Kristof Teelen. Consensus sets for affine transformation uncertainty polytopes. Computers and Graphics, 30 (1):77–85, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930500213X.

Voinea:2007:VDM

[VT07] Lucian Voinea and Alexandru Telea. Visual data mining and analysis of software repositories. Computers and Graphics, 31(3): 410–428, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000623.

Vrotsou:2022:FSS

[VT22] Katerina Vrotsou and Cagatay Turkay. Foreword to the special section on visual analytics. Computers and Graphics, 103(??): A3–A4, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000401.

vanToll:2021:PSC

[vTCB^{+}21] Wouter van Toll, Thomas Chatagnon, Cédric Braga, Barbara Solenthaler, and Julien Pettré. SPH crowds: Agent-based crowd simulation up to extreme densities using fluid dynamics. Computers and Graphics, 98 (??):306–321, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001205.

vanToll:2020:SNA

[vTP20] Wouter van Toll and Julien Pettré. Synchronizing navigation algorithms for crowd simulation via topological strategies. Computers and Graphics, 89(??): 24–37, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300418.

vanToll:2020:CNM

[vTTK^{+}20] Wouter van Toll, Roy Triesscheijn, Marcelo Kallmann, Ramon Oliva, Nuria Pelechano, Julien Pettré, and Roland Geraerts. Comparing navigation meshes: Theoretical analysis and practical metrics. Computers and Graphics, 91 (??):52–82, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300959.

denHerrewegen:2023:SSL

[VTW23] Jarne Van den Herrewegen, Tom Tourwé, and Francis Wyffels. Self-supervised learning for robust object retrieval without human annotations. Computers and Graphics, 115 (??):13–24, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001231.

VanOosterom:1989:OOA

[VV89] Peter Van Oosterom and Jan Van Den Bos. An object-oriented approach to the design of geographic information systems. Computers and Graphics, 13(4): 409–418, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vieira:2012:RVC

[VVCN12] Roberto C. Cavalcante Vieira, Creto A. Vidal, and Joaquim Bento Cavalcante-Neto. Reproducing virtual characters. Computers and Graphics, 36(2): 80–91, April 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311001737.

Voltoline:2021:MVC

[VW21] Raphael Voltoline and Shin-Ting Wu. Multimodal visualization of complementary color-coded FA map and tensor glyphs for interactive tractography ROI seeding. Computers and Graphics, 96(??):24–35, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000315.

vanWezel:2023:VRT

[vWdlHFK23] Chris S. van Wezel, Willard A. Verschoore de la Houssaije, Steffen Frey, and Jirí Kosinka. Virtual Ray Tracer 2.0. Computers and Graphics, 111(??):89– 102, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000067.

vanWijk:2004:IBR

[vWS04] J. J. van Wijk and D. Saupe.

Image based rendering of iterated function systems. Computers and Graphics, 28 (6):937–943, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Vanian:2022:IPD

[VZP22] Vazgen Vanian, Georgios Zamanakos, and Ioannis Pratikakis. Improving performance of deep learning models for 3D point cloud semantic segmentation via attention mechanisms. Computers and Graphics, 106 (??):277–287, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001133.

Wood:1975:CTM

[WA75] P. M. Wood and D. M.

Austin. CARTE: a thematic mapping program. Computers and Graphics, 1(2– 3):239–250, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wu:2002:CNO

[WA02] Shin-Ting Wu and Osmar Aléssio. Complete and non-overlapping marching along a closed regular intersection curve. Computers and Graphics, 26(6): 853–864, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walker:1989:BOE

[Wal89] Marshall Walker. Boolean operations with enriched octtree structures. Computers and Graphics, 13(4): 487–495, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walter:1993:CGS

[Wal93a]  David Walter.  Chaos

and graphics: Systemised serendipity for producing computer art. Computers and Graphics, 17(6):699– 700, November–December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walter:1993:CAR

[Wal93b] David John Walter. Computer art representing the behavior of the Newton– Raphson method. Computers and Graphics, 17(4): 487–488, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walker:1994:IHB

[Wal94a]  Marshall Walker.  Inter-

polation with hybrid Bspline surfaces. Computers and Graphics, 18(4): 525–530, July–August 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walter:1994:CAN

[Wal94b] David Walter. Computer art from Newton’s, secant, and Richardson’s methods. Computers and Graphics, 18(1): 127–131, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walsh:2006:RBU

[Wal06] James A. Walsh. Reverse bifurcations in a unimodal queueing model. Computers and Graphics, 30(4):650– 657, August 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000884.

Wretborn:2017:ACP

[WAM17] Joel Wretborn, Rickard Armiento, and Ken Museth.

Animation of crack propagation by means of an extended multi-body solver for the material point method. Computers and Graphics, 69(??):131–139, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301711.

Wang:2004:CIM

[Wan04] Charlie C. L. Wang. Cyber-Tape: an interactive measurement tool on polyhedral surface. Computers and Graphics, 28(5):731– 745, October 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2023:NIS

[Wan23] Cuilan Wang. A novel isosurface segmentation method using common boundary tests. Computers and Graphics, 110(??):86–97, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002266.

Warner:1976:DAM

[War76] J. R. Warner. Design applications of the MIDAS graphics system. Computers and Graphics, 2(1):15–22, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Woodworth:2024:VCV

[WB24] Jason W. Woodworth and Christoph W. Borst. Visual cues in VR for guiding attention vs. restoring attention after a short distraction. Computers and Graphics, 118(??):194– 209, February 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003023.

Walther:2016:TMC

[WBA16] Jeppe U. Walther, J. Andreas Bærentzen, and Henrik Aanæs. Tangible 3D modeling of coherent and themed structures. Computers and Graphics, 58 (??):53–65, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300474.

Wand:2008:PIE

[WBB^{+}08] Michael Wand, Alexander Berner, Martin Bokeloh, Philipp Jenke, Arno Fleck, Mark Hoffmann, Benjamin Maier, Dirk Staneker, Andreas Schilling, and Hans-Peter Seidel. Processing and interactive editing of huge point clouds from 3D scanners. Computers and Graphics, 32(2): 204–220, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000253.

Wagner:2021:PII

[WBJ^{+}21] Sebastian Wagner, Julia Belger, Fabian Joeres, Angelika Thöne-Otto, Christian Hansen, Bernhard Preim, and Patrick Saalfeld. iVRoad: Immersive virtual road crossing as an assessment tool for unilateral spatial neglect. Computers and Graphics, 99 (??):70–82, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100128X.

Wan:1998:TSB

[WBK98] Ming Wan, Steve Bryson, and Arie Kaufman. Technical section — boundary cell-based acceleration for volume ray casting. Computers and Graphics, 22(6):715–722, December 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
6/604.pdf.

White:1997:TDG

[WBL^{+}97] Martin White, Mike Bassett, Dairsie Latimer, Shaun Mc-Cann, Alex Makris, Marcus Waller, Graham Dunnett, Joachim Binder, and Paul Lister. The TAYRA 3-D graphics raster processor. Computers and Graphics, 21(2):129–142, March–April 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=2&aid=9600076.

Wang:1992:NPR

[WBP92] Mingfu Wang, Hujun Bao, and Qunsheng Peng. A new progressive radiosity algorithm through the use of accurate form-factors. Computers and Graphics, 16(3): 303–309, Fall 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Waldin:2016:PCM

[WBRV16] Nicholas Waldin, Matthias Bernhard, Peter Rautek, and Ivan Viola. Personalized 2D color maps. Computers and Graphics, 59 (??):143–150, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300772.

Welstead:1989:CPT

[WC89] Stephen T. Welstead and Thomas L. Cromer. Coloring periodicities of twodimensional mappings. Computers and Graphics, 13(4): 539–543, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wei:2011:IBC

[WCA^{+}11] Yiyi Wei, Stéphane Cotin, Jérémie Allard, Le Fang, Chunhong Pan, and Songde Ma. Interactive blood-coil simulation in real-time during aneurysm embolization. Computers and Graphics, 35(2):422–430, April 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000197.

Wu:1998:SVP

[WCdA98] Shin-Ting Wu, Sidney P. Campos, and Marcus A. M. de Aguiar. Scientific visualization of Poincaré maps. Computers and Graphics, 22(2–3):209–216, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/cag/sub/1998/22/
                                        2-3/543.pdf.

Wang:2023:TTD

[WCF^{+}23] Yi Wang, Jing-Song Cheng, Qiao Feng, Wen-Yuan Tao, Yu-Kun Lai, and Kun Li. TSNeRF: Text-driven stylized neural radiance fields via semantic contrastive learning. Computers and Graphics, 116(??):102–114, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001796.

Wang:2011:SAD

[WCH^{+}11] Tinghuai Wang, John Collomosse, Rui Hu, David Slatter, Darryl Greig, and Phil Cheatle. Stylized ambient displays of digital media collections. Computers and Graphics, 35(1):54– 66, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001731.

Wei:2024:HHP

[WCH^{+}24] Jiangxia Wei, Lan Cheng, Zhimin Hu, Mifeng Ren, Xinying Xu, and Gaowei Yan. HPNet: High precision point cloud registration using feature pyramid and hybrid position encoding. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000232.

Wang:2022:CAM

[WCHM22] Yong Wang, Jin Chen, Yujuan Han, and Duoqian Miao. Combining attention mechanism and Retinex model to enhance low-light images. Computers and Graphics, 104(??): 95–105, May 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000541.

Wang:2023:NMS

[WCL23] Meng Wang, Jiaxing Chen, and Haipeng Liu. A novel multi-scale architecture driven by decoupled semantic attention transfer for person image generation. Computers and Graphics, 111(??):24–36, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000043.

Wang:2021:PAA

[WCLT21] Yechao Wang, Jinming Cao, Yangyan Li, and Changhe Tu. APM: Adaptive permutation module for point cloud classification. Computers and Graphics, 97 (??):217–224, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000741.

Wu:2014:TSD

[WCLZ14] Kan Wu, Li Chen, Jing Li, and Yanheng Zhou. Tooth segmentation on dental meshes using morphologic skeleton. Computers and Graphics, 38(??): 199–211, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001696.

Wang:2023:PTV

[WCS23] Liwen Wang, Shaoyu Cai, and Christian Sandor. Perceptual thresholds of visual size discrimination in augmented and virtual reality. Computers and Graphics, 117(??):105–113, December 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002388.

Wang:2024:SSN

[WCW^{+}24] Haocheng Wang, Yanlong Cao, Xiaoyao Wei, Yejun Shou, Lingfeng Shen, Zhijie Xu, and Kai Ren.

Structerf-SLAM: Neural im-

plicit representation SLAM for structural environments. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000207.

Wu:2020:CSC

[WCY^{+}20] Zizhao Wu, Xingyu Chen, Lingyun Yu, Alexandru Telea, and Jirí Kosinka. Coskeletons: Consistent curve skeletons for shape families. Computers and Graphics, 90(??):62–72, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300571.

Wither:2009:AOA

[WDH09] Jason Wither, Stephen Di-Verdi, and Tobias Höllerer. Annotation in outdoor augmented reality. Computers and Graphics, 33(6): 679–689, December 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000867.

Weber:1984:MGD

[Web84]  H. R. Weber.  Models

for graphics dialogue programming. Computers and Graphics, 8(2):207–210, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weber:1985:MMM

[Web85] H. R. Weber. Meditation on man-machine interfaces or our personal role in graphics dialogue programming. Computers and Graphics, 9 (3):237–245, 1985. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weber:1987:ZDC

[Web87] H. R. Weber. ZGDV — the Darmstadt computer graphics center. Computers and Graphics, 11(1):51–53, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weck:1979:GSU

[Wec79] G. Weck. Graphic system utilizing an associative data structure. Computers and Graphics, 4(1):43–50, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weeks:2021:BCC

[Wee21] Jeff Weeks. Body coherence in curved-space virtual reality games. Computers and Graphics, 97(??): 28–41, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000443.

Weissflog:1984:EDI

[Wei84a] Uwe Weissflog. Experience in design and implementation of an IGES translator. Computers and Graphics, 8(3):269–273, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weitek:1984:WSM

[Wei84b]  Weitek.  Weitek solids

modeling engine. Computers and Graphics, 8(4): 437, ???? 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0097849384900426.

Weichsel:1999:PFU

[Wei99] Joel I. Weichsel. Pattern formation under various tiling rules. Computers and Graphics, 23(5):751–762, October 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/32/
38/article.pdf;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/34/38/abstract.
html.

Welsch:1976:GDA

[Wel76] R. E. Welsch. Graphics for data analysis. Computers and Graphics, 2(1):31–37, 1976. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wenz:1984:SPM

[Wen84]  H. Wenz.  Space-plot-

modelling in a volume system. Computers and Graphics, 8(2):199–202, 1984. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

West:1994:AIR

[Wes94] Thomas G. West. Advanced interaction: a return to mental models and learning by doing. Computers and Graphics, 18(5):685–689, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Waller:1999:TSE

[WEWL99] Marcus D. Waller, Jon P.

Ewins, Martin White, and Paul F. Lister. Technical section — efficient primitive traversal using adaptive linear edge function algorithms. Computers and Graphics, 23(3):365– 375, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/688.pdf.

Weiss:1988:CAT

[WF88] Gunter Weiss and Peter Furtner. Computer-aided treatment of developable surfaces. Computers and Graphics, 12(1):39–51, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wei:2011:RTR

[WF11] Feifei Wei and Jieqing Feng. Real-time ray casting of algebraic B-spline surfaces. Computers and Graphics, 35 (4):800–809, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931100104X.

Wyvill:2022:CBF

[WF22] Brian Wyvill and Bianca Falcidieno. A conversation with Bianca Falcidieno: SMI 2021 Tosiyasu Kunii Achievement Award winner. Computers and Graphics, 105(??):A4–A8, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000838.

Wallraven:2009:CAC

[WFC^{+}09] Christian Wallraven, Roland Fleming, Douglas Cunningham, Jaume Rigau, Miquel Feixas, and Mateu Sbert. Categorizing art: Comparing humans and computers. Computers and Graphics, 33 (4):484–495, August 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000612.

Wang:2014:EBS

[WFC14] Ming Wang, Jie-Qing Feng, and Wei Chen. Efficient boundary surface reconstruction from heterogeneous volumetric data via tri-prism decomposition. Computers and Graphics, 38 (??):212–221, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001751.

Wang:2003:SRP

[WFG03] Yigang Wang, Bernd Fröhlich, and Martin Göbel. Selective refinement of progressive meshes using vertex hierarchies. Computers and Graphics, 27(2):179– 188, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wazzan:1982:SBP

[WFS^{+}82] G. Wazzan, W. R. Franklin, W. R. Spillers, A. Greenwood, H. Chu, and T. F. Garrity. Simulation of buried power transmission systems: Some computer graphics options. Computers and Graphics, 6(1):7–14, 1982. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wodniok:2017:CBV

[WG17] Dominik Wodniok and Michael Goesele. Construction of bounding volume hierarchies with SAH cost approximation on temporary subtrees. Computers and Graphics, 62(??):41–52, ???? 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301376.

Wu:2000:AAP

[WGLS00] Huayi Wu, Jianya Gong, Deren Li, and Wenzhong Shi. An algebraic algorithm for point inclusion query. Computers and Graphics, 24 (4):517–522, August 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/32/29/article.
pdf.

Wimmer:1999:FWI

[WGS99] Michael Wimmer, Markus Giegl, and Dieter Schmalstieg. Fast walkthroughs with image caches and ray casting. Computers and Graphics, 23(6):831– 838, December 1999. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/24/34/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/24/35/36/article.
pdf.

Wang:2018:SSSa

[WGS^{+}18a] Pengyu Wang, Yuan Gan, Panpan Shui, Fenggen Yu, Yan Zhang, Songle Chen, and Zhengxing Sun. 3D shape segmentation via shape fully convolutional networks. Computers and Graphics, 70(??):128–139, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301231.

See corrigendum [WGS^{+}18c].

Wang:2018:SSSb

[WGS^{+}18b] Pengyu Wang, Yuan Gan, Panpan Shui, Fenggen Yu, Yan Zhang, Songle Chen, and Zhengxing Sun. 3D shape segmentation via shape fully convolutional networks. Computers and Graphics, 76(??):182–192, November 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301158.

Wang:2018:CSS

[WGS^{+}18c] Pengyu Wang, Yuan Gan, Panpan Shui, Fenggen Yu, Yan Zhang, Songle Chen, and Zhengxing Sun. Corrigendum to “3D shape segmentation via shape fully convolutional networks” [Computers and Graphics 70 (2018) 128– 139]. Computers and Graphics, 76(??):181, November 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849318301146. See

[WGS^{+}18a].

Westermann:1996:NID

[WH96] Birgit Westermann and Rolf Hauser. Non-invasive 3-D patient registration for image-guided skull base surgery. Computers and Graphics, 20(6):793–799, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600049.

Wang:2016:BSP

[WHFL16] Chunxue Wang, Xin Hu, Xiaoming Fu, and Ligang Liu. Bijective spherical parametrization with low distortion. Computers and Graphics, 58(??):161– 171, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300486.

Wu:2006:CAP

[WHH06] Qingbiao Wu, Xinxian Huang, and Yujuan Han. A clipping algorithm for parabola segments against circular windows. Computers and Graphics, 30 (4):540–560, August 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000732.

Wan:2021:IAL

[WHH21] Yong Wan, Holly A. Holman, and Charles Hansen. Interactive analysis for large volume data from fluorescence microscopy at cellular precision. Computers and Graphics, 98(??):138– 149, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000935.

Whittington:2004:PEC

[Whi04] Jana Whittington. The process of effective critiques. Computers and Graphics, 28(3):401–407, June 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2009:GAD

[WHL^{+}09] Hongyu Wang, Ying He, Xin Li, Xianfeng Gu, and Hong Qin. Geometryaware domain decomposition for T-spline-based manifold modeling. Computers and Graphics, 33(3): 359–368, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000211.

Wang:2022:DDO

[WHW^{+}22] Hua Wang, Yifan He, Zheng Wang, Guanfeng Li, Yanqiu Xiao, and Mingliang Xu. Data-driven online traffic reconstructions: Interactively optimizing in virtual reality. Computers and Graphics, 105(??):85–93, June 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000528.

Wang:2018:MSG

[WHZ^{+}18] Xiaochao Wang, Jianping Hu, Dongbo Zhang, Lixin Guo, Hong Qin, and Aimin Hao. Multi-scale geometry detail recovery on surfaces via Empirical Mode Decomposition. Computers and Graphics, 70(??): 118–127, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301176.

Wong:2000:VBM

[WI00] Helena T. F. Wong and Horace H. S. Ip. Virtual brush: a model-based synthesis of Chinese calligraphy. Computers and Graphics, 24 (1):99–113, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/36/article.
pdf.

Willmert:1975:OMH

[Wil75] K. D. Willmert. Occupant model for human motion. Computers and Graphics, 1(1):123–128, May 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wild:1986:VAY

[Wil86] S. Wild. Voyaging to Australia: 30,000 years ago. Computers and Graphics, 10 (3):207–212, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wiltsche:2003:CSI

[Wil03]  Albert Wiltsche.  C^{1}-

and C^{2}-continuous splineinterpolation of a regular triangular net of points. Computers and Graphics, 27(6): 917–930, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wald:2008:FPA

[WIP08] Ingo Wald, Thiago Ize, and Steven G. Parker. Fast, parallel, and asynchronous construction of BVHs for ray tracing animated scenes. Computers and Graphics, 32(1):3–13, February 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307002014.

Wirth:1980:BDU

[Wir80] J. L. Wirth. Binary display using pseudo-random ensembles with selected distributions. Computers and Graphics, 5(2–4):99–101, 1980. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wisskirchen:1986:TOO

[Wis86]  Peter Wisskirchen.  To-

wards object-oriented graphics standards. Computers and Graphics, 10(2): 183–187, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wisskirchen:1987:CGW

[Wis87] P. Wisskirchen. Computer graphics within the AGF. Computers and Graphics, 11 (2):203–213, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wittek:1984:SMS

[Wit84] D. Wittek. Solid modeling and system design. Computers and Graphics, 8(4): 423–431, 1984. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:1991:AOS

[WJ91] Ronghang Wang and Wenhai Jiang. An algorithm of the offset shape. Computers and Graphics, 15(3): 435–439, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2009:DAK

[WJD^{+}09] Xiaoyu Wang, Dong Hyun Jeong, Wenwen Dou, Seok-Won Lee, William Ribarsky, and Remco Chang. Defining and applying knowledge conversion processes to a visual analytics system. Computers and Graphics, 33 (5):616–623, October 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000909.

Wagner:2019:DFV

[WJG^{+}19] Sebastian Wagner, Fabian Joeres, Mareike Gabele, Christian Hansen, Bernhard Preim, and Patrick Saalfeld. Difficulty factors for VR cognitive rehabilitation training

— crossing a virtual road. Computers and Graphics, 83 (??):11–22, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301037.

Weber:2015:MMR

[WJGG15] Yoann Weber, Vincent Jolivet, Guillaume Gilet, and Djamchid Ghazanfarpour. A multiscale model for rain rendering in real-time. Computers and Graphics, 50 (??):61–70, August 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000461.

Wang:2023:ICG

[WJW^{+}23] Jingqiang Wang, Shiyu Jia, Guodong Wang, Zhenkuan Pan, and Xiaokang Yu. An improved CPU–GPU parallel framework for real-time interactive cutting simulation of deformable objects. Computers and Graphics, 114(??):59–72, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000687.

Wallner:2014:PVA

[WK14] G. Wallner and S. Kriglstein.

PLATO: a visual analytics system for gameplay data. Computers and Graphics, 38 (??):341–356, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001891.

Weiler:2004:DVR

[WKE04] Manfred Weiler, Thomas Klein, and Thomas Ertl. Direct volume rendering in OpenSG. Computers and Graphics, 28(1):93–98, February 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Winnemoller:2012:XED

[WKO12] Holger Winnemöller, Jan Eric Kyprianidis, and Sven C. Olsen. XDoG: an eXtended difference-of-Gaussians compendium including advanced image stylization. Computers and Graphics, 36 (6):740–753, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200043X.

Weidenhausen:2003:LLW

[WKS03] Jens Weidenhausen, Christian Knoepfle, and Didier Stricker. Lessons learned on the way to industrial augmented reality applications, a retrospective on ARVIKA. Computers and Graphics, 27 (6):887–891, December 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2021:SBM

[WKT21] Jieying Wang, Jirí Kosinka, and Alexandru Telea. Splinebased medial axis transform representation of binary images. Computers and Graphics, 98(??):165– 176, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001114.

Wu:2016:SIM

[WKW16] Jun Wu, Lou Kramer, and Rüdiger Westermann. Shape interior modeling and mass property optimization using ray-reps. Computers and Graphics, 58(??):66– 72, August 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300462.

Wang:2015:CHV

[WL15] Jia Wang and Robert Lindeman. Coordinated hybrid virtual environments: Seamless interaction contexts for effective virtual reality. Computers and Graphics, 48(??):71–83, May 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000151.

Wang:2023:CCH

[WLB^{+}23] Xiaoyuan Wang, Yang Li, Adnane Boukhayma, Changbo Wang, and Marc Christie. Contact-conditioned handheld object reconstruction from single-view images. Computers and Graphics, 114(??):150–157, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300078X.

Walker:1988:HLD

[WLC88] Marshall Walker, Reng Song Lo, and S. F. Cheng. Hidden line detection in polytree representations. Computers and Graphics, 12 (1):65–69, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2011:RBM

[WLDB11] Kai Wang, Guillaume Lavoué, Florence Denis, and Atilla Baskurt. Robust and blind mesh watermarking based on volume moments. Computers and Graphics, 35(1): 1–19, February 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001433.

Wurmlin:2004:VFD

[WLG04] Stephan Würmlin, Edouard Lamboray, and Markus Gross. 3D video fragments: dynamic point samples for real-time freeviewpoint video. Computers and Graphics, 28(1):3– 14, February 2004. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weiss:2019:PBR

[WLJT19] Tomer Weiss, Alan Litteneker, Chenfanfu Jiang, and Demetri Terzopoulos. Position-based real-time simulation of large crowds. Computers and Graphics, 78 (??):12–22, February 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301699.

Wang:2012:AHF

[WLL^{+}12] Xiaochao Wang, Xiuping Liu, Linfa Lu, Baojun Li, Junjie Cao, Baocai Yin, and Xiquan Shi. Automatic hole-filling of CAD models with feature-preserving. Computers and Graphics, 36(2):101–110, April 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001749.

Wu:2021:DTC

[WLL^{+}21a] Huisi Wu, Yifan Li, Xueting Liu, Chengze Li, and Wenliang Wu. Deep texture cartoonization via unsupervised appearance regularization. Computers and Graphics, 97(??):99–107, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000571.

Wu:2021:TAS

[WLL21b] Jiang Wu, Chunxiao Liu, and Biao Li. Texture-aware and structure-preserving superpixel segmentation. Computers and Graphics, 94 (??):152–163, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301825.

Wang:2022:FPM

[WLL22] Chunxue Wang, Zheng Liu, and Ligang Liu. Featurepreserving Mumford–Shah mesh processing via nonsmooth nonconvex regularization. Computers and Graphics, 106(??):222–236, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200108X.

Wang:2024:ECS

[WLL^{+}24] Zhuang Wang, Zhen Liu, Tingting Liu, Yumeng Zhao, and Yanjie Chai. An emotional crowd simulation method based on audiovisual linkage for terrorist attacks. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000190.

Wang:2014:EBM

[WLP^{+}14] Pengjie Wang, Rynson W. H. Lau, Zhigeng Pan, Jiang Wang, and Haiyu Song. An Eigen-based motion retrieval method for real-time animation. Computers and Graphics, 38 (??):255–267, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001878.

Wei:2018:IBR

[WLQC18] Wei Wei, Hui-Yi Liu, Su-Bin Qian, and Shuang-Shuang Chen. Imagebased relighting using image segmentation and bootstrap strategy. Computers and Graphics, 77(??): 217–226, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301766.

Wang:2005:PPT

[WLW05] Wencheng Wang, Jing Li, and Enhua Wu. 2D pointin-polygon test by classifying edges into layers. Computers and Graphics, 29(3): 427–439, June 2005. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wu:2021:EPP

[WLX^{+}21] You Wu, Guiqing Li, Chuhua Xian, Xiaofeng Ding, and Yunhui Xiong. Extracting POP: Pairwise orthogonal planes from point cloud using RANSAC. Computers and Graphics, 94(??): 43–51, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301539.

Wei:2023:NRS

[WLX^{+}23] Housheng Wei, Yanli Liu, Guanyu Xing, Zhisheng Yan, and Yanci Zhang. Noreference shadow detection quality assessment via reference learning and multimode exploring. Computers and Graphics, 116 (??):13–23, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001541.

Wang:2023:BTB

[WLY23] Yong Wang, Bo Li, and Xin-

lin Yuan. BrightFormer:

a transformer to brighten the image. Computers and Graphics, 110(??):49– 57, February 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002242.

Wang:2019:DRG

[WLYH19] Zihao Wang, Hongwei Lin, Xiaofeng Yu, and Yusuf Fatihu Hamza. A dimensional reduction guiding deep learning architecture for 3D shape retrieval. Computers and Graphics, 81(??): 82–91, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300433.

Wang:2023:SST

[WLZ^{+}23] Yong Wang, Yangyang Liu, Pengbo Zhou, Guohua Geng, and Qi Zhang. SparseFormer: Sparse transformer network for point cloud classification. Computers and Graphics, 116 (??):24–32, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001693.

Wang:2018:OTU

[WLZL18] Fasheng Wang, Baowei Lin, Junxing Zhang, and Xucheng Li. Object tracking using Langevin Monte Carlo particle filter and locality sensitive histogram based likelihood model. Computers and Graphics, 70 (??):214–223, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301164.

Wilf:1989:TPA

[WM89] Itzhak Wilf and Yehuda Manor. Tracking parameterized algebraic curves on raster displays. Computers and Graphics, 13(3): 355–359, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walton:1990:CS

[WM90] D. J. Walton and D. S. Meek. Clothoidal splines. Computers and Graphics, 14 (1):95–100, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Walton:1995:PNI

[WM95] D. J. Walton and D. S. Meek. Point normal interpolation for stereolithography modelling. Computers and Graphics, 19(3): 345–353, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=3&aid=9500005.

Walton:2005:CCS

[WM05] D. J. Walton and D. S. Meek. A controlled clothoid spline. Computers and Graphics, 29(3):353–363, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wu:2024:VBP

[WM24] Hang Wu and Yubin Miao. Voting-based patch sequence autoregression network for adaptive point cloud completion. Computers and Graphics, 118(??):111–122, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323003060.

Wu:2008:RWM

[WMDR08] Xiaomao Wu, Lizhuang Ma, Zhuoqun Dong, and Lionel Revéret. Robust watermarking motion data with DL-STDM. Computers and Graphics, 32(3): 320–329, June 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000447.

Weber:1989:MGM

[WMFR89] Jacques Weber, Pierre-Yves Morgantini, Peter Fluekiger, and Michel Roch. Molecular graphics modeling of organometallic reactivity. Computers and Graphics, 13 (2):229–235, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weber:2015:DSU

[WMRA^{+}15] Daniel Weber, Johannes Mueller-Roemer, Christian Altenhofen, André Stork, and Dieter Fellner. Deformation simulation using cubic finite elements and efficient p-multigrid methods. Computers and Graphics, 53 (part B)(??):185– 195, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931500103X.

Wright:2013:UVV

[WMW13] H. Wright, C. Mathers, and J. P. R. B. Walton. Using visualization for visualization: an ecological interface design approach to inputting data. Computers and Graphics, 37(3):202–213, May 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000150.

Wang:2022:MAG

[WMZW22] Shaodong Wang, Shuai Ma, Hui Zhao, and Wencheng Wang. A multigrid approach for generating harmonic measured foliations. Computers and Graphics, 102 (??):380–389, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002144.

Wolfe:2000:BIC

[Wol00] Rosalee Wolfe. Bringing the introductory computer graphics course into the 21st century. Computers and Graphics, 24(1):151– 155, February 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/40/article.
pdf.

Wolfe:2002:ETV

[Wol02a] Rosalee Wolfe. Education:

Teaching visual aspects in an introductory computer graphics course. Computers and Graphics, 26(1):163– 168, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/42/abstract.
html.

Wollensak:2002:CMI

[Wol02b] Andrea Wollensak. Curricular modules: 3D and immersive visualization tools for learning. Computers and Graphics, 26(4):599–602, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/39/abstract.
html.

Woodward:1987:CSD

[Woo87] Charles D. Woodward. Crosssectional design of B-spline surfaces. Computers and Graphics, 11(2):193–201, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Woodwark:1989:CEO

[Woo89] J. R. Woodwark. Comments on “extended octtrees”. Computers and Graphics, 13(4):529, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Woodbury:1990:VSD

[Woo90] Robert F. Woodbury. Variations in solids. A declarative treatment. Computers and Graphics, 14(2): 173–188, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Willmert:1977:IHD

[WP77] K. D. Willmert and T. E. Potter. Improved human display model for occupant crash simulation programs. Computers and Graphics, 2 (2):51–54, 1977. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2023:EHD

[WPB^{+}23] Xiaofang Wang, Stéphanie Prévost, Adnane Boukhayma, Eric Desjardin, Céline Loscos, Benoit Morisset, and Franck Multon. Evaluation of hybrid deep learning and optimization method for 3D human pose and shape reconstruction in simulated depth images. Computers and Graphics, 115(??):158– 166, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001346.

Wang:2014:PAV

[WPH^{+}14] Rui Wang, Minghao Pan, Xiang Han, Weifeng Chen, and Hujun Bao. Parallel and adaptive visibility sampling for rendering dynamic scenes with spatially varying reflectance. Computers and Graphics, 38 (??):374–381, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001775.

Woo:2023:SDL

[WPJP23] Taeyun Woo, Wonjung Park, Woohyun Jeong, and Jinah Park. A survey of deep learning methods and datasets for hand pose estimation from hand-object interaction images. Computers and Graphics, 116(??): 474–490, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002364.

Wang:2023:RSB

[WPL^{+}23] Weijia Wang, Wei Pan, Xiao Liu, Kui Su, Bernard Rolfe, and Xuequan Lu. Random screening-based feature aggregation for point cloud denoising. Computers and Graphics, 116 (??):64–72, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001826.

Wang:2017:SFF

[WQL^{+}17] Weiming Wang, Sicheng Qian, Liping Lin, Baojun Li, Baocai Yin, Ligang Liu, and Xiuping Liu. Supportfree frame structures. Computers and Graphics, 66 (??):154–161, August 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300717.

Warman:1979:ETC

[WR79] E. A. Warman and F. J. Reader. Education and training for CAD — an industrial viewpoint. Computers and Graphics, 4(1):1–4, 1979. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wiens:2002:GET

[WR02] Andrea L. Wiens and Brian J. Ross. Gentropy: evolving 2D textures. Computers and Graphics, 26 (1):75–88, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/13/20/68/27/35/abstract.
html.

Wright:1975:PCG

[Wri75] T. J. Wright. Practical computer graphics for scientific users. philosophy and implementation. Computers and Graphics, 1(2–3): 157–160, September 1975. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2023:ABD

[WRLZ23] Yadong Wang, Teng Ran, Yuan Liang, and Guoquan Zheng. An attentionbased and deep sparse priori cascade multi-view stereo network for 3D reconstruction. Computers and Graphics, 116(??):383– 392, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001863.

Wan:2012:GBF

[WS12] Guowei Wan and Andrei Sharf. Grammar-based 3D facade segmentation and reconstruction. Computers and Graphics, 36(4): 216–223, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000192.

Wyvill:2022:CSY

[WS22] Brian Wyvill and Melina Skouras. A conversation with SMI 2021 Young Researcher Award winner Melina Skouras. Computers and Graphics, 105(??): A9–A11, June 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932200084X.

Wecker:2010:MAI

[WSG10] L. Wecker, F. Samavati, and M. Gavrilova. A multiresolution approach to iris synthesis. Computers and Graphics, 34(4):468– 478, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000798.

Wang:2022:VIS

[WSHY22] Fangjun Wang, Yanzhi Song, Zhangjin Huang, and Zhouwang Yang. Vectorized instance segmentation using periodic B-splines based on cascade architecture. Computers and Graphics, 102 (??):592–600, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001862.

Wilson:2024:TMU

[WSJJ24] Ethan Wilson, Frederick Shic, Sophie Jörg, and Eakta Jain. Towards mitigating uncann(eye)ness in face swaps via gaze-centric loss terms. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000153.

Wang:2022:IIM

[WSK^{+}22] Jieying Wang, Dennis J.

Silva, Jiří Kosinka, Alexandru Telea, Ronaldo F. Hashimoto, and Jos B. T. M. Roerdink. Interactive image manipulation using morphological trees and splinebased skeletons. Computers and Graphics, 108 (??):61–73, November 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001637.

Wong:1992:TPW

[WSL92] Isaac W. S. Wong, D. A.

Swayne, and D. C.-L. Lam. A tight package wrapping for planar point-sets. Computers and Graphics, 16(3): 265–271, Fall 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2019:SRU

[WSL^{+}19] Wen Wang, Tianyun Su, Haixing Liu, Xinfang Li, Zhen Jia, Lin Zhou, Zhuanling Song, and Ming Ding. Surface reconstruction from unoriented point clouds by a new triangle selection strategy. Computers and Graphics, 84(??): 144–159, November 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931930130X.

Wu:2014:ISC

[WSWL14] Zizhao Wu, Ruyang Shou, Yunhai Wang, and Xinguo Liu. Interactive shape co-segmentation via label propagation. Computers and Graphics, 38(??):248– 254, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300188X.

Wuhrer:2012:PIS

[WSX12] Stefanie Wuhrer, Chang Shu, and Pengcheng Xi. Posture-invariant statistical shape analysis using Laplace operator. Computers and Graphics, 36 (5):410–416, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000659.

Wither:2011:IAR

[WTA11] Jason Wither, Yun-Ta Tsai, and Ronald Azuma. Indirect augmented reality. Computers and Graphics, 35 (4):810–822, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001130.

Watanabe:1995:ETD

[WTF95] Takashi Watanabe, Akiko Tashiro, and Seizo Fujii. Estimation of threedimensional objects from orthographic views with inconsistencies. Computers and Graphics, 19(6):815–829, November–December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500052.

Wang:2011:EIP

[WTL^{+}11] Lu Wang, Tiow-Seng Tan, Chi-Wan Lim, Xiang-Xu Meng, Cheng Du, and Zhen fang Ji. Efficient implementation of point set reconstruction by multi-layer peeling algorithm. Computers and Graphics, 35(2): 208–217, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310001901.

Wang:2012:FRB

[WTM12] Kai Wang, Fakhri Torkhani, and Annick Montanvert. A fast roughness-based approach to the assessment of 3D mesh visual quality. Computers and Graphics, 36 (7):808–818, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001203.

Wang:2023:DOI

[WTW^{+}23] Bei Wang, Yan Tian, Jialei Wang, Jiayu Hu, Dongsheng Liu, and Zhaocheng Xu. Detect occluded items in X-ray baggage inspection. Computers and Graphics, 115 (??):148–157, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001425.

Wang:2018:ASC

[WTWT18] Tongtong Wang, Min Tang, Zhendong Wang, and Ruofeng Tong. Accurate self-collision detection using enhanced dual-cone method. Computers and Graphics, 73(??): 70–79, June 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300475.

Wu:1989:NCM

[Wu89] Shin-Ting Wu. A new combinatorial model for boundary representations. Computers and Graphics, 13(4): 477–486, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wu:2002:GEI

[Wu02] Shin-Ting Wu. Guest editor’s introduction: (computer graphics in Brazil). Computers and Graphics, 26 (6):837–839, December 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Weinmann:2015:DFA

[WUH^{+}15] M. Weinmann, S. Urban, S. Hinz, B. Jutzi, and C. Mallet. Distinctive 2D and 3D features for automated large-scale scene analysis in urban areas. Computers and Graphics, 49(??): 47–57, June 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000072.

Wuthrich:1998:MCR

[Wüt98] Charles A. Wüthrich. A model for curve rasterization in n-dimensional space. Computers and Graphics, 22(2–3):153–160, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/535.pdf.

Wu:2016:VIL

[WVY16] Shin-Ting Wu, Raphael Voltoline, and Clarissa Lin Yasuda. A view-independent line-coding colormap for diffusion tensor imaging. Computers and Graphics, 60 (??):66–75, November 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301005.

Wang:2006:SPS

[WW06] Chung-Ming Wang and Peng-Cheng Wang. Steganography on point-sampled geometry. Computers and Graphics, 30(2):244–254, April 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600032X.

Wang:2008:INS

[WW08] Xing-Yuan Wang and Shu-Guo Wang. An improved no-search fractal image coding method based on a modified gray-level transform.

Computers and Graphics, 32 (4):445–450, August 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000320.

Wimmer:2010:E

[WW10] Michael Wimmer and Peter Wonka. Editorial. Computers and Graphics, 34(4): 293, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000774.

Wu:2019:DCN

[WWCZ19] Wenyan Wu, Xingzhe Wu, Yici Cai, and Qiang Zhou. Deep coupling neural network for robust facial landmark detection. Computers and Graphics, 82(??): 286–294, August 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300937.

White:1995:GAD

[WWD^{+}95] Martin White, Marcus D.

Waller, Graham J. Dunnett, Paul F. Lister, and Richard L. Grimsdale. Graphics ASIC design using VHDL. Computers and Graphics, 19(2):301–308, March–April 1995. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=2&aid=9400156.

Wu:2018:FSD

[WWF^{+}18] Zhongke Wu, Xingce Wang, Yan Fu, Junchen Shen, Qianqian Jiang, Yuanshuai Zhu, and Mingquan Zhou. Fitting scattered data points with ball B-spline curves using particle swarm optimization. Computers and Graphics, 72(??):1–11, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300062.

Wu:2023:DGH

[WWF^{+}23] Xian Wu, Chen Wang, Hongbo Fu, Ariel Shamir,

                                        and Song-Hai Zhang. DeepPortraitDrawing:

Generating human body images from freehand sketches. Computers and Graphics, 116(??):73–81, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0097849323001723.

Wang:2021:AGL

[WWH^{+}21] Hua Wang, Yue Wu, Xu Han, Mingliang Xu, and Weizhe Chen. Automatic generation of large-scale 3D road networks based on GIS data. Computers and Graphics, 96 (??):71–81, May 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000261.

Wientapper:2011:CFM

[WWK11] Folker Wientapper, Harald Wuest, and Arjan Kuijper. Composing the feature map retrieval process for robust and ready-to-use monocular tracking. Computers and Graphics, 35 (4):778–788, August 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001117.

Wang:2012:ALD

[WWL^{+}12] Shandong Wang, Enhua Wu, Youquan Liu, Xuehui Liu, and Yanyun Chen. Abstract line drawings from photographs using flowbased filters. Computers and Graphics, 36(4): 224–231, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000349.

Wang:2016:BAP

[WWL16] Wentao Wang, Wenke Wang, and Sikun Li. Batch advection for the piecewise linear vector field on simplicial grids. Computers and Graphics, 54(??):75– 83, February 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001193.

Wang:2023:IVA

[WWO^{+}23] Zhongwei Wang, Yating Wei, GongChang Ou, Han Gao, Haitao Yang, Yue Wang, Chen Cao, Minfeng Zhu, and Wei Chen. Interactive visual analytics of parallel training strategies for DNN models. Computers and Graphics, 115(??):392– 403, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001607.

Wu:2013:UCS

[WWS^{+}13] Zizhao Wu, Yunhai Wang, Ruyang Shou, Baoquan Chen, and Xinguo Liu. Unsupervised co-segmentation of 3D shapes via affinity aggregation spectral clustering. Computers and Graphics, 37 (6):628–637, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313000897.

Wu:2022:AEL

[WWWW22] Xian-Tao Wu, Yi Wang, Yi Wan, and Wen Wu. Annotation is easy: Learning to generate a shadow mask. Computers and Graphics, 104(??):152–161, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000553.

Wang:2006:FEB

[WWY06] Yu Wang, Charlie C. L.

Wang, and Matthew M. F. Yuen. Fast energy-based surface wrinkle modeling. Computers and Graphics, 30 (1):111–125, February 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849305002153.

Wu:2014:ICU

[WX14] Hao Wu and Dan Xu. Image compositing using dominant patch transformations. Computers and Graphics, 38 (??):277–285, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931300157X.

Wang:2023:TBR

[WXC^{+}23] Dong Wang, Wenjun Xie, Youcheng Cai, Xinjie Li, and Xiaoping Liu. Transformerbased rapid human pose estimation network. Computers and Graphics, 116(??): 317–326, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002248.

Wang:2014:CDI

[WXG^{+}14] Linbo Wang, Tianchen Xia, Yanwen Guo, Ligang Liu, and Jue Wang. Confidencedriven image co-matting. Computers and Graphics, 38 (??):131–139, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001556.

Wang:2023:SHM

[WXL^{+}23] Tiancheng Wang, Yanrui Xu, Ruolan Li, Haoping Wang, Yuege Xiong, and Xiaokun Wang. Simulating hyperelastic materials with anisotropic stiffness models in a particlebased framework. Computers and Graphics, 116(??): 437–447, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002303.

Wang:2018:STU

[WXZ^{+}18] Hua Wang, Mingliang Xu, Fubao Zhu, Zhigang Deng, Yafei Li, and Bing Zhou. Shadow traffic: a unified model for abnormal traffic behavior simulation. Computers and Graphics, 70 (??):235–241, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300973.

Wang:2003:FES

[WY03] Charlie C. L. Wang and Matthew M. F. Yuen. Freeform extrusion by sketched input. Computers and Graphics, 27(2):255–263, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Wang:2011:QCS

[WY11a] Jun Wang and Zeyun Yu.

Quadratic curve and surface fitting via squared distance minimization. Computers and Graphics, 35(6): 1035–1050, December 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001464.

Wang:2011:SFB

[WY11b] Jun Wang and Zeyun Yu.

Surface feature based mesh segmentation. Computers and Graphics, 35(3): 661–667, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000550.

Wu:2023:PCT

[WYC^{+}23] Yushuang Wu, Zizheng Yan, Shengcai Cai, Guanbin Li, Xiaoguang Han, and Shuguang Cui. PointMatch: a consistency training framework for weakly supervised semantic segmentation of 3D point clouds. Computers and Graphics, 116(??): 427–436, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323002297.

Wang:2022:RIR

[WYXM22] Huidi Wang, Lele Yuan, Jing-Jing Xiong, and Jun Mao. The relaxed implicit randomized algebraic reconstruction technique for curve and surface reconstruction. Computers and Graphics, 102(??):9–17, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002454.

Wan:2011:TPO

[WYZ^{+}11] Shenghua Wan, Zhao Yin, Kang Zhang, Hongchao Zhang, and Xin Li. A topology-preserving optimization algorithm for polycube mapping. Computers and Graphics, 35(3): 639–649, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000574.

Wu:2020:PSF

[WYZ20] Yan Wu, Jun Yang, and Jinlong Zhao. Partial 3D shape functional correspondence via fully spectral eigenvalue alignment and upsampling refinement. Computers and Graphics, 92 (??):99–113, November 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301382.

Wang:2009:CAM

[WZ09] Kun-Peng Wang and Cai-Ming Zhang. Content-aware model resizing based on surface deformation. Computers and Graphics, 33(3): 433–438, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000260.

Wang:2021:RTO

[WZ21] Xin Wang and Risong Zhang. Rendering transparent objects with caustics using real-time ray tracing. Computers and Graphics, 96(??):36–47, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784932100039X.

Wang:2021:FCN

[WZC^{+}21] Xinyi Wang, Zhenyang Zhu, Xiaodiao Chen, Kentaro Go, Masahiro Toyoura, and Xiaoyang Mao. Fast contrast and naturalness preserving image recolouring for dichromats. Computers and Graphics, 98(??): 19–28, August 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000698.

Wu:2024:TCB

[WZC^{+}24] Yichun Wu, Huihuang Zhao, Wenhui Chen, Yunfei Yang, and Jiayi Bu. TextStyler: a CLIP-based approach to text-guided style transfer. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000141.

Wang:2022:PBD

[WZL^{+}22] Xianyu Wang, Ke Zhang, Haoyu Li, Hua Luo, and Jingyu Wang. ProbNet: Bayesian deep neural network for point cloud analysis. Computers and Graphics, 104(??):106–115, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849322000565.

Wang:2019:HML

[WZLQ19] Changbo Wang, Shenfan Zhang, Chen Li, and Hong Qin. Hybrid modeling of Lagrangian–Eulerian method for high-speed fluid simulation. Computers and Graphics, 78(??):87– 96, February 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849318301936.

Wang:2019:PSI

[WZS19] Kai Wang, Jianmin Zheng, and Hock Soon Seah. Progressive sketching with instant previewing. Computers and Graphics, 81(??): 9–19, June 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300408.

Wang:1997:CTB

[WZT97] Lazhu Wang, Xinxiong Zhu, and Zesheng Tang. Coons type blended Bspline (CNSBS) surface and its conversion to NURBS surface. Computers and Graphics, 21(3):297–303, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700006.

Wang:1997:ATV

[WZW97] Wen-Cheng Wang, Ding-Hong Zhou, and En-Hua Wu. Accelerating techniques in volume rendering of irregular data. Computers and Graphics, 21(3): 289–295, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700005.

Wu:2018:JAS

[WZZ^{+}18] Zizhao Wu, Yunhui Zhang, Ming Zeng, Feiwei Qin, and Yigang Wang. Joint analysis of shapes and images via deep domain adaptation. Computers and Graphics, 70 (??):140–147, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301061.

Wu:2021:SRD

[WZZ^{+}21] Wen Wu, Shuping Zhang, Kai Zhou, Jie Yang, Xiantao Wu, and Yi Wan. Shadow removal via dual module network and low error shadow dataset. Computers and Graphics, 95 (??):156–163, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000273.

Wang:2018:DSE

[WZZZ18] Xiaotian Wang, Bin Zhou, Yu Zhang, and Yifan Zhao. Deep style estimator for 3D indoor object collection organization and scene synthesis. Computers and Graphics, 74(??): 76–84, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300670.

Xia:2019:ESC

[XCL^{+}19] Qing Xia, Chengju Chen, Jiarui Liu, Shuai Li, Aimin Hao, and Hong Qin. Efficient 4D shape completion from sparse samples via cubic spline fitting in linear rotation-invariant space. Computers and Graphics, 82 (??):129–139, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300755.

Xu:2009:QET

[XCW^{+}09] Kai Xu, Zhi-Quan Cheng, Yanzhen Wang, Yueshan Xiong, and Hao Zhang. Quality encoding for tetrahedral mesh optimization. Computers and Graphics, 33(3):250–261, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000375.

Xiao:2023:MPD

[XCXC23] Yanyang Xiao, Juan Cao, Shaoping Xu, and Zhonggui Chen. Meshless power diagrams. Computers and Graphics, 114(??):247–256, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001073.

Xiao:2016:BTS

[XCZ^{+}16] Zhoufang Xiao, Jianjun Chen, Yao Zheng, Jianjing Zheng, and Desheng Wang. Booleans of triangulated solids by a boundary conforming tetrahedral mesh generation approach. Computers and Graphics, 59 (??):13–27, October 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300358.

Xin:2014:MLG

[XCZP14] Shi-Qing Xin, Shuangmin Chen, Jieyu Zhao, and Zhigeng Pan. Measuring length and girth of a tubular shape by quasi-helixes. Computers and Graphics, 38(??): 392–398, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001787.

Xi:2008:NRG

[XD08] Yongjian Xi and Ye Duan. A novel region-growing based iso-surface extraction algorithm. Computers and Graphics, 32(6):647–654, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308001167.

Xu:2018:HMQ

[XGC18] Kaoji Xu, Xifeng Gao, and Guoning Chen. Hexahedral mesh quality improvement via edge-angle optimization. Computers and Graphics, 70 (??):17–27, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931730095X.

Xian:2011:AAD

[XGZ11] Chuhua Xian, Shuming Gao, and Tianming Zhang. An approach to automated decomposition of volumetric mesh. Computers and Graphics, 35(3):461–470, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000811.

Xue:1988:NDS

[XH88] Jing-Ling Xue and Xian-Long Hong. A new data structure for representing cell hierarchy in layout design. Computers and Graphics, 12(3–4):341–348, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Xu:2012:BNS

[XHGL12] Yin Xu, Ruizhen Hu, Craig Gotsman, and Ligang Liu. Blue noise sampling of surfaces. Computers and Graphics, 36(4): 232–240, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000222.

Xi:2016:LSU

[XhKKL16] Zhonghua Xi, Yun hyeong Kim, Young J. Kim, and Jyh-Ming Lien. Learning to segment and unfold polyhedral mesh from failures.

Computers and Graphics, 58 (??):139–149, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300656.

Xia:2024:EPC

[XHL^{+}24] Zi-Xun Xia, Jian-Yu Hao, Kang Li, Ao-Xiang Tian, and Zheng-Jun Du. Edit propagation via color palettes. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000025.

Xiong:2010:SBA

[XL10] Yiyan Xiong and Joseph J.

LaViola, Jr. A ShortStrawbased algorithm for corner finding in sketch-based interfaces. Computers and Graphics, 34(5):513– 527, October 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310001044.

Xu:2011:CCD

[XLGG11] Yin Xu, Ligang Liu, Craig Gotsman, and Steven J. Gortler. Capacity-constrained Delaunay triangulation for point distributions. Computers and Graphics, 35(3): 510–516, June 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000707.

Xu:2016:UIS

[XLGS16] Qiuying Xu, Songrun Liu, Yotam Gingold, and Karan Singh. Using isophotes and shadows to interactively model normal and height fields. Computers and Graphics, 59(??):130– 142, October 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300139.

Xu:2021:FCL

[XLHH21] Chenkai Xu, Hongwei Lin, Hui Hu, and Yaqi He. Fast calculation of Laplace– Beltrami eigenproblems via subdivision linear subspace. Computers and Graphics, 97(??):236–247, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000613.

Xu:2018:MHB

[XLL^{+}18] Haocan Xu, Jituo Li, Guodong Lu, Haijin Deng, Dongliang Zhang, and Juntao Ye. Modeling 3D human body with a smart vest. Computers and Graphics, 75 (??):36–43, October 2018.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318301043.

Xu:2020:PRM

[XLL^{+}20] Haocan Xu, Jituo Li, Guodong Lu, Dongliang Zhang, and Juncai Long. Predicting ready-made garment dressing fit for individuals based on highly reliable examples. Computers and Graphics, 90 (??):135–144, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300911.

Xiong:2012:CAB

[XLM12] Yunhui Xiong, Guiqing Li, and Aihua Mao. Convergence analysis for Bspline geometric interpolation. Computers and Graphics, 36(7):884–891, November 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200129X.

Xing:2012:PAR

[XLQP12] Guanyu Xing, Yanli Liu, Xueying Qin, and Qunsheng Peng. A practical approach for real-time illumination estimation of outdoor videos. Computers and Graphics, 36 (7):857–865, November 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312001343.

Xiao:2022:LMI

[XLSW22] Dong Xiao, Siyou Lin, Zuoqiang Shi, and Bin Wang. Learning modified indicator functions for surface reconstruction. Computers and Graphics, 102(??):309– 319, February 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321002272.

Xing:2024:SEF

[XLW^{+}24] Jialu Xing, Jianping Liu, Jian Wang, Lulu Sun, Xi Chen, Xunxun Gu, and Yingfei Wang. A survey of efficient fine-tuning methods for vision-language models — prompt and adapter. Computers and Graphics, 119(??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000128.

Xiao:2011:CMM

[XLXG11] Dong Xiao, Hongwei Lin, Chuhua Xian, and Shuming Gao. CAD mesh model segmentation by clustering. Computers and Graphics, 35(3):685–691, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311000598.

Xu:2012:PMI

[XM12] Ling Xu and David Mould.

A procedural method for irregular tree models. Computers and Graphics, 36(8): 1036–1047, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001410.

Xu:2012:RTI

[XMD^{+}12] Shibiao Xu, Xing Mei, Weiming Dong, Zhiyi Zhang, and Xiaopeng Zhang. Realtime ink simulation using a grid-particle method. Computers and Graphics, 36(8): 1025–1035, December 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312001392.

Xu:1990:ARM

[XPL90] Hau Xu, Qun-Sheng Peng, and You-Dong Liang. Accelerated radiosity method for complex environments. Computers and Graphics, 14 (1):65–71, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Xu:1997:PBS

[XSQ^{+}97] Yingqing Xu, Cheng Su, Dongxu Qi, Hua Li, and Shenquan Liu. Physically based simulation of water currents and waves. Computers and Graphics, 21(3): 277–280, May–June 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=3&aid=9700003.

Xiao:2023:ADR

[XSW23] Dong Xiao, Zuoqiang Shi, and Bin Wang. Alternately denoising and reconstructing unoriented point sets. Computers and Graphics, 116(??):139–149, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001784.

Xu:2008:PCI

[Xu08] Lin Xu. Parametric curve interpolation by combination of two conic sections. Computers and Graphics, 32 (6):655–659, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001234.

Xu:2020:ATF

[XWF^{+}20] Zheng Xu, Michael Wilber, Chen Fang, Aaron Hertzmann, and Hailin Jin. Adversarial training for fast arbitrary style transfer. Computers and Graphics, 87(??): 1–11, April 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300030.

Xu:2018:VPB

[XWW^{+}18] Xiang Xu, Beibei Wang, Lu Wang, Yanning Xu, and Tamy Boubekeur. Vectorized point based global illumination on Intel MIC architecture. Computers and Graphics, 70(??):206– 213, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301048.

Xu:2019:TDB

[XWW^{+}19] Xiang Xu, Beibei Wang, Lu Wang, Yanning Xu, Chenglei Yang, and Xiangxu Meng. A task and data balanced distributed photon mapping method. Computers and Graphics, 82 (??):214–221, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300640.

Xiao:2021:PBD

[XWW^{+}21] Qinjie Xiao, You Wu, Dinghong Wang, Yong-Liang Yang, and Xiaogang Jin.

                                        Beauty3DFaceNet:  Deep

geometry and texture fusion for 3D facial attractiveness prediction. Computers and Graphics, 98 (??):11–18, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000650.

Xu:2021:PPP

[XWWK21] Hongbin Xu, Lvequan Wang, Qiuxia Wu, and Wenxiong Kang. PVLNet: Parameterizedview-learning neural network for 3D shape recognition. Computers and Graphics, 98(??):71–81, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000789.

Xu:2015:WGS

[XWY15] Wenzhuo Xu, Bin Wang, and Dong-Ming Yan. Wall grid structure for interior scene synthesis. Computers and Graphics, 46(??): 231–243, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001204.

Xia:1994:PPG

[XXC94] Weixin Xia, Guorong Xuan, and Peiqi Chai. Periodic pattern graphics. Computers and Graphics, 18(1):81– 86, January–February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Xu:2021:RRN

[XXHM21] Ziqi Xu, Chao Xu, Jing Hu, and Zhaopeng Meng. Robust resistance to noise and outliers: Screened Poisson surface reconstruction using adaptive kernel density estimation. Computers and Graphics, 97(??): 19–27, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000479.

Xu:2018:TGO

[XXT18] Minfeng Xu, Shiqing Xin, and Changhe Tu. Towards globally optimal normal orientations for thin surfaces. Computers and Graphics, 75 (??):25–35, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S009784931830092X.

Xu:2023:ASS

[XXX^{+}23] Yanrui Xu, Yuanmu Xu, Yuege Xiong, Dou Yin, Xiaojuan Ban, Xiaokun Wang, Jian Chang, and Jian Jun Zhang. Anisotropic screen space rendering for particlebased fluid simulation. Computers and Graphics, 110 (??):118–124, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002308.

Xiao:2000:FBS

[XZ00] Yingcai Xiao and John P. Ziebarth. FEM-based scattered data modeling and visualization. Computers and Graphics, 24(5):775– 789, October 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/36/article.
pdf.

Xu:2009:DHF

[XZCOX09] Kai Xu, Hao Zhang, Daniel Cohen-Or, and Yueshan Xiong. Dynamic harmonic fields for surface processing. Computers and Graphics, 33(3):391–398, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000442.

Xiong:1988:IGK

[XZL88] Yihua Xiong, Hongwen Zhang, and Shenquan Liu. An implementation of the graphical kernel system (GKS). Computers and Graphics, 12(3–4):329–333, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Xu:2022:MSD

[XZL^{+}22] YuJie Xu, YongJun Zhang, Zhi Li, ZhongWei Cui, and YiTong Yang. Multiscale dehazing network via high-frequency feature fusion. Computers and Graphics, 107(??):50–59, October 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001182.

Xiang:2021:EFT

[XZPG21] Jun Xiang, Ning Zhang, Ruru Pan, and Weidong Gao. Efficient fine-texture image retrieval using deep multi-view hashing. Computers and Graphics, 101 (??):93–105, December 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002156.

Xing:2021:STE

[XZY^{+}21] Yu Xing, Yu Zhou, Xin Yan, Haisen Zhao, Wenqiang Liu, Jingbo Jiang, and Lin Lu. Shell thickening for extrusion-based ceramics printing. Computers and Graphics, 97(??): 160–169, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932100073X.

Xu:2023:NSD

[XZY^{+}23] Zhe Xu, Buwei Zhou, Zhou Yang, Xiaohui Yuan, Yankong Zhang, and Qiang Lu. NeatSankey: Sankey diagrams with improved readability based on node positioning and edge bundling. Computers and Graphics, 113(??):10–20, June 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000481.

Yildiz:2023:MAC

[YAKE23] Tolga Yildiz, Ergun Akleman, Vinayak Krishnamurthy, and Matthew Ebert. A modular approach for creation of any bi-axial woven structure with congruent tiles. Computers and Graphics, 114(??):357– 367, August 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001127.

Yamazaki:1994:ILM

[Yam94] Shigeru Yamazaki. ISDB: a layered model for integrated services digital broadcasting. Computers and Graphics, 18(1):7–10, January– February 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yao:2024:DDE

[YBD^{+}24] Zhiping Yao, Jiang Bi, Wei Deng, Wenlin He, Zihan Wang, Xu Kuang, Mi Zhou, Qinquan Gao, and Tong Tong. DEUNet: Dual-encoder UNet for simultaneous denoising and reconstruction of single HDR image. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000098.

Yu:2023:SMT

[YBTB23] Yuncong Yu, Tim Becker, Le Minh Trinh, and Michael Behrisch. SAXRegEx: Multivariate time series pattern search with symbolic representation, regular expression, and query expansion. Computers and Graphics, 112(??):13–21, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000316.

Yang:2010:TCD

[YC10] Sang Wook Yang and Young Choi. Triangulation of CAD data for visualization using a compact array-based triangle data structure. Computers and Graphics, 34 (4):424–429, August 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000233.

Yang:2018:CIS

[YCF18] Yang Yang, Shuangming Chai, and Xiao-Ming Fu. Computing interior supportfree structure via hollowto-fill construction. Computers and Graphics, 70 (??):148–156, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300985.

Yang:2021:FLC

[YCL^{+}21] Qinqi Yang, Zhuxin Chen, Yanli Liu, Guanyu Xing, and Yanci Zhang. Foveated light culling. Computers and Graphics, 97(??): 200–207, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000637.

Yamac:2023:LIG

[YCO23] Goksu Yamac, Jackey J. K.

Chang, and Carol O’Sullivan. Let it go! Point of release prediction for virtual throwing. Computers and Graphics, 110(??):11– 18, February 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001996.

You:2004:PBS

[YCZ04] L. H. You, P. Comninos, and Jian J. Zhang. PDE blending surfaces with C^{2} continuity. Computers and Graphics, 28 (6):895–906, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yuan:2000:HLR

[YD00] Xiaobu Yuan and Xiaomin Dong. Hidden-line rendering with a dynamic P-buffer. Computers and Graphics, 24(3):443–451, June 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/30/39/article.
pdf.

Ye:2002:ACO

[Ye02] Ruisong Ye. Another choice for orbit traps to generate artistic fractal images. Computers and Graphics, 26(4):629–633, August ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/55/44/abstract.
html.

Yang:2009:SMA

[YF09] Wenwu Yang and Jieqing Feng. 2D shape morphing via automatic feature matching and hierarchical interpolation. Computers and Graphics, 33(3): 414–423, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309000387.

Yu:2012:PDF

[YGS12] Hui Yu, Oliver G. B. Garrod, and Philippe G. Schyns. Perception-driven facial expression synthesis. Computers and Graphics, 36(3): 152–162, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001695.

Yuan:2021:ISB

[YH21] Qi Yuan and Yongjian Huai. Immersive sketch-based tree modeling in virtual reality. Computers and Graphics, 94 (??):132–143, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301813.

Young:1993:ICP

[YHHS93] Bong Kim Young, Sub Kim Hoi, Oh Kim Hong, and Yong Shin Sung. Infinitecorner-point fractal image generation by Newton’s method for solving exp(−αζ + Z/ζ − Z) − 1 =

0. Computers and Graphics, 17(6):705–711, November– December 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

You:2022:RTV

[YHNC22] Jiawei You, Yongjian Huai, Xiaoying Nie, and Yuanyuan Chen. Real-time 3D visualization of forest fire spread based on tree morphology and finite state machine. Computers and Graphics, 103(??):109–120, April 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000097.

Yao:2023:GDL

[YHW23] Siyuan Yao, Jun Han, and Chaoli Wang. GMT: a deep learning approach to generalized multivariate translation for scientific data analysis and visualization. Computers and Graphics, 112(??):92–104, May 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000390.

Yu:2010:RSF

[YHX10] Wei-Wei Yu, Fei He, and Ping Xi. A rapid 3D seed-filling algorithm based on scan slice. Computers and Graphics, 34(4):449– 459, August 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000725.

Yoon:2009:VBN

[YIL09] Mincheol Yoon, Ioannis Ivrissimtzis, and Seungyong Lee. Variational Bayesian noise estimation of point sets. Computers and Graphics, 33(3): 226–234, June 2009. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000363.

Yin:2004:REN

[Yin04] Zhongwei Yin. Reverse engineering of a NURBS surface from digitized points subject to boundary conditions. Computers and Graphics, 28(2):207–212, April 2004.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yu:1999:WNW

[YJC99] Young-Jung Yu, Ho-Youl Jung, and Hwan-Gue Cho. WSCG ’98 — a new water droplet model using metaball in the gravitational field. Computers and Graphics, 23(2):213– 222, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/674.pdf.

Yu:2021:CIT

[YJLZ21] Ying-Ying Yu, Ye Ji, Jing-Gai Li, and Chun-Gang Zhu. Conditions for injectivity of toric volumes with arbitrary positive weights. Computers and Graphics, 97(??):88–98, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000686.

Young:2018:GAG

[YK18] Gavin Young and Adarsh Krishnamurthy. GPUaccelerated generation and rendering of multi-level voxel representations of solid models. Computers and Graphics, 75(??): 1–10, October 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931830102X.

Yilmaz:2023:IGI

[YK23] Doga Yilmaz and Furkan Kiraç. Illumination-guided inverse rendering benchmark: Learning real objects with few cameras. Computers and Graphics, 115 (??):107–121, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001310.

Yang:1996:RCC

[YL96] Ching-Yung Yang and Ja-Chen Lin. RWM-cut for color image quantization. Computers and Graphics, 20(4):577–588, July– August 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=9600028.

Yan:2023:RTD

[YL23] Jiacong Yan and Hongwei Lin. Reasonable thickness determination for implicit porous sheet structure using persistent homology. Computers and Graphics, 115 (??):236–245, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001528.

Yuan:2018:HFG

[YLH^{+}18] Jingjing Yuan, Xinglong Liu, Fei Hou, Hong Qin, and Aimin Hao. Hybrid-feature-guided lung nodule type classification on CT images. Computers and Graphics, 70 (??):288–299, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301139.

Yang:2021:AIL

[YLS^{+}21] Bailin Yang, Liuliu Li, Chao Song, Zhaoyi Jiang, and Yun Ling. Automatic interior layout with userspecified furniture. Computers and Graphics, 94 (??):124–131, February 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320301801.

Ying:2018:APS

[YLT^{+}18] Jianming Ying, Lin Lu, Lihao Tian, Xin Yan, and Baoquan Chen. Anisotropic porous structure modeling for 3D printed objects. Computers and Graphics, 70 (??):157–164, February 2018.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301012.

Yang:2013:USC

[YLYJ13] Ben Yang, Youquan Liu, Lihua You, and Xiaogang Jin. A unified smoke control method based on signed distance field. Computers and Graphics, 37(7): 775–786, November 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000757.

Yong:2012:IPM

[YMYH12] Liu Yong, Zhang Mingmin, Jiang Yunliang, and Zhao Haiying. Improving procedural modeling with semantics in digital architectural heritage. Computers and Graphics, 36(3): 178–184, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000040.

Yilmaz:2011:RMS

[YMYI11] Erdal Yilmaz, Eray Molla, Cansin Yildiz, and Veysi Isler. Realistic modeling of spectator behavior for soccer videogames with CUDA. Computers and Graphics, 35(6):1063– 1069, December 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849311001476.

Yoo:2015:MRU

[YMZ^{+}15] Innfarn Yoo, Michel Abdul Massih, Illia Ziamtsov, Raymond Hassan, and Bedrich Benes. Motion retiming by using bilateral time control surfaces. Computers and Graphics, 47(??): 59–67, April 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001393.

Yoshimura:1994:DMI

[YNS94] Tetsuya Yoshimura, Yasuhiro Nakamura, and Masataka Sugiura. 3D direct manipulation interface: Development of the Zashiki-Warashi system. Computers and Graphics, 18(2): 201–207, March–April 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Youngblood:1989:AIC

[You89] Michael S. Youngblood.

AT&T image capture board (ICB) and TIPS software. Computers and Graphics, 13 (1):101–??, 1989. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yu:2005:RSC

[YP05] Jinhui Yu and Qunsheng Peng. Realistic synthesis of cao shu of Chinese calligraphy. Computers and Graphics, 29(1):145–153, February 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yang:2019:PPQ

[YPLL19] Jingru Yang, Hao Peng, Lin Liu, and Lin Lu. 3D printed perforated QR codes. Computers and Graphics, 81(??):117–124, June 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300524.

Yin:2001:RMW

[YPSZ01] Kangkang Yin, Zhigeng Pan, Jiaoying Shi, and David Zhang. Robust mesh watermarking based on multiresolution processing. Computers and Graphics, 25(3): 409–420, June 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/32/article.
                                        pdf.

Yang:2023:GIW

[YPZ^{+}23] Qiuxia Yang, Yuanyuan Pu, Zhengpeng Zhao, Dan Xu, and Siqi Li. W2GAN: Importance weight and wavelet feature guided image-to-image translation under limited data. Computers and Graphics, 116(??): 115–127, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300184X.

Yining:1990:CGB

[YQY90] Zhu Yining, Peng Qunsheng, and Liang Youdong. Computers and Graphics Best Paper Award (1988–1989). Computers and Graphics, 14 (1):1–??, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yao:1996:BDI

[YR96] Chengfu Yao and Jon G.

Rokne. Bi-directional incremental linear interpolation. Computers and Graphics, 20(2):295–305, March–April 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500131.

Yao:1998:RLS

[YR98] Chengfu Yao and Jon G. Rokne. Run-length slice algorithms for the scanconversion of ellipses. Computers and Graphics, 22(4): 463–477, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/4/579.pdf.

Yoo:2024:CER

[YRD^{+}24] Soojeong Yoo, João Ramalhinho, Thomas Dowrick, Murali Somasundaram, Kurinchi Gurusamy, Brian Davidson, Matthew J. Clarkson, and Ann Blandford. Can engineers represent surgeons in usability studies? Comparison of results from evaluating augmented reality guidance for laparoscopic surgery. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849324000086.

Yadav:2018:CBP

[YRS^{+}18] Sunil Kumar Yadav, Ulrich Reitebuch, Martin Skrodzki, Eric Zimmermann, and Konrad Polthier. Constraintbased point set denoising using normal voting tensor and restricted quadratic error metrics. Computers and Graphics, 74(??):234– 243, August 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300797.

Yuan:1997:BHL

[YS97] Xiaobu Yuan and Hanqiu Sun. P -buffer: a hidden-line algorithm in image-space. Computers and Graphics, 21(3):359– 366, May–June 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1997&volume=21&
issue=3&aid=9700013.

Yanik:2015:ALS

[YS15] Erelcan Yanik and Tevfik Metin Sezgin. Active learning for sketch recognition. Computers and Graphics, 52(??):93– 105, November 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001260.

Yesilbek:2017:SRF

[YS17] Kemal Tugrul Yesilbek and T. Metin Sezgin. Sketch recognition with few examples. Computers and Graphics, 69(??):80–91, December 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301516.

See corrigendum [YS21].

Yesilbek:2021:CSR

[YS21] Kemal Tugrul Yesilbek and T. Metin Sezgin. Corrigendum to “Sketch recognition with few examples” [computers and graphics 69 (2017) 80–91]. Computers and Graphics, 94(??): 191, February 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320301758.

See [YS17].

Yamamoto:2023:TCM

[YS23] Takumi Yamamoto and Yuta Sugiura. Turning carpets into multi-image switchable displays. Computers and Graphics, 111(??): 190–198, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000286.

Yue:2013:NUI

[YSD13] Tao Yue, Jinli Suo, and Qionghai Dai. Non-uniform image deblurring using an optical computing system. Computers and Graphics, 37(8):1039–1050, December 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001520.

Yurtoglu:2024:PPH

[YSDG24] Ayda Yurtoglu, Sinan Sonlu, Yalim Dogan, and Ugur Güdükbay. Personality perception in human videos altered by motion transfer networks. Computers and Graphics, 119(??): ??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932400013X.

Yagel:1996:BVE

[YSW^{+}96] Roni Yagel, Don Stredney, Gregory J. Wiet, Petra Schmalbrock, Louis Rosenberg, Dennis J. Sessanna, and Yair Kurzion. Building a virtual environment for endoscopic sinus surgery simulation. Computers and Graphics, 20(6):813–823, November–December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=6&aid=9600051.

Yong:2009:P

[YSWP09] Jun-Hai Yong, Michela Spagnuolo, Wenping Wang, and Program Chairs of IEEE International Conference on Shape Modeling and Applications (SMI). Preface. Computers and Graphics, 33(3):193–194, June 2009. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849309000570.

Ye:2022:HFR

[YSZ22] Yuping Ye, Zhan Song, and Juan Zhao. High-fidelity 3D real-time facial animation using infrared structured light sensing system. Computers and Graphics, 104(??):46–58, May 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000395.

Yamaguchi:1983:SMS

[YT83] F. Yamaguchi and T. Tokieda. A solid modelling system: Freedom II. Computers and Graphics, 7(3– 4):225–232, 1983. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yuen:1987:ESP

[YT87] S. Yuen and J. D. Taylor. Enhancements of small peaks using the Moiré technique. Computers and Graphics, 11(1):33–34, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yu:1999:TSE

[Yu99] Yizhou Yu. Technical section — efficient visibility processing for projective texture mapping. Computers and Graphics, 23(2):245– 253, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/677.pdf.

Yuen:1986:DIC

[Yue86] W. Y. D. Yuen. Design and implementation of a CAD system for cold roll forming. Computers and Graphics, 10(3):213–217, 1986. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yang:2022:PSI

[YWC22] Yang Yang, Yu Wu, and Yi Cao. Probabilistic summarization via importancedriven sampling for largescale patch-based scientific data visualization. Computers and Graphics, 106 (??):119–129, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001066.

Yuan:2016:STM

[YWH^{+}16] Yazhen Yuan, Rui Wang, Jin Huang, Yanming Jia, and Hujun Bao. Simplified and tessellated mesh for realtime high quality rendering. Computers and Graphics, 54 (??):135–144, February 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001144.

Yang:2003:IHD

[YWR03] Jing Yang, Matthew O. Ward, and Elke A. Rundensteiner. Interactive hierarchical displays: a general framework for visualization and exploration of large multivariate data sets. Computers and Graphics, 27(2):265–283, April 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Yeh:1988:AIB

[YY88] Zhenghao Yeh and Daoning Ying. An automated interface between CAD and CAM. Computers and Graphics, 12(3–4):349–357, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

You:1996:RCM

[YY96] Chun-Fong You and Shih-Shing Yang. Reconstruction of curvilinear manifold objects from orthographic views. Computers and Graphics, 20(2): 275–293, March–April 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=2&aid=9500130.

Yatagawa:2014:TCV

[YY14] Tatsuya Yatagawa and Yasushi Yamaguchi. Temporally coherent video editing using an edit propagation matrix. Computers and Graphics, 43(??): 1–10, October 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314000533.

Yoshiyasu:2016:SAE

[YYG16] Yusuke Yoshiyasu, Eiichi Yoshida, and Leonidas Guibas. Symmetry aware embedding for shape correspondence. Computers and Graphics, 60(??):9– 22, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316300826.

Yang:2019:MFC

[YYY19] Liuming Yang, Meng Yang, and Gang Yang. Modeling fractures and cracks on tree branches. Computers and Graphics, 80(??): 63–72, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300299.

Yang:2017:ARP

[YZ17] Hao Yang and Hui Zhang. Automatic 3D reconstruction of a polyhedral object from a single line drawing under perspective projection. Computers and Graphics, 65(??):45–59, June 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300419.

Yang:2023:DIS

[YZC^{+}23] HeLiang Yang, YongJun Zhang, ZhongWei Cui, Yu-Jie Xu, and YiTong Yang. DGRN: Image super-resolution with dual gradient regression guidance. Computers and Graphics, 110(??): 141–150, February 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932200228X.

Yang:2024:ARR

[YZWJ24] Jieyin Yang, Mingyang Zhao, Yingrui Wu, and Xiaohong Jia. Accurate and robust registration of low overlapping point clouds. Computers and Graphics, 118 (??):146–160, February 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002911.

Zachmann:2015:F

[Zac15] Gabriel Zachmann. Foreword. Computers and Graphics, 53 (part B)(??): A2, December 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315001363.

Zalik:2001:MSP

[Žal01] B. Žalik. Merging a set of polygons. Computers and Graphics, 25(1):77– 88, February 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/57/26/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/57/26/33/article.
pdf.

Zayer:2012:NSA

[Zay12]  Rhaleb Zayer.  A non-

linear static approach for curve editing. Computers and Graphics, 36(5):514– 520, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000635.

Zeng:2000:MMP

[ZBM00] Jianchao Zeng, John J. Bauer, and Seong K. Mun. Modeling and mapping of prostate cancer. Computers and Graphics, 24(5):683– 694, October 2000. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/33/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/33/29/article.
pdf.

Zhang:2018:KAI

[ZBP^{+}18] Liang Zhang, Guido Brunnett, Katharina Petri, Marco Danneberg, Steffen Masik, Nicole Bandow, and Kerstin Witte. KaraKter: An autonomously interacting Karate Kumite character for VR-based training and research. Computers and Graphics, 72(??): 59–69, May 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300086.

Zalik:1999:TSU

[ŽC99] Borut Žalik and Gordon J.

Clapworthy. Technical section — a universal trapezoidation algorithm for planar polygons. Computers and Graphics, 23(3):353– 363, June 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
3/687.pdf.

Zhu:2007:SVE

[ZC07] Weizhong Zhu and Chaomei Chen. Storylines: Visual exploration and analysis in latent semantic spaces. Computers and Graphics, 31(3): 338–349, June 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307000568.

Zhou:2018:CFD

[ZC18] Jie Zhou and Xuejin Chen.

Convertible furniture design. Computers and Graphics, 70 (??):165–175, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301267.

Zhang:2016:BSS

[ZCC^{+}16] Yuhua Zhang, Juan Cao, Zhonggui Chen, Xin Li, and Xiao-Ming Zeng. Bspline surface fitting with knot position optimization. Computers and Graphics, 58(??):73–83, August 2016. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931630053X.

Zhang:2013:PCN

[ZCL^{+}13] Jie Zhang, Junjie Cao, Xiuping Liu, Jun Wang, Jian Liu, and Xiquan Shi. Point cloud normal estimation via lowrank subspace clustering. Computers and Graphics, 37 (6):697–706, October 2013. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313000824.

Zhang:2018:MCC

[ZCL^{+}18] Yu-Wei Zhang, Yanzhao Chen, Hui Liu, Zhongping Ji, and Caiming Zhang. Modeling Chinese calligraphy reliefs from one image. Computers and Graphics, 70 (??):300–306, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301152.

Zhu:2022:CBL

[ZCL^{+}22] Xiaoqiang Zhu, Qi Chen, Sihu Liu, Chenjie Fan, Chenze Song, Junjie Zhang, Dan Zeng, and Xiaogang Jin. Controllable blending of line and polygon skeleton-based convolution surfaces with finite support kernels. Computers and Graphics, 106 (??):98–109, August 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000887.

Zhou:1995:EAD

[ZCT95] Yong Zhou, Weihai Chen, and Zesheng Tang. An elaborate ambiguity detection method for constructing isosurfaces within tetrahedral meshes. Computers and Graphics, 19(3): 355–364, May–June 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=3&aid=9500006.

Zou:2018:COD

[ZCZ^{+}18] Beiji Zou, Changlong Chen, Chengzhang Zhu, Xuanchu Duan, and Zailiang Chen. Classified optic disc localization algorithm based on verification model. Computers and Graphics, 70 (??):281–287, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301243.

Zaharia:2004:MVP

[ZD04] M. D. Zaharia and L. Dorst. Modeling and visualization of 3D polygonal mesh surfaces using geometric algebra. Computers and Graphics, 28(4):519–526, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhang:2023:HLD

[ZDC^{+}23] Jia-Qi Zhang, Hao-Bin Duan, Jun-Long Chen, Ariel Shamir, and Miao Wang.

                                        HoughLaneNet: Lane de-

tection with deep Hough transform and dynamic convolution. Computers and Graphics, 116(??):82– 92, November 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001814.

Zhang:2023:VPF

[ZDD23] Yalan Zhang, Zirui Dong, and Feilong Du. Visual perception of fluid viscosity: Toward realistic fluid simulation. Computers and Graphics, 117(??):13– 19, December 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002418.

Zhang:2019:PRT

[ZDL^{+}19] Song-Hai Zhang, Xin Dong, Hui Li, Ruilong Li, and Yong-Liang Yang. PortraitNet: Real-time portrait segmentation network for mobile device. Computers and Graphics, 80(??): 104–113, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300305.

Zhou:2022:LPH

[ZDL22] Chi Zhou, Zhetong Dong, and Hongwei Lin. Learning persistent homology of 3D point clouds. Computers and Graphics, 102 (??):269–279, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002399.

Zhao:2024:GTS

[ZDL^{+}24] Yan Zhao, Jiahui Deng, Feihong Liu, Wen Tang, and Jun Feng. GO: a two-step generative optimization method for point cloud registration. Computers and Graphics, 119(??):??, April 2024. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000372.

Zhang:2007:CCB

[ZDT07] Yunfeng Zhang, Qi Duan, and E. H. Twizell. Convexity control of a bivariate rational interpolating spline surfaces. Computers and Graphics, 31 (5):679–687, October 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784930700115X.

Zarrinmehr:2017:IAS

[ZEK^{+}17] Saied Zarrinmehr, Mahmood Ettehad, Negar Kalantar, Alireza Borhani, Shinjiro Sueda, and Ergun Akleman. Interlocked Archimedean spirals for conversion of planar rigid panels into locally flexible panels with stiffness control. Computers and Graphics, 66(??): 93–102, August 2017. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317300596.

Zhao:2020:EVP

[ZFG^{+}20] Xi Zhao, Issei Fujishiro, Kentaro Go, Masahiro Toyoura, Kenji Kashiwagi, and Xiaoyang Mao. Enhancing visual performance of hemianopia patients using overview window. Computers and Graphics, 89(??): 59–67, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932030039X.

Zyda:1990:MGW

[ZFJ90] Michael J. Zyda, Mark A.

Fichten, and David H. Jennings. Meaningful graphics workstation performance measurements. Computers and Graphics, 14(3–4): 519–526, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhongke:2003:TPV

[ZFS03] Wu Zhongke, Lin Feng, and Seah Hock Soon. Topology preserving voxelisation of rational Bézier and NURBS curves. Computers and Graphics, 27(1):83– 89, February 2003. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhongke:2004:EDB

[ZFSY04] Wu Zhongke, Lin Feng, Seah Hock Soon, and Chan Kai Yun. Evaluation of difference bounds for computing rational Bézier curves and surfaces. Computers and Graphics, 28 (4):551–558, August 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zanni:2015:AIB

[ZGC15] C. Zanni, M. Gleicher, and M.-P. Cani. N-ary implicit blends with topology control. Computers and Graphics, 46(??):1– 13, February 2015. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001009.

Zheng:2019:GBE

[ZGC^{+}19] Liping Zheng, Zhiqiang Gui, Ruiwen Cai, Yue Fei, Gaofeng Zhang, and Benzhu Xu. GPU-based efficient computation of power diagram. Computers and Graphics, 80(??): 29–36, May 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300342.

Zuffo:1996:PEH

[ZGdDL^{+}96] Marcelo Knörich Zuffo, Andrew J. Grant, Roseli de Deus Lopes, Eduardo Toledo Santos, and João Antonio Zuffo. A programming environment for highperformance volume visualization applications. Computers and Graphics, 20(3): 385–394, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=9600007.

Zhang:2017:TIR

[ZGS17] Xingzi Zhang, Michael Goesele, and Alexei Sourin. Tangible images of real life scenes. Computers and Graphics, 64(??):62– 73, May 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317300195.

Zhang:2016:SAE

[ZGW^{+}16] Yuhe Zhang, Guohua Geng, Xiaoran Wei, Shunli Zhang, and Shanshan Li. A statistical approach for extraction of feature lines from point clouds. Computers and Graphics, 56(??): 31–45, May 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316300097.

Zhao:2016:RTE

[ZGWP16] Hanli Zhao, Dandan Gao, Ming Wang, and Zhigeng Pan. Real-time edge-aware weighted median filtering on the GPU. Computers and Graphics, 61(??):11– 18, December 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849316301121.

Zhang:2016:CCB

[ZGZ^{+}16] Sen Zhang, Jianwei Guo, Hui Zhang, Xiaohong Jia, Dong-Ming Yan, Junhai Yong, and Peter Wonka. Capacity constrained blue-noise sampling on surfaces. Computers and Graphics, 55(??): 44–54, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315001971.

Zhong:2022:SDS

[ZGZS22] Yue Zhong, Yulia Gryaditskaya, Honggang Zhang, and Yi-Zhe Song. A study of deep single sketchbased modeling: View/ style invariance, sparsity and latent space disentanglement. Computers and Graphics, 106(??):237– 247, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322001078.

Zhang:1996:PGC

[Zha96]  Hansong Zhang.  Pat-

tern generation with color map Gouraud shading.

Computers and Graphics, 20(1):157–162, January– February 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1&aid=9500076.

Zhang:1998:LDB

[Zha98a] Jian J. Zhang. Least distorted bump mapping onto surface patches. Computers and Graphics, 22(2–3):233– 242, March 6, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1998/22/
2-3/546.pdf.

Zhang:1998:SFC

[Zha98b] Yuefeng Zhang. Space-filling curve ordered dither. Computers and Graphics, 22(4): 559–563, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/4/577.pdf.

Zanni:2011:WBH

[ZHC11] C. Zanni, E. Hubert, and M.-P. Cani. Warp-based helical implicit primitives. Computers and Graphics, 35(3):517–523, June 2011.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000665.

Zou:2021:PGL

[ZHG^{+}21] Lu Zou, Zhangjin Huang, Naijie Gu, Fangjun Wang, Zhouwang Yang, and Guoping Wang. GMDN: a lightweight graph-based mixture density network for 3D human pose regression. Computers and Graphics, 95(??):115–122, April 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000108.

Zhao:2019:CSB

[ZHP^{+}19] Long Zhao, Fangda Han, Xi Peng, Xun Zhang, Mubbasir Kapadia, Vladimir Pavlovic, and Dimitris N. Metaxas. Cartoonish sketchbased face editing in videos using identity deformation transfer. Computers and Graphics, 79(??): 58–68, April 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300147.

Zhu:1991:VEV

[Zhu91] Qiuming Zhu. Virtual edges, viewing faces, and boundary traversal in line drawing representation of objects with curved surfaces. Computers and Graphics, 15(2): 161–173, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zou:2021:PCC

[ZHW^{+}21] Lu Zou, Zhangjin Huang, Fangjun Wang, Zhouwang Yang, and Guoping Wang. CMA: Cross-modal attention for 6D object pose estimation. Computers and Graphics, 97(??):139– 147, June 2021. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000601.

Zhukov:2000:BNM

[ZI00] S. Zhukov and A. Iones.

Building the navigational maps for intelligent agents. Computers and Graphics, 24 (1):79–89, February 2000. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/13/20/47/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/13/20/47/27/34/article.
pdf.

Zhang:2019:SSI

[ZIP^{+}19] Qimeng Zhang, Jaeho Im, Minju Park, MyungJin Choi, Chang-Hun Kim, and Yoonsik Shim. Shrubbery-shell inspired 3D model stylization. Computers and Graphics, 82(??):13–21, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319300561.

Zyda:1987:SCP

[ZJH87] Michael J. Zyda, Allan R. Jones, and Patrick G. Hogan. Surface construction from planar contours. Computers and Graphics, 11 (4):393–408, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zeng:2014:SSF

[ZjLW^{+}14] Long Zeng, Yong jin Liu, Jin Wang, Dong liang Zhang, and Matthew Ming-Fai Yuen. Sketch2Jewelry: Semantic feature modeling for sketch-based jewelry design. Computers and Graphics, 38 (??):69–77, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001581.

Zeng:2022:DMS

[ZJSB22] Zheng Zeng, Xiaohong Jia, Liyong Shen, and Pengbo Bo. Developable mesh segmentation by detecting curve-like features on Gauss images. Computers and Graphics, 109(??):42– 54, December 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322001819.

Zalik:2003:PTS

[ŽJŽ03] B. Žalik, A. Jezernik, and K. Rizman Žalik. Polygon trapezoidation by sets of open trapezoids. Computers and Graphics, 27 (5):791–800, October 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zewe:1995:MVA

[ZK95] R. Zewe and H.-J. Koglin.

A method for the visual assessment of overhead lines. Computers and Graphics, 19(1):97–108, January– February 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1&aid=9400125.

Zhao:1998:GDW

[ZK98] Jian Zhao and Eckhard Koch. A generic digital watermarking model. Computers and Graphics, 22(4): 397–403, August 1, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/cag/sub/1998/
                                        22/4/563.pdf.

Zhang:2007:CBR

[ZK07] Haitao Zhang and Arie Kaufman. A classificationbased rendering method for point models. Computers and Graphics, 31(5):730– 736, October 2007. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307001380.

Zingsheim:2024:LSS

[ZK24] Domenic Zingsheim and Reinhard Klein. Learning subsurface scattering solutions of tightly-packed granular media using optimal transport. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000220.

Zuiderveld:1996:MVM

[ZKS^{+}96] Karel J. Zuiderveld, Anton H. J. Koning, Rik Stokking, J. B. Antoine Maintz, Fred J. R. Appelman, and Max A. Viergever. Multimodality visualization of medical volume data. Computers and Graphics, 20(6):775–791, November– December 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/free/1996/20/
6/fig12.jpg.

Zhang:2015:SSH

[ZLG^{+}15] Huijuan Zhang, Chong Li, Leilei Gao, Sheng Li, and Guoping Wang. Shape segmentation by hierarchical splat clustering. Computers and Graphics, 51 (??):136–145, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931500059X.

Zhang:2010:MRM

[ZLGH10] Lei Zhang, Ligang Liu, Craig Gotsman, and Hua Huang. Mesh reconstruction by meshless denoising and parameterization. Computers and Graphics, 34(3):198–208, June 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849310000452.

Zhang:2015:FIP

[ZLL^{+}15] Zhibang Zhang, Guiqing Li, Huina Lu, Yaobin Ouyang, Mengxiao Yin, and Chuhua Xian. Fast as-isometric-as-possible shape interpolation. Computers and Graphics, 46 (??):244–256, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314000934.

Zhou:2020:PBB

[ZLL^{+}20] Jun Zhou, Yuanpeng Liu, Jinshan Liu, Qian Xie, Yuqi Zhang, Xusheng Zhu, and Xiao Ding. BOLD3D: a 3D BOLD descriptor for 6Dof pose estimation. Computers and Graphics, 89(??): 94–104, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300595.

Zhang:2021:FIC

[ZLL^{+}21] Jie Zhang, Jian Liu, Xiuping Liu, Jiang Wei, Junjie Cao, and Kewei Tang. Feature interpolation convolution for point cloud analysis. Computers and Graphics, 99 (??):182–191, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321001308.

Zhang:2023:FEE

[ZLL^{+}23a] Wenbo Zhang, Weidong Liu, Le Li, Huifeng Jiao, Yanli Li, Liwei Guo, and Jingming Xu. A framework for the efficient enhancement of nonuniform illumination underwater image using convolution neural network. Computers and Graphics, 112 (??):60–71, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784932300033X.

Zhang:2023:IMD

[ZLL^{+}23b] Yucun Zhang, Tao Li, Qun Li, Xianbin Fu, and Tao Kong. Image motion deblurring via attention generative adversarial network. Computers and Graphics, 111(??):122–132, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000092.

Zhi:2018:TRT

[ZLLG18] Shuaifeng Zhi, Yongxiang Liu, Xiang Li, and Yulan Guo. Toward realtime 3D object recognition: A lightweight volumetric CNN framework using multitask learning. Computers and Graphics, 71 (??):199–207, April 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849317301735.

Zhang:2021:PDD

[ZLLL21] Shufang Zhang, Jiang Liu, Yuhong Liu, and Nam Ling.

DIMNet: Dense implicit function network for 3D human body reconstruction. Computers and Graphics, 98(??):1–10, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000777.

Zou:2006:OTR

[ZLLY06] Yuru Zou, Wenxia Li, Jian Lu, and Ruisong Ye. Orbit trap rendering method for generating artistic images with cyclic or dihedral symmetry. Computers and Graphics, 30(3): 470–473, June 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849306000707.

Zhang:2015:SIB

[ZLM^{+}15] Yan Zhang, Zicheng Liu, Zheng Miao, Wentao Wu, Kai Liu, and Zhengxing Sun. Singe image-based data-driven indoor scenes modeling. Computers and Graphics, 53 (part B)(??): 210–223, December 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931500165X.

Zhang:1998:VDI

[ZLS98] Shouqing Zhang, Ling Li, and Hocksoon Seah. Vectorization of digital images using algebraic curves. Computers and Graphics, 22(1): 91–101, February 25, 1998. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/cag/sub/1998/
22/1/534.pdf.

Zhang:1999:TSF

[ZLS99] Shouqing Zhang, Ling Li, and Hocksoon Seah. Technical section — fine-tuning in vectorization using algebraic curves. Computers and Graphics, 23(2):269– 276, April 1, 1999. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/cag/sub/1999/23/
2/679.pdf.

Zhong:2020:CPS

[ZLZ^{+}20] Fanchao Zhong, Wenqiang Liu, Yu Zhou, Xin Yan, Yi Wan, and Lin Lu. Ceramic 3D printed sweeping surfaces. Computers and Graphics, 90(??):108– 115, August 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300583.

Zhang:2012:CCM

[ZLZG12] Min Zhang, Yinghua Li, Wei Zeng, and Xianfeng Gu. Canonical conformal mapping for high genus surfaces with boundaries. Computers and Graphics, 36 (5):417–426, August 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000453.

Zheng:1991:LPM

[ZM91] J. L. Zheng and C. B. Millham. Linear programming method for ray-convex polyhedron intersection. Computers and Graphics, 15(2): 195–204, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zheng:1992:LPM

[ZM92] J. L. Zheng and C. B. Millham. A linear pivoting method for detecting and tracing planar section curves of free-form surfaces. Computers and Graphics, 16 (4):411–420, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhou:2007:SBA

[ZM07] Hao Zhou and S. P. Mudur. 3D scan-based animation techniques for Chinese opera facial expression documentation. Computers and Graphics, 31(6):788–799, December 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849307001537.

Zhao:2023:LSP

[ZMH^{+}23] Yiqiang Zhao, Xingyi Ma, Bin Hu, Qi Zhang, Mao Ye, and Guoqing Zhou. A large-scale point cloud semantic segmentation network via local dual features and global correlations. Computers and Graphics, 111(??):133–144, April 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000122.

Zhou:2018:HSB

[ZMK18] Wei Zhou, Caiwen Ma, and Arjan Kuijper. Hough-space-based hypothesis generation and hypothesis verification for 3D object recognition and 6D pose estimation. Computers and Graphics, 72(??):122–134, May 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0097849318300116.

Zeng:2011:VCW

[ZMKG11] Wei Zeng, Joseph Marino, Arie Kaufman, and Xianfeng David Gu. Volumetric colon wall unfolding using harmonic differentials.

Computers and Graphics, 35(3):726–732, June 2011. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000471.

Zhou:2018:FFI

[ZML^{+}18] Wei Zhou, Caiwen Ma, Shenghui Liao, Jinjing Shi, Tong Yao, Peng Chang, and Arjan Kuijper. Feature fusion information statistics for feature matching in cluttered scenes. Computers and Graphics, 77 (??):50–64, December 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318301468.

Zyda:1990:RTT

[ZMM^{+}90] Michael J. Zyda, Robert B. McGhee, Corrine M. Mc-Conkle, Andrew H. Nelson, and Ron S. Ross. A real-time, three-dimensional moving platform visualization tool. Computers and Graphics, 14(2):321–333, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zeng:2006:FEO

[ZMYH06] Wei Zeng, XiangXu Meng, ChengLei Yang, and Lei Huang. Feature extraction for online handwritten characters using Delaunay triangulation. Computers and Graphics, 30 (5):779–786, October 2006. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306001270.

Zimmermann:2008:SC

[ZNA08] Johannes Zimmermann, Andrew Nealen, and Marc Alexa. Sketching contours. Computers and Graphics, 32 (5):486–499, October 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000563.

Zanabria:2016:IIS

[ZNGN16] Germain Garcia Zanabria, Luis Gustavo Nonato, and Erick Gomez-Nieto. iStar (i*): an interactive star coordinates approach for high-dimensional data exploration. Computers and Graphics, 60(??):107– 118, November 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849316301054.

Zhang:2018:EVF

[ZNT^{+}18] Lei Zhang, Duong Nguyen, David Thompson, Robert Laramee, and Guoning Chen. Enhanced vector field visualization via Lagrangian accumulation. Computers and Graphics, 70(??):224– 234, February 2018. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849317301073.

Zhang:2007:OBA

[ZO07] Jinghua Zhang and Charles B. Owen. Octree-based animated geometry compression. Computers and Graphics, 31(3):463–479, June 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849307000027.

Zomorodian:2010:FCV

[Zom10] Afra Zomorodian. Fast construction of the Vietoris– Rips complex. Computers and Graphics, 34(3): 263–271, June 2010. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849310000464.

Zhou:1992:SPB

[ZP92] Yong Zhou and Qunsheng Peng. The super-plane buffer: An efficient formfactor evaluation algorithm for progressive radiosity. Computers and Graphics, 16 (2):151–158, Summer 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zwicker:2004:PBC

[ZP04] Matthias Zwicker and Mark Pauly. Point-based computer graphics. Computers and Graphics, 28(6): 799–800, December 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhang:2007:DBI

[ZP07] Yanci Zhang and Renato Pajarola. Deferred blending: Image composition for single-pass point rendering. Computers and Graphics, 31(2):175–189, April 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784930600238X.

Zhong:2023:EFF

[ZPIS23] Shengze Zhong, Parinya Punpongsanon, Daisuke Iwai, and Kosuke Sato. Estimation of fused-filament-fabrication structural vibroacoustic performance by modal impact sound. Computers and Graphics, 115 (??):137–147, October 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001371.

Zhu:1988:PPE

[ZPL88] Yining Zhu, Qunsheng Peng, and Youdong Liang. PERIS: a programming environment for realistic image synthesis. Computers and Graphics, 12(3–4):299–307, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zou:2015:SBD

[ZPL^{+}15] Changqing Zou, Xiaojiang Peng, Hao Lv, Shifeng Chen, Hongbo Fu, and Jianzhuang Liu. Sketch-based 3-d modeling for piecewise planar objects in single images. Computers and Graphics, 46 (??):130–137, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849314001198.

Zhang:2021:MMI

[ZPN^{+}21] Heng Zhang, Yuanyuan Pu, Rencan Nie, Dan Xu, Zhengpeng Zhao, and Wenhua Qian. Multi-modal image synthesis combining contentstyle adaptive normalization and attentive normalization. Computers and Graphics, 98(??):48–57, August 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000728.

Zhang:2023:ACM

[ZPP^{+}23] Chao Zhang, Romain Pinquié, Arnaud Polette, Gregorio Carasi, Henri De Charnace, and Jean-Philippe Pernot. Automatic 3D CAD models reconstruction from 2D orthographic drawings. Computers and Graphics, 114(??):179–189, August 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000766.

Zhou:2012:BMP

[ZQ12] Pei Zhou and Wen-Han Qian. G^{n}-blending of multiple parametric normal ringed surfaces by adding implicit closings G^{n}-continuous with the surfaces. Computers and Graphics, 36(4): 297–304, June 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000404.

Zhao:2015:DAH

[ZQL15] Yong Zhao, Hailong Qian, and Shengjie Lu. A deformation-aware hierarchical framework for shapepreserving editing of static and time-varying mesh data. Computers and Graphics, 46 (??):80–88, February 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849314001010.

Zedler:1997:MIM

[ZR97] Jörg Zedler and Marwan Ramadan. i-Media: An integrated media server and media database as a basic component of a cross media publishing system. Computers and Graphics, 21(6):693–702, November– December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/cag/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=21&
                                        issue=6&aid=9700047.

Zhuo:2012:CBO

[ZR12] Wei Zhuo and Jarek Rossignac. Curvature-based offset distance: Implementations and applications. Computers and Graphics, 36(5):445– 454, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000520.

Zhao:1994:RSO

[ZS94] Xiuwei Zhao and Jiaguang Sun. Reconstruction of a symmetrical object from its perspective image. Computers and Graphics, 18(4): 463–467, July–August 1994.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhou:2002:BPC

[ZS02] Ji Zhou and Jiaoying Shi.

Best papers of CAD & CG 2001: a robust algorithm for feature point matching. Computers and Graphics, 26(3):429–436, June ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/54/32/abstract.
html.

Zhang:2014:TMS

[ZS14] Jingqiao Zhang and Zhe Shi.

Triangulation of molecular surfaces based on extracting surface atoms. Computers and Graphics, 38 (??):291–299, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001866.

Zhou:2012:TFC

[ZSH12] Liang Zhou, Mathias Schott, and Charles Hansen. Transfer function combinations. Computers and Graphics, 36 (6):596–606, October 2012. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000246.

Zhou:1995:SMS

[ZSK95] Chen Zhou, Renben Shu, and Mohan S. Kankanhalli. Selectively meshed surface representation. Computers and Graphics, 19 (6):793–804, November– December 1995. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/cag/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=6&aid=9500050.

Zhang:2008:TSB

[ZSL08] Yan Zhang, Zhengxing Sun, and Wenhui Li. Texture synthesis based on Direction Empirical Mode Decomposition. Computers and Graphics, 32(2):175– 186, April 2008. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849308000034.

Zhao:2012:RWA

[ZSL12] Ling Zhao, Xukun Shen, and Xiang Long. Robust wrinkle-aware non-rigid registration for triangle meshes of hand with rich and dynamic details. Computers and Graphics, 36(5):577– 583, August 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009784931200074X.

Zorzal:2019:AST

[ZSM^{+}19] Ezequiel R. Zorzal, Maurício Sousa, Daniel Mendes, Rafael Kuffner dos Anjos, Daniel Medeiros, Soraia Figueiredo Paulo, Pedro Rodrigues, José João Mendes, Vincent Delmas, Jean-Francois Uhl, José Mogorrón, Joaquim Armando Jorge, and Daniel Simões Lopes. Anatomy Studio: a tool for virtual dissection through augmented 3D reconstruction. Computers and Graphics, 85(??):74– 84, December 2019. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319301578.

Zhu:2018:BIR

[ZSS^{+}18] Hengliang Zhu, Bin Sheng, Zhiwen Shao, Yangyang Hao, Xiaonan Hou, and Lizhuang Ma. Better initialization for regression-based face alignment. Computers and Graphics, 70(??): 261–269, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301292.

Zheng:2020:CRI

[ZSS20] Yufei Zheng, Hatsu Shi, and Suguru Saito. Creating reference image of realistic cloth folded surface using sketchbased interactive modeling. Computers and Graphics, 91 (??):108–117, October 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300984.

Zhang:2008:NNR

[ZSW08] Yanci Zhang, Hanqiu Sun, and Enhua Wu. NBS: a new representation for point surfaces based on genetic clustering algorithm: CAD and graphics. Computers and Graphics, 32(6): 639–646, December 2008. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849308001180.

Zhang:2016:FSS

[ZT16] Hao (Richard) Zhang and Anthony Tang. Foreword to special section on Graphics Interface 2015. Computers and Graphics, 55(??): A1–A2, April 2016. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849315002095.

Zamanakos:2021:CSL

[ZTAP21] Georgios Zamanakos, Lazaros Tsochatzidis, Angelos Amanatiadis, and Ioannis Pratikakis. A comprehensive survey of LIDAR-based 3D object detection methods with deep learning for autonomous driving. Computers and Graphics, 99(??):153–181, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001321.

Zhao:2022:GDO

[ZTF^{+}22] Yan Zhao, Wen Tang, Jun Feng, Taoruan Wan, and Long Xi. General discriminative optimization for point set registration. Computers and Graphics, 102 (??):521–532, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321002417.

Zuckerberger:2002:PSD

[ZTS02] Emanoil Zuckerberger, Ayellet Tal, and Shymon Shlafman. Polyhedral surface decomposition with applications. Computers and Graphics, 26(5):733–743, October ??, 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/56/36/abstract.
html.

Zubiaga:1988:ITP

[Zub88] Ricardo Berlanga Zubiaga. On the intersection of two planar polygons. Computers and Graphics, 12(3–4): 401–403, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zyda:1988:DNS

[ZW88] Michael J. Zyda and Robert A. Walker. Design notes on a single board multiprocessor for real-time contour surface display generation. Computers and Graphics, 12(1):91–97, 1988. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhang:2020:TVD

[ZW20] Huayan Zhang and Chunxue Wang. Total variation diffusion and its application in shape decomposition. Computers and Graphics, 90 (??):95–107, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300753.

Zhou:2022:MIO

[ZWL^{+}22] Ruqin Zhou, Hanyun Wang, Xixing Li, Yulan Guo, Chenguang Dai, and Wanshou Jiang. MaskNet++: Inlier/outlier identification for two point clouds. Computers and Graphics, 103(??):

90–100, April 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322000085.

Zyda:1993:NOD

[ZWP^{+}93] Michael J. Zyda, Kalin P.

Wilson, David R. Pratt, James G. Monahan, and John S. Falby. NPSOFF: An object description language for supporting virtual world construction. Computers and Graphics, 17(4): 457–464, July–August 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Zhao:2006:MFM

[ZWQ^{+}06] Ye Zhao, Lujin Wang, Feng Qiu, Arie Kaufman, and Klaus Mueller. Melting and flowing in multiphase environment. Computers and Graphics, 30(4):519– 528, August 2006. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849306000847.

Zhao:2018:PFP

[ZWR^{+}18] Haiming Zhao, Jufeng Wang, Xiaoyu Ren, Jingyuan Li, Yong-Liang Yang, and Xiaogang Jin. Personalized food printing for portrait images. Computers and Graphics, 70 (??):188–197, February 2018.

CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009784931730105X.

Zadeh:2019:IIU

[ZWS19] Shekoufeh Gorgi Zadeh, Maximilian W. M. Wintergerst, and Thomas Schultz. Intelligent interaction and uncertainty visualization for efficient drusen and retinal layer segmentation in Optical Coherence Tomography. Computers and Graphics, 83 (??):51–61, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849319301074.

Zhang:2022:FSS

[ZWSW22] Yunbo Zhang, Emily Whiting, Cynthia Sung, and Charlie C. L. Wang. Foreword to the special section on computational fabrication. Computers and Graphics, 102(??):A6–A7, February 2022. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000188.

Zhang:2023:HRG

[ZWWC23] Gege Zhang, Luping Wang, Liang Wang, and Zengping Chen. Hand-raising gesture detection in classroom with spatial context augmentation and dilated convolution. Computers and Graphics, 110(??):151– 161, February 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849322002047.

Zhou:2014:IBR

[ZXD^{+}14] Xuehong Zhou, Guanyu Xing, Zhipeng Ding, Yanli Liu, Junjun Xiong, and Qunsheng Peng. Image-based relighting from a sparse set of outdoor images. Computers and Graphics, 38 (??):230–238, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001714.

Zhang:2012:VNM

[ZXH^{+}12] Mingmin Zhang, Mingliang Xu, Lizhen Han, Yong Liu, Pei Lv, and Gaoqi He. Virtual Network Marathon with immersion, scientificalness, competitiveness, adaptability and learning. Computers and Graphics, 36(3): 185–192, May 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849312000076.

Zeng:2021:RTI

[ZXL^{+}21] Wei Zeng, Bo Xia, Hui Li, Guanyu Xing, Yanli Liu, and Yanci Zhang. Realtime indirect illumination by virtual planar area lights. Computers and Graphics, 97(??):183–190, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849321000649.

Zhu:2023:MDE

[ZXLF23] Tianyu Zhu, Zeng-Hao Xu, Ligang Liu, and Xiao-Ming Fu. Modeling with discrete equivalence classes of planar quads. Computers and Graphics, 115(??):404– 411, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323001619.

Zhang:2001:CSU

[ZY01] Dongliang Zhang and Matthew M. F. Yuen. Cloth simulation using multilevel meshes. Computers and Graphics, 25(3):383–389, June 2001. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/13/20/57/32/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/13/20/57/32/29/article.
                                        pdf.

Zhang:2002:PBS

[ZY02] Jian J. Zhang and Lihua You. PDE based surface representation — vase design. Computers and Graphics, 26(1):89–98, February 2002. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
13/20/68/27/36/abstract.
html.

Zhu:2023:SPC

[ZYML23] Mengting Zhu, Meng Yang, Weiliang Meng, and Ping Li. Sand painting conversion based on detail preservation. Computers and Graphics, 115(??):371– 381, October 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323001401.

Zheng:2021:NCM

[ZYW^{+}21] Liping Zheng, Yuyou Yao, Wenming Wu, Benzhu Xu, and Gaofeng Zhang. A novel computation method of hybrid capacity constrained centroidal power diagram. Computers and Graphics, 97(??):108–116, June 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000492.

Zheng:2023:DNS

[ZYW23] Qiumei Zheng, Tao Yu, and Fenghua Wang. DCU-NET: Self-supervised monocular depth estimation based on densely connected Ushaped convolutional neural networks. Computers and Graphics, 111(??): 145–154, April 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323000237.

Zhang:2019:RTE

[ZYX^{+}19] Congyi Zhang, Lei Yang, Liyou Xu, Guoping Wang, and Wenping Wang. Realtime editing of man-made mesh models under geometric constraints. Computers and Graphics, 82 (??):174–182, August 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849319300901.

Zhao:2015:ICR

[ZZ15] Xuan-Yi Zhao and Chun-Gang Zhu. Injectivity conditions of rational Bézier surfaces. Computers and Graphics, 51(??):17–25, October 2015. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849315000643.

Zheng:2014:GCC

[ZZC^{+}14] Liping Zheng, Jianming Zhao, Yajun Cheng, Haibo Chen, Xiaoping Liu, and Wenping Wang. Geometryconstrained crowd formation animation. Computers and Graphics, 38(??): 268–276, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849313001763.

Zhang:2020:PDF

[ZZC20a] Yuzhe Zhang, Jianmin Zheng, and Yiyu Cai. Proxydriven free-form deformation by topology-adjustable control lattice. Computers and Graphics, 89(??): 167–177, June 2020. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849320300649.

Zhao:2020:AMP

[ZZC^{+}20b] Junhao Zhao, Chen Zong, Luming Cao, Shuangmin Chen, Guozhu Liu, Jian Xu, and Shiqing Xin. Automatically modeling piecewise planar furniture shapes from unorganized point cloud. Computers and Graphics, 90 (??):116–125, August 2020. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849320300704.

Zheng:2014:SBH

[ZZCL14] Jiaxiang Zheng, Ming Zeng, Xuan Cheng, and Xinguo Liu. SCAPE-based human performance reconstruction. Computers and Graphics, 38 (??):191–198, February 2014. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001647.

Zhang:2022:SSM

[ZZCY22] Wenjing Zhang, Jianmin Zheng, Yiyu Cai, and Anders Ynnerman. Seamless simplification of multichart textured meshes with adaptively updated correspondence. Computers and Graphics, 106(??):77– 87, August 2022. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849322000930.

Zielasko:2019:VBE

[ZZD^{+}19] Daniel Zielasko, Xiaoqing Zhao, Ali Can Demiralp, Torsten W. Kuhlen, and Benjamin Weyers. Voxelbased edge bundling through direction-aware kernel smoothing. Computers and Graphics, 83(??):87–96, October 2019. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0097849319301025.

Zhu:2010:QAD

[ZZDZ10] Qing Zhu, Junqiao Zhao, Zhiqiang Du, and Yeting Zhang. Quantitative analysis of discrete 3D geometrical detail levels based on perceptual metric. Computers and Graphics, 34 (1):55–65, February 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849309001307.

Zhang:2023:DSD

[ZZF^{+}23] Yucun Zhang, Jiawei Zhang, Xianbin Fu, Nanhe Jiang, and Qun Li. Dynamic scene deblurring via receptive field attention generative adversarial network. Computers and Graphics, 116(??): 354–362, November 2023. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849323002273.

Zhou:2023:MPV

[ZZH^{+}23] Wei Zhou, Xiaodan Zhang, Xingxing Hao, Dekui Wang, and Ying He. Multi Point-Voxel Convolution (MPV-Conv) for deep learning on point clouds. Computers and Graphics, 112(??):

72–80, May 2023. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849323000377.

Zhang:2024:OEG

[ZZH^{+}24] Fukai Zhang, Lulu Zhang, Tiancheng He, Yiran Sun, Shan Zhao, Yanmei Zhang, Xueliang Zhao, and Weiye Zhao. An overlap estimation guided feature metric approach for real point cloud registration. Computers and Graphics, 119 (??):??, April 2024. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849324000104.

Zhao:2014:MBC

[ZZJ^{+}14] Shifeng Zhao, Mingquan Zhou, Taorui Jia, Pengfei Xu, Zhongke Wu, Yun Tian, Yu Peng, and Jesse S. Jin. Multi-branched cerebrovascular segmentation based on phase-field and likelihood model. Computers and Graphics, 38(??):239– 247, February 2014. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849313001830.

Zhao:2018:EID

[ZZJ18] Hanli Zhao, Haining Zhang, and Xiaogang Jin. Efficient image decolorization with a multimodal contrastpreserving measure. Computers and Graphics, 70 (??):251–260, February 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849317301024.

Zhang:2021:PUI

[ZZL21] Chi Zhang, Wei Zeng, and Ligang Liu. UrbanVR: an immersive analytics system for context-aware urban design. Computers and Graphics, 99 (??):128–138, October 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321001370.

Zhang:2021:EPS

[ZZLZ21] Yun Zhang, Fang-Lue Zhang, Yu-Kun Lai, and Zhe Zhu. Efficient propagation of sparse edits on 360^{∘} panoramas. Computers and Graphics, 96(??):61–70, May 2021. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0097849321000418.

Zhang:2011:ISV

[ZZQW11] Yingping Zhang, Dengming Zhu, Xianjie Qiu, and Zhaoqi Wang. Importance sampling for volumetric illumination of flames. Computers and Graphics, 35(2): 312–319, April 2011. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849311000021.

Zhou:2018:PNR

[ZZXT18] Shizhe Zhou, Chengfeng Zhou, Yi Xiao, and Guanghua Tan. PatchSwapper: A novel real-time single-image editing technique by regionswapping. Computers and Graphics, 73(??):80– 87, June 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849318300359.

Zhou:2012:EAS

[ZZY12] Hailing Zhou, Jianmin Zheng, and Xunnian Yang. Euler arc splines for curve completion. Computers and Graphics, 36(6):642– 650, October 2012. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0097849312000891.